A Complete Bibliography of Lecture Notes in Mathematics (1970–1974)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

13 October 2017 Version 1.09

Title word cross-reference

(0, 1) [Hed71]. (AC)^{o} [Luc73g]. (B, N) [Cur70]. (C, α) [Tre73e, Tre73d]. (k − 1) [Lan72]. (ω AC) [Luc73g]. (ωAC) [Luc73g]. (p, q) [Ram74c]. (p − 2) [Chi71c]. (\underline{V}) [Luc73h, Luc73g]. (Λ)^{o} [Luc73g]. 0 [HMS74e, Tim71]. \{0, 1\} [Hec74b]. 1 [And74a, Gib72, cH73, Lit74a, Mat73a, Mey70e, Ray70j, Sac73]. 1ère [Bre72c]. 2 [But74d, Dev73i, Dwo73, Fre71b, GM74, Gra74e, Gra74d, Gra74h, Gra74j, GM71c, HR70, KS74b, MS73a, Qui73c, Str74d, Yam74f]. 2^{(Im)} [Won70]. 2ième [Bre72d]. 2^{x} [Fle70]. 2k [Lan72]. 3 [Bin74, Háj73, Hak74, Kin72, Lor74, Mar74a, Mar71d, Orl72a, Roi71]. 4 [Bar71a, Bar72, BS71]. 8 [Qui73c]. [p] [vdK73h]. ^{∗} [Wei73]. ^{e} [Die72]. _{0} [Grz70]. _{1} [Pop70m]. A(G, τ) [HC70j]. A_{∞} [May72a, May72n]. A_{n} [Sta70a]. A_{n}(−1) [FS70e]. A_{n}, D_{n}, E_{n} [Tit74g]. A_{p} [Eym71]. A_{q}(Γ) ⊂ B_{q}(Γ) [Leh73b]. α [Bru74b, GU71i, GU71k, GU71l, Orl72b, Ulm71]. argL(12 + it, x) [Mon71h]. B [BS71]. B(k) [Lor70c]. B_{n}(−1) [FS70e]. B_{P L} [Mil71]. B_{T OP} [Mil71]. B_{X} [Sta70r]. m̄ [KMT74c]. N̄ [Coh74g, Coh74j, Coh74m, Coh74h]. ¯∂ [Ker73, Sat71]. ¯∂^{±} [Ati74h]. BI_{D} [Luc73j]. mod3 [Har74f]. modp [Har74g, Sta74c, Tod71, Wil74b, Cur71, Mis71b]. modp > 2 [Sta71d]. BP_{∗}(BP) [Liu71]. C [Coh74e, Coh74a, Coh74b, Coh74i, Coh74f, HN74, Pop70i, Pop70e, Pop70h, Pop70o, Sch74b]. C(α) [Chi73]. C(X) [Bin73]. C∗ [Beh72a, BDF73, Hof70a, Hof70b, Hof70f, Hof70l, Hof70m, Hof70n, Hof70o, OW72]. c = 1/2 [Mur72j]. C^{(g)} [Ive70b]. C^{1} [DD74a]. C^{∞} [LB70j, Mat71b, Poe74k, Wal71b]. C^{n} [Roo74, Sic74]. C^{ω} [LB70a, LB70k]. C^{p} [Kel74d]. C^{p,q} [Lof70]. C_{∗} [May72f]. C_{n} [Tit74i, Tit74j, Tit74h]. Con(ZF) → Con(ZF C + GCH + SH) [DJ74a]. D [Mur72f, Mur72g, Mur72j, Mur72k, Bea70d]. d − σ − G [Bea70i]. d′d′′ [Bar74b]. d_{p} [Sch72f]. Δ [GU71c, Tak70g, Tak70h, Tak70j]. ◇ [DJ74c]. e [Del71a]. e[u] = 0 [Gil74a]. e^{iv−ρ(H(x))} [Coh74k]. E^{n} [Ner71l]. E^{n}(A) [Gat73b]. E_{∞} [May72a, May72b]. ℓ [Lan73b, SD73, Ver74]. ℓ^{2} [Gre74]. ϵ [Che74f, Fit72]. Ext(X) [KS73a]. F [Adl73]. f(λ) + f(ωλ)f(ω^{−1}λ) = 1, (ω^{5} = 1) [SC73]. F (z_{xx}, z_{xy}, z_{yy}) = 0 [Bul72]. F_{4} [Har74f, Tit74k]. F_{f} [HC70h]. F_{J} [Coh74n, Coh74o]. F_{J}(ν) [Coh74q]. F_{J}(ν|n̄)(b) [Coh74l]. F_{n} [Mil74c]. F_{n}(X) → ΩF_{n−1}(ΣX) [Mil74j]. Fun(A, B) [Gra74i]. G [Aus73b, Bur74a, Col73b, Ill72f, MS73b, MS73c, MS73d, MS73f, MS73g, MS73h, MS73i, Sil70b, Zag72a, Zag72b, vdK73g, vdK73h, vdK73i, Pop70g]. G∗ [vdK73j, vdK73n]. G = S^{1} [Ati74g, Ati74h]. G = T^{n} [Ati74j]. g_{R}^{∗} [vdK73a]. g_{Z}^{∗} [vdK73a]. G_{0} [SE70i]. G_{1} [SE70b]. G_{δ} [Bur74b]. g_{F}: k_{2}F → B(F) [EL73]. g_{p} [Sch72f]. ≥ 2 [Pop70i]. ≥ 3 [Tit74n]. GL(2) [Del73e, JL70a, Jac72a]. GL(2) × GL(2) [Jac72c, Jac72d]. Gl_{2} [Cas73b, Rob73k]. GL_{2}(k) [Cas73a]. H [Bec71d, CD71, Cur71, Har74g, HR70, Jam71, Kle72, Mey73d, Mis71a, Mis71b, Por71c, Rat73, Roi71, Sha71l, Sig71a, Sig71b, Sta70m, Sta70l, Sta70e, Sta70f, Sta70q, Sta70g, Sta71a, Sta71d, Sta71b, Sta74a, Wil74b, Zab70, Zab71, Ber74c, Ber74d, Mat73i, Mey74a, Wil74c]. H_{1}(I, M) [FK72a]. H_{N}/H_{S+N} [OG74a]. H_{p} [Her74e]. Âd [vdK73a]. I(R_{T}: S_{T}) [OG74a]. I_{N}∗ [Sca71h]. j [Kan71f, Kan71j]. J_{S}(X) [Maa71f]. J_{T} (N, S + N) [OG74a]. K [And73a, AKW73, Bas73a, Bas73b, Bas73c, Bas73d, Bas73e, Bec74b, Ber71n, Ber71l, Blo73a, Blo73b, BG73, Cap73b, Ger73a, Ger73b, Ger73c, GL70, HR74c, cH73, Jou73a, Jou73b, Kar70b, KGLZ70, Kar70c, Kar73b, Kar73c, Keu73, Lic73, Pet73, Qui73b, Rob73m, Sha73a, Sha73b, Sil70b, Swa70, Swa73, Wag73, Zis70, AS71q, Chu74a, Poe74k]. K(π, n) [Mil74b]. K(Z, n) [Mil74l]. K = PGL(2, 0) [Sil70c]. K^{.} [Ber71m]. K_{0} [MP73b, Ped73, Rob73n, SE70a, SE70i]. K_{1} [MP73b, SE70b]. K_{2} [Bas72, Car73a, DS73, Den73a, Hat73, Sha73c]. K_{2}Z [Hum73]. K_{i} [Qui73a]. kerπ [vdK73i]. L [Ben73, Boh74, Del73l, Deu73e, Mon71d, Mon71k, Ran73b, Wal73b, Zag72c, Zag72d]. L^{1} [Rei71b, Rei71a, Rei71g, Rei72, Lep72]. L^{2}(G) [CW71b]. L^{2}(R_{n}^{+}) [Sch71x]. L^{p} [Eym71, Gru74a]. L_{∞} [PBR71]. L_{p} [Sim72a]. L_{V} [vdK73b]. λ [Ber71k, Fei70, Knu73c, Knu73g]. ≤ 8 [Fou71]. m [Bar71a]. M(D) [Beh74a]. M_{(P)} [Cho73d]. M_{+}(X) [Top70m]. M_{α} [Maa71b]. M_{i} [BL74b]. C^{2} [Ste72]. C^{K} [Kuj71a]. C^{n} [Sko72]. P_{□} [Lem74d]. P_{n} − P_{k} [Bar74b]. E(G) [Poé74a]. H^{√} [Jan71]. ∑_{r=0}^{n} c_{r}(pD_{1} + qD_{2})^{r}u = 0 [Fen74]. μ

[Aze70f, Dev73l]. N [AC74, Hay74, Jan74a, JW73a, JW73f, Smi74, Ter70, Wie74, Gro70b, Ive70e, Maa71i, Maa71j, Mar71c, Mil74p, NS72, Pop70f, Sai71a, Väi71c]. n = 1 [Sch73-31]. n = 2 [Sch73-31]. ≠ [Ste70c]. n ≥ [Gil74c]. ν [Bla74, Coh74h]. O [Pop70p]. O^{+} [Ahm71]. Ω_{∗} [May72f]. ω_{1} [BS73c, Büc73a]. ω_{X} [AK70j]. ⊗ [Riv72a]. p [BK72n, BK72c, Bre72c, Bre72d, Bru74c, Bru74a, CR73b, Chi71c, Cor74, Dem72d, Dem72a, Dem72c, Gru70h, HC70d, Kat73g, PŠ70, Pop70g, Pop70f, Ram74c, Sch72f, Ser71a, Ser73k, Sil70a, Van72, Yam74g, Zab74]. P (t)u_{tt} + M(t)u_{t} + N(t)u = C [Lev73]. p > [Ber74i]. p > 0 [Ber74i, Pop70g, Pop70f]. P_{1} [Rei71k, Rei71q]. PGL(2, Ω) [Sil70f]. PGL_{2} [Sil70a]. ϕ(∗)Φ(Y) [AS71a]. Φ_{ϕ}(x) [AS71a]. ϕ: G∗ → G [vdK73f, vdK73e]. Π [Bas73f, Pop70m]. Π_{1}(X − C) [Pop70o]. π_{1} = Z^{k} [Hsi71]. Π_{1}^{(z)}(P^{n} − C) [Pop70e, Pop70e]. Π_{1}^{(z)}(X − C) [Pop70i, Pop70h]. Π_{1}^{1} [Cut72]. π_{i}(a; MrelX) [ABK71d, ABK71g]. π_{i}(B, a; MrelX) [ABK71a]. π_{i+1}(B, a; MrelX) [ABK71d]. p ≠ 2 [Yam74g]. Pseud(A, B) [Gra74i]. ψ^{q} − 1 [Hus73]. Q [Gel74, Coh74t, Mey74a, Ram74c]. qf [Luc73h, Luc73g]. R [BK72p, BK72o, BK72d, BK72j, BK72q, Mat73d, Mat73c]. R^{3} [Ess74c]. R^{n} [KR73]. r: g′Z → gZ [vdK73d]. ρ [ST71a]. ⇒ [Poe74t]. S [Gui72g, Poe72]. S4 [Fin72]. S^{1} [KK72, Orl72d, Wan72, Kaw72a, Kaw72b]. S^{3} [KK72, Orl72e, Rec71, Wan72]. S^{p+q} [Lod73]. S^{p} × S^{q} [Lod73]. S_{∗} [May72f]. S_{u}^{0} [Sil70d]. Σ [GU71j, vdK73b, Hor70a, Loe74a, OS70, Sch73z]. σ(∞, (S^{1}, M^{n})) [KR72]. SK_{1}(ZΠ) = 0 [Bas73f]. sk_{⊥}(ZZπ) [ADS73]. SL(2, R) [Low73]. SOX_{R}(−) [BR73b]. Sp(n, l) [Rei74a]. SU(2) [CW71c, CW71e]. su(3) [MM72]. T [Aze70e, Ber71j, Luc73b, Luc73h]. T^{n} [McG72]. T_{B} [Bea70c]. T_{C} [Bea70j]. T ∪ BR [Luc73n, Luc73a, Luc73g]. θ [Maz72]. θ(δ) [Fer70g]. θ = 0 [HC70f, HC70g, HC70h]. Δ̃_{1} [oT72]. u [Sil70d]. U_{p} [Dwo73]. \underline{lim} [Jen72f]. V [Dub70f, Dub70e, Sta73c]. W_{o}^{m,p}(G)(1 < p < ∞) [Sim72e]. W_{T} [OG74a]. X [Pop70i, Pop70h, Pop70o, Iwa71a]. x_{K} [Jen72i]. ξ [Top70h, Top70p]. X log^{+} X [GC70]. X → Ω^{n}Σ^{n}X [Mil74f]. y(x + 1) = y(x) + 1 + λ/y(x) [Kim73]. y′′ = 6y^{2} + x [Ree74]. Y′′ = q(t)Y [Bor74]. z [Fer70g]. Z_{2} [Sch72d]. Z_{2k} [Kaw72a]. Z_{ν} [Spi74b]. ζ [Kat73d]. |D|^{− \frac{1}{2} −ε} [HC70f].

- [Luc73h, Luc73g, Roi71, Beh72a, Luc73g]. -actions [Kaw72b]. -Adic [But74d, Dem72d, Lan73b, HC70d, PŠ70, SD73, Van72, Yam74g, Yam74f, Kat73g]. -adics [Sil70a]. -adique [CR73b, Ver74]. -adiques [Del71a, Ser73k]. -Algebras [Kan71f, Kan71j, Rei71a, BDF73, Hof70a, Hof70f, Hof70m, Rei71b, Rei71g, Rei72]. -Algebren [Lep72]. -analysis [Luc73g]. -analytiques [Col73b]. -Anneaux [Ber71k]. -Bigebras [Hof70l, Hof70n]. -body [Mil74p, Hay74, Jan74a, Smi74, Wie74]. -Bounded [Tre73e, Tre73d]. -calculus [Fit72]. -cascades [Sha71l]. -categoricity [Cut72]. -Categories [Gra74h, Gra74e, Gra74d, KS74b, Riv72a]. -category [Str74d]. -cell [Bin74].

-circuits [Kle72]. -classes [Zag72c, Zag72d]. -cobordisme [Poe72]. -Cogebras [Hof70b]. -cohomologie [Bar74b]. -Cohomology [Sat71, GM71c]. -color [BS71]. -comma [Gra74j]. -completion [BK72p, BK72o, BK72d, BK72j]. -Completions [BK72n, BK72c]. -complets [Mey74a]. -complexes [Sta71a]. -concaves [Ram74c]. -connected [Lan72, vdK73b]. -continuous [GU71i, GU71j]. -convexes [Ram74c]. -convexes-concaves [Ram74c]. -convexity [ST71a]. -definability [oT72]. -diagonal [Bur74b]. -Dimensional [HMS74e, Mat73a, Väi71c, Gro70b, Mar74a, Tim71]. -Dirichlet [Wei73]. -Divisible [Dem72c, Mat73i, Dem72a]. -domains [Rat73]. -embedded [Bla74]. -embedding [HN74]. -Entropie [Che74f]. -erzeugbare [GU71k]. -expansions [Adl73]. -factorizations [And74a]. -factors [Gib72, Lit74a]. -fields [Sch73z]. -Fines [Bru74c, Bru74b]. -finite [Hor70a, Loe74a, OS70]. -fold [Ive70e]. -Function [Coh74b, Coh74e, Coh74i, MS73b, MS73c, MS73g, MS73h, MS73i]. -functions [Ber71j, Deu73e, MS73d, MS73f, Mon71d, Mon71k]. -functors [Ger73b]. -generated [Ulm71]. -gons [Bar71a]. -groups [Gru70h]. -harmonic [Aze70f]. -harmoniques [Aze70f]. -induced [Aus73b]. -invariant [Orl72b, vdK73h]. -localization [GM74]. -localizations [BK72q]. -Manifolds [Háj73, Fre71b, Kin72, Lan72, Orl72a]. -maps [Sig71a, Sta70q]. -matrices [Hed71]. -Matrizen [Boh74]. -module [vdK73g, vdK73i]. -monads [Dub70e]. -monomorphisms [Dub70f]. -normal [Cor74]. -order [Chi71c]. -pairs [Cur70]. -parameter [Ter70]. -parti [Ber74c, Ber74d, Mey74a]. -particle [JW73a, JW73f]. -Piecewise-Polynomial [DD74a]. -präsentierbare [GU71l, GU71c]. -presentable [Ulm71]. -primen [Pop70g, Pop70f]. -problem [Top70h, Top70p]. -producible [GU71k]. -radonifiantes [Sch72f]. -regular [Bar71a]. -relator [Sac73]. -Ring [Coh74f, Coh74a]. -Rings [Knu73c, Bar72, Knu73g]. -schémas [Ill72f]. -sets [BS71]. -signature [Zag72a, Zag72b]. -sommantes [Sch72f]. -space [Lor74, Mis71a, Por71c, Sta70l, Sta71b, Sta74a]. -Spaces [Sig71b, Wil74c, Har74g, Wil74b, Bec71d, CD71, Cur71, HR70, Jam71, Mis71b, Roi71, Sta70m, Sta70e, Sta70f, Sta70g, Sta70a, Sta71d, Zab70, Zab71]. -spheres [Hak74]. -stability [Wal71b]. -stetiger [GU71i, GU71j]. -structures [vdK73h]. -suffisants [Poe74k]. -th [NS72]. -Théorie [Jou73b, KGLZ70, Kar73b, Jou73a, Kar70b, Kar70c, Zis70]. -Theories [Bas73c, And73a, AKW73]. -Theory [Bas73a, Bas73b, Sim72a, Swa70, Bas73d, Bas73e, Bec74b, Blo73a, Blo73b, BG73, Cap73b, Ger73a, Ger73c, HR74c, cH73, Jan71, Kar70c, Kar73c, Keu73, Lic73, Pet73, Qui73b, Ran73b, Rob73m, Sha73a, Sha73b, Swa73, Wag73, Wal73b]. -Torsion [Ser71a]. -torus [Sai71a, Sch72d]. -ultradistributions [Cho73d]. -Uniform [Gel74]. -variation [Bru74a, Bre72c, Bre72d].

/ [Sca71h].

1 [Dev73c, Tac73k, Tur71]. 10e [Azr72]. 11.-4 [AT72]. 13 [Tac73h]. 13th [Fil73]. 14th [Fil73]. 1963 [Ser73d]. 1966 [Bre71p]. 1967/68 [Tem71b]. 1968 [BCC^{+}70, LLNS70, Moi70]. 1968/69 [Ano71a, BCC^{+}70]. 1969 [BCdV^{+}70, Chi71a, DS70a, Fle70, HLM71, Lel70b, Yor70]. 1969/70 [Ano71b, Tem71b]. 1970 [Bro71a, Bru71c, Din71b, Eic71e, Gra71b, Her71a, Hor70b, Hor71, HS71, Kui71a, LLN74, Lel71c, Maa71r, Mor71c, Pet70a, Sig71b, Suc70]. 1970/71 [Hof72c]. 1970/71 [Ano72]. 1971 [AT72, GG72, GL72, Kom73c, KMSS72a, KMSS72b, Law72a, Lel72b, MR73, Mor71b, Rob73j, Ura71c]. 1971/72 [DE73, Rob73j]. 1972 [Aku73, ALW72, AHiN74, Lel73b, AT73, Bas73a, Bas73b, Bas73c, BRC74, BR73c, DK73b, ES72c, GW73, HS73b, HL74c, Kel74i, Kno73b, Kui73a, KS73b, Nac74a, SJS73, Vid74]. 1972/1973 [Kel74i]. 1972/73 [Ano74]. 1973 [AT74, BH74a, BH74b, DF74b, Fil73, GUW74, HS74b, Lel74d, McA74a, New74b, Nor74a, Wat74, dFP74]. 1974 [CG74, SM74b].

2 [Art72a]. 29th [BCdV^{+}70].

3 [Art73j, Tac73e, Tac73a, Tac73k, Tac73i, Tac73j, Tac73m]. 30th [Pet70a]. 347 [Ano71a]. 347-363 [Ano71a]. 363 [Ano71a]. 381 [Ano71b]. 399 [Ano72].

417 [DE73]. 435 [Ano74]. 4th [Pet70a].

5th [BCdV^{+}70].

’70 [Hod72].

A. [Del72a, Mey73h]. Abbildungen [Sch74c]. abelian [Ake74, Bar71c, Bas73f, Bro73a, Dem72d, FM71, Gri71b, Hau73, Rei71b, Rei71n, SM74a, War73, Mes72c, Sta73c]. abélien [Ray70-33]. abéliennes [Tat71]. abeliens [AdL72, Gat73a, Ray70q]. abgeschlossenem [Lor70g]. abnormally [Lew72a]. Abschätzung [GU71a]. abschliessende [Pop70a]. Absolute [Fel70, Kat70a, Eis71a, Gri73, ST71j]. Absolutely [Kap70, Bea70k]. Abstract [Bre71a, Fic72, Gro70a, Kre73a, PE73, Ste74a, Got71, Kel72a, Ōuc73, Top70i]. Abstracts [Yat72a]. abstraite [Hir72c]. Abstraites [Ler71]. abstraits [Hir74a]. AC [Luc73h, Jec71f]. Accelerated [KS70, You74]. according [Fra70a, Mal73, Mey73e]. Accroissements [Bre73, Bre71t, Bre72d, Dup74b]. accuracy [BS74a, Eng70]. Actes [Aku73, Bru71c, Sig71b, dFP74]. acting [JW73g]. Action [Puk72, Ray70d, Fár72, Gru70i, OW72, vdK73a]. Actions [Hec74b, KMT74b, Ati74i, BV73e, Bre72a, Car72b, CR72a, Hus74b, KR72, Kaw72b, Kaw72a, Kin72, KK72, MY72, Orl72a, Orl72e, Pet70c, Sch72e, Sch72d, Wan72, Wan74]. active [LRV70a]. acycliques [Art73e]. Adams [Mil74n]. adaptive [vdH73]. Addenda [Jón72a]. additifs [Dez74]. additional [OG74b]. Additive [BN73, Lam71a, Lam71f, ER70, Gal72a, Mon71i, Rya72, Top70g]. additives [Del70h, DMW73a]. adhérentes [Man73b]. Adic [But74d, Dem72d, Lan73b, HC70d, PŠ70, SD73, Van72, Yam74g, Yam74f, Kat73g]. adics [Sil70a]. adique [CR73b, Ver74]. adiques [Del71a, Ser73k]. Adjoint [Gra74e, Bun71, Day74c, Mei72, Mil73a, Pal71, Tak70k]. adjoint-functor [Day74c]. Adjointness [Gra74h, Eas74, Gra74k]. Adjoints [Ler71, Pul70]. adjunction [Kel74g]. Admissibility [Gol73, Noh71]. Admissible [Ler73, Sto70a, vdK73b]. admit [Kle72, Law73]. Advanced [BCC^{+}70]. Advances [Bec73a, Mey72h]. Affine [AS74a, Con72, Har70a, Mur72a, Ste74b, Con70, Hir71c, KMT74d, Kan71a, Wil70a]. Affinely [Mur72b]. Affines [Gro70d, Gro70g, Art73i, Ray70i, Riv72h]. after [Sch73c, Van72]. Aftermath [Wal72b]. Ahlfors [Ess74c]. Aléatoire [Fau70]. aléatoires [Art72b, Bre72c, Bre72d, Car70b, DC70, Del69, Del71c, Kat73h, MM74b, MM74a, MM74c, Mey74b, Mey74c, Mey74d, Bad70g, BC74a]. Alef [Grz70]. Algebra [BR73c, Ive73b, Jón72f, Pet70a, BGR73b, BE73, Cla71b, Coh70g, Hai73, Hic73, Liu71, Mur72f, Rin72b, RHD70b, Ste71a, Tak70b, Tak70m, Yam74d, Yam74k]. Algebraic [AS74c, AS71b, Bas73a, Blo73a, BV73d, BCC^{+}70, BGJ72, BG73, BF74, Coh74a, DeV72a, Deu73c, Eic70, Elg71, Hal70b, Har70b, Har70d, Hil71, Ive70d, KMT74a, Knu71b, Knu71c, Law72a, McA74a, Pop74b, Ran73b, Sta70l, Ste74c, Wac74a, Wag71, Zim72c, AG71c, AdK74a, AKW73, AS71a, Ban71, Bas73d, Bas73e, Ber71i, BV73h, Bom73, Buc72, Cla71a, Gri71a, cH73, iI71, KMT74h, KMT74j, Kar73c, Keu73, Knu71f, Knu71h, Kuj71a, Lic73, Lip74f, Mac73, Mat71c, May70, Puk72, Qui73a, Qui73b, Rob73n, Rob73m, Ste74b, Ste74a, Ste74e, Swa73, Tit74d, Uen74, Wal73b, Whi70, Zim72a, Zim72b, Zim72n, GU71b, McA74a]. Algebraic-geometry [Wac74a]. Algebraische [GU71b, LL74a, Wil70b]. algebraischen [Pop70n, Pop70o]. algebraischer [Pop70l, Pop74a]. Algebras [DI71e, Fer70e, GJ72c, Hof70c, Hof72b, JL70d, Jón72b, Kan71f, Rei71a, Ste70b, Ste70j, Tak70p, Wae71b, dlH72a, Adl74, All73, And71b, Ara71, AR73, Bar70, Beh72a, BZ71, BGR73d, BGR73f, BDF73, Bur74c, DI71b, Dom73, Gel74, Gol74a, Gui72a, HS73a, Hof70a, Hof70f, Hof70m, Hus74a, Jam70d, Jec71r, Jec71u, Kad70, Kan71j, Kel74f, Lof70, Luc71d, Mic72, Moo71, MP71, Mur72g, Omo74g, Pet70b, Ray70-30, Rei71b, Rei71c, Rei71d, Rei71e, Rei71g, Rei71i, Rei71h, Rei71l, Rei71n, Rei71m, Rei71o, Rei71p, Rei71r, Rei72, Ric74, Rin72a, RHD70e, Roi74, Sai72a, Seh73, Sin73, Smi72, Ste71a, Ste70k, Ste70n, Tak70a, Tak70c, Tak70f, Tak70k, Tak70n, Tak70o, Tak72, Tel71, Ves70, Wei73, Yam74c, Yam74e, dlH72g, dlH72b, dlH72c]. Algébre [Ill71e, Wag70, Fer70f, Fer70g, Lel70a, Lic74, MM72]. Algebren [BGR73d, Lep72]. Algèbres [Ber73, Dem70d, Gro70h, Kar70a, Leb73a, Lem74b, Lem74f, Lem74c, Dix71, Dix74, Fer70h, Gab71, Gat73a, Kah73, KO74d, Ray70o, Ray70-30, Sko73, KO74c, Eym71, Fer70e, KO74e, Lem74a]. Algébrique [Gro72f, Art72a, Art73d, Art73j, Bru74m, Che74c, GR71, Jou73b, Ray70-32, Riv72d, Wal74f, DK73a, Ray71b]. algébriquement [Mey72c]. Algébriques [Dem70k, Gab70d, Gro70h, GS70, JG71, Ray70b, And74e, Art73i, Dem71, Lel72d, Lod73]. Algol [Iga71]. Algol-like [Iga71]. Algorithm [Ber71f, Sch73-29, PBR71, Sch73-28, Ber72b]. Algorithmic [Eng71b, Sch71v]. algorithmische [Sch71v]. algorithms [Che70]. Algorithmus [Sac74a]. allgemeinen [End74]. Allocution [Lau70]. allure [Bre74a]. Almost [Car71e, Fin74a, Fin74b, Sei71, Cor74, Fin74c, Fin74i, Fin74o, Ken73, Yos71]. almost-product [Cor74]. along [Har70e]. Alphabet [Bon72b]. alternating [Ker71a]. alternative [FNSS73c]. alternées [FS70e]. amalgamation [Cas73c]. ambiante [Mou72j]. Ambits [Got71]. Ambrose [Pap70]. Ambrose-Kakutani [Pap70]. American [BK74]. Amherst [KMSS72a, KMSS72b]. among [And73a]. Ampère [Sty74]. Ample [Har70c, Har70d]. amples [Ray70r, Ray70s]. Amsterdam [Kui71a]. analog [PE73]. analog-generability [PE73]. analogue [Sch72b]. analogues [AR72]. Analyse [Lel73b, AdL72, CW71a, Far74, Lel70b, Lel71c, Lel72b, Lel74d, Mey70f, Poé74h, Car73b]. Analysis [Bac72, Ben71c, FNSS73g, GUW74, GL72, HC70d, HL74c, LLN74, Luc73e, Mor71b, Mor71c, Nac74a, Sei72g, Sil70a, Taa70b, Tro73a, AB72, Ben71i, Ces71, Ces74a, Dzh74, Fer74c, Gar74b, HO73, Har72b, Hel72, Her71d, How74, Kos70a, Kos70b, Kre73j, Luc73o, Luc73i, Luc73h, Luc73g, Luc73f, Oka72, PS72a, Poé74h, Sat73d, Sca71a, SP73, Sta73h, Taa70a, Tho74b, Väi71b, Van72, Zie74]. Analytic [AS71b, Beh74a, Coh74b, Gar72a, Gar72b, Gun71, Har71a, Har71b, Her71b, Iwa71a, ST71e, BS73a, BZ71, BŁ71b, Cla71b, Col73c, Fic74, Gil74d, Gil74g, Har70b, Kad70, Kaw73a, KS71a, Kow73, Kub71, Kuo71, Lel71b, Maa71k, Mar71a, Ram74b, Sch74e, Shi71a, Sic72, Sic74, ST71d, Siu71b, Sul71a, Thi70, Väi71a, Wae74b, Wen72, Wen74a, Coe70, Her71a, HS71]. Analytical [NS72, Dzh74]. Analytically [BD72a, BD72b]. analyticité [Cou73]. analyticity [Her71e]. Analytique [Ray71b, Bea71, Blo70, Blo72, Car74b, Coe70, Coe74, Col70, Lel72c, Maz71, Ray74, Sch71a, Sib74]. analytiquement [Dos71]. Analytiques [Cha73, Del72e, Del72f, Car71a, Cho73a, Cho73g, Col73b, Del71g, Dol72, Dou71, Exb70, Gom70, Hir71a, Kah73, Kan72, Laz73, Lel72c, Mar70e, Mar71c, Nor74d, Nov70, Rug70, Sko74]. Analytische [GR70, Gru74b]. analytischen [ZVC70g]. analytischer [LL74a]. analyze [ADT72]. and-spaces [May72m]. Anfangs [AH70c]. Anfangs- [AH70c]. Anfangskurve [Haa74]. Anfangswertaufgabe [Rau74]. Anfangswertaufgaben [AH70f, Rau72, Tau74, AH70a, AH70e, AH70g, Has72, Has74, Kre74c]. Angeles [Hof72b]. angle [Fuc74]. angular [Fri72]. Anhang [FK72c, GU71c]. anneau [Jen72j, Ray70d]. Anneaux [Ber71k, Ray70e, Ray70f, Jen72b, Jen72a]. Année [Lel73b, Lel70b, Lel71c, Lel72b, Lel74d]. Anneles [Ill71d]. Announcement [Bar71c]. annulation [Jen72j, Jen72c]. Anosov [God71b]. Ansätze [Sch71l]. answer [Gol74c]. Antitone [GQ71]. Antwerp [DK73b, Kui73a, KS73b]. Anwendungen [Sei72b, Pit72e, Pop70b, Pop70a, Pop70o]. août [Sig71b]. Appendice [Poe74i]. Appendix [Shu71a, Tit74b, Tit74a, FK72c, Poe74i]. Application [Bad74a, BKW74, Ben73, Ber71f, Coh74d, GS70, HMS74a, Ker71a, Lem74b, Mar70c, MS73b, Rev70, Tou74, dM70, Adl74, Bal70, Ben74b, Bre71j, Coh74c, Col74f, ESS74a, GM71b, Hir72b, Jan74b, KMT74c, Lel72d, Lip74a, Man72, Mok72, Mon71q, Par72g, TP71j, dM72, DM74, FL70, Ray70a, Ray70-27]. Applications [ALW72, AHiN74, Ati74c, Bad70a, Bas73b, BD72a, Bre71b, Cha73, DI71a, Del70a, Eic71a, Fug72a, Gar72c, GUW74, HMS74b, HMS74e, HN71a, HN71m, Ive70d, Jen72b, Jen72a, Jen72f, KT71, Laz73, Lio70, Lod73, Mar73c, MS73e, Meg71a, Mes72c, Mey73b, Mor71b, Nac74a, Par72i, Pet70a, Ray70g, Sha71a, Smi70a, Smo73, Taa70b, Tak70p, Tre73a, TP71f, Zag72a, Zyg71a, Art73f, Azé72, Boo74a, Bre71i, Bru74l, Buc72, Cai70, Col73b, Dob70d, DD74d, Dun72, Dzh74, ER70, Ess74b, Fer70f, Fri73c, Fuc74, Gou74a, HC70i, Inf71, Jec71u, Knu73b, Kos70a, Kos70b, Lel71a, MS73h, Mes72e, Mit74a, Moo74, MV70c, Mot73, Nob71, PT73f, Par72k, Pet73, Pie72, Poe74j, Poe74t, Rei71k, Sat73j, Sch72f, Sim72e, Sto71, Sug71, Taa70a, Tel71, Bre71n, Gro70d, GV72a, Poé74g]. Applied [Cur72, Kno73a]. appreciation [Whi70]. approach [BD74, Hoa71, Hop71, Kel72a, May70, Paw70, Tho70, Top70i, vW73]. approaches [Gan71d]. Approchées [Bru74b]. approchés [DDD71]. Approximate [Hen73, Mos70, Nas74, Rei71c, Rei71d, Rei71e, TP71a, Yev70, BD72e, Coo71b, Rei71i]. approximated [Luc73p]. Approximation [And70, Ang71, CS70, Cha70, Col72, DeV72b, DeV72c, Gre74, Hol72b, Mor70, NW74, Nov74a, Sha71b, Sha71c, Sha71j, Tem71a, Tre73e, AJ73, Col71, Col74a, Col74b, CG71b, CG71a, DeV72d, DD74d, Fic74, Fre74b, Gam71, GW74a, Gar72c, Hol72e, May72c, PBR71, Sch73-33, Sha71d, Sha71e, Sha71f, Str74a, Sum74, Top70d, Wac71b, Wid71a, dB74, rd74e]. approximations [AH70b, Bom73, Tho73c, Wid71b, vdC73]. approximationstheoretische [AT73]. approximative [Jaw70]. April [Fil73, GG72, Hor70b, Hor71, HS71, Pet70a]. Äquivalenzsatz [Kre74c]. Äquivariante [tD72a]. Araki [Kri70]. Arbitrary [MV70d, Aur74, Ben74a, Cop71b, KMT74h]. arbres [Car73b]. arc [Bel74a]. archimédienne [Ray74]. areas [OG74b]. arguments [Gal72b, Ler73]. arising [Tac73b]. Arithmetic [Bas73a, Cas73b, GG72, HR74c, Mon71a, MS70d, Tro73a, Ben71e, Ber72b, ES72b, Gai72, Goo73a, Hen71, Hir74b, Jec71d, Luc73h, Par72a, Rya72, Smi72, Sub72, Tro73b]. arithmetical [Rao72a]. arithmetische [Sch74c]. Arnold [Bec73a, Bri73]. Art [Kol72]. article [Mey71d]. Artin [AR73, Del73n, Gol74c, Ray71a]. Artinian [Tac73a, Mat73b]. Aspects [Dev73a, Fra70a, FPC74, Hof71, Kha71, Maa71d, Sha71f, Wal72e, Fra70a]. assertions [Lon71]. assignment [dB71a]. Associated [Mes72c, Bea70e, BdS74, Ell73, EG74, Gil74a, Mes72b, Sat71, Spe74, Tay74a]. associative [Mal74, Sha71m]. Associativity [Sta70b, Sta70h]. Associé [Wag70]. associée [Ber74h]. associées [Car70b, Leb73a]. Associés [Bad74f, DC70, San74, Sko74]. Assortment [Cas73a]. assumptions [Bea70l]. astronomical [Ste70o]. Asymptotic [Bha70, Chi73, Fin74c, HO73, HL74b, WW74a, Coh74e, Cov71, Gol73, Kat70c, Kri70, NS71, Pit72g]. Asymptotically [WW74a]. asymptotique [Vau73]. asymptotische [Pit72g, Pit72a, Pit73]. Atiyah [KR72]. Atkin [Dwo73]. attached [Maa71u]. attack [Mar73b, Qui73c]. Attention [Bic73b]. attraction [LKR73]. au-dessus [MM74d]. AU-spaces [BD72c]. Aufgaben [AG73b]. auflösbarer [BGR73d]. Aufs [Bae74]. augmentation [Coh72a]. August [Bas73a, Bas73b, Bas73c, Chi71a, DK73b, Gra71b, Kui73a, KS73b, Kui71a, MR73, New74b]. Ausgangsdaten [AG73b]. ausgeschlossenen [Sch71e, Sch71n]. aussi [Mor71a]. Austin [Bet74b]. Australian [Hol74b, New74b]. autoadjointes [Ber73]. automata [Sch71j]. Automaten [Sch71j]. AUTOMATH [dB70]. Automatic [Cha74a, LLNS70, Kow70, Kre70a, MW71, Paw70, RW70, Wan70]. automorphes [God71c, Rob73k]. Automorphic [Bai70, Gar74a, JL70a, Jac72a, Wei71e, Kub71, Lan70]. Automorphism [ABK71b, Maa71a, FGQ71, Hai73, Hau73, Kan71a, Law73, Mur72a, Tak70e, Tak70l]. Automorphismen [ZVC70a]. Automorphismes [Dem70a, Lic74, KO74d]. Automorphisms [Mur72h, Bro73a, HS73a, Mat70, Puk72, Smo71j, Ste70p, Tak70h, Tay73, ZVC70a]. Autre [PMN70]. Autres [CW71b]. aux [Bad74f, Del74a, DDM70, Gro70d, GS70, GV72a, Jen72b, Kré74a, Lem70, Lio73d, Mar70c, Mey73b, Nor74c, Ray70g, Ray70a, Rev70, Sko74, Spi74b, Tem71a, Zyg71a, Zyg71e, dM70]. Avant [Man73a]. Avant-Propos [Man73a]. Average [BW72]. averages [Han70]. Averaging [Fin74d, Sei71, Yev70]. Ax [Ami72]. Axiom [Jec71a, Jec71b, RW70, dB71a, Dev73b, DJ74h, Gar74b, Kun73]. axiomatic [Bre71h, Bre71m, Fug72b, Han71a, Hoa71]. Axiomatique [PP74b, Bru74j]. axiomatizability [CD72b]. Axiomatization [BS73c]. Axioms [Jec71c, Sch73f, Aul74, Con73, Kow70, Sch73i].

b [DeV72g, Che74f]. B. [Mos72]. Back [BK72a, CR73a, Del73a, DJ74b, Ill72a, Leh73a, MS70a, Rob73a, Ser73a, Tit74c, Zuc73a, vL71a, vL73a, Pop70p]. Background [Väi71b]. bad [Mas74a]. Baer [Ste70c]. Bahnenräume [BGR73d]. Baire [Top70i]. Baker [Ser71b, Wal74a]. Balanced [DR72, De 73]. balayage [Bre71b, CD74]. ball [BP72]. Banach [LT96, AdK74b, Bur74c, Cla71b, DC74, Dom73, Dun72, Exb70, Fer74c, Her71b, Her74c, Hou72b, Hus74a, Kah73, Leb73a, Lel71a, LT73, Nov72, Nov74a, Rei71c, Rei71d, Sch74e, Sha71a, Sha71c, Tre73e, Ves70, Wag70, dlH72a, dlH72b]. Banach-valued [Dun72]. banachique [Car71a, Maz71, Her74d]. Bang [dV70]. Bank [Sch71t]. bar [BV73a, Luc73o, Luc73a, Luc73j, Mac70, Sta70c, Sta70d, Sta70n]. Barban [Mon71l]. barinduction [Sca71j]. barrelledness [De 73]. Barsotti [Mes72b, Mes72c, Mes72h]. Base [Bor72c, HN71b, Rob73n, Art73h, Art73g, Art73f, Aul74, Bab73, Nov72, Pet74a, Ray70j]. base-r [Bab73]. Based [DD74a, Sim72a, BV73g, Csö73, Sch73a]. bases [Tul70b]. Basic [Ber71b, Fro74, Mey72e, Omo74a, Omo74b, Sch74d, Sch73m, Ste70d, Top70a, Top70b, Wei71a, BD72b, CM74e, KMT74g, Mon71o, Top70f, Pit72d]. Basis [Eic73, Wac71b]. Battelle [Bas73a, Bas73b, Bas73c, Pet70a, SJS73]. be [Bre72b]. beams [Kom72a, Kom72c]. bear [Ste70e]. Begriff [Sch71g, Sch71s]. Begründung [Sch71v]. Behandlung [AT73, AT74, Gen74]. behavior [Fri72, HC70a, Wal73c, WW74a]. behaviour [Bre71d, Coh74e, Set72, Woo74a]. Beispiele [Pop70a, GU71h, GU71n, GU71o, Pop70b]. belonging [Mur72k]. Bemerkung [Opf74]. Bemerkungen [Pop70a, Has74, Lor70a, tD72a]. Bendixson [Sai71b]. Benefits [MS70b]. Benveniste [Mey73h]. berandeten [Gru74b]. Berechnung [Gek73]. Berge [Cun74]. Bernoulli [Orn70, Smo71i]. Beschreibung [BGR73d, ZVC70f]. Besicovitch [Bru74d]. Besonderheiten [Pop70a]. Bessel [GK72]. Best [Hol72b, Sha71d, Sha71e, Ben70, Hol72e, Mil73b, Wid71a]. Betti [Fár72]. Between [Boo74b, AKW73, Con73, HB70, Mes72h, OG74a, OG74l, OG74q, OG74r, Sig71a, Tur73a]. Beurling [Gat73a]. Beweismethoden [HRS^{+}73a]. bi [Ars72, Sur72]. bi-orthogonal [Ars72]. bi-unitary [Sur72]. bibliographie [MM74b]. Bibliography [Jam71, MM74b]. Bicategories [Gra74f]. bichromatiques [Las74]. bicubic [BBF74]. Biemann [Sch74d]. Biextensions [Gro72b, Gro72c]. Bifibration [Bun71]. bifunctors [Sha71m]. bifurcate [Jos73]. Bifurcation [Lan73a, Ree73, Sat73c, Sat73b, Sat73i, Her73, Jos73, Tho71a]. Bigebras [Hof70d, Hof70l, Hof70n]. bilinear [Ade70, Maa71a]. Billingsley [Sch73n]. binaires [Bru74o]. binärer [Gey74]. Binary [But71, But74c, SM74a]. Bing [Bur74a]. Binghamton [McA74a]. biochemical [Wal73c]. Biology [SJS73, Nic73, Tho73b]. Bipolaires [Poé74e]. Birational [AS74b]. Birthday [Pet70a]. Bishop [GM72]. Blaine [Ros72]. block [Lor74, McG73, Sto72d, Tay74a, Wel71b, vL74p, vL74k]. Blocks [Bra74, Mic72]. blocs [Che74d]. Blowing [Lev71e]. ‘BMO’ [Mey73d]. BN [Tit74f, Tit74n]. BN-Pairs [Tit74f, Tit74n]. Bochner [Bad70e, BB70, Lan70]. Bockstein [Moo71]. body [AC74, GK70, Hay74, Jan74a, Mil74p, Smi74, Wie74]. Bohr [Hur74]. Bois [Art72a, Art73j, GR71]. Bois-Marie [Art72a, Art73j, GR71]. Bombieri [Lel72d, Mon71b]. Boolean [Adl74, But74b, Gui72a, Jec71m, Jec71r, Jec71u]. Boolean-valued [Jec71m, Jec71u]. bord [Sch70g]. Bordered [Fay73a]. bordism [HO72, Mil71]. Bordismen [BtD70e]. Bordismen-Homologie-Theorie [BtD70e]. Bordismengruppen [BtD70d]. Bordismus [BtD70a]. Borel [Kan71c, Loe74a, Sch72b, Sch73o, Top70i]. borné [And70]. bornée [Bru74a]. bornées [AdL72, Lie74, Rud73, Sib74]. Bornés [Che74g, Fer70b]. Bornological [HN71g]. Bornologie [Hir71a, HN73]. Bornologies [HN71b, HN71c, HN71m, HN71n, HN71d, HN71d]. Bornologique [HN71g]. Bornologiques [HN71a, HN71f, HN71h, Jac72b, Laz73]. both [Bre71g]. Bott [Sch72b]. bound [JW73e, Mon71j]. Boundaries [Bre71p, Kaz71]. Boundary [Beh74b, Bra72b, Gil74a, Kre74b, Las70, O’M74, Sim72a, AJ73, BD74, Bre71a, Bre71d, Bre71r, Ces71, Cod71, Eva71, GP74, GL74, Gor71, Gou74a, GH73, Hab74, Har73, Hes74b, Hop71, HK71i, Jan71, Kaz71, Kom72a, Kor70, Lio73d, Mei72, Net74, Nob71, Per74, Sat73g, Sib71a, Tak70e, Tay74c, Vel72, Wal74m, Wen72, You74, vdC73]. boundary-value [Mei72, Vel72]. Bounded [Har74i, Hsi73, Iwa71b, Kan71d, Tre73e, BMN74, How71, Kno72, Luc71b, Sum74, Tre73d, dlH72b]. Boundedness [Car74c, Mor72]. Bounding [Sch74e, Gar73]. Bounds [Coo71b, Baz74, Eis71a, Goo71, Nar70, Wal72a, Wid71a]. Bourbaki [Ano71a, Ano71b, Ano72, Ano74, DE73, Tem71b]. Bouts [Poé74a, Poé74c, Poé74d, Bor72a]. Bowdoin [DS70a]. Braids [Gol74c]. branch [IP70]. Branching [Lou73, Nag72, Gre73b]. Brasil [Nac74a]. Brauer [DI71b, Yam73, Yam74a, Yam74j]. Breaking [Zee71]. Bref [Dez74]. Breiman [Smo71l]. Britannique [Bru71c]. broadening [vW73]. Brouwer [Sch71c]. Browder [Fan70]. Brownian [DMW73b, Kni71, Kni74]. brownienne [Far73]. browniens [Car71b]. Bruign [vL74f]. bruits [Dez74]. Bubnov [Ste74g]. Büchi [MS70d]. Buffalo [Fle70]. building [Tit74d]. Buildings [Tit74e, Tit74f, Tit74i, Tit74j, Tit74h, Tit74g, Tit74k, Wag73]. bumps [Sto70h]. bundle [AB72, Omo74l]. bundles [AS71t, Gia72, Gre73a, Gun71, HR71, HMR71, Oka72, Sta71d, Sto72d]. Burkholder [Mey72b, Rao72b]. Burnside [Qui73c]. butterfly [Woo74c]. bzw [GU71n].

C [DeV72g, Lat72, Van72]. C. [Ben74a, Che74f, Del70g, DMW73a, Har70d, Lat72]. cône [Lel73a, Mok72]. cônes [Mok71]. Calcul [Ber71m, Ber74g, LM70, LT70, MM72, Poé74a, Fer70c, Fer74d, Glo70, Gol74b]. calculated [Smi74]. Calculating [Mil74a]. Calculation [VCLN74, Coh70e, Her73, Yev70]. Calculations [Bir73, Mil74b, Mil74e, Mil74l]. Calculs [Ber71n, Ber71l, CR73b, Jou73a]. Calculus [Kel74c, Yam74p, Coh74m, Fit72, Kel72c, Lam72, Luc71f, Luc73a, Luc73b, Wae74a]. Calibres [Del72c]. Cambridge [MR73, MR73]. Cambridge/England [MR73]. Campbell [Coh74c]. can [Bre72b, Mar73b]. Cancellation [SE70c, EH73, Fuc72, HMR71, Lip73, Wil74c]. Canonical [Rei71f, HK71a, Mat73c]. Cantelli [Csö73, Sch73o]. Capability [Del72i]. Capacitaires [Del72i]. Capacités [Del71f, Del72d, Del72e]. capacities [Ang71, Bre71b]. Capacity [Fug72b, Gar72b, Gar72a, Geh74]. Capital [BH74b]. Caractères [Spr74]. Caractérisation [Aze70a, Gro70c, Poe74j]. Caractéristique [Ber74i]. caractéristiques [Art73a, Art71, Hae73, Mar74b, Ver74]. Carathéodory [Bru74d, Tit71b].

Cardinal [Jec71d, Jec71e]. Cardinality [BW74]. Cardinality-dependent [BW74]. cardinals [AR72, Dev73j, Dev73h, Dev73i, Dev73k, Dev73o, Dev73p, Dev73r, Dev73t, Hec74a, Jec71j]. Carleson [Ass71]. carré [DD71b]. Cartan [Gol74b, Gro70g, Nov73, Ste70f, Ste70g, Wil73c]. Cartesian [Lam72]. Cas [Art73b, Aze70b, Ray70h, Ray70j, Ray70i, Ser73c, Ser73h, Mey70c, Mey73d, Rev70, Sch72b, Zyg71b, Zyg71h, Zyg71g]. cascade [Rén73]. cascades [Sha71l]. Case [HLM71, Bre71g, Coh72c, Coh72e, HC70c, Kow70, MS73c, Sum74, Wei71b, Wei71j]. cases [Ber71h, MS73i]. Cassels [Lor70e]. catastrophe [Woo74b, Woo74c, Woo74e, Woo74d]. Catastrophes [WP74b, Was74c, WP74c, WP74d, WP74e, Woo74a]. Categorical [Cla71a, Coh70g, Grz70, Hag74, May72d]. Categorically [DS70a]. categoricity [Cut72]. Categories [BGvO71, GU71b, Gra74g, Gra74h, KLLM72, Ste71b, Ill71c, Ill72b, Ler71, Sch72a, Bar71b, Day70, Day74d, Day74a, Day74b, Dub70a, Dub70c, FM71, GU71k, GU71m, Gra74e, Gra74d, Gra74j, Gri71b, Gri71c, Gru70f, Gru70g, Hof70p, KS74b, Lam72, Mal74, May72k, Mei74, MP71, Pul70, Riv72a, Świ70, Tac73i, Ulm71, Van71, Riv72g, Ver72b, GU71i, GU71j, Riv72b]. Category [Dub70d, Gra71b, Gra71a, Gra74h, Kel74i, MAB^{+}70, Fro74, Hof70c, Hof70e, Hof70o, HMS74c, Pal71, Str74d, vdK73b, Kel74i]. Catherine [Dol70b]. Cauchy [BS73b, Col73c, Gar72d, Lev73, Lie74, Ōuc73, Roo74]. cause [Mar73b]. Cayley [Kan71b, LR72]. Cebysev [LZ73]. Cebysev-Entwicklung [LZ73]. Cech [Dob70a, Dob70e, Dob70d]. Celebrate [Pet70a]. celestial [Sma71a]. Cell [Cha74b, Gre71b, Bin74, Ray71d, vdK73m]. Cell-like [Cha74b]. celui [Hei74a]. census [BP72]. Center [Bas73a, Bas73b, Bas73c]. centers [Mic72]. Central [DI71b, PS72d, Sta73a, vdK73l, Bor74, KMT74c, Sta73f, Was71, vdK73p, Hey70l, vdK73p]. centrale [MM72]. centralizer [Tac73d]. Centralizers [Iwa70, KP72]. Centrally [Phi73]. Centre [Aku73, Cur72, Gro70g]. Cerf [Poe71b]. Certain [Bru74i, Coh74g, PT73i, Rei71g, Rei71h, Ban71, BD72e, Coh74r, Coh74s, EG74, Fre71a, HB70, HN74, HK71i, iI71, Kha71, Kur71, Mil73a, Moo71, Nov74a, OG74q, Ped73, Rao72a, Sat71, Str71, SW74b, Wal74j]. Certaines [CW71a, Dix74, Lem74b, Lem74f, Car70b, Kah73, San74]. Certains [CR73b, CW71a, Eke70, Rav71, Aut72, Dup74b, Hei73a, Hei74a, Ray70-27, dM72]. CH [DJ74m]. Chain [And71a, Rat73, Gus73a]. chained [Col74a]. chains [KŌ73a, Rah74]. Champs [Aut72, Sad74, Ber73, Car72c, Car74b, God74a]. chance [Sch71j]. Chandra [Coh74b, Kan71h, Van72]. change [Rob73n]. Changement [Lio73c, Lio73e, Lio73i, Art73h, Art73g, Art73f]. Changing [Kro74]. channel [Iha73]. Chapitre [Bru74l, Bru74k, Bru74m, Bru74o, Bru74p, Poe74i, Bru74j, Bru74n]. Chapter [Lip74e, Lip74c, Lip74d, Lip74g, Yam74b, Lip74b, Lip74f, Poe74i]. char [HC70f, HC70g, HC70h]. character [GM71a, HMS74d, Mon71h]. Characterisation [Bar74c, Rob74]. Characteristic [Bec74b, Bry74a, KK72, Mar73a, Rob73h, Rob73i, Sán71, Sch74f, Ahm71, Ber74i, Bot72, Gia72, LW72, Pop70p, BtD70b]. characteristics [Kru73]. characterization [Ben71a, Cla71b, For74, KŌ73a, Mas71]. Characterizations [BL74b]. Characters [Bra74, Spr70a, HC70a, HC70b, HC70c, HC70g, Hir72a, Maa71h, Sil70b, Sil70f, Wil73c]. Charakterisierung [GU71d, Sch71d, Sch71c, Sch71e, Sch71b]. Charakteristik [Pop70g, Pop70f, Pop70p]. Charakteristische [BtD70b, BtD70c]. Chern [GM71a, Ill71f, LB70b]. Chevalley [Bor70, Iwa70, Spr74, vdK73l]. Chirurgie [Lat72, Mey73c]. choice [Con73, Jec71a, Par72l, Kun73]. Choquet [Bre71r, KN73, Mat72]. Chow [Moi74]. chromatic [BE74, Eis71a, EHR73, Kro72, Ste74h]. Chromatically [Bar74a]. chu [Hey73, Hey70a]. Chu-Dualität [Hey70a]. Chung’ [Ass71]. Chvatal [Ste74i]. Circle [Orl72a, Bre72a, MY72, Sch72e, Tay74j]. circuits [Kle72]. circulaires [Mar70b]. Circular [Hen71]. City [Bro71a, CFK71]. Class [RU73, Shi71b, Sta73i, Zim72a, Zim72b, And74a, BGR73f, Bro73b, CM71, Gia72, Gou72b, HO73, IW73, Kaw73b, Kel74d, Lip74a, Lor74, OTF70, Puk72, Set72, Wil70f, Wil72]. Class-Numbers [Sta73i]. Classe [Ber74h, Geb74, Lel74b]. Classes [BC74b, Cos74, Ill71f, LB70b, Mar74b, Ste74c, Tat71, Zag72a, Azu73, Bec74b, Bot72, BtD70b, Car70a, Cho73a, DeV72b, GM71c, Hae73, Her74e, Hir71c, Jec71i, Lip73, Mou74, Sch74e, ST71k, SS70, Top70r, Zab71, Zag72c, Zag72d]. Classical [Bas73a, Bre71c, Bre71d, Bre71f, Bre71e, Gru70a, Her71c, LT73, Luc73e, Mon71c, dlH72a, dlH72b, dlH72c, Bas73d, Ber70a, Bre71i, Bre71o, Dan74, Din71a, Kre73b, Luc73o, Luc73i, Luc73h, Luc73g, Luc73f, Luc73p, Sca71a, Sca71b, Ste71d, Wei71b, dlH72g, Bre71g, LT96]. Classification [Dem72a, Gro70c, Kom72a, Pop74b, ABK71a, ABK71f, CM72, Con74, HR70, Mou72h, Roi71, Sta74b, Uen74, Wal71b, Was74c]. Classifying [Bec71d, Gir72, Sto72a, BV73e, Wil74d]. classique [Art73b, Glo70]. Cleveland [HLM71]. Clifford [Kar70a, Lof70]. clique [HS72]. cliques [Mey74a]. Closed [BC74b, KM72, Rei71i, Aus70a, Dav73, Day70, Day74d, Day74a, Day74b, Dom73, Glo73, GM71b, Har74e, Hea74b, Lam72, Lew72b, Mat73d, Wae74b]. Closedness [ST71b]. Closure [Pie72, Mat73j]. clubs [Kel74h]. clusters [Wie74]. co [Dom73, KMT74c]. co-dimension [Dom73]. coût [Pet74a, Lio73a]. coûts [Pet74a]. Coalgebras [Pet70b, Moo70]. cobar [Smi70a]. Cobordism [Boa72, BtD70h, Gia72, Ker71f, Lan72, Orl72b, Sch70h, Sto72a]. Cobordisme [Kar73a, Poe72]. coboundary [ST71b]. cocycles [PS72a]. Codes [vL71c, vL71e, vL71g, vL71h, vL73c, vL73e, vL73g, vL73h]. codimension [AS71q, BV74, cH73, Ker71f, Lut74]. codimensions [Har70f]. Coding [vL71b, vL73b, Iha73]. codissipatifs [HR74b]. coefficient [Fin74j]. Coefficients [SD73, Bai70, Bor74, Bra72a, Del70g, DDDM70, Dol70b, Eis71a, Fer74a, Gil74d, Gil74g, Kan73b, NS72, O’M74, Sim72c, Tho74c, Wen74b]. Cofaserungen [tDKP70a]. Cofibrations [May72e, tDKP70a]. Cogebras [Hof70b, Hof70e]. Cohen [HK71a, Fel71a, HK71a, Hoc73, Mat73a].

Coherence [KLLM72, Kel74f, Lap72a, Lew72b, CLL73, Kel72a, KM72, Kuh74, Lap72b]. Coherent [Hig74, ST71e, Siu71b, ST71k, Thi70]. cohérentes [Kan70]. cohérents [Dou71]. Cohomological [Coh72f, Dob70a, Dob70b, Gru70b, Gru70c, Dob70c, Git72, Gru70k]. cohomologicas [Git72]. Cohomologie [AGV72, Art72a, Art73a, Art73d, Art73j, Ber74j, Ber74i, Blo70, Del73c, Del73b, Dem70k, Ill72c, Ill72b, LB70d, LB70c, Sch70a, Ser72, Ser73c, Ser73e, Ser73b, Ser73j, Ver72a, Ver73, Art73b, Bar74b, Ber74h, Ill73, Kat73b, KO74f, Nor74d, Rob71a, Ver74, Ser73d, Ser73h, Ser73i]. Cohomologies [God74a]. Cohomologique [Ray71c, Art73c, Art73i, Gro72d, Jen72i, Jen72d, Kat73a, Nor71, SD72]. Cohomology [Bar74c, Coh72b, Han71b, Mar71a, MM74g, Nar70, Rin72a, Sat71, Zab70, Ati74k, Ber74e, Ber74i, BG73, Dem72d, Gar74a, GM71c, Gru70d, Kom73e, Lic73, Mil74c, Sch70a, Ser73e, ST71k, ST71j, Ste71a, Tel71, Zab71, dlH72g]. Coincidence [Fre74a, Sch73c]. colimit [FM71]. colimits [BV73b, Gri71b]. Collaboration [Har70d]. collapsing [Jec71j]. collectionwise [Boo74a]. College [DS70a, GL72, Hor70b, Hor71, Ser73d]. Collinear [Sch73b, Sch73c]. collinearity [Sch73g]. collineation [Ost70b]. Collocation [DD74a, Esp74, DD74d, Sch74a]. Colloque [Vid74]. Colloquium [LLN74, Moi70]. color [BS71]. coloration [Ber74c]. coloring [BS74b]. colors [Chu74a]. Columbus [Pet70a]. combinaisons [Dez74]. combinations [vL74b]. Combinator [Sta73h]. Combinatorial [Hof71, Sul71a, vL74a, Bry74c, Dev73c, DJ74c, Jan74b, Hol74b]. Combinatorics [BH74b, WWS72a, Kle73b]. combinatory [Lam72]. comma [Gra74j]. commande [Lob72]. Commentaires [Del71f]. Comments [Gru70e, Hol72a, Hea74b, Mró70]. Commutatif [Ser73c, Ser73h]. Commutatifs [Ray71c, Car71d]. commutation [FM71, Tak70a]. commutations [Mar70a]. Commutative [BR73c, CW71a, DI71e, Dob70a, Ive73b, Kap73, Ser73e, BGR73b, Cla71b, Dom73, Hof70f, Hof70m, Moo70, Zab70, Ser73i]. commutativity [KMT74i, Sta70i]. commutator [Mis74, Pop70e]. commutators [Bro73b]. Compact [Ati74a, Bac72, Hey70c, Hof70l, HMS74e, HS70, KMSS72a, KMSS72b, Pop74b, Sto70k, Yam74l, AB72, Aus73b, Aus73a, Bha70, Bro73a, BDF73, Cha74d, Chu71a, Dev73t, EHR72, Geb74, Gun71, Hey72, Hof70i, Hof70k, Hof70j, HMS74b, HMS74c, HMS74d, KN73, Law73, Ram74a, Rei71b, Rei71r, Ric71, Sch72b, Sic74, Tel71, Uen74, Vau73, Wel71a, You70, dlH72g, dlH72b, LB70l]. Compactes [LB70l, BB73]. Compactification [Meg71c, Meg71b, Ray72a, Bre71j, Mey74g]. compactifications [Hur74]. compactifié [Meg71f]. Compactness [Kee70, Top70c, BF74]. Compactological [Luc71a]. Compacts [Aze70d, Bor72d, AdL72, Car71c, CW71g, CW71f, Gat73a]. comparables [Mok70a]. Comparaison [Art73b, Del73d, Dez74]. comparative [SW74b]. Comparison [Bre71g, GM71c, Moo74, Nak73, OG74a, Rob73m, Sch73p]. Compatible [Mat73e]. complémentaire [Che74c, Fou74, Rob71a]. complémentaires [Jen72d]. complementary [AM71, Vel72]. Complements [Gro72c, Ray70a, Che74e, Eym72a, Rob73b, Rob73c, Bei71, Che74e, Rob73b, Rob73c, Ray70k]. complet [Ber74c, Ber74d, Fou74]. Completant [Luc71b]. Complete [Dub70a, Fer70e, BE74, CT74, Din73, GM71b, Jec71r, Kei74, Pri73b, Ric71, Sch73e, Sin73]. completely [Jon74]. complètement [BI70c, Buc73b]. Completeness [Bea70a, Dub70b, Fra70a, MS70c]. complétes [Kat73b, Del73c, Fer70h, Fer70e]. Completion [Nov74b, BK72p, BK72o, BK72d, BK72j, GH74b]. Completions [AK70a, BK72m, Mat73f, Gon74, BK72n, BK72c]. Complets [HN71h, HN71l, Jen72b, Mey74a]. complétude [Fra70a, Sun70b]. Complex [Bro71a, Gan71b, Her71d, Hor70b, Hor71, Kre73c, Pop74b, Rob73d, Rob73e, Shi73a, Sta73i, Sto70k, Wil73e, Wol71, AS74f, Aur74, Ben74a, Bom70, Car72b, Dzh74, Fer74c, Gir73b, Ham71, Her70, Hir72a, Ill72e, Kor70, Lem71, Luc71h, Moi74, Nar70, Pet72, Sib71a, Uen74, Wil73d, dlH72g, rd74c, Ill71h, Ill72d]. Complexe [Aku73, Cho73c, Ill71g, Ill71h, Ill71i, Ill72d, Lit74b, Car71a, Lel72c, Lit74c]. Complexes [Ill71f, Ill72b, Nor74a, Tit74l, Bro74b, Cho73g, DM74, Dol72, Lel71a, Mar71a, Nor74c, Nor74d, Nov70, Sad74, Sta70g, Sta71a, Tit74m]. component [AC74]. components [Mat73d, vdK73b]. Comportement [Vau73]. composables [BS74c]. composition [May72p]. Compostela [Vid74]. Compound [vdC73]. compte [Dez74]. computability [PE73]. Computable [MS73c, Gat73b]. Computation [BKW74, Bro74a, HK71i, Sil70b, Hum73, Luc73c, Wac74a, Luc73c]. Computational [CG74, Zim72c]. computations [ABK71g, DS73, Eng70, Hay74]. computers [Fra74]. Computing [Sch74f, Bal70]. con [DJ74m]. concaves [Ram74c]. Concentration [Bad74b, Bad74c, HT74]. concept [HR74a]. concepts [Ber71b, Dub70b, HN71k, Wei71a]. concernant [Mey74a]. Concerning [McA70, McA71, Kur70a, May72o, Mon71e, Tak71c, Tit71b]. Concluding [Ben71b, Sch73d]. Conclusions [OG74b]. Concordance [ABK71b, Kau74]. Concordance-Homotopy [ABK71b]. concrete [Fre70]. condensation [Dev73b]. condensers [Geh74]. condition [Bet74a, Fou74, Hei73a, Hön74b, OS70, Ste70l, Tac73d, Tak70e]. Conditional [PP74a, Sch73q, Smo71a, Smo71c]. Conditionally [Gui72b, Sch73e]. conditionals [dB71b]. conditionnelles [dSL70]. Conditions [Bad74b, Bad74c, Con70, GV72a, Ill71a, Ill71c, Ing71, Kur71, Bru70, Dwo73, Eva71, Fin74c, Gil73, Gor71, Gru73a, Jen72c, Kom72a, Kre74b]. conduction [LRV70a]. cone [Gow74]. Conference [ALW72, AHiN74, BH74b, Bas73a, Bas73b, Bas73c, Bec73a, Bet74b, CFK71, DF74b, Fil73, GUW74, GG72, HLM71, Hod72, Hol74b, Hor70b, Hor71, HMR73, HS71, Kom73c, KMSS72a, KMSS72b, McA74a, New74b, Pet70a, SM74b, Suc70, Fle70, BR73c, ES72c, GW73, GL72, Mor71b, Wat74]. configurations [Hig74]. confluent [Lel74a]. conformal [Gai74, Wal74j, Wil74e]. Conforming [Cia74]. confunctor [Hof70f]. Congruence [Jón72c, Kat73d, Wil70f]. Congruences [Ser73f, SD73]. Conjecture [Ami72, Leh73b, Chu73, Del73r, Har74d, Mar71e, Ros74, Sch73-31, Ste74i, Tac73b, Tak71b, Tit71b, Ass71]. Conjectures [Tac73b, EE73, Kar73c, Mon71e, Rat73]. Conjugacy [Car70a, SS70, Ste74c, App73, Lip73]. Conjugate [MK70, BD74]. connected [Cad71, Cer71, Ken73, Law73, Lan72, vdK73b]. Connection [Oku71, Was71]. Connections [Bas73a]. Connectivity [GH71, AM71, Cap72, Kee74]. connexe [Lat72]. Connexes [Lem74a, Lem74e]. Connexions [Del70b, LB70e]. connexité [Ray70n]. Conséquences [Ber71n, Luc73i, vdK73o]. conservation [CM74g]. conservative [Rob71b]. consistance [Kri71]. Consistency [Jec71f, Luc73c, Luc73e, Dev73b, DJ74h, Sca71h, Sca71i]. Constant [Deu73a, Fin74j, Kan73b, Sim72c, Tho74c]. Constantes [Del73l]. Constantinescu [Bre71j]. Constrained [Zie74]. constraints [BD72f]. Constructibility [Dev73a, Dev73d, Dev73b, Dev73q]. Constructible [Fel71b, Jec71g, Dev73e, Dev73m]. constructibles [Art73d]. Construction [Art70, Aze70c, Gab70a, Ive70a, Knu73a, Ray74, Top70d, AHV74, BV73a, Geo74, Han73, Lac73, Mac70, May72d, Smi70a, Sta70c, Sta70d, Sta70j, Sta70n, Sto70c, Sto70b, Top70a, Top70b, Bru74l]. constructions [Cla71a, Goo73a, MM74e, vL74p]. Constructive [Gil74b, Sco70, Bau71h, Gil74e, GM72, Sta73h, CG74]. Contact [Mar71d, Lut74]. contain [Gon74]. containing [Fin72, Rah74, Sca71j]. Contensou [Mar70a]. contents [dlH72d]. context [Han73]. contingent [Kik71]. continu [MR72, Mey73d]. continua [Bel74a]. continuation [Col73c, Her71b, Mil73a, Sic74]. Continued [Fra74, Zim72d, GM71i]. continues [Buc73b, Mey70i]. Continuité [Che74g, Hei73a, Hei74a]. continuity [Fel70, Gri73, Loe74b, Top70j]. Continuous [Bec70, Bec71a, Bec71b, Bec71c, DeV72c, Gra73, Gui72c, Gui72i, Par72b, Par72d, Par72c, PS72d, Sco72, You70, Bea70d, Bea70k, Dol70c, EHM70, GU71i, GU71j, Gus73a, Hag74, Itô72, Jac70, Kni71, Kur70a, McA71, Mic70, PS72c, Pro74, Sei72h, Sim72e, SR73, Sei72i]. continuous-time [Jac70]. Continuously [Kel74b, LRV70a]. continus [Ber73, Del70g, DDDM70, Dol70b]. continuum [Cur74, Dev73b, Dev73d, Jec71j, Kno73a, Par72d, Tie72]. Contrôlable [PMN70]. Contrôle [Eke70, Lio73b, Lio73a, Lio73d, dM70]. Contractibilité [Mou72a]. contractible [Kin72]. contraction [Nad70]. contractive [Fra70b]. contraintes [Cés74b]. Contre [DC74, HN71e, Ray70l, Ros74, DDD71, DMW73a, Mar71e]. Contre-exemple [DC74, Ros74, DDD71, DMW73a, Mar71e]. Contre-Exemples [HN71e, Ray70l]. contributed [Yat72a]. Contribution [Mar70b]. Contributions [Dre73, Suc70]. Control [Cur72, Kom72d, Bal70, BJ70, Con72, Dun70b, Kat70a, Kno72, Kom72a, Kom72e, Kom72c, Lio73d, Man72, Mar73b, Neu70, Yev70]. controlled [Wal73c]. controls [Kom72f]. convection [CGS71]. Conventions [Mes72a, vdK73c, KMT74g, dlH72h]. Convergence [AG74, BI70a, Ber71c, Bin73, Hel73, HN71i, Kow73, Kru73, Moz71b, Sch73r, Tho73a, Tho74b, Bor73, Chu71b, CT74, Din71a, Han70, Kli70, Ste74g, SR73, Top70e, Top70n, Wei71c]. Convergent [And74b]. converse [BŁ71a, FNSS73a]. Convex [Kel74c, De 73, Din73, HM74b, Hor73, Sun70a, vL74j]. Convexe [Gru72, LM70, Art72b, Nov74b]. Convexes [HN71n, LM70, Ram74c]. Convexite [Fer72, And74e, Hir72c, Din72, Nov72]. Convexity [Kno73b, Kno73a, Lev73, ST71a]. convoluteurs [Eym71]. Convolution [BD72a, Cho73c, Gan71c, BD72e, BD73, Ber74a, Bor72b, Che74a, CD74, DeV72h, Mey71h]. Convolutions [Ner71a, Sub72]. coordinate [Bab73]. coplanaires [Mar70b]. coplanar [HB70, Mos70]. coproducts [Hof70m]. Cornea [Bre71j]. corps [Bas72, Tat71, Fou71]. correct [Lon71]. Correction [Del71c, PBR71]. corrections [Gru70e]. correctness [JL71, Man71]. corrélation [Car74a]. correspondances [Ray70g]. correspondences [Swa73]. corresponding [Gri73, Maa71c]. Corrigenda [Bur90]. Cosimplicial [BK72b]. cosmoi [Str74c]. Cotangent [Ill71g, Ill71h, Ill71i, Ill72e, Ill72d]. cotorsion [Mat73i]. cotriples [AT70]. Couches [Lio73f, Lio73h]. Countable [Fri73a, Hic73, MS73j, Coh72c]. counter [AS71c]. counterexample [Jos74]. Counterexamples [HN71e]. Counting [Bom74, McA74b]. couplage [Bru74n]. Couples [Ray70m, Art72b]. Courants [Dol71a, Lel72c, Lel73a, Lel74b]. courbe [Art73d, Che74c]. Courbes [DR73, Maz73, Pha71, Ray72a, Ray74, Ser71a, Trá74a]. courbure [NF74]. Cours [Ser73d]. Course [BK74, Eic71e, Her71a, Hol72b, Maa71r, Bre71p]. Cousinverteilungen [LL74c, LL74b]. covariance [OG74j]. covering [BCM74, Fox74, Hav74]. coverings [Fay73b, Fay73e, GM71d, GM71f, GM71k, GM71j]. covers [ST71d]. Coxeter [Tit74m]. Craig [Gab72]. Crash [BK74]. Cremona [Hir73]. crible [Del73q]. cristallin [Ber74m]. Cristalline [Ber74i, Ill72b, Ber74j]. Cristaux [Ber74k]. critère [Bru74o, Trá74b, Wal74h, Ray70-30]. Critères [Gro70d]. Criteria [Top70e, BW72, Tre73d, Wil72]. criterion [Ray70-30]. criterium [Trá74b]. critical [Baz74, Coh74h, KP71, Tim71]. Croissance [Sko72]. croissant [Car70b]. Crossing [Guy72, Guy71]. Crossings [GM71g, Mor71d]. Crystalline [MM74g, Ber74i]. Crystals [Mes72c, MM74f, Mes72b, Mes72f, Ber74k]. cube [Cha74d, Cha74c, Ray70w]. cubes [GP72, May72j]. cubic [Cad71, Dav71, Mur74, Zim72j]. cure [Mar73b]. Current [Lip74h, Gui72d]. currents [Geo74]. curvature [Kui71b, Mas74b]. Curvatures [Spe74]. curve [Väi71e]. Curved [MM74h]. Curves [AS74c, AK70b, Cas73b, Ive70d, Lew72a, PS73, Rob73d, Rob73e, Rob73h, Rob73i, Rob73j, Bom74, Ive70a, Oor74, Pop70p]. Cusp [Spr70b, WP74a, Coh74j]. cuspoids [WP74c, Woo74a, Woo74b, Woo74c, Woo74e, Woo74d]. cusps [Hir72d]. cut [Kel72b, Tho74a]. cut-elimination [Kel72b]. cuts [Pop70e]. cycle [Ber74h]. Cycles [Rug70, Bar74b, Del73j, Gri71a, HL74a, Mar73b]. Cyclic [Bau71b, Fay73b, Ste74j, vL71c, vL71e, vL73c, Chu74a, Kin72, vL73e]. Cyclotomic [Yam74c]. cyclotomy [SW74a]. Cylinders [Sch73s]. cylindrical [LS72]. Cylindrique [Bad74a, Bad74b, Bad74c]. Cylindriques [Bad70g, Bad74e, BC74a, Bad70c, Bad70f, BBB70].

D. [Del73n, Dol72, DDDM70, God71b]. Dafermos [Hed72]. d’algèbres [Gat73a, God74a, Lem74e, Ray70p]. Dalhousie [Fil73, Law72a]. d’Amitsur [KO74f]. Damping [Kom72d]. d’analyse [Lit74c]. d’analyticité [Hir71b]. d’anneau [Ray70y]. d’anneaux [Ray70z]. dans [Aku73, And74e, Céa70, Cha72, Cho73a, CW71b, DC74, DM74, Fer70g, Fer70h, God72, Gro70e, Gru73b, Gru72, HN71b, HN71f, Ill71c, Kah73, KR73, Laz73, Lem70, Lio73d, Lod73, Mar70a, Mar70c, Mey70c, Nov70, Nov72, Rei74a, Rev70, Roo74, Rou74, Sch71a, Sch72b, Ser73j, Sko73, Spi74b, Ste72, Ver72a, Ver73, Zyg71h, Zyg71g]. d’applications [Poe74j, Poe74k]. d’approximation [DC74]. d’après [Ami72, Azr72, Bas72, Bri73, Car71a, Col70, Del73n, Del70g, Dol72, DDDM70, Dol70b, Dou73, Gru74a, GC70, Hir73, Lat72, Lel72d, Lit74c, Mal73, Mey70a, Mey71b, Mey73c, Mey73e, Mor71a, Ros72, Tat71]. d’arrêt [Del73p, Mey72c]. darstellbare [Sch71j]. Darstellung [BtD70d, Lep72, Sch71f]. Darstellungen [Hey70b, Hey70h, Hey70m]. d’Auteurs [Tem71b]. Davis [Azr72, Mey72b]. d’Azéma [Mey74e, Mey73g]. d’Azumaya [KO74d, KO74e]. d’E. [Mar71e]. December [AHiN74, LLNS70]. Décembre [Nor74a]. Décentralisation [LT70]. Decidability [MS70c]. Decidable [MS70d, MS73j, Sca71j]. Decision [Grz70, Beh73, Gir73b, MS70b]. decomposability [Jam70a]. Decomposition [Bru74f, Cai71, cH73, KMT74c, Sil74a, AK70a, Chu74b, Kaz72b, Mat73m, Rao72b, Ray71d, ST71i, Sta74c, Ste70f, Sto72j, Sze74, Tod71, Kos71, Tem71a]. decompositions [Gre73a, Gui72a, Har74g, Ste74d, Maz72, Wil74b]. decreasing [Gan71f]. Deductive [Lam72]. Defect [Bec74a]. deficiency [Dev74, KP71]. definability [Jec71o, oT72]. define [Goo73b]. defined [Beh74b, Col74d, Dol70c, Tak70h]. definies [Nor74b]. Definite [Gui72i, PS72d, Gui72b, Gui72d, Gui72g, PS72b, PS72c]. Definition [BD72b, Bru74j, Hof70g, Ill71g, Mou72c, Mou72b, Sch71l, Top70f, Sto70d]. Definitional [Paw70]. Definitions [Mes72d, Sto72b, Yam74m, AR72, ABK71e, Luc73k, Mey72e, Rei71c, Sch73m, Ste70d, Top70k, Zuc73c]. Deformation [Kra74, Pop70m, Mes72e, Rim72, DM74, Gra74c]. Deformations [Ill71j, Pha71, Ill71h, Ill72d, Ill72f, Ill71h, Ill72e, Ill72d, Mor71d]. d’égales [Art73a]. Degenerate [Fay73c, Gou74a, vdK73d, Fri74, Gou72a]. degeneration [Nam74]. Degree [Sha71f, Bel70, Dan74, GW74a, HS70, Maa71i, Maa71j, Sat73j]. degrees [Lac72, Yat72b]. delay [Gro70b, Yor71]. Delivered [BGJ72, Bre71p]. Dellacherie [Del70g, DMW73a]. Demazure [Hir73]. Demonstration [LLNS70, Ass71, Mey71a, Kow70, Paw70, Art73g, Del71g, Del73s, Fer74b, Mey71h]. démonstrations [Arn70, Bor72a]. Denombrable [HN71b, Jen72h, Lel72c]. dense [Bur74c]. d’Ensembles [Che74g, Ray71c]. Densité [Mok70a]. Density [Wal72a, GU71e]. d’entiers [Kat73h]. d’entrelacement [Ver71]. Dependence [AS71b, AS71d]. dependent [BW74, Set72]. Déploiements [Dem70g]. Depth [AK70c]. d’équations [Kré74a, Tem71a]. d’équisingularité [Trá74b]. Derivable [Rah74]. derivables [Cho73a]. derivation [Tul70b, Hei74a, Mey74h]. derivations [Ste70p]. Derivative [Jen72f, Gor71]. derivatives [AdK74a, Bic73d, O’M74]. Derived [Bei71, CR72a, Keu73]. Derivées [Ill71c, Lem70, Zyg71e, Kré74a, Tem71a]. Derivés [Bru74e, Jen72f, Kar70b]. descendingly [Pri73b]. Descent [GM71d]. Descente [KO74e, Ray70-27, SD72, KO74b]. Description [BGR73d, Bea70c, Sil70c]. design [Kor74, Lor74, Mor74]. designs [Lor74, Tay74a, Wel71b, vL74k, vL74p]. d’espace [BB73]. d’Espaces [Lem74b, Lel71a]. dessus [MM74d]. Detailed [dlH72d]. détaillée [CW71c]. d’Eté [BH74a, BCM73]. Detection [OG74l]. determinant [Bro73b, Ive70c]. Determinants [Ive70d]. Determination [Gai74, Lem74b, Hen71]. determined [BŁ71b, Was74a]. d’Euler [FS70b]. deux [Ben73, Mok70a, Mok72, Poé74a]. development [Hul74]. developments [Gou74b]. Développement [Pis71]. Développements [Den73b]. d’évolution [Lio73c, Lio73e]. d’excédances [FS70f]. d’Existence [Dem70i, Del70g, God72, Kat73e]. d’exponentielles [Ami72]. d’extension [Poe74s]. d’extremalité [Bru74o]. d’holomorphie [Nac71]. d’homotopie [God72, Sie74]. diagonal [Bur74b, Mac72]. Diagonalisables [Gro70f]. diagonalizable [Ste74d]. Diagonals [Shu71a]. Diagrammes [Ill72c]. diagrams [Ber74i, LR72, Sco71, Ste70h]. Dichotomies [Sel73]. dichotomy [Fin74e]. dicht [End74]. dicht-vollständigen [End74]. Dichte [GU71e]. Dictionary [Del70c]. Dictionnaire [Del70c]. Dieudonné [Dem72f]. diffeomorphism [cHyH70, Tur74]. Diffeomorphismes [Mou72d, Hec74b, God71b, Gra74a]. diffeomorphisms [Bru71b, Lan72, Omo74e, Omo74h]. Difference [Bau71b, Bau71c, Stu74, Sug71, Tho73c, Bau71e, Coh74i, Fre73, Gol71c, Gor71, Gou72b, Har71a, Kim73, Mor74, Osb71, Per74, Sac72, Spi72, Ste74j, SW74a, SW74b, Tho71d, Tho74b, Wid71b, AH70b, Tau74, vL74l]. differences [Bea70h, Kro74]. Different [Gan71d]. Differentiability [Yam74n, Yam74o, Sei72h, Sei72i, Sei72j, Sei72a]. Differentiable [HLM71, Luc71c, MY72, Cer71, Dzh74, Kel74b, KK72, Lev71d, Poé71a, Yam74s, Gla71b]. Différentiables [CGN71, HN71a, Poe74t, Gla71a].

Differential [AT72, Bet74b, BGJ72, Chi71a, CG74, ES72c, Fin74a, GH73, HS73b, Her74a, HS71, Kel74c, Kom73c, Kre73d, LRV70e, SM74b, Tay74h, Ura71c, Wat74, Yam74p, Yor70, Ara71, Ars72, Ban71, BJ70, BS74a, Bea74, Bec73b, BP74, Ben70, Ben71l, Bor71b, Bor74, Bra72a, BdS74, Bun72, But74a, Car71e, Car74c, Caz74b, Ces74a, Cha74a, Coh73b, Coh74d, Coh74k, Coh74j, Col73a, Col74b, Col74c, CG71b, CG71a, Coo71b, Dav71, Dun70a, Dwy74, ESS74a, Esp74, Ess74b, EG74, Fen74, Fre71a, Fre71c, Fri72, Gil74d, Gil74f, Gil74e, Gil74g, Gil74l, Gob74, Gor71, Gou72a, Gou72b, Gro70b, Hal71, Har71b, Hen73, Hön74b, Hop71, How74, Hul74, Iwa71b, JLP73, Kan73b, Kap70, Kat70b, Kaw73a, Kha71, Kom73a, Kom73b, Kom73e, Kom73f, Kom72f, Kre74b, Kre73e, Kre73i, Kro74, Kur70b, Lam74]. differential [Las70, Maa71b, Mar71a, Nak73, NS71, O’M74, Oku71, Osb71, Par72c, Pay73, Ple72, Poh71, Por71a, PBR71, Rox71, Sel73, Set72, Spe71, Spe74, Ste74j, Str71, Sty74, Tay74e, Tay74g, Tho73c, Trè74, TP71g, TP71h, TP71i, Ura71b, Wal72e, Wal74m, Was71, Wid71b, Wil74a, Yor71, dJ72, vdH73, Del70d, Poé74h]. Differential- [AT72, Her74a]. differential-delay [Gro70b]. Differentialgeometrie [HRS^{+}73a]. Differentialgleichungen [AT74, Gen74, Haa74, Mic74, NW74, Gek72, Rau74, Sch72c, Sei72b, Tör73]. Differentialoperatoren [Sac74a]. Differentials [Sta70d, Itô72]. Differentiation [Sei72c, Sio73a, Por71a]. différentiel [Ber74g, Gér74, Leb73a]. Différentielle [Poé74h, Vid74, Gri72]. Différentielles [Del70d, Lem74f, Pri74, Lem74c]. différentiels [Blo72, Gru73b, Kan72, Kan74, Kum74]. Differenzapproximationen [AH70b]. Differenzenoperatoren [Sac72]. Differenzenverfahren [AH70c, Has72, Has74, Kol72, AH70d, Gor73, Tau74, vF73, vFD74]. Differenzierbarkeit [Sei72a, Sei72h, Sei72i, Sei72j]. difficulties [Mil74p]. Diffusion [Pri74, SV71a, SV71b, KR73, MM74d]. Diffusions [Del70g, DDDM70, Dol70b, IW73]. Diffusionsgleichungen [vF73, vFD74]. digraphs [GH71, Hem72]. d’image [Maz74]. Dimension [Art73c, Art73i, Coh72f, Mar71d, Poé74g, Pop70i, AK70c, Bab73, BMRH73, Coe74, Col73b, CD71, Din72, Dob70b, Dob70e, Dob70c, Dob70d, Dom73, Fre74a, Gil74c, GM71b, Gru70b, Hir71a, Hir71b, Jen72i, Jen72d, Kau72, Mou71, Nov73, Nov74b, Ohm73, Ray70j, Rug70, Sal73, Sch73n, Sch73x, Sch73-28, Str74a, Tac73c, Vas73]. Dimensional [CM74a, HS74b, Her71a, HMS74e, Mat73a, MM74g, Omo74f, Väi71c, Bre72a, CM74e, Col74e, Gil74h, Gro70a, Gro70b, Har74i, Her71d, HMS74b, HMS74c, KMT74c, KMT74i, KO73b, Mar74a, Nac74b, Nak73, Ram74b, Sch74a, Sch73j, Sch73l, Tay74a, Tim71, Wae74a, dlH72e]. Dimensionen [Gen74]. Dimensions [Dob70a, Gan71h, Her74b, Jen72a, Kor74, Mat71b]. Dinatural [DS70b]. d’indépendance [Wal74f]. d’indices [Jen72h]. d’Information [Hun74, Lan74]. d’instabilité [Rou74]. diophantiennes [Azr72]. Diophantine [Ber71d, Zim72e]. Dirac [Jör72]. Direct [Gol71c, Mac72, BS74b, BK72k, Hil74b, Siu71a, Ste74h, Tak70o]. directe [Maz74, Ray70-31]. Directed [Gri71b, CL74b]. directes [Dou73, Kuh74]. directs [Eym72f]. Dirichlet [Maa71r, Aut72, Bli71a, Bre71d, CR73b, For74, Fre71a, GH74a, Hel73, Maa71c, Meg71h, Sil74a, Sim72a, Wei71d, Wei71e, Wei73]. Dirichletproblems [Tör73]. disc [Col74d]. Discoherently [Sha71m]. Disconjugacy [Cop71a, Har71c]. Disconnected [Ben71c, Ben71j, Hof70j]. discontinuities [Col74c]. discontinuous [Tau74, ZVC70b]. Discret [Aum70]. Discrete [Hof70h, Mey72f, SR73, Bor72a, AG74, GK72, HC70b, HMS74c, HMS74d, Lip74a, Par71b, Tho73a, TP71d, Poé74b]. Discretization [Col74a, Gre71b, Mei72]. discretizations [Ste71h]. Discrets [Poé74b, Bor71a, Ser72]. discriminant [Gir73b]. Discussion [Smo71b]. disjoint [BCM74]. Disjointness [Wu71]. disk [Beh74b, BP72, Hel72]. diskontinuierliche [ZVC70b, ZVC70d]. diskontinuierlichen [ZVC70g]. diskret [Opf74]. diskret-harmonischer [Opf74]. Diskretisierung [Ess73, GS74a, Mei72]. Diskretisierungen [Tör73]. d’isogénie [Tat71]. d’isomorphisms [Con74]. d’isotopie [Mou72j]. dispersions [Bor74]. disque [Cho73a, Ste72]. dissections [BP72]. dissipatifs [HR74b]. distance [Csö73]. Distances [Har74j]. distillation [Ran73a]. distinct [Chi71c, vL74g]. distinguished [Mac72]. Distribués [Lio73b]. distributed [Ben72b]. Distribution [Gal72a, Mon71d, Sto70k, Csö73, Sto70j]. Distributions [HN71f, Ner71b, Dol70c, Gir73b, Poe74n, Vau73]. distributive [Kel74f]. distributivity [Lap72a, Lap72b]. Divergence [OG74c, Mil73a, OG74f, Sac72]. Divergenzform [Sac72]. divers [Eym72a, Sch70e]. Diverses [Hir71b]. Divided [Mes72f, Ber74n, Kro74]. divisées [Ber74n]. diviseurs [Ray70-29, Szp73]. divisibility [GM71a, Mot73]. Divisible [Dem72c, Gui72i, Mat73i, Dem72a, Gui72b, Hey72, Mat73b, Mat73o, Mat73n]. Division [Rob73f, Rob73g, HS73a, Poe74n, Poe74o]. Divisor [GM71g, GM71h, GM71i, Lip74a]. divisorielles [Ray70g]. divisors [Har70c, Kuj71a, Sur72, Wal72a]. do [Mar73b]. Doctrinal [Kel74g]. doctrines [Kel74h]. Dolbeault [Lit74b]. Dold [Geo74]. Doléans [Dol70b]. domain [Aur74, Dzh74, Gil73, Kub73, LKR73, Maa71e, Wel71a, Wil72]. Domaine [Aku73]. domaines [Con70]. Domains [Gil70, Kan71d, Mur72h, AS74a, AS74e, Gai74, Har74i, Hei73b, Iwa71b, Jat70d, Kan71a, Kan71b, Kan71g, Kan71i, Mat70, Mur72a, Mur72b, Mur72d, Mur72f, Mur72g, Mur72i, Rat73, Ric74, Sch74d]. Dominant [Tac73c]. donné [Dez74]. donnée [Sko72]. Données [Dem70b]. Dont [Bad74e]. Doob [Chu71b, Rao72b, Smo71a]. d’opérateur [Ray70r]. d’opérateurs [Ber73, CW71b]. d’optimalité [Con70]. d’optimisation [Bru74p]. d’Ordre [Fou71, Los74, Gru73b]. d’Ornstein [Con74]. Double [WP74a, Fay73e, Pal71, Tac73d]. doubles [Cai71]. doubly [Hay74]. d’ouverts [Hir71b]. d’ouverture [Lau70]. dozen [Hec74a]. droite [GC74]. Droites [LB70a]. DSN [LR71]. Dual [Col74d, AdK74a, Mey73d, Puk72, Sha71a]. Duales [LM70]. Dualität [Hey70a, Hey70c]. Dualitätssatz [Hey70d, Hey70e, Hey70f]. Dualitätssätze [Hey70g]. Dualité [Ber74l, Gru72, HN71g, Las71, Riv72c, DC70, Del73g, Mok72, Sch71a, Ser73j, Ver73]. Dualities [Hof70i, Hof70j]. Duality [AK70d, AK70g, Hof70l, HMS74e, ST71c, Tak72, Ake74, Buc72, Coh70d, Hey70a, Hey73, Hof70k, Hof70g, Hof70h, HMS74a, HMS74b, HN71g, Las71, Tac73f, Hey70c]. Dubins [Mey70a, Mey71d]. Dudley [Fer74b]. d’un [Arn70, Art70, Art73b, Ben73, Ber71n, Ber71m, Ber74m, Blo70, Car70b, Céa70, Del70f, Gér74, Glo70, Gro71a, Gro72e, Jen72i, Jen72d, Kum74, MM74d, Mey71a, Mou72a, PMN70, Ray70d, Ray70k, Ray70a, Rev70]. Dundee [ES72c, Mor71b, Mor71c, SM74b, Wat74]. Dundee/Scotland [ES72c, Mor71b, Mor71c]. d’une [Art70, Art73d, Aus70b, BGM71b, BGM71c, Bru74e, Bru74g, Bru74h, Car71a, Che74c, Fau70, Fer70f, Gra74a, Gri72, HHP71, LM70, Lel72d, Ler71, Mey71h, Rob71a, Tho74a, Zyg71h]. Dunford [Cas70]. d’Uniforme [Che74g]. durch [BGR73d, NW74, Sch71d, Sch71c, Sch71e, Sch71b, Sch71f, Sch71l, Sch71j]. durée [Mar70b]. Dwork [Kat73f]. dynamic [IP70, LRV70c, OTF70, Pro70]. Dynamical [Chi71a, Cur72, Par71a, Sai71a, Yor70, Aus70a, Aus73a, Baj73, Bha70, Got73, Inf71, Kub73, McC71, Par71b, Sma71c, Ura71a]. Dynamics [Bec73a, HLM71, Bow73, Chu71a, Háj70]. Dynamique [Aum70]. dynamiques [Con74]. Dynkin [Ste70h]. d’Yves [Sch71w].

E. [Gol74c, Lel72d, Ver71]. e.a.d.s [Poe74k, Poé74f]. ebene [ZVC70d, ZVC70b]. ebenen [ZVC70g]. ebener [ZVC70a]. échantillon [Car70b]. Échantillons [Art72b]. Echte [GU71f]. Éclate [Ber71m]. Éclatement [LT70]. Éclatements [LB70f]. École [BH74a, BCM73]. ecological [FPC74]. Edge [Mor73]. edges [Chv74a, FGQ71]. Edinburgh [Kno73b]. Edinburgh/Scotland [Kno73b]. edition [Bre71p]. Effective [Con73, JLP73]. effects [Smi74]. Efficiency [WW74b, SW74b]. Eichler [Ber74e, Gar74a]. eigenfunction [Ben74b]. Eigenschaften [LL74a]. eigensystem [DD74d]. eigenvalue [Gol71b, Osb71, Sat73b]. eigenvalues [Gre73b, Ram74a, Sch73a]. Eilenberg [And71b, Coh70a, Lin71, Smi70b, Smi70c]. Einbettungseigenschaften [Bae74]. eindeutigen [Wil70d]. Eindeutigkeitssätze [Pit72b]. Eindhoven [vL74a]. einem [FK72c]. einfach [Pop70i]. Einführung [HRS^{+}73e]. Einhüllenden [BGR73d]. Einige [Pop70j, Ros71, Sch71k]. einiger [Her74a]. Einiges [Pop70k]. Einleitung [BGR73a, Sch71r, Wil70c]. einschliessende [GS74a]. Einschliessungsaussagen [Sch72c]. Einwände [Sch71s]. Eisenstein [HC72, Maa71k, Maa71n]. ELAM [BGJ72]. Elastic [Kom72d]. element [Cia74, Kin74, MM74h, SW74b, Wac74a, Wid71a, Zie74, Zlá74, Mit74b]. élémentaire [Zis70, Sch72a]. Elementare [Hey70i]. Elementary [Bic73a, DK74, Str74c, WP74b, Adl74, Bic73f, Eng71a, Was74c, WP74c, WP74d, WP74e, Zag72b, Zis70]. Elemente [Lor70n]. Elements [Sta73b, Bor71b, KS74b, MM74h, PT73e, Phi73, Ste74e, VCLN74, Bru74q, Lel73a, Gro70h, GS70]. Elimination [Luc73d, Ray70n, Kel72b]. Elliptic [Ati74a, Gil74b, Gil74c, Gil74d, Gil74f, Gil74e, Rob73d, Rob73e, Rob73h, Rob73i, Tay74c, Tay74b, AJ73, Ati74d, Ati74b, BD74, Col73c, Com71, Edm74, ESS74a, Fre73, GH74a, Gil74g, Gil74h, Gil74i, GW74b, Gil74j, GL74, Gol71c, Gou74a, Hes74a, Hes74b, LRV70b, Lio73g, MS73a, Mil73a, Omo74k, Sac72, Sat73g, Sim72c, Sim72d, Stu72, Wac71a, WP74d, You74, dJ72, Rob73j]. Elliptiques [DR73, Lio73f, Lio73g, Lio73h, Maz73, Blo72, Gou71, Sch70c, Ser71a]. Elliptiques-Elliptiques [Lio73f, Lio73h]. elliptische [Sac72]. elliptischen [Mit74b]. Elliptischer [Tör73, Ade72, Gen74]. embeddable [Tit74j]. embedded [Bla74]. Embedding [Lac72, Lam71b, Bar71c, Chu73, Day74a, GM74, GP72, HN74, Lin74, Zas73]. Embeddings [KS74a, KKMSD73c, KKMSD73a, Nar74a, Tho70, Tit74i]. Emphasis [Jec71l]. empirical [Csö73]. endliche [NW74, Sch71j]. Endlichen [Bae74]. endomorphism [Fai72]. endomorphismes [Ver74]. Ends [Coh72d]. energy [Baz74]. Engelken [HK71g]. Enlarged [Bre71p]. Enlargements [Gon74]. enough [Chi71c]. Enriched [Dub70d]. ensemble [Col70, Del74b, Jen72i, Jen72d, Jen72h]. Ensembles [Che74f, Del69, Del72e, Del72f, Fer70f, Mai71, MM74b, MM74a, MM74c, Mey74b, Mey74c, Mey74d, Car71a, Del71g, Del73r, Lel70a, Lel72c, MM70, Mey70c, Mey70h, Sko74, Del71c]. entière [Mar71c]. entières [Gru73b, Lel71a, Sko72]. entire [Bol74, SC73, Ves70]. Entre [HN71a, Ray70b, Mar70b]. Entropie [Che74f]. entropies [BN73]. Entropy [Bab73, OG74d, Smo71c, Smo71d, Smo71g, Smo71e, Goo71, Goo73b]. Entwicklung [LZ73]. Entwicklungen [Pit72a]. enumerable [Lac72, Lac73]. enumerating [Lit74a]. Enumeration [vL71i, Har74a, vL73i]. envelope [Aur74, BGR73d]. envelope-resolvable [BGR73d]. enveloping [BGR73f]. Épaisseur [Che74f]. Épaisseurs [Del72g]. Epimorphismen [GU71f]. Epinglés [Dem70h]. equal [OG74j]. equality [Kow70]. Equation [Pat74, AG74, Bor74, Bra72b, But74a, Fen74, Har72b, Hön74b, Kaz72c, Ker73, Kim73, LRV70c, LRV70d, Lev71a, Mey73a, Neu70, Ran73a, Sán71, SC73, Tho74b, TP71a, TP71e, TP71j, Tur71, Wen74a, Cho73c, Glo70, Gri72, Ree74, LR71]. Equations [BD72a, Bet74a, Bet74b, Chi71a, CG74, Cop71b, DD74a, ES72c, Fin74a, Gil74b, HS73b, HS71, Kom73c, LRV70e, SM74b, TP71f, Ura71c, Wat74, Yor70, Ars72, Auc73, Auc74, Ban71, BJ70, BS74a, Bea74, Bec73b, Ben70, BD72e, BD73, Ber71d, Bor71b, Bra72a, Bro70, Bun72, Car71e, Car74c, Caz74b, Ces74a, Cha74a, Chu72, Coh73b, Coh74d, Coh74k, Coh74j, Coh74i, Col73a, Col71, Col74b, Col74c, CG71b, CG71a, Col73c, Col74f, Com71, Coo71b, Cop71e, Dav71, Del70d, Dun70a, Dun70b, Dwy74, Dye73, Edm74, ESS74a, Esp74, Fin74f, Fin74i, Fin74j, Fin74l, Fre71a, Fre73, Fri72, Fri74, Gan71c, GH74a, Gil74c, Gil74d, Gil74f, Gil74e, Gil74g, Gil74i, Gil74j, Gil74l, Gob74, Gol73, Gol71c, Gor71, Gou72b, GH73, Gro70b, Gup74, Hal71, HO73, Har71a, Hed72, Hen73]. equations [Her72a, Hop71, How74, Hsi73, Hul74, Iwa71b, Jos73, JLP73, Kan73b, Kap70, Kat70b, Kaw73a, Kha71, Kik71, Kom73a, Kom73b, Kom73e, Kom73f, Kom72f, Kre74b, Kre73d, Kre73e, Kre73g, Kre73i, Kro74, Kur70b, Kur71, Lam74, Las70, LRV70b, Mil73a, Nak73, NS71, Nas74, New74a, Nob71, Noh71, O’M74, Oku71, Osb71, Par72c, Pay73, Pim73, Sel73, Sim72c, Sim72d, Ste74j, Sty74, Tho73c, Trè74, TP71d, Wal72e, Wal74m, Was71, Wid71b, Wil74a, Yor71, Zim72e, Zlá74, vdH73, Lie74, Zyg71e, Del70d, Del73l, Lem70, Lio73c, Mar70c, Pri74]. Équicontinues [HN71c]. Equidistribution [Sto70e]. equilibria [Shu71a, Sma71a]. equilibrium [Bro70, Leb73b]. equisingularity [Trá74b]. Équisingulières [Pha71]. Equivalence [Lev71c, Spi72, AR73, BDF73, Dan74]. equivalences [Mat70, Zab74]. equivalent [Bar74a, Rei71q, Sca71b, Sch73w]. équivalentes [BI70d]. Equivariant [KKMSD73a, LW72, Orl72b, Orl72c, OW72, Zag72a, Kar70c, Pet73]. équivariante [Kar70c]. Équivariantes [Ill72f]. Ergebnisse [NW72, Ros71]. Ergodic [Aus73b, Leb73b, Sch73t, Smo71g, Smo71f, Suc70, Bru70, HK70, Kak73, KN73, Kow73, Sch73u, Ter70, Wei73]. Ergodicity [Kub73, Sch73v, Jac70]. ergodigity [Fur73]. ergodique [Con74]. Erhaltung [Pop70p]. Errata [Fri73b, Ill71k, Ill72g, Ill71k, Ill72g]. Erratum [BK72c, BK72d, BK72e, Ser73g, Ser73h, Ser73i, Ser73j, Zuc73b]. error [Coo71b, DD74b, Spi74a, Ste74g, Wid71a]. errors [Smi74]. Erste [Sch71l]. erster [Haa74]. Erweiterung [Lor70a]. Erweiterungen [GU71h, GU71d]. erzeugbare [GU71k, GU71n]. Erzeugende [Pop70e]. Erzeugendensystem [Pop70i]. erzeugte [Sei72g]. erzeugter [Sei72d, Sei72e]. Erzeugungs [GU71a]. Erzeugungs- [GU71a]. Espace [Bar74b, Céa70, Aze70c, Aze70a, BI70c, Blo70, Blo72, Buc73b, Exb70, Gra74a, Har74j, HR74b, MM74d, Mey70i, Mou72a]. Espaces [Aze70d, BI70b, BC74a, Car71c, CW71a, Del71d, Dos71, Eym72b, Eym72c, Gru72, HN71a, HN71b, HN71f, HN71h, HN71l, Jac72b, LB70g, Lem74a, Meg71b, Mor70, And74f, Cho73d, DC70, DC74, DM74, Dol72, Geb74, Gru73b, Hir71a, Hir72c, HN73, Kan72, Laz73, Meg71c, Nov70, Nov72, Ram74c, Ray70h, Ray70i, Ray70s, Sad74]. Espagne [Vid74]. espérances [dSL70]. Esquisse [Gro71a]. essential [Eas74, JW73a, JW73b, JW73e, JW73k]. essentially [JW73k]. Est [Bad74e, Mey73d, Ray70j]. Estension [ST71e]. estimate [BS73a]. Estimates [Sch73a, Coh74u, DD74b, DD74e, Ker73, Sim72c, Spi74a, Tho73a]. Estimating [Cap72]. estimation [CW71b, FL70, LKR73]. estimator [WW74c, WW74d]. Estimators [WW74e, WW74b, WW74f]. Etale [Edw74, Ive73a, Knu71d, Lic73, AK70e, AGV72, Art72a, Art73j]. étalée [Mou72j]. etalées [Mou72c, Mou72b, Mou72f]. Etaler [Pop70m]. étales [GR71, Gro72e, Ray70p, Ray70t, Mou72d, Ray70-30]. etc [Mey71f, Mor71a, Sta70l]. étude [Ben73, Car71b, Gro72d, Mar70b, Sko74, CW71a, CW71c, Gab70b, Gab70c, Kat73a, Mar70a, Mar70c, Mey70c, Ray70o, Zyg71b]. Euclidean [Hel72, Mac73]. Euler [And72, Ohm73, tD72a]. Euler-klassen [tD72a]. Eulerian [Car72a, FS70g, Str74b]. Eulériens [FS70g, God71c, FS70a, FS70d]. evaluated [Fra74]. evaluation [CM72]. évanescents [Del73j]. evasion [Sak71]. even [Kre73d, Wal74k]. events [OG74k]. Everitt [Dev74]. Every [CL74b, Ste70k]. everything [Bec71d]. Evidence [OG74l, OG74i, OG74f, OG74h, OG74k, OG74p]. Exact [Bar71b, BGvO71, Coh70b, DI71f, Ger73c, Mil74d, Mil74k]. exacte [Jou73b]. Exactness [ABK71c]. examination [DS70a]. Example [HK70, AS71c, Bur74a, Mil74d, Wer71]. Examples [BS74a, Bea70b, BD72c, Dun70a, JW73c, Kak73, Kaz72a, Knu71a, Mur72c, Rei71j, Sil74b, Smo71h, Wei71f, dlH72e, Abh73, Bre71j, Gar72g, Mes72d, Omo74i, Bre71g]. excellents [Art73a]. Exceptional [Bac73, Bli71b, Kau72]. exceptionnels [Szp73]. excessives [Mey71c]. excision [Ati74e]. excitation [Wed72]. excitations [LS72]. exemple [DC74, Del70i, DDD71, DMW73a, Mar71e, Ros74]. Exemples [Che74g, HN71e, Ray70p, Ray70q, Riv72d, CW71b, Del70a, Ray70l, Ray70t, Ray70z]. Existence [Bau71d, Baz74, HC70b, HC70c, Ill71b, Kni74, Mit74a, Mou72e, Mou72f, Sch73w, Sto72c, tD72b, Bau71h, Ber70b, Bru70, Cov71, Kaw73a, OS70, Sim72e, Sim72d, Wal74m, vdK73e, Mit74a]. Existenz [Ade72]. Existenzsätze [Pit72c]. exit [Chu74b]. exotiques [Mou74]. expanding [Sac74b]. expansion [Bab73, Ben71l, Fai74]. Expansions [GS74b, Tre73e, Adl73, Ben74b, Chi73, Fra74, Tre73a, Tre73d]. expansive [Law73]. expectations [Sch73q, Smo71a, Tur73b]. Expected [OG74e, OG74f, OG74g, OG74h, OG74i, OG74o, OG74p]. expenditure [Mos70]. Experience [Lon71]. experiments [CA74, Hed72]. Explicit [MM74e]. explicites [Jen72h]. Exponent [Ady74, EG73, Hal73, Qui73c]. Exponential [Fin74e]. exponentials [Mes72f]. exponentiel [Mar71c, Nac71]. exponentielle [FS70c, Wal74f]. exponentieller [Sch71u]. exponentiels [Wal74i]. exponents [Fin74k]. exposants [Cou73]. expose [Mey73b, Mey73h, Del70g, DMW73a, Dol70b, Lel72a]. Exposés [Ano71a, Ano71b, Ano72, Ano74, DE73, Tem71b, Del72a]. exposition [Lam71c]. Expressing [OG74i]. expressions [Iwa71a, OG74o]. extended [AG73a]. Extending [Shi71a]. ’extension [Pis74, Bre71h, GM71e, Gru70f, Gru70g, HHP71, Ler71, MV70d, Mor70, ST71d, Thi70, AS71a, BP74, Bic73g, Coe70, Dho74, Ell73, Gre74, Gru70a, Hea74c, cH73, Kan73b, Lit74a, Omo74k, Sat71, Siu71b, ST71l, ST71m, Wei71g, vdK73d, vdK73f]. Extensional [Luc73e]. extensionality [Luc73d]. Extensions [Dem70c, Fan70, Ill71i, MM74g, Ray70b, Sta73c, vdK73g, vdK73l, BDF73, Deu73a, Dub70d, Eis71b, Fel71a, Hea74b, HN74, KMT74c, KN73, KM74, Lip74d, Luc71h, Mat73e, Mat73p, MM74e, MM74f, Par71b, Ped73, Ran73b, Ric74, RHD70a, Sch73-32, Sta73a, Zim72g, dB70, vdK73p]. external [JW73a]. Extrapolation [Kin74, Sto74, Her72b, Sha71i]. Extremal [Erd74, Caz74b, Din71c, Geh74, MK70, Sha71a, Sha71b, Sha71d, Sim74, Vel72, vL74i]. Extremalprobleme [Vel72]. Extremalproblemen [Ess73]. Extremaux [Bru74c, Lel73a]. Extreme [Bli74, Sun70a].

F. [Fan70]. factor [Pop70e]. Factorisation [Rud73, Day74c].

Factorisations [MM74d]. Factorizable [Gui72d]. Factorization [Zim72f, KM74]. factorizations [And74a]. factorized [Bre72b]. Factors [Won70, Cha74d, Gib72, Lit74a]. facts [KMT74g]. Fading [Col73a]. faibles [Kaz73, Kré74a]. Faisceau [Wag70, Sch74b]. Faisceaux [Art73d, Dem70l, Ray70r, Ray70s, Ray71c, Dou71, Gro72b, Ray70-29, Sch70a, Ver72b, Ver72c]. faithfulness [BR73b]. Faktorkommutatorgruppe [Pop70e]. Fall [Pop70d]. falls [Pop70h]. false [DJ74m]. Familie [Pop70o]. Families [Bau71e, MR71, Sta71a, Bru71b, Chv74a, Fin74o, Mar71a, PS72b, Sto70a]. famille [DD71a]. Family [Ive70d, And73b, BBF74, Ive70a, Väi71e]. fantasia [Lin71]. far [AC74]. Faraut [Den73b]. Faserungen [tDKP70b]. fastperiodische [Hey70i, Hey70j]. February [CG74]. Federal [Nac74a]. feedback [Ben70, Har72b, Iha73, Kat70a, Wil72]. Fehlerabschatzung [Gor73]. Fehlerabschätzungen [AG73b]. Feller [KŌ73a]. fermé [Sun70b, HN71j]. fermées [BB73, Ber71a, Gat73a]. fermés [Kah73, Lel72c, Lel73a]. Fermi [Bra72b]. Fermiane [Wei71e]. feuilletage [HHP71]. Feuilletages [God74b, Ros72, God72, Hae73, Mou74, Rou74]. few [CR72d]. fewer [Col72]. Fiber [AB72, Sto71]. Fibering [Orl72d, CR72b]. Fibonacci [CHS72]. Fibrations [Str74d, BK72t, HR71, Hus70, May72e, Smi70d, Sta74b, tDKP70b]. Fibre [Ber71n, BK72f, Hus73, Jam70a, Mou72c]. Fibres [LB70g, LB70a, Ray70i, GV72a, Gro72d, Dem70d]. Field [Col74b, Nar74a, Zim72g, AS71e, BT73, Chu72, CA74, Col71, Deu73a, Gli70, KMT74h, Kel74a, Luc71h, Yam74g, Yam74f, Yam74i, Yam74h]. Fields [MV70d, Sta73i, AG71c, Ber71i, Bom74, Car73a, Coh72j, Deu73b, Dob70b, Ede70, Gor70, Gra73, Kei74, Łoj71, Lor70h, MS73a, MV70b, Mur72k, Omo74g, Oor74, PŠ70, Sch73z, Seg70, Shi71b, Tho71c, Wal71e, Wei71h, Wei71l, Zim72a, Zim72b, Zim72h, Zim72n]. filter [Ber70a]. filtered [AS74a]. filtering [Ben72b, Gri73]. Filters [Glo73]. filtrations [Coh74s, Riv72e]. Fin [Art73g]. Final [Maa71d, DS70a]. Finding [Alb74]. fine [Bre71a, Bre71c, Bre71k, Bre71l, Dev73f, Bre71g]. Finely [Fug72c, Fug72d]. Fines [Bru74c, Bru74b]. fini [Jen72j, Ray70d, Tat71]. finis [Gro72d, Ray70x, Spr74, Tit71a]. Finite [BCC^{+}70, Bra74, Cos74, Gru70f, Gru70h, Her71a, Hof70m, Jec71h, MS73d, Ost70a, Qui73a, Swa70, Tit74f, Tit74n, Zim72h, Zlá74, AG73a, AS71a, Ati74i, Bas73f, BN73, Bom74, Cia74, Coh73a, Con73, CD71, Cur70, Dom73, Fai72, Fin72, Har74g, Her71e, Hor70a, cHyH70, Iwa70, Kin74, Kre70b, Lem71, Lin74, Loe74a, Mis74, MM74h, Orl72e, OS70, Per74, Pet70c, Rin72b, Sch71j, Spi72, Spr70b, Sta73h, Sta70g, Sta71a, Ste74j, SW74b, Tac73j, Tho74b, Wac74a, Wid71b, Wid71a, Zie74, dlH72c, Mit74b, vL74m]. finite-difference [Spi72, Ste74j]. finite-element [Wac74a, Mit74b]. Finitely [Bau74a, Top70g, Was74a, BŁ71b, Coh72e, Lem71]. Finiteness [Knu71e, SE70d, Gru73a, Tak70l]. Finitude [Ill71a, Ill71c, Kle71, Trá74a, Art73i, GV72a, Ram74c, Ray72b]. First [CFK71, Fin74f, Ste74d, Sto70f, Suc70, Top70h, Bre71j, Fit72, Gil74f, Gil74e, Mat73g, PS72a, dJ72]. first-order [Fit72, dJ72]. Fisher [Agg74]. Fitting [BC74b, Ste70c]. five [CS74]. five-spoked [CS74]. fixe [Eym72e, GC74]. Fixed [Bel70, Bon72a, Fin74g, Gru70i, Rho74, Fan70, GQ71, Gus73a, cHyH70, Jaw70, Pat70, Pow70, Sch72e, tD72b]. Fixed-point [Bon72a]. Fixpunktsätze [Sei72b]. Fläche [Pop70o, Pop70h, Gru74b]. Flächen [Pop70o, ZVC70g, ZVC70d, ZVC70c]. Flächenkomplexe [ZVC70f]. Flat [AK70f, Spe71, Gun71]. flats [Rob74]. flight [Bal70, IDS70]. flots [dM72]. Flour [BH74a]. flow [Bec73b, Bha70, Fur73, Kow73, Sai71c, Sco71, Sim73]. Flows [Ahm71, Bra71, Aus73b, Bec70, Bec71a, Bec71b, Bec71c, Car71e, Fre71b, Gre71a, Jac70, Orn70, Rén73, Ric71, Sha71k]. Fock [Dwy74]. Fokker [Har72b]. fold [Ive70e]. Folgen [Sch71d, Sch71n, Sch71t, Sch71m, Sch71q]. foliations [Bot72, Rei74b]. Föllmer [Mey72a]. following [Bom73]. Foncteur [Kle71, RK71, Ray70g, Ray70-31]. Foncteurs [Gro72d, Jen72f, Kar70b, Ler71, Jen72e]. Fonction [Aus70b, Bru74e, Bru74g, Bru74h, Far73, LM70, Lio73a, Ber74l, Kat73d, Mar71c, Wal74f]. fonctionnel [Fer74d]. fonctionnelle [Nov72]. Fonctionnelles [CGN71, Del73l, Gir73a, Mar71c, Del70h, DMW73a, Kré74a, Mar70e]. Fonctions [Art71, Bad70g, BC74a, Bru74c, Bru74b, Bru74a, Buc73b, Del73l, Del72f, Exb70, Fer70g, Fer70a, FS70a, Gla71a, Lel70a, Man73b, Meg71d, Nor74b, Nov70, Rei74a, And70, Aze70f, Bre74a, Bre72c, Bre72d, Cho73a, Cho73g, DM74, Fer74d, Gom70, Gru73b, Gru74a, GC74, Hir74a, Hir71a, Kah73, Mey70i, Mey71c, Nor74c, Rud73, Sch70g, Ser73k, Sib74, Sko72, Sko73, Nor74a]. Fonctorialité [Ver72b]. fondamental [Che74c, GR71, Ray72b]. Fondamentales [HN71k, Mou72c, Mou72b]. fondamentaux [Del73o, Mou72d]. force [AC74]. forced [GS71]. forces [JW73a]. Forcing [Jec71l, DJ74g, Jec71m]. Foreword [FNSS73b]. form [Coh74j, Fay73d, Gan71g, Mil73a, Sac72, Sil74a, Ste70i, Sto70c, Sto70d]. Formal [Gra74h, GM71g, Har70e, Knu71f, Rim72, Dem72b, Dem72e, Gra73, GM71e, GM71h, GM71i, GM71j, Knu71a, Lip74a, Luc73f, Mes72h, BtD70f]. Formal-rational [Har70e]. Formale [BtD70f]. formalisation [Arn70]. formalism [Buc72]. Formalisme [God71a, JG71, Ray70t, Del73j, Gol74b]. Formalization [BW71, GM72]. formally [Ple72, Sca71a, Sca71b]. format [Lov70]. Formelles [Ill72b]. formels [Car71d, God74a, Kar73a]. Formen [Lor70d, Lor70g, Lor70h, Lor70i, Lor70j, Lor70k, Lor70m, Gey74]. Formes [Del71a, Del73e, God71c, LB70d, Mar71d, Rob73k, Ser73k, BB73, Dol71a, Ser73f]. Forms [Eic71e, Eic73, JL70a, Jac72a, KMT74d, KMT74e, Kat73g, Lan73b, Maa71r, Shi73b, SD73, Wei71e, Bai70, Bli71a, Cap73a, Eic70, Eic71a, For74, Fuk73, Gar74a, HC70k, KS71a, Kub71, Lan70, Lor70h, Maa71a, Maa71c, Maa71h, Maa71i, Maa71m, Maa71u, SM74a, Spr70b, Stu72, Zim72m]. formula [Ati74k, Ber72a, Coh74c, Fuc74, cH73, Luc71g, Man72, Sha71g]. formulae [Bet74a, vL74c]. Formulaire [Bor72b]. Formulas [Jec71i, Sca71c, Sil70d]. Formulation [Sko73, Tak70b, VCLN74]. formulations [Mon71a]. formule [Del73g, Del73h, Del73i, FS70c, Kat73d]. Forschungsinstitut [AT72, AT73, AT74]. fortement [Fou74]. fortes [Lan74]. Fortsetzung [Gru74b]. Fortsetzungssätze [LL74c, LL74d]. Foundation [Kom73a, Sch71v]. Foundations [AG71a, Sch73f, Wal73b, Gli70, Wac74a]. Four [But74d, EG73, Hal73]. Fourier [BBB70, Bai70, Ben71i, Ben71l, Cho73c, Fin74h, Gli72, GW72, GK72, Mal73, Moz71b, Ner71b, Ner71c, Rub73, Sha71i, Sil70a, Sil70d, Tre73e, Trè74, Vau73, Zyg71a]. Fournier [Cun74]. Fourth [Kre73e]. fraction [Fra74]. fractions [Coh72j, Zim72d]. Fraenkel [Dev73u, Fel71c, Jec71c]. Fragen [Pop70j]. framed [Rei74b]. France [Ser73d]. Francis [Den73b]. Francisco [BK74]. Franco [Bru71c]. Frankel [Myh73]. Fréchet [Bon72a, DM74, Luc71d, PT73b, PT73i]. Fredholm [FNSS73c, Kar70a, Mou72c, Mou72b, Mou72g, Mou72h, Sch73a, TP71d, TP71j]. Free [Bry74a, Dix73, Gru70j, Orl72e, Str72, WWS72a, Wan72, CGS71, Cas73c, CLL73, Coh73a, Col74b, Gru70i, HR70, JW73b, JW73m, Kaw72b, Kro72, KK72, Mon71r, MY72, Pet70c, SW74a, Wal73a]. Freedman [Mey70a]. Freie [ZVC70e]. freier [Sac74a]. French [Agg74, AK70e, Ami72, And70, And74e, And74g, And74f, AdL72, Arn70, Art70, Art73b, Art73a, Art73c, Art73d, Art73e, Art73h, Art73g, Art73f, Art73i, Art71, Art72b, Ass71, Aum70, Aus70b, Aut72, Azé72, AM74, Aze70c, Aze70a, Aze70b, Aze70e, Aze70f, Azr72, BI70a, BI70b, BI70c, Bad70b, Bad70c, Bad70d, Bad70f, Bad70e, BI70d, BBB70, BB70, Bad74a, Bad74b, Bad74c, BH74a, Bad74e, Bad74d, Bad74f, Bar74b, Bas72, Bea71, Ben73, BB73, Ber74a, Ber74c, BGM71a, BGM71b, BGM71c, Ber71a, BGM71d, Ber74d, Ber73, Ber71n, Ber71k, Ber71m, Ber71l, Ber71o, Ber74h, Ber74j, Ber74i, Ber74g, Ber74k, Ber74l, Ber74m, Ber74o, Ber74n, Ber70d, BS74c, Blo70, Blo72, Boc72, Bom71, BS73b, Bor71a, Bor72b, Bor72c, Bor72a, Bou74a, Bre74a, Bre71t, Bre72c]. French [Bre72d, BCM73, Bre73, Bre74b, Bri73, Bru74c, Bru74b, Bru74a, Bru74f, Bru74d, Bru74e, Bru74g, Bru74h, Bru74j, Bru74l, Bru74k, Bru74m, Bru74n, Bru74o, Bru74p, Buc73b, Cai70, Cai71, Car71a, Car70b, Car71c, Car71b, Car71d, Car72c, CR73b, Car73b, Car74a, Car74b, Cas70, Céa70, CGN71, Cés74b, Cha72, Cha73, Cha74e, Che74a, Che74b, Che74c, Che74d, Che74e, Che74f, Che74g, Cho73a, CD74, Cho73d, Cho73e, Cho73g, Cho73f, Cho73b, Coe70, Coe74, CW71b, CW71c, CW71d, CW71e, CW71g, CW71f, Col70, Col73b, Con70, Con74, Cou73, DC70, DC74, Del70a, Del70b, Del70c, Del70d, Del70e, Del71a, Del71b, Del72a, Del72b, Del73d, Del73l, Del73c, Del73j, Del73g, Del73h, Del73e, Del73i, Del73f, Del73m, DR73, Del73k, Del73n, Del69, Del70f, Del70g]. French [Del70i, Del70j, Del70h, Del71c, Del71g, Del71d, Del71f, Del71e, Del72d, Del72c, Del72g, Del72f, Del72h, Del72i, Del73r, Del73q, Del73s, DMW73a, Del73p, Del73o, Del74a, Dem70a, Dem70b, Dem70c, Dem70d, Dem70g, Dem70e, Dem70f, Dem70h, DG70b, Dem70k, Dem70j, Dem70i, Dem70l, Dem71, Den73b, DM74, Dez74, Die72, Din72, Dix71, Dix74, Dol71a, Dol71b, Dol72, DDDM70, DDM70, DD71a, DD71b, Dol70b, Dos71, Dou71, Dou73, Dup74b, Eke70, Exb70, Eym71, Eym72a, Eym72b, Eym72c, Eym72d, Eym72e, Eym72f, Eym72g, Far73, Fau70, Fer74b, Fer70c, Fer70f, Fer70g, Fer70a, Fer70b, Fer70h, Fer70e, Fer70d, Fer72, Fer74d, FS70c, FS70a, FS70b, FS70d, FS70f, FS70e, FS70g, Fou71, Fou74, Fra70a, FL70, Gab70a, Gab70b, Gab70c]. French [Gab70d, Gab71, Gat73a, Geb74, Gér74, Gir73a, Gla71a, Gla71b, Glo70, God71b, God72, God74a, God74b, God71c, Gol74b, Gom70, Gou71, Gra74a, Gri72, Gro70c, Gro70d, Gro70h, GS70, Gro70f, Gro70e, Gro70g, Gro71a, Gro71b, Gro72b, Gro72c, GV72a, Gro72d, GR72, GV72b, Gro72e, Gru73b, Gru74a, Gru72, GC74, GC70, Hae73, Har74j, HHP71, Hec74b, Hei73a, Hei74a, Her74d, Her74e, HR74b, Hir74a, Hir71a, Hir71b, Hir72c, Hir73, HN71a, HN71b, HN71c, HN71d, HN71e, HN71g, HN71f, HN71h, HN71i, HN71k, HN71l, HN71m, HN71n, HN71j, HN73, Hou72a, Hou72b, Hun74, Ill71e, Ill71f, Ill71g, Ill71h, Ill71i, Ill71a, Ill71j, Ill71k, Ill71b, Ill71c, Ill71d, Ill72d, Ill72c, Ill72b, Ill72f, Ill72g]. French [Ill73, Jac72b, Jen72b, Jen72j, Jen72a, Jen72c, Jen72i, Jen72f, Jen72e, Jen72d, Jen72h, Jen72g, Jou73a, Jou73b, JG71, Kah73, Kan70, Kan71k, Kan72, Kan74, Kar70a, Kar70b, Kar70c, Kar73a, Kar73b, KR73, Kat73a, Kat73d, Kat73b, Kat73e, Kat73f, Kat73c, Kat73h, Kaz73, KF73, Kle71, KO74d, KO74b, KO74a, KO74e, KO74f, KO74g, Kos71, Kré74a, Kri71, Kuh74, Kui73b, Kum74, Lan74, Las71, Las74, LR71, LB70b, LB70d, LB70c, LB70f, LB70g, LB70a, LB70h, LB70i, LB70k, LB70l, LB70m, LB70j, Lat72, Lau70, LM70, dSL70, Laz73, Leb73a, Lel70a, Lel71a, Lel72a, Lel72c, Lel72d, Lel73a, Lel74b, Lem70, Lem74b, Lem74f, Lem74c, Lem74e, Lem74d, Ler71, Lic74, Lie74, LT70, Lio70, Lio71, Lio73b, Lio73a, Lio73c, Lio73e, Lio73f, Lio73g, Lio73h, Lio73i, Lio73d, Lit74b]. French [Lit74c, Lob72, Lod73, Los74, Luc71d, Lut74, MM70, Mai71, Mai72, MM74b, MM74a, MM74c, Mal73, MM72, MM74d, Man73a, Man73b, Mar70a, Mar70b, Mar70c, Mar71b, Mar70e, Mar71c, Mar71e, Mar71d, Mar74b, Maz71, Maz72, Maz74, Maz73, Meg71a, Meg71c, Meg71f, Meg71d, Meg71e, Meg71h, Mey70c, Mey70i, Mey70d, Mey70a, Mey70g, Mey70e, Mey70h, Mey71d, Mey71h, Mey71a, Mey71b, Mey71c, Mey71e, Mey72a, Mey72d, Mey72b, MW72, MR72, Mey72c, Mey73b, Mey73c, Mey73d, Mey73e, Mey73h, Mey73i, Mey73f, Mey73g, MT73, Mey74b, Mey74c, Mey74d, Mey74g, Mey74a, Mey74e, Mey74h, Mey74f, Mit74a, Mok70a, Mok70b, Mok71, Mok72, Mor70, Mor71a, Mou71, Mou72a, Mou72d, Mou72c, Mou72b, Mou72e, Mou72f, Mou72j, Mou72h, Mou72i, Mou74, Nac71, NF74, Nor71, Nor74b, Nor74c, Nor74d, Nov70, Nov72]. French [Nov73, Nov74b, PP74b, Pet74a, Pha71, Pis71, Pis74, Pit72f, Poe71b, Poe72, Poe74i, Poé74h, Poe74j, Poé74a, Poe74n, Poe74k, Poé74b, Poe74m, Poé74e, Poé74f, Poe74t, Poé74c, Poe74o, Poe74s, Poe74p, Poe74q, Poé74d, Poe74r, Poé74g, PMN70, Pri74, Ram74c, Rav71, Ray70g, Ray70d, Ray70e, Ray70f, Ray70l, Ray70h, Ray70m, Ray70j, Ray70i, Ray70k, Ray70a, Ray70p, Ray70o, Ray70n, Ray70q, Ray70r, Ray70-31, Ray70t, Ray70b, Ray70c, Ray70u, Ray70v, Ray70x, Ray70y, Ray70z, Ray70-28, Ray70-27, Ray70-29, Ray70-30, Ray70w, Ray70-32, Ray70-33, Ray71b, Ray71c, Ray71a, RK71, Ray72a, Ray72b, Ray74, Ree74, Rei74a, Rev70, Riv72c, Riv72d, Riv72e, Riv72g, Riv72h, Rob73b, Rob73c, Rob73k, Rob71a, Roo74, Ros72, Ros74, Rou74, Rud73, Rug70, Sad74, SD72, Sal74, San74, Sch70a, Sch70b]. French [Sch70c, Sch70e, Sch70d, Sch70f, Sch70g, Sch71a, Sch74b, Sch72b, Sch72a, Sch71w, Sch72f, Ser71b, Ser71a, Ser72, Ser73f, Ser73e, Ser73b, Ser73j, Ser73k, Sib74, Sie74, Sko72, Sko73, Sko74, Spi74b, Spr74, Ste72, Ste74i, Sun70b, Szp73, Tat71, Tem71a, Tem71b, Tho74a, Tit71a, Trá74b, Trá74a, Trá74c, Vau73, Ver72a, Ver72b, Ver72c, Ver73, Ver74, Ver71, Wag70, Wal74h, Wal74a, Wal74b, Wal74c, Wal74d, Wal74e, Wal74f, Wal74g, Wal74i, Wei70, Zis70, Zyg71a, Zyg71b, Zyg71d, Zyg71c, Zyg71e, Zyg71f, Zyg71h, Zyg71g, dF74a, dM70, dM72]. frequency [Chi71b, Gol71a, Wil72]. Frobenius [Moo71, MP71, Omo74b, Omo74c, SE70e, Tac73k, Ver74]. Front [BK72g, BK72h, BK72i, BtD70g, BGG^{+}72, DJ74d, MS70e, Rob73l, Ser73l, Tit74o, vL71d, vL73d, And74d, Dio74, Tay74f]. front-sets [And74d]. frontière [Bre74a, Cho73a, Con70]. Fuchsian [Bur90, Bur74d]. fuel [Mos70]. fugue [Lin71]. Full [AS71f, Bar71c]. fully [Jat70a]. Function [Coh74b, Deu73b, Dun70a, Wei71h, AG71c, AdK74b, Beh74c, BZ71, Bre71n, Bru74g, Bru74h, Coh74e, Coh74h, Coh74i, Csö73, Deu73e, Fay73f, Gir73b, Hön74b, KŌ73a, MS73a, MS73b, MS73c, MS73g, MS73h, MS73i, Ric74, Sca71b, Sca71k, Smi72, Wac71b, Wal72a, WW74a, WW74f, Zie74]. Function-fields [Wei71h]. Functional [Bli71b, Ces71, Ces74a, Hal71, Luc73e, Luc73h, Luc73g, Sat73d, Ura71c, BJ70, Bli71a, Col73a, Fre71a, Gar74b, GH73, Hen73, Kat70b, Kur70b, Luc71f, Luc73i, Luc73l, SC73, Sim72c, Sim72d, Tho73c, Wae74a, GUW74, Nac74a]. functional-differential [Col73a, GH73]. functionals [Bea70h, CGN71, Dud70, Luc73a, Luc73b, Sim72e]. Functions [AS71b, DeV72c, DK73b, Deu73c, Fay73g, Fug72d, GG72, GJ72c, Gui72i, Her71a, Kel74d, Kui73a, KS73b, MS73e, Ogg73, PS73, Sil70a, AS71j, AS74f, Aze70f, Bab73, Beh73, Beh74b, Bel70, Ber71j, Ber72b, BL74a, Bli74, Bol74, Bre71h, CHS72, CM72, Coh74g, Coh74l, Col72, Col74d, Csö73, DeV72b, Deu73e, Dho74, Din74a, Dwo73, Dzh74, ES72b, Ess74b, Fef72, Fic74, Fin74c, Fin74b, Fra74, Fuc74, Fug72c, Gal72a, GP74, Gui72b, Gui72d, Gui72g, Hag74, Har70e, Hir74b, Jac70, Kan73a, Kel74b, KS71a, Kor70, Lel71b, Lel74c, Lic73, Luc71c, Maa70, Maa71q, Maa71u, Mac73, MS73d, MS73f, MC70, Mon71d, Mon71k, Nov74a, OG74j, PŠ70, Ple72, Pro74, Rub73, Rya72, Sch74a, Sch74e, Shi71b, Sio73b, Tul70a, Ura71b, Ves70, Wae74b, Wen72, Wen74a]. functions [Wid71a, Wil72, vL74d, vdH73]. Functor [Dub70c, Ben74a, BR73b, Day74c, DS73, Lew72b, Mil74i, vdK73j]. functorial [Kel72c]. functors [And71a, And74b, Day70, Day74d, Fef72, GU71e, GU71i, GU71j, Ger73b, Jen72f, Jen72e, Keu73, Mac70, SE70e]. Fundamental [Coh74m, DeV72d, Fri74, GM71g, Kan73b, Net74, Orl72f, BD72d, GM71b, GM71h, GM71i, Ive73f, Knu73d, Maa71e, Ste70q, Yam74m, HN71k]. Fundamentalgruppe [Pop70d, Pop70g, Pop70f]. Fundamentalgruppen [Pop70l]. Funktionalgleichungen [AT73, Pit73]. Funktionen [Opf74, LL74b, Mic74, Pit72a, Sch74a]. Funktionenkörpern [Lor70g]. funktionentheoretische [End74]. funktionentheoretischen [Opf74]. Funktoren [GU71e, GU71i, GU71j]. Further [Bre71i, Gil74e, Gil74g, Knu73b, Luc73i, Mil74e, Mon71e, Orl72g].

G. [Ben73, Che74f]. Gårding [Sim72a, Tay74i]. Gårding’s [Sim72e]. Gaddum [CM71]. Gaifman [Glo73]. Galerkin [Dup74a, Ste74g, Tho73a, Tho74c, Wen74b]. Galois [DI71c, Ser73e, Zim72i]. Galoisienne [Ser73c, Ser73d, Ser73e, Ser73h, Ser73i]. Galton [KF73]. games [BL70, Par72g, Pro70, Rox71]. Gap [Dev73g, Dev73h, ST71e, Dev73i, ST71i, ST71j]. gap- [Dev73i]. Gap-1 [Dev73g, Dev73h]. Gap-Sheaves [ST71e, ST71i, ST71j]. Garben [GU71g]. Garland [Bas72]. gas [Dup74a]. Gaschütz [Bar74c]. gauges [Bic73g]. Gaussian [Che74e, Gir73b, Gui72e, Gui72i, Iha73, KK73, KO73b, Kot73, OG74a, OG74i, OG74j, OG74s, Sha71g]. Gaussienne [Bad74a]. Gaussiennes [BC74a, Bad70c, Che74e]. gaussiens [Hei73a, Hei74a]. gauss’schen [Eng73]. GCH [Jec71f]. Gear [Dan74]. gegen [Sch71s]. Gelfand [Luc71e, Wlo73]. Gel’fond [Wal74b]. gender [Pop70p]. Genera [Rog70a]. generability [PE73]. General [AHiN74, Bor73, Bre71j, Bre71k, Bre71l, Can74, CL74a, Gil74f, Gru70g, Omo74d, PT73a, Par72e, Sha71f, Sil74c, Ste71c, Tre73b, TP71b, Bre71c, CM74f, Coh72j, Csö73, Fre74b, Gan71j, Got73, HC70c, Luc73o, May70, Mes71a, Pay73, Rei71p, Ric74, WW74f, DG70b]. générale [Del70i, Del73o, Mey70h, Mey73g]. Générales [DG70b, FS70f, Gro72c]. Generalisation [Sai71b, Bru74f]. Généralisations [Her74e]. généralisé [Cas70, Del73q]. Generalised [Mur72d, Dev73b, Dev73j, Dev73k, Dev73t, Par72c]. Généralisée [Los74, dF74a, Che74a, Cou73]. généralisés [Geb74]. Généralités [Ber71k, Ber74m, Ber70d, Dem70e, Fer70b, Gab70d, Ill71c, Jac72b]. generalities [Sil70e]. Generalization [Har70f, Luc73j, Sim72a, Caz74a, Dob70c, Dro71, Kan71c, Maa71f, Sim72e]. Generalizations [Siu71a, Tac73k]. Generalized [BdS74, Bru71a, Bur74e, Eis71a, Luc73k, MS73e, Rao70, Smi72, Tak70c, Ark74, BG73, Ess74c, Har72a, Har74c, Hei74b, Hig74, Inf71, Kle72, Nar74a, Nar74b, OS70, OG74i, Rah74, Sta70n, Tak70k, Wal73a, Wen72, Wen74a]. generated [Coh72e, cHyH70, Lem71, Ulm71]. Generating [vL74d]. generation [Fuk73, Low73, Luc73j, Qui73a]. Generations [Sai72a]. Generatoren [GU71n]. Generators [MV70a, Pop70e, Kre70b, Qui73c, Sal73, Smo71d, vdK73o, Tit74b]. génératrices [FS70a]. Généraux [Aum70, Gro71a, DG70a]. Generic [Ive73b, Rob71b]. genericity [Sma71c]. générique [Lob72]. Gentzen [Sca71i]. Genus [Wil74c, AS74b, Mis71a]. Geodätische [FK72c, FK72b]. Geodesic [Gre71a, FK72c]. géodésiques [BB73, Ber71a]. Geometric [AS74c, ABK71b, Bas73b, Bea70c, Bea70d, Bor71b, Kha71, Lit71, May72f, May72g, Por71a, Sul71c, Hop71, KR72, Pet73, Wal71f, FS70g]. Geometrical [McA74a, WP74b, WP74a]. Geometrie [FK72d, Gro72f, LR71, GR71, Art72a, Art73j, Bea71, Maz71, Ray74, Riv72d, Sch71a, Pop70n, Car73b, DK73a, Ray71b, Vid74]. Geometrien [Wil70g, Wil70d]. geometries [Bry74c, DK74, Wil70f, Wil70g, vL74m]. Géométrique [FS70g, Bru74k]. géométriquement [Ray70f]. Géométriques [Bad74c]. Geometrische [Wil70e]. Geometry [Law72a, Maa71g, May72h, AS74a, AS74b, AS74d, AS74e, BV74, Blo73a, Buc72, Har70b, KMT74d, LRV70f, Maa71s, Mat71c, Poh71, Por71a, Ram74b, Wac74a, WP74c, WP74d, WP74e]. George [BH74b]. German [Ade72, AS71g, AH70a, AH70b, AH70d, AH70e, AH70g, AG73b, Bae74, Boh74, BGR73a, BGR73d, BtD70e, BtD70a, BtD70b, BtD70c, BtD70d, BtD70f, BtD70h, BtD70i, BtD70j, BtD70k, BtD70l, BtD70n, BtD70m, Bul72, End74, Eng73, Ess73, FK72b, FK72d, FK72c, FK72a, GU71a, GU71b, GU71c, GU71d, GU71e, GU71f, GU71g, GU71h, GU71i, GU71j, GU71k, GU71l, GU71m, GU71n, GU71o, GU71q, GU71p, Gek72, Gek73, Gen74, Gey74, Gor73, GS74a, GR70, Gra74c, Gru74b, Haa74, Has72, Has74, Her74a, HK71a, HK71d, HK71e, HK71f, HK71b, HK71c, HK71g, HK71h, Hey70a, Hey70b, Hey70c, Hey70d, Hey70e, Hey70f, Hey70g, Hey70h, Hey70j, Hey70k, Hey70m, Hey70l, Hir74c, HRS^{+}73a, HRS^{+}73e, HRS^{+}73b, HRS^{+}73c, HRS^{+}73d, Ive73d, Kol72, Kre74c, LL74a, LL74b, LL74d, LL74e, LL74f, Lep72, LZ73]. German [Lor70a, Lor70d, Lor70f, Lor70c, Lor70g, Lor70h, Lor70i, Lor70j, Lor70k, Lor70l, Lor70m, Lor70e, Lor70b, Lor70n, Meg71g, Mei72, Meu73, Mic74, Mit74b, NW72, NW74, Opf74, Pit72a, Pit72c, Pit72b, Pit72d, Pit72e, Pit72g, Pit73, Pop70b, Pop70a, Pop70j, Pop70k, Pop70i, Pop70d, Pop70e, Pop70h, Pop70g, Pop70f, Pop70m, Pop70n, Pop70c, Pop70o, Pop70p, Pop74a, Rau72, Rau74, Rog70d, Ros71, Sac72, Sac74a, Sch74a, Sch74c, Sch71d, Sch71c, Sch71e, Sch71b, Sch71j, Sch71f, Sch71l, Sch71k, Sch71m, Sch71n, Sch71o, Sch71p, Sch71g, Sch71h, Sch71i, Sch71r, Sch71q, Sch71s, Sch71u, Sch71t, Sch71v, Sch72c, Sei72a, Sei72b, Sei72d, Sei72e, Sei72f, Sei72h, Sei72i, Sei72j, Tau74, Tör73, Vel72, Wac72, Wac74b, Wil70b, Wil70a, Wil70c, Wil70d, Wil70e, Wil70f, Wil70g, Wil70h, Wil70i, ZVC70a, ZVC70b, ZVC70c]. German [ZVC70e, ZVC70g, ZVC70f, tDKP70a, tDKP70b, tDKP70d, tDKP70c, tDKP70e, tD72a, vF73, vFD74]. Germes [Poe74k, Poe74j]. germs [BŁ71b, Mat71c, Was74a]. Gerschgorin [Hof71]. Gerschgorins [Gor73]. Geschlecht [Pop70g]. Geschlechts [Pop70p]. Gesetzes [Sch71f]. Gewinnung [Opf74]. gewisser [Ade72]. gewöhnliche [Tau74]. gewöhnlichen [Sei72b]. Give [Coo74]. Given [Coo74, Eic71e, Her71a, Maa71r, Rob73j, BK74, OTF70, Ste70b]. Gleason [Bea70b, Bea70g]. Gleichung [NW72]. glimpse [BK72j, Nac74b]. Glivenko [Csö73]. Global [DD74b, GJ72a, JL70b, Kur70b, Str71, BT73, Ber70b, Kaw73a, Kru73, Rox71, Vas73, Wil72, HLM71]. globale [Del73g, Ray70-27]. Globales [Ill71b]. Globalisierung [Wac74b]. Globalization [AS71h, ST71f]. globaux [Bas72]. Gödel [Luc73e]. Gödel’s [Luc73l]. Goethals [Di 71]. Goldie [Jat70a, Jat70b]. Goldman [Sto72j]. gons [Bar71a]. Good [Mas74a, dB74]. Gordon [Bre72b]. Gorenstein [Mat73h]. governed [Ben70]. Graded [Eic71b, Eic71c]. gradiant [Dat70]. Gradient [rd74a, BD74, KS70, MC70, Ree73, Wal71e]. Graduations [Riv72e]. Graph [BH74b, Bon72b, CFK71, Alb74, BCM74, Cap72, Chi71b, Chi71c, CL74b, Chv74b, Gal72b, HR74a, Kai74, Sch74f, Wae74b, ALW72]. Graphe [Ber74d, HN71j, Sun70b]. Graphen [ZVC70e]. Graphical [CM71, Har74a]. graphiques [Bru74o]. Graphs [Hed71, Tay74a, AM71, ADT72, Bar74a, Beh72b, Bei71, BBF74, Bur74e, But72, Cad71, CS71, EHR73, Erd74, FGQ71, Gib72, GS71, GP72, Gra74b, Guy72, Har72a, HS74a, HS72, How71, Kro72, Lit74a, Pec74, Rén73, Sze74, BH74b]. Grassmann [BV74]. gratitude [Lan70]. Grauert [Dou73, Siu71a]. gravitational [AC74, CA74, Jan74a, Mil74p]. gravity [Ede70]. greater [HL74a]. Greek [Bon72b]. Green [Ura71b, Del70h, GP74, Hön74b, Kan73a, Ple72]. grid [Bea74]. Griffiths [Del71b, Kat73c]. Gromoll [Ber71a]. Gromov [Hae71a]. Groot [McD74]. gros [dM72]. grosse [Pop70h]. großen [Sch71f, HRS^{+}73a, Maa71h]. Größen-characters [Maa71h]. Grothendieck [AK70g, Del72a, Gus73b, Har70g, Ill71d, Rog70b]. ground [KMT74h]. Groundwater [Sim73]. Group [Boo74b, Bur74d, But74d, Can74, Coh73a, Coo74, Dem72b, Fár72, GW73, Gre73a, GM71g, Gru70c, Hil74c, HMR73, Knu73g, Lip74h, Par71b, Sta73d, Yam74j, Bas73f, Ber74f, Bro73a, Car70a, DI71b, Ell73, FGQ71, GM71b, GM71h, GM71i, Gru73a, Hai73, HC70i, Hir72d, Ill72e, KMT74b, KMT74c, KMT74f, KMT74j, KN73, Lip74a, Lip74d, Maa71e, Maa71j, Maa71n, Mic72, Mot73, Pop70e, Ros73, Sat71, Seh73, Tak70e, Tak70l, Tit74d, War73, Wil73d, Yam73, vdK73j, vdK73n, Bur90]. Group-like [Gre73a]. Group-schemes [Dem72b]. Groupe [Art70, Gra74a, Gro70g, Che74c, CW71c, GR71, Hir73, Mou72a, Ray70d, Ray70r, Ray70-32, Ray72b]. Groupes [Aze70b, Aze70d, Aze70e, Ber71l, Ber70d, Bor72d, DK73a, Dem70a, Dem70g, Dem70e, Dem70f, Dem70h, DG70a, DG70b, Dem70k, Dem70j, DG70c, Gab70b, Gab70d, Gom70, Gro70c, Gro70d, Gro70h, GS70, Gro70f, Gro70e, Gro70g, Gro72f, Ill71d, Ill72f, Poé74a, Poé74b, Poé74e, Poé74f, Poé74d, Ray70a, Ray70b, Ray70c, Ray71c, Ser73b, Ser73j, Tit71a, Ver73, AdL72, Ber74a, Bor71a, Bri73, Car71c, Car71d, Che74a, CW71g, CW71f, Gat73a, Gro72b, Gro72c, Hec74b, Ill73, Kar73a, Kos71, Mar71b, Ray70h, Ray70i, Ray70k, Ray70s, Ray70-27, Rei74a, Riv72h, San74, Ser72, Spr74]. Groupes-Quotients [Dem70g]. Groups [Ady74, ABK71b, Ati74a, Bar74c, Bau74a, Bau74b, BKW74, BK74, BCC^{+}70, Bra74, Bru74i, Bry74a, Bry74b, Cap73a, CL74a, Coh72f, Cos74, DG70b, Dem72c, Gab70c, HC70d, Hey70c, Hey70l, Hey73, KMT74a, Ker71c, KMSS72a, KMSS72b, Mes72c, New74b, Omo74e, Omo74f, Rob73o, Ste74c, Swa70, Wil73e, Zag72a, dlH72a, vdK73l, AG73a, ABK71d, App73, Ati74i, Beh73, Ben74a, Ber70c, Bli74, Bor70, BK72p, Bou74b, BtD70f, Cas73c, CM74g, Chu71a, CR72c, Cur70, Dem72a, Dem72f, Dix73, EG73, EHR72, Gat73b, Gru70d, Gru70f, Gru70j, Gru70h, Gui72d, Hal73, HO72, Hau73, Her72b, Hey72, Hil74b, Hir72a, Hof70i, cHyH70, iI71, Iwa70, KMT74c, KMT74d, KMT74e, KMT74i, KMT74h, KMT74k, Kan71a, Kei71, Ker71a, Ker71e, Kla73, Knu73f, Lam73, Law73, Lip73, Lip74e, Lip74f, Lip74g, Maa71a]. groups [Maa71l, Mar74a, Mas74a, Mau73, Mes72b, Mes72h, Mil74a, Mil74l, Mis74, Mur72a, Omo74c, Omo74d, Omo74h, Omo74j, Omo74l, Orl72e, Orl72f, Ost70b, PS72a, Pet70c, Phi73, Poé74b, Pri73a, Puk72, Qui73a, Qui73c, RS74, RU73, Rei71b, Rei71n, Rei71o, Rei71r, Ric71, Rog70b, Roi74, Sac73, Sha71m, Spr70b, Spr70a, Sta73e, Sta73g, Sta70b, Sta74c, Ste74b, Ste74a, Ste74d, Ste74e, Ste74f, Tay73, Tay74a, Tod71, Tur74, Van72, Wal73a, ZVC70a, ZVC70b, Zim72a, Zim72i, dlH72b, vdK73k]. growth [Dwo73, Str71]. Grundbegriffe [Wil70b, Wil70e]. Grundeigenschaften [Pit72d]. Grundkörper [Lor70g]. Grundkörpers [Lor70a]. Grundlagen [tDKP70e]. Gruppen [Hey70i, Hey70j, ZVC70d, ZVC70e, ZVC70f, BtD70f, Hey70b, Hey70c, Hey70l, ZVC70a, ZVC70b, ZVC70g]. Grushko [Poé74g]. Guide [Ive73d]. Gundy [Mey72b]. Gustav [Bec73a].

H [Bur74a, Coh74h, FK72c]. H. [Azr72, Bas72, FK72c, Mey72a, Ros72]. Hadamard [WWS72a, Wal72c, vL74o]. Hahn [Sha71a]. halblineare [Has72, Has74]. Halblinearer [AH70f]. Half [Shi73b]. Halifax [Fil73, Law72a]. Hall [Rah74]. Hamilton [Gol74b]. Hamiltonian [CM74b, CM74d, CM74a, CM74e, CM74h, Cop71d, GH74b, JW73a, JW73b, JW73n, JW73m, JW73l, Pec74]. Hamiltonians [JW73f]. Hammerstein [Fes74, Gup74, Her72a, Her73]. Hammerstein-equations [Her72a]. Hamonique [CW71a]. hand [Tau74]. Hanson [ES72a]. Hardy [Her74e, Mor70]. Harish [Coh74b, Kan71h, Van72]. Harish-Chandra [Coh74b, Kan71h, Van72]. Harmonic [Ben71c, Dzh74, Fug72d, GL72, HC70d, Hel72, Oka72, Van72, AB72, Aze70f, Bli74, Bre71h, Fug72c, Han71b, Jan71, Mit74a, Tel71, Meg71a]. Harmonicity [Wei71i, Wei71j]. harmonics [Ner71m]. harmonieux [Mey70c, Mey70h]. Harmonique [Aku73, Far74, Mey70f, AdL72]. Harmoniques [LB70d, Meg71b, Aze70f, Bre74a, GC74, Meg71a, Meg71c, Mit74a]. harmonisables [Meg71d]. harmonischer [Opf74]. Harris [OS70]. Harrop [Sca71c]. Hartree [Baz74]. hasard [MT73]. Hasse [Lor70d, Zim72j]. Hasse-Prinzip [Lor70d]. Hauptvermutung [Mor71a]. Hausdorff [Bab73, Ben74b, Coh74c, Del72e, Del72h, Dup74b, Gar72e, HN74, Kau72, Sch73x, Sch73y]. having [AJ73, Chv74a, FGQ71, Kan73a, Lor74]. heart [Mar73b]. heat [BD72f, LRV70a]. Hecke [All73, Eic73, Wei71k]. Hedlund [Bec73a]. Heidelberg [KGLZ70]. Heins [Ess74c]. Heinz [HK71d, Jam72b]. Heisenberg [AB72]. Held [BCdV^{+}70, BCC^{+}70, ES72c, Kno73b, LLNS70, MR73, Moi70, Mor71b, Mor71c, Ura71c, Bas73a, Bas73b, Bas73c, Bec73a, CFK71, Fle70, HLM71, Hor70b, Hor71, Pet70a, SM74b, Suc70]. Helson [Ben71d, McG72]. Hemi [Bou74a]. Hemi-Questionnaires [Bou74a]. Henkine [Lit74c]. Henselian [Ray70e]. hensélien [Ray70y]. henséliens [Ray70e, Ray70m, Ray70z]. Hensélisation [Ray70u]. Herbrand [Fra70a]. Hereditarily [Jon74, Bel74a]. hereditary [Chv74a]. Heriot [Kno73b]. hermitesch [Sch74c]. Hermitian [Bas73b, Cap73a, Cap73b, Sha73a, Sha73b]. hermitienne [Kar73b]. Herrera [Col70, Dol72]. Herzog [HK71e, HK71f]. Heuristic [Pit70]. Heyting [Luc73h, Luc73g]. Heyting-analysis [Luc73h, Luc73g]. hierarchy [Dev73f, Dev73m]. high [Bet74a, Gol71a]. Higher [Ger73a, Gil74h, Gil74i, Qui73b, AKW73, Ben74a, CD72b, Ger73b, Ger73c, Gil74g, Har70f, Mei72, Wag73, Bas73c, Git72]. higher-order [CD72b, Mei72]. highest [O’M74]. Higman [Gat73b, RHD70a]. Hilbert [Azr72, Bad74a, Ben72a, Cha74d, Cha74c, Chu73, FK72c, Gui72a, Gui72h, Gui72i, Hir72d, Kre70a, Lev73, Mil73a, Mou72a, Ner71d, PT73c, Pat74, Sha71b, Tak70a, Tak70b, Tak70c, Tak70f, Tak70k, Tak70o, Tak70p, Wei73, Zyg71h, Zyg71g, dlH72a, dlH72e, rd74c, rd74d]. hilbertienne [Sko73, Mou72b, Mou72h]. Hilbertiennes [Mou72g, Dix71, Har74j]. Hilbertmannigfaltigkeiten [FK72c, FK72d]. Hilbertschen [Hir74c]. Hironaka [Poe74n]. Hirsch [Den73b]. Historical [ST71g, BV73f, Sta71c]. historique [FS70b]. Hochheben [Pop70m, Pop70p]. Hochschild [vdK73k]. Hoheisel [Mon71n]. höherer [Mei72]. holomorphe [And74e, Din72]. holomorpher [LL74b]. holomorphes [And70, Fer74d, Lel70a, NF74, Nor74b, Rud73, Sad74, Sch70g, Sib74, Sko73]. Holomorphic [CR72b, Din74a, Ric74, Sto70k, AS71f, AS71n, AS71r, BMN74, Dwy74, Gre74, GK72, Luc71f, Nov74a, Sto70j, Wae74a]. Holomorphically [Car72b, Din73]. Holomorphy [HS74b, Aur74, Col74e, Nac74b, Ric74, Wae74a]. Holonomy [Rei74b]. Homeomorphism [Nar74b]. homeomorphisms [CR72d]. homogene [HR74b, Har74j]. Homogeneous [DJ74e, Kan71d, AS74e, Gus73a, Har74i, Mur72b, Mur72f, Oka72, Orl72h, Sha71c, MM74b]. Homogènes [CW71a, Eym72b, Eym72c, MM74b, MM74a, MM74c, Mey74b, Mey74c, Mey74d, Ray70h, Ray70i, Ray70s]. Homological [Bou74b, Ohm73, RHD70b, KS73a]. Homologie [BtD70e, BtD70j, Lem74a, Lem74c]. homologiques [Jen72a]. Homology [Sta73d, And71a, Coh70b, Dol70a, Gla71c, Gru70d, HH73, Hir71c, Hus73, Moo70, Orl72h, Sta73b, Świ70, Zag72c]. homomorphism [Ati74g, Ati74j, Coh74n, Coh74o, Coh74q, EL73]. Homomorphismes [Gro70e]. homomorphisms [Aus71, BV73c, Ste74a]. homotopie [HHP71]. Homotopiegruppen [BtD70d, tDKP70c]. Homotopiemengen [tDKP70c]. Homotopien [Wac72]. Homotopietheorie [tDKP70d, tDKP70e]. Homotopique [Ill71e]. Homotopy [ABK71b, BV73b, BV73c, BV73d, BK72k, BK72l, BK72m, Coh70b, Fre70, Gol74c, Hae71b, Hil74c, Mal74, Mil74o, Poé71a, Sta70h, Sta70i, Sta70f, AHV74, Bec71d, Ben74a, Bre72a, Bro71b, Bro74b, Cha74d, Coh70f, Coh70g, Coh70e, Edw74, Gan71a, Gre74, Hak74, Har74f, KK72, Lem71, Mil74b, Mil74l, MY72, Pet70c, Pet72, Pri73a, Rec71, RS74, Roi71, Roi74, Sch72d, Sta70e, Wan72, Wan74, Zab70, Zab74, tDKP70d]. homotopy-everything [Bec71d]. Honda [Tat71]. honor [Bec73a]. Hopf [And71b, Ara71, Jam72b, Moo71, Pet70b, Sta70j]. hopscotch [Gou72b, Gou74b]. horizontale [MM74d]. Hörmander [Gru74a, Mal73]. horocycle [Fur73]. hulls [HM74a, HM74b]. Hunt [Ben73, Del70f, Han71a, Mey72d]. hydrodynamic [Sat73f]. hyper [Kel74a]. hyper-quantum [Kel74a]. hyperalgebra [KMT74f]. hyperbolic [Bow73, Bré74c, Geh74, Hed72, Kom72f, Kre74b, New74a, SW74b, Tho74b, WP74d, Tay74d]. hyperbolique [BB73, Cha74e, Cho73f]. Hyperboliques [Rav71]. hyperbolischen [Haa74]. Hyperelliptic [Oor74]. hyperfläche [Pop70i]. Hyperfonctions [BS73b, Kan70, Sch70b, Sch70f, Sch71a]. hyperfunction [Mor73, Ōuc73]. Hyperfunctions [Kom73b, Kom73d, Kom73h, Sch70f, Kom73c, Sch70b]. Hypergeometric [MS73e, Fra74]. Hypergraph [Fab74, BE74, Sim74, BRC74]. hypergraphe [Ber74c, Ber74d, Fou74]. hypergraphes [Las74, Mey74a]. hypergraphs [BS74b, Ber74b, Chv74a, Erd74, HL74a, Lov74, Rad74, Ste74h]. hyperharmonic [Fug72c]. hyperspace [Cur74]. hyperspaces [Kee70]. Hyperstability [Hal70a]. hypersurface [Rob71a, Sko72]. hypersurfaces [Bom71, Hir71c]. Hyperzufällige [Sch71m, Sch71n]. hyperzufälliger [Sch71d]. hypoelliptic [Dio74, Net74, Tay74b]. Hyponormal [Sai72b]. hypothèse [Ben73]. hypotheses [WW74h, Ray70n]. hypothesis [Dev73b, Dev73j, Dev73k, Dev73n, Dev73s, Dev73t, DJ74l, Jec71j, Tie72].

I. [Azr72]. iceux [Ray70i]. idea [Gou74b]. Ideal [Jat70a, Coh72a, Jat70c, Mat73c, RHD70a]. ideals [Arn73, BGR73d, BGR73e, BGR73c, Bur74c, DG73, Dom73, Eic71c, Lam71e, Pet74e, Rei71g, Rei71i, Rei71h, Rei71f, Rog70d, Sal73, Ber74n]. Idéaux [Fer74d, Kah73, Pha71, Ber74n]. Identification [Fau70, Åst72, Bel71b, May72q]. identité [Mey71h]. identities [Hir72b]. identity [Bea70i, Hof70m]. ie104 [Dev73c]. ie104-1 [Dev73c]. ie145 [Tur71]. ie145-1 [Tur71]. if [Pop70e]. ihre [Sch71l]. II [DK73b, Hor71, Mey74c, Yor70, Del71c, Bau71a, Bec71c, BKW73, BGJ72, Bou74b, Bru74k, Car74b, Com71, CR72c, Cur71, Day74d, Del69, Gib72, Hei73b, Ill72d, Jac72a, KO74b, Kos70b, Lep72, Lev71b, Lip74c, MS73j, Omo74b, Poe74i, Rho74, Rog70c, SV71b, Taa70a, Tit74i, Wal71d, dB74, vL74c]. II-modules [Bou74b]. III [BH74a, KS73b, MM74a, Bru74l, EHR72, Kan71g, KO74c, Lam72, Lev71c, Lip74d, Ran73b, Taa70b, Tit74j, vL74d]. ILB [Omo74c, Omo74d, Omo74l]. ILLIAC [Wil74a]. im [AT72, AT73, AT74, HRS^{+}73a, Lor70n, Sch71s]. Image [Bad74e, Ray70-31, Siu71a]. images [Dou73, Har74e, Kuh74, McA71, Rob73o]. imaginary [Yam74i]. Imbedding [Orn70, Bel71a]. imbeddings [Kan71c, Kan71h]. Immersions [Ber71m, Sch70h, Wil74e]. immune [Pim73]. impairs [Zyg71g]. implementation [JL71]. Implications [Sch73g, Con73, Yat72b]. implicit [BR73a]. implicites [DM74]. Important [vL71e, vL73e]. imprimitivity [Fel72]. improperly [Nas74, Pay73]. improving [BS74a]. impulse [HB70]. impulsionnels [Mar70b]. inaccessibles [Del73p]. including [OG74j, OG74s]. inclusion [Mil74f]. Increasing [Sch73z, De 73]. increments [Gus73a]. indecomposable [Bel74a, Rog70e]. indefiniment [Cho73a]. Indépendance [Ver74]. Indépendants [Bre73, Art72b, Bre71t, Bre72d, Dup74b]. Independence [Jec71j, Mas71]. Independent [Ben71e, Ben71f, Alb74, Gil74i, Gus73a, Kru73]. index [Ati74g, Ati74j, Ati74d, Ati74f, Bel70, BE74, Coh73a, Dev74, Fre71b, Gra74b, Hai73, HH73, KR72, MS73a]. Indexmethode [HRS^{+}73b]. Indiana [CG74]. indices [Cai70, Cai71]. individual [Hay74]. induced [AdK74b, And74a, Aus73b, Bun71, Dre73, Hir72a, Lip74c, Sta70q, Wei73, Wil73b, Wil73c]. Induction [Pet73, Luc73o, Luc73j, Sca71b, Sca71k]. Inductive [AR72, Fer74c, Fre74a, Lon71, Luc73j, Luc73k, Zuc73c]. industrial [IP70]. Ineffable [Dev73k]. inégalité [Cai70]. inégalités [Mey70a, Car74a]. Inequalities [Din71a, EG74, Geh74, Bur71, Gob74, GH73, Rao72b, Sta72, Sug71]. Inequality [Sim72a, Ben74b, Ess74b, Sim72e, Tay74i]. Inf [Hun74]. infini [Gru73b, Lel72c]. infinie [Coe74, Col73b, Din72, Hir71a, Hir71b, Mou71, Nov73, Nov74b, Rug70]. infinies [Dix74]. Infinitary [Fef72, Kle73b, CD72b]. Infinite [CM74a, HS74b, Her71a, Omo74f, Sta70k, Sta71c, Poé74d, BE74, CM74e, Col74e, Gro70a, Gru73a, HK70, Har74i, Her71d, Her74b, Kin72, Kli70, MS73d, May72b, Nac74b, Rad74, Ram74b, Rho74, Ste71h, Wae74a, Wei71i, dlH72e, Sei72j]. Infinitely [Gui72i, Hey72, Dzh74, Gui72b]. infinitesimal [PS72b]. Infinitesimale [Gab70b, Gab70c, Poe74t]. Infinitesimales [Dem70c, Ill71i]. Infinitesimally [vdK73l]. infinity [JW73h]. influence [GK70]. Information [Agg74, BKW73, Los74, OG74k, OG74l, dF74a, dFP74, HT74, Lan74, OG74e, OG74f, OG74g, OG74o]. Informations [BS74c, Sal74]. Initial [Gil74j, Kaz71, Yat72b, AH70e, AH70g, BR73a, Bax74, Bre72b, Spi74a, Sto74, Stu74, Tau74, Tay74d, Tho73a, Tho74c, Wen74b, Wid71b]. Initial-boundary [Kaz71]. initial-value [Tho73a, Tho74c, Wen74b]. Injective [CR72c, Car72b, Ste71f]. injectives [Bac72]. Injectivity [Coh74o]. Inner [Bea70e, Dev73l, Tay73]. inputs [Jam72a]. insbesondere [AT73]. Inseparable [MV70d]. instabilities [Mor74]. Instability [Shu71b, CGS71, Sat73a, Wed72]. Institute [Bas73a, Bas73b, Bas73c, BCC^{+}70, DF74b, Pet70a, SJS73, DS70a]. integer [Sur72]. integers [Qui73a]. intégrabilité [GC70]. integrability [Hae71b, Kur71]. integrable [Kni71, Mey71f, Mey71g, Sio73b, DD71b]. intégrables [And70]. Integral [Bea70f, CG74, Col74f, HH73, Shi73b, TP71f, Bai70, Ben71l, Bre71f, Cho70, Col71, Dun70b, Fer74a, Geo74, GL74, Gil73, Gil70, Gol73, Gup74, HO73, HS74a, HC72, Hei73b, Hof72a, Kaz72c, KP72, LRV70f, Lew72a, Mat73j, Nas74, Ner71i, Nob71, OG74i, Ran73a, Sel73, Tay74j, Top70b, Top70k, Trè74, TP71a, TP71e, TP71j, Wen74a]. intégrale [Glo70, Mey71c, Roo74]. Intégrales [CW71a, Zyg71c, Nor74b, Nor74c, Ver71, Zyg71a, CW71d, Del74a, DDM70, DD71a]. Integralformelmethode [HRS^{+}73c]. Integralformeln [HRS^{+}73c]. Integralgleichungen [Her74a]. Integrally [Dav73, Kap70]. Integrals [MS73f, Mey72g, Ner71j, Ber74e, Bic73f, Ner71k, Ner71l, Zyg71c]. Integration [AC74, Bic73b, Sei72c, Smi74, BI70c, Col70, BS74a, Bea74, Ben71i, BL74a, FPC74, HL74b, How74, Jan74a, Kro74, Moo74, Sto71, Wie74, vdH73]. intégraux [CW71b]. Integro [NW74, Esp74]. integro-differential [Esp74]. Integro-Differentialgleichungen [NW74]. Integrodifferential [AT74]. Integrodifferential-und [AT74]. Integrodifferentialgleichungen [AT72]. interconnection [ADT72]. interest [Pit70]. intérieure [Cho73b]. Intermediate [MV70b, Lam73]. internal [JW73b]. International [AHiN74, Cur72, DK73b, GUW74, Hor70b, Hor71, Kui73a, KS73b, New74b]. Internationale [Aku73]. interpolation [CR73b, Dho74, Dru72, Gab72, Gla71a, Sha71g]. interpolationsquadraturen [Eng73]. Interpretation [Luc73e, Luc73i, Luc73h, Luc73g, Luc73l]. Interprétations [Gir73a, KR72]. Intersecting [Chv74a]. Intersection [Bur74d, LM70, Bur90]. Intersections [BGI71, Del73f, Gro71a, Gro71b, JG71, Del73c, Hei73b, Kat73b]. Intertwining [BS73a]. intervals [Ste71h]. Introduction [AK70g, AS71i, Bau71f, Ber71e, Ber74o, Bic73c, Car71b, Coh74p, Cop71c, Del70e, Eic71d, Eym72d, FS70b, Fug72e, Gar72f, Gil74k, Gui72f, HC70e, Ive73c, Jen72e, JW73d, Kan71k, Ker71b, Knu71g, Knu73e, Kre73f, Law72b, Luc73m, MM74b, Meg71e, Mes72g, Mey70d, Mil74g, Mur72e, Ner71e, Neu74, Nor71, Nor74c, OG74m, Par72f, Poe74l, Poe74m, Ray70v, Riv72f, Sat73e, Sca71d, Sch73h, Sha71h, Sil70e, Spi74b, Sto70g, SE70f, Tak70d, Tre73c, Uen74, Wal71a, Wid71b, Wil73a, Yam74b, Yam74q, Zim72k, Zyg71d, dlH72f, rd74b, vL71f, vL73f, Bec74a, Kom73d, PT73a, TP71b, Ber74o, Del70e, Eym72d, Jen72e, Lip74b, MM74b, Mey70d, Poe74m, Ray70v, Zyg71d]. Introductory [Kom72b]. intuitionism [vD73]. intuitionist [Lam72]. Intuitionistic [Sca71g, Tro73a, Fri73c, Luc73p, Myh73, Sca71a, Sca71e, Tro73b]. intuitonistic [Sca71b]. Invariance [Beh73, Coh74q, HM74a, Beb74, CS70, Sie74]. Invariant [Bel71a, BV73d, DP73, Bro73b, Bru70, For74, Hor70a, Kau74, KR72, Kre70b, Lan73a, Orl72b, OS70, Sac74b, Sch73w, Sch73t, vdK73h]. Invarianten [Gey74]. Invariantes [Eym72g, Art71, Eym72e, Eym72f, Har74j]. Invariants [BGM71a, Kaw72b, Kum74, Die72, Gan71a, HR74b, KK72, Rei74b, Sul71a, Sul71c, Wag71]. Invarianzeigenschaften [Sch71d, Sch71b, Sch71o]. invers [GS74a]. invers-isotone [GS74a]. Inverse [DeV72e, LRV70a, LRV70b, LRV70e, Yam74r, Beh74c, BK72l, But74b, LRV70c, LRV70d, Osb71]. inverses [Sta70l]. inversibilité [Cho73e]. inversibles [Ray70-29]. inversion [Sil70d, vL74c]. Invertible [Ive70b]. Investigation [Tro73a, IDS70].

Involutions [Pet72, AHV74, Boa72, Iwa70]. Involutive [Hof70n, Hof70h, dlH72b, dlH72c]. involving [Aul74, Bra72b]. Ipli [Jat70a]. Ipli-rings [Jat70a]. IR [Riv72g]. Irreducibility [KS72, Kin72]. irreducible [GM71h, GM71i, Hir72a, Sil70c, Sil70f, Tit74n]. irreduzible [Pop70h, Hey70h]. irreduziblen [Pop70g]. irregular [BD74, Hsi73, Iwa71a, Kub73, Sib71b]. Irregularité [Gér74]. isolated [Gli72, Ham71]. isolating [McG73]. Isolement [Cou73]. Isometric [CS71]. isometries [HO72]. isomorphe [Fou74]. Isomorphism [Ber74b, Smo71i]. isomorphisms [Ura71a]. isotone [GS74a]. isotopie [Mou72f, Poe71b]. isotopies [Wag71]. isotopy [Cer71, Hat73]. isotriviaux [Ray70-32]. isotropic [Kan73a]. isotropy [Ati74i]. Itô [Mey71b]. Itô’s [Man72]. Iterated [AT70, BV73e, DJ74g, May72h, May72i, Mil74h, Zuc73c, HS72, Mil74m, Sta70o]. iterates [ES72b]. Iteration [Wac71a]. iterations [Rho74]. Iterationsverfahren [GS74a, Sac72, Wac72]. Iterative [CG71a, GW74b, Pat74, rd74c, rd74d, Sac72]. itself [Bin74]. IV [MM74c, Mey70b, Bru74m, EG73, KO74d, Lip74e, MAB^{+}70, Wil74a, vL74e]. Iwasawa [Kan71e]. IX [vL74f].

J [HK71e, HK71f, Sch71c]. J. [Azr72, Bas72, Car71a, Kos71, McD74, Poe71b]. J.-P. [Car71a]. Jacobi [Ber71f, Ber72b, Sch73-28, Sch73-29]. Jacobian [Ray70-30]. Jacobien [Ray70-30]. Jacobiens [Pha71]. Jacques [Den73b]. Jamaica [CFK71]. James [Ami72]. January [BK74, Law72a]. Japan [MP73a, Ura71c]. Jensen [Dev73f, Dev73i, Dev73m, DJ74f, DJ74g]. jets [Boc72, Poe74k]. Jeux [Lem70, MT73]. John [CFK71]. Joint [Sch70h, Top70j]. Jordan [Ste74d]. JPA [Ber71c, Ber71g, Ber71h]. Juin [dFP74]. July [BCdV^{+}70, Chi71a, Cur72, DK73b, Kui73a, KS73b, Nac74a, SJS73]. June [BCdV^{+}70, BH74b, Bec73a, CFK71, HLM71, Moi70]. Jyväskylä [LLN74].

K. [Mey71b]. Kac [Csö73]. Kahlerian [LB70l]. Kähleriennes [LB70c, LB70l, LB70m]. Kakutani [Pap70, Pat70]. Kalamazoo [HS71]. Kampen [Che74c]. Kan [Dub70d]. kanonische [HK71a]. Kansas [BR73c]. Kan’sche [GU71h]. Kanschen [GU71d]. Kanter [Fer74b]. Karcher [FK72c]. Kas [AS71c]. Kasteleyn [Lit74a]. Katata [Kom73c]. Kategorien [GU71b, GU71i, GU71j, GU71k, GU71l, GU71m, GU71q, Sei72d, Sei72e]. Kategorientheoretische [tDKP70e]. Kategorische [Sei72f]. Kelly [Lin71]. Kempf [Szp73]. Kentucky [HS74b]. kernel [CL74b, Sta73c]. Kernels [PS72d, Bea70f, Del72i, KŌ73a, PS72b, PS72c]. Kesten [Bre71t]. Kiehl [Dou73]. killed [DJ74f]. Killing [Ste70i]. kinds [Gon74]. Kinematic [Sch73f, Sch73i, Fin74e, LRV70d]. kinematics [Sch73j, Sch73l]. Kirby [Mor71a]. Klassen [BtD70b, Wil70h, Wil70i, tD72a]. Klassenzahlrelationen [Hir74c]. klassische [Pop70d]. Kleene [Fri73c]. Kleiman [Szp73]. Klein [AG71a, AG71b, Bre72b, But74d]. Kleinian [BK74, Ber70c, Ber74f, Mar74a, Mas74a]. Kneser [Poé74g]. Knight [DMW73a, Mey71a, Mey73c]. Knot [Ker71f, App73]. knots [DD74e, dB74]. known [Ost70c]. Kobayashi [Roy71]. Kobordismentheorie [BtD70h, BtD70l]. Kobordismus [BtD70a]. Kodaira [LB70i]. Kohomologie [BtD70m, BtD70j]. Kohomologiegruppen [GR70]. Kollektive [Sch71l]. Kollektivs [Sch71g]. Kollokation [Sch74a]. Kolmogoroff [Sch71i]. Kolmogorov [Har72b, Smo71j]. Kolmogorov-automorphisms [Smo71j]. Kombinatorische [ZVC70f]. Kompakt [Sei72g, Sei72d, Sei72e]. Kompakten [FK72b]. kompakter [Gra74c, Hey70b]. komplementäre [Vel72]. komplex [ZVC70g]. komplex-analytischen [ZVC70g]. Komplexen [GR70]. komplexer [Gra74c]. Kongruenzklassengeometrien [Wil70f]. Konstantenerweiterung [Pop70m]. Konstruktion [End74, Wac72]. konstruktive [Sch71c]. Konvergenz [AH70c, Has72]. Konvergenzbegriffe [AH70d]. Koordinatisierung [Wil70a]. Körpern [Lor70h, Lor70d, Lor70i, Lor70j, Lor70k]. Körpers [Lor70l]. Krein [Hey70g]. Krickeberg [Kaz72b]. Krieger [Smo71d]. Kripke [Smo73]. Kritik [Sch71l, Sch71p]. Kronecker [Ben71g]. Krull [BMRH73]. Kubo [Tak70e]. Kuiper [BŁ71a]. Kullback [FL70]. Kummer [GM71f]. Kunneth [Smi70b]. Kunz [HK71b, HK71c]. Kunze [Ver71]. Kuo [BŁ71a]. Kurepa [Dev73k, Dev73n]. Kurve [Pop70h, Pop70g, Pop70f, Pop70o]. Kurven [Pop70p, Hir74c, Pop70a, Pop70k]. Kutta [Bet74a]. Kuzmin [Sch73-27, Sch73-32]. Kyoto [MP73a, Ura71c]. Kyoto/Japan [Ura71c].

L [Dev73l, Sch71c]. Lacets [Lem74a, Lem74b]. Lacunarity [EHR72]. lacunary [Kau72]. lag [TP71i]. Lagrangian [CM74c]. Lagrangien [God71a]. Laksov [Szp73]. l’Algèbre [Poé74a]. L’Analyse [Aku73]. langage [Arn70]. language [dB70]. Languages [Eng71b, CD72b]. Laplacien [BB73]. laplaciens [DDD71, Geb74]. Large [Dev73o, Mon71f, PT73i, AR72, Dev73p, Dev73r, Fre74a, Mon71a, Mon71e, Mon71g, PT73c, PT73b, PT73g, PT73h]. last [Chu74b]. lateral [Hön74b]. Latest [Guy71]. latin [Lin74, vL74n]. Lattice [Gar73, Gow74, HMS74a, Kei71, Sco71]. lattice-ordered [Kei71]. Lattices [RHD70c, Rog70c, Hag74, Jam70b, Jam70c, Lac72, Rog70a, RHD70a, Rog70e, Sco72, vdK73b]. Laurent [cH73, Ran73b]. Lausanne [Rob73j]. Lawrence [BR73c]. Laws [PT73c, PT73b, PT73i, CM74g, Kel74f, PT73g, PT73h, Sch71j]. Lawson [Ros72]. Lax [Meu73, Kel74f]. layer [O’M74, Wal74m]. leading [Din71a]. Least [Mon71h, FGQ71]. Lebesgue [Bru74f, Hei74a, Jec71k, Mey74h]. Lecture [Ati74c, Ati74k, Ati74e, Ati74g, Ati74j, Ati74d, Ati74f, Ati74h, Ati74i, Ati74b, HK71d, HK71e, HK71f, HK71b, HK71c, HK71h]. Lectures [Bea70g, BK74, BGJ72, Bot72, Dem72c, Deu73c, Hae71a, Jec71l, KT71, Kos70a, Kos70b, Rin72b, Rob73j, Smi70c, Taa70a, Taa70b, Tay74e, Väi71c, Wal71b, vD73, Bre71p, Sat73a, Yor70, Hof72b, Hof72c, Pit72g]. Lées [Mor71a]. Lefschetz [Del73i, Har70g, Kat73a, Kat73e, Pow70, Trá74c]. Left [Jat70b, Jat70c, Tac73e, Jat70a, Tac73a]. Leitfaden [Ive73d, Meg71g]. Leja [Lel72a]. Lemma [AS71s, Moi74, Dev73b, Mon71i, Sch73o, Sic72, ST71m, Str74d]. lemmas [ABK71e, BK72e, BK72f, BK72s, Lin71, Rei71d, Wei71c]. lemme [Del70j, Lel72a]. length [AS74d, AK70a, HL74a]. lengths [Geh74]. letters [AS71g]. leurs [Cha73, Jen72f]. Level [Sto70h, Dwo73, Gus73a, Mes71b]. Levi [Her74d, Jos74]. Lévy [BB70, Ben73, Mey73h]. Lewy [And74g]. Lezioni [Wei71e]. Liapunov [Kat71]. Libres [Lem74f, Lem74c, Kos71]. Lie [BGR73d, Beh73, BGR73d, Dem70d, Dix74, Gab71, God74a, God74b, Gro70h, HC70i, Hir72a, Kum74, Lic74, Mes72h, Mur72f, Mur72g, Omo74c, Omo74d, Omo74f, Omo74g, Omo74l, RS74, Roi74, Seh73, Sta74c, Ste70j, Ste70k, Tod71, Wil73d, Wil73e, dlH72a, dlH72g, dlH72b, dlH72c]. Lie-Algebren [BGR73d]. Lieberman [Dol72]. Liée [Mor70]. lies [Dez74, Pet74a]. Lifting [Pop70p]. Liftings [Bic73d, Tul70b]. light [AS71n]. like [Cha74b, Gil73, Gre73a, Iga71]. likelihood [WW74a, WW74b, WW74c]. Limes [BtD70m]. Limit [Day74b, FM71, Hon74a, PS72b, PS72d, Hof72a, Mar73b]. Limit-colimit [FM71]. Limit-operators [Hon74a]. limitation [GK70]. limite [GV72a]. limited [Ede70]. Limites [Bad70b, Lio73f, Lio73h, Lio73d, Mey73e, MM74d]. Limits [BK72m, BK72k, BK72l, Bre71g, Bre71l, Din74a, Fer74c, Hil74b, Mey73e]. Lindelöf’s [Her74b]. Line [HS72, Hem72, AHK74, AM71, AS71t, Gia72, KMT74d, Wal72e, Wal74m]. line-connectivity [AM71]. linéaire [Bru74j, Glo70, Mou72a]. Linéaires [Bad74f, Rav71, Bad70f, Lit74c, Mar71c, Riv72h, Bad70g]. Linear [Bea70h, CL74a, CM74d, Cop71d, DeV72c, Dho74, Fin74i, Fin74j, Ive70d, Jam70b, Jam70c, Omo74h, Ost70b, PT73i, Pat74, Tur73a, Yam74p, vL71g, vL73g, AH70g, BJ70, Ben70, Bor70, Bor71b, Bra72a, BE73, Dav71, DeV72d, Dix73, Dol70c, Dud70, GH74a, Gil74i, Gol71a, Gup74, Har71b, HL74b, Hau72, Hes74a, Hön74b, Ive70c, Jam70e, Jam70f, Jam70g, Kan73b, Kat70b, Kaw73a, Kaw73b, Kom73b, Koz72, Lel71b, LS72, Lov70, May72q, Mei72, NS72, NS71, Nas74, Oku71, PT73g, PT73h, Pet74d, Sim72c, Sim72e, Stu74, Tho74b, Tho73b, Trè74, Was71, Wed72, Zim72m, dB71b]. lineare [AH70c, Has74]. linearen [Haa74, Mit74b]. Linearer [AH70g, GS74a]. Linearisierung [Bul72]. Linearized [LRV70c, LRV70d]. Linearly [Pet74b]. L’inégalité [FL70]. lines [Aur74, DJ74e, DJ74j, Lor74, vW73]. link [Kau74]. Liouville [Eva71, Fai74, Leb73a, Rab71]. Lipschitz [Jac70]. lisse [Art73b, Art73f]. Lisses [Gro70g, Ray70a, Ray70i]. List [OG74n]. little [May72j]. Littlewood [Ste70a]. Liverpool [Wal71c, Wal71d]. Ljusternik [FNSS73a, FNSS73d]. Local [AS74d, DD74c, GJ72b, Ive73b, JL70c, Jac72c, Jac72d, McC71, Ter70, Tho71b, Ura71a, Beh74c, Car73a, Cla71b, Fef72, GM71b, HC70f, HC70g, Hic73, IW73, Itô72, Kaz72a, Kaz72b, Kee74, Ray70e, Ray70y, ST71j, Trè74, GU71k, GU71m, Ray70-30]. locale [Gri72, Ray70-30, Trá74a]. Localement [Aze70d, Bor72d, AdL72, Car71c, DD71b, Ray70q, Ray70-32]. locales [DDM70]. Localization [Hil74c, Hil74a, Sta73e, Ati74f, GM74]. Localizations [BK72m, Mat73k, BK72q, Bou74b]. localizes [Ben74a]. Locally [Hey70c, Hor73, Kel74c, Ken73, Ulm71, AG73a, De 73, Din73, HM74b, Knu71h, Ric71, Tel71, You70]. location [Ben72b]. Locaux [Mey71e, MM70, Mai71, Ray70e]. Locus [Tho74a]. Loewner [Tit71b]. Löf [Sch71s]. Logarithmic [Kno73a, Kno73b, Lev73, MS73c]. Logic [Boo74b, Hod72, Law72a, MR73, Coo71a, Lam72]. Logics [Fin72, Gab72]. Loi [Art70]. lois [Fer74b, MM74d]. Łojasiewicz [Poe74q]. lokal [GU71q, GU71k, GU71l, GU71m]. lokalen [Lor70i, Pop70c]. lokaler [Wac74b]. Lokalkompakter [Hey70c]. London [Hod72]. long [Coh70b, Wan70]. long-range [Wan70]. look [Fel72]. Loop [May72h, Rec71, Sta70b, BV73e, May72i, May72b, Mil74h, Mil74i, Mil74m, Moo70, Sta70k, Sta70o, Sta71c]. lorsque [Pet74a]. loss [WW74f]. Lösung [AT72, Gek72, Her74a, Kre74c, Mic74, NW74, Sac74a, vF73, Rau74]. Lösungen [Ade72, Gek73, Pit73]. lower [Eis71a, JW73e, Mon71j, Sta73f]. lowest [Baz74]. Lucas [CHS72]. Luminy [dFP74]. Lurie [Mey73a]. Lusin [Del71e, Del73s]. lusiniens [BI70b]. Lyapunov [Auc73, Wil72].

M [Bom73, Fer74b, Ver71]. M. [Azr72, Col70, Csö73, Del73n, Dol72, Mey73i, Ray71a, Sch74b]. Macaulay [Abh73, HK71a, Hoc73, Mat73a]. Macaulay-Rings [HK71a]. machinery [Gru70k]. machines [Str74b]. Mackey [Fel72, HN71i, HN71l, HN71j]. Mackey-Complets [HN71l]. Mackey-Convergence [HN71i]. MacLane [Coh70a, And71b]. macroscopic [Smi74]. Madras [GUW74]. Mahlo [Glo73]. Mai [Bru71c]. Main [Sto70f, Coh72g, Moz71a, Sch73-28, Ray70x]. Maine [DS70a]. mais [DD71b]. major [Bar71a, Bar72, OG74s]. Majorations [Gru74a]. majorization [New74a]. Mal [Bru74b]. Malgrange [Poe74p, Bol74, Łoj71, Mat71a, Nir71, Pis74]. Malliavin [Ben71h]. Mandelbrojt [Mey73i]. maneuvers [OTF70]. manifold [Cha74d, McG73, Wel71a]. Manifolds [Bro71b, CR72d, CRW72, Háj73, Hsi71, Kui71a, Orl72j, Pop74b, Sto70k, AB72, BV74, Cer71, Cha74c, FK72c, Fre71b, Gil70, GL70, Hir71c, KR72, Kin72, KS71b, Lan72, Mah74, MR71, Orl72a, Orl72h, PŠ70, Sul71c, Tak71c, Wol71, Woo74b, Woo74c, Woo74e, Woo74d, Zag72c, Kui71a]. Mannigfaltigkeit [FK72a, Pop70i]. Mannigfaltigkeiten [Pop70l, FK72b, Pop74a]. manoeuvrabilité [Con70]. Many [Kel72c, Lew72a]. Many-variable [Kel72c]. map [AS71r, Bin74, Bom70, Mat71c, Mil74j, Sac74b]. map-germs [Mat71c]. Mapping [Bel74a, Väi71d, Gup74, HC70h, Sch73a, Yam74r]. Mappings [Fle70, Väi71c, And74c, BMN74, BL74b, Caz74b, Cha74b, Ear74, Gar73, GS74b, GH74c, Hea74b, Kel74e, Kur70a, Lel74a, Lel71b, Lev71b, Lev71d, Mat71b, Nad70, Sic72, SR73, Väi71a, Väi71f, War70, Yam74l, Yam74n, Yam74s]. Maps [Sta70m, Sta70l, Sto70k, Tim71, Ade70, AS71f, AS71n, AS71o, AS71p, AS71q, AS71u, Bar71a, Bar72, CRW72, DK74, Fra70b, GQ71, Hal70b, Ive70b, Mah74, Ohm73, Por71b, Pow70, Sig71a, Sta70q, Sto70j, Tho71a, Wil74d]. March [DF74b, ES72c, Kno73b, Mor71b, Pet70a, SM74b, Suc70]. Marek [Fer74b]. Margret [HK71g]. Marie [Art72a, Art73j, GR71]. Markoff [Ros71]. Markoff-prozesses [Ros71]. Markov [Doo71, Fuk73, Geb74, Gus73a, Kan73a, KŌ73a, Kot73, MM74b, Mey73b, Mey73c, Mey73f, Mey74f, Nag72, Sca71f, Sil74e, Spi74b, dM72]. Markovian [DMW73b, KO73b]. Markovienne [Fau70]. markoviens [MM74b, MM74a, MM74c, Mey74b, Mey74c, Mey74d]. Marseille [dFP74]. Martin [Tak70e, Bre71d, Bre71f, Bre71e, Bre71m, DJ74h, Jan71, Sch71s]. Martin-Löf [Sch71s]. Martineau [Pis74]. Martingale [Bur71, Din71a, DD71b, Mey72f, Smo71a, Sch71f]. martingale-convergence [Din71a]. Martingales [Cha73, GC70, Mey72g, Cai70, DDM70, Itô72, Kaz72a, Kaz72b, Kaz73, Kni71, Mey70a, Mey71f, Mey71g, Rao70, SV71a, SV71b, Din71b]. Maryland [Eic71e, GL72, Her71a, Hor70b, Hor71, Maa71r, Yor70]. Maslov [Mal73]. mass [GK70, Sch71p]. Maß-Wahrscheinlichkeitstheorie [Sch71p]. Massachusetts [KMSS72a, KMSS72b]. Massey [Sta70n, Tho70]. Mathematical [BK74, Gli70, Hod72, Hor70b, Hor71, LLN74, Man71, MR73, Mes71a, Mes71b, Nic73, PT73d, Gol71d, Sat73f, Seg70, Tur73b, dB70]. Mathematics [Hol74b, GM72]. mathématiques [Arn70, Car72c, Car74b]. Mathematischen [AT72, AT73, AT74]. Mather [Che74b, Łoj71, Poe74o, Poe74p]. Matiasevitch [Azr72]. Matrices [WWS72a, Wal72c, But74b, Hed71, Hof72a, KP72, Maa71p, Osb71, Rho74, Wal74k, vL74o]. matriciels [CW71e]. Matrix [Gol71d, ESS74a, KS71a, Pra70, Wel71a]. Matrizen [Boh74]. matroid [Las71]. Matroïdes [Bru71c, Fou71, Las71, Bru74j, Bru74l, Bru74k, Bru74m, Bru74o, Bru74n, Bru74o, Bru74p, Bru74q]. Matroids [Wel71b, Din71c, Ing71]. Matter [BK72a, BK72g, BK72h, BK72i, BtD70g, BGG^{+}72, CR73a, Del73a, DJ74b, DJ74d, Ill72a, Leh73a, MS70a, MS70e, Rob73a, Rob73l, Ser73a, Ser73l, Tit74c, Tit74o, Zuc73a, vL71a, vL71d, vL73a, vL73d]. Maximal [Bar71a, DG73, Hey70i, RHD70d, Ste71d, SE70g, CWW74, Rog70d, SE70a, Zas73, Hey70j]. Maximaux [Gro70g]. maximization [BD72f]. Maximum [Dun70b, WW74f, WW74e, And73b, Dan74, Sty74, WW74b, WW74c, WW74d, dV70]. maximum-minimum [dV70]. Maximummethode [HRS^{+}73d]. May [ALW72, Din71b, Fle70, GG72, HS73b, HS71, Sta74a]. Mayer [Cap73b, Jou73b, RU73]. Mazur [Mou72h, Luc71e]. Mazurkiewicz [Del71e]. MCC [LZ73]. MCC-Verfahren [LZ73]. McMillan [Smo71l]. Mean [Mon71k, Mon71c, Mon71e, Mon71l, Mon71m, Tul70a]. meaning [Eng71a]. Measurability [Bic73e, Gow74, Jec71k]. Measurable [Par72g, Sio73b, BN73, Par72l, Par72k]. Measure [Gar72b, Top70k, Top70q, Bic73a, Gar72e, Goo71, Gro70a, Gui72j, HK70, Kak73, Kli70, Kre70b, Kri70, Lel74c, OS70, Rya72, Sai71c, Sch73w, Sch73-33, Top70a]. measure-preserving [Sai71c]. measure-theoretic [Goo71]. Measures [Bic73b, OG74l, Sio73d, Ang71, Bea70j, Bea70k, Ben71j, Bru70, GW72, Gri73, Gui72e, Hey72, Hor70a, KN73, Kli70, Lan74, Loe74a, OG74q, OG74r, Sac74b, Sch73t, Sch73y, Sch73-30, Sio73c, Top70i, Top70d, Top70l]. mécanique [Car74a, Geb74]. Mechanics [Cur72, Kel74a, Kno73a, Leb73b, Oli71, Rue70, Sma71a, Sma71b, Tho73b]. Medial [Mey73e]. médiales [Mey73e]. Medium [Dev73p]. Meeting [Din71b, BK74, Gra71b]. mehrdimensionalen [Sch74a]. Mehrfache [Sei72h]. Meijer [MS73g, MS73i]. Melbourne [Hol74b]. Memorial [Bas73a, Bas73b, Bas73c, Pet70a]. Memory [MC70, Col73a]. Mergelyan [Fre71c]. méromorphe [Lel72d]. méromorphes [Dol71a]. Meromorphic [AS71b, AS71j, AS74f, AS71t, Bec74a, Beh74b, Fuc74, JLP73, Kim73, Rub73]. mesurable [Del74b]. Mesure [Bad74a, Bad74e, Bad74d, Dup74b, Gér74, dF74a, Del70j, Kri71, Mey72a]. Mesures [Bad70c, Bad70g, Bad74e, BC74a, Bad74f, Bru74d, Del72e, Del72h, Hun74, AdL72, Art71, BI70a, Bad70b, Bad70d, Bad70f, BBB70, Buc73b, Fer74b, Lan74, Mok72]. Metabelian [Bau74a, Bry74b]. Metamathematical [Tro73a]. Metastable [GM74, Mil74d, Mil74k, Mil74m]. metatheorems [Pit70]. Method [Jec71l, AG74, CG71b, DD74d, Hay74, Hed72, Jan74a, Jec71m, Kin74, Lac73, Lit74a, Mei72, MC70, MM74h, Rub73, Sei71, Ste74g, Wal72e, Wal74m, Wen74a, Yev70, You74]. Methode [Gor73, Lio70, Sko74, Wal74a, Wal74b, Wal74c, Rav71]. Méthodes [LM70, Tem71a]. Methods [CG74, DD74a, Gil74b, McA74a, Moi70, Pat74, Zim72c, Auc73, BR73a, Bea74, But74a, Cia74, CG71a, Dan74, Dup74a, Eic70, ESS74a, ER70, Fin74g, Fre73, Fri73c, Gil74e, GL74, Gol71c, Gol71d, Gou72b, Gri71a, Gru73a, Hag74, Her72a, Kad70, Lam74, Mil73b, O’M74, Per74, Rue70, Rya72, Sac72, Sim73, Spi72, Ste74j, Sto74, Stu74, SW74b, Tau74, Tho73a, Tho74c, Tre73f, Wen74b, Wid71a, Zie74, Zlá74, rd74a, rd74c, rd74d, rd74e, vdH73]. metric [Bea70a, Har74e, Hav74, Nar74a, Nar74b, PT73e, Par72d, Ric71, Roy71, Sin73, SR73, You70]. Metrical [Sch73-29]. Metrics [Roy74]. metrizability [Aul74]. Metrization [Har74e, Hod74]. Meyer [Ber71a, Sch71w]. Michigan [ALW72, GG72, HS71]. microfunctions [Kom73g]. Midwest [MAB^{+}70, Gra71b, Gra71a]. Midwestern [Suc70]. Milgram [Mac70]. Milnor [BT73, Del73h, Hum73]. Minimal [Bar72, Sha71i, Wil74e, Aus71, Bal70, Bec74a, Bha70, Bre71a, Bre71e, Bre71m, Eis71b, Ken73, Mar70d, Pet74a, Pet74e, Sha71k, Ste70l, Wu71]. minimales [Bom71]. minimality [Chv74b, Háj73]. Minimax [Lov74]. Minimisantes [Eke70]. Minimisation [CGN71, Fau70]. Minimization [CGN71, MC70]. Minimum [LM70, Cho73g, HB70, Mos70, dV70]. minimum-impulse [HB70]. Minkowski [Sch73f, Sch73i]. Minneapolis [HS73b]. Minnesota [HS73b]. minus [GM71b]. Miscellaneous [Gan71e, Omo74i]. Miscellany [Bau71g]. mise [Mey73h]. Mises [Sch71g]. Mitiaguine [Lit74c]. Mitteln [Opf74]. Mixing [Smo71j, Kak73]. Mixte [Che74f, Cha74e]. Mobile [PMN70]. modal [Gab72]. Model [DMW73b, Adl74, Dev73l, Fin72, Jam72a, Kun73, Luc73o, Luc73n]. model-reference [Jam72a]. Modèles [GR72]. modeling [Dup74a]. Models [Fel71d, Hir74b, BD72f, Cut72, Fel71a, Fel71c, Fri73a, Gai72, IP70, Jec71m, Jec71n, Jec71u, Jec71t, Par72a, Smo73, Tho71b]. Modern [Taa70b, Kos70a, Kos70b, Taa70a]. Modifications [Moi74]. modified [Csö73, Gol71b, Kro74, Moo71]. modifizierte [Sch71k]. Modul [HK71a]. Modulaires [Del73e, Maz73, Del71a, Dix71, Ser73f, Ser73k]. Modular [Bir73, CL74a, Cas73b, Cur70, Eic71e, Eic73, Kat73g, Lan73b, Maa71i, Maa71r, Ogg73, PS73, Shi73b, SD73, Tak70f, Tak70p, Dwo73, Eic70, Eic71a, Hir72d, Maa71c, Maa71e, Maa71j, Maa71n, PŠ70, Seh73, Shi71b, Tak70a, Tak70b, Tak70e, Tak70g, Tak70h, Tak70j, Tak70l, Tak70o, DK73b, Kui73a, KS73b]. module [Cho73g, Gru70l, HK71a, MP71, Ray72a, vdK73g, vdK73i]. Modules [DR73, Fai72, Fin74k, Hof72c, Ill71j, Jen72f, Pro74, Ste71e, Zim72l, Bou74b, Dem72f, Eic71b, EE73, Fuc72, Gai74, Hoc73, Jen72b, Jen72j, Jen72a, Koh72, Lam71d, Mat73b, Mat73i, Mat73o, Mat73n, RHD70f, ST71b, Ste71b, Tac73i, Rog70d]. Modulflächen [Hir74c]. Moduli [Abi74a, Mon71e, Mon71f, Mon71k]. Moduln [LL74a]. modulo [BDF73, Mey70e]. Modulräume [Pop74a]. modulus [Väi71e]. Mokobodzki [Ben73, Mey71c, Mey73e]. Mokobodzky [Rev70]. Moments [Kri73, HK71i]. Monadic [MS70d, MS73j, BS73c, Büc73a]. monads [Dub70e, May72l]. Monge [Sty74]. monochromatic [GS71]. Monodromie [DK73a, Gro72f, GR72, Trá74a]. monogenic [HO72]. monograph [WW74g]. monoid [Por71c]. Monoidal [May72k]. monomorphisms [Dub70f]. Monothetic [Bro73a]. Monoton [GS74a]. Monotone [Kei74]. monotoner [Kol72]. monotonically [Hea74c]. Monotonicity [Col74c]. Monotonieeigenschaften [Tör73]. montées [FS70f]. Montpellier [Aku73]. Moore [But74b, Smi70b, Smi70c]. morasses [Dev73g]. Morita [Han73, RHD70e, Tac73f]. morphisme [Art73h, Art73g, Art73f, Art73i]. Morphismes [Art73e, Ler71, Ray70x, Gro72d, Ray70t]. Morphisms [Ive73b, AK70e, AK70f, AK70i, Gra74e, Ive73a, Ive73e, Ive73g, Tho71b]. Morrey [Fre73]. Morse [FNSS73e, Kui71b]. Moser [ES72a]. Mostowski [Fel71c]. Motifs [Dem71]. Motion [DMW73b, GK70, Kni71, Kni74]. Motivation [BV73f]. mouvements [Car71b]. movability [Kee74]. moving [Kaz71]. moyen [Pet74a]. moyennables [Eym72b, Eym72c]. moyenne [Mar70b]. Moyennes [Eym72e, Eym72f, Eym72g]. MP [Luc73h, Luc73g]. MU [Laz72]. Multi [Bea74, Bro74c, Cod71, Fai74, IP70, KO73b, Mar70b, Mes71b, Nad70, Pow70, Sch74a]. multi-branch [IP70]. multi-dimensional [KO73b, Sch74a]. multi-impulsionnels [Mar70b]. multi-level [Mes71b]. Multi-off-grid [Bea74]. Multi-parameter [Bro74c, Fai74]. multi-step [Bea74]. Multi-valued [Cod71, Nad70, Pow70]. Multidimensional [LRV70e, Kot73]. multilinear [Kel74e, Lin71]. Multiple [BL70, Coh73b, Gre73b, Kat70a, Sei72h]. Multiple-stage [BL70]. multiples [Cai70]. Multiplicateurs [CW71e, CW71d]. Multiplicatif [DG70a, Gro70c, Gro70d, Gro70e, Ray70b]. Multiplication [Shi73a, Hof70c, Hof70e, Hof70o]. multiplications [Sta70l]. Multiplicative [Hof70p, Mon71p, Smi70d, Ati74e, Gal72a, Hau72, BtD70i]. Multiplicities [Mat73l]. multiplicity [Ben71f, Cha70]. Multiplier [Tre73d]. Multipliers [Bau71h, Tre73e, Sán71, Tur71]. multiplikative [Lor70m, BtD70i]. multiplying [O’M74]. multivariable [May72q]. Mumford [Del73n]. Müntz [Fer74a]. Müntz’s [GW74a]. Musili [Har70d]. mutual [OG74e, OG74f, OG74g, OG74o]. mutually [Bea70k]. My [Tur73b]. Mycielski [Chv74b].

N [Dev74, Pet70a]. nach [Pop70p, Sch71c, Sch71g, Sch71i]. Näherungsverfahren [Rau72]. naive [Sch71g]. Nakayama [Tac73b]. narrower [Luc73h, Luc73g, Luc73l]. National [Aku73, New74b]. natural [KM72]. naturality [ABK71c, Ati74f]. nature [Sma71a]. naturelles [Arn70]. Navier [Dye73, Jos73]. Near [Hur74, AC74, Bha70, Bra71, Iwa71a]. Nearly [MP71, Mil73b]. nearly-best-possible [Mil73b]. negatifs [NF74]. negation [Kun73]. negative [Hof72a, Azr72]. neighborhood [Mat73g]. Neighbourhood [GM71g, GM71h, GM71i]. Néron [GR72]. nets [Ray70t]. nettes [Ray70o, Ray70-30]. Network [BD72f]. networks [ADT72]. Neuchâtel [Sig71b]. Neumann [Beh72a, Par72l, Rin72b, Sai72a, Tak70n, Tak72, Tel71]. newsletter [Sta74a]. Neyman [Chi73]. Nice [BCdV^{+}70, Knu73a, Mat71b]. nicht [GS74a, Lor70j, Pop70h]. nicht-linearer [GS74a]. nichtlineare [AH70c, Kol72, Sac72, Wac72]. nichtlinearen [Mei72, NW72, Vel72]. nichtlinearer [Ade72, AT72, AT74, Gek72, Mic74]. Nikodým [AdK74a]. nilmanifolds [Par71a]. Nilpotent [BKW74, Mis74, Roi74, BK72n, BK72c, BK72j, BK72q, Hil74b, Hil74a, Sta73e, Ste70n, Lip74e]. nilpotente [Lor70n]. Nirenberg [Gri72, Mat71a]. niveau [Kat73b]. no [AHK74, CRW72, HL74a]. no-three-in-line [AHK74]. Noether [Del73k, Mar71e]. Noetherian [Gol72, Hei73b, Knu71h, Tac73a]. noethérien [Jen72j]. noise [Hau72, OG74a, OG74j, OG74s]. nombre [Dez74]. Nombres [Ber74c, Bru74b, Bru74e, Mey70f, Sch71w, Wal74d, FS70b, Mar71b]. noms [Tem71b]. Non [Ake74, Aus73a, Bac72, Bar71c, BGR73b, CGN71, Coh74r, CW71a, CD72b, Dav71, Háj73, Kno73b, Maa71k, Mar70d, Mey71f, Mil73a, Mor71d, Ray71c, ADS73, AdL72, BN73, Ber70a, Ber73, BL74a, Ber74h, BK72j, Bre71g, Dan74, Del73n, Din71c, DD71b, DP73, For74, Gan71c, Geb74, Glo70, Gri71b, HB70, Hel72, HMR71, Hof72a, Jon74, KM74, Kri70, Lat72, Leb73b, Mar70b, Mar71c, Mei72, Mil73b, Mon71h, Mos70, NF74, Rav71, Ray70q, Ray74, Sch72b, Ser73e, Ser73i, Tho73b, Tit74j]. Non-abelian [Ake74, Bar71c, Gri71b]. non-additive [BN73]. Non-analytic [Maa71k]. Non-axiomatizability [CD72b]. non-cancellation [HMR71]. non-classical [Ber70a]. Non-Commutative [CW71a, BGR73b]. Non-compact [Aus73a]. non-completely [Jon74]. non-coplanar [HB70, Mos70]. Non-Différentiables [CGN71]. Non-embeddable [Tit74j]. non-equilibrium [Leb73b]. non-equivalence [Dan74]. non-Euclidean [Hel72]. Non-linear [Dav71, Mei72, Tho73b]. Non-minimality [Háj73]. non-negatifs [NF74]. non-negative [Hof72a]. non-nilpotent [BK72j]. non-quasitriangular [DP73]. non-residues [Mon71h]. non-separable [Din71c]. Non-singular [Mor71d, Kri70]. Non-square [Mey71f]. Non-Standard [Bac72, BL74a, KM74]. non-symmetric [For74]. non-tangential [Bre71g]. Non-trivial [Mar70d, Gan71c]. non-triviality [ADS73]. Non-unique [Mil73a]. Non-vanishing [Coh74r]. Non-Well-Posed [Kno73b, Mil73b]. Noncommutative [Ser73e]. nonconforming [Cia74]. Nondegenerate [KP71]. nondifferentiable [CGN71]. nondistributive [Lac72]. Nonlinear [Col71, Edm74, FNSS73g, Hes74a, Lam74, Sat73g, SP73, AJ73, BD74, Bun72, CM74f, Coh73b, ESS74a, Fes74, FNSS73c, Gri73, Gup74, HO73, Har72b, Hed72, Hes74b, Hop71, Hsi73, Iwa71b, Lev71a, Per74, Rab71, Ran73a, Sac72, Spi72, Sty74, Vel72, vdC73, SJS73]. nonrandom [OG74j]. Nonselfadjoint [Kre73g]. nonsingular [Ade70, VCLN74]. Nonstandard [Bel74b, HL74c, HM74a, HM74b, Loe74a]. Nordhaus [CM71]. Nordsieck [Dan74]. norm [BS73a]. Normal [GM71g, KS71a, Cor74, Hau73, Hea74c, Mor71d, Rei71m]. Normale [Bad74a, Pop70i, Pop70e]. normaler [Pop70d]. normality [Boo74a, Rob73o]. Normalizers [CWW74]. Normalteilern [Bae74]. Norman [Whi70]. Normed [PT73i, Bea70e, PT73g, PT73h]. normés [Hir72c]. normierten [End74]. norms [Sha71d, Yam74o]. not-so-long [Coh70b]. Notation [Ner71f, Smo71k, GU71p]. Notations [Bor72b, KMT74g, Bea70l, Ray70-28, dlH72h]. Note [Chu74b, Kaz72c, Sim74, BR73b, Bur74a, Cov71, Gai72, Hor70a, Kaz73, Loe74b, Mey72d, Mey74g, Ram74a]. Notes [Hal73, Har70d, Rob73j, Tro73b, ST71g]. Notion [Bre71n, Ray70y, Bea71, Her74e]. Notions [Bor72c, Bre71o, HN71k, Bre71k, Bre71q, Col74e, Got73, GM71e, Hir71b]. Nouvelle [Sko74, AM74]. Nova [Fil73]. Novikov [Mau73]. Novosibirsk [Moi70]. Novosibirsk/USSR [Moi70]. noyau [Ben73, Zyg71b]. Noyaux [CD74, Del72i, Zyg71g]. Nucléaires [HN71d]. nuclear [Bol74, SP73, HN71d]. Nuffic [Kui71a]. Nul [Ray70c]. null [HR74a]. null-graph [HR74a]. Nullmengen [Sch71c]. Nullstellensatz [Maz71, Ros73, Sko73]. Nullteiler [Lor70n]. Number [Dob70d, Mon71p, Wei71l, Zag72a, Zim72c, Ber71i, But72, CS74, ER70, EHR73, FGQ71, Her71e, Kro72, Kub71, Mon71i, Mon71n, Oor74, Rog70e, Sal73, Sca71e, Ste74h, Sur72, Zag72b, Zim72a, Zim72b, Zim72n]. Number-fields [Wei71l]. number-theoretical [Kub71]. Numbers [PT73i, Sta73i, BS74b, Bom73, Car72a, Chu74a, Guy71, Guy72, Jec71e, Jec71p, Jec71q, LW72, PT73c, PT73b, PT73g, PT73h, Wal74d, Zim72b]. numeric [Kru73]. Numerical [Bet74b, CA74, Eva71, FPC74, Her72a, Her73, Mil74p, Mor71b, Mor71c, Wat74, Auc74, BR73a, BS74a, ESS74a, GP74, Gan71a, GL74, Hed72, How74, Jan74a, Moo74, Nob71, Sim73, Spi74a, Wac71a, Wie74]. Numérique [Rav71, Céa70, Glo70]. numériques [CR73b]. Numerische [LZ73, Mic74, NW72, AT72, AT73, AT74, Pit72e]. numerischen [Gen74, Her74a, Kre74c]. nur [Pop70i, Pop70e]. nutrient [FPC74]. Nyström [Bet74a].

Oberwolfach [AT72, AT73, AT74, Din71b]. Objekte [GU71n, GU71c]. observations [Ste70o]. Obstructions [Sto72d]. occasion [Bec73a]. October [Bet74b, McA74a]. Octobre [Nor74a]. Odd [Ady74, Bre72a, Cob73]. ODE’s [Mil73a]. off [Bea74]. offene [Gru74b, Pop70j]. Ohio [BRC74, HLM71, Pet70a, Suc70]. Oka [Nov73]. One [Bau74b, Bru74i, Bru71b, Coh72f, DK73b, Deu73c, Kui73a, KS73b, MM74g, Ogg73, Sch73j, vdH73, KP71, Mar73b, McA71, Nak73, Pet74e, Ric71, Rog70d, Str74a, Tak70h]. One-dimensional [Sch73j, Nak73]. One-parameter [Bru71b, Ric71, Tak70h]. One-Relator [Bau74b, Bru74i]. one-sided [Pet74e, Rog70d, Str74a]. One-step [vdH73]. one-to-one [McA71]. only [JW73g, Pop70e]. onto [Bel74a, Bin74]. Open [Sto72e, Beh74b, Har70a, Kin72, McA70, Ram74b, Sch74d, vdK73m]. Opening [Lau70]. Operaciones [Git72]. Operads [May72l, May72m, May72j, May72a]. opérant [Hir74a]. Opérateur [Cho73f, Leb73a]. Opérateurs [Che74a, HR74b, Kan74, Kar70a, Mal73, Zyg71e, Gou71, Mok70b, NF74, Rev70, Blo72, Gru73b, Kan72, Sch70c]. Operation [Coo74, Glo73]. operational [Ben70]. Operationen [BtD70k, BtD70l]. Opérationnelles [HN71f]. Operations [Sta70o, Bad70d, Bic73a, BtD70k, CR72c, Git72, Gow74, May70, Ste71a, Zab70]. Operator [Fil73, Hof72b, Pat74, Baz74, BdS74, Dwo73, Fif73, Gel74, GL74, Her73, Hof72c, Jör72, Kad70, Maa71b, Nas74, Neu70, Rin72a, Tak70g, Tak70h, Tak70j]. Operatoren [Rau72]. Operators [Ati74a, DeV72c, Eic73, FNSS73g, JW73g, JW73l, Tay74h, dlH72a, Ati74d, Ati74h, Ati74b, BP74, Ben71l, BdS74, Bre72b, Bro73b, BDF73, Car72a, Cob73, Cod71, Col74f, DeV72a, DeV72d, DeV72h, DeV72i, Dio74, Dol70c, DP73, Eas74, EG74, Fes74, FNSS73c, Gou72a, HH73, Hes74a, Hon74a, JW73k, JW73j, Mal73, Ner71i, Omo74k, OG74i, Ple72, Ram74a, Sac72, Sai72b, Spe71, Spe74, Tay74c, Tay74b, Tay74d, Tay74e, Tay74g, Tay74j, Ter70, Tho71d, Trè74, Tur73a, Ura71b, Wei71k, dJ72, dlH72g, dlH72b, dlH72c]. Optimal [BJ70, Eke70, Iha73, Kom72e, Kom72d, Lio73b, Lio73a, Lio73d, Mar73b, Neu70, dM70, Bal70, Ber70a, Che70, Hal70a, IDS70, Kno72, Kom72a, Kom72c, Mar70a, Pfe70, Yev70]. optimale [Sch71t]. Optimales [Mar70c]. optimalization [IP70]. optimaux [Mar70b]. Optimization [BCdV^{+}70, Ben72b, Hol72b, Moi70, Bal70, Erm70, Hol72f, OTF70]. Optimum [Ede70, Céa70, GK70]. Orbit [Zag72a, BGR73e, Ede70]. orbital [Moo74, Mos70, Pfe70]. orbites [Mar70b]. orbits [HB70, Ste74b, VCLN74]. orden [Git72]. Order [MS70d, MS73j, Bet74a, Bor71b, BS73c, Büc73a, But74a, But72, Car74c, Chi71c, CD72b, Cop71b, Cop71e, Fin74f, Fin74l, Fit72, FGQ71, Gil74f, Gil74e, Gil74g, Gil74h, Gil74i, Git72, Iwa71a, Jan74a, JLP73, Kre73d, Kre73e, Las70, LRV70b, Mei72, NS72, PS72a, Per74, Sat73g, SE70a, Tou74, Tro73b, Wal74k, Zas73, dJ72]. Ordered [Jam70d, Jam70e, Jam70f, Jam70g, Kei71, War70]. ordering [EHM70]. Orders [HMR73, RHD70c, Rog70c, SE70h, Swa70, Low73, RU73, RHD70d, Rog70d, Rog70e, Rog70f, SE70g, Zim72l, Rog70d]. Ordinal [Jec71o, Jec71p, Fef72]. Ordinals [MS73j, And73b, Ler73, Zuc73c]. Ordinary [Bet74b, ES72c, HS73b, Ura71c, Bea74, BP74, Ben71l, Bor71b, Coo71b, Hop71, How74, Hul74, Iwa71b, Lam74, O’M74, Ple72, Ste74j, Tau74, SM74b]. ordinates [AG74]. Ordnung [Gen74, Haa74, Mei72, Sch71h, GS74a, Sch71u]. ordonné [Jen72d]. Organized [Cur72]. orientables [Bru74o]. Orlicz [Tur73a]. Ornstein [Mar73c, Smo71i]. orthogonal [Ars72, HR71, vL74n]. Oscillation [Gro72a, Kre73h, Dia72, Kre73a, Kre73c]. oscillations [Hal70a, Kni74]. oscillator [Gol71a]. Osgood [Siu71b]. Other [MS73h, OG74o, Ste70m, Fin74c, Kno73a, Koh72, Mil73b, Sha71d, Sta70l]. Ouelques [Gir73a]. Outer [Sch73-30, Sio73c]. Outline [Sim72b]. ouvert [And70, CD74]. ouvertes [Ray70l]. Ouverts [Gro71b, Zyg71f]. Overdetermined [Sch71x]. overrings [Gil73].

p [Bru74b]. P. [BB70, Car71a]. Painlevé [Ree74]. pair [Zyg71b]. Paire [Ler71]. pairs [Bun71, Cov71, Cur70, Dav73, Ple72, Tit74n, Tit74f]. Paley [Ste70a]. Papakyriakopoulos [Poé74g]. paper [Bur90, Shu71a]. papers [Yat72a]. Parabolic [DD74a, WP74e, Auc73, Col74f, Fri74, Gor71, Mil73a, Tay74c, Wal72e, Wal74m, Zlá74]. Paraboliques [Dem70j]. parabolische [Kol72, Rau72]. Parabolischen [Gek73]. parabolischer [Gek72, Haa74, Sac74a]. Parallel [Sta74b]. parallels [Sch73k]. parameter [Bro74c, Bru71b, Fai74, Hop71, KO73b, Kot73, Ric71, Set72, Tak70h, Ter70]. Parameters [OTF70, GK70, Hoa71, Lor74, Sca71b, Sca71k, Wac71a]. Paramètre [Spi74b]. paramètres [Car71b]. parametric [Wed72]. Parametrization [May72q]. Parametrizing [Wel71a]. Paramodulation [WR70]. Parfaits [Ill71f]. Park [Bro71a, GL72, Hor70b, Hor71]. Parkside [GW73]. Part [Ste74d, Ste74f, Bea70a, JW73g, Jac72a]. parti [Ber74c, Ber74d, Mey74a]. Partial [Caz74b, Dwy74, ES72c, Kre73i, SM74b, But72, Ces74a, Coh73b, Col74c, CG71b, CG71a, ESS74a, EG74, Gil74d, Gil74l, Gob74, Gou72b, Kom73b, Kom72f, Lin74, Man71, Pay73, Tho71d, Trè74, Wid71b, Wil74a]. partially [War70]. particle [JW73a, JW73f, Kub73]. particles [Sch73e, Sch73b]. Particular [Jec71l, MS73i, Tre73f, CG71b, Tre73a]. partie [Bre72c, Bre72d]. partiellen [Haa74]. partieller [AT72]. Partielles [Lem70, Tem71a, Zyg71e]. Partir [Art70, Poe74n]. partition [BE74, CM72, Chi71b]. partitions [BN73, vL74e]. Parts [Bea70g, Bea70b, Bea70c, Bea70e, Bea70j, GW72]. passage [GV72a, Gus73a]. paths [BCM74, MK70]. Paved [Del71d]. pavés [Del71d]. Peano [Cur74]. Pénalisation [Lio70]. penalty [Kin74, Zie74].

1973 [Kel74i]. 68 [Tem71b]. 69 [Ano71a, BCC^{+}70]. 70 [Ano71b, Tem71b]. 71 [Ano72]. 72 [DE73, Rob73j]. 73 [Ano74]. England [MR73]. France [LLNS70]. Japan [Ura71c]. NJ [BCC^{+}70]. Scotland [ES72c, Kno73b, Mor71b, Mor71c]. USSR [Moi70]. Penrose [But74b]. Perfect [vL71h, vL73h, GH74c, Tac73l]. period [Wel71a]. Périodes [Aze70f]. Periodic [Ady74, Fin74a, Ves70, Bor74, Bun72, Car71e, CR72d, CRW72, Eva71, Fin74c, Fin74b, Fin74i, Fin74o, Jos73, Ken73, Lan73a, Lou73, Sat73c, Sei71, Wal73c, Wal74m, Yos71, FK72c]. Périodicité [Kar73b]. Periodicity [Ber71g, Sul71c]. périodiques [Man73b]. Periodische [FK72c, FK72b]. Periodischer [Gek73]. periods [Hur74, Aze70f]. permanence [Ray70z]. Permutation [Ker71c]. Permutations [vL74b]. Pernambuco [Nac74a]. Perron [Ber72b, Ber71f, Sch73-29]. Perturbation [Ber70b, Sib71b, Com71, Fif73, Sim73]. Perturbations [Jör72, JW73h, Lio73d, Noh71, Bré74c, Bro74a, Hab74, Hes74a, HS70, Stu72, dJ72]. perturbed [Bax74, Har73, Str71]. Peter [CW71g]. Petrowsky [Tay74c]. Pettis [Cas70]. Pfaff [Mar74b]. Pfaffian [Gib72]. Phase [How74]. Phenômes [Rou74]. phenomena [HMR71]. Philipp [Sch73o]. Physical [MS73e, SJS73, Mor74]. physics [Hol73]. Picard [Del73i, Ive70a, Ive70d, Kle71, Lip74a, Ray70g, RK71]. Piecewise [DD74a, Str74a, Wid71a]. Pinceaux [Kat73e, Kat73a]. pincées [Kui73b]. Pincherle [Pis71]. pipelines [Dup74a]. Pisot [Mey70f, Sch71w]. Pittenger [Mey73c]. Pittsburgh [AHiN74]. PL [Mil71]. place [Wei71i]. Planar [ZVC70b, Alb74, Lit74a]. Planarity [LR72]. Planck [Har72b]. plane [And74c, Bec70, Bec71a, Bec71b, Bec71c, Che74c, Sib71a, Sta70p, ZVC70a]. Planes [Ost70a, Pha71, Art71, KS74a, Ost70c, Rah74, Trá74a]. planet [IDS70]. planing [Del71d]. Plasmaphysik [vFD74]. plate [Cia74]. plates [Kom72a, Kom72e]. Pli [Jat70d, Jat70e]. Pli-domains [Jat70d]. pli-rings [Jat70e]. plongée [Tho74a]. Plongement [And74f, Mou72f]. Plongements [Ste72]. plumbing [Orl72c]. pluricanonical [Bom70]. plurisousharmoniques [Gru74a, Lel70a]. Plurisubharmonic [Her71a, Lel74c, Lel71b, Nov74a]. Plurisuperharmonicity [Her71e]. plus [Luc73h, Luc73g]. Plusieurs [Gla71a, Nor74a, Car71b, Nor74c, Zyg71g]. poids [Fer74d]. Poincaré [Ber74l, Fre71b]. Point [Mil74o, Sta70f, BCM74, Bon72a, Eym72e, Fan70, Fin74g, GP74, Gér74, GQ71, Gru70i, HK71i, KS73a, Kri73, Mey73h, Pat70, Pow70, RS74, Rho74, Sch72e, Sib71b, Sta70e, vdC73]. pointless [HR74a]. Points [Aus70b, Bru74c, Del70d, Rob73f, Rob73g, Art72b, Bel70, BCM74, Bom74, Chi71c, Coh74h, Cov71, Her73, Hof72a, Jaw70, KP71, Lew72a, Lor74, MK70, Sun70a, Tit71b, Was71, tD72b]. Pointwise [Moz71b, Boo74a]. Poisson [Aze70c, Aze70a, Aze70d, Bro74a, Car71c, Del74a, Mey71b, Pap70]. Poissonian [Gui72i]. polaires [Del73r, Lel70a]. Polar [Lel74c, Tit74i, Tit74j, Tit74h]. Polarisations [Riv72g]. polarization [Luc71g]. polonais [BI70b]. Polya [Mar71c]. Polycyclic [Ros73]. polygons [Hig74]. Polyhedra [Sto72f]. Polyhedral [Sze74, Knu73a]. Polynômes [FS70g, FS70a, FS70d, Lel71a, Ver74, Wal74i]. polynômial [Lel72a]. Polynomial [Ars72, DD74a, Gam71, Ste71a, Bra72a, BMRH73, Coh74g, Dan74, DG73, DeV72a, DeV72i, EE73, Fre74b, Hen71, Hir72b, MP73b, Ped73, Sch74f, Sic72, Str74a, Wid71a]. polynomials [Coh74r, Eis71a, FS70g, Gir74, Zim72f]. Polytechnic [DF74b]. ponctuels [Mey71b]. Pontrjagin [Hey70d, Hir71c, Zag72a]. Pontryagin [Ake74, HMS74e, Lem74b, Mar70a]. portés [Art72b]. Posed [Kno73b, Mil73b, Nas74, Pay73]. positifs [Lel72c, Lel73a, Mok70b]. Positive [Beb74, DeV72c, Gui72g, Gui72i, PS72c, PS72d, Bea70h, DeV72d, Gui72b, Gui72d, Iwa71a, Maa71p, Maa71m, PS72b]. positives [GC74]. possibility [Top70l]. possible [Mil73b, Wid71a]. Postgraduate [Rob73j]. Potential [Bau71a, Bre71p, Bre71o, Fug72b, Gro70a, Han71a, Hea74a, Her71c, KŌ73a, Sne71]. potentials [Sto70b, Sto70i]. Potentiel [Far74, Den73b]. Potentiels [Del70h, DMW73a, Hir74a, Mok70a, Mok71, Mok72, Rev70, Sko74]. pour [Art73h, Art73g, Art73i, Cai70, Con70, DC70, Fer74b, Fer70g, Fou74, Gro71a, Gru74a, Hei73a, Jen72j, Jen72c, Jen72h, Jen72g, Kat73d, Kle71, LR71, LM70, Lel71a, Lel74b, MM70, Mar71c, Ram74c, Sko74]. Poursuivi [PMN70]. power [Arn73, Gra73]. powers [Ber74n, Iwa71a, Lip73, Mac72, Mes72f]. practical [Sto74]. Präliminarien [Lor70f]. präsentierbare [GU71c, GU71l, GU71m, GU71n]. präsentierbaren [GU71q]. Präsentierungszahlen [GU71a]. Pre [Poé74e]. Pre-Groupes [Poé74e]. Précapacités [Lan74]. precapacities [Lan74]. precedent [Mey73b, Lel72a]. precise [Gan71h]. Précision [LR71, VCLN74]. predator [Pim73]. predator-prey [Pim73]. predicates [Sca71j]. predistributions [Loe74b]. Preface [AK70h, AS71k, Ner71g, Man73a]. Préfaisceaux [GV72b]. pregeometry [Rob74]. Préliminaires [Poé74f, KO74a, Wal74e]. Preliminaries [Ben71i, Coh70c, DI71d, DeV72f, DJ74i, FNSS73f, Fug72f, HMS74f, Hol72c, Jam70h, Jón72d, Ner71h, RHD70f, ST71h, Tac73g, Tak70i, TP71c, WWS72b, Was74b, Wil73b, PT73d, Sca71d, KO74a, LL74f]. Preliminary [ABK71e, Maa71l, KMT74g]. Premières [Ill71g]. Premiers [Art73c, Del72a]. preparation [Łoj71, Mat71a, Nir71, Poe74p, Wal71a, Gla71b]. Preparations [AS71l, BtD70n]. Prerequisites [Fel71e]. Préschéma [Ray70a]. Préschémas [Ber70d, Gab70a, Art73a]. prescribed [Gir74]. presentable [GU71m, Ulm71]. Présentation [Arn70, Mey73f]. Presentations [Gru70m, Lem74e, Hak74]. Presented [Bau74a]. Preservation [Coh74s, Top70n, Pop70p]. preserving [HK70, Kak73, Sai71c]. Presheaves [GV72b]. presque [Hei73a, Hei74a, Man73b, Mok70b]. pressure [vW73]. prevalence [Jac70]. prévisibles [Mey72c]. prey [Pim73]. primaires [Kah73, Maz72]. Primary [Hau73, ST71i, AK70a, Her73, Mat73m, Sto72j]. Prime [Arn73, BGR73c, Coh72j, Fay73d, Hir72b, Mon71i, Mon71n, BGR73d]. prime-form [Fay73d]. primen [Pop70g, Pop70f]. primes [Bac73]. Primideale [Lor70c, BGR73d]. Primitive [BGR73e, Omo74j, Luc73b, Wil70h, Wil70i]. Princeton [BCC^{+}70].

Princeton/NJ [BCC^{+}70]. Principal [Wil73e, Jat70c, KS72, Wil73d]. principales [Dol72]. Principe [DC70, Cha72]. principle [BD72d, Her74b, Kan73b, May72o, May72n, Sca71f, Sty74, Zim72j, dV70]. principles [Dev73c, Mon71o, Zie74]. Prinzip [Lor70d, Sch71e, Sch71n]. priori [Gob74, Sim72c]. priority [Lac73, Ler73]. prisms [Kle72]. pro [Gom70]. pro-analytiques [Gom70]. probabiliste [Mey71h]. probabilistic [Sne71]. Probabilités [CD72a, Che74e, DMW73c, DMW74, Mey70b, Cha72, DD71a, BH74a, BCM73, DE71]. probabilities [Che74e, PP74a, Smo71a]. Probability [BKW73, MP73a, OG74l, PS72d, Suc70, WW74e, ER70, Gal72b, Hey72, OG74q, OG74r, Sch71v, WW74d, WW74f]. Problem [DJ74k, Eic73, Sim72a, AHK74, AC74, App73, Ben72b, Ben71l, Bre71d, Bre72b, BE73, CGS71, Chu73, Cia74, Col73c, Din71c, Dun70b, EH73, Eva71, FPC74, GH74a, GH74b, Hab74, Hei74b, Her74c, Jan74a, Jan74b, Jec71s, Jos74, Kim71, LRV70c, LRV70d, Lev73, Mil74p, Osb71, Pat70, Qui73c, Sib71a, Sim74, Sim73, Sum74, SW74b, Tay74d, Top70h, Top70o, Top70p, Was71, Wei71g, Wie74]. Problème [And74g, Azr72, BS73b, PMN70, Bre74b, Cha74e, DDD71, DM74, Dez74, Dol70b, Fer70h, Glo70, Her74d, Mey70g, Mey73i, Meg71h]. Problèmes [Car72c, Car74b, Eke70, God72, Gro71b, Lio73e, Lio73f, Lio73g, Lio73h, Lio73i, Lio73d, Lit74c, Rav71, Bru74p, Mey74a, Zyg71f, Aut72, Lio71]. Problems [Bau74b, Ger73b, Her72b, Kno73b, Kuj71b, Lan70, LRV70e, Sma71a, SJS73, Tak71c, Zim72c, Abi74b, AJ73, AH70e, AH70g, BR73a, BD74, Bas73d, Bax74, Ber74b, Ber70a, Bra72b, Bro74c, Caz74b, Ces71, Cod71, Col71, DS73, Dun70a, ER70, ES72a, Erd74, Fic72, GP74, Gan71a, Gil74a, Gil74h, Gil74j, GL74, Gli70, Gol71b, Gou74a, GH73, Har73, Hes74b, Hir72d, Hol73, Hol72a, Hop71, HK71i, IW73, IP70, Jam72b, Kar73c, Kau72, Kaz71, Kno72, Kno73a, Kur70a, Las70, LRV70a, LRV70b, LRV70f, Lel74c, Lio73g, Lio73d, Mat72, Mei72, Mil73b, Net74, Neu70, Nic73, Nob71, Oku71, Osb71, Ōuc73, Pay73, Per74, Ram74b, Rox71, Rue70, Sak71, Sat73f, Sat73g, Sch74d, Sha73b, Sha71a, Sha71b, Sha71d, Spi74a, SP73, Sta71b, Sto74, Stu74, Sug71, Tau74, Tay74c]. problems [Tho73a, Tho74c, Tho73b, Väi71d, Vel72, Wac71a, Wal74m, Wen72, Wen74b, Wid71b, You74, vL74i, vdC73]. procedure [Grz70, MS70b, Pra70]. Procedures [Hoa71, Kre70a]. Proceedings [ALW72, BH74b, Bas73a, Bas73b, Bas73c, Bec73a, CFK71, Cur72, DK73b, Fle70, GG72, HLM71, Hol74b, Hor70b, Hor71, HS71, Kel74i, Kom73c, Kui73a, KS73b, Kui71a, Nac74a, Pet70a, SM74b, SJS73, Suc70, Bet74b, Chi71a, Fil73, HS74b, HMR73, KMSS72a, KMSS72b, MP73a, New74b, Wal71c, Wal71d]. process [Ben71k, Csö73, Gus73a, Pap70]. Processes [Gui72i, Sil74e, Bor73, Con72, Doo71, Dun72, Fel70, Fuk73, Gri73, Kan73a, KK73, Kaw73b, KO73b, Kot73, Kri73, Leb73b, Luc73j, Mar73c, Nag72, OG74i, OG74f, OG74j, OG74q, OG74s, Ran73a, SV71a, SV71b]. Processus [Bad70f, Bad74f, Bre72d, BCM73, Bre73, Fau70, KR73, KF73, Mey71b, Mey73f, Mey74f, Pri74, Spi74b, Ben73, Bre71t, Cai71, Del70f, Del70i, Del73o, Del74a, Dup74b, Geb74, Hei73a, Hei74a, Mey73b, Mey73c, Mey73g, Wei70, dM72]. proches [Mar70b]. producible [GU71k]. Product [AS71m, BC74b, Aur74, BS74b, Bre72b, CLL73, Cor74, Mac70, Pec74, Ste74h, Tak70o, Wil73d]. Products [Baj73, Gui72i, PS72d, Wil73e, Zag72a, Zim72m, Bre72a, Cas73c, Gui72c, Hir72b, Hou72a, Ive70e, Ker71d, Ker71e, May72p, PS72c, Sta70n, Wal73a, Zag72d]. produit [Céa70]. Produits [Hou72a, Sch72f, Eym72f, God71c, Kos71]. Produkten [Pop70m]. Professor [Bec73a]. profile [Bal70]. Profinis [Ser73b, Ser73j, Ver73]. Program [Can74, Coo71a, Elg71, MW71]. Programmation [Aum70]. Programme [Gro71a, Kre70a]. programming [Erm70, Gol71d]. Programmkomplexität [Sch71i]. programs [Coo71a, Eng71a, Lon71]. progressions [Ben71e]. progressivement [Del74b]. Projectif [Ber71n, And74f]. projectifs [Jen72j, Ray70q, Sad74]. projection [KS70, ST71m]. Projective [AS74e, Che74c, KS74a, Lor74, Mil74e, Pet72, Sta70p, Tac73i, Eic71e, Tit74i]. projectives [Bad70b]. projectivité [Ray70-27]. projektive [Pop70i, Pop70h]. projektiven [Pop70g, Pop70f]. proliferation [Mon71r]. Prolongations [Háj70]. Prolongement [Car74b, Coe70, Coe74, Dou71, Sib74, Nov70]. Prolongements [Del71f]. Proof [ABK71f, Bea70i, Eng70, JW73i, Luc73e, Sca71g, Sch73-31, Chu71b, Din71a, Kre70a, Luc73c, Mat71a, Moz71c, Moz71d, Moz71e, Moz71f, Nir71, Pra70, Sca71h, Sca71i, Tit71b]. proofs [Sne71]. Propagation [Tay74f]. Proper [AS71n, Bro74b, Mau73, AS71p, AS71q, AS71u, Bar72]. Properties [Bor70, CM74a, DG70b, Fei70, Fle70, Gra74i, Gra74j, JW73l, Kat73g, Sto70i, Yam74d, Ars72, Ati74e, Aul74, Azu73, Bea70d, BW71, BD72b, BW74, Bre71c, Bro70, CM74e, CM74f, Coh74a, Coh74q, DeV72d, Dye73, Fef72, Gil74l, Gre74, Kee70, Moo70, Mor72, Myh73, Pit72d, Pit72f, Rei71q, Rob71b, Seh73, Smi74, Ste71c, Top70f, Tou74, Väi71a, Wal73c, WP74a, Yam74m]. Property [Bar74c, CLL73, Chv74a, DJ74c, Fes74, Fin72, Fuc72, HL74a, Hav74, KO73b, Kot73, Nag72, Rei71k, dB71b]. Propos [Man73a, Ree74]. propositional [Lam72]. Propositions [Sal74]. propre [Art73h, Art73g, Art73i, Maz74]. Propres [JG71, Del73b]. Propriété [GC74, Lem74d, Ray71c, DC74]. Propriétés [Bou74a, DG70b, Ill71g, Eym72b, Eym72e, Lob72, Mok70b, Mou72c, Mou72b, FS70f, Gro72c, Maz72, Ray70z, Ray72b, Wal74f]. prospects [Wan70]. Protorsion [Lam71d]. providing [Top70l]. Proving [JL71, Lon71, RW70, Wan70]. proximity [Sto70c, Sto70d]. prozesses [Ros71]. Prüfer [Gil73]. Prüfer-like [Gil73]. Pseudo [Ben71j, Hat73, HN71n, Kom73c, Mok72, Rav71, Tay74g, Tay74h, Bel74a, Cer71, Kom73a, Kom73f, MY72, Nov74b, Poe71b, Tay74e, Wag71]. pseudo-arc [Bel74a]. pseudo-convexe [Nov74b]. Pseudo-Convexes [HN71n]. Pseudo-Differential [Kom73c, Kom73a, Kom73f]. pseudo-free [MY72]. pseudo-isotopie [Poe71b]. pseudo-isotopies [Wag71]. Pseudo-isotopy [Hat73, Cer71]. Pseudo-measures [Ben71j]. Pseudo-quotient [Mok72]. Pseudoaffine [Wil70g]. Pseudoconcave [AS71o]. pseudoconcaves [And74f]. pseudoconvex [MR71]. pseudodifferential [Trè74]. pseudogroupe [Kum74]. pseudoperiodic [Ura71b]. puissance [Jen72i]. puissances [Ber74n]. Punkt [GS74a]. Punkten [Pop70f]. Punktierten [Pop70f]. Punktmengen [NW74]. Purely [MV70d]. Puritz [Mat72]. Purpose [Can74, WW74g]. pursuit [Sak71]. Putnam [Azr72].

QF [Han73, Tac73e, Tac73a, Tac73k, Tac73h, Tac73i, Tac73j, Tac73m]. QF-1 [Tac73k]. QF-13 [Tac73h]. QF-3 [Tac73e, Tac73a, Tac73k, Tac73i, Tac73j, Tac73m]. Quadratic [Sta73i, Maa71h, Maa71m, Maa71u, Shi71b, SM74a, Lor70h]. quadratische [Lor70d, Lor70g, Lor70h, Lor70i, Lor70j, Lor70k]. Quadratsummen [Lor70l]. quadrature [Sha71g]. Quadrics [Del73m]. quadrilaterals [Gai74]. Quadriques [Del73m]. quantique [Car72c, Car74b]. quantitative [BL70]. Quantization [Kos70c]. Quantum [Kel74a, Gli70, Hol73, Oli71, Seg70]. Quasi [AS71p, AS71q, Bre71q, Gra74k, Tac73k, Wei70, Cou73, Eym72c, GH74a, GH74c, HR71, Itô72, Jos73, Mar70b, May72e, Mil73b, Mit74b, Ray70x, Sch71q, Ray70-27]. Quasi-adjointness [Gra74k]. quasi-analyticité [Cou73]. quasi-circulaires [Mar70b]. quasi-fibrations [HR71, May72e]. quasi-finis [Ray70x]. Quasi-Frobenius [Tac73k]. quasi-linear [GH74a]. quasi-linearen [Mit74b]. quasi-martingales [Itô72]. quasi-perfect [GH74c]. quasi-periodic [Jos73]. Quasi-processus [Wei70]. Quasi-projectivité [Ray70-27]. Quasi-proper [AS71p, AS71q]. quasi-regulières [Eym72c]. quasi-rekursive [Sch71q]. quasi-reversibility [Mil73b]. Quasi-topological [Bre71q]. Quasicoherent [Knu71h]. Quasicomplemented [Hus74a]. Quasiconformal [Ear74, Väi71c, Väi71f, And74c, Caz74b, Sch73a, Väi71a]. quasiconformality [Caz74a]. quasigroups [Lin74]. quasilinear [Fif73, AH70e]. Quasilineare [AH70e]. Quasilinearer [Tör73, vF73, Gen74, Kre74c]. Quasisimple [Koh72]. quasitriangular [DP73]. Quaternion [JL70d, EL73]. quelconque [Lut74]. Quelques [Aum70, Ber71l, Bor72c, Bre74a, Che74g, Del71f, Jou73a, Lob72, Mey70a, Mey74a, Mok70b, Ray70-28, Zyg71f, CW71f]. question [Gol74c, KMT74i]. Questionnaires [Bou74a, Cés74b, PP74b, Pet74a]. Questions [Fer72, Bau71d, Bel71b, GV72a, Pet74a, Ray70l, Sto72e, Top70o, Tou74]. qui [Gat73a]. Quillen [Blo73b, Ill73, Laz72]. Quotient [Gab70a, Mok72, Rei71o]. Quotienten [Sch74c]. Quotients [Dem70g, Lam71f, Ste74f, Ste71e, Kan74, Lam71e, Ste71b, Ste71c, Ste71d, War73]. qu’un [Fou74].

R [DDDM70, HK71h, Bab73]. R. [Bur74a, Fer74b, Kri71, Ver71]. rabotages [Del71d]. Rademacher [Bre72c]. radiation [Hei74b]. radical [Ste70e]. radicals [Ste70m]. Radicielles [Dem70b]. radii [Hof72a]. Radon [AdK74a, Bad70b, Bad70d, Bad74e, Bad74f, Che74e]. radoniens [BI70b]. radonifiantes [Bad70a, Sch72f]. Rahmenaussagen [Wil70h]. Raides [Lio73g, Lio73h]. Ramanujan [And72]. Ramified [Fay73e, GM71d, GM71k, GM71j]. Ramis [Car71a]. Ramsey [Bur74e, Chu74a, CS74, Har72a, Har74c, vL74h]. Randkomponenten [Gru74b]. Random [Dud70, Erm70, PT73e, TP71d, TP71f, Bor73, Csö73, HK71i, Jam72a, Kru73, MM74b, OG74f, OG74q, PT73c, TP71a, TP71e, Wed72, Del69]. Randomness [Sch71v]. Randwertaufgaben [GS74a, Ade72, Vel72]. Randwertprobleme [Sac74a, Kol72]. Randwertproblemen [Gek73, Mei72, Mit74b]. Randwerttechnik [Gek72]. Rang [Dem70f, Ray70c, Cés74b]. range [GW72, Mil74m, Wan70]. Rank [Bur74d, AS71r, Bur90, HR70, Roi71, Tit74n, dlH72c]. Rapidly [Gan71f]. Rappels [Ray70-29, Sch70d]. rapport [Del74a, DDM70, DD71a, Mey74h, Mok72, Ver74]. Rate [OG74p, Kow73, OG74g]. rates [Han70, Str71]. ratio [Kli70, Kri70, Kuo71]. rational [Bre72a, Fic74, Har70e, Hir71c, Ive70b, Lem71, PBR71, Roi74, Wac71b, Zag72c]. rationnalité [Ber74l]. Rationnelle [Art70]. rationnelles [Del73n]. ratios [Fra74]. Räume [Gra74c, Sch74c, Sei72d]. Räumen [End74, GR70]. Ray [MW72, Mey74g]. Razumikhin [Kat71]. reactor [SP73]. Real [AG71c, Jec71q, Kaw73a, Kub71, Sat73h, Shi71b, Väi71b, Yam74h, Yam74n, rd74d]. Real-algebraic [AG71c]. Realizability [Kle73a, Sta73h]. realization [May72f, May72g, Sto74]. Realizations [Kan71g]. reals [McA71]. réarrangement [Car70b]. récents [Den73b, Mey72b, Ray71a]. Recessive [Bra72a]. Recherche [Aku73, Aus70b, Céa70]. recherches [Bre74a]. rechter [Tau74]. Recife [Nac74a]. recognition [May72o, May72n]. récollement [Poe74q]. reconstructible [Chi71c]. Reconstructing [Bry74c]. Reconstruction [Rad74, Fab74, Har74d]. reconstructions [Woo74a, Woo74b, Woo74c, Woo74d]. reconstuctions [Woo74e]. recurrence [Bha70]. recurrent [KŌ73a, Rev70]. recursions [vL74d]. recursive [Hir74b, Luc73o, Luc73a, Luc73b, Sca71b, Sca71k]. recursively [Lac72, Lac73]. Reduced [WP74a, Bre71n]. reducible [Ves70]. Reductif [Gro70g]. Reductifs [Dem70a, Dem70e, Dem70h, Dem70j, Dem70g, Dem70f, DG70c]. Reduction [PŠ70, Tit74p, Wil73e, KKMSD73b, Kni71, Maa71m, Pra70, Glo70]. Reductive [HC70d, Ste74e, Van72]. Réduites [MT73]. réel [Mey70c]. reell [Lor70g]. reellen [Lor70d, Lor70j, Lor70k]. reeller [Sei72e]. reference [Jam72a]. Refinement [Luc70, DD74c]. refinements [Fro74]. Reflections [Wal71e, cHyH70]. reflective [Hon74a]. reflexivity [HM74b]. regard [Tou74]. régénératifs [MM70, Mai71]. Regions [Wed72, BD74, Mon71r]. Regular [Gri71c, Bar71a, Bar72, Bli71a, Del70d, HC70a, Hus74b, Jon74, Mei74, Sch74e, Ste74e, Van71]. reguläre [GU71f, Pop70i, Pop70h]. regulären [Pop70g]. regularité [Cho73b, Aut72, Bom71]. Regularity [CT74, Gou72a, Sim72d, Jec71b]. regularized [Nas74]. régulier [BI70c, Buc73b, Ray70j]. Regulières [Ber71m, Del73f, Eym72c, Del70b]. Réguliers [Del70d, Gro70h, GS70]. Reihen [Pit72g]. Reiter [Eym72b]. rekursive [Sch71q]. Related [BCC^{+}70, Gui72i, Hil74c, HMR73, Kre73j, WW74e, Bel71b, Ber70a, CHS72, Caz74b, Hir72d, IW73, Kee70, Osb71, Pet72, Sai72b, Top70o]. relatifs [CR73b]. relating [Ran73a]. Relation [Los74, BK72p, But74b, Hof70n, Mes72h, Rob73o]. Relational [Bar70, Jón72b, Pul70]. Relationen [Pop70e]. Relations [AKW73, Azr72, Bir73, Bre71r, But74c, Mei74, vdK73m, Bec74a, Ben72a, Ber71b, BF74, But71, Jón72c, Kui71b, Pop70e, Tit74b, vdK73o]. Relationship [And73a, Pat70]. Relationships [OG74q]. Relative [Dev73q, Ill71e, Kom73e, RK71, Sca71h, Gus73b, Mok70a, Shu71a, ST71i, Sma71a]. Relatives [Ill71a]. relativiste [Geb74]. relativistic [Oli71]. Relativity [Sch73f]. Relativization [Rei71q]. Relator [Bau74b, Bru74i, Sac73]. relevant [Fuc74]. Relèvement [Boc72]. Relèvements [Car71d]. Remainder [Gan71g, Gan71i, Gan71j]. remainders [Gan71f]. Remark [Wal71f, Kor70, Omo74k]. Remarks [BŁ71b, Cho70, Gre73b, Jec71r, Jos73, Kan73a, Kan71h, Ker73, Lau70, May72o, Roy71, Ben71b, DI71a, Gan71e, Hea74c, Kom72b, Maa70, Maa71l, Par72j, Pay73, Sch73d, Sim73, Tho71c]. remarquables [Mok70b]. remarque [Azé72, Bre74b]. Remarques [Lel72a, Lel74b, Lio70, Nor74d, KO74g, Mey73h]. Removal [CGS71]. remplissage [MR72]. Rencontre [Bru71c]. Rencontres [dFP74]. rendez [Mar70b]. rendez-vous [Mar70b]. rendezvous [HB70]. rendu [Dez74]. Répartition [Mey71e, Mey70e]. Report [Din71b, Lel74a, Maa71n]. Reports [Gra71a, MAB^{+}70]. Representabilité [Gro70d, RK71]. representability [Ing71]. representable [Sch71j]. representatif [Ber74d]. Representation [Fes74, Har74h, Ive70e, Knu73g, Pet70c, Fau70, Fou71, Bre71f, Coh74t, Dol70c, Kei71, Ker71a, Knu73f, Loe74a, Maa71g, Rei71f, Sim72e, Tac73j, Tak70n, AM74, Mar71c, Bru74k, Bru74m, Mey71c]. Representations [But74d, CL74a, Cas73a, Cas73b, Ker71c, Ker71d, Lan73b, Lip74h, Sil70a, Ste70n, SD73, Wil73c, Wil73e, Del73e, Eym72f, Eym72g, Gab71, Riv72h, AS71m, Bor70, CHS72, Cur70, Dre73, Gil74g, HC70a, HC70g, HC70k, Hir72a, iI71, Kos70c, Lip74c, MS73c, Sil70c, Sil70f, Wil73b, Wil73d, Wol71, Del71a, Dix74, Eym72c, Nor74c, Lip74d, Lip74f, Bor72d, CW71g, CW71f]. Representatives [vdK73n, vL74g]. Representing [Bea70j, Bea70k]. Research [Bas73a, Bas73b, Bas73c, OG74b]. Reserve [HLM71]. résidu [Bea71]. residually [AG73a]. Residue [BGR73f]. Residues [Her70, Dol70a, Mon71h]. résidus [Dol71a, Dol71b, Dol72, Kan71k, Nor71]. résolubles [Gab71]. Resolution [Glo70, Orl72i, Rav71, Hir72d, Lov70, Luc70, OW72, Poe74n, Gri72]. resolutions [Gru70m, Mil74i, Ill71b]. resolvable [BGR73d]. resolvante [Mey74h]. résolvantes [MW72]. resolvent [Tak70j]. respect [Fre74b, Her71e, Sca71b, Sca71h, Sca71j, Sca71k]. response [Pim73]. restricted [Gia72, Qui73c]. restrictions [Neu70]. result [Di 71, Lap72b, Top70a, Top70b]. résultat [MW72]. Résultats [Aum70, Bor72c, Bor72a, Bre71t, Gir73a, Mey73g, Art73c, Del70g, Dez74, Jen72d, Mey71c, Mey72b, Jen72h, Sch70e]. Results [Cas73a, SE70i, Zim72c, Bar71c, Bre71l, CD72b, CG71b, CG71a, Dos74, Dud70, Gil74e, Gol74a, Guy71, Han70, Har72a, Har74a, Har74b, Hod74, How71, Inf71, Kai74, Kom72f, Kub71, Lel71b, Mon71e, Mur74, Nad70, OG74s, Ram74b, Rei71p, Sch74d, Sch73r, Sta72, Wid71a, Lip74c]. Résumé [Tho74a, Del72a]. Réticentes [Mar70c]. retirement [Bec73a]. Retour [Mey70g, Azé72]. retournement [Mey74e]. retracts [Jaw70]. Retrake [GU71o]. retrospective [Kle73a]. réunion [Sig71b]. Rev{êtements [GR71]. reversibility [Mil73b]. Review [KS74b, Omo74k, Sca71i]. Revisited [Sta71a, Adl73, Boa72, Har73]. Rham [Ber74j, Blo70, Ill72b, Tel71]. Riccati [Bro70, Sán71]. Riemann [Abi74a, BGI71, Ber71o, Deu73d, Fay73a, Fay73c, Fay73f, Fay73g, Gun71, Kim71, LB70h, Lie74]. Riemannian [BGM71c, FK72c, Gre73a, Maa71o, Maa71p, Maa71t]. riemannien [MM74d]. Riemannienne [BGM71b, Far73, BGM71c]. Riemanniennes [BGM71d, Kui73b, LB70j]. Riemanniens [BGM71a]. Riemannsche [FK72a, FK72c]. Riemannschen [FK72b, FK72d, Gru74b, ZVC70g]. Riesz [Bic73f, Jam70b]. right [Pul70, Tau74]. right-hand [Tau74]. Rigid [DJ74j]. Ring [Coh74f, Gru73a, Hof72b, KT71, BT73, Coh72j, Coh74a, Fai72, Gai74, GM71b, HK71a, Hir74b, Hof72c, Koh72, Mat73g, Mat73p, Pie72, War73, Lor70n]. Ringen [LL74b]. Rings [DI71e, Dob70a, Hof72c, HMR73, Knu73c, Lam71f, Mat73a, Pet74c, Ste71e, Tac73k, Vas73, Arn73, Bar72, BMRH73, CLL73, DG73, DR72, Dob70e, Dob70c, EH73, Eic71c, EE73, Ger73a, Ger73c, Gol72, Gus73b, Han73, Hir72b, Jat70a, Jat70b, Jat70c, Jat70e, Kap73, Kei71, Lam71e, Lip74a, Mac73, Mat73h, Mil71, Moo71, MP73b, Pet74b, Pet74d, Pet74e, PŠ70, Qui73a, Ray70e, RHD70f, Ros73, Ste71c, Ste71d, Ste71f, Tac73e, Tac73a, Tac73h, Tac73i, Tac73j, Tac73l, Tac73m, War73, HK71a, Knu73g]. Rise [Coo74]. Robinson [Azr72, Loe74b]. Roch [BGI71, Ber71o, Deu73d]. rocket [IDS70, MK70]. Rogers [And72]. role [Gal72b, cH73]. Ronde [Aku73]. Room [WWS72a, Wal72d]. rootless [KP72]. roots [Ste70q]. rotating [Auc74]. Rotation [And74c]. rotations [Art71]. Roth [Mon71j]. RUCA [DK73b, Kui73a, KS73b]. Ruelle [Lan73a]. Runde [Lor70m]. Runge [Bet74a]. Russell [KMT74d, KMT74e, Mos72].

S [DDDM70, Gor73]. S. [DDDM70, God71b]. S4 [Fit72]. sûre [Hei73a]. Saarbrücken [KGLZ70]. Saint [BH74a]. Saint-Flour [BH74a]. Salem [Mey70f]. Salomon [Lan70]. same [Lor74]. sample [Csö73]. San [BK74]. s’annulant [Sko72]. Sans [Lio73c, Ben73]. Santiago [Vid74]. Sard [FNSS73e]. satellite [VCLN74]. satisfied [Coh74k, Coh74j, Coh74i]. Sato [Sch74b]. Saturation [DeV72g, DeV72h]. Satz [Lor70e, Lor70b]. Sätze [Pop70a, Pop70n]. sauts [Del70f]. Schanuel [Ami72]. Schéma [Gro70e, Ber71m, Art73b, Ber74m, Gro72e, MR72, Ray70j, Ray70k, Ray70-33]. Schémas [AGV72, Art72a, Art73j, Ber74i, DR73, DG70a, DG70b, DG70c, Gab70b, Gab70c, Gro70d, Gro70g, Gro71a, JG71, LR71, Art73i, Gro72c, Ill72f, Jou73a, Ray70i, Ray70q, Ray70r, Ray70s, Ray70-27, Riv72h]. Schemata [Pop70d]. scheme [Ive70a, KMT74f, KMT74j, Tho74b, GM71g, Ive70d]. Schemes [DG70b, Dem72e, Kat73g, Mes72c, Coo71a, Dem72b, Elg71, GW74b, Gor71, GM71e, GM71k, GM71j, Ill72e, KMT74b, KMT74c, Knu71d, Lip74a, Mor74]. Schliessungsaussagen [Wil70i]. Schluss [Bae74]. Schmidt [Bom73, Sch73o]. Schmitt [HK71d]. Schneider [Wal74c]. Schnirelmann [FNSS73a, FNSS73d]. Schnitte [Pop70i, Pop70e]. School [Chi71a, DK73b, Kui73a, KS73b, Kui71a, MR73, Wae73]. Schrödinger [Eas74, JW73j]. Schur [Yam73, Yam74d, Yam74e, Yam74g, Yam74f, Yam74j, Yam74i, Yam74h, Yam74k]. Schwach [AG73b, Gek72]. Schwartz [HC70i, HN71c]. Schwarz [AS71s]. Schwinger [Tak70e]. Sciences [MS73e, SJS73]. Scientifique [Aku73]. Scotia [Fil73]. Scotland [SM74b]. search [Che70, Kre70a]. Seattle [Bas73a, Bas73b, Bas73c, SJS73]. Second [AHiN74, Cop71e, Fin74l, Hol74b, KMSS72a, KMSS72b, MP73a, MS70d, MS73j, New74b, Ste74f, Top70p, Bor71b, BS73c, Büc73a, Car74c, JLP73, Las70, LRV70b, Sat73g, Sca71i, Tro73b]. second-order [Bor71b, LRV70b]. Sections [AS71t, Ive70e, Kei71, McC71]. Segal [Bur74c, Gol74a, Rei71a, Rei71e, Rei71i, Rei71h, Rei71l, Rei71n, Rei71m, Rei71o, Rei71p, Rei71r, Rei72]. segments [Yat72b]. Seidel [Di 71]. Seifert [CR72b, Orl72j, Ros74]. seite [Tau74]. Selberg [Maa70, Maa71q, Mon71n]. selbstadjungierten [Mei72]. Selected [Hol72d, Yor71]. Selection [Bea70k, Par72i, Bea70d, Par72j, Pat70, Sun70a]. Selections [Fle70, EHM70, McA70, Mic70, Par72b, Par72d, Par72c, Par72g, Par72k, You70]. selectors [Par72e]. Self [Ste71f, Tak70k, Wil74d, Bro73b, Eas74, Mei72, Mil73a]. Self-adjoint [Tak70k, Mei72, Mil73a]. self-adjointness [Eas74]. Self-injective [Ste71f]. Self-maps [Wil74d]. Selfadjoint [BP74, JW73k, Ple72]. Selle [Aus70b]. selon [Fra70a, Gri72]. Semantics [Eng71b, Iga71, Lam71f, Knu71a, Lam71a]. Sémantique [Ler71]. Semi [AS71u, Aze70b, DMW73b, Dem70f, HM74b, Hus74b, KKMSD73b, Mat73n, Tac73l, Tak70l, Baj73, Ber74a, Bor71a, Che74a, CL74b, Col70, Con70, Del73r, Dix74, Dol71a, Eym72f, Hir74b, iI71, Kaw72b, Kur70a, Mar71b, Ric71, Tho73a, Tho74b, Tit74d, Wan72, Wil73d, Yam74o]. semi-affine [Con70]. semi-analytique [Col70]. semi-continuous [Kur70a]. semi-directs [Eym72f]. semi-discrete [Tho73a]. semi-dynamical [Baj73]. Semi-finiteness [Tak70l]. semi-flows [Ric71]. semi-free [Kaw72b, Wan72]. semi-groupes [Ber74a, Che74a]. semi-groups [Ric71]. semi-kernel [CL74b]. semi-linear [Tho74b]. Semi-Markovian [DMW73b]. semi-méromorphes [Dol71a]. semi-norms [Yam74o]. Semi-perfect [Tac73l]. semi-polaires [Del73r]. Semi-proper [AS71u]. Semi-reflexivity [HM74b]. Semi-regular [Hus74b]. semi-ring [Hir74b]. Semi-Simple [Dem70f, Mat73n, iI71, Tit74d, Wil73d]. Semi-Simples [Aze70b, Bor71a, Dix74, Mar71b]. Semi-stable [KKMSD73b]. Semifree [Sch72e, Wan74]. Semigroup [Sio73c, Sio73d]. Semigroups [Bel71a, Hof70l, Pet74c, Pet74d, CM74f, Hof70k, Hof70h, Hof70j, Ter70]. Semilattices [HMS74e, HMS74b, HMS74c, HMS74d]. semilinear [Gil74j]. Séminaire [Art72a, Art73j, GR71, Nor74a, Tem71b, Ano71a, Ano71b, Ano72, Lel73b, Ano74, Bad70g, CD72a, DMW73c, DMW74, DE71, DE73, Fer70e, Hou72b, KGLZ70, Lel70b, Lel71c, Lel72b, Lel74d, Mey70b]. Seminar [Bau71a, BRC74, Gra71b, Gra71a, Kel74i, MAB^{+}70, Ura71c, Yor70, vL74a, Hou72b, BCC^{+}70, Fer70e, Yor70]. semipolar [Tul70a]. Semiprime [Pet74e]. Semisimple [Wil73e, HC70i, Hir72a, Lip74b, Ste74e]. sens [Mar70a]. sense [Luc73h, Luc73g, Luc73l, Ōuc73]. sensor [Ben72b]. Separable [DI71e, DI71b, Din71c, Jon74, PT73c, PT73e, RHD70e, You70]. separables [KO74c]. separate [Her71e]. Separated [Fin74m, Knu71h, Phi73]. Séparation [HN71l, AC74, Del71g]. September [Bas73a, Bas73b, Bas73c, Chi71a, Mor71c, Ura71c]. septembre [Aku73]. Sequence [LT96, Smi70c, Bab73, DI71f, Mil74d, Mil74k, Mil74n, RU73, Smi70b, Sta70c]. Sequences [Fra70b, Cap73b, Coh70b, Coh70e, De 73, Din74b, Ger73c, Jen72g, Mar73a, May72b, Moo71, vL74f]. Sequential [GH74c, LKR73]. Série [Bor72a]. Series [Maa71r, Moz71b, Wei71e, Wil73e, Arn73, Bro74a, Fin74h, Fre71a, Gra73, GK72, HC70b, Hel73, Kau72, KS72, Maa71c, Maa71k, Maa71n, MS73d, Pit72g, Rub73, Sta73f, Wei71d, Wil73d, CR73b, Man73b, Bre72c]. ses [Cai70]. sesquilinear [Stu72]. session [BK74]. Set [Jaw70, Jec71l, War70, Alb74, Bel70, Bha70, Dev73u, EHR73, Fra70b, Fri73a, Gil73, GM71b, Hal70b, HC70a, Kri70, Kur70a, Myh73, Ray71d, oT72, Tim71, WR70, Fel71d, Fle70]. Set-valued [Jaw70, War70, Fra70b, Hal70b, Kur70a, Fle70]. Sets [Bau71b, Ben71c, Str72, WWS72a, Yam74s, And74d, Aus70a, Aus71, BS71, BMN74, Bau71c, Bau71e, Beh74a, Ben71a, Ben71d, Ben71e, Ben71f, Ben71g, Ben71j, Bla74, Bli71b, BK72r, De 73, Del69, Dev73m, Dio74, Dru72, Eis71b, Fel71b, Gam71, Jec71g, Jec71h, Kau72, Ken73, Lac73, Lel74c, MM74b, Mar70d, McG72, PP74a, Par72h, Sch74e, Sch72e, SW74a, Sun70a, Tay74f, Tho71b, Tul70a, Wal71g, Wu71, vL74l]. seven [MY72, Was74c]. Several [Bro71a, AJ73, Gan71h, Kor70, Hor70b, Hor71]. SGA [GR71]. SH [DJ74m, DJ74h]. Shadows [Tit74a]. Shannon [Smo71l]. Shape [Fox74, Edw74]. sharp [Ray70-30, Tay74i]. Sheaf [Tie72, BG73, Fár72, Sch70a]. Sheaves [BGvO71, KT71, ST71e, ST71j, Van71, Gri71b, Ive70b, Kei71, Knu71h, Kom73e, Siu71b, ST71i, ST71k, GU71g]. shells [LS72]. shifts [Orn70, Smo71i]. Shimura [Del72b]. short [Bre71m]. should [Goo73b]. sided [Pet74e, Rog70d, Spi74a, Str74a]. sides [Tau74]. Siebenmann [Mor71a]. siècle [Die72]. Siegel [Kan71a, Kan71d, Kan71g, Maa71r, Mat70, Mur72a, Mur72b, Mur72h, Mur72d, Mur72f, Mur72g, Mur72i]. Sierpiński [Del71e]. sieve [Gal72b, Mon71a, Mon71g, Mon71q]. Sieves [And72]. Signal [OG74l]. signals [OG74a, OG74j, OG74s]. signature [Hir71c, Zag72a, Zag72b]. similarity [Fin74e]. Simple [Cha74d, Dem70f, GJ72c, Kom72d, Mat73o, Por71b, Chu71b, Fer74b, iI71, Mat73n, Sat73b, Sie74, Ste74a, Tho74b, Tit74d, Wil73d, dB71a]. simplement [Lat72]. Simples [Aze70b, Bor71a, Dix74, Mar71b, Tit71a]. Simplicial [BK72r, Bro74b, May72g]. simplifiée [Mey71a]. simply [Cer71]. simulating [Pim73]. Simultaneous [Bom73, Hea74b, McC71]. Sinai [Smo71i]. Singer [KR72]. Single [DD74a]. Singular [Bré74c, Com71, Fif73, Hab74, Ner71i, Ner71k, Ner71j, SM74a, Stu72, Tay74j, dJ72, BR73a, BdS74, Cap73a, Del70d, Kha71, Kri70, Mor71d, Sib71b, Sim73, Sto70b, Sto70i, Lio73d, Ner71l, Zyg71c]. singulärer [Her74a]. Singularitäten [Pop70i, Pop70e, Pop70h]. singularités [Kan74, Poe74n]. Singularities [Lev71b, Lev71d, Oli71, Poh71, ST71k, Sul71b, Wal71c, Wal71d, Ham71, Hir72d, Lev71e, Lit71, OW72, Orl72i, Poé71a, Pop70e, Por71b, Tay74f, Wal71f]. Singularity [Gil74l, OG74l, OG74r, Bra71, Hsi73, Iwa71a, Kan73a, OG74q]. Singularly [Bax74, Har73]. singulier [Ber74h, Gér74]. Singulière [Lio73a]. Singulières [CW71a, Lio73d, Mar70c, Zyg71c, CW71d, Zyg71a]. Singuliers [Del70d, Lio73b, CW71b, Zyg71e]. Sinne [Sch71s]. Site [Gro72e]. sites [GV72a]. Six [Sat73a, DI71f]. six-term [DI71f]. Sixtieth [Pet70a]. size [Csö73]. Skew [Coh72j]. Skorohod [Mai72]. Skorokhod [AM74, Bre74b, Hea74a]. Smale [God71b, Shu71a]. Small [Dev73r, Gol71a, Fre74a, GS71, Hec74a, Hop71, JW73h, Kni74, Lip73, O’M74]. small-cancellation [Lip73]. smash [May72p]. Smith [Chu73]. Smooth [AK70i, Ive73e, Omo74k, Wan72]. smoothed [DD74d]. Smoothness [Yam74t]. Sobolev [Ner71k]. Society [BK74]. software [Hul74]. soit [Fou74]. solar [Bro74a]. Solovay [Gar74b, Kri71]. Soluble [Cos74, Gru73a]. Solution [Bet74b, Mat72, Pat74, Wat74, Auc74, Azr72, BR73a, Cha74a, ESS74a, Eva71, GP74, GL74, Nob71, SC73, Spi74a, Sto74, Tak71a, Top70h, Top70p, Wac71a]. Solutions [BS73b, Eke70, Kré74a, Lie74, Mar70c, AJ73, Ars72, Aut72, Ban71, Bec73b, BD72e, Bra72a, Bro70, Bun72, Car71e, Car74c, Coh73b, CG71b, CG71a, Col73c, Coo71b, Dye73, Esp74, Fin74m, Fin74n, Fre71a, Fri72, Fri74, Gan71c, Gil74g, Gil74l, Gol71a, Hen73, Hsi73, HK71i, Iwa71a, Iwa71b, Jos73, JLP73, Kan73b, Kaw73a, Kha71, Kim73, Kom73e, Kur70b, Lan73a, Lou73, Nak73, Nas74, Net74, Par72c, Sat73c, Sei71, Sim72c, Str71, TP71a, Wal74m]. solvability [Trè74]. solvable [Ste74f, Lip74g]. Solving [Wil74a, Cia74, Gol71c, Hul74, You74]. solvmanifolds [Aus73b]. Some [Abi74b, AHK74, Aul74, Azu73, Bas73d, Bau74b, Ber70a, Ber71h, Bir73, BZ71, Boo74a, Buc72, CHS72, CM74e, CM74f, Chu72, Coh74u, Dos74, Dud70, Dun72, Dye73, Fri73c, Gan71a, Gan71h, Gar72g, Ger73c, Gol71b, Got73, Gri71a, Hak74, Han70, HC70i, Hea74b, Hea74c, Hed72, Hod74, Inf71, Ive73f, Kad70, Kai74, Kar73c, KS71b, Kub71, Kur70a, LRV70f, Lip74h, Maa70, Mil74l, Moo70, MV70c, Mos72, Mró70, Mur74, Myh73, Nad70, Nob71, Ost70c, PT73f, Pay73, Ram74b, Rei71p, Rei71q, Rox71, Sch73-32, Sha73b, Ste70o, Tho71c, Tho73b, Wal72e, Wid71a, Yam74k, dlH72h, Bel70, Bel71b, CM72, Col74e, ER70, ES72b, Gal72b, GL74, Gou72a, Gri71b, GM71e, Hag74, Hes74b, Hir71c, KMT74g, Kau72, Kom72f, Mil74e, PS72a, Sto71, Sub72, Tho73c, Zag72b, dB70, Bre71i]. Some [vL74i]. sommantes [Sch72f]. sommes [FS70e]. son [Fou74]. sont [Gat73a, Pet74a]. sources [KO74g, LRV70a]. Sous [Bor71a, Car71a, Dem70g, Dem70j, Gro70d, Gro70g, Ray70a, Cha72, Fer70f, Fer70g, Gat73a, Ray70i, Ray70k, Ray70-32, Ray70-33]. sous-algèbre [Fer70f, Fer70g]. sous-algèbres [Gat73a]. Sous-ensembles [Car71a]. Sous-Groupes [Dem70g, Dem70j, Gro70d, Gro70g, Bor71a]. sous-suites [Cha72]. Sous-Tores [Ray70a, Ray70k]. Souslin [Del73s, Dev73j, Dev73s, Dev73t, DJ74f, DJ74e, DJ74j, DJ74l, DJ74k]. sousliniens [BI70b]. Space [AS74c, DD74a, Pat74, Sch73f, dlH72a, Bel71a, Ben72a, Blo73b, BK72o, BK72d, Bre71f, Bre71e, Bre71m, Doo71, Fre73, Gui72a, Gui72h, Gui72j, Gun71, HK70, HC70j, HC70i, HM74a, HM74b, How74, Kri70, Lev73, Lor74, Maa71g, Maa71o, Maa71p, Mas74b, Mil74i, Mil73a, Mis71a, Mró70, Por71c, Roy74, Sch73i, Sin73, Sta70l, Sta71b, Sta74a, Tho71a, Top70l, Woo74a]. space-time [Sch73i]. Spaces [BD72a, BD72e, Ber70c, BV73d, Bur74b, DD74a, Gui72i, Har74g, Her71a, Kel74c, Kel74e, Knu71c, LT96, May72h, PT73i, Sig71b, Sta70q, Tre73e, Wae71b, Wil74b, Yam74p, Zag72a, AdK74b, AS74f, Bea70e, Bec71d, Bel74b, BD72b, BD72c, BW74, Bic73a, Bic73f, Bla74, Bli71a, Bli71b, BV73e, BV73g, Bol74, Bon72a, BL74b, BGR73d, BGR73e, BK72p, BK72n, BK72b, BK72c, BK72j, BK72q, BK72r, Bou74b, CT74, Coh70a, CD71, Cur71, D’A74, Day74b, De 73, Del71d, Din73, Dos74, Dwy74, Fer74c, Fox74, Fro74, Gui72e, GH74c, Hag74, Ham71, Han71b, Har74e, Har74i, Har74h, Hav74, Hea74c, Her70, Her71b, Her71d, Her74c, HR70, Hil74a, Hor73, HN74, Jam70a, Jam71, Jam70b, Jam70e, Jam70f, Jam70g, Jan71, Jon74, Kli70, Knu71b, Knu71f, Knu71h, Kra74, Lel71b, Lel74c]. spaces [LT73, Luc71a, Maa71t, Mas71, May72g, May72i, May72m, May72n, May72b, Mil74e, Mil74h, Mil74m, Mis71b, Mor71d, Nar70, Nar74a, Nar74b, Ner71k, Nov74a, PT73c, PT73b, PT73e, PT73g, PT73h, Par72h, Pet74b, Pet72, Pie72, Ric71, Roi71, Sat71, Sch74e, Sha71b, Sha71c, Sic72, Sta70b, Sta70k, Sta70m, Sta70o, Sta70e, Sta70f, Sta70g, Sta70a, Sta71c, Sta71d, Str74a, SR73, Sul71a, Sul71b, Tel71, Tit74i, Tit74j, Tit74h, Tur73a, Uen74, Wae71a, War70, Wil74c, Wil74d, You70, Zab70, Zab71, dlH72e, rd74c, rd74d, vL74j, LT96, Wae73]. span [Day74b]. Spanier [Coh70d]. Spanish [Git72]. sparse [ESS74a]. Spatiales [Mar70c]. spec [Hof70f]. Special [ABK71g, Bic73b, Rog70f, Sch73f, Bau71c, Ber71h, BK74, Hak74, Hol72d, PS72a, Spr70a, Wei71j]. species [Luc73o, Luc73j]. Specker [Fel71c]. Spectra [JW73k, And74b, BGR73c, Coh70a, HS74a]. Spectral [Coh70e, FNSS73g, JW73l, Smi70c, Cha70, Dom73, Her72b, Hof72a, Mey72h, Mil74n, Moo71, Poe74r, Smi70b, Sta70c, Tak70m, Wlo73, vW73, Jen72g]. spectrale [Gou71, Poe74r]. spectrales [Fer70g, Jen72g]. spectraux [Fer70f]. Spectre [BGM71b, BGM71c, Fer70h, Fer70d]. spectrum [Aur74, BGM71c, Coh72j, GM71b, JW73a, JW73b, JW73e, JW73m]. Spektren [BtD70j]. spezielle [Rau72, Eng73]. spezieller [Pit72a]. Sphere [HMR71, Sta71d, Lod73]. spheres [AHV74, Bre72a, Cob73, Gla71c, Gor70, Hak74, HMR71, Kaw72a, KK72, Mah74, MY72, Pet70c, Sch72d, Sig71a, Sta71d, Wan72, Wan74, Ros72]. Spherical [Hus70, Ner71m, Sil70a, Tit74f, Zab71]. Sphérique [LR71]. sphériques [Rei74a]. Spielsystem [Sch71e, Sch71n]. Spielsysteme [Sch71k]. Spline [Tho74c, Wen74b, Sch74a, Mic74, Sch74a]. Spline-Funktionen [Mic74, Sch74a]. Spline-Galerkin [Tho74c, Wen74b]. splines [dB74]. split [Cur70]. Splitting [Coh72h, NS71]. spoked [CS74]. Sponsored [CFK71, Cur72]. sporadiques [Tit71a]. SPRAYS [Sch73c]. spread [Ray70-30]. Spring [Eic71e, Her71a]. sprouts [Mos72]. SQ [Sac73]. SQ-Universal [Sac73]. Square [Mey71g, Kni71, Mey71f]. square-integrable [Kni71]. Squares [WWS72a, Wal72d, Lin74, Pal71, vL74n]. squashed [GP72]. SSOR [You74]. St [CFK71]. Stabile [BtD70k]. Stabilität [Has74]. Stabilité [LR71, Che74b, Mou72i, Poe74t, Rou74, Poe74k]. Stability [Den73a, Fri72, Gro70b, Hol74a, Jam72a, Koz72, LS72, Mat71b, Mey73a, Sat73i, Sma71c, Ste71h, Wal73c, Was74d, Yos71, Aus70a, BW72, Cho70, Dat70, Gil70, Gra74b, Jos73, Kat70a, Kat70c, Kus72, Lev71c, Sat73f, Sug71, Tak71b, Tho71d, Wil72, vdH73, Cur72, Wal71b]. Stabilization [Hau72, Hal70a]. Stabilized [Mil73b]. Stable [AR73, Bun72, Coh70f, Coh70g, Fin74n, Mat71c, Mil74o, Ric71, Was74e, BGR73c, Coh70e, Iwa71b, Kap70, KKMSD73b, McA74b, McG73, Mil74b, Mil74l, Pri73a, BtD70k]. stables [Fer74b, Poe74j]. stage [BL70]. Stallings [Kos71, Poé74a, Poé74d, Poé74g]. Stammfunktion [LZ73]. Standard [Bac72, BL74a, KM74, Pop70e, Tak70n]. Standing [Bea70l]. star [Ste70b, Wie74, Woo74d]. starke [Wil70i]. starken [Sch71f]. stars [Auc74]. start [DJ74m]. starter [And74a]. Startlösungen [Wac72]. State [BRC74, DF74b, Doo71, Fle70, Suc70, Kno72]. statements [Iga71, dB71a]. States [Ura71c]. stationary [KO73b, Kot73, Mar73c]. Stationnaires [Lio73i, Dup74b]. Statistical [Hol73, Leb73b, Rue70, Tho73b]. Statistics [MS73e, CA74, MS73b, WW74a]. statistique [Car74a]. Steenrod [Pet70a, BtD70l, May70, Whi70]. Steenrod-Operationen [BtD70l]. Stein [Lit74b, Moz71g, Ver71, Wel71a]. Steinberg [Kla73]. Steinhaus [Lel71a]. step [Bea74, Hay74, Per74, vdH73]. stepsize [Kro74]. Stereographic [Woo74a, Woo74b, Woo74c, Woo74e, Woo74d]. stetige [Sei72h, Sei72i]. stetiger [GU71i, GU71j]. Stetigkeitseigenschaften [GU71d]. Stiefel [GL70]. Stieltjes [Gli72, GW72, Hir74a]. stiff [Lam74]. Stochastic [Cur72, Gui72i, Itô72, Kus72, Mey72g, TP71f, BW72, Bun72, Erm70, Fel70, Fri72, Gri73, Har72b, Kaz72c, Koz72, LS72, Man72, Mor72, NS72, Nak73, Par72g, Pfe70, Set72, TP71g, TP71h, TP71i, TP71j, Wil72]. Stochastique [Aum70]. Stochastiques [BCM73, Pri74, Bad70f, Del74a, DDM70, DD71a]. Stokes [Dye73, Jos73, Tur71]. Stolzenberg [Wer71]. Stone [D’A74, dSL70].

Stopping [Din74b, Hea74a]. stoppverteilungen [Ros71]. storage [BW71]. Straight [Lio73g]. Strange [Bre72a]. Strasbourg [CD72a, DMW74, DE71, KGLZ70, Mey70b]. strategy [Pfe70]. stratifications [Kuo71, Sto72a, Wag73]. Stratified [Sto72f, Tho71b, Wal71g, Ray71d]. strict [Jac70]. Strong [PT73g, Bre71n, DK74, Ede70, Kin72, KŌ73a, Kre70b, Omo74c, Omo74d, Omo74l, Sca71a, Bre71n, Lan74]. Strongly [Mat73p, Kak73, MR71]. Stroock [DDDM70]. Structural [Bro70, Wag70, Pit72f]. Structure [Aus71, Bra74, DG70a, DG70c, Eng71a, Hir72a, Ive73b, Jat70e, Kom73f, Ler71, Mok71, MV70d, Mur72j, Ray70-30, Sil74d, Sta70r, All73, Bel71a, BS73a, Cla71b, Coh72i, Dev73f, Gol72, Hau73, HMS74b, IW73, Law73, Mah74, Mil74m, Pop70e, Sha71l, Sha71k, Sta70l, Tac73m, Top70m, Wal74j, War73, Yat72b, vdK73a]. Structures [BV73d, BV73g, Dem70k, Lut74, Mou72g, Poé74e, AB72, Bin73, CM74h, CD72b, Cor74, DM74, Fer70b, Gun71, Jón72b, Luc71b, Mou72c, Mou72b, Mou72h, Rec71, vdK73h]. Struktur [Pop70e, Pop70h, Pop70g, Pop70f, Sac74a]. Struktureigenschaften [Pit72f]. Strukturierten [AG73b]. Struktursätze [Hey70j]. studied [Jam72b]. Studies [Nam74]. Study [AK70j, BCC^{+}70, WP74b, Bre71i, Dat70, Gri71a, NS72, Rya72, Trè74]. studying [Mor74]. Stufe [Lor70l]. Sturm [Eva71, Fai74, Leb73a, Rab71]. style [DJ74g]. sub [Sch73c, Sin73]. sub-algebras [Sin73]. sub-SPRAYS [Sch73c]. subalgebra [Ste70k]. subalgebras [Ste70g, Tak70k]. subcategories [Hon74a]. subclasses [Won70]. subcontinua [Cur74]. Subdirect [BC74b]. subdivisions [Knu73a]. subdominantly [How71]. Subfields [MS73a]. subgraphs [Chi71c]. Subgroup [Yam74j, Wil73c, Yam73, Yam74g, Yam74f, Yam74i, Yam74h]. Subgroups [Bry74a, Bur74d, But74c, Bur90, Coh73a, Dix73, Kan71e, Mis74, Omo74d, Rei71m]. subharmonic [Ess74b]. subideal [Ste70k]. Subideals [Ste70p, Ste70l]. subjected [LS72]. Subjective [dF74a]. submanifold [cH73]. submanifolds [Poh71, Wel71a]. submodels [Gai72]. subordinateurs [Mai71]. subregular [Ste74e]. subset [Tho71a]. subsets [Har70a, Hea74b, Mró70, Sic74]. Subsheaf [ST71l]. Subsheaves [ST71e, Thi70]. subspaces [DP73]. substantielle [Rau74]. Substructures [Jón72e]. Subvarieties [Har70d, Shi71a]. Subvariety [ST71m, Har70e]. Successive [rd74e, vdC73]. Successor [MS70d, Dev73j]. Suées [dM70]. sufficient [WW74a]. suffisante [Hei73a]. suffisants [Poe74k]. Suisse [Sig71b]. Suite [GS70, Jou73b, DC70]. Suites [Eke70, Kat73h, Cha72, DC70, Jen72g]. sujet [Del70f]. Sum [Str72, SW74a, WWS72a, And73b, Tak70o, Wal72a]. Sum-Free [Str72, WWS72a, SW74a]. summability [HC70f, HC70g]. Summary [How71, OG74s, And74c, Kom72f]. summation [Ber72a, Tre73f]. Summer [Chi71a, DK73b, Kui73a, KS73b, Kui71a, MR73, SJS73, DS70a, Wae73]. sums [Kru73, Rao72a, vdK73d]. SUNY [McA74a]. Super [Tul70a]. Super-mean-valued [Tul70a]. Superconvergence [DD74e, DD74c].

Supercusp [HC70k]. supercuspidal [HC70a, HC70g, HC70k]. supérior [Git72]. Supersquares [Wal74l]. Supplementary [Par72j]. supplements [Eym72a]. support [Lel74b, Vau73, WR70]. supported [Ben71j]. supports [Del73b, Gro72d, Mar70e]. sure [Hei74a]. surface [Art72b, Bom70, Gra74a, Gun71]. Surfaces [AG71a, Fay73g, LB70h, Abi74a, AG71b, Bec74a, Del73f, Fay73a, Fay73c, Kim71, SM74a, Zim72j]. Surgery [Mil71, Sha73c, Tay73, Mau73, Pet70c]. surjective [Din74a]. surmartingales [Mey72a]. Survey [Lip74h, Ogg73, Pat74, BV73f, Bre71m, Bur74e, DS73, Har74c, Har74d, Kle73a, Lin74, Luc73m, Mar74a, Mar70d, Mey71g, Mic70, OG74s, PE73, Rub73, Sta71c, Tur74, Wal71g]. Suslin [Jec71s]. Suspension [HS73a]. suspensions [Gla71c]. sustainable [Dun70b]. swallowtall [Woo74b]. Sydney [Kel74i]. symbol [EL73]. Symboles [Maz73]. Symbolic [Bow73]. symbolique [Fer70c, MM72]. symbols [Gra73, OG74n, Smo71k]. symétrique [MM74d]. symétriques [Geb74]. Symmetric [Ben72a, CL74a, Gui72i, HB70, Knu73g, Rei71r, Sil74e, Zag72a, For74, Fuk73, Gui72h, Gui72j, Har74i, Hof72a, Ive70e, Ker71a, Maa71t, Pri73a, Zag72d, dlH72e]. Symmetries [JW73n]. symmetrischer [Sch74c]. Symmetry [CM74g, Sto72g]. Symplectic [CM74h, Maa71s]. symplectiques [Lic74]. Symposium [Chi71a, CG74, Cur72, Eng71b, Hil71, HL74c, MP73a, Nac74a, Wal71c, Wal71d, BCdV^{+}70, Bro71a, HS73b, Kno73b, LLNS70, Mor71c]. synthèse [Mey70g, Poe74r]. synthesis [Dom73, Her72b, MW71, Mey72h, Poe74r]. System [Åst72, Fit72, JW73a, JW73b, JW73m, Koz72, Kub73, Luc73f, NS72, OTF70, Sca71a, Sca71e, Sca71j, Sty74, Wed72]. Systématique [PP74b]. système [Gér74, Sal74, Haa74]. Systemen [End74]. Systèmes [Lio73b, Lio73a, Con74, FS70f, Jen72j, Lob72, Mar74b, Mey73h]. Systems [Bru74i, CM74a, Chi71a, Cur72, Kom72d, Sca71g, Ste70q, TP71f, Yor70, Aus70a, Aus73a, Baj73, Bel71b, Ben70, Ber70b, Bow73, Bré74c, CM74b, CM74d, CM74c, CM74e, CM74h, Cop71d, EHR73, Fef72, Fri73c, Gil74f, Gil74e, GW74b, Gob74, Har71b, Har72b, HL74b, Hau72, Inf71, Jam72a, Kat70a, Kom73f, Lam72, Lam74, Man72, May72q, McC71, Mes71a, Mes71b, Mor72, Oku71, Par71a, Par71b, Pul70, Ree73, RW70, Rob71b, Sai71a, Sch71x, Set72, Sma71c, Smi74, Str71, TP71g, TP71h, TP71i, TP71j, Ura71a, Wal73c, Wil72, Yos71, dB71a, dV70]. Szasz [Fer74a].

T [Lat72]. T. [Tat71]. Table [Aku73, Tem71b, dlH72d]. tables [CM72]. Tagung [AT72, AT73, AT74]. Takahashi [Hey70e]. Takens [Lan73a]. Tamagawa [Mar71b]. Tame [GM71g, GM71h, GM71i]. Tamely [GM71k, GM71j, GM71d]. tangent [Blo73b]. tangential [Bre71g]. Tangents [Dem70d]. Tannaka [Hey70g, Riv72c]. Tannakiennes [Riv72b, Riv72g]. Tate [Bas72, Mes72b, Mes72c, Mes72h]. Tatsuuma [Hey70f]. Tauberian [Gan71b, Gan71d, Gan71i]. technique [Bal70, New74a]. Techniques [HN73, SD72, Bec71b, Bec71c, Jan74b, JL71, Kno73a, Moo74]. Technology [vL74a]. Teichmüller [Gra74a, Mas74b, Roy74]. Teils [Pop70g, Pop70f]. telegraph [LRV70c]. tempérées [Fer70a]. Temps [Aum70, Del73p, MM70, Mey71e, Mey72c, Azé72, Mai71, MR72, Mey74e]. Tensor [Gui72i, PS72d, Tak70o, Wil73e, Gui72a, Mac70, PS72c, Wil73d, Hou72a]. Tensoriels [Hou72a, Sch72f]. tenue [Aku73]. term [DI71f]. Terminologie [GU71p]. terms [OG74i]. test [Kuo71]. Testing [WW74h]. tests [Bau71h, Chi73]. Texas [Bet74b]. th [NS72]. Theôrie [Gro71a]. their [Bea70e, BD72a, Bli71b, CM72, Col74f, Hau73, Hof70n, Jen72f, MP71, Par72i, Sil70f, ST71k, Smi74, Ste74b, Sto74, Sug71, Ura71a]. theme [Ste70a]. themes [Lin71]. Theorem [Knu71e, Sch73-27, Sto70f, Zag72a, ABK71a, ABK71f, Ark74, Ati74f, BS71, Bea70d, Beh74c, Ben71g, Ben71h, BŁ71a, Bol74, BE74, Cer71, Chu71b, Coh72i, Coh74m, Day74a, Deu73d, Dev74, Dev73h, Dev73i, Din71a, Dro71, Eis71b, Ess74c, Fai74, Fel72, Fer74a, Fra70a, Fre71c, FNSS73e, Gab72, Gan71j, GW74a, Gat73b, Hae71a, Han71a, Hof71, cHyH70, JW73i, Kel72b, Knu73d, Kow73, Laz72, Łoj71, Luc71e, Mat71a, May72c, McG73, Mon71b, Mon71l, Mor73, Moz71a, Moz71c, Moz71d, Moz71e, Moz71f, Moz71g, Nir71, Omo74b, Omo74c, Omo74k, Pap70, Par72l, Pat70, Poe74r, Rab71, Ray70x, RW70, Sch73u, Sch73-28, Sch73-32, Sha71a, Siu71a, Siu71b, Smo71a, Smo71d, Smo71f, Smo71l, Tak70a, Tel71, Wae74b, Wal71a, Wan70, Was74c, Yam74a, Zag72b, vL74h, vdK73o, Beb74, LL74e]. theorem-proving [Wan70]. Théorème [Art70, Art73h, Art73g, Art73i, BGI71, Bru74f, Fra70a, Kat73e, LB70i, Lel71a, Mou72j, Mou72h, Mou72i, Poé74g, Bad70e, Ber71a, Ber71o, Cas70, Che74c, Con74, Del73k, Del71g, Del73s, Dem70i, DM74, Dou73, Fer74b, HN71j, Kat73c, Lel74b, Maz74, Mey71a, Mey71e, Mey72d, Mey74h, Nov73, Poe72, Poe74o, Poe74s, Poe74p, Poe74q, Poé74d, Poe74r, Ray70w, RK71, Rev70, Sch72b, Sun70b, Trá74c, Art73f, CW71g, Gla71b, dSL70, Poé74a]. Théorèmes [Kle71, Poé74g, Ram74c, Pis74, BB70, Ben73, Del71e, Del73o, Nov70, Sch71a, Kri71, Hei74a]. Theorems [Gan71i, PS72d, Par72i, AK70d, And72, Beh74b, Bon72a, Bor73, CM71, Coh72g, Coh72h, Col74d, Csö73, Cun74, DeV72e, Dia72, Fan70, Gan71b, Gan71d, Gan71g, Gan71h, Gro72a, Har70g, Hea74c, Hic73, Ive73f, Jaw70, KMT74c, Kat71, Kel74f, KS71b, KS72, Kru73, Lam71b, Lin74, Lov74, Luc70, Mon71c, Mon71f, Mon71m, Mon71n, Nak73, Omo74a, Omo74b, Par72e, PS72b, Par72j, Pow70, Rad74, RHD70e, Sch73p, Sim72e, Sim72d, Smo71i, Sne71, Spi72, Sun70a, SE70c, SE70d, Ter70, Yam74k, Yam74r]. theoretic [Dob70d, ER70, Goo71, Gru73a, KM74, Rya72]. Theoretical [Cur72, Kub71, Nic73]. Theoretische [NW72]. Theorie [AH70f, Hey70i, ZVC70g, Art72a, Art73j, BGI71, FS70g, Gro71b, Jen72f, Jou73b, KGLZ70, Kar73b, Las71, Poé74f, Poé74c, Sch70f, Sch72a, dF74a, Car72c, Car74b, Cha72, Con74, Del73d, Del70i, Del70j, Del73o, Den73b, Die72, FL70, Gou71, HN73, Jen72a, Jou73a, Kan71k, Kar70b, Kar70c, Kri71, Mar70a, Mey72a, Mey73g, Nor71, Zis70, AH70g, BtD70e, AGV72, Bru71c, Bru74q, Che74d, Dol71b, Far74, HN71m, KO74e, Mey70h]. Theories [Bas73c, Lam71f, MS70d, MS73j, And71a, And73a, AKW73, BV73a, BV73h, Bre71h, Bre71m, Dob70e, Elg71, Fri73a, Grz70, Lam71c, Luc73p, Sto72a, KO74b, dFP74]. Theory [AS74c, ALW72, AG71a, AK70g, BH74b, Bas73a, Bas73b, Bau71a, BKW73, Ber71f, Bic73b, Bon72b, Boo74b, Bre71p, Can74, CFK71, Coh74b, Cur72, Deu73c, Dol70a, Dub70d, ES72c, Fel71d, Fil73, GW73, GG72, GJ72a, GJ72b, Gra74h, Gru70c, Hil74c, Hof72b, Hol72e, Hol72f, HS71, JL70b, JL70c, Jec71l, KT71, Kel74i, Knu73g, Kom73g, Kom72d, Kre73h, MP73a, Mon71p, Mor71c, MS73j, New74b, PS72d, Sat73i, Sca71g, Sch73k, Sch73-29, Sha71j, Sim72a, Sio73d, Smo71g, Sta73d, Suc70, Swa70, Tak70p, Tel71, Tre73e, Zag72a, Zim72c, vL71b, vL73b, vL74a, Adl74, Bas73d, Bas73e, Bec74b, Ben71i, BL74a, Bic73g, Blo73a, Blo73b, Bor71b, Bre71o, BtD70h, Bro71b, BG73, Bro74b, Bru71a, Bru70, BS73c, Büc73a, Bur74e, Cap73b, Cha74d, Cha74c, Coh72b, DI71c, Dev73u, Dob70d, Dre73, Dzh74]. theory [Edw74, Eic70, Eng70, Fai74, Fin74h, FS70g, Fox74, Fre73, FNSS73a, FNSS73d, Fug72b, Gal72b, Gel74, Ger73a, Ger73c, Gil74f, GL70, Gli70, Gol73, Goo73a, Gri73, Gro70a, GM71e, Gru70a, Gru70g, Gru70l, Gru73a, Gus73b, Han71a, Har72a, Har74c, HC72, Har71a, Har71b, Hea74a, Her71c, Hod74, Hof72c, HMS74a, HMS74b, HMS74d, HS70, HR74c, cH73, Jac72c, Jac72d, Jan71, Jat70a, Jec71u, Jen72f, Kad70, Kai74, KMT74h, Kar70c, Kar73c, Kel74a, Ker71a, Keu73, KN73, Knu73f, Koh72, Kom73a, Kom73d, Kom72a, Kom72e, Kre73a, Kre73c, Kre73b, Lan70, Las71, Leb73b, Lic73, LKR73, Lip74b, Luc70, Maa71m, Man71, Mar73b, Mar73c, Mes71a, Mes71b, Mes72e, Mey72f, Mon71i, Myh73, Omo74d, Pet70c, Pet73, Pie72, Poé71a, Pri73a, Pro70, Qui73b, Ran73b, Ree73, Rim72, Rob73m]. theory [Sai71b, Sat73j, Sca71b, Sca71e, Sch71v, Sch73-28, Seg70, Sha73a, Sha73b, Sha71b, Sha71c, Shi71b, Sil74c, Sil74d, Sne71, Sta73b, Sta72, Ste71g, Swa73, oT72, Tho71d, Tie72, Top70g, Tre73b, TP71j, Uen74, Wag73, Wal73b, Wal74m, Wlo73, Zag72b, Zis70, tDKP70d, vW73, AH70g, HN71m, Sch70f]. Theta [Fay73g, Fay73f]. thin [Gam71, Kom72a, Kom72e]. thinness [Bre71a, Bre71e, Bre71i, Bre71k, Bre71m, Bre71n, Bre71s]. thinnesses [Bre71g]. Thom [Che74b, Geo74, Ray71d, Tho70]. Thomas [Bra72b]. Thomist [Por71a]. thoughts [AHK74]. Three [EE73, Mon71o, Sch73l, AHK74, FGQ71, Tay74a]. Three-dimensional [Sch73l, Tay74a]. threefolds [Mur74]. thrust [Ede70]. thrust-limited [Ede70]. Thullen [Nov73]. tightness [Kui71b]. Time [Sch73f, Bal70, Bel71a, Jac70, Sch73i, TP71i]. tires [Riv72d]. Toccata [Lin71]. Toeplitz [Cob73]. tolerance [Tak71b]. Tome [Art72a, Art73j]. Tomita [Dix71, Tak70p]. tool [Dat70]. Topics [Beh72a, Gru70c, Gui72i, Hil74c, Jón72f, Lip74h, Mon71p, Sat73i, Sha71j, Tho71d, WW74e, Hol72d, Koh72, Kre73j, Orl72g, Pet72, Sai72b, Yor71, HMR73, LLN74]. TOPO [AHiN74]. Topological [Bec73a, BV73d, BV73h, Chu71a, EHM70, Fle70, Goo71, Kaw72a, Luc71h, McA74a, Sat73j, Wae71a, Wae71b, Wae73, Yam74p, Adl74, AdK74a, Bel74b, BW74, Bla74, BK72r, Bre71q, Din73, Dos74, Goo73b, Gui72e, Háj70, Jam70g, Jan74b, Kai74, KS71b, Lel71b, Lel74c, Maa71l, Par72h, Rob73m, Sic72, Sta70b]. Topological-algebraic [BV73h]. Topologies [AdK74b, BI70d, Bre71p, Dem70l, HN71n, Lel72c, Mai72, Ver72c]. topologique [Che74b, HHP71, Sie74]. Topologiques [HN71b, Jac72b, HN73, Jen72c, Nov70]. topologized [Pet74b]. Topology [AHiN74, BGJ72, Ham71, Hil71, McA74a, Sma71b, Top70q, Bre71a, Bre71c, Bre71k, BF74, Cla71a, Fei70, Hal70b, cH73, Jam70c, Knu71d, Mar74a, Sha73a, Top70c, Top70f, Top70l, Whi70, DF74b]. topology-general [Bre71c]. Topos [AGV72, Art72a, Art73j, Gro72e, GV72c, Ill71d, KM74, Ber74m, Gir72, Ver72a, GV72a]. Topos-theoretic [KM74]. Toposes [Law72a]. Toral [Ati74i, Car72b, CR72c, KR72]. tore [Lod73]. tores [Ray70k, Gro70g, Ray70a]. Tori [KKMSD73a, CWW74, Lan73a]. Toroidal [KKMSD73c]. Torseurs [Ray70-32, Ray70-33]. Torsion [Lam71e, Lam71f, Ser71a, Ste71g, Car73a, HR70, Lam71c]. torsion-free [HR70]. Torsionselemente [Lor70n]. torsors [Ill72e]. torus [Kow73, Sai71a, Sai71c, Sch72d]. Total [Beh72b, Kat70c, Bru74g, Bru74h]. Totale [Bru74g, Bru74h]. totalement [BS74c, Del73p]. Totally [BR73a, Ben71c, Ben71j, Hof70j]. tout [CD74]. Tower [BK72e, BK72s]. Towers [BK72t]. trace [BS73a, Bro73b, Rin72b]. Traces [Eic73, HH73]. track [BGR73d]. trajectoire [Dup74b]. Trajectoires [Bad74f, Mar70c, Sal74, dM70, Hei73a, Hei74a, Sad74]. transcendance [Ami72, Wal74h, Wal74g]. transcendante [Del73d]. Transcendants [Wal74d]. transcendental [Gri71a, Wal74d]. Transfer [Pri73a, BD72f, Ede70, Pfe70]. transfers [Mos70]. Transfinite [Sca71k, Sca71b]. Transform [Ner71d, Gar72d, Wei73]. Transformation [Cho73c, Hir74a, KMSS72a, KMSS72b, Mey74f, Smo71e, Zag72a, Chu71a, HK70, KM72, Lam73, Omo74j, Sch73v, Sil70d, Zyg71a]. Transformationen [BtD70i]. Transformations [Kre70b, Omo74f, BS74a, Bor71b, BtD70i, DS70b, Kak73, Kri70, Pet74d]. Transformée [Zyg71h, Zyg71g]. Transformées [BBB70, Vau73]. transforms [Gli72, GW72, GK72, Kan71b, Ner71b, Ner71c, Sha71i]. transition [KŌ73a, Sat73a]. Transitive [Jec71t, Jec71n, Rob73o]. Transitivity [Bru74i]. Transitivity-Systems [Bru74i]. Translation [Luc73p, Ost70a, BdS74, Che74a, For74, Ost70c]. Transonic [Kor74]. Transport [LR71, AG74, Sta74b]. transversal [Bru71a, DK74, Mas71]. Transversales [Bru74n]. Transversality [Sto72h, Ing71, Sto72g]. Transversally [Ati74b, Ati74d]. transverse [HHP71]. Travaux [Che74b, Del71b, Del72b, Den73b, God71b, Ill73, Kat73f, Kos71, Poe71b, Ray71a, Ser71b, Szp73, Mey74e, Sch71w]. treatment [Her72a]. tree [DJ74f]. trees [Cad71, DJ74e, DJ74j, Har74b, Luc73j, McA74b, Sca71b, Zuc73c]. Trends [CFK71]. très [Kui73b]. tresses [Bri73]. Trèves [Gri72]. Triads [Świ70]. triangle [Kro72]. triangle-free [Kro72]. triangleless [HS72]. triangles [GS71]. triangular [Chu74a]. triangulation [Mor71a]. trick [vdK73p]. tridimensionelles [San74]. Trigonometric [DeV72i]. Tripartite [ADT72]. Tripel [GU71q]. triplets [Wlo73]. trivalent [Cad71, Sze74]. trivial [Gan71c, Mar70d]. triviality [ADS73]. Trois [Poé74g, Azé72]. truncated [Mil74e]. TTF [Azu73]. TTF-classes [Azu73]. tubulaires [Mou72e, Mou72j]. Tulane [Hof72c]. turning [Was71]. Tutte [BS71]. Twisted [Ran73b, Sch72e]. Two [Dia72, Jam72b, Jec71u, Par72k, Spi74a, BS74b, Dev73h, Dev73i, Fan70, GP74, Gil74d, Gil74i, GM71b, GM71c, HL74a, HK71i, KMT74c, KMT74i, Ker71f, Kor74, OG74r, Sne71, Vas73, vdC73]. two-dimensional [KMT74c, KMT74i]. two-point [GP74]. Two-sided [Spi74a]. typ [Eng73]. Type [DG70a, Gro70c, Gro70d, Gro70e, Gru72, Hun74, Lio73c, Lio73e, Lio73i, Ray70b, Tit74f, Wal74g, Ara71, Auc73, AM74, Aze70e, BB70, Bau71c, Csö73, Geb74, Gru74a, Gui72g, Gup74, Har74f, Hei74a, Hic73, Hsi73, Iwa71a, Jen72j, KMT74d, KMT74e, Kan71g, Kat71, Mai72, Mar71c, Mey73a, Nac71, Puk72, Rec71, Sie74, Siu71b, Sta73h, Sty74, Tac73j, Tit74i, Tit74j, Tit74h, Tit74g, Tit74k, Tit74n, Trá74c, TP71e, TP71j, Zab74]. types [De 73, Nar74b, Rog70f, Roi74]. typical [Gan71j]. Typus [Haa74].

Überlageruncen [Pop70c]. Überlagerungen [Pop70a, Pop70k, Pop70m]. Ultimate [Set72]. ultra [Hir72b]. ultra-products [Hir72b]. Ultradistributions [Kom73h, Cho73d]. ultrafilters [Pri73b]. Umbilic [WP74e, Tit71b]. umbilics [WP74d]. Unabhängigkeit [Pop70m]. unbounded [BL74a, Gup74]. uncountable [Hec74a]. underlying [KMT74j]. Unendliche [Bae74, Sei72j]. Unfoldings [Was74d, Was74e, Was74f]. unibranches [Ray70f]. Unicité [Dem70h]. Uniform [BD72a, Gel74, NS71, Rén73, BD72b, Fro74, Hag74, HM74a, PP74a, Sha71e]. uniforme [DC74]. uniformément [DD71b]. uniformes [Dos71]. uniformisation [Par72h]. Uniformité [Nac71]. Uniformity [Top70r, Top70o]. uniformization [Ear74]. Uniformly [Fin74o, Kat70c, Hag74, Mil73a, PS72b, Sim72d]. unilateral [Fic72]. unilatéraux [Lio71]. Union [Cur72]. Unions [Dru72]. Unipotent [KMT74a, Ray70c, KMT74b, KMT74c, KMT74f, KMT74i, KMT74h, KMT74j, KMT74k, Ste74d]. Unipotents [Ray70b]. unique [Fur73, Mil73a]. Uniqueness [Bec73b, Sto72i, Ben71a]. unirationnelles [Del73n]. uniserial [Mat73n]. unit [Beh74b, Col74d]. Unitaires [Eym72g]. Unitary [Bas73e, BDF73, HO72, Kos70c, Sha73c, Sil70f, Sur72, Tay74a, Wol71]. United [Ura71c]. Units [Ber71i, Zim72n, Rei71c, Rei71d, Rei71e, Rei71i]. Universal [Jón72f, Kan71j, Kan71i, MM74f, MM74g, Sac73, Was74f, vdK73p, Chu72, Kan71b, MM74e, OTF70, Por71c]. universe [Dev73e]. Universidade [Nac74a]. Université [CD72a, DMW74, Mey70b, DE71]. University [ALW72, BH74b, Bec73a, BRC74, Bet74b, CFK71, Chi71a, CG74, Cur72, DK73b, DF74b, Eic71e, Fil73, Fle70, GW73, GG72, HLM71, HS74b, Her71a, Hol74b, HS71, HL74c, Kno73b, KMSS72a, KMSS72b, Kui73a, KS73b, Law72a, Maa71r, New74b, Suc70, Yor70, vL74a, Hof72c]. Unramified [Ive73g, Fay73b, Mat73p]. unsolved [Abi74b]. Unstable [GM71a, Mil74o, Mil74n]. unstetiger [Tau74]. Untergarben [LL74d]. Unterringe [LL74e]. unthinness [Bre71n]. unusual [Bin74]. Upper [Bic73g]. USA [Hof72b]. usage [dB70]. Use [GP74, Rya72]. useful [BBF74]. Usefulness [Trè74]. using [Dom73, Kow70, Kro74, Pit70, Rho74, Wil74a]. USSR [MP73a]. Utah [Bro71a]. utative [KMT74c]. Utilisation [Mar70c, Sch71a]. utilization [FPC74].

V [DE71, vL74g, Bru74n, Gra71a, KO74f, Lip74f, Mey74d]. V. [Bri73, Ste74i]. vague [Top70l]. Valeurs [Dol72, Lel72d, Sch70g]. validity [Sco70]. valuations [Deu73b]. Value [Sim72a, Sto70j, Sto70k, AJ73, AH70e, AH70g, BR73a, BD74, Bax74, Beh74b, Bra72b, Bre72b, BW72, Ces71, Cod71, GP74, Gil74a, Gil74j, GL74, Gli72, Gou74a, Hab74, Har73, Hes74b, Hop71, HK71i, Kaz71, Las70, Lio73d, Mei72, Mon71c, Mon71l, Mon71m, Net74, Nob71, Per74, Sat73g, Sib71a, Spi74a, Sto74, Stu74, Tau74, Tay74c, Tay74d, Tho73a, Tho74c, Vel72, Wen72, Wen74b, Wid71b, You74, vdC73]. Valued [Sio73d, Bel70, Cod71, Dun72, Fra70b, Hal70b, Jaw70, Jec71m, Jec71u, KS71a, Kur70a, Luc71c, Nad70, Pow70, Sio73c, Tul70a, War70, Fle70]. Values [Lic73, Eis71a, Kor70]. vanishing [Coh74r, dJ72]. Varadhan [DDDM70]. Variable [DK73b, Deu73c, DD74a, Kui73a, KS73b, Ogg73, Per74, Dat70, GK70, Jan74a, Kel72c, Wen74b, Yam74n, Zyg71h, dB74]. Variablen [Haa74]. Variables [Bro71a, Gla71a, Hor70b, Hor71, Nor74a, Car70b, Col72, DC70, Gil74d, Gil74i, Her71e, JW73g, Kno72, Kor70, Kru73, Mar71c, Nor74c, PT73c, Zyg71g]. Variante [Meu73]. Variation [Bru74g, Bru74h, Bru74a, KK73, Bre72c, Bre72d]. variational [Ber70a, Her72a, Sta72, Zie74]. variationnels [Aut72]. Variations [Hum73, LT70, Mar70c, Ste70a, Glo70, Gol74b]. Variété [BGM71c, Car71a, Far73, Lel72c, Mou72b, Mou72h, Tho74a, BGM71b]. Variétés [Del73n, LB70c, LB70k, LB70l, LB70m, LB70j, Poé74g, And74e, Dem71, Kui73b, Lit74b, Mor71a, Mou72f, Nor74d, Tat71, BGM71d, Mar71d, Mou71, Mou72g]. Varieties [Bry74b, Eic71e, Har70d, Pop74b, Sta73g, Bac73, Dem72d, LB70l, Moi74, Pet70c, Rob73n, San74, Ste74b, Uen74]. variety [BGM71c]. variiert [Pop70o]. Various [Ber71j, De 73, Col74e, Gon74, Eym72a]. vecteurs [Aut72, God74a, Sad74]. Vector [Bic73b, Gor70, Hag74, Mur72k, Omo74l, Wae71b, Wae73, Bel74b, Din73, Dos74, Gui72e, Gun71, Hab74, Her71d, KMT74e, Lel74c, Luc71c, Oka72, Omo74g, Pet74b, Sic72, Tho71c, Wae71a, Wal71e]. vector-valued [Luc71c]. vectoriel [Mou72c]. Vectoriels [Jac72b, LB70g, HN73, Lel71a, Nov70]. Vectories [HN71b]. Vektorräume [Sei72g, Sei72e]. Verallgemeinerter [Lep72]. Verbindungsgeraden [Wil70d]. Verdier [Dou73]. Verfahren [Mei72, Mit74b, LZ73, Wac74b]. Verfahrens [Meu73]. Verhalten [Pop70c, Pop70o]. Versailles [LLNS70]. Versailles/France [LLNS70]. Verschwindungssätze [GR70]. version [Fre71c]. versus [PE73]. Verträglichkeit [BtD70m]. Verwendung [Mic74]. verzweigter [Pop70c]. Verzweigung [Pop70a]. VI [CD72a, Bru74o, KO74g, Lip74g, Mat71b, vL74h]. via [Hea74a, McG73]. vibrating [Kom72c]. Vibrations [Kom72d]. vicinity [IDS70]. Victoria [HL74c, HL74c]. Vier [Pop70n]. Vietoris [Cap73b, Jou73b, RU73]. View [Mil74o, Sta70f, KS73a, Por71a, RS74, Sta70e]. VII [Bru74p, DMW73c, vL74i]. VIII [DMW74, vL74j]. Ville [Sch71l]. Vinogradov [Mon71b]. Virginia [DF74b]. vis [Maz72]. vis-a-vis [Maz72]. Viscosité [Rav71]. Vlasov [NW72]. Vlasov-Gleichung [NW72]. voisinages [Mou72e, Mou72j]. vol [Ano71a, Ano71b, Ano72, Ano74, DE73]. vollständigen [End74]. Volterra [Lev71a, Noh71, TP71a, TP71d, TP71e]. Volume [Sch73-33]. vom [Bae74, Eng73, Haa74, Pop70g, Sch71e, Sch71n]. VOP [VCLN74]. Voraussagbare [Sch71q]. Vorbemerkungen [Sei72f]. Vorbereitungen [BtD70n, Hey70k, LL74f]. Vorlesungen [Pit72g]. Voronoï [Ber72a]. Vortrag [HK71d, HK71e, HK71f, HK71b, HK71c, HK71g, HK71h]. Vorträge [AT72, AT73, AT74]. Vorwort [Sch71r]. vous [Mar70b]. vs [Sei72c].

W [DDDM70]. W. [Ben74a, Bom73, Dev74]. Waelbroeck [Fer70d]. Wahrscheinlichkeit [Sch71v]. Wahrscheinlichkeitstheorie [Sch71v, Sch71p]. Waldi [HK71h]. wall [Mau73, Lat72, Mor71a]. wall-Novikov [Mau73]. walls [Kub73]. Warwick [Chi71a, Cur72]. washing [Str74b]. Washington [BH74b]. Watson [KF73]. Watt [Kno73b]. Wave [And74d, Dio74, LRV70d, Tay74f]. waves [Zee71]. Wazewski [Beb74]. Weak [Bre71s, Kli70, PT73h, Sic74, Wae74b, Wei73, Top70e, Top70c, Top70f, Top70n]. weaker [Fin74c]. Weakly [Dev73t, Maa71t, Bun72, Kak73, Lel74a]. wedge [Mor73]. Weierstrass [D’A74, Her74c, Poe74p]. Weight [GK70, OG74l, Shi73b, vL71i, vL73i, BD72f, OG74i, OG74f, OG74h, OG74k, OG74p]. weighted [Han70, Mon71q, Orl72h, Sum74]. weights [Bre71b, Fre74b]. Weil [Sch72b, Art70]. Weiss [Moz71g]. Weitere [Sch71s]. Well [Kno73b, EHM70, Mil73b, Ost70c]. well-ordering [EHM70]. wellfounded [Sca71b]. wellorderings [Sca71k]. Wendroff [Meu73]. Wendroff-Verfahrens [Meu73]. wenn [Pop70i, Pop70e, Pop70o]. wertaufgaben [AH70c]. Western [ALW72, GG72, HLM71, HS71]. Weyl [Car70a, CW71g, Gro70g, vdK73n]. wheel [CS74]. Wheels [Tut71]. Which [Coo74, HS74a, Kle72, Ben74a, Bre72b, Jos73, Kak73, Kap70, Law73, Sta70g, Ste70k]. whirls [Tut71]. white [Iha73]. Whitehead [Cas73c, Coh70d, Dro71, Wal73a]. Whitney [Kuo71, Łoj71, Poe74s, Poe74r]. whose [Lem71, Rob73o]. widerspruchsfreien [Sch71l]. Wiener [BC74a, Bad74d, Ben71k, Del74a, Gan71g, Gro70a, MM72, Meg71f]. wigwam [Woo74e]. wilder [Pop70a]. Williamson [Wal74k]. Wisconsin [GW73]. without [Fin72, Gan71c, Kre70b, Sca71k]. Witt [Dem72f, Lor70b, War73, Yam74a]. Wittschen [Lor70n]. Woods [Kri70]. Words [Coo74]. work [Lan73a, Whi70]. works [McD74]. world [Sat73h]. Wound [KMT74k]. Wreath [Ker71e, Ker71d]. written [Har70d].

Wronskians [Har71c].

X [Pop70m, vL74k]. XI [vL74l]. XII [vL74m]. XIII [vL74n]. XIV [vL74o]. XIX [Die72]. XV [vL74p]. XVI [Kun73].

Yale [Bec73a]. year [Hof72c]. yields [Dun70b]. Yoneda [Lin71, Str74d]. York [CFK71, Fle70, CFK71]. Young [Ben74b].

zêta [Ser73k]. Zabrodsky [Ark74]. Zahlen [Sch71f, BtD70c]. zahm [Pop70c]. Zariski [Ray70x]. Zeeman [Tak71b]. Zeilensummenbedingung [Boh74]. Zentrale [Hey70l]. Zermelo [Dev73u, Jec71c, Myh73]. Zero [Mon71r, Gir74, HMS74b, HMS74c, Rob73h, Rob73i, Sal73, Fer74b]. zero-dimensional [HMS74c]. Zero-free [Mon71r]. zeros [Hen71, Kni74, Mon71d, Mon71r, NF74, Wal74i]. Zeta [Deu73e, GJ72c, Maa71u, PS73, Ber74l, Lic73, Maa70, Maa71q]. Zeta-Functions [PS73, Lic73]. ZF [Fel71a, Fel71d]. ZF-models [Fel71a]. ZF-Set [Fel71d]. ZFC [DJ74m]. ZTi [Sca71h]. ZTN [Sca71h]. Zufälligkeit [Sch71v]. Zufallsfolge [Sch71s]. Zufallsfolgen [Sch71c, Sch71e, Sch71b, Sch71o, Sch71t]. Zufallsgesetze [Sch71j, Sch71u]. Zufallsgesetzes [Sch71h]. Zur [vF73, Has74, Mei72, NW72, Opf74, Sac74a, Sch71l, Wac72, Ade72, Ess73, FK72d, Gek72, Gek73, Gen74, Her74a, Sac74a, ZVC70g]. Zürich [Gra71b]. Zurück [Pop70o, Pop70p]. zusammenhängende [Pop70i]. Zwei [GS74a, Haa74]. Zwei-Punkt-Randwertaufgaben [GS74a]. Zweischritt [Meu73]. Zweischritt-Lax [Meu73]. zweiter [GS74a]. Zyklische [Hey70m]. Zylindersymmetrie [vF73].

References

Auslander:1972:FBS

[AB72] Louis Auslander and Jonathan Brezin. Fiber bundle structures and harmonic analysis of compact Heisenberg manifolds. Lecture Notes in Mathematics, 266:17–21, 1972. CODEN LN-MAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

            link.springer.com/chapter/10.1007/BFb0059633/.

Abhyankar:1973:ME

[Abh73] Professor Shreeram Abhyankar. On Macaulay’s examples. Lecture Notes in Mathematics, 311:1–16, 1973. CODEN LN-MAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

            link.springer.com/chapter/10.1007/BFb0068914/.

Abikoff:1974:MRS

[Abi74a] William Abikoff. Moduli of Riemann surfaces. Lecture Notes in Mathematics, 400:79–93, 1974. CODEN LNMAA2. ISBN 3-540-06840-6 (print), 3-540-37776-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065677/.

Abikoff:1974:SUP

[Abi74b] William Abikoff. Some unsolved problems. Lecture Notes in Mathematics, 400:124–130, 1974. CODEN LNMAA2. ISBN 3-540-06840-6 (print), 3-540-37776-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065681/.

Antonelli:1971:CT

[ABK71a] Peter L. Antonelli, Dan Burghelea, and Peter J. Kahn. A classification theorem for π_{i}(B, a; MrelX). Lecture Notes in Mathematics, 215:32–46, 1971. CODEN LNMAA2. ISBN 3-540-05560-6 (print), 3-540-36880-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061181/.

Antonelli:1971:CHG

[ABK71b] Peter L. Antonelli, Dan Burghelea, and Peter J. Kahn. The Concordance-Homotopy Groups of Geometric Automorphism Groups, volume 215 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05560-6 (print), 3-540-36880-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 140 pp. LCCN QA3 .L28 no. 215. URL http:/

/link.springer.com/book/10.1007/BFb0061176;   http://
www.springerlink.com/content/978-3-540-36880-9.

Antonelli:1971:EN

[ABK71c] Peter L. Antonelli, Dan Burghelea, and Peter J. Kahn. Exactness and naturality. Lecture Notes in Mathematics, 215:14–27, 1971. CODEN LNMAA2. ISBN 3-540-05560-6 (print), 3-540-36880-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061179/.

Antonelli:1971:G

[ABK71d] Peter L. Antonelli, Dan Burghelea, and Peter J. Kahn. The groups π_{i}(a; MrelX) and π_{i+1}(B, a; MrelX). Lecture Notes in Mathematics, 215:12–14, 1971. CODEN LNMAA2. ISBN 3-540-05560-6 (print), 3-540-36880-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061178/.

Antonelli:1971:PDL

[ABK71e] Peter L. Antonelli, Dan Burghelea, and Peter J. Kahn. Preliminary definitions and lemmas. Lecture Notes in Mathematics, 215: 1–12, 1971. CODEN LNMAA2. ISBN 3-540-05560-6 (print), 3-540-36880-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061177/.

Antonelli:1971:PCT

[ABK71f] Peter L. Antonelli, Dan Burghelea, and Peter J. Kahn. Proof of the classification theorem. Lecture Notes in Mathematics, 215: 47–78, 1971. CODEN LNMAA2. ISBN 3-540-05560-6 (print), 3-540-36880-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061182/.

Antonelli:1971:SC

[ABK71g] Peter L. Antonelli, Dan Burghelea, and Peter J. Kahn. Special computations of π_{i}(a; MrelX). Lecture Notes in Mathematics, 215:28–32, 1971. CODEN LNMAA2. ISBN 3-540-05560-6 (print), 3-540-36880-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061180/.

Ahmad:1974:IBG

[AC74] Afaq Ahmad and Leon Cohen. Integration of the N body gravitational problem by separation of the force into a near and a far component. Lecture Notes in Mathematics, 362:313–336, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066599/.

Adem:1970:NBM

[Ade70] José Adem. On nonsingular bilinear maps. Lecture Notes in Mathematics, 168:11–24, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058515/.

Ade:1972:ELG

[Ade72] Hans Ade. Zur Existenz von Lösungen gewisser nichtlinearer elliptischer Randwertaufgaben. (German) []. Lecture Notes in Mathematics, 267:3–14, 1972. CODEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061610/.

Alo:1974:TDA

[AdK74a] Richard A. Alò and André de Korvin. The topological dual, the algebraic dual and Radon–Nikodým derivatives. Lecture Notes in Mathematics, 399:22–32, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063564/.

Alo:1974:TIB

[AdK74b] Richard A. Alò and André de Korvin. Topologies induced by Banach function spaces. Lecture Notes in Mathematics, 378:16– 40, 1974. CODEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068457/.

Argabright:1972:AHM

[AdL72] Loren N. Argabright and Jesús Gil de Lamadrid. Analyse harmonique des mesures non bornées sur les groupes abeliens localement compacts. (French) []. Lecture Notes in Mathematics, 266: 1–16, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059632/.

Adler:1973:ER

[Adl73] Roy L. Adler. F -expansions revisited. Lecture Notes in Mathematics, 318:1–5, 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061717/.

Adler:1974:AEM

[Adl74] Andrew Adler. An application of elementary model theory to topological Boolean algebras. Lecture Notes in Mathematics, 369:

1–4, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065993/.

Alperin:1973:NT

[ADS73] Roger C. Alperin, R. Keith Dennis, and Michael R. Stein. The non-triviality of sk_{⊥}(ZZπ). Lecture Notes in Mathematics, 353: 1–7, 1973. CODEN LNMAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059258/.

Anderson:1972:TGA

[ADT72] William N. Anderson, Jr., Richard J. Duffin, and George E. Trapp. Tripartite graphs to analyze the interconnection of networks. Lecture Notes in Mathematics, 303:1–12, 1972. CODEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067351/.

Adyan:1974:PGO

[Ady74] S. I. Adyan. Periodic groups of odd exponent. Lecture Notes in Mathematics, 372:8–12, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21571-5_
2/;   http://link.springer.com/chapter/10.1007/BFb0065156/

.

Alling:1971:FTK

[AG71a] Norman L. Alling and Newcomb Greenleaf. Foundations of the Theory of Klein Surfaces, volume 219 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05577-0 (print), 3-540-36885-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 117 pp. LCCN QA3 .L28 no. 219. URL http:/

/link.springer.com/book/10.1007/BFb0060987;   http://
www.springerlink.com/content/978-3-540-36885-4.

Alling:1971:KS

[AG71b] Norman L. Alling and Newcomb Greenleaf. Klein surfaces. Lecture Notes in Mathematics, 219:1–82, 1971. CODEN LN-MAA2. ISBN 3-540-05577-0 (print), 3-540-36885-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060988/.

Alling:1971:RAF

[AG71c] Norman L. Alling and Newcomb Greenleaf. Real-algebraic function fields. Lecture Notes in Mathematics, 219:83–111, 1971. CO-DEN LNMAA2. ISBN 3-540-05577-0 (print), 3-540-36885-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060989/.

Allenby:1973:LER

[AG73a] R. B. J. T. Allenby and R. J. Gregorac. On locally extended residually finite groups. Lecture Notes in Mathematics, 319:9– 17, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058924/.

Ansorge:1973:FAS

[AG73b] R. Ansorge and C. Geiger. Fehlerabschätzungen Bei Aufgaben Mit Schwach Strukturierten Ausgangsdaten. (German) []. Lecture Notes in Mathematics, 333:21–43, 1973. CODEN LN-MAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060685/.

Anselone:1974:CDO

[AG74] P. M. Anselone and A. G. Gibbs. Convergence of the discrete ordinates method for the transport equation. Lecture Notes in Mathematics, 430:1–27, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066263/.

Aggarwal:1974:IFF

[Agg74] Nand Lal Aggarwal. Sur l’information de Fisher. (French) []. Lecture Notes in Mathematics, 398:111–117, 1974. CODEN LN-MAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065772/.

Artin:1972:TTCa

[AGV72] Michael Artin, A. (Alexandre) Grothendieck, and Jean Louis Verdier. Théorie des Topos et Cohomologie Étale des Schémas, volume 269 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05896-6 (print), 3-540-37549-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 518 (est.) pp. LCCN QA3 .L28 no. 269. URL http:/

/link.springer.com/book/10.1007/BFb0081551;   http://
www.springerlink.com/content/978-3-540-37549-4.

Ansorge:1970:AG

[AH70a] R. Ansorge and R. Hass. Anfangswertaufgaben. (German) []. Lecture Notes in Mathematics, 159:1–13, 1970. CODEN LN-MAA2. ISBN 3-540-05187-2 (print), 3-540-36370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060287/.

Ansorge:1970:DGD

[AH70b] R. Ansorge and R. Hass. Differenzapproximationen. (German) [Difference approximations]. Lecture Notes in Mathematics, 159: 14–34, 1970. CODEN LNMAA2. ISBN 3-540-05187-2 (print), 3-540-36370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060288/.

Ansorge:1970:KDL

[AH70c] R. Ansorge and R. Hass. Konvergenz von Differenzenverfahren für lineare und nichtlineare Anfangs- wertaufgaben, volume 159 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05187-2 (print), 3-540-36370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 139 (est.) pp. LCCN QA3 .L28 no. 159. URL http:/

/link.springer.com/book/10.1007/BFb0060286;   http://
www.springerlink.com/content/978-3-540-36370-5.

Ansorge:1970:KDG

[AH70d] R. Ansorge and R. Hass. Konvergenzbegriffe Bei Differenzenverfahren. (German) []. Lecture Notes in Mathematics, 159:35– 56, 1970. CODEN LNMAA2. ISBN 3-540-05187-2 (print), 3-540-36370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060289/.

Ansorge:1970:QAG

[AH70e] R. Ansorge and R. Hass. Quasilineare Anfangswertaufgaben. (German) [Quasilinear initial value problems]. Lecture Notes in Mathematics, 159:108–139, 1970. CODEN LNMAA2. ISBN 3-540-05187-2 (print), 3-540-36370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060292/.

Ansorge:1970:THA

[AH70f] R. Ansorge and R. Hass. Theorie halblinearer anfangswertaufgaben. Lecture Notes in Mathematics, 159:85–107, 1970. CODEN LNMAA2. ISBN 3-540-05187-2 (print), 3-540-36370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060291/.

Ansorge:1970:TLA

[AH70g] R. Ansorge and R. Hass. Theorie Linearer Anfangswertaufgaben. (German) [Theory of linear initial value problems]. Lecture Notes in Mathematics, 159:57–84, 1970. CODEN LNMAA2. ISBN 3-540-05187-2 (print), 3-540-36370-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060290/.

Alo:1974:TGT

[AHiN74] Richard A. Alò, Robert W. Heath, and Jun iti Nagata, editors. TOPO 72 — General Topology and its Applications: Second Pittsburgh International Conference December 18–22, 1972, volume 378 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 378. URL http:/

/link.springer.com/book/10.1007/BFb0068455;   http://
www.springerlink.com/content/978-3-540-38323-9.

Adena:1974:STN

[AHK74] Michael A. Adena, Derek A. Holton, and Patrick A. Kelly. Some thoughts on the no-three-in-line problem. Lecture Notes in Mathematics, 403:6–17, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057371/.

Ahmad:1971:FC

[Ahm71] Shair Ahmad. Flows of characteristic O^{+}. Lecture Notes in Mathematics, 235:1–22, 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059189/.

Alexander:1974:CIH

[AHV74] J. P. Alexander, G. C. Hamrick, and J. W. Vick. A construction for involutions on homotopy spheres. Lecture Notes in Mathematics, 375:1–5, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064005/.

Allgower:1973:ASN

[AJ73] E. L. Allgower and M. M. Jeppson. The approximation of solutions of nonlinear elliptic boundary value problems having several solutions. Lecture Notes in Mathematics, 333:1–20, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060684/.

Altman:1970:CPD

[AK70a] Allen Altman and Steven Kleiman. Completions, primary decomposition and length. Lecture Notes in Mathematics, 146:15– 37, 1970. CODEN LNMAA2. ISBN 3-540-04935-5 (print), 3-540-36309-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060935/.

Altman:1970:C

[AK70b] Allen Altman and Steven Kleiman. Curves. Lecture Notes in Mathematics, 146:163–181, 1970. CODEN LNMAA2. ISBN 3-540-04935-5 (print), 3-540-36309-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060941/.

Altman:1970:DD

[AK70c] Allen Altman and Steven Kleiman. Depth and dimension. Lecture Notes in Mathematics, 146:38–66, 1970. CODEN LN-MAA2. ISBN 3-540-04935-5 (print), 3-540-36309-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060936/.

Altman:1970:DT

[AK70d] Allen Altman and Steven Kleiman. Duality theorems. Lecture Notes in Mathematics, 146:67–81, 1970. CODEN LN-MAA2. ISBN 3-540-04935-5 (print), 3-540-36309-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060937/.

Altman:1970:EMF

[AK70e] Allen Altman and Steven Kleiman. Étale morphisms. (French) []. Lecture Notes in Mathematics, 146:102–127, 1970. CODEN LNMAA2. ISBN 3-540-04935-5 (print), 3-540-36309-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060939/.

Altman:1970:FM

[AK70f] Allen Altman and Steven Kleiman. Flat morphisms. Lecture Notes in Mathematics, 146:82–101, 1970. CODEN LN-MAA2. ISBN 3-540-04935-5 (print), 3-540-36309-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060938/.

Altman:1970:IGD

[AK70g] Allen Altman and Steven Kleiman. Introduction to Grothendieck Duality Theory, volume 146 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04935-5 (print), 3-540-36309-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 185 pp. LCCN QA3 .L28 no. 146.

URL http://link.springer.com/book/10.1007/BFb0060932;
 http://www.springerlink.com/content/978-3-540-36309-
5.

Altman:1970:P

[AK70h] Allen Altman and Steven Kleiman. Preface. Lecture Notes in Mathematics, 146:1–2, 1970. CODEN LNMAA2. ISBN 3-540-04935-5 (print), 3-540-36309-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060933.

Altman:1970:SM

[AK70i] Allen Altman and Steven Kleiman. Smooth morphisms. Lecture Notes in Mathematics, 146:128–162, 1970. CODEN LN-MAA2. ISBN 3-540-04935-5 (print), 3-540-36309-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060940/.

Altman:1970:S

[AK70j] Allen Altman and Steven Kleiman. Study of ω_{X}. Lecture Notes in Mathematics, 146:5–14, 1970. CODEN LNMAA2. ISBN 3-540-04935-5 (print), 3-540-36309-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060934/.

Akemann:1974:NAP

[Ake74] Charles A. Akemann. Non-abelian Pontryagin duality. Lecture Notes in Mathematics, 399:18–21, 1974. CODEN LN-MAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063563/.

Akutowicz:1973:LHD

[Aku73] E. J. Akutowicz, editor. L’Analyse Harmonique dans le Domaine Complexe: Actes de la Table Ronde Internationale du Centre National de la Recherche Scientifique tenue à Montpellier du 11 au 15 septembre 1972, volume 336 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 336. URL http:/

/link.springer.com/book/10.1007/BFb0065780;   http://
www.springerlink.com/content/978-3-540-47170-7.

Anderson:1973:RBH

[AKW73] D. Anderson, M. Karoubi, and J. Wagoner. Relations between higher algebraic K-theories. Lecture Notes in Mathematics, 341: 73–81, 1973. CODEN LNMAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067052/.

Albertson:1974:FIS

[Alb74] Michael O. Albertson. Finding an independent set in a planar graph. Lecture Notes in Mathematics, 406:173–179, 1974. CO-DEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066439/.

Allan:1973:SHA

[All73] Nelo D. Allan. On the structure of Hecke algebras. Lecture Notes in Mathematics, 319:1–8, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058923/.

Alavi:1972:GTA

[ALW72] Y. Alavi, D. R. Lick, and A. T. White, editors. Graph Theory and Applications: Proceedings of the Conference at Western Michigan University, May 10–13, 1972, volume 303 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 303.

URL http://link.springer.com/book/10.1007/BFb0067350;
 http://www.springerlink.com/content/978-3-540-38114-

3. Sponsored jointly by Western Michigan University and the U.S. Army Research Office–Durham, under Grant Number DA-ARO-D-31-124-72-G155.

Alavi:1971:CLC

[AM71] Yousef Alavi and John Mitchem. The connectivity and lineconnectivity of complementary graphs. Lecture Notes in Mathematics, 186:1–3, 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059416/.

Azema:1974:NRD

[AM74] J. Azéma and P.-A. Meyer. Une nouvelle représentation du type de Skorokhod. (French) []. Lecture Notes in Mathematics, 381: 1–10, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057250/.

Amice:1972:CST

[Ami72] Yvette Amice. Conjecture de Schanuel sur la transcendance d’exponentielles (d’après James Ax). (French) []. Lecture Notes in Mathematics, 244:1–10, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058693/.

Andre:1970:AFH

[And70] B. André. Approximation des fonctions holomorphes intégrables sur un ouvert borné de \begin{matrix}c \\ ◯ \end{matrix}. (French) []. Lecture Notes in Mathematics, 164:62–68, 1970. CODEN LNMAA2. ISBN 3-540-05191-0 (print), 3-540-36376-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069605/.

Anderson:1971:CFH

[And71a] D. W. Anderson. Chain functors and homology theories. Lecture Notes in Mathematics, 249:1–12, 1971. CODEN LNMAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060890/.

Andre:1971:HEM

[And71b] Michel André. Hopf and Eilenberg–Maclane algebras. Lecture Notes in Mathematics, 195:1–28, 1971. CODEN LNMAA2. ISBN 3-540-05442-1 (print), 3-540-36548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072305/.

Andrews:1972:STE

[And72] George E. Andrews. Sieves for theorems of Euler, Rogers, and Ramanujan. Lecture Notes in Mathematics, 251:1–20, 1972. CO-DEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058782/.

Anderson:1973:RAT

[And73a] D. W. Anderson. Relationship among K-theories. Lecture Notes in Mathematics, 341:57–72, 1973. CODEN LNMAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067051/.

Anderson:1973:MSF

[And73b] John A. H. Anderson. The maximum sum of a family of ordinals. Lecture Notes in Mathematics, 337:419–438, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066783/.

Anderson:1974:CSI

[And74a] B. A. Anderson. A class of starter induced 1-factorizations. Lecture Notes in Mathematics, 406:180–185, 1974. CODEN LN-MAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066440/.

Anderson:1974:CFS

[And74b] D. W. Anderson. Convergent functors and spectra. Lecture Notes in Mathematics, 418:1–5, 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070636/.

Anderson:1974:RPQ

[And74c] G. D. Anderson. Rotation of plane quasiconformal mappings — a summary. Lecture Notes in Mathematics, 419:1–3, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064705/.

Andersson:1974:WFS

[And74d] Karl G. Andersson. Wave front-sets. Lecture Notes in Mathematics, 384:129–138, 1974. CODEN LNMAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057544/.

Andreotti:1974:CHV

[And74e] Aldo Andreotti. Convexité holomorphe dans les variétés algébriques. (French) []. Lecture Notes in Mathematics, 409:327– 334, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068107/.

Andreotti:1974:PPE

[And74f] Aldo Andreotti. Plongement projectif des espaces pseudoconcaves. (French) []. Lecture Notes in Mathematics, 409:204–211, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068101/.

Andreotti:1974:PLF

[And74g] Aldo Andreotti. Problème de Lewy. (French) []. Lecture Notes in Mathematics, 409:212–222, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068102/.

Anger:1971:ACM

[Ang71] Bernd Anger. Approximation of capacities by measures. Lecture Notes in Mathematics, 226:152–170, 1971. CODEN LN-MAA2. ISBN 3-540-05638-6 (print), 3-540-36912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060483/.

Anonymous:1971:SBVa

[Ano71a] Anonymous, editor. Séminaire Bourbaki vol. 1968/69 Exposés 347-363, volume 179 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 179. URL http:/

/link.springer.com/book/10.1007/BFb0058801;   http://
www.springerlink.com/content/978-3-540-36451-1.

Anonymous:1971:SBVb

[Ano71b] Anonymous, editor. Séminaire Bourbaki vol. 1969/70 Exposés 364–381, volume 180 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 180. URL http:/

/link.springer.com/book/10.1007/BFb0058820;   http://
www.springerlink.com/content/978-3-540-36452-8.

Anonymous:1972:SBV

[Ano72] Anonymous, editor. Séminaire Bourbaki vol. 1970/71 Exposés 382–399, volume 244 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 244. URL http:/

/link.springer.com/book/10.1007/BFb0058692;   http://
www.springerlink.com/content/978-3-540-37094-9.

Anonymous:1974:SBV

[Ano74] Anonymous, editor. Séminaire Bourbaki vol. 1972/73 Exposés 418–435, volume 383 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 383. URL http:/

/link.springer.com/book/10.1007/BFb0057298;   http://
www.springerlink.com/content/978-3-540-38450-2.

Appel:1973:CPK

[App73] K. I. Appel. On the conjugacy problem for knot groups. Lecture Notes in Mathematics, 319:18, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058925.

Aczel:1972:IDA

[AR72] Peter Aczel and Wayne Richter. Inductive definitions and analogues of large cardinals. Lecture Notes in Mathematics, 255: 1–9, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059534/.

Auslander:1973:SEA

[AR73] Maurice Auslander and Idun Reiten. Stable equivalence of Artin algebras. Lecture Notes in Mathematics, 353:8–71, 1973. CODEN LNMAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059259/.

Araki:1971:TDH

[Ara71] Shôrô Araki. On a type of differential Hopf algebras. Lecture Notes in Mathematics, 196:74–84, 1971. CODEN LN-MAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067750/.

Arkowitz:1974:GZT

[Ark74] Martin Arkowitz. The generalized Zabrodsky theorem. Lecture Notes in Mathematics, 418:6–12, 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070637/.

Arnold:1970:PLF

[Arn70] André Arnold. Présentation d’un langage de formalisation des démonstrations mathématiques naturelles. (French) []. Lecture Notes in Mathematics, 125:6–28, 1970. CODEN LNMAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060622/.

Arnold:1973:PIP

[Arn73] Jimmy T. Arnold. Prime ideals in power series rings. Lecture Notes in Mathematics, 311:17–25, 1973. CODEN LN-MAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068915/.

Arscott:1972:PSD

[Ars72] F. M. Arscott. Polynomial solutions of differential equations with bi-orthogonal properties. Lecture Notes in Mathematics, 280:202– 206, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066929/.

Artin:1970:TWC

[Art70] Michael Artin. Théorème de Weil sur la construction d’un groupe à partir d’une loi rationnelle. (French) []. Lecture Notes in Mathematics, 152:632–653, 1970. CODEN LNMAA2. ISBN 3-540-05180-5 (print), 3-540-36337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059016/.

Artzner:1971:FCM

[Art71] Ph. Artzner. Fonctions caractéristiques et mesures planes invariantes par rotations. (French) []. Lecture Notes in Mathematics, 191:1–16, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058842/.

Artin:1972:TTCb

[Art72a] M. Artin. Théorie des Topos et Cohomologie Étale des Schémas: Séminaire de géométrie algébrique du Bois-Marie 1963. Tome 2, volume 270 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06012-X (print), 3-540-37987-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 417 pp. LCCN QA3 .L28 no. 270. URL http:/

/link.springer.com/book/10.1007/BFb0061319;   http://
www.springerlink.com/content/978-3-540-37987-4.

Artzner:1972:ECI

[Art72b] Philippe Artzner. Échantillons et couples indépendants de points aléatoires portés par une surface convexe. (French) []. Lecture Notes in Mathematics, 258:1–34, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059458/.

Artin:1973:CPE

[Art73a] M. Artin. Cohomologie des préschémas excellents d’égales caractéristiques. (French) []. Lecture Notes in Mathematics, 305: 588–638, 1973. CODEN LNMAA2. ISBN 3-540-06118-5 (print), 3-540-38124-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070725/.

Artin:1973:CAC

[Art73b] M. Artin. Comparaison avec la cohomologie classique: Cas d’un schéma lisse. (French) []. Lecture Notes in Mathematics, 305: 64–78, 1973. CODEN LNMAA2. ISBN 3-540-06118-5 (print), 3-540-38124-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070717/.

Artin:1973:DCP

[Art73c] M. Artin. Dimension cohomologique: Premiers résultats. (French) []. Lecture Notes in Mathematics, 305:43–63, 1973. CO-DEN LNMAA2. ISBN 3-540-06118-5 (print), 3-540-38124-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070716/.

Artin:1973:FCC

[Art73d] M. Artin. Faisceaux constructibles cohomologie d’une courbe algébrique. (French) []. Lecture Notes in Mathematics, 305:1– 42, 1973. CODEN LNMAA2. ISBN 3-540-06118-5 (print), 3-540-38124-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070715/.

Artin:1973:MAF

[Art73e] M. Artin. Morphismes acycliques. (French) []. Lecture Notes in Mathematics, 305:168–205, 1973. CODEN LNMAA2. ISBN 3-540-06118-5 (print), 3-540-38124-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070721/.

Artin:1973:TCBc

[Art73f] M. Artin. Théorème de changement de base par un morphisme lisse, et applications. (French) []. Lecture Notes in Mathematics, 305:206–249, 1973. CODEN LNMAA2. ISBN 3-540-06118-5 (print), 3-540-38124-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070722/.

Artin:1973:TCBb

[Art73g] M. Artin. Théorème de changement de base pour un morphisme propre: Fin de la démonstration. (French) []. Lecture Notes in Mathematics, 305:132–144, 1973. CODEN LNMAA2. ISBN 3-540-06118-5 (print), 3-540-38124-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070719/.

Artin:1973:TCBa

[Art73h] M. Artin. Théorème de changement de base pour un morphisme propre. (French) []. Lecture Notes in Mathematics, 305:

79–131, 1973. CODEN LNMAA2. ISBN 3-540-06118-5 (print), 3-540-38124-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070718/.

Artin:1973:TFP

[Art73i] M. Artin. Théorème de finitude pour un morphisme propre; dimension cohomologique des schémas algébriques affines. (French) []. Lecture Notes in Mathematics, 305:145–167, 1973. CODEN LNMAA2. ISBN 3-540-06118-5 (print), 3-540-38124-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070720/.

Artin:1973:TTC

[Art73j] M. Artin. Théorie des Topos et Cohomologie Étale des Schémas: Séminaire de géométrie algébrique du Bois-Marie 1963. Tome 3, volume 305 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06118-5 (print), 3-540-38124-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 640 pp. LCCN QA3 .L28 no. 305. URL http:

//link.springer.com/book/10.1007/BFb0070714;   http:/
/www.springerlink.com/content/978-3-540-38124-2.

Andreotti:1971:FAE

[AS71a] Aldo Andreotti and Wilhelm Stoll. Φ_{ϕ}(x) as a finite algebraic extension of ϕ(∗)Φ(Y). Lecture Notes in Mathematics, 234:184– 207, 1971. CODEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058607/.

Andreotti:1971:AAD

[AS71b] Aldo Andreotti and Wilhelm Stoll. Analytic and Algebraic Dependence of Meromorphic Functions, volume 234 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CO-DEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 390 pp. LCCN QA3 .L28 no. 234. URL http:/

/link.springer.com/book/10.1007/BFb0058595;   http://
www.springerlink.com/content/978-3-540-36995-0.

Andreotti:1971:CEK

[AS71c] Aldo Andreotti and Wilhelm Stoll. A counter example by kas. Lecture Notes in Mathematics, 234:365–384, 1971. CODEN LN-MAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058615/.

Andreotti:1971:D

[AS71d] Aldo Andreotti and Wilhelm Stoll. Dependence. Lecture Notes in Mathematics, 234:83–116, 1971. CODEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058602/.

Andreotti:1971:F

[AS71e] Aldo Andreotti and Wilhelm Stoll. The field. Lecture Notes in Mathematics, 234:127–153, 1971. CODEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058604/.

Andreotti:1971:FHM

[AS71f] Aldo Andreotti and Wilhelm Stoll. Full holomorphic maps. Lecture Notes in Mathematics, 234:219–241, 1971. CODEN LN-MAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058609/.

Andreotti:1971:GL

[AS71g] Aldo Andreotti and Wilhelm Stoll. German letters. Lecture Notes in Mathematics, 234:5, 1971. CODEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058597.

Andreotti:1971:G

[AS71h] Aldo Andreotti and Wilhelm Stoll. Globalization. Lecture Notes in Mathematics, 234:242–265, 1971. CODEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058610/.

Andreotti:1971:I

[AS71i] Aldo Andreotti and Wilhelm Stoll. Introduction. Lecture Notes in Mathematics, 234:6–13, 1971. CODEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058598/.

Andreotti:1971:MF

[AS71j] Aldo Andreotti and Wilhelm Stoll. Meromorphic functions. Lecture Notes in Mathematics, 234:59–82, 1971. CODEN LN-MAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058601/.

Andreotti:1971:Pa

[AS71k] Aldo Andreotti and Wilhelm Stoll. Preface. Lecture Notes in Mathematics, 234:1–3, 1971. CODEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058596/.

Andreotti:1971:Pb

[AS71l] Aldo Andreotti and Wilhelm Stoll. Preparations. Lecture Notes in Mathematics, 234:290–298, 1971. CODEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058613/.

Andreotti:1971:PR

[AS71m] Aldo Andreotti and Wilhelm Stoll. Product representations. Lecture Notes in Mathematics, 234:43–58, 1971. CODEN LN-MAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058600/.

Andreotti:1971:PLH

[AS71n] Aldo Andreotti and Wilhelm Stoll. Proper, light, holomorphic maps. Lecture Notes in Mathematics, 234:117–126, 1971. CO-DEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058603/.

Andreotti:1971:PM

[AS71o] Aldo Andreotti and Wilhelm Stoll. Pseudoconcave maps. Lecture Notes in Mathematics, 234:299–364, 1971. CODEN LN-MAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058614/.

Andreotti:1971:QPMa

[AS71p] Aldo Andreotti and Wilhelm Stoll. Quasi-proper maps. Lecture Notes in Mathematics, 234:166–183, 1971. CODEN LN-MAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058606/.

Andreotti:1971:QPMb

[AS71q] Aldo Andreotti and Wilhelm Stoll. Quasi-proper maps of codimension k. Lecture Notes in Mathematics, 234:208–218, 1971. CODEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058608/.

Andreotti:1971:RHM

[AS71r] Aldo Andreotti and Wilhelm Stoll. The rank of a holomorphic map. Lecture Notes in Mathematics, 234:14–42, 1971. CODEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058599/.

Andreotti:1971:SL

[AS71s] Aldo Andreotti and Wilhelm Stoll. The Schwarz lemma. Lecture Notes in Mathematics, 234:266–280, 1971. CODEN LN-MAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058611/.

Andreotti:1971:SML

[AS71t] Aldo Andreotti and Wilhelm Stoll. Sections in meromorphic line bundles. Lecture Notes in Mathematics, 234:281–289, 1971. CO-DEN LNMAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058612/.

Andreotti:1971:SPM

[AS71u] Aldo Andreotti and Wilhelm Stoll. Semi-proper maps. Lecture Notes in Mathematics, 234:154–165, 1971. CODEN LN-MAA2. ISBN 3-540-05670-X (print), 3-540-36995-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058605/.

Abhyankar:1974:AGF

[AS74a] Prof. Dr. Shreeram Shankar Abhyankar and Prof. Dr. Avinash Madhav Sathaye. Affine geometry or filtered domains. Lecture Notes in Mathematics, 423:234–256, 1974. CODEN LN-MAA2. ISBN 3-540-06969-0 (print), 3-540-37280-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066170/.

Abhyankar:1974:BGG

[AS74b] Prof. Dr. Shreeram Shankar Abhyankar and Prof. Dr. Avinash Madhav Sathaye. Birational geometry or genus. Lecture Notes in Mathematics, 423:180–233, 1974. CODEN LN-MAA2. ISBN 3-540-06969-0 (print), 3-540-37280-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066169/.

Abhyankar:1974:GTA

[AS74c] Prof. Dr. Shreeram Shankar Abhyankar and Prof. Dr. Avinash Madhav Sathaye. Geometric Theory of Algebraic Space Curves, volume 423 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06969-0 (print), 3-540-37280-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 256 (est.) pp. LCCN QA3 .L28 no. 423. URL http:/

/link.springer.com/book/10.1007/BFb0066166;   http://
www.springerlink.com/content/978-3-540-37280-6.

Abhyankar:1974:LGL

[AS74d] Prof. Dr. Shreeram Shankar Abhyankar and Prof. Dr. Avinash Madhav Sathaye. Local geometry of length. Lecture Notes in Mathematics, 423:1–65, 1974. CODEN LNMAA2. ISBN 3-540-06969-0 (print), 3-540-37280-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066167/.

Abhyankar:1974:PGH

[AS74e] Prof. Dr. Shreeram Shankar Abhyankar and Prof. Dr. Avinash Madhav Sathaye. Projective geometry or homogeneous domains. Lecture Notes in Mathematics, 423:66–179, 1974. CODEN LNMAA2. ISBN 3-540-06969-0 (print), 3-540-37280-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066168/.

Andreotti:1974:MFC

[AS74f] Aldo Andreotti and Wilhelm Stoll. Meromorphic functions on complex spaces. Lecture Notes in Mathematics, 409:279–309, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068105/.

Assouad:1971:DCC

[Ass71] Patrice Assouad. Démonstration de la ‘Conjecture de Chung’ par Carleson. (French) []. Lecture Notes in Mathematics, 191: 17–20, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058843/.

Aastrom:1972:SI

[Åst72] K. J. Åström. System identification. Lecture Notes in Mathematics, 294:35–55, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064933/.

Applegate:1970:IC

[AT70] H. Applegate and M. Tierney. Iterated cotriples. Lecture Notes in Mathematics, 137:56–99, 1970. CODEN LNMAA2. ISBN 3-540-04926-6 (print), 3-540-36292-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060440/.

Ansorge:1972:NLN

[AT72] R. Ansorge and W. Törnig, editors. Numerische Lösung nichtlinearer partieller Differential- und Integrodifferentialgleichungen: Vorträge einer Tagung im Mathematischen Forschungsinstitut Oberwolfach, 28. 11.-4. 12. 1971, volume 267 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 267.

URL http://link.springer.com/book/10.1007/BFb0061609;
 http://www.springerlink.com/content/978-3-540-37540-
1.

Ansorge:1973:NIA

[AT73] R. Ansorge and W. Törnig, editors. Numerische, insbesondere approximationstheoretische Behandlung von Funktionalgleichungen: Vorträge einer Tagung im Mathematischen Forschungsinstitut Oberwolfach, 4.–8. 12. 1972, volume 333 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 333.

URL http://link.springer.com/book/10.1007/BFb0060683;
 http://www.springerlink.com/content/978-3-540-46986-
5.

Ansorge:1974:NBN

[AT74] R. Ansorge and W. Törnig, editors. Numerische Behandlung nichtlinearer Integrodifferential-und Differentialgleichungen: Vorträge einer Tagung im Mathematischen Forschungsinstitut Oberwolfach, 2. 12.–7. 12. 1973, volume 395 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 395.

URL http://link.springer.com/book/10.1007/BFb0060660;
 http://www.springerlink.com/content/978-3-540-37771-
9.

Atiyah:1974:EOC

[Ati74a] Michael Francis Atiyah. Elliptic Operators and Compact Groups, volume 401 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06855-4 (print), 3-540-37811-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 93 pp. LCCN QA3 .L28 no. 401. URL http:/

/link.springer.com/book/10.1007/BFb0057821;   http://
www.springerlink.com/content/978-3-540-37811-2.

Atiyah:1974:LTE

[Ati74b] Michael Francis Atiyah. Lecture 1: Transversally elliptic operators. Lecture Notes in Mathematics, 401:3–8, 1974. CODEN LNMAA2. ISBN 3-540-06855-4 (print), 3-540-37811-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057822/.

Atiyah:1974:LA

[Ati74c] Michael Francis Atiyah. Lecture 10: Applications. Lecture Notes in Mathematics, 401:81–93, 1974. CODEN LNMAA2. ISBN 3-540-06855-4 (print), 3-540-37811-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057831/.

Atiyah:1974:LIT

[Ati74d] Michael Francis Atiyah. Lecture 2: The index of transversally elliptic operators. Lecture Notes in Mathematics, 401:9–17, 1974. CODEN LNMAA2. ISBN 3-540-06855-4 (print), 3-540-37811-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057823/.

Atiyah:1974:LEM

[Ati74e] Michael Francis Atiyah. Lecture 3: The excision and multiplicative properties. Lecture Notes in Mathematics, 401:18–26, 1974. CODEN LNMAA2. ISBN 3-540-06855-4 (print), 3-540-37811-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057824/.

Atiyah:1974:LNI

[Ati74f] Michael Francis Atiyah. Lecture 4: The naturality of the index and the localization theorem. Lecture Notes in Mathematics, 401: 27–35, 1974. CODEN LNMAA2. ISBN 3-540-06855-4 (print), 3-540-37811-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057825/.

Atiyah:1974:LIHa

[Ati74g] Michael Francis Atiyah. Lecture 5: The index homomorphism for G = S^{1}. Lecture Notes in Mathematics, 401:36–43, 1974. CODEN LNMAA2. ISBN 3-540-06855-4 (print), 3-540-37811-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057826/.

Atiyah:1974:LO

[Ati74h] Michael Francis Atiyah. Lecture 6: The operators ¯∂^{±} for G = S^{1}. Lecture Notes in Mathematics, 401:44–54, 1974. CODEN LNMAA2. ISBN 3-540-06855-4 (print), 3-540-37811-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057827/.

Atiyah:1974:LTA

[Ati74i] Michael Francis Atiyah. Lecture 7: Toral actions with finite isotropy groups. Lecture Notes in Mathematics, 401:55–64, 1974. CODEN LNMAA2. ISBN 3-540-06855-4 (print), 3-540-37811-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057828/.

Atiyah:1974:LIHb

[Ati74j] Michael Francis Atiyah. Lecture 8: The index homomorphism for G = T^{n}. Lecture Notes in Mathematics, 401:65–68, 1974. CODEN LNMAA2. ISBN 3-540-06855-4 (print), 3-540-37811-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057829/.

Atiyah:1974:LCF

[Ati74k] Michael Francis Atiyah. Lecture 9: The cohomology formula. Lecture Notes in Mathematics, 401:69–80, 1974. CODEN LN-MAA2. ISBN 3-540-06855-4 (print), 3-540-37811-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057830/.

Auchmuty:1973:LME

[Auc73] J. F. G. Auchmuty. Lyapunov methods and equations of parabolic type. Lecture Notes in Mathematics, 322:1–14, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060559/.

Auchmuty:1974:NSE

[Auc74] J. F. G. Auchmuty. The numerical solution of the equations for rotating stars. Lecture Notes in Mathematics, 430:28–60, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066264/.

Aull:1974:SPI

[Aul74] C. E. Aull. Some properties involving base axioms and metrizability. Lecture Notes in Mathematics, 378:41–45, 1974. CODEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068458/.

Aumasson:1970:QRG

[Aum70] Claude Aumasson. Quelques résultats généraux sur la programmation dynamique stochastique à temps discret. (French) []. Lecture Notes in Mathematics, 132:1–36, 1970. CODEN LN-MAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066670/.

Aurich:1974:SEH

[Aur74] Volker Aurich. The spectrum as envelope of holomorphy of a domain over an arbitrary product of complex lines. Lecture Notes in Mathematics, 364:109–122, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069010/.

Auslander:1970:SCS

[Aus70a] Joseph Auslander. On stability of closed sets in dynamical systems. Lecture Notes in Mathematics, 144:1–4, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059913/.

Auslender:1970:RPS

[Aus70b] A. Auslender. Recherche des points de selle d’une fonction. (French) []. Lecture Notes in Mathematics, 132:37–52, 1970. CO-DEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066671/.

Auslander:1971:SHM

[Aus71] Joseph Auslander. Structure and homomorphisms of minimal sets. Lecture Notes in Mathematics, 235:23–30, 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059190/.

Auslander:1973:NCD

[Aus73a] Joseph Auslander. Non-compact dynamical systems. Lecture Notes in Mathematics, 318:6–11, 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061718/.

Auslander:1973:EIF

[Aus73b] Louis Auslander. Ergodic G-induced flows on compact solvmanifolds. Lecture Notes in Mathematics, 318:12–22, 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061719/.

Authier:1972:CVR

[Aut72] M. Authier. Champs de vecteurs et régularité des solutions de certains problèmes de Dirichlet variationnels. (French) []. Lecture Notes in Mathematics, 280:207–211, 1972. CODEN LN-MAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066930/.

Azencott:1970:CEPb

[Aze70a] Robert Azencott. Caractérisation de l’espace de Poisson. (French) []. Lecture Notes in Mathematics, 148:28–54, 1970. CO-DEN LNMAA2. ISBN 3-540-04937-1 (print), 3-540-36311-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059354/.

Azencott:1970:CGS

[Aze70b] Robert Azencott. Cas des groupes semi-simples. (French) []. Lecture Notes in Mathematics, 148:55–70, 1970. CODEN LN-MAA2. ISBN 3-540-04937-1 (print), 3-540-36311-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059355/.

Azencott:1970:CEPa

[Aze70c] Robert Azencott. Construction de l’espace de Poisson. (French) []. Lecture Notes in Mathematics, 148:1–27, 1970. CODEN LNMAA2. ISBN 3-540-04937-1 (print), 3-540-36311-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059353/.

Azencott:1970:EPG

[Aze70d] Robert Azencott. Espaces de Poisson des Groupes Localement Compacts, volume 148 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04937-1 (print), 3-540-36311-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 138 (est.) pp. LCCN QA3 .L28 no. 148.

URL http://link.springer.com/book/10.1007/BFb0059352;
 http://www.springerlink.com/content/978-3-540-36311-
8.

Azencott:1970:GTF

[Aze70e] Robert Azencott. Groupes de type (T). (French) []. Lecture Notes in Mathematics, 148:100–138, 1970. CODEN LNMAA2. ISBN 3-540-04937-1 (print), 3-540-36311-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059357/.

Azencott:1970:PFH

[Aze70f] Robert Azencott. Périodes des fonctions μ-harmoniques. (French) [Periods of μ-harmonic functions]. Lecture Notes in Mathematics, 148:71–99, 1970. CODEN LNMAA2. ISBN 3-540-04937-1 (print), 3-540-36311-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059356/.

Azema:1972:RTR

[Azé72] J. Azéma. Une remarque sur les temps de retour trois applications. (French) []. Lecture Notes in Mathematics, 258:35–50, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059459/.

Azra:1972:RDS

[Azr72] Jean-Pierre Azra. Relations diophantiennes et la solution négative du 10e Problème de Hilbert (d’après M. Davis, H. Putnam, J. Robinson et I. Matiasevitch). (French) []. Lecture Notes in Mathematics, 244:11–28, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058694/.

Azumaya:1973:SPT

[Azu73] Gorô Azumaya. Some properties of TTF-classes. Lecture Notes in Mathematics, 353:72–83, 1973. CODEN LNMAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059260/.

Baba:1973:EHD

[Bab73] Yoshikazu Baba. Entropy and Hausdorff dimension of a sequence of coordinate functions in base-r expansion. Lecture Notes in Mathematics, 330:1–6, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061477/.

Bacsich:1972:CIN

[Bac72] Paul D. Bacsich. Compact injectives and non-standard analysis. Lecture Notes in Mathematics, 255:10–28, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059535/.

Bachmuth:1973:EPV

[Bac73] S. Bachmuth. Exceptional primes in varieties. Lecture Notes in Mathematics, 319:19–25, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058926/.

Badrikian:1970:AR

[Bad70a] A. Badrikian. Applications radonifiantes. Lecture Notes in Mathematics, 139:196–219, 1970. CODEN LNMAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067905/.

Badrikian:1970:LPM

[Bad70b] A. Badrikian. Limites projectives de mesures de Radon. (French) []. Lecture Notes in Mathematics, 139:95–111, 1970. CODEN LNMAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067900/.

Badrikian:1970:MCG

[Bad70c] A. Badrikian. Mesures cylindriques gaussiennes. (French) []. Lecture Notes in Mathematics, 139:180–195, 1970. CODEN LN-MAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067904/.

Badrikian:1970:OMR

[Bad70d] A. Badrikian. Opérations sur les mesures de Radon. (French) []. Lecture Notes in Mathematics, 139:83–94, 1970. CODEN LNMAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067899/.

Badrikian:1970:TBF

[Bad70e] Albert Badrikian. Le théorème de Bochner. (French) []. Lecture Notes in Mathematics, 139:37–47, 1970. CODEN LN-MAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067896/.

Badrikian:1970:PSL

[Bad70f] Albert Badrikian. Processus stochastiques linéaires et mesures cylindriques. (French) []. Lecture Notes in Mathematics, 139: 1–15, 1970. CODEN LNMAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067894/.

Badrikian:1970:SFA

[Bad70g] Albert Badrikian. Séminaire sur les Fonctions Aléatoires Linéaires et les Mesures Cylindriques, volume 139 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 221 pp. LCCN QA3 .L28 no. 139. URL http:/

/link.springer.com/book/10.1007/BFb0067893;   http://
www.springerlink.com/content/978-3-540-36296-8.

Badrikian:1974:AMC

[Bad74a] Albert Badrikian. Application à la mesure cylindrique Gaussienne normale sur un Hilbert. (French) []. Lecture Notes in Mathematics, 379:179–221, 1974. CODEN LNMAA2. ISBN 3-540-06843-0 (print), 3-540-37796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060499/.

Badrikian:1974:CCC

[Bad74b] Albert Badrikian. Conditions de concentration cylindrique. (French) []. Lecture Notes in Mathematics, 379:90–136, 1974. CODEN LNMAA2. ISBN 3-540-06843-0 (print), 3-540-37796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060497/.

Badrikian:1974:CGC

[Bad74c] Albert Badrikian. Conditions géométriques de concentration cylindrique. (French) []. Lecture Notes in Mathematics, 379:35– 89, 1974. CODEN LNMAA2. ISBN 3-540-06843-0 (print), 3-540-37796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060496/.

Badrikian:1974:MWF

[Bad74d] Albert Badrikian. Mesure de Wiener. (French) []. Lecture Notes in Mathematics, 379:1–11, 1974. CODEN LNMAA2. ISBN 3-540-06843-0 (print), 3-540-37796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060494/.

Badrikian:1974:MCD

[Bad74e] Albert Badrikian. Mesures cylindriques dont l’image est une mesure de Radon. (French) []. Lecture Notes in Mathematics, 379: 12–34, 1974. CODEN LNMAA2. ISBN 3-540-06843-0 (print), 3-540-37796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060495/.

Badrikian:1974:TPL

[Bad74f] Albert Badrikian. Trajectoires des processus linéaires associés aux mesures de Radon. (French) []. Lecture Notes in Mathematics, 379:137–178, 1974. CODEN LNMAA2. ISBN 3-540-06843-0 (print), 3-540-37796-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060498/.

Baer:1974:ENS

[Bae74] Reinhold Baer. Einbettungseigenschaften von Normalteilern: Der Schluss vom Endlichen Aufs Unendliche. (German) []. Lecture Notes in Mathematics, 372:13–62, 1974. CO-DEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (ebook), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21571-5_3/;   http://link.springer.
com/chapter/10.1007/BFb0065157/.

Baily:1970:AFI

[Bai70] Walter L. Baily, Jr. Automorphic forms with integral Fourier coefficients. Lecture Notes in Mathematics, 155:1–8, 1970. CODEN LNMAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060313/.

Bajaj:1973:PSD

[Baj73] Prem N. Bajaj. Products of semi-dynamical systems. Lecture Notes in Mathematics, 318:23–29, 1973. CODEN LN-MAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061720/.

Balakrishnan:1970:NCT

[Bal70] A. V. Balakrishnan. On a new computing technique in optimal control and its application to minimal time flight profile optimization. Lecture Notes in Mathematics, 112:1–20, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060194/.

Bank:1971:CSA

[Ban71] Steven Bank. On certain solutions of algebraic differential equations. Lecture Notes in Mathematics, 183:165–169, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060416/.

Barr:1970:RA

[Bar70] Michael Barr. Relational algebras. Lecture Notes in Mathematics, 137:39–55, 1970. CODEN LNMAA2. ISBN 3-540-04926-6 (print), 3-540-36292-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060439/.

Bari:1971:MGR

[Bar71a] Ruth A. Bari. Maximal m-gons in 4-regular major maps. Lecture Notes in Mathematics, 186:5–8, 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059417/.

Barr:1971:EC

[Bar71b] Michael Barr. Exact categories. Lecture Notes in Mathematics, 236:1–120, 1971. CODEN LNMAA2. ISBN 3-540-05678-5 (print), 3-540-36999-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058580/.

Barr:1971:NAF

[Bar71c] Michael Barr. Non-abelian full embedding; announcement of results. Lecture Notes in Mathematics, 195:205–208, 1971. CODEN LNMAA2. ISBN 3-540-05442-1 (print), 3-540-36548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072312/.

Bari:1972:MRM

[Bar72] Ruth A. Bari. Minimal regular major maps with proper 4-rings. Lecture Notes in Mathematics, 303:13–20, 1972. CODEN LN-MAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067352/.

Bari:1974:CEG

[Bar74a] Ruth A. Bari. Chromatically equivalent graphs. Lecture Notes in Mathematics, 406:186–200, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066441/.

Barlet:1974:ECC

[Bar74b] Daniel Barlet. Espace des cycles et d′d′′-cohomologie de P_{n} −P_{k}. (French) []. Lecture Notes in Mathematics, 409:98–123, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068095/.

Barnes:1974:CGG

[Bar74c] D. W. Barnes. Characterisation of the groups with the Gaschütz cohomology property. Lecture Notes in Mathematics, 372:63–64, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
662-21571-5_4;   http://link.springer.com/chapter/10.
1007/BFb0065158.

Bass:1972:CGD

[Bas72] Hyman Bass. K_{2} des corps globaux [d’après J. Tate, H. Garland, …]. (French) []. Lecture Notes in Mathematics, 244:233– 255, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058705/.

Bass:1973:CAE

[Bas73a] H. Bass, editor. “Classical” Algebraic K-Theory, and Connections with Arithmetic: Proceedings of the Conference held at the Seattle Research Center of the Battelle Memorial Institute, from August 28 to September 8, 1972, volume 342 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 342.

URL http://link.springer.com/book/10.1007/BFb0073717;
 http://www.springerlink.com/content/978-3-540-37770-
2.

Bass:1973:HEG

[Bas73b] H. Bass, editor. Hermitian K-Theory and Geometric Applications: Proceedings of the Conference held at the Seattle Research Center of the Battelle Memorial Institute, from August 28 to September 8, 1972, volume 343 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 343. URL http:/

/link.springer.com/book/10.1007/BFb0061366;   http://
www.springerlink.com/content/978-3-540-37773-3.

Bass:1973:HEP

[Bas73c] H. Bass, editor. Higher K-Theories: Proceedings of the Conference held at the Seattle Research Center of the Battelle Memorial Institute, from August 28 to September 8, 1972, volume 341 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 341. URL http:/

/link.springer.com/book/10.1007/BFb0067048;   http://
www.springerlink.com/content/978-3-540-37767-2.

Bass:1973:SPC

[Bas73d] Hyman Bass. Some problems in ‘classical’ algebraic K-theory. Lecture Notes in Mathematics, 342:1–73, 1973. CODEN LN-MAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073718/.

Bass:1973:UAT

[Bas73e] Hyman Bass. Unitary algebraic K-theory. Lecture Notes in Mathematics, 343:57–265, 1973. CODEN LNMAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061370/.

Bass:1973:FAG

[Bas73f] Professor Hyman Bass. Is SK_{1}(ZΠ) = 0 for Π a finite abelian group. Lecture Notes in Mathematics, 353:84, 1973. CODEN LNMAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059261.

Bauer:1971:SPT

[Bau71a] Heinz Bauer. Seminar on Potential Theory, II, volume 226 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05638-6 (print), 3-540-36912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 170 pp. LCCN QA3 .L28 no. 226. URL http:/

/link.springer.com/book/10.1007/BFb0060478;   http://
www.springerlink.com/content/978-3-540-36912-7.

Baumert:1971:CDS

[Bau71b] Leonard D. Baumert. Cyclic Difference Sets, volume 182 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05368-9 (print), 3-540-36453-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 166 pp. LCCN QA3 .L28 no. 182. URL http:/

/link.springer.com/book/10.1007/BFb0061260;   http://
www.springerlink.com/content/978-3-540-36453-5.

Baumert:1971:DSS

[Bau71c] Leonard D. Baumert. Difference sets of special type. Lecture Notes in Mathematics, 182:77–98, 1971. CODEN LN-MAA2. ISBN 3-540-05368-9 (print), 3-540-36453-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061264/.

Baumert:1971:EQ

[Bau71d] Leonard D. Baumert. Existence questions. Lecture Notes in Mathematics, 182:11–53, 1971. CODEN LNMAA2. ISBN 3-540-05368-9 (print), 3-540-36453-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061262/.

Baumert:1971:FDS

[Bau71e] Leonard D. Baumert. Families of difference sets. Lecture Notes in Mathematics, 182:99–142, 1971. CODEN LNMAA2. ISBN 3-540-05368-9 (print), 3-540-36453-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061265/.

Baumert:1971:I

[Bau71f] Leonard D. Baumert. Introduction. Lecture Notes in Mathematics, 182:1–10, 1971. CODEN LNMAA2. ISBN 3-540-05368-9 (print), 3-540-36453-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061261/.

Baumert:1971:M

[Bau71g] Leonard D. Baumert. Miscellany. Lecture Notes in Mathematics, 182:143–158, 1971. CODEN LNMAA2. ISBN 3-540-05368-9 (print), 3-540-36453-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061266/.

Baumert:1971:MCE

[Bau71h] Leonard D. Baumert. Multipliers and constructive existence tests. Lecture Notes in Mathematics, 182:54–76, 1971. CODEN LNMAA2. ISBN 3-540-05368-9 (print), 3-540-36453-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061263/.

Baumslag:1974:FPM

[Bau74a] Gilbert Baumslag. Finitely presented metabelian groups. Lecture Notes in Mathematics, 372:65–74, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (ebook), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21571-5_5/;   http://link.springer.
com/chapter/10.1007/BFb0065159/.

Baumslag:1974:SPO

[Bau74b] Gilbert Baumslag. Some problems on one-relator groups. Lecture Notes in Mathematics, 372:75–81, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (ebook), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21571-5_6/;   http://link.springer.
com/chapter/10.1007/BFb0065160/.

Baxley:1974:SPI

[Bax74] John V. Baxley. Singularly perturbed initial value problems. Lecture Notes in Mathematics, 415:15–22, 1974. CODEN LN-MAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065507/.

Bazley:1974:EBL

[Baz74] N. W. Bazley. Existence and bounds for the lowest critical energy of the Hartree operator. Lecture Notes in Mathematics, 415:

23–34, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065508/.

Badrikian:1970:TTB

[BB70] A. Badrikian and P. Bernard. Les théorèmes de type Bochner et P. Lévy. (French) []. Lecture Notes in Mathematics, 139:152– 179, 1970. CODEN LNMAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067903/.

Berard-Bergery:1973:LGF

[BB73] Lionel Bérard-Bergery. Laplacien et géodésiques fermées sur les formes d’espace hyperbolique compactes. (French) []. Lecture Notes in Mathematics, 317:107–122, 1973. CODEN LN-MAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069279/.

Badrikian:1970:TFM

[BBB70] Monsieur A. Badrikian, B. Brunet, and P. Bernard. Transformées de Fourier des mesures cylindriques. (French) []. Lecture Notes in Mathematics, 139:16–36, 1970. CODEN LNMAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067895/.

Boreham:1974:UFB

[BBF74] T. G. Boreham, I. Z. Bouwer, and R. W. Frucht. A useful family of bicubic graphs. Lecture Notes in Mathematics, 406:213– 225, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066443/.

Badrikian:1974:MCE

[BC74a] Albert Badrikian and Simone Chevet. Mesures Cylindriques, Espaces de Wiener et Fonctions Aléatoires Gaussiennes, volume 379 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06843-0 (print), 3-540-37796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 383 pp. LCCN QA3 .L28 no. 379. URL http:

//link.springer.com/book/10.1007/BFb0060493;   http:/
/www.springerlink.com/content/978-3-540-37796-2.

Bryce:1974:SPC

[BC74b] R. A. Bryce and John Cossey. Subdirect product closed fitting classes. Lecture Notes in Mathematics, 372:158–164, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-662-21571-5_13/;   http://link.
springer.com/chapter/10.1007/BFb0065167/.

Borel:1970:SAG

[BCC^{+}70] A. Borel, R. Carter, C. W. Curtis, N. Iwahori, T. A. Springer, and R. Steinberg, editors. Seminar on Algebraic Groups and Related Finite Groups: Held at The Institute for Advanced Study, Princeton/NJ, 1968/69, volume 131 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04920-7 (print), 3-540-36272-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 131. URL http:/

/link.springer.com/book/10.1007/BFb0081541;   http://
www.springerlink.com/content/978-3-540-36272-2.

Balakrishnan:1970:SOH

[BCdV^{+}70] A. V. Balakrishnan, M. Contensou, B. F. de Veubeke, P. Krée, J.-L. Lions, and N. N. Moiseev, editors. Symposium on Optimization: Held in Nice, June 29th–July 5th, 1969, volume 132 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 132. URL http:/

/link.springer.com/book/10.1007/BFb0066669;   http://
www.springerlink.com/content/978-3-540-36275-3.

Bretagnolle:1973:EEP

[BCM73] J. L. Bretagnolle, S. D. Chatterji, and P.-A. Meyer. École d’Été de Probabilités: Processus Stochastiques. (French) [], volume 307 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06126-6 (print), 3-540-38321-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 198 (est.) pp. LCCN QA3 .L28 no. 307. URL http:/

/link.springer.com/book/10.1007/BFb0059706;   http://
www.springerlink.com/content/978-3-540-38321-5.

Boesch:1974:CPG

[BCM74] F. T. Boesch, S. Chen, and J. A. M. McHugh. On covering the points of a graph with point disjoint paths. Lecture Notes in Mathematics, 406:201–212, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066442/.

Berenstein:1972:AUS

[BD72a] Carlos A. Berenstein and Milos A. Dostal. Analytically Uniform Spaces and their Applications to Convolution Equations, volume 256 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05746-3 (print), 3-540-37163-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 130 pp. LCCN QA3 .L28 no. 256. URL http:

//link.springer.com/book/10.1007/BFb0070913;   http:/
/www.springerlink.com/content/978-3-540-37163-2.

Berenstein:1972:DBP

[BD72b] Carlos A. Berenstein and Milos A. Dostal. Definition and basic properties of analytically uniform spaces. Lecture Notes in Mathematics, 256:1–25, 1972. CODEN LNMAA2. ISBN 3-540-05746-3 (print), 3-540-37163-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070914/.

Berenstein:1972:EAS

[BD72c] Carlos A. Berenstein and Milos A. Dostal. Examples of AUspaces. Lecture Notes in Mathematics, 256:26–68, 1972. CODEN LNMAA2. ISBN 3-540-05746-3 (print), 3-540-37163-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070915/.

Berenstein:1972:FP

[BD72d] Carlos A. Berenstein and Milos A. Dostal. The fundamental principle. Lecture Notes in Mathematics, 256:90–121, 1972. CODEN LNMAA2. ISBN 3-540-05746-3 (print), 3-540-37163-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070917/.

Berenstein:1972:SAS

[BD72e] Carlos A. Berenstein and Milos A. Dostal. Spaces of approximate solutions to certain convolution equations. Lecture Notes in Mathematics, 256:69–89, 1972. CODEN LNMAA2. ISBN 3-540-05746-3 (print), 3-540-37163-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070916/.

Bhargava:1972:NMM

[BD72f] S. Bhargava and R. J. Duffin. Network models for maximization of heat transfer under weight constraints. Lecture Notes in Mathematics, 303:41–42, 1972. CODEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067355.

Berenstein:1973:CE

[BD73] C. A. Berenstein and M. A. Dostal. On convolution equations I. Lecture Notes in Mathematics, 336:79–94, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065789/.

Bartels:1974:CGA

[BD74] Richard Bartels and James W. Daniel. A conjugate gradient approach to nonlinear elliptic boundary value problems in irregular regions. Lecture Notes in Mathematics, 363:1–11, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069120/.

Brown:1973:UEM

[BDF73] L. G. Brown, R. G. Douglas, and P. A. Fillmore. Unitary equivalence modulo the compact operators and extensions of C∗- algebras. Lecture Notes in Mathematics, 345:58–128, 1973. CO-DEN LNMAA2. ISBN 3-540-06496-6 (print), 3-540-37808-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058917/.

Braaksma:1974:GTO

[BdS74] B. L. J. Braaksma and H. S. V. de Snoo. Generalized translation operators associated with a singular differential operator. Lecture Notes in Mathematics, 415:62–77, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065512/.

Buchsbaum:1973:PLA

[BE73] David A. Buchsbaum and David Eisenbud. On a problem in linear algebra. Lecture Notes in Mathematics, 311:50–56, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068918/.

Bonnet:1974:CII

[BE74] R. Bonnet and P. Erdős. The chromatic index of an infinite complete hypergraph: a partition theorem. Lecture Notes in Mathematics, 411:54–60, 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066178/.

Bear:1970:CPM

[Bea70a] Herbert S. Bear. Completeness of the part metric. Lecture Notes in Mathematics, 121:38–41, 1970. CODEN LNMAA2. ISBN 3-540-04910-X (print), 3-540-36255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071125/.

Bear:1970:EGP

[Bea70b] Herbert S. Bear. Examples of Gleason parts. Lecture Notes in Mathematics, 121:2–4, 1970. CODEN LNMAA2. ISBN 3-540-04910-X (print), 3-540-36255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071117/.

Bear:1970:GDP

[Bea70c] Herbert S. Bear. Geometric description of parts in T_{B}. Lecture Notes in Mathematics, 121:9–15, 1970. CODEN LNMAA2. ISBN 3-540-04910-X (print), 3-540-36255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071119/.

Bear:1970:GPC

[Bea70d] Herbert S. Bear. Geometric properties of d and a continuous selection theorem. Lecture Notes in Mathematics, 121:36–38, 1970. CODEN LNMAA2. ISBN 3-540-04910-X (print), 3-540-36255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071124/.

Bear:1970:IPT

[Bea70e] Herbert S. Bear. Inner parts and their associated normed spaces. Lecture Notes in Mathematics, 121:18–24, 1970. CODEN LN-MAA2. ISBN 3-540-04910-X (print), 3-540-36255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071121/.

Bear:1970:IK

[Bea70f] Herbert S. Bear. Integral kernels. Lecture Notes in Mathematics, 121:30–35, 1970. CODEN LNMAA2. ISBN 3-540-04910-X (print), 3-540-36255-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071123/.

Bear:1970:LGP

[Bea70g] Herbert S. Bear. Lectures on Gleason Parts, volume 121 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CO-DEN LNMAA2. ISBN 3-540-04910-X (print), 3-540-36255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 47 pp. LCCN QA3 .L28 no. 121. URL http:/

/link.springer.com/book/10.1007/BFb0071115;   http://
www.springerlink.com/content/978-3-540-36255-5.

Bear:1970:LFD

[Bea70h] Herbert S. Bear. Linear functionals as differences of positive functionals. Lecture Notes in Mathematics, 121:41–46, 1970. CODEN LNMAA2. ISBN 3-540-04910-X (print), 3-540-36255-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071126/.

Bear:1970:PI

[Bea70i] Herbert S. Bear. Proof of the d − σ − G identity. Lecture Notes in Mathematics, 121:4–9, 1970. CODEN LNMAA2. ISBN 3-540-04910-X (print), 3-540-36255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071118/.

Bear:1970:RMP

[Bea70j] Herbert S. Bear. Representing measures and parts in T_{C}. Lecture Notes in Mathematics, 121:15–18, 1970. CODEN LN-MAA2. ISBN 3-540-04910-X (print), 3-540-36255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071120/.

Bear:1970:SMA

[Bea70k] Herbert S. Bear. Selection of mutually absolutely continuous representing measures. Lecture Notes in Mathematics, 121:24– 29, 1970. CODEN LNMAA2. ISBN 3-540-04910-X (print), 3-540-36255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071122/.

Bear:1970:SAN

[Bea70l] Herbert S. Bear. Standing assumptions and notations. Lecture Notes in Mathematics, 121:1–2, 1970. CODEN LNMAA2. ISBN 3-540-04910-X (print), 3-540-36255-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071116.

Beauville:1971:NRG

[Bea71] A. Beauville. Une notion de résidu en géométrie analytique. (French) []. Lecture Notes in Mathematics, 205:183–203, 1971. CODEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059839/.

Beaudet:1974:MGM

[Bea74] Dr. Paul R. Beaudet. Multi-off-grid methods in multi-step integration of ordinary differential equations. Lecture Notes in Mathematics, 362:128–148, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066589/.

Bebernes:1974:PIW

[Beb74] J. Bebernes. Positive invariance and a Wazewski Theorem. Lecture Notes in Mathematics, 415:35–39, 1974. CODEN LN-MAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065509/.

Beck:1970:CFP

[Bec70] Anatole Beck. Continuous flows in the plane. Lecture Notes in Mathematics, 160:1–6, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060640/.

Beck:1971:CFPa

[Bec71a] A. Beck. Continuous flows on the plane. Lecture Notes in Mathematics, 206:12–13, 1971. CODEN LNMAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070123.

Beck:1971:CFPb

[Bec71b] A. Beck. Continuous flows on the plane: techniques I. Lecture Notes in Mathematics, 206:19–21, 1971. CODEN LN-MAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070129/.

Beck:1971:CFPc

[Bec71c] A. Beck. Continuous flows on the plane: techniques II. Lecture Notes in Mathematics, 206:21–23, 1971. CODEN LN-MAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070130/.

Beck:1971:CSH

[Bec71d] Jon Beck. Classifying spaces for homotopy-everything H-spaces. Lecture Notes in Mathematics, 196:54–62, 1971. CODEN LN-MAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067747/.

Beck:1973:RAT

[Bec73a] Anatole Beck, editor. Recent Advances in Topological Dynamics: Proceedings of the Conference on Topological Dynamics, held at Yale University, June 19–23, 1972, in honor of Professor Gustav Arnold Hedlund on the occasion of his retirement, volume 318 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 318. URL http:/

/link.springer.com/book/10.1007/BFb0061716;   http://
www.springerlink.com/content/978-3-540-38414-4.

Beck:1973:UFS

[Bec73b] Anatole Beck. Uniqueness of flow solutions of differential equations. Lecture Notes in Mathematics, 318:30–50, 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061721/.

Beckenbach:1974:DRM

[Bec74a] E. F. Beckenbach. Defect relations for meromorphic minimal surfaces: an introduction. Lecture Notes in Mathematics, 419: 18–41, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064707/.

Becker:1974:CCT

[Bec74b] J. C. Becker. Characteristic classes and K-theory. Lecture Notes in Mathematics, 428:132–143, 1974. CODEN LNMAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070537/.

Behncke:1972:TNA

[Beh72a] Horst Behncke. Topics in C∗- and von Neumann algebras. Lecture Notes in Mathematics, 247:1–54, 1972. CODEN LNMAA2. ISBN 3-540-05729-3 (print), 3-540-37117-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058551/.

Behzad:1972:TG

[Beh72b] Mehdi Behzad. Total graphs. Lecture Notes in Mathematics, 303: 21–23, 1972. CODEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067353/.

Behara:1973:IDF

[Beh73] M. Behara. Invariance of decision functions under Lie groups I. Lecture Notes in Mathematics, 296:216–223, 1973. CODEN LNMAA2. ISBN 3-540-06211-4 (print), 3-540-38485-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059826/.

Behrens:1974:AS

[Beh74a] Michael Behrens. Analytic sets in M(D). Lecture Notes in Mathematics, 369:5–22, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065994/.

Behrens:1974:BVT

[Beh74b] Michael Behrens. Boundary value theorems for meromorphic functions defined in the open unit disk. Lecture Notes in Mathematics, 369:23–33, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065995/.

Behrens:1974:LIF

[Beh74c] Michael Behrens. A local inverse function theorem. Lecture Notes in Mathematics, 369:34–36, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065996/.

Beineke:1971:DGD

[Bei71] Lowell W. Beineke. Derived graphs with derived complements. Lecture Notes in Mathematics, 186:15–24, 1971. CODEN LN-MAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059419/.

Bell:1970:FPI

[Bel70] Harold Bell. Fixed points, index, and degree for some set valued functions. Lecture Notes in Mathematics, 171:1–7, 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069711/.

Bellman:1971:IIS

[Bel71a] Richard Bellman. Invariant imbedding: Semigroups in time, space and structure. Lecture Notes in Mathematics, 228:9– 18, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069444/.

Bellman:1971:ISS

[Bel71b] Richard Bellman. On the identification of systems and some related questions. Lecture Notes in Mathematics, 228:1–8, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069443/.

Bellamy:1974:MHI

[Bel74a] David P. Bellamy. Mapping hereditarily indecomposable continua onto a pseudo-arc. Lecture Notes in Mathematics, 375:6–14, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064006/.

Bellenot:1974:NTV

[Bel74b] Steven F. Bellenot. Nonstandard topological vector spaces. Lecture Notes in Mathematics, 369:37–39, 1974. CODEN LN-MAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065997/.

Bensoussan:1970:BLF

[Ben70] A. Bensoussan. On the best linear feedback for systems governed by differential operational equations. Lecture Notes in Mathematics, 132:53–57, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066672/.

Benedetto:1971:CUS

[Ben71a] John Benedetto. A characterization of uniqueness sets. Lecture Notes in Mathematics, 202:46–68, 1971. CODEN LN-MAA2. ISBN 3-540-05488-X (print), 3-540-36659-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069053/.

Benedetto:1971:CR

[Ben71b] John Benedetto. Concluding remarks. Lecture Notes in Mathematics, 202:159–160, 1971. CODEN LNMAA2. ISBN 3-540-05488-X (print), 3-540-36659-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069058.

Benedetto:1971:HAT

[Ben71c] John Benedetto. Harmonic Analysis on Totally Disconnected Sets, volume 202 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05488-X (print), 3-540-36659-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 261 pp. LCCN QA3 .L28 no. 202. URL http:/

/link.springer.com/book/10.1007/BFb0069050;   http://
www.springerlink.com/content/978-3-540-36659-1.

Benedetto:1971:HS

[Ben71d] John Benedetto. Helson sets. Lecture Notes in Mathematics, 202:141–158, 1971. CODEN LNMAA2. ISBN 3-540-05488-X (print), 3-540-36659-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069057/.

Benedetto:1971:ISA

[Ben71e] John Benedetto. Independent sets and arithmetic progressions. Lecture Notes in Mathematics, 202:69–92, 1971. CODEN LN-MAA2. ISBN 3-540-05488-X (print), 3-540-36659-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069054/.

Benedetto:1971:ISM

[Ben71f] John Benedetto. Independent sets of multiplicity. Lecture Notes in Mathematics, 202:124–140, 1971. CODEN LNMAA2. ISBN 3-540-05488-X (print), 3-540-36659-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069056/.

Benedetto:1971:KTK

[Ben71g] John Benedetto. Kronecker’s theorem and Kronecker sets. Lecture Notes in Mathematics, 202:93–123, 1971. CODEN LN-MAA2. ISBN 3-540-05488-X (print), 3-540-36659-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069055/.

Benedetto:1971:MT

[Ben71h] John Benedetto. Malliavin’s theorem. Lecture Notes in Mathematics, 202:207–250, 1971. CODEN LNMAA2. ISBN 3-540-05488-X (print), 3-540-36659-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069060/.

Benedetto:1971:PFA

[Ben71i] John Benedetto. Preliminaries from Fourier analysis and integration theory. Lecture Notes in Mathematics, 202:1–19, 1971. CODEN LNMAA2. ISBN 3-540-05488-X (print), 3-540-36659-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069051/.

Benedetto:1971:PMS

[Ben71j] John Benedetto. Pseudo-measures supported by totally disconnected sets. Lecture Notes in Mathematics, 202:20–45, 1971. CO-DEN LNMAA2. ISBN 3-540-05488-X (print), 3-540-36659-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069052/.

Benedetto:1971:WP

[Ben71k] John Benedetto. The Wiener process. Lecture Notes in Mathematics, 202:161–206, 1971. CODEN LNMAA2. ISBN 3-540-05488-X (print), 3-540-36659-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069059/.

Benzinger:1971:FIE

[Ben71l] Harold E. Benzinger. The Fourier integral and the expansion problem for ordinary differential operators. Lecture Notes in Mathematics, 183:23–27, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060407/.

Bennewitz:1972:SRH

[Ben72a] Ch. Bennewitz. Symmetric relations on a Hilbert space. Lecture Notes in Mathematics, 280:212–218, 1972. CODEN LN-MAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066931/.

Bensoussan:1972:OSL

[Ben72b] Alain Bensoussan. Optimization of sensors’ location in a distributed filtering problem. Lecture Notes in Mathematics, 294: 62–84, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064935/.

Benveniste:1973:ADT

[Ben73] A. Benveniste. Application de deux théorèmes de G. Mokobodzki à l’étude du noyau de Lévy d’un processus de Hunt sans hypothèse (L). (French) []. Lecture Notes in Mathematics, 321: 1–24, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071389/.

Bendersky:1974:FWL

[Ben74a] Martin Bendersky. A functor which localizes the higher homotopy groups of an arbitrary C. W. complex. Lecture Notes in Mathematics, 418:13–21, 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070638/.

Benzinger:1974:AHY

[Ben74b] Harold E. Benzinger. An application of the Hausdorff–Young inequality to eigenfunction expansions. Lecture Notes in Mathematics, 415:40–46, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065510/.

Berkovitz:1970:SNC

[Ber70a] Leonard D. Berkovitz. Some non-classical variational problems related to optimal filter problems. Lecture Notes in Mathematics, 132:58–60, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066673/.

Bernfeld:1970:PSG

[Ber70b] Stephen Bernfeld. Perturbation of systems with global existence. Lecture Notes in Mathematics, 144:16–21, 1970. CODEN LN-MAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059915/.

Bers:1970:SKG

[Ber70c] Lipman Bers. Spaces of Kleinian groups. Lecture Notes in Mathematics, 155:9–34, 1970. CODEN LNMAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060314/.

Bertin:1970:GPG

[Ber70d] J. E. Bertin. Généralités sur les préschémas en groupes. (French) []. Lecture Notes in Mathematics, 151:318–410, 1970. CODEN LNMAA2. ISBN 3-540-05179-1 (print), 3-540-36336-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059000/.

Berger:1971:TGM

[Ber71a] Marcel Berger. Le théorème de Gromoll–Meyer sur les géodésiques fermées. (French) []. Lecture Notes in Mathematics, 180:1–17, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058821/.

Bernstein:1971:BCR

[Ber71b] Leon Bernstein. Basic concepts and relations. Lecture Notes in Mathematics, 207:11–18, 1971. CODEN LNMAA2. ISBN 3-540-05497-9 (print), 3-540-36665-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069407/.

Bernstein:1971:CJ

[Ber71c] Leon Bernstein. Convergence of JPA. Lecture Notes in Mathematics, 207:19–29, 1971. CODEN LNMAA2. ISBN 3-540-05497-9 (print), 3-540-36665-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069408/.

Bernstein:1971:DE

[Ber71d] Leon Bernstein. Diophantine equations. Lecture Notes in Mathematics, 207:134–156, 1971. CODEN LNMAA2. ISBN 3-540-05497-9 (print), 3-540-36665-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069413/.

Bernstein:1971:I

[Ber71e] Leon Bernstein. Introduction. Lecture Notes in Mathematics, 207:1–10, 1971. CODEN LNMAA2. ISBN 3-540-05497-9 (print), 3-540-36665-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069406/.

Bernstein:1971:JPA

[Ber71f] Leon Bernstein. The Jacobi–Perron Algorithm Its Theory and Application, volume 207 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05497-9 (print), 3-540-36665-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 161 pp. LCCN QA3 .L28 no. 207.

URL http://link.springer.com/book/10.1007/BFb0069405;
 http://www.springerlink.com/content/978-3-540-36665-
2.

Bernstein:1971:PJ

[Ber71g] Leon Bernstein. Periodicity of JPA. Lecture Notes in Mathematics, 207:30–48, 1971. CODEN LNMAA2. ISBN 3-540-05497-9 (print), 3-540-36665-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069409/.

Bernstein:1971:SSC

[Ber71h] Leon Bernstein. Some special cases of JPA. Lecture Notes in Mathematics, 207:49–71, 1971. CODEN LNMAA2. ISBN 3-540-05497-9 (print), 3-540-36665-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069410/.

Bernstein:1971:UAN

[Ber71i] Leon Bernstein. Units in algebraic number fields. Lecture Notes in Mathematics, 207:102–133, 1971. CODEN LNMAA2. ISBN 3-540-05497-9 (print), 3-540-36665-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069412/.

Bernstein:1971:VF

[Ber71j] Leon Bernstein. Various T -functions. Lecture Notes in Mathematics, 207:72–101, 1971. CODEN LNMAA2. ISBN 3-540-05497-9 (print), 3-540-36665-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069411/.

Berthelot:1971:GAF

[Ber71k] P. Berthelot. Généralités sur les λ-anneaux. (French) []. Lecture Notes in Mathematics, 225:297–364, 1971. CODEN LN-MAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066289/.

Berthelot:1971:QCG

[Ber71l] P. Berthelot. Quelques calculs de groupes K. (French) []. Lecture Notes in Mathematics, 225:498–518, 1971. CODEN LN-MAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066293/.

Berthelot:1971:IRC

[Ber71m] Pierre Berthelot. Immersions regulières et calcul du K^{.} d’un schèma Éclate. (French) []. Lecture Notes in Mathematics, 225: 416–465, 1971. CODEN LNMAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066291/.

Berthelot:1971:FPC

[Ber71n] Pierre Berthelot. Le K d’un fibre projectif: Calculs et conséquences. (French) []. Lecture Notes in Mathematics, 225: 365–415, 1971. CODEN LNMAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066290/.

Berthelot:1971:TRR

[Ber71o] Pierre Berthelot. Le théorème de Riemann–Roch. (French) []. Lecture Notes in Mathematics, 225:466–497, 1971. CODEN LN-MAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066292/.

Berndt:1972:VSF

[Ber72a] Bruce C. Berndt. The Voronoï summation formula. Lecture Notes in Mathematics, 251:21–36, 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058783/.

Bernstein:1972:AFJ

[Ber72b] Leon Bernstein. The arithmetic functions of the Jacobi–Perron Algorithm. Lecture Notes in Mathematics, 251:37–63, 1972. CO-DEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058784/.

Bernard:1973:ANA

[Ber73] Alain Bernard. Algèbres non autoadjointes de champs continus d’opérateurs. (French) []. Lecture Notes in Mathematics, 336: 15–19, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065783/.

Berg:1974:SGC

[Ber74a] Christian Berg. Sur les semi-groupes de convolution. (French) []. Lecture Notes in Mathematics, 404:1–26, 1974. CODEN LNMAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060607/.

Berge:1974:IPH

[Ber74b] Claude Berge, C. N. R. S. Isomorphism problems for hypergraphs. Lecture Notes in Mathematics, 411:1–12, 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066174/.

BergeCNRS:1974:NCH

[Ber74c] Claude Berge, C. N. R. S. Nombres de coloration de l’hypergraphe h-parti complet. (French) []. Lecture Notes in Mathematics, 411:13–20, 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066175/.

Bermond:1974:GRH

[Ber74d] Jean-Claude Bermond, C. N. R. S. Graphe representatif de l’hypergraphe h-parti complet. (French) []. Lecture Notes in Mathematics, 411:34–53, 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066177/.

Bers:1974:ECE

[Ber74e] Lipman Bers. On Eichler cohomology and on Eichler integrals. Lecture Notes in Mathematics, 419:53–59, 1974. CODEN LN-MAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064709/.

Bers:1974:WKG

[Ber74f] Lipman Bers. What is a Kleinian group? Lecture Notes in Mathematics, 400:1–14, 1974. CODEN LNMAA2. ISBN 3-540-06840-6 (print), 3-540-37776-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065672/.

Berthelot:1974:CDF

[Ber74g] Prof. Pierre Berthelot. Calcul différentiel. (French) []. Lecture Notes in Mathematics, 407:81–177, 1974. CODEN LN-MAA2. ISBN 3-540-06852-X (print), 3-540-37807-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068639/.

Berthelot:1974:CCA

[Ber74h] Prof. Pierre Berthelot. Classe de cohomologie associée à un cycle non singulier. (French) []. Lecture Notes in Mathematics, 407: 381–520, 1974. CODEN LNMAA2. ISBN 3-540-06852-X (print), 3-540-37807-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068643/.

Berthelot:1974:CCS

[Ber74i] Prof. Pierre Berthelot. Cohomologie Cristalline des Schémas de Caractéristique p > 0. (French) [Crystalline cohomology of diagrams of characteristic p > 0], volume 407 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CO-DEN LNMAA2. ISBN 3-540-06852-X (print), 3-540-37807-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 586 (est.) pp. LCCN QA3 .L28 no. 407. URL http:/

/link.springer.com/book/10.1007/BFb0068636;   http://
www.springerlink.com/content/978-3-540-37807-5.

Berthelot:1974:CCC

[Ber74j] Prof. Pierre Berthelot. Cohomologie cristalline et cohomologie de de Rham. (French) []. Lecture Notes in Mathematics, 407: 287–380, 1974. CODEN LNMAA2. ISBN 3-540-06852-X (print), 3-540-37807-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068642/.

Berthelot:1974:CFC

[Ber74k] Prof. Pierre Berthelot. Cristaux. (French) [Crystals]. Lecture Notes in Mathematics, 407:225–286, 1974. CODEN LN-MAA2. ISBN 3-540-06852-X (print), 3-540-37807-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068641/.

Berthelot:1974:DPR

[Ber74l] Prof. Pierre Berthelot. Dualité de Poincaré et rationnalité de la fonction zeta. (French) []. Lecture Notes in Mathematics, 407: 521–586, 1974. CODEN LNMAA2. ISBN 3-540-06852-X (print), 3-540-37807-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068644/.

Berthelot:1974:GTC

[Ber74m] Prof. Pierre Berthelot. Généralités sur le topos cristallin d’un schéma. (French) []. Lecture Notes in Mathematics, 407:179– 223, 1974. CODEN LNMAA2. ISBN 3-540-06852-X (print), 3-540-37807-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068640/.

Berthelot:1974:IPD

[Ber74n] Prof. Pierre Berthelot. Idèaux à puissances divisées. (French) [Ideals with divided powers]. Lecture Notes in Mathematics, 407: 26–79, 1974. CODEN LNMAA2. ISBN 3-540-06852-X (print), 3-540-37807-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068638/.

Berthelot:1974:IFI

[Ber74o] Prof. Pierre Berthelot. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 407:7–25, 1974. CODEN LN-MAA2. ISBN 3-540-06852-X (print), 3-540-37807-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068637/.

Bettis:1974:ECH

[Bet74a] Dale G. Bettis. Equations of condition for high order Runge– Kutta–Nyström formulae. Lecture Notes in Mathematics, 362: 76–91, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066586/.

Bettis:1974:PCN

[Bet74b] Dale G. Bettis, editor. Proceedings of the Conference on the Numerical Solution of Ordinary Differential Equations: 19,20 October 1972, The University of Texas at Austin, volume 362 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CO-DEN LNMAA2. ISBN 0-387-06602-0, 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 362; QA3 .L35; QA3 .L28 no. 362; QA372; QA3 .L28; QA1 .L471; QA3 .L4; QA372 .C765p 1972.

URL http://link.springer.com/book/10.1007/BFb0066582;
 http://www.springerlink.com/content/978-3-540-37911-
9.

Bulman-Fleming:1974:ACR

[BF74] Sydney Bulman-Fleming. Algebraic compactness and its relations to topology. Lecture Notes in Mathematics, 378:89–94, 1974. CODEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068461/.

Brown:1973:ATG

[BG73] Kenneth S. Brown and Stephen M. Gersten. Algebraic K-theory as generalized sheaf cohomology. Lecture Notes in Mathematics, 341:266–292, 1973. CODEN LNMAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067062/.

Bucur:1972:FM

[BGG^{+}72] I. Bucur, J. Giraud, N. Goodman, J. Myhill, L. Illusie, J. Lambek, D. S. Scott, and M. Tierney. Front matter. Lecture Notes in Mathematics, 274:i–vi, 1972. CODEN LNMAA2. ISBN 3-540-05920-2 (print), 3-540-37609-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-37609-5/1.pdf.

Berthelot:1971:TIT

[BGI71] Pierre Berthelot, A. (Alexandre) Grothendieck, and Luc Illusie. Théorie des Intersections et Théorème de Riemann–Roch, volume 225 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 700 pp. LCCN QA3 .L28 no. 225; QA3 .L28 no. 225; QA564.

URL http://link.springer.com/book/10.1007/BFb0066283;

 http://www.springerlink.com/content/978-3-540-36936-
3.

Bott:1972:LAD

[BGJ72] R. Bott, S. Gitler, and I. M. James. Lectures on Algebraic and Differential Topology: Delivered at the II. ELAM, volume 279 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05944-X (print), 3-540-37616-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 174 pp. LCCN QA3 .L28 no. 279. URL http:

//link.springer.com/book/10.1007/BFb0058508;   http:/
/www.springerlink.com/content/978-3-540-37616-3.

Berger:1971:IRF

[BGM71a] Marcel Berger, Paul Gauduchon, and Edmond Mazet. Invariants riemanniens. (French) []. Lecture Notes in Mathematics, 194: 10–140, 1971. CODEN LNMAA2. ISBN 3-540-05437-5 (print), 3-540-36547-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064645/.

Berger:1971:SVRa

[BGM71b] Marcel Berger, Paul Gauduchon, and Edmond Mazet. Le spectre d’une Variété riemannienne. (French) []. Lecture Notes in Mathematics, 194:141–241, 1971. CODEN LNMAA2. ISBN 3-540-05437-5 (print), 3-540-36547-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064646/.

Berger:1971:SVRb

[BGM71c] Marcel Berger, Paul Gauduchon, and Edmond Mazet. Le Spectre d’une Variété Riemannienne. (French) [The spectrum of a Riemannian variety], volume 194 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05437-5 (print), 3-540-36547-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 241 (est.) pp. LCCN QA3 .L28 no. 194.

URL http://link.springer.com/book/10.1007/BFb0064643;
 http://www.springerlink.com/content/978-3-540-36547-

Berger:1971:VRF

[BGM71d] Marcel Berger, Paul Gauduchon, and Edmond Mazet. Variétés riemanniennes. (French) []. Lecture Notes in Mathematics, 194: 1–9, 1971. CODEN LNMAA2. ISBN 3-540-05437-5 (print), 3-540-36547-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064644/.

Borho:1973:EG

[BGR73a] Walter Borho, Peter Gabriel, and Rudolf Rentschler. Einleitung. (German) []. Lecture Notes in Mathematics, 357:1–5, 1973. CO-DEN LNMAA2. ISBN 3-540-06561-X (print), 3-540-37859-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069766/.

Borho:1973:NCA

[BGR73b] Walter Borho, Peter Gabriel, and Rudolf Rentschler. Noncommutative algebra. Lecture Notes in Mathematics, 357:6– 47, 1973. CODEN LNMAA2. ISBN 3-540-06561-X (print), 3-540-37859-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069767/.

Borho:1973:PIS

[BGR73c] Walter Borho, Peter Gabriel, and Rudolf Rentschler. Prime ideals and stable spectra. Lecture Notes in Mathematics, 357:132– 173, 1973. CODEN LNMAA2. ISBN 3-540-06561-X (print), 3-540-37859-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069770/.

Borho:1973:PEA

[BGR73d] Walter Borho, Peter Gabriel, and Rudolf Rentschler. Primideale in Einhüllenden auflösbarer Lie-Algebren: Beschreibung durch Bahnenräume. (German) [Prime ideals in envelope-resolvable Lie algebras: Description by track spaces], volume 357 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CO-DEN LNMAA2. ISBN 3-540-06561-X (print), 3-540-37859-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 182 pp. LCCN QA3 .L28 no. 357. URL http:/

/link.springer.com/book/10.1007/BFb0069765;   http://
www.springerlink.com/content/978-3-540-37859-4.

Borho:1973:PIO

[BGR73e] Walter Borho, Peter Gabriel, and Rudolf Rentschler. Primitive ideals and orbit spaces. Lecture Notes in Mathematics, 357:88– 131, 1973. CODEN LNMAA2. ISBN 3-540-06561-X (print), 3-540-37859-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069769/.

Borho:1973:RCA

[BGR73f] Walter Borho, Peter Gabriel, and Rudolf Rentschler. Residue class algebras of enveloping algebras. Lecture Notes in Mathematics, 357:48–87, 1973. CODEN LNMAA2. ISBN 3-540-06561-X (print), 3-540-37859-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069768/.

Barr:1971:ECC

[BGvO71] Michael Barr, Pierre A. Grillet, and Donovan H. van Osdol. Exact Categories and Categories of Sheaves, volume 236 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05678-5 (print), 3-540-36999-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 239 (est.) pp. LCCN QA3 .L28 no. 236. URL http:/

/link.springer.com/book/10.1007/BFb0058579;   http://
www.springerlink.com/content/978-3-540-36999-8.

Badrikian:1974:EEP

[BH74a] A. Badrikian and P.-L. Hennequin. École d’Eté de Probabilités de Saint-Flour III — 1973. (French) [], volume 390 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06816-3 (print), 3-540-37765-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 189 pp. LCCN QA3 .L28 no. 390. URL http:/

/link.springer.com/book/10.1007/BFb0082432;   http://
www.springerlink.com/content/978-3-540-37765-8.

Bari:1974:GCP

[BH74b] Prof. Ruth A. Bari and Prof. Frank Harary, editors. Graphs and Combinatorics: Proceedings of the Capital Conference on Graph Theory and Combinatorics at the George Washington University June 18–22, 1973, volume 406 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 406. URL http:/

/link.springer.com/book/10.1007/BFb0066428;   http://
www.springerlink.com/content/978-3-540-37809-9.

Bhatia:1970:ARD

[Bha70] Nam P. Bhatia. Asymptotic recurrence and dynamical flow near a compact minimal set. Lecture Notes in Mathematics, 144:22– 29, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059916/.

Badrikian:1970:CMF

[BI70a] A. Badrikian and B. Ivol. Convergence des mesures. (French) []. Lecture Notes in Mathematics, 139:68–82, 1970. CODEN LNMAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067898/.

Badrikian:1970:EPL

[BI70b] A. Badrikian and B. Ivol. Espaces polonais, lusiniens, sousliniens et radoniens. (French) []. Lecture Notes in Mathematics, 139: 112–139, 1970. CODEN LNMAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067901/.

Badrikian:1970:IEC

[BI70c] A. Badrikian and B. Ivol. Intégration sur un espace complètement régulier. (French) []. Lecture Notes in Mathematics, 139:48– 67, 1970. CODEN LNMAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067897/.

Badrikian:1970:TEF

[BI70d] A. Badrikian and B. Ivol. Topologies équivalentes. (French) []. Lecture Notes in Mathematics, 139:140–151, 1970. CODEN LN-MAA2. ISBN 3-540-04928-2 (print), 3-540-36296-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067902/.

Bichteler:1973:EOM

[Bic73a] Klaus Bichteler. Elementary operations on measure spaces. Lecture Notes in Mathematics, 315:223–299, 1973. CODEN LN-MAA2. ISBN 3-540-06158-4 (print), 3-540-38369-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069351/.

Bichteler:1973:ITS

[Bic73b] Klaus Bichteler. Integration Theory (with Special Attention to Vector Measures), volume 315 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06158-4 (print), 3-540-38369-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 357 pp. LCCN QA3 .L28 no. 315.

URL http://link.springer.com/book/10.1007/BFb0069346;
 http://www.springerlink.com/content/978-3-540-38369-
7.

Bichteler:1973:I

[Bic73c] Klaus Bichteler. Introduction. Lecture Notes in Mathematics, 315:1–11, 1973. CODEN LNMAA2. ISBN 3-540-06158-4 (print), 3-540-38369-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069347/.

Bichteler:1973:LD

[Bic73d] Klaus Bichteler. Liftings and derivatives. Lecture Notes in Mathematics, 315:300–342, 1973. CODEN LNMAA2. ISBN 3-540-06158-4 (print), 3-540-38369-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069352/.

Bichteler:1973:M

[Bic73e] Klaus Bichteler. Measurability. Lecture Notes in Mathematics, 315:165–222, 1973. CODEN LNMAA2. ISBN 3-540-06158-4 (print), 3-540-38369-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069350/.

Bichteler:1973:RSE

[Bic73f] Klaus Bichteler. Riesz spaces and elementary integrals. Lecture Notes in Mathematics, 315:12–72, 1973. CODEN LN-MAA2. ISBN 3-540-06158-4 (print), 3-540-38369-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069348/.

Bichteler:1973:UGE

[Bic73g] Klaus Bichteler. Upper gauges and extension theory. Lecture Notes in Mathematics, 315:73–164, 1973. CODEN LN-MAA2. ISBN 3-540-06158-4 (print), 3-540-38369-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069349/.

Binz:1973:CS

[Bin73] E. Binz. Convergence structures on C(X). Lecture Notes in Mathematics, 331:203–210, 1973. CODEN LNMAA2. ISBN 3-540-06367-6 (print), 3-540-46977-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068231/.

Bing:1974:UMC

[Bin74] R. H. Bing. An unusual map of a 3-cell onto itself. Lecture Notes in Mathematics, 375:15–33, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064007/.

Birch:1973:SCM

[Bir73] B. J. Birch. Some calculations of modular relations. Lecture Notes in Mathematics, 320:175–186, 1973. CODEN LNMAA2. ISBN 3-540-06219-X (print), 3-540-38509-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38509-7_6/.

Banks:1970:OCL

[BJ70] H. T. Banks and Marc Q. Jacobs. Optimal control and linear functional differential equations. Lecture Notes in Mathematics, 144:5–15, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059914/.

Bousfield:1972:BM

[BK72a] Aldridge K. Bousfield and Daniel M. Kan. Back matter. Lecture Notes in Mathematics, 304:341–348, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bbm:978-3-540-38117-4/1.pdf.

Bousfield:1972:CSb

[BK72b] Aldridge K. Bousfield and Daniel M. Kan. Cosimplicial spaces. Lecture Notes in Mathematics, 304:265–286, 1972. CODEN LN-MAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-38117-4_
10/.

Bousfield:1972:ECN

[BK72c] Aldridge K. Bousfield and Daniel M. Kan. Erratum to: p-Completions of nilpotent spaces. Lecture Notes in Mathematics, 304:349, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38117-4_15. See [BK72n].

Bousfield:1972:ECS

[BK72d] Aldridge K. Bousfield and Daniel M. Kan. Erratum to: The R-completion of a space. Lecture Notes in Mathematics, 304: 349, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38117-4_13. See [BK72o].

Bousfield:1972:ETL

[BK72e] Aldridge K. Bousfield and Daniel M. Kan. Erratum to: Tower lemmas. Lecture Notes in Mathematics, 304:349, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-38117-4_

14. See [BK72s].

Bousfield:1972:FL

[BK72f] Aldridge K. Bousfield and Daniel M. Kan. Fibre lemmas. Lecture Notes in Mathematics, 304:48–69, 1972. CODEN LNMAA2.

ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38117-4_2/.

Bousfield:1972:FMa

[BK72g] Aldridge K. Bousfield and Daniel M. Kan. Front matter. Lecture Notes in Mathematics, 304:i–v, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-38117-4/1.pdf.

Bousfield:1972:FMb

[BK72h] Aldridge K. Bousfield and Daniel M. Kan. Front matter. Lecture Notes in Mathematics, 304:1–9, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-540-38117-4/1/1.pdf.

Bousfield:1972:FMc

[BK72i] Aldridge K. Bousfield and Daniel M. Kan. Front matter. Lecture Notes in Mathematics, 304:224–227, 1972. CODEN LN-MAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bfm:978-3-540-38117-4/
2/1.pdf.

Bousfield:1972:GCN

[BK72j] Aldridge K. Bousfield and Daniel M. Kan. A glimpse at the Rcompletion of non-nilpotent spaces. Lecture Notes in Mathematics, 304:202–223, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-38117-4_7/.

Bousfield:1972:HDL

[BK72k] Aldridge K. Bousfield and Daniel M. Kan. Homotopy direct limits. Lecture Notes in Mathematics, 304:325–340, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-38117-4_

Bousfield:1972:HIL

[BK72l] Aldridge K. Bousfield and Daniel M. Kan. Homotopy inverse limits. Lecture Notes in Mathematics, 304:287–324, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38117-4_11/.

Bousfield:1972:HLC

[BK72m] Aldridge K. Bousfield and Daniel M. Kan. Homotopy Limits, Completions and Localizations, volume 304 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 348 pp. LCCN QA3 .L28

no. 304. URL http://link.springer.com/book/10.1007/978-
3-540-38117-4;   http://www.springerlink.com/content/
978-3-540-38117-4.

Bousfield:1972:CNS

[BK72n] Aldridge K. Bousfield and Daniel M. Kan. p-Completions of nilpotent spaces. Lecture Notes in Mathematics, 304:163–201, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38117-4_6/. See erratum [BK72c].

Bousfield:1972:CSa

[BK72o] Aldridge K. Bousfield and Daniel M. Kan. The R-completion of a space. Lecture Notes in Mathematics, 304:10–47, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-38117-4_

1/. See erratum [BK72d].

Bousfield:1972:CGR

[BK72p] Aldridge K. Bousfield and Daniel M. Kan. An R-completion of groups and its relation to the R-completion of spaces. Lecture Notes in Mathematics, 304:99–125, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38117-4_4/.

Bousfield:1972:LNS

[BK72q] Aldridge K. Bousfield and Daniel M. Kan. R-localizations of nilpotent spaces. Lecture Notes in Mathematics, 304:126–162, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38117-4_5/.

Bousfield:1972:SST

[BK72r] Aldridge K. Bousfield and Daniel M. Kan. Simplicial sets and topological spaces. Lecture Notes in Mathematics, 304:228– 248, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38117-4_8/.

Bousfield:1972:TL

[BK72s] Aldridge K. Bousfield and Daniel M. Kan. Tower lemmas. Lecture Notes in Mathematics, 304:70–98, 1972. CODEN LNMAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38117-4_3/. See

erratum [BK72e].

Bousfield:1972:TF

[BK72t] Aldridge K. Bousfield and Daniel M. Kan. Towers of fibrations. Lecture Notes in Mathematics, 304:249–264, 1972. CODEN LN-MAA2. ISBN 3-540-06105-3 (print), 3-540-38117-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-38117-4_
9/.

Bers:1974:CCK

[BK74] Lipman Bers and Irwin Kra, editors. A Crash Course on Kleinian Groups: Lectures given at a special session at the January 1974 meeting of the American Mathematical Society at San Francisco, volume 400 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06840-6 (print), 3-540-37776-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 400. URL http:/

/link.springer.com/book/10.1007/BFb0065671;   http://
www.springerlink.com/content/978-3-540-37776-4.

Behara:1973:PIT

[BKW73] M. Behara, K. Krickeberg, and J. Wolfowitz. Probability and Information Theory II, volume 296 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 0-387-06211-4, 3-540-06211-4 (print), 3-540-38485-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 223 pp. LCCN QA3 .L28 no. 296; QA3 .L28 no. 296 QA273. URL http:/

/link.springer.com/book/10.1007/BFb0059817;   http://
www.springerlink.com/content/978-3-540-38485-4.

Bayes:1974:CNG

[BKW74] A. J. Bayes, J. Kautsky, and J. W. Wamsley. Computation in nilpotent groups (application). Lecture Notes in Mathematics, 372:82–89, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21571-5_7/;
http://link.springer.com/chapter/10.1007/BFb0065161/.

Blaquiere:1970:MSQ

[BL70] A. Blaquière and G. Leitmann. Multiple-stage quantitative games. Lecture Notes in Mathematics, 112:51–79, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060197/.

Bochnak:1971:CKK

[BŁ71a] J. Bochnak and S. Łojasiewicz. A converse of the Kuiper–Kuo theorem. Lecture Notes in Mathematics, 192:254–261, 1971. CO-DEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066825/.

Bochnak:1971:RFD

[BŁ71b] Jacek Bochnak and Stanislaw Łojasiewicz. Remarks on finitely determined analytic germs. Lecture Notes in Mathematics, 192: 262–270, 1971. CODEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066826/.

Bernstein:1974:NSI

[BL74a] Allen R. Bernstein and Peter A. Loeb. A non-standard integration theory for unbounded functions. Lecture Notes in Mathematics, 369:40–49, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065998/.

Borges:1974:CMS

[BL74b] C. R. Borges and D. J. Lutzer. Characterizations and mappings of M_{i} spaces. Lecture Notes in Mathematics, 375:34–40, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064008/.

Blair:1974:EST

[Bla74] Robert L. Blair. On ν-embedded sets in topological spaces. Lecture Notes in Mathematics, 378:46–79, 1974. CODEN LN-MAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068459/.

Bliedtner:1971:DFR

[Bli71a] Jürgen Bliedtner. Dirichlet forms on regular functional spaces. Lecture Notes in Mathematics, 226:15–62, 1971. CODEN LN-MAA2. ISBN 3-540-05638-6 (print), 3-540-36912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060480/.

Bliedtner:1971:FST

[Bli71b] Jürgen Bliedtner. Functional spaces and their exceptional sets. Lecture Notes in Mathematics, 226:1–14, 1971. CODEN LN-MAA2. ISBN 3-540-05638-6 (print), 3-540-36912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060479/.

Bliedtner:1974:EHF

[Bli74] Jürgen Bliedtner. Extreme harmonic functions on groups. Lecture Notes in Mathematics, 404:27–34, 1974. CODEN LN-MAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060608/.

Bloom:1970:CRE

[Blo70] Thomas Bloom. Cohomologie de de Rham d’un espace analytique. (French) []. Lecture Notes in Mathematics, 116:65– 74, 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059246/.

Bloom:1972:ODE

[Blo72] Thomas Bloom. Opérateurs différentiels elliptiques sur un espace analytique. (French) []. Lecture Notes in Mathematics, 275: 158–165, 1972. CODEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058499/.

Bloch:1973:ATA

[Blo73a] S. Bloch. Algebraic K-theory and algebraic geometry. Lecture Notes in Mathematics, 341:259–265, 1973. CODEN LN-MAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067061/.

Bloch:1973:TSQ

[Blo73b] S. Bloch. On the tangent space to Quillen K-theory. Lecture Notes in Mathematics, 341:205–210, 1973. CODEN LN-MAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067058/.

Barroso:1974:BSH

[BMN74] Jorge Alberto Barroso, Mário C. Matos, and Leopoldo Nachbin. On bounded sets of holomorphic mappings. Lecture Notes in Mathematics, 364:123–134, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069011/.

Brewer:1973:KDP

[BMRH73] J. W. Brewer, P. R. Montgomery, E. A. Rutter, and W. J. Heinzer. Krull dimension of polynomial rings. Lecture Notes in Mathematics, 311:26–45, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068916/.

Behara:1973:ANA

[BN73] M. Behara and P. Nath. Additive and non-additive entropies of finite measurable partitions. Lecture Notes in Mathematics, 296: 102–138, 1973. CODEN LNMAA2. ISBN 3-540-06211-4 (print), 3-540-38485-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059821/.

Boardman:1972:CIR

[Boa72] J. M. Boardman. Cobordism of involutions revisited. Lecture Notes in Mathematics, 298:131–151, 1972. CODEN LN-MAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070037/.

Bochnak:1972:RJF

[Boc72] J. Bochnak. Relèvement des jets. (French) []. Lecture Notes in Mathematics, 275:106–118, 1972. CODEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058494/.

Bohl:1974:ZMG

[Boh74] Erich Bohl. Über eine Zeilensummenbedingung bei L-Matrizen. (German) []. Lecture Notes in Mathematics, 395:247–262, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060676/.

Boland:1974:MTE

[Bol74] Philip J. Boland. Malgrange theorem for entire functions on nuclear spaces. Lecture Notes in Mathematics, 364:135–144, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069012/.

Bombieri:1970:PMC

[Bom70] Enrico Bombieri. The pluricanonical map of a complex surface. Lecture Notes in Mathematics, 155:35–87, 1970. CODEN LN-MAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060315/.

Bombieri:1971:RHM

[Bom71] Enrico Bombieri. Régularité des hypersurfaces minimales. (French) []. Lecture Notes in Mathematics, 179:111–121, 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058808/.

Bombieri:1973:SAA

[Bom73] Enrico Bombieri. Simultaneous approximations of algebraic numbers [following W. M. Schmidt]. Lecture Notes in Mathematics, 317:1–20, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069273/.

Bombieri:1974:CPC

[Bom74] Enrico Bombieri. Counting points on curves over finite fields. Lecture Notes in Mathematics, 383:234–241, 1974. CODEN LN-MAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057311/.

Bona:1972:FPT

[Bon72a] J. Bona. Fixed-point theorems for Fréchet spaces. Lecture Notes in Mathematics, 280:219–222, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066932/.

Bondy:1972:GTG

[Bon72b] J. A. Bondy. The ‘graph theory’ of the Greek alphabet. Lecture Notes in Mathematics, 303:43–54, 1972. CODEN LN-MAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067356/.

Boone:1974:SAP

[Boo74a] James R. Boone. Some applications of pointwise collectionwise normality. Lecture Notes in Mathematics, 378:80–88, 1974. CO-DEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068460/.

Boone:1974:BLG

[Boo74b] William W. Boone. Between logic and group theory. Lecture Notes in Mathematics, 372:90–102, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (ebook), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21571-5_8/;   http://link.springer.
com/chapter/10.1007/BFb0065162/.

Borel:1970:PLR

[Bor70] Armand Borel. Properties and linear representations of Chevalley groups. Lecture Notes in Mathematics, 131:1–55, 1970. CODEN LNMAA2. ISBN 3-540-04920-7 (print), 3-540-36272-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081542/.

Borel:1971:SGD

[Bor71a] Armand Borel. Sous-groupes discrets de groupes semi-simples. (French) []. Lecture Notes in Mathematics, 179:199–215, 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058813/.

Boruvka:1971:GET

[Bor71b] O. Borůvka. Geometric elements in the theory of transformations of ordinary second-order linear differential equations. Lecture Notes in Mathematics, 206:15–17, 1971. CODEN LN-MAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070127/.

Borel:1972:SDR

[Bor72a] Armand Borel. La série discrète (résultats; bouts de démonstrations). (French) []. Lecture Notes in Mathematics, 276:69–94, 1972. CO-DEN LNMAA2. ISBN 3-540-05926-1 (print), 3-540-37611-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058410/.

Borel:1972:NFC

[Bor72b] Armand Borel. Notations. Formulaire de convolution. (French) []. Lecture Notes in Mathematics, 276:1–4, 1972. CODEN LN-MAA2. ISBN 3-540-05926-1 (print), 3-540-37611-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058408/.

Borel:1972:QNR

[Bor72c] Armand Borel. Quelques notions et résultats de base. (French) []. Lecture Notes in Mathematics, 276:5–67, 1972. CODEN LNMAA2. ISBN 3-540-05926-1 (print), 3-540-37611-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058409/.

Borel:1972:RGL

[Bor72d] Armand Borel. Représentations de Groupes Localement Compacts, volume 276 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05926-1 (print), 3-540-37611-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 98 pp. LCCN QA3 .L28 no. 276. URL http:/

/link.springer.com/book/10.1007/BFb0058407;   http://
www.springerlink.com/content/978-3-540-37611-8.

Borovkov:1973:GTC

[Bor73] A. A. Borovkov. General theorems of convergence for random processes. Lecture Notes in Mathematics, 330:7–15, 1973. CO-DEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061478/.

Boruvka:1974:CDD

[Bor74] O. Borůvka. On central dispersions of the differential equation Y′′ = q(t)Y with periodic coefficients. Lecture Notes in Mathematics, 415:47–61, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065511/.

Bott:1972:LCC

[Bot72] Raoul Bott. Lectures on characteristic classes and foliations. Lecture Notes in Mathematics, 279:1–94, 1972. CODEN LN-MAA2. ISBN 3-540-05944-X (print), 3-540-37616-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058509/.

Bouchon:1974:PHQ

[Bou74a] Bernadette Bouchon. Propriétés des hemi-questionnaires. (French) []. Lecture Notes in Mathematics, 398:152–164, 1974. CODEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065775/.

Bousfield:1974:HLS

[Bou74b] A. K. Bousfield. Homological localizations of spaces, groups, and II-modules. Lecture Notes in Mathematics, 418:22–30, 1974. CO-DEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070639/.

Bowen:1973:SDH

[Bow73] Rufus Bowen. Symbolic dynamics for hyperbolic systems. Lecture Notes in Mathematics, 318:51–58, 1973. CODEN LN-MAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061722/.

Beineke:1972:CBD

[BP72] L. W. Beineke and R. E. Pippert. A census of ball and disk dissections. Lecture Notes in Mathematics, 303:25–40, 1972. CODEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067354/.

Bennewitz:1974:SEO

[BP74] Christer Bennewitz and Åke Pleijel. Selfadjoint extension of ordinary differential operators. Lecture Notes in Mathematics, 419: 42–52, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064708/.

Barnes:1973:TIM

[BR73a] E. R. Barnes and D. L. Russell. Totally implicit methods for numerical solution of singular initial value problems. Lecture Notes in Mathematics, 312:164–187, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060054/.

Brewer:1973:NFF

[BR73b] J. W. Brewer and E. A. Rutter. A note on the faithfulness of the functor SOX_{R}(−). Lecture Notes in Mathematics, 311:46– 49, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068917/.

Brewer:1973:CCA

[BR73c] James W. Brewer and Edgar A. Rutter, editors. Conference on Commutative Algebra: Lawrence, Kansas 1972, volume 311 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 311. URL http:/

/link.springer.com/book/10.1007/BFb0068913;   http://
www.springerlink.com/content/978-3-540-38340-6.

Braun:1971:FNS

[Bra71] Martin Braun. Flows near a singularity. Lecture Notes in Mathematics, 235:31–34, 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059191/.

Braaksma:1972:RSL

[Bra72a] B. L. J. Braaksma. Recessive solutions of linear differential equations with polynomial coefficients. Lecture Notes in Mathematics, 280:1–15, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066915/.

Bradley:1972:BVP

[Bra72b] J. S. Bradley. Boundary value problems involving the Thomas– Fermi equation. Lecture Notes in Mathematics, 280:223–226, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066933/.

Brauer:1974:SBC

[Bra74] Richard Brauer. On the structure of blocks of characters of finite groups. Lecture Notes in Mathematics, 372:103–130, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-21571-5_9/;   http://link.springer.
com/chapter/10.1007/BFb0065163/.

Berge:1974:HSO

[BRC74] Claude Berge and Dijen Ray-Chaudhuri, editors. Hypergraph Seminar: Ohio State University 1972, volume 411 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 411.

URL http://link.springer.com/book/10.1007/BFb0066173;
 http://www.springerlink.com/content/978-3-540-37803-
7.

Brelot:1971:AMT

[Bre71a] Marcel Brelot. Abstract minimal thinness, minimal boundary, minimal fine topology. Lecture Notes in Mathematics, 175:99– 105, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060365/.

Brelot:1971:ABW

[Bre71b] Marcel Brelot. Applications to balayage, weights and capacities. Lecture Notes in Mathematics, 175:61–72, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060361/.

Brelot:1971:CFT

[Bre71c] Marcel Brelot. Classical fine topology-general properties. Lecture Notes in Mathematics, 175:54–60, 1971. CODEN LN-MAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060360/.

Brelot:1971:CMB

[Bre71d] Marcel Brelot. Classical Martin boundary Dirichlet problem and boundary behaviour. Lecture Notes in Mathematics, 175:135– 147, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060369/.

Brelot:1971:CMSb

[Bre71e] Marcel Brelot. Classical Martin space and minimal thinness. Lecture Notes in Mathematics, 175:122–134, 1971. CODEN LN-MAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060368/.

Brelot:1971:CMSa

[Bre71f] Marcel Brelot. Classical Martin space the Martin integral representation. Lecture Notes in Mathematics, 175:111–121, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060367/.

Brelot:1971:CBT

[Bre71g] Marcel Brelot. Comparison of both thinnesses. Fine limits and non-tangential limits. (Classical case. Examples). Lecture Notes in Mathematics, 175:148–159, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060370/.

Brelot:1971:EAT

[Bre71h] Marcel Brelot. Extension to axiomatic theories of harmonic functions. Lecture Notes in Mathematics, 175:90–97, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060364/.

Brelot:1971:FSC

[Bre71i] Marcel Brelot. Further study of classical thinness. Some applications. Lecture Notes in Mathematics, 175:73–85, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060362/.

Brelot:1971:GCC

[Bre71j] Marcel Brelot. General compactification of Constantinescu– Cornea first examples of application. Lecture Notes in Mathematics, 175:106–110, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060366/.

Brelot:1971:GNT

[Bre71k] Marcel Brelot. General notions of thinness and fine topology. Lecture Notes in Mathematics, 175:1–8, 1971. CODEN LN-MAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060354/.

Brelot:1971:GRF

[Bre71l] Marcel Brelot. General results on fine limits. Lecture Notes in Mathematics, 175:18–21, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060356/.

Brelot:1971:MSM

[Bre71m] Marcel Brelot. Martin space and minimal thinness in axiomatic theories — short survey. Lecture Notes in Mathematics, 175: 160–164, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060371/.

Brelot:1971:NRF

[Bre71n] Marcel Brelot. Notion of reduced function. Applications. Strong thinness and strong unthinness. Lecture Notes in Mathematics, 175:9–17, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060355/.

Brelot:1971:NCP

[Bre71o] Marcel Brelot. Notions in classical potential theory. Lecture Notes in Mathematics, 175:37–53, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060359/.

Brelot:1971:TBP

[Bre71p] Marcel Brelot, editor. On Topologies and Boundaries in Potential Theory: Enlarged edition of a course of lectures delivered in 1966, volume 175 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 175. URL http:/

/link.springer.com/book/10.1007/BFb0060353;   http://
www.springerlink.com/content/978-3-540-36447-4.

Brelot:1971:QTN

[Bre71q] Marcel Brelot. Quasi-topological notions. Lecture Notes in Mathematics, 175:22–31, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060357/.

Brelot:1971:RCB

[Bre71r] Marcel Brelot. Relations with the Choquet boundary. Lecture Notes in Mathematics, 175:86–89, 1971. CODEN LN-MAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060363/.

Brelot:1971:WT

[Bre71s] Marcel Brelot. Weak thinness. Lecture Notes in Mathematics, 175:32–36, 1971. CODEN LNMAA2. ISBN 3-540-05327-1 (print), 3-540-36447-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060358/.

Bretagnolle:1971:RKP

[Bre71t] Jean Bretagnolle. Résultats de Kesten sur les processus à accroissements indépendants. (French) []. Lecture Notes in Mathematics, 191:21–36, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058844/.

Bredon:1972:SCA

[Bre72a] Glen E. Bredon. Strange circle actions on products of odd dimensional spheres, and rational homotopy. Lecture Notes in Mathematics, 298:291–299, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070049/.

Bresters:1972:IVP

[Bre72b] D. W. Bresters. The initial value problem for operators which can be factorized into a product of Klein–Gordon operators. Lecture Notes in Mathematics, 280:227–231, 1972. CODEN LN-MAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066934/.

Bretagnolle:1972:VFAa

[Bre72c] Jean Bretagnolle. p-Variation de fonctions aléatoires 1ère partie: Séries de Rademacher. (French) []. Lecture Notes in Mathematics, 258:51–63, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059460/.

Bretagnolle:1972:VFAb

[Bre72d] Jean Bretagnolle. p-Variation de fonctions aléatoires 2ième partie: Processus à accroissements indépendants. (French) []. Lecture Notes in Mathematics, 258:64–71, 1972. CODEN LN-MAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059461/.

Bretagnolle:1973:PAI

[Bre73] J. L. Bretagnolle. Processus à accroissements indépendants. (French) []. Lecture Notes in Mathematics, 307:1–26, 1973. CO-DEN LNMAA2. ISBN 3-540-06126-6 (print), 3-540-38321-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059707/.

Brelot:1974:QRA

[Bre74a] M. Brelot. Quelques recherches sur l’allure à la frontière des fonctions harmoniques. (French) []. Lecture Notes in Mathematics, 419:60–66, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064710/.

Bretagnolle:1974:RPS

[Bre74b] J. Bretagnolle. Une remarque sur le probléme de Skorokhod. (French) []. Lecture Notes in Mathematics, 381:11–19, 1974. CO-DEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057251/.

Brezis:1974:SPH

[Bré74c] Haïm Brézis. Singular perturbations of hyperbolic systems. Lecture Notes in Mathematics, 384:168–176, 1974. CODEN LN-MAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057547/.

Brieskorn:1973:GTD

[Bri73] Egbert Brieskorn. Sur les groupes de tresses [d’après V. I. Arnol’d]. (French) []. Lecture Notes in Mathematics, 317:21– 44, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069274/.

Brockett:1970:SPE

[Bro70] Roger W. Brockett. Structural properties of the equilibrium solutions of Riccati equations. Lecture Notes in Mathematics, 132: 61–69, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066674/.

Brooks:1971:SSC

[Bro71a] R. M. Brooks, editor. Symposium on Several Complex Variables, Park City, Utah, 1970, volume 184 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 184. URL http:/

/link.springer.com/book/10.1007/BFb0059274;   http://
www.springerlink.com/content/978-3-540-36455-9.

Browder:1971:MHT

[Bro71b] William Browder. Manifolds and homotopy theory. Lecture Notes in Mathematics, 197:17–35, 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068609/.

Brown:1973:MAC

[Bro73a] James R. Brown. Monothetic automorphisms of a compact abelian group. Lecture Notes in Mathematics, 318:59–77, 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061723/.

Brown:1973:DIO

[Bro73b] Lawrence G. Brown. The determinant invariant for operators with trace class self commutators. Lecture Notes in Mathematics, 345:210–228, 1973. CODEN LNMAA2. ISBN 3-540-06496-6 (print), 3-540-37808-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058920/.

Broucke:1974:CSP

[Bro74a] R. Broucke. Computation of solar perturbations with Poisson series. Lecture Notes in Mathematics, 362:237–259, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066594/.

Brown:1974:PHT

[Bro74b] Edward M. Brown. Proper homotopy theory in simplicial complexes. Lecture Notes in Mathematics, 375:41–46, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064009/.

Browne:1974:MPP

[Bro74c] Patrick J. Browne. Multi-parameter problems. Lecture Notes in Mathematics, 415:78–84, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065513/.

Brunel:1970:NCE

[Bru70] A. Brunel. New conditions for existence of invariant measures in ergodic theory. Lecture Notes in Mathematics, 160:7–17, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060641/.

Brualdi:1971:GTT

[Bru71a] Richard A. Brualdi. Generalized transversal theory. Lecture Notes in Mathematics, 211:5–30, 1971. CODEN LNMAA2. ISBN 3-540-05545-2 (print), 3-540-36877-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061072/.

Brunovksy:1971:OPF

[Bru71b] P. Brunovksý. One-parameter families of diffeomorphisms. Lecture Notes in Mathematics, 206:29–33, 1971. CODEN LN-MAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070134/.

Bruter:1971:TMR

[Bru71c] C. P. Bruter, editor. Théorie des Matroïdes: Rencontre Franco– Britannique Actes 14–15 Mai 1970, volume 211 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05545-2 (print), 3-540-36877-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 211.

URL http://link.springer.com/book/10.1007/BFb0061070;
 http://www.springerlink.com/content/978-3-540-36877-
9.

Bruneau:1974:FVB

[Bru74a] Michel Bruneau. Fonctions à p-variation bornée. (French) []. Lecture Notes in Mathematics, 413:1–18, 1974. CODEN LN-MAA2. ISBN 3-540-06952-6 (print), 3-540-37905-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066142/.

Bruneau:1974:FPF

[Bru74b] Michel Bruneau. Fonctions p,α-fines et nombres mal approchées. (French) []. Lecture Notes in Mathematics, 413:68–99, 1974. CO-DEN LNMAA2. ISBN 3-540-06952-6 (print), 3-540-37905-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066144/.

Bruneau:1974:FFP

[Bru74c] Michel Bruneau. Fonctions p-fines et points extremaux. (French) []. Lecture Notes in Mathematics, 413:19–67, 1974. CODEN LNMAA2. ISBN 3-540-06952-6 (print), 3-540-37905-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066143/.

Bruneau:1974:MCM

[Bru74d] Michel Bruneau. Mesures de Carathéodory et mesures de Besicovitch. (French) []. Lecture Notes in Mathematics, 413:100– 144, 1974. CODEN LNMAA2. ISBN 3-540-06952-6 (print), 3-540-37905-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066145/.

Bruneau:1974:NDF

[Bru74e] Michel Bruneau. Sur les nombres derivés d’une fonction. (French) []. Lecture Notes in Mathematics, 413:248–263, 1974. CODEN LNMAA2. ISBN 3-540-06952-6 (print), 3-540-37905-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066148/.

Bruneau:1974:GDT

[Bru74f] Michel Bruneau. Une généralisation du théorème de decomposition de Lebesgue. (French) []. Lecture Notes in Mathematics, 413: 231–247, 1974. CODEN LNMAA2. ISBN 3-540-06952-6 (print), 3-540-37905-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066147/.

Bruneau:1974:VTFa

[Bru74g] Michel Bruneau. Variation totale d’une fonction. (French) [Total variation of a function]. Lecture Notes in Mathematics, 413: 145–230, 1974. CODEN LNMAA2. ISBN 3-540-06952-6 (print), 3-540-37905-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066146/.

Bruneau:1974:VTFb

[Bru74h] Michel Bruneau. Variation Totale d’une Fonction. (French) [Total variation of a function], volume 413 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06952-6 (print), 3-540-37905-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). 263 (est.) pp. LCCN QA3 .L28 no. 413. URL http://

link.springer.com/book/10.1007/BFb0066141;   http://
www.springerlink.com/content/978-3-540-37905-8.

Brunner:1974:TSC

[Bru74i] A. M. Brunner. Transitivity-systems of certain one-relator groups. Lecture Notes in Mathematics, 372:131–140, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-662-21571-5_10/;   http://link.
springer.com/chapter/10.1007/BFb0065164/.

Bruter:1974:CAL

[Bru74j] Claude P. Bruter. Chapitre I. l’axiomatique linéaire: Definition des matroïdes. (French) []. Lecture Notes in Mathematics, 387: 4–18, 1974. CODEN LNMAA2. ISBN 3-540-06789-2 (print), 3-540-38397-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057564/.

Bruter:1974:CIRa

[Bru74k] Claude P. Bruter. Chapitre II. Représentation géométrique des matroïdes. (French) []. Lecture Notes in Mathematics, 387:19– 27, 1974. CODEN LNMAA2. ISBN 3-540-06789-2 (print), 3-540-38397-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057565/.

Bruter:1974:CIC

[Bru74l] Claude P. Bruter. Chapitre III. Construction des matroïdes et applications. (French) []. Lecture Notes in Mathematics, 387: 28–42, 1974. CODEN LNMAA2. ISBN 3-540-06789-2 (print), 3-540-38397-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057566/.

Bruter:1974:CIRb

[Bru74m] Claude P. Bruter. Chapitre IV. Représentation algébrique des matroïdes. (French) []. Lecture Notes in Mathematics, 387:43– 71, 1974. CODEN LNMAA2. ISBN 3-540-06789-2 (print), 3-540-38397-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057567/.

Bruter:1974:CVMa

[Bru74n] Claude P. Bruter. Chapitre V. Matroïdes de couplage: Transversales. (French) []. Lecture Notes in Mathematics, 387:72–79, 1974. CODEN LNMAA2. ISBN 3-540-06789-2 (print), 3-540-38397-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057568/.

Bruter:1974:CVMb

[Bru74o] Claude P. Bruter. Chapitre VI. Matroïdes binaires, matroïdes orientables, matroïdes graphiques, le critère d’extremalité. (French) []. Lecture Notes in Mathematics, 387:80–90, 1974. CODEN LNMAA2. ISBN 3-540-06789-2 (print), 3-540-38397-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057569/.

Bruter:1974:CVMc

[Bru74p] Claude P. Bruter. Chapitre VII. Matroïdes et problémes d’optimisation. (French) []. Lecture Notes in Mathematics, 387: 91–109, 1974. CODEN LNMAA2. ISBN 3-540-06789-2 (print), 3-540-38397-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057570/.

Bruter:1974:ETM

[Bru74q] Claude P. Bruter. Eléments de la Théorie des Matroïdes, volume 387 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06789-2 (print), 3-540-38397-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 141 pp. LCCN QA3 .L28 no. 387. URL http:

//link.springer.com/book/10.1007/BFb0057563;   http:/
/www.springerlink.com/content/978-3-540-38397-0.

Bryant:1974:CSF

[Bry74a] Roger M. Bryant. Characteristic subgroups of free groups. Lecture Notes in Mathematics, 372:141–149, 1974. CO-DEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-662-21571-5_11/;   http://link.
springer.com/chapter/10.1007/BFb0065165/.

Bryce:1974:MVG

[Bry74b] R. A. Bryce. Metabelian varieties of groups. Lecture Notes in Mathematics, 372:150–157, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
662-21571-5_12/;   http://link.springer.com/chapter/
10.1007/BFb0065166/.

Brylawski:1974:RCG

[Bry74c] Thomas H. Brylawski. Reconstructing combinatorial geometries. Lecture Notes in Mathematics, 406:226–235, 1974. CODEN LN-MAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066444/.

Barnette:1971:TTC

[BS71] David W. Barnette and Sherman K. Stein. On a theorem of Tutte, a 4-color theorem and B-sets. Lecture Notes in Mathematics, 186:9–14, 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059418/.

Berger:1973:IAS

[BS73a] C. A. Berger and B. I. Shaw. Intertwining, analytic structure, and the trace norm estimate. Lecture Notes in Mathematics, 345: 1–6, 1973. CODEN LNMAA2. ISBN 3-540-06496-6 (print), 3-540-37808-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058914/.

Bony:1973:SHD

[BS73b] Jean-Michel Bony and Pierre Schapira. Solutions hyperfonctions du problème de Cauchy. (French) []. Lecture Notes in Mathematics, 287:82–98, 1973. CODEN LNMAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068146/.

Buchi:1973:AMS

[BS73c] J. Richard Büchi and Dirk Siefkes. Axiomatization of the monadic second order theory of ω_{1}. Lecture Notes in Mathematics, 328:129–217, 1973. CODEN LNMAA2. ISBN 3-540-06345-5 (print), 3-540-46946-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082722/.

Baumgarte:1974:ETI

[BS74a] J. Baumgarte and E. Stiefel. Examples of transformations improving the numerical accuracy of the integration of differential equations. Lecture Notes in Mathematics, 362:207–236, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066593/.

Berge:1974:CND

[BS74b] C. Berge and M. Simonovits. The coloring numbers of the direct product of two hypergraphs. Lecture Notes in Mathematics, 411: 21–33, 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066176/.

Bertoluzza:1974:ITC

[BS74c] Carlo Bertoluzza and Michel Schneider. Informations totalement composables. (French) []. Lecture Notes in Mathematics, 398: 90–98, 1974. CODEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065770/.

Bass:1973:MRG

[BT73] H. Bass and J. Tate. The Milnor ring of a global field. Lecture Notes in Mathematics, 342:347–446, 1973. CODEN LN-MAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073733/.

Brocker:1970:BKG

[BtD70a] Theodor Bröcker and Tammo tom Dieck. Bordismus und Kobordismus. (German) []. Lecture Notes in Mathematics, 178:184– 191, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-16378-8_12/.

Brocker:1970:CKG

[BtD70b] Theodor Bröcker and Tammo tom Dieck. Charakteristische Klassen. (German) [Characteristic classes]. Lecture Notes in Mathematics, 178:83–107, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-16378-8_6/.

Brocker:1970:CZG

[BtD70c] Theodor Bröcker and Tammo tom Dieck. Charakteristische Zahlen. (German) []. Lecture Notes in Mathematics, 178:145– 169, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-16378-8_10/.

Brocker:1970:DBH

[BtD70d] Theodor Bröcker and Tammo tom Dieck. Darstellung von Bordismengruppen als Homotopiegruppen. (German) []. Lecture Notes in Mathematics, 178:24–41, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-16378-8_3/.

Brocker:1970:BHT

[BtD70e] Theodor Bröcker and Tammo tom Dieck. Die Bordismen-Homologie-Theorie. (German) []. Lecture Notes in Mathematics, 178:10–23, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-16378-8_2/.

Brocker:1970:FGG

[BtD70f] Theodor Bröcker and Tammo tom Dieck. Formale Gruppen. (German) [Formal groups]. Lecture Notes in Mathematics, 178: 108–122, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-16378-8_7/.

Brocker:1970:FM

[BtD70g] Theodor Bröcker and Tammo tom Dieck. Front matter. Lecture Notes in Mathematics, 178:i–xvi, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/content/pdf/bfm:978-3-662-16378-8/1.pdf.

Brocker:1970:KGC

[BtD70h] Theodor Bröcker and Tammo tom Dieck. Kobordismentheorie. (German) [Cobordism theory], volume 178 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 191 pp. LCCN QA3 .L28

no. 178. URL http://link.springer.com/book/10.1007/978-
3-662-16378-8;   http://www.springerlink.com/content/
978-3-662-16378-8.

Brocker:1970:MTG

[BtD70i] Theodor Bröcker and Tammo tom Dieck. Multiplikative Transformationen. (German) [Multiplicative transformations]. Lecture Notes in Mathematics, 178:123–132, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-16378-8_8/.

Brocker:1970:SHK

[BtD70j] Theodor Bröcker and Tammo tom Dieck. Spektren, Homologie und Kohomologie. (German) []. Lecture Notes in Mathematics, 178:42–68, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-16378-8_4/.

Brocker:1970:SOG

[BtD70k] Theodor Bröcker and Tammo tom Dieck. Stabile Operationen. (German) [Stable operations]. Lecture Notes in Mathematics, 178:170–183, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-16378-8_11/.

Brocker:1970:SOK

[BtD70l] Theodor Bröcker and Tammo tom Dieck. Steenrod-Operationen in der Kobordismentheorie. (German) []. Lecture Notes in Mathematics, 178:133–144, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-16378-8_9/.

Brocker:1970:VKL

[BtD70m] Theodor Bröcker and Tammo tom Dieck. Verträglichkeit der Kohomologie mit dem Limes. (German) []. Lecture Notes in Mathematics, 178:69–82, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-16378-8_5/.

Brocker:1970:VGP

[BtD70n] Theodor Bröcker and Tammo tom Dieck. Vorbereitungen. (German) [Preparations]. Lecture Notes in Mathematics, 178:1– 9, 1970. CODEN LNMAA2. ISBN 3-540-05341-7 (print), 3-662-16378-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-16378-8_1/.

Bucur:1972:SAF

[Buc72] I. Bucur. Some applications of the formalism of duality in algebraic geometry. Lecture Notes in Mathematics, 274:137– 158, 1972. CODEN LNMAA2. ISBN 3-540-05920-2 (print), 3-540-37609-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073968/.

Buchi:1973:MSO

[Büc73a] J. Richard Büchi. The monadic second order theory of ω_{1}. Lecture Notes in Mathematics, 328:1–127, 1973. CODEN LN-MAA2. ISBN 3-540-06345-5 (print), 3-540-46946-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082721/.

Buchwalter:1973:FCM

[Buc73b] Henri Buchwalter. Fonctions continues et mesures sur un espace complètement régulier. (French) []. Lecture Notes in Mathematics, 331:183–202, 1973. CODEN LNMAA2. ISBN 3-540-06367-6 (print), 3-540-46977-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068230/.

Bulirsch:1972:LG

[Bul72] Roland Bulirsch. Über eine Linearisierung von F (z_{xx}, z_{xy}, z_{yy}) = 0. (German) []. Lecture Notes in Mathematics, 267:15–21, 1972. CODEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061611/.

Bunge:1971:BIA

[Bun71] Marta C. Bunge. Bifibration induced adjoint pairs. Lecture Notes in Mathematics, 195:70–122, 1971. CODEN LNMAA2. ISBN 3-540-05442-1 (print), 3-540-36548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0072308/.

Bunke:1972:SPS

[Bun72] Helga Bunke. Stable periodic solutions of weakly nonlinear stochastic differential equations. Lecture Notes in Mathematics, 294:283–290, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064948/.

Burkholder:1971:MI

[Bur71] D. L. Burkholder. Martingale inequalities. Lecture Notes in Mathematics, 190:1–8, 1971. CODEN LNMAA2. ISBN 3-540-05396-4 (print), 3-540-36515-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065884/.

Burke:1974:NRH

[Bur74a] Dennis K. Burke. A note on R. H. Bing’s example G. Lecture Notes in Mathematics, 375:47–52, 1974. CODEN LN-MAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064010/.

Burke:1974:SD

[Bur74b] Dennis K. Burke. Spaces with a G_{δ}-diagonal. Lecture Notes in Mathematics, 378:95–100, 1974. CODEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068462/.

Burnham:1974:SAD

[Bur74c] James T. Burnham. Segal algebras and dense ideals in Banach algebras. Lecture Notes in Mathematics, 399:33–58, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063565/.

Burns:1974:RIS

[Bur74d] R. G. Burns. On the rank of the intersection of subgroups of a Fuchsian group. Lecture Notes in Mathematics, 372: 165–187, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21571-5_14/;
http://link.springer.com/chapter/10.1007/BFb0065168/.

See corrigenda [Bur90].

Burr:1974:GRT

[Bur74e] Stefan A. Burr. Generalized Ramsey theory for graphs — a survey. Lecture Notes in Mathematics, 406:52–75, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066435/.

Burns:1990:CPR

[Bur90] R. G. Burns. Corrigenda to the paper ‘On the rank of the intersection of subgroups of a Fuchsian group’. Lecture Notes in Mathematics, 1456:198, 1990. CODEN LNMAA2. ISBN 3-540-53475-X (print), 3-540-46900-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0100742. See [Bur74d].

Butler:1971:BR

[But71] Kim Ki-Hang Butler. Binary relations. Lecture Notes in Mathematics, 186:25–47, 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059420/.

Butler:1972:NPO

[But72] Kim Ki-Hang Butler. The number of partial order graphs. Lecture Notes in Mathematics, 303:55–63, 1972. CODEN LN-MAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067357/.

Butcher:1974:ODE

[But74a] J. C. Butcher. The order of differential equation methods. Lecture Notes in Mathematics, 362:72–75, 1974. CODEN LN-MAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066585/.

Butler:1974:MPI

[But74b] Kim Ki-Hang Butler. A Moore–Penrose inverse for Boolean relation matrices. Lecture Notes in Mathematics, 403:18–28, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057372/.

Butler:1974:SBR

[But74c] Kim Ki-Hang Butler. Subgroups of binary relations. Lecture Notes in Mathematics, 372:188–196, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (ebook), 3-662-21571-3 (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-662-21571-5_15/;   http://link.
springer.com/chapter/10.1007/BFb0065169/.

Butler:1974:ARK

[But74d] M. C. R. Butler. The 2-adic representations of Klein’s four group. Lecture Notes in Mathematics, 372:197–203, 1974. CO-DEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-662-21571-5_16/;   http://link.
springer.com/chapter/10.1007/BFb0065170/.

Boardman:1973:BCT

[BV73a] J. M. Boardman and R. M. Vogt. The bar construction for theories. Lecture Notes in Mathematics, 347:72–94, 1973. CODEN LNMAA2. ISBN 3-540-06479-6 (print), 3-540-37799-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068550/.

Boardman:1973:HC

[BV73b] J. M. Boardman and R. M. Vogt. Homotopy colimits. Lecture Notes in Mathematics, 347:218–227, 1973. CODEN LN-MAA2. ISBN 3-540-06479-6 (print), 3-540-37799-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068554/.

Boardman:1973:HH

[BV73c] J. M. Boardman and R. M. Vogt. Homotopy homomorphisms. Lecture Notes in Mathematics, 347:95–154, 1973. CODEN LN-MAA2. ISBN 3-540-06479-6 (print), 3-540-37799-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068551/.

Boardman:1973:HIA

[BV73d] J. M. Boardman and R. M. Vogt. Homotopy Invariant Algebraic Structures on Topological Spaces, volume 347 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06479-6 (print), 3-540-37799-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 257 pp. LCCN QA3 .L28 no. 347. URL http:/

/link.springer.com/book/10.1007/BFb0068547;   http://
www.springerlink.com/content/978-3-540-37799-3.

Boardman:1973:ILS

[BV73e] J. M. Boardman and R. M. Vogt. Iterated loop spaces and actions on classifying spaces. Lecture Notes in Mathematics, 347: 174–217, 1973. CODEN LNMAA2. ISBN 3-540-06479-6 (print), 3-540-37799-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068553/.

Boardman:1973:MHS

[BV73f] J. M. Boardman and R. M. Vogt. Motivation and historical survey. Lecture Notes in Mathematics, 347:1–26, 1973. CODEN LNMAA2. ISBN 3-540-06479-6 (print), 3-540-37799-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068548/.

Boardman:1973:SBS

[BV73g] J. M. Boardman and R. M. Vogt. Structures on based spaces. Lecture Notes in Mathematics, 347:155–173, 1973. CODEN LN-MAA2. ISBN 3-540-06479-6 (print), 3-540-37799-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068552/.

Boardman:1973:TAT

[BV73h] J. M. Boardman and R. M. Vogt. Topological-algebraic theories. Lecture Notes in Mathematics, 347:27–71, 1973. CODEN LNMAA2. ISBN 3-540-06479-6 (print), 3-540-37799-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068549/.

Barth:1974:GCG

[BV74] W. Barth and A. Van de Ven. On the geometry in codimension 2 of Grassmann manifolds. Lecture Notes in Mathematics, 412: 1–35, 1974. CODEN LNMAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066152/.

Bekic:1971:FSP

[BW71] H. Bekić and K. Walk. Formalization of storage properties. Lecture Notes in Mathematics, 188:28–61, 1971. CODEN LN-MAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059692/.

Brockett:1972:AVC

[BW72] Roger W. Brockett and Jan C. Willems. Average value criteria for stochastic stability. Lecture Notes in Mathematics, 294: 252–272, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064946/.

Bernstein:1974:CDP

[BW74] Allen R. Bernstein and Frank Wattenberg. Cardinalitydependent properties of topological spaces. Lecture Notes in Mathematics, 369:50–59, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065999/.

Birtel:1971:SAF

[BZ71] F. Birtel and W. Zame. Some analytic function algebras. Lecture Notes in Mathematics, 184:1–9, 1971. CODEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059275/.

Cohen:1974:NES

[CA74] Leon Cohen and Afaq Ahmad. Numerical experiments on the statistics of the gravitational field. Lecture Notes in Mathematics, 362:337–359, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066600/.

Cadogan:1971:CCG

[Cad71] C. C. Cadogan. On connected cubic graphs and trivalent trees. Lecture Notes in Mathematics, 186:49–54, 1971. CODEN LN-MAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059421/.

Cairoli:1970:IPM

[Cai70] R. Cairoli. Une inégalité pour martingales à indices multiples et ses applications. (French) []. Lecture Notes in Mathematics, 124:

1–27, 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059329/.

Cairoli:1971:DPI

[Cai71] R. Cairoli. Decomposition de processus à indices doubles. (French) []. Lecture Notes in Mathematics, 191:37–57, 1971. CO-DEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058845/.

Cannon:1974:GPG

[Can74] John Cannon. A general purpose group theory program. Lecture Notes in Mathematics, 372:204–217, 1974. CO-DEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic). URL http://link.springer.com/
chapter/10.1007/978-3-662-21571-5_17/;   http://link.
springer.com/chapter/10.1007/BFb0065171/.

Capobianco:1972:ECG

[Cap72] Michael Capobianco. Estimating the connectivity of a graph. Lecture Notes in Mathematics, 303:65–74, 1972. CODEN LN-MAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067358/.

Cappell:1973:GSH

[Cap73a] Sylvain E. Cappell. Groups of singular Hermitian forms. Lecture Notes in Mathematics, 343:513–525, 1973. CODEN LN-MAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061377/.

Cappell:1973:MVS

[Cap73b] Sylvain E. Cappell. Mayer–Vietoris sequences in Hermitian Ktheory. Lecture Notes in Mathematics, 343:478–512, 1973. CO-DEN LNMAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061376/.

Carter:1970:CCW

[Car70a] R. Carter. Conjugacy classes in the Weyl group. Lecture Notes in Mathematics, 131:297–318, 1970. CODEN LNMAA2. ISBN 3-540-04920-7 (print), 3-540-36272-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081548/.

Cartier:1970:CVA

[Car70b] P. Cartier. Sur certaines variables aléatoires associées au réarrangement croissant d’un échantillon. (French) []. Lecture Notes in Mathematics, 124:28–36, 1970. CODEN LN-MAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059330/.

Cartan:1971:SEA

[Car71a] Henri Cartan. Sous-ensembles analytiques d’une variété banachique complexe (d’après J.-P. Ramis). (French) []. Lecture Notes in Mathematics, 179:123–138, 1971. CODEN LN-MAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058809/.

Cartier:1971:IEM

[Car71b] P. Cartier. Introduction à l’étude des mouvements browniens à plusieurs paramètres. (French) []. Lecture Notes in Mathematics, 191:58–75, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058846/.

Cartier:1971:EPG

[Car71c] Pierre Cartier. Espaces de Poisson des groupes localement compacts. (French) []. Lecture Notes in Mathematics, 180:107–127, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058827/.

Cartier:1971:RGF

[Car71d] Pierre Cartier. Relèvements des groupes formels commutatifs. (French) []. Lecture Notes in Mathematics, 179:217–230, 1971.

CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058814/.

Cartwright:1971:APS

[Car71e] M. L. Cartwright. Almost periodic solutions of differential equations and flows. Lecture Notes in Mathematics, 235:35–43, 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059192/.

Carlitz:1972:ENO

[Car72a] L. Carlitz. Eulerian numbers and operators. Lecture Notes in Mathematics, 251:65–70, 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058785/.

Carrell:1972:HIC

[Car72b] James B. Carrell. Holomorphically injective complex toral actions. Lecture Notes in Mathematics, 299:205–236, 1972. CODEN LNMAA2. ISBN 3-540-06078-2 (print), 3-540-38066-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066765/.

Cartier:1972:PMT

[Car72c] Pierre Cartier. Problèmes mathématiques de la théorie quantique des champs. (French) []. Lecture Notes in Mathematics, 244:107–122, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058699/.

Carroll:1973:TLF

[Car73a] Joseph E. Carroll. On the torsion in K_{2} of local fields. Lecture Notes in Mathematics, 342:464–473, 1973. CODEN LN-MAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073735/.

Cartier:1973:GAA

[Car73b] Pierre Cartier. Géométrie et analyse sur les arbres. (French) []. Lecture Notes in Mathematics, 317:123–140, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069280/.

Cartier:1974:ICM

[Car74a] Pierre Cartier. Inégalités de corrélation en mécanique statistique. (French) []. Lecture Notes in Mathematics, 383:242–264, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057312/.

Cartier:1974:PMT

[Car74b] Pierre Cartier. Problèmes mathématiques de la théorie quantique des champs II: Prolongement analytique. (French) []. Lecture Notes in Mathematics, 383:1–33, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057299/.

Cartwright:1974:BSS

[Car74c] M. L. Cartwright. Boundedness of solutions of second order differential equations. Lecture Notes in Mathematics, 419:67–70, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064711/.

Castaing:1970:TDP

[Cas70] Charles Castaing. Le théorème de Dunford–Pettis généralisé. (French) []. Lecture Notes in Mathematics, 116:133–151, 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059251/.

Casselman:1973:ARR

[Cas73a] W. Casselman. An assortment of results on representations of GL_{2}(k). Lecture Notes in Mathematics, 349:1–54, 1973. CODEN LNMAA2. ISBN 3-540-06558-X (print), 3-540-37855-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-37855-6_
1/;   http://link.springer.com/content/pdf/bfm:978-3-
540-37855-6/1.pdf.

Casselman:1973:RAM

[Cas73b] W. Casselman. On representations of Gl_{2} and the arithmetic of modular curves. Lecture Notes in Mathematics, 349:107– 141, 1973. CODEN LNMAA2. ISBN 3-540-06558-X (print), 3-540-37855-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-37855-6_3/.

Casson:1973:WGF

[Cas73c] A. J. Casson. Whitehead groups of free products with amalgamation. Lecture Notes in Mathematics, 342:144–154, 1973. CODEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073725/.

Cazacu:1974:GQ

[Caz74a] Cabiria Andreian Cazacu. A generalization of the quasiconformality. Lecture Notes in Mathematics, 419:4–17, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064706/.

Cazacu:1974:PDE

[Caz74b] Cabiria Andreian Cazacu. Partial differential equations related to extremal problems for quasiconformal mappings. Lecture Notes in Mathematics, 415:1–14, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065506/.

Curjel:1971:SFD

[CD71] C. R. Curjel and R. R. Douglas. On H-spaces of finite dimension. Lecture Notes in Mathematics, 196:42, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067745.

Cartier:1972:SPV

[CD72a] P. Cartier and C. Dellacherie. Séminaire de Probabilités VI Université de Strasbourg, volume 258 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LN-MAA2. ISBN 0-387-05773-0, 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 253 pp. LCCN QA3 .L28 no. 258. URL http:

//link.springer.com/book/10.1007/BFb0059457;   http:/
/www.springerlink.com/content/978-3-540-37183-0.

Cole:1972:NAR

[CD72b] J. C. Cole and M. A. Dickmann. Non-axiomatizability results in infinitary languages for higher-order structures. Lecture Notes in Mathematics, 255:29–41, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059536/.

Choquet:1974:NCB

[CD74] Gustave Choquet and Jacques Deny. Noyaux de convolution et balayage sur tout ouvert. (French) []. Lecture Notes in Mathematics, 404:60–112, 1974. CODEN LNMAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060610/.

Cea:1970:RNO

[Céa70] Jean Céa. Recherche numérique d’un optimum dans un espace produit. (French) []. Lecture Notes in Mathematics, 112:33– 50, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060196/.

Cerf:1971:PIT

[Cer71] Jean Cerf. The pseudo-isotopy theorem for simply connected differentiable manifolds. Lecture Notes in Mathematics, 197:76– 82, 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068612/.

Cesari:1971:FAB

[Ces71] Lamberto Cesari. Functional analysis and boundary value problems. Lecture Notes in Mathematics, 183:177–194, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060418/.

Cesari:1974:FAP

[Ces74a] Lamberto Cesari. Functional analysis and partial differential equations. Lecture Notes in Mathematics, 419:71–75, 1974. CO-DEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064712/.

Cesari:1974:QAC

[Cés74b] Yves Césari. Questionnaires avec contraintes de rang. (French) []. Lecture Notes in Mathematics, 398:186–193, 1974. CODEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065777/.

Capobianco:1971:RTG

[CFK71] M. Capobianco, J. B. Frechen, and M. Krolik, editors. Recent Trends in Graph Theory: Proceedings of the First New York City Graph Theory Conference held on June 11, 12, and 13, 1970 Sponsored by St. John’s University, Jamaica, New York, volume 186 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 186. URL http:/

/link.springer.com/book/10.1007/BFb0059415;   http://
www.springerlink.com/content/978-3-540-36508-2.

Colton:1971:NRAb

[CG71a] David Colton and Robert P. Gilbert. New results on the approximation of solutions to partial differential equations: Iterative methods. Lecture Notes in Mathematics, 183:213–220, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060422/.

Colton:1971:NRAa

[CG71b] David Colton and Robert P. Gilbert. New results on the approximation of solutions to partial differential equations: The method of particular solutions. Lecture Notes in Mathematics, 183:207– 212, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060421/.

Colton:1974:CCM

[CG74] Prof. Dr. David Lem Colton and Prof. Dr. Robert Pertsch Gilbert, editors. Constructive and Computational Methods for Differential and Integral Equations: Symposium, Indiana University February 17–20, 1974, volume 430 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 430.

URL http://link.springer.com/book/10.1007/BFb0066262;
 http://www.springerlink.com/content/978-3-540-37302-
5.

Cea:1971:MFN

[CGN71] J. Céa, R. Glowinski, and J. C. Nédélec. Minimisation de fonctionnelles non-différentiables. (French) [minimization of nondifferentiable functionals]. Lecture Notes in Mathematics, 228:19– 38, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069445/.

Caretto:1971:RIF

[CGS71] L. S. Caretto, A. D. Gosman, and D. B. Spalding. Removal of an instability in a free convection problem. Lecture Notes in Mathematics, 228:264–269, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069460/.

Hsiang:1973:DFL

[cH73] Wu chung Hsiang. Decomposition formula of Laurent extension in algebraic K-theory and the role of codimension 1 submanifold in topology. Lecture Notes in Mathematics, 342:308–327, 1973. CODEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073730/.

Chacon:1970:ASM

[Cha70] R. V. Chacon. Approximation and spectral multiplicity. Lecture Notes in Mathematics, 160:18–27, 1970. CODEN LN-MAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060642/.

Chatterji:1972:PSS

[Cha72] S. D. Chatterji. Un principe de sous-suites dans la théorie des probabilités. (French) []. Lecture Notes in Mathematics, 258:72– 89, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059462/.

Chatterji:1973:MLA

[Cha73] S. D. Chatterji. Les martingales et leurs applications analytiques. (French) []. Lecture Notes in Mathematics, 307:27–164, 1973. CODEN LNMAA2. ISBN 3-540-06126-6 (print), 3-540-38321-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059708/.

Chang:1974:ASD

[Cha74a] Y. F. Chang. Automatic solution of differential equations. Lecture Notes in Mathematics, 430:61–94, 1974. CODEN LN-MAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066265/.

Chapman:1974:CLM

[Cha74b] T. A. Chapman. Cell-like mappings. Lecture Notes in Mathematics, 428:230–240, 1974. CODEN LNMAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070541/.

Chapman:1974:THC

[Cha74c] T. A. Chapman. On the theory of Hilbert cube manifolds. Lecture Notes in Mathematics, 378:101–107, 1974. CODEN LN-MAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068463/.

Chapman:1974:SHT

[Cha74d] T. A. Chapman. Simple homotopy theory for compact Hilbert cube manifold factors. Lecture Notes in Mathematics, 375:53– 63, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064011/.

Chazarain:1974:PMH

[Cha74e] Jacques Chazarain. Le problème mixte hyperbolique. (French) []. Lecture Notes in Mathematics, 383:265–285, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057313/.

Chernousko:1970:OAS

[Che70] F. L. Chernousko. On optimal algorithms for search. Lecture Notes in Mathematics, 112:95–103, 1970. CODEN LN-MAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060199/.

Chebli:1974:OTG

[Che74a] Houcine Chebli. Opérateurs de translation généralisée et semigroupes de convolution. (French) []. Lecture Notes in Mathematics, 404:35–59, 1974. CODEN LNMAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060609/.

Chenciner:1974:TTM

[Che74b] Alain Chenciner. Travaux de Thom et Mather sur la stabilité topologique. (French) []. Lecture Notes in Mathematics, 383:116– 140, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057305/.

Cheniot:1974:TVK

[Che74c] Denis Cheniot. Le théorème de Van Kampen sur le groupe fondamental du complémentaire d’une courbe algébrique projective plane. (French) []. Lecture Notes in Mathematics, 409: 394–417, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068112/.

Chevalley:1974:TBF

[Che74d] Claude Chevalley. Théorie des blocs. (French) []. Lecture Notes in Mathematics, 383:34–49, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057300/.

Chevet:1974:CPR

[Che74e] Simone Chevet. Compléments sur les probabilités de Radon Gaussiennes. (French) [Complements of Gaussian Radon probabilities]. Lecture Notes in Mathematics, 379:337–376, 1974. CO-DEN LNMAA2. ISBN 3-540-06843-0 (print), 3-540-37796-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060502/.

Chevet:1974:GBE

[Che74f] Simone Chevet. G. B. ensembles, G. C. ensembles. Épaisseur mixte et ϵ-entropie. (French) []. Lecture Notes in Mathematics, 379:243–336, 1974. CODEN LNMAA2. ISBN 3-540-06843-0 (print), 3-540-37796-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060501/.

Chevet:1974:QEE

[Che74g] Simone Chevet. Quelques exemples d’ensembles bornés et d’ensembles d’uniforme continuité. (French) []. Lecture Notes in Mathematics, 379:222–242, 1974. CODEN LNMAA2. ISBN 3-540-06843-0 (print), 3-540-37796-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060500/.

Chillingworth:1971:PSD

[Chi71a] David Chillingworth, editor. Proceedings of the Symposium on Differential Equations and Dynamical Systems: University of Warwick September 1968–August 1969 Summer School, July 15– 25, 1969, volume 206 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 206. URL http:/

/link.springer.com/book/10.1007/BFb0070117;   http://
www.springerlink.com/content/978-3-540-36662-1.

Chinn:1971:FPG

[Chi71b] Phyllis Zweig Chinn. The frequency partition of a graph. Lecture Notes in Mathematics, 186:69–70, 1971. CODEN LN-MAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059424.

Chinn:1971:GPE

[Chi71c] Phyllis Zweig Chinn. A graph with p points and enough distinct (p − 2)-order subgraphs is reconstructible. Lecture Notes in Mathematics, 186:71–73, 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059425/.

Chibisov:1973:AEN

[Chi73] D. M. Chibisov. Asymptotic expansions for Neyman’s C(α) tests. Lecture Notes in Mathematics, 330:16–45, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061479/.

Chow:1970:RIS

[Cho70] Shui-Nee Chow. Remarks on integral stability. Lecture Notes in Mathematics, 144:30–33, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059917/.

Chollet:1973:CFA

[Cho73a] Anne-Marie Chollet. Sur des classes de fonctions analytiques dans le disque et indefiniment derivables à la frontière. (French) []. Lecture Notes in Mathematics, 336:144–147, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065795/.

Chou:1973:RIF

[Cho73b] Chin-Cheng Chou. La regularité intérieure. (French) []. Lecture Notes in Mathematics, 325:101–119, 1973. CODEN LN-MAA2. ISBN 3-540-06301-3 (print), 3-540-46910-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060983/.

Chou:1973:TFC

[Cho73c] Chin-Cheng Chou. La Transformation de Fourier Complexe et L’équation de Convolution, volume 325 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CO-DEN LNMAA2. ISBN 3-540-06301-3 (print), 3-540-46910-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 137 pp. LCCN QA3 .L28 no. 325. URL http:/

/link.springer.com/book/10.1007/BFb0060979;   http://
www.springerlink.com/content/978-3-540-46910-0.

Chou:1973:EUF

[Cho73d] Chin-Cheng Chou. Les espaces de M_{(P)}-ultradistributions. (French) []. Lecture Notes in Mathematics, 325:1–45, 1973. CO-DEN LNMAA2. ISBN 3-540-06301-3 (print), 3-540-46910-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060980/.

Chou:1973:LF

[Cho73e] Chin-Cheng Chou. L’inversibilité. (French) []. Lecture Notes in Mathematics, 325:60–100, 1973. CODEN LNMAA2. ISBN 3-540-06301-3 (print), 3-540-46910-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060982/.

Chou:1973:OHF

[Cho73f] Chin-Cheng Chou. Opérateur hyperbolique. (French) []. Lecture Notes in Mathematics, 325:120–134, 1973. CODEN LN-MAA2. ISBN 3-540-06301-3 (print), 3-540-46910-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060984/.

Chou:1973:MMF

[Cho73g] Chin-Cheng Chou. Sur le module minimum des fonctions analytiques complexes. (French) []. Lecture Notes in Mathematics, 325: 46–59, 1973. CODEN LNMAA2. ISBN 3-540-06301-3 (print), 3-540-46910-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060981/.

Carlitz:1972:SFR

[CHS72] L. Carlitz, V. E. Hoggatt, and Richard Scoville. Some functions related to Fibonacci and Lucas representations. Lecture Notes in Mathematics, 251:71–102, 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058786/.

Chu:1971:TDC

[Chu71a] Hsin Chu. Topological dynamics and compact transformation groups. Lecture Notes in Mathematics, 235:44–53, 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059193/.

Chung:1971:SPD

[Chu71b] K. L. Chung. A simple proof of Doob’s convergence theorem. Lecture Notes in Mathematics, 191:76, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058847.

Chung:1972:SUF

[Chu72] Kai Lai Chung. Some universal field equations. Lecture Notes in Mathematics, 258:90–97, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059463/.

Chu:1973:EPH

[Chu73] Hsin Chu. On the embedding problem and the Hilbert–Smith conjecture. Lecture Notes in Mathematics, 318:78–85, 1973. CO-DEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061724/.

Chung:1974:TCR

[Chu74a] Fan Chung. On triangular and cyclic Ramsey numbers with k colors. Lecture Notes in Mathematics, 406:236–242, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066445/.

Chung:1974:NLE

[Chu74b] K. L. Chung. Note on last exit decomposition. Lecture Notes in Mathematics, 381:20–21, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057252.

Chvatal:1974:IFE

[Chv74a] V. Chvátal. Intersecting families of edges in hypergraphs having the hereditary property. Lecture Notes in Mathematics, 411: 61–66, 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066179/.

Chvatal:1974:MMG

[Chv74b] V. Chvátal. The minimality of the Mycielski graph. Lecture Notes in Mathematics, 406:243–246, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066446/.

Hsiang:1970:FTF

[cHyH70] Wu chung Hsiang and Wu yi Hsiang. A fixed theorem for finite diffeomorphism groups generated by reflections. Lecture Notes in Mathematics, 168:90–106, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058520/.

Ciarlet:1974:CNF

[Cia74] P. G. Ciarlet. Conforming and nonconforming finite element methods for solving the plate problem. Lecture Notes in Mathematics, 363:21–31, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069122/.

Carter:1974:MRG

[CL74a] R. W. Carter and G. Lusztig. On the modular representations of the general linear and symmetric groups. Lecture Notes in Mathematics, 372:218–220, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
662-21571-5_18/;   http://link.springer.com/chapter/
10.1007/BFb0065172/.

Chvatal:1974:EDG

[CL74b] V. Chvátal and L. Lovász. Every directed graph has a semikernel. Lecture Notes in Mathematics, 411:175, 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066192.

Clark:1971:CCA

[Cla71a] Allan Clark. Categorical constructions in algebraic topology. Lecture Notes in Mathematics, 196:68–73, 1971. CODEN LN-MAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067749/.

Clayton:1971:LCA

[Cla71b] Dennis Clayton. A local characterization of analytic structure in a commutative Banach algebra. Lecture Notes in Mathematics, 184:10–36, 1971. CODEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059276/.

Choo:1973:FPR

[CLL73] K. G. Choo, K. Y. Lam, and E. Luft. On free product of rings and the coherence property. Lecture Notes in Mathematics, 342: 135–143, 1973. CODEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073724/.

Chartrand:1971:GTN

[CM71] Gary Chartrand and John Mitchem. Graphical theorems of the Nordhaus–Gaddum class. Lecture Notes in Mathematics, 186: 55–61, 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059422/.

Chawla:1972:CES

[CM72] L. M. Chawla and John E. Maxfield. On the classification and evaluation of some partition functions and their tables. Lecture Notes in Mathematics, 251:103–112, 1972. CODEN LN-MAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058787/.

Chernoff:1974:PID

[CM74a] Dr. Paul Robert Chernoff and Dr. Jerrold Eldon Marsden. Properties of Infinite Dimensional Hamiltonian Systems, volume 425 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-07011-7 (print), 3-540-37287-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 160 pp. LCCN QA3 .L28 no. 425. URL http:

//link.springer.com/book/10.1007/BFb0073665;   http:/
/www.springerlink.com/content/978-3-540-37287-5.

Chernoff:1974:HS

[CM74b] P. R. Chernoff and J. E. Marsden. Hamiltonian systems. Lecture Notes in Mathematics, 425:72–99, 1974. CODEN LN-MAA2. ISBN 3-540-07011-7 (print), 3-540-37287-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073670/.

Chernoff:1974:LS

[CM74c] P. R. Chernoff and J. E. Marsden. Lagrangian systems. Lecture Notes in Mathematics, 425:100–124, 1974. CODEN LN-MAA2. ISBN 3-540-07011-7 (print), 3-540-37287-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073671/.

Chernoff:1974:LHS

[CM74d] P. R. Chernoff and J. E. Marsden. Linear Hamiltonian systems. Lecture Notes in Mathematics, 425:16–55, 1974. CODEN LNMAA2. ISBN 3-540-07011-7 (print), 3-540-37287-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073668/.

Chernoff:1974:SBP

[CM74e] P. R. Chernoff and J. E. Marsden. Some basic properties of infinite dimensional Hamiltonian systems. Lecture Notes in Mathematics, 425:1–2, 1974. CODEN LNMAA2. ISBN 3-540-07011-7 (print), 3-540-37287-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073666.

Chernoff:1974:SGP

[CM74f] P. R. Chernoff and J. E. Marsden. Some general properties of nonlinear semigroups. Lecture Notes in Mathematics, 425:56– 71, 1974. CODEN LNMAA2. ISBN 3-540-07011-7 (print), 3-540-37287-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073669/.

Chernoff:1974:SGC

[CM74g] P. R. Chernoff and J. E. Marsden. Symmetry groups and conservation laws. Lecture Notes in Mathematics, 425:125–153, 1974. CODEN LNMAA2. ISBN 3-540-07011-7 (print), 3-540-37287-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073672/.

Chernoff:1974:SSH

[CM74h] P. R. Chernoff and J. E. Marsden. Symplectic structures and Hamiltonian systems. Lecture Notes in Mathematics, 425:3– 15, 1974. CODEN LNMAA2. ISBN 3-540-07011-7 (print), 3-540-37287-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073667/.

Coburn:1973:TOO

[Cob73] L. A. Coburn. Toeplitz operators on odd spheres. Lecture Notes in Mathematics, 345:7–12, 1973. CODEN LNMAA2. ISBN 3-540-06496-6 (print), 3-540-37808-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058915/.

Coddington:1971:MVO

[Cod71] Earl A. Coddington. Multi-valued operators and boundary value problems. Lecture Notes in Mathematics, 183:1–8, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060405/.

Coeure:1970:PAF

[Coe70] Gérard Coeuré. Prolongement analytique. (French) [Analytic extension]. Lecture Notes in Mathematics, 116:128–132, 1970. CO-DEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059250/.

Coeure:1974:PAD

[Coe74] G. Coeuré. Prolongement analytique en dimension infinie. (French) []. Lecture Notes in Mathematics, 384:1–19, 1974. CO-DEN LNMAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057539/.

Cohen:1970:EMS

[Coh70a] Joel M. Cohen. Eilenberg–MacLane spaces and spectra. Lecture Notes in Mathematics, 165:31–51, 1970. CODEN LN-MAA2. ISBN 3-540-05192-9 (print), 3-540-36393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059786/.

Cohen:1970:HHL

[Coh70b] Joel M. Cohen. Homotopy and homology not-so-long exact sequences. Lecture Notes in Mathematics, 165:5–30, 1970. CODEN LNMAA2. ISBN 3-540-05192-9 (print), 3-540-36393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059785/.

Cohen:1970:P

[Coh70c] Joel M. Cohen. Preliminaries. Lecture Notes in Mathematics, 165:1–4, 1970. CODEN LNMAA2. ISBN 3-540-05192-9 (print), 3-540-36393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059784/.

Cohen:1970:SWD

[Coh70d] Joel M. Cohen. Spanier–Whitehead duality. Lecture Notes in Mathematics, 165:52–75, 1970. CODEN LNMAA2. ISBN 3-540-05192-9 (print), 3-540-36393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059787/.

Cohen:1970:SSC

[Coh70e] Joel M. Cohen. Spectral sequences and calculation of stable homotopy. Lecture Notes in Mathematics, 165:141–192, 1970. CO-DEN LNMAA2. ISBN 3-540-05192-9 (print), 3-540-36393-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059789/.

Cohen:1970:SH

[Coh70f] Joel M. Cohen. Stable homotopy, volume 165 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CO-DEN LNMAA2. ISBN 3-540-05192-9 (print), 3-540-36393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 194 pp. LCCN QA3 .L28 no. 165. URL http:/

/link.springer.com/book/10.1007/BFb0059783;   http://
www.springerlink.com/content/978-3-540-36393-4.

Cohen:1970:SHC

[Coh70g] Joel M. Cohen. Stable homotopy in categorical algebra. Lecture Notes in Mathematics, 165:76–140, 1970. CODEN LN-MAA2. ISBN 3-540-05192-9 (print), 3-540-36393-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059788/.

Cohen:1972:AI

[Coh72a] Daniel E. Cohen. The augmentation ideal. Lecture Notes in Mathematics, 245:57–66, 1972. CODEN LNMAA2. ISBN 3-540-05759-5 (print), 3-540-37180-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069649/.

Cohen:1972:CT

[Coh72b] Daniel E. Cohen. Cohomology theory. Lecture Notes in Mathematics, 245:1–16, 1972. CODEN LNMAA2. ISBN 3-540-05759-5 (print), 3-540-37180-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069646/.

Cohen:1972:CC

[Coh72c] Daniel E. Cohen. The countable case. Lecture Notes in Mathematics, 245:69–73, 1972. CODEN LNMAA2. ISBN 3-540-05759-5 (print), 3-540-37180-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069651/.

Cohen:1972:E

[Coh72d] Daniel E. Cohen. Ends. Lecture Notes in Mathematics, 245: 17–40, 1972. CODEN LNMAA2. ISBN 3-540-05759-5 (print), 3-540-37180-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069647/.

Cohen:1972:FGC

[Coh72e] Daniel E. Cohen. The finitely generated case. Lecture Notes in Mathematics, 245:67–68, 1972. CODEN LNMAA2. ISBN 3-540-05759-5 (print), 3-540-37180-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069650.

Cohen:1972:GCD

[Coh72f] Daniel E. Cohen. Groups of Cohomological Dimension One, volume 245 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05759-5 (print), 3-540-37180-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 99 pp. LCCN QA3 .L28 no. 245. URL http:/

/link.springer.com/book/10.1007/BFb0069645;   http://
www.springerlink.com/content/978-3-540-37180-9.

Cohen:1972:MT

[Coh72g] Daniel E. Cohen. The main theorems. Lecture Notes in Mathematics, 245:79–84, 1972. CODEN LNMAA2. ISBN 3-540-05759-5 (print), 3-540-37180-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069653/.

Cohen:1972:STb

[Coh72h] Daniel E. Cohen. Splitting theorems. Lecture Notes in Mathematics, 245:74–78, 1972. CODEN LNMAA2. ISBN 3-540-05759-5 (print), 3-540-37180-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069652/.

Cohen:1972:STa

[Coh72i] Daniel E. Cohen. The structure theorem. Lecture Notes in Mathematics, 245:41–56, 1972. CODEN LNMAA2. ISBN 3-540-05759-5 (print), 3-540-37180-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069648/.

Cohn:1972:SFF

[Coh72j] P. M. Cohn. Skew fields of fractions, and the prime spectrum of a general ring. Lecture Notes in Mathematics, 246:1–71, 1972.

CODEN LNMAA2. ISBN 3-540-05760-9 (print), 3-540-37181-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059563/.

Cohen:1973:GFS

[Coh73a] Daniel E. Cohen. Group with free subgroups of finite index. Lecture Notes in Mathematics, 319:26–44, 1973. CODEN LN-MAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058927/.

Cohen:1973:MSN

[Coh73b] Donald S. Cohen. Multiple solutions of nonlinear partial differential equations. Lecture Notes in Mathematics, 322:15–77, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060560/.

Cohn:1974:APR

[Coh74a] Dr. Leslie Cohn. Algebraic properties of the C-ring. Lecture Notes in Mathematics, 429:94–96, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064352/.

Cohn:1974:ATH

[Coh74b] Dr. Leslie Cohn. Analytic Theory of the Harish-Chandra C-Function, volume 429 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 154 pp. LCCN QA3 .L28 no. 429. URL http:/

/link.springer.com/book/10.1007/BFb0064335;   http://
www.springerlink.com/content/978-3-540-37298-1.

Cohn:1974:ACH

[Coh74c] Dr. Leslie Cohn. An application of the Campbell–Hausdorff formula. Lecture Notes in Mathematics, 429:17–20, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064339/.

Cohn:1974:ADE

[Coh74d] Dr. Leslie Cohn. Application of the differential equations. Lecture Notes in Mathematics, 429:39–41, 1974. CODEN LN-MAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064345/.

Cohn:1974:ABF

[Coh74e] Dr. Leslie Cohn. The asymptotic behaviour of the C-function. Lecture Notes in Mathematics, 429:107–122, 1974. CODEN LN-MAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064355/.

Cohn:1974:R

[Coh74f] Dr. Leslie Cohn. The C-ring. Lecture Notes in Mathematics, 429: 8–11, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064337/.

Cohn:1974:CPFa

[Coh74g] Dr. Leslie Cohn. Certain polynomial functions on N̄. Lecture Notes in Mathematics, 429:12–16, 1974. CODEN LN-MAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064338/.

Cohn:1974:CPFb

[Coh74h] Dr. Leslie Cohn. The critical points of the function ν (h( n̄ )). Lecture Notes in Mathematics, 429:102–106, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064354/.

Cohn:1974:DESc

[Coh74i] Dr. Leslie Cohn. The difference equations satisfied by the Cfunction. Lecture Notes in Mathematics, 429:97–101, 1974. CO-DEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064353/.

Cohn:1974:DESb

[Coh74j] Dr. Leslie Cohn. The differential equations on N̄ satisfied by the cusp form. Lecture Notes in Mathematics, 429:37–38, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064344.

Cohn:1974:DESa

[Coh74k] Dr. Leslie Cohn. The differential equations satisfied by e^{iv−ρ(H(x))}. Lecture Notes in Mathematics, 429:33–36, 1974. CO-DEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064343/.

Cohn:1974:F

[Coh74l] Dr. Leslie Cohn. The functions F_{J}(ν|n̄)(b). Lecture Notes in Mathematics, 429:42–50, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064346/.

Cohn:1974:FTC

[Coh74m] Dr. Leslie Cohn. The “fundamental theorem of calculus” on N̄. Lecture Notes in Mathematics, 429:21–22, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064340.

Cohn:1974:H

[Coh74n] Dr. Leslie Cohn. The homomorphism F_{J}. Lecture Notes in Mathematics, 429:51–54, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064347/.

Cohn:1974:IH

[Coh74o] Dr. Leslie Cohn. Injectivity of the homomorphism F_{J}. Lecture Notes in Mathematics, 429:64–73, 1974. CODEN LN-MAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064349/.

Cohn:1974:I

[Coh74p] Dr. Leslie Cohn. Introduction. Lecture Notes in Mathematics, 429:1–7, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064336/.

Cohn:1974:IPH

[Coh74q] Dr. Leslie Cohn. Invariance properties of the homomorphism F_{J}(ν). Lecture Notes in Mathematics, 429:74–90, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064350/.

Cohn:1974:NVC

[Coh74r] Dr. Leslie Cohn. Non-vanishing of certain polynomials. Lecture Notes in Mathematics, 429:91–93, 1974. CODEN LN-MAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064351/.

Cohn:1974:PCF

[Coh74s] Dr. Leslie Cohn. Preservation of certain filtrations. Lecture Notes in Mathematics, 429:55–63, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064348/.

Cohn:1974:RQ

[Coh74t] Dr. Leslie Cohn. The representation q. Lecture Notes in Mathematics, 429:31–32, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064342.

Cohn:1974:SE

[Coh74u] Dr. Leslie Cohn. Some estimates. Lecture Notes in Mathematics, 429:23–30, 1974. CODEN LNMAA2. ISBN 3-540-07017-6 (print), 3-540-37298-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064341/.

Colombe:1970:IES

[Col70] Yves Colombé. Intégration sur un ensemble semi-analytique (d’après M. Herrera.) (French) []. Lecture Notes in Mathematics, 116:39–55, 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059244/.

Collatz:1971:NIE

[Col71] L. Collatz. Nonlinear integral equations and field approximation problems. Lecture Notes in Mathematics, 228:39–55, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069446/.

Collatz:1972:AFF

[Col72] L. Collatz. Approximation by functions of fewer variables. Lecture Notes in Mathematics, 280:16–31, 1972. CODEN LN-MAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066916/.

Coleman:1973:FMF

[Col73a] Bernard D. Coleman. Fading memory and functional-differential equations. Lecture Notes in Mathematics, 322:78–86, 1973. CO-DEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060561/.

Colombeau:1973:AAA

[Col73b] J.-F. Colombeau. Sur les applications G-analytiques et analytiques en dimension infinie. (French) []. Lecture Notes in Mathematics, 332:48–58, 1973. CODEN LNMAA2. ISBN 3-540-06372-2 (print), 3-540-46980-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060906/.

Colton:1973:CPA

[Col73c] David Colton. Cauchy’s problem and the analytic continuation of solutions to elliptic equations. Lecture Notes in Mathematics, 316:55–66, 1973. CODEN LNMAA2. ISBN 3-540-06159-2 (print), 3-540-38370-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069623/.

Collatz:1974:DCA

[Col74a] L. Collatz. Discretization and chained approximation. Lecture Notes in Mathematics, 363:32–43, 1974. CODEN LN-MAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069123/.

Collatz:1974:FAF

[Col74b] L. Collatz. Field approximation and free approximation for differential equations. Lecture Notes in Mathematics, 399:59–74, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063566/.

Collatz:1974:MDP

[Col74c] L. Collatz. Monotonicity with discontinuities in partial differential equations. Lecture Notes in Mathematics, 415:85–102, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065514/.

Collingwood:1974:DTF

[Col74d] E. F. Collingwood. Dual theorems for functions defined in the unit disc. Lecture Notes in Mathematics, 419:76–79, 1974. CO-DEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064713/.

Colombeau:1974:SVN

[Col74e] J. F. Colombeau. On some various notions of infinite dimensional holomorphy. Lecture Notes in Mathematics, 364:145–149, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069013/.

Colton:1974:IOP

[Col74f] David Colton. Integral operators for parabolic equations and their application. Lecture Notes in Mathematics, 430:95–111, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066266/.

Comstock:1971:SPE

[Com71] Craig Comstock. Singular perturbation of elliptic equations II. Lecture Notes in Mathematics, 183:200–206, 1971. CODEN LN-MAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060420/.

Contensou:1970:COP

[Con70] Pierre Contensou. Conditions d’optimalité pour les domaines de manoeuvrabilité à frontière semi-affine. (French) []. Lecture Notes in Mathematics, 112:139–158, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060203/.

Conti:1972:ACP

[Con72] R. Conti. Affine control processes. Lecture Notes in Mathematics, 280:32–50, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066917/.

Conway:1973:EIB

[Con73] J. H. Conway. Effective implications between the ‘finite’ choice axioms. Lecture Notes in Mathematics, 337:439–458, 1973. CO-DEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066784/.

Conze:1974:TIO

[Con74] Jean-Pierre Conze. Le théorème d’isomorphisms d’Ornstein et la classification des systèmes dynamiques en théorie ergodique. (French) []. Lecture Notes in Mathematics, 383:50–68, 1974. CO-DEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057301/.

Cooper:1971:PSP

[Coo71a] D. C. Cooper. Program schemes, programs and logic. Lecture Notes in Mathematics, 188:62–70, 1971. CODEN LN-MAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059693/.

Cooper:1971:BEA

[Coo71b] G. J. Cooper. Bounds for the error in approximate solutions of ordinary differential equations. Lecture Notes in Mathematics, 228:270–276, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069461/.

Cooper:1974:WWG

[Coo74] C. D. H. Cooper. Words which give rise to another group operation for a given group. Lecture Notes in Mathematics, 372:221–225, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21571-5_19/;
http://link.springer.com/chapter/10.1007/BFb0065173/.

Coppel:1971:D

[Cop71a] W. A. Coppel. Disconjugacy, volume 220 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05584-3 (print), 3-540-36905-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 148 pp. LCCN QA3 .L28 no. 220. URL http:/

/link.springer.com/book/10.1007/BFb0058618;   http://
www.springerlink.com/content/978-3-540-36905-9.

Coppel:1971:EAO

[Cop71b] W. A. Coppel. Equations of arbitrary order. Lecture Notes in Mathematics, 220:81–135, 1971. CODEN LNMAA2. ISBN 3-540-05584-3 (print), 3-540-36905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058622/.

Coppel:1971:I

[Cop71c] W. A. Coppel. Introduction. Lecture Notes in Mathematics, 220: 1–3, 1971. CODEN LNMAA2. ISBN 3-540-05584-3 (print), 3-540-36905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058619/.

Coppel:1971:LHS

[Cop71d] W. A. Coppel. Linear Hamiltonian systems. Lecture Notes in Mathematics, 220:34–80, 1971. CODEN LNMAA2. ISBN 3-540-05584-3 (print), 3-540-36905-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058621/.

Coppel:1971:SOE

[Cop71e] W. A. Coppel. Second order equations. Lecture Notes in Mathematics, 220:4–33, 1971. CODEN LNMAA2. ISBN 3-540-05584-3 (print), 3-540-36905-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058620/.

Cordero:1974:NAP

[Cor74] Luis A. Cordero. On p-normal almost-product structures. Lecture Notes in Mathematics, 392:1–9, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064103/.

Cossey:1974:CFS

[Cos74] John Cossey. Classes of finite soluble groups. Lecture Notes in Mathematics, 372:226–237, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
662-21571-5_20/;   http://link.springer.com/chapter/
10.1007/BFb0065174/.

Coulomb:1973:IEQ

[Cou73] Geneviève Coulomb. Isolement des exposants et quasi-analyticité généralisée. (French) []. Lecture Notes in Mathematics, 336:104– 110, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065791/.

Coven:1971:NEA

[Cov71] Ethan M. Coven. A note on the existence of asymptotic pairs of points. Lecture Notes in Mathematics, 235:54–58, 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059194/.

Conner:1972:DA

[CR72a] P. E. Conner and Frank Raymond. Derived actions. Lecture Notes in Mathematics, 299:237–310, 1972. CODEN LNMAA2. ISBN 3-540-06078-2 (print), 3-540-38066-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066766/.

Conner:1972:HSF

[CR72b] P. E. Conner and Frank Raymond. Holomorphic Seifert fibering. Lecture Notes in Mathematics, 299:124–204, 1972. CODEN LNMAA2. ISBN 3-540-06078-2 (print), 3-540-38066-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066764/.

Conner:1972:IOT

[CR72c] P. E. Conner and Frank Raymond. Injective operations of the toral groups II. Lecture Notes in Mathematics, 299:109–123, 1972. CODEN LNMAA2. ISBN 3-540-06078-2 (print), 3-540-38066-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066763/.

Conner:1972:MFP

[CR72d] P. E. Conner and Frank Raymond. Manifolds with few periodic homeomorphisms. Lecture Notes in Mathematics, 299:1–75, 1972. CODEN LNMAA2. ISBN 3-540-06078-2 (print), 3-540-38066-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066760/.

Cartier:1973:BM

[CR73a] P. Cartier and Y. Roy. Back matter. Lecture Notes in Mathematics, 350:350, 1973. CODEN LNMAA2. ISBN 3-540-06483-4 (print), 3-540-37802-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-37802-0/1.pdf.

Cartier:1973:CCN

[CR73b] P. Cartier and Y. Roy. Certains calculs numériques relatifs à l’interpolation p-adique des séries de Dirichlet. (French) []. Lecture Notes in Mathematics, 350:269–349, 1973. CODEN LN-MAA2. ISBN 3-540-06483-4 (print), 3-540-37802-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-37802-0_
5/.

Conner:1972:MNP

[CRW72] P. E. Conner, Frank Raymond, and Peter J. Weinberger. Manifolds with no periodic maps. Lecture Notes in Mathematics, 299: 81–108, 1972. CODEN LNMAA2. ISBN 3-540-06078-2 (print), 3-540-38066-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066762/.

Chacon:1970:AI

[CS70] R. V. Chacon and T. Schwartzbauer. Approximation and invariance. Lecture Notes in Mathematics, 160:28–36, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060643/.

Chartrand:1971:IG

[CS71] Gary Chartrand and M. James Stewart. Isometric graphs. Lecture Notes in Mathematics, 186:63–67, 1971. CODEN LN-MAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059423/.

Chvatal:1974:RNF

[CS74] V. Chvátal and A. Schwenk. On the Ramsey number of the fivespoked wheel. Lecture Notes in Mathematics, 406:247–261, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066447/.

Csorgo:1973:GCT

[Csö73] Miklós Csörgő. Glivenko–Cantelli type theorems for distance functions based on the modified empirical distribution function of M. Kac and for the empirical process with random sample size in general. Lecture Notes in Mathematics, 296:149–164, 1973. CODEN LNMAA2. ISBN 3-540-06211-4 (print), 3-540-38485-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059823/.

Cochran:1974:RCR

[CT74] A. C. Cochran and R. B. Trail. Regularity and complete regularity for convergence spaces. Lecture Notes in Mathematics, 375: 64–70, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064012/.

Cunningham:1974:TBF

[Cun74] W. H. Cunningham. On theorems of Berge and Fournier. Lecture Notes in Mathematics, 411:67–74, 1974. CODEN LN-MAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066180/.

Curtis:1970:MRF

[Cur70] C. W. Curtis. Modular representations of finite groups with split (B, N)-pairs. Lecture Notes in Mathematics, 131:57–95, 1970. CODEN LNMAA2. ISBN 3-540-04920-7 (print), 3-540-36272-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081543/.

Curtis:1971:SI

[Cur71] Morton Curtis. H-spaces modp (II). Lecture Notes in Mathematics, 196:11–19, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067740/.

Curtain:1972:SSD

[Cur72] Ruth F. Curtain, editor. Stability of Stochastic Dynamical Systems: Proceedings of the International Symposium Organized by “The Control Theory Centre”, University of Warwick, July 10–14, 1972 Sponsored by the “International Union of Theoretical and Applied Mechanics”, volume 294 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 294.

URL http://link.springer.com/book/10.1007/BFb0064928;
 http://www.springerlink.com/content/978-3-540-38000-
9.

Curtis:1974:HSP

[Cur74] D. W. Curtis. The hyperspace of subcontinua of a Peano continuum. Lecture Notes in Mathematics, 378:108–118, 1974. CODEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068464/.

Cutland:1972:MC

[Cut72] Nigel Cutland. Π_{1}^{1} models and Π_{1}^{1}-categoricity. Lecture Notes in Mathematics, 255:42–62, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059537/.

Coifman:1971:AHN

[CW71a] Ronald R. Coifman and Guido Weiss. Analyse Hamonique Non-Commutative sur Certains Espaces Homogènes: Étude de Certaines Intégrales Singulières, volume 242 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CO-DEN LNMAA2. ISBN 3-540-05703-X (print), 3-540-37076-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 160 pp. LCCN QA3 .L28 no. 242. URL http:/

/link.springer.com/book/10.1007/BFb0058946;   http://
www.springerlink.com/content/978-3-540-37076-5.

Coifman:1971:AEO

[CW71b] Ronald R. Coifman and Guido Weiss. Autres exemples d’opérateurs intégraux singuliers et estimation dans L^{2}(G). (French) []. Lecture Notes in Mathematics, 242:145–154, 1971. CODEN LNMAA2. ISBN 3-540-05703-X (print), 3-540-37076-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058952/.

Coifman:1971:EDD

[CW71c] Ronald R. Coifman and Guido Weiss. Étude détaillée du groupe SU(2). (French) []. Lecture Notes in Mathematics, 242:101– 130, 1971. CODEN LNMAA2. ISBN 3-540-05703-X (print), 3-540-37076-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058950/.

Coifman:1971:ISM

[CW71d] Ronald R. Coifman and Guido Weiss. Intégrales singulières et multiplicateurs. (French) []. Lecture Notes in Mathematics, 242: 66–100, 1971. CODEN LNMAA2. ISBN 3-540-05703-X (print), 3-540-37076-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058949/.

Coifman:1971:MMF

[CW71e] Ronald R. Coifman and Guido Weiss. Multiplicateurs matriciels sur SU(2). (French) []. Lecture Notes in Mathematics, 242:131– 144, 1971. CODEN LNMAA2. ISBN 3-540-05703-X (print), 3-540-37076-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058951/.

Coifman:1971:RQG

[CW71f] Ronald R. Coifman and Guido Weiss. Représentations de quelques groupes compacts. (French) []. Lecture Notes in Mathematics, 242:24–65, 1971. CODEN LNMAA2. ISBN 3-540-05703-X (print), 3-540-37076-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058948/.

Coifman:1971:RGC

[CW71g] Ronald R. Coifman and Guido Weiss. Représentations des groupes compacts. Théorème de Peter–Weyl. (French) []. Lecture Notes in Mathematics, 242:1–23, 1971. CODEN LNMAA2. ISBN 3-540-05703-X (print), 3-540-37076-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058947/.

Curtis:1974:NMT

[CWW74] Morton Curtis, Alan Wiederhold, and Bruce Williams. Normalizers of maximal tori. Lecture Notes in Mathematics, 418:31– 47, 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070640/.

DAristotle:1974:SWS

[D’A74] Anthony J. D’Aristotle. On Stone–Weierstrass spaces. Lecture Notes in Mathematics, 378:119–131, 1974. CODEN LN-MAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068465/.

Danchick:1974:NEM

[Dan74] Roy Danchick. On the non-equivalence of maximum polynomial degree Nordsieck–Gear and classical methods. Lecture Notes in Mathematics, 362:92–106, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066587/.

Datko:1970:VGT

[Dat70] R. Datko. The variable gradiant as a tool in the study of stability. Lecture Notes in Mathematics, 144:34–43, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059918/.

Davies:1971:NLC

[Dav71] T. V. Davies. Non-linear cubic differential equations. Lecture Notes in Mathematics, 206:13–14, 1971. CODEN LN-MAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070124.

Davis:1973:ICP

[Dav73] Edward D. Davis. Integrally closed pairs. Lecture Notes in Mathematics, 311:103–106, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068923/.

Day:1970:CCF

[Day70] Brian Day. On closed categories of functors. Lecture Notes in Mathematics, 137:1–38, 1970. CODEN LNMAA2. ISBN 3-540-04926-6 (print), 3-540-36292-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060438/.

Day:1974:ETC

[Day74a] Brian Day. An embedding theorem for closed categories. Lecture Notes in Mathematics, 420:55–64, 1974. CODEN LN-MAA2. ISBN 3-540-06966-6 (print), 3-540-37270-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063099/.

Day:1974:LSC

[Day74b] Brian Day. Limit spaces and closed span categories. Lecture Notes in Mathematics, 420:65–74, 1974. CODEN LNMAA2. ISBN 3-540-06966-6 (print), 3-540-37270-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063100/.

Day:1974:AFF

[Day74c] Brian Day. On adjoint-functor factorisation. Lecture Notes in Mathematics, 420:1–19, 1974. CODEN LNMAA2. ISBN 3-540-06966-6 (print), 3-540-37270-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063097/.

Day:1974:CCF

[Day74d] Brian Day. On closed categories of functors II. Lecture Notes in Mathematics, 420:20–54, 1974. CODEN LNMAA2. ISBN 3-540-06966-6 (print), 3-540-37270-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063098/.

deBruijn:1970:MLA

[dB70] N. G. de Bruijn. The mathematical language AUTOMATH, its usage, and some of its extensions. Lecture Notes in Mathematics, 125:29–61, 1970. CODEN LNMAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060623/.

deBakker:1971:ASS

[dB71a] J. W. de Bakker. Axiom systems for simple assignment statements. Lecture Notes in Mathematics, 188:1–22, 1971. CODEN LNMAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059690/.

deBakker:1971:PLC

[dB71b] J. W. de Bakker. A property of linear conditionals. Lecture Notes in Mathematics, 188:23–27, 1971. CODEN LNMAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059691/.

deBoor:1974:GAS

[dB74] Carl de Boor. Good approximation by splines with variable knots. II. Lecture Notes in Mathematics, 363:12–20, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069121/.

Dacunha-Castelle:1970:PDP

[DC70] D. Dacunha-Castelle. Principe de dualité pour des espaces de suites associés à une suite de variables aléatoires. (French) []. Lecture Notes in Mathematics, 124:47–59, 1970. CODEN LN-MAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059332/.

Dacunha-Castelle:1974:CEP

[DC74] Didier Dacunha-Castelle. Contre-exemple à la propriété d’approximation uniforme dans les espaces de Banach. (French) []. Lecture Notes in Mathematics, 383:286–293, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057314/.

Doleans-Dade:1971:ISP

[DD71a] C. Doléans-Dade. Intégrales stochastiques par rapport à une famille de probabilités. (French) []. Lecture Notes in Mathematics, 191:141–146, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058855/.

Doleans-Dade:1971:MUI

[DD71b] C. Doléans-Dade. Une martingale uniformément intégrable mais non localement de carré intégrable. (French) []. Lecture Notes in Mathematics, 191:138–140, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058854/.

Douglas:1974:CMP

[DD74a] Jim Douglas, Jr. and Todd Dupont. Collocation Methods for Parabolic Equations in a Single Space Variable: Based on C^{1}-Piecewise-Polynomial Spaces, volume 385 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CO-DEN LNMAA2. ISBN 3-540-06747-7 (print), 3-540-38337-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 147 pp. LCCN QA3 .L28 no. 385. URL http:/

/link.springer.com/book/10.1007/BFb0057337;   http://
www.springerlink.com/content/978-3-540-38337-6.

Douglas:1974:GEE

[DD74b] Jim Douglas, Jr. and Todd Dupont. Global error estimates. Lecture Notes in Mathematics, 385:1–62, 1974. CODEN LN-MAA2. ISBN 3-540-06747-7 (print), 3-540-38337-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057338/.

Douglas:1974:LSL

[DD74c] Jim Douglas, Jr. and Todd Dupont. Local superconvergence by local refinement. Lecture Notes in Mathematics, 385:99– 108, 1974. CODEN LNMAA2. ISBN 3-540-06747-7 (print), 3-540-38337-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057340/.

Douglas:1974:SCM

[DD74d] Jim Douglas, Jr. and Todd Dupont. A smoothed collocation method and applications to eigensystem approximation. Lecture Notes in Mathematics, 385:109–143, 1974. CODEN LN-MAA2. ISBN 3-540-06747-7 (print), 3-540-38337-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057341/.

Douglas:1974:SEK

[DD74e] Jim Douglas, Jr. and Todd Dupont. Superconvergence estimates at the knots. Lecture Notes in Mathematics, 385:63–98, 1974. CODEN LNMAA2. ISBN 3-540-06747-7 (print), 3-540-38337-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057339/.

Dellacherie:1971:CEAb

[DDD71] C. Dellacherie and C. Doléans-Dade. Un contre-exemple au problème des laplaciens approchés. Lecture Notes in Mathematics, 191:127–137, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058853/.

Doleans-Dade:1970:DCC

[DDDM70] Catherine Doléans-Dade, Claude Dellacherie, and Paul-André Meyer. Diffusions à coefficients continus d’après D. W. Stroock et S. R. S. Varadhan. (French) []. Lecture Notes in Mathematics, 124:240, 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059346.

Doleans-Dade:1970:ISP

[DDM70] C. Doléans-Dade and P.-A. Meyer. Intégrales stochastiques par rapport aux martingales locales. (French) []. Lecture Notes in Mathematics, 124:77–107, 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059337/.

DeWilde:1973:VTB

[De 73] M. De Wilde. Various types of barrelledness and increasing sequences of balanced and convex sets in locally convex spaces. Lecture Notes in Mathematics, 331:211–217, 1973. CODEN LN-MAA2. ISBN 3-540-06367-6 (print), 3-540-46977-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068232/.

Dold:1971:SPV

[DE71] A. Dold and B. Eckmann. Séminaire de Probabilités V Université de Strasbourg, volume 191 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 372 pp. LCCN QA3 .L28 no. 191.

URL http://link.springer.com/book/10.1007/BFb0058841;
 http://www.springerlink.com/content/978-3-540-36517-
4.

Dold:1973:SBV

[DE73] A. Dold and B. Eckmann, editors. Séminaire Bourbaki vol. 1971/ 72 Exposés 400–417, volume 317 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 317. URL http:/

/link.springer.com/book/10.1007/BFb0069272;   http://
www.springerlink.com/content/978-3-540-38403-8.

Dellacherie:1969:EAI

[Del69] C. Dellacherie. Ensembles aléatoires II. (French) [Random sets II]. Lecture Notes in Mathematics, 88:115–136, 1969. CODEN LNMAA2. ISBN 3-540-04607-0 (print), 3-540-36097-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081629/. See cor-

rection [Del71c].

Deligne:1970:AEF

[Del70a] Pierre Deligne. Applications et exemples. (French) []. Lecture Notes in Mathematics, 163:122–127, 1970. CODEN LN-MAA2. ISBN 3-540-05190-2 (print), 3-540-36373-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061198/.

Deligne:1970:CRF

[Del70b] Pierre Deligne. Connexions régulières. (French) []. Lecture Notes in Mathematics, 163:41–121, 1970. CODEN LNMAA2. ISBN 3-540-05190-2 (print), 3-540-36373-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061197/.

Deligne:1970:DFD

[Del70c] Pierre Deligne. Dictionnaire. (French) [Dictionary]. Lecture Notes in Mathematics, 163:3–40, 1970. CODEN LNMAA2. ISBN 3-540-05190-2 (print), 3-540-36373-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061196/.

Deligne:1970:EDP

[Del70d] Pierre Deligne. Équations Différentielles à Points Singuliers Réguliers. (French) [Differential equations with singular points regular], volume 163 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05190-2 (print), 3-540-36373-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 133 pp. LCCN QA3 .L28 no. 163. URL http:/

/link.springer.com/book/10.1007/BFb0061194;   http://
www.springerlink.com/content/978-3-540-36373-6.

Deligne:1970:IFI

[Del70e] Pierre Deligne. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 163:1–2, 1970. CODEN LNMAA2. ISBN 3-540-05190-2 (print), 3-540-36373-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061195.

Dellacherie:1970:ASS

[Del70f] C. Dellacherie. Au sujet des sauts d’un processus de Hunt. (French) []. Lecture Notes in Mathematics, 124:71–72, 1970. CO-DEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059334.

Dellacherie:1970:DCC

[Del70g] C. Dellacherie. Diffusions à coefficients continus: résultats d’existence d’après un exposé de C. Dellacherie. (French) []. Lecture Notes in Mathematics, 124:252–261, 1970. CODEN LN-MAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059348/.

Dellacherie:1970:PGF

[Del70h] C. Dellacherie. Potentiels de Green et fonctionnelles additives. (French) []. Lecture Notes in Mathematics, 124:73–75, 1970. CO-DEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059335/.

Dellacherie:1970:ETG

[Del70i] C. Dellacherie. Un exemple de la théorie générale des processus. (French) []. Lecture Notes in Mathematics, 124:60–70, 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059333/.

Dellacherie:1970:LTM

[Del70j] C. Dellacherie. Un lemme de la théorie de la mesure. (French) []. Lecture Notes in Mathematics, 124:76, 1970. CODEN LN-MAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059336.

Deligne:1971:FMR

[Del71a] Pierre Deligne. Formes modulaires et représentations e-adiques. (French) []. Lecture Notes in Mathematics, 179:139–172, 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058810/.

Deligne:1971:TGF

[Del71b] Pierre Deligne. Travaux de Griffiths. (French) []. Lecture Notes in Mathematics, 180:213–237, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058833/.

Dellacherie:1971:CEAa

[Del71c] C. Dellacherie. Correction à ‘Ensembles aléatoires II’. (French) []. Lecture Notes in Mathematics, 191:86, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0058850. See [Del69].

Dellacherie:1971:EPR

[Del71d] C. Dellacherie. Espaces pavés et rabotages. (French) [Paved spaces and planing]. Lecture Notes in Mathematics, 191:103– 126, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058852/.

Dellacherie:1971:TMS

[Del71e] C. Dellacherie. Les théorèmes de Mazurkiewicz–Sierpiński et de Lusin. (French) []. Lecture Notes in Mathematics, 191:87– 102, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058851/.

Dellacherie:1971:QCP

[Del71f] C. Dellacherie. Quelques commentaires sur les prolongements de capacités. (French) []. Lecture Notes in Mathematics, 191: 77–81, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058848/.

Dellacherie:1971:DDT

[Del71g] C. Dellacherie. Une démonstration du théorème de séparation des ensembles analytiques. (French) []. Lecture Notes in Mathematics, 191:82–85, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058849/.

Deligne:1972:RPE

[Del72a] P. Deligne. Résumé des premiers exposés de A. Grothendieck. (French) []. Lecture Notes in Mathematics, 288:1–24, 1972. CO-DEN LNMAA2. ISBN 3-540-05987-3 (print), 3-540-37984-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068689/.

Deligne:1972:TSF

[Del72b] Pierre Deligne. Travaux de Shimura. (French) []. Lecture Notes in Mathematics, 244:123–165, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058700/.

Dellacherie:1972:CFb

[Del72c] Claude Dellacherie. Calibres. (French) []. Lecture Notes in Mathematics, 295:41–50, 1972. CODEN LNMAA2. ISBN 3-540-06076-6 (print), 3-540-38059-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060709/.

Dellacherie:1972:CFa

[Del72d] Claude Dellacherie. Capacités. (French) []. Lecture Notes in Mathematics, 295:19–40, 1972. CODEN LNMAA2. ISBN 3-540-06076-6 (print), 3-540-38059-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060708/.

Dellacherie:1972:EAC

[Del72e] Claude Dellacherie. Ensembles Analytiques Capacités Mesures de Hausdorff, volume 295 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06076-6 (print), 3-540-38059-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 123 pp. LCCN QA3 .L28 no. 295.

URL http://link.springer.com/book/10.1007/BFb0060706;
 http://www.springerlink.com/content/978-3-540-38059-
7.

Dellacherie:1972:EFA

[Del72f] Claude Dellacherie. Ensembles et fonctions analytiques. (French) []. Lecture Notes in Mathematics, 295:1–18, 1972. CODEN LNMAA2. ISBN 3-540-06076-6 (print), 3-540-38059-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060707/.

Dellacherie:1972:EF

[Del72g] Claude Dellacherie. Épaisseurs. (French) []. Lecture Notes in Mathematics, 295:65–81, 1972. CODEN LNMAA2. ISBN 3-540-06076-6 (print), 3-540-38059-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060711/.

Dellacherie:1972:MHF

[Del72h] Claude Dellacherie. Mesures de Hausdorff. (French) []. Lecture Notes in Mathematics, 295:82–101, 1972. CODEN LN-MAA2. ISBN 3-540-06076-6 (print), 3-540-38059-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060712/.

Dellacherie:1972:NCF

[Del72i] Claude Dellacherie. Noyaux capacitaires. (French) [Capability kernels]. Lecture Notes in Mathematics, 295:51–64, 1972. CO-DEN LNMAA2. ISBN 3-540-06076-6 (print), 3-540-38059-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060710/.

Deligne:1973:BM

[Del73a] P. Deligne. Back matter. Lecture Notes in Mathematics, 349: 598, 1973. CODEN LNMAA2. ISBN 3-540-06558-X (print), 3-540-37855-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-37855-6/1.pdf.

Deligne:1973:CSP

[Del73b] P. Deligne. Cohomologie a supports propres. Lecture Notes in Mathematics, 305:250–480, 1973. CODEN LNMAA2. ISBN 3-540-06118-5 (print), 3-540-38124-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070723/.

Deligne:1973:CIC

[Del73c] P. Deligne. Cohomologie des intersections complètes. (French) []. Lecture Notes in Mathematics, 340:39–61, 1973. CODEN LNMAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060507/.

Deligne:1973:CAT

[Del73d] P. Deligne. Comparaison avec la théorie transcendante. (French) []. Lecture Notes in Mathematics, 340:116–164, 1973. CODEN LNMAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060510/.

Deligne:1973:FMR

[Del73e] P. Deligne. Formes modulaires et représentations de GL(2). (French) []. Lecture Notes in Mathematics, 349:55–105, 1973. CODEN LNMAA2. ISBN 3-540-06558-X (print), 3-540-37855-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

URL http://link.springer.com/chapter/10.1007/978-3-
540-37855-6_2/.

Deligne:1973:ISR

[Del73f] P. Deligne. Intersections sur les surfaces regulières. (French) []. Lecture Notes in Mathematics, 340:1–38, 1973. CODEN LN-MAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060506/.

Deligne:1973:FDG

[Del73g] P. Deligne. La formule de dualité globale. (French) []. Lecture Notes in Mathematics, 305:481–587, 1973. CODEN LN-MAA2. ISBN 3-540-06118-5 (print), 3-540-38124-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070724/.

Deligne:1973:FMF

[Del73h] P. Deligne. La formule de Milnor. (French) []. Lecture Notes in Mathematics, 340:197–211, 1973. CODEN LNMAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060512/.

Deligne:1973:FPL

[Del73i] P. Deligne. La formule de Picard–Lefschetz. (French) []. Lecture Notes in Mathematics, 340:165–196, 1973. CODEN LN-MAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060511/.

Deligne:1973:FCE

[Del73j] P. Deligne. Le formalisme des cycles évanescents. (French) []. Lecture Notes in Mathematics, 340:82–115, 1973. CODEN LN-MAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060509/.

Deligne:1973:TNF

[Del73k] P. Deligne. Le théorème de Noether. (French) []. Lecture Notes in Mathematics, 340:328–340, 1973. CODEN LNMAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060515/.

Deligne:1973:CEF

[Del73l] P. Deligne. Les constantes des Équations fonctionnelles des fonctions L. (French) []. Lecture Notes in Mathematics, 349:501– 597, 1973. CODEN LNMAA2. ISBN 3-540-06558-X (print), 3-540-37855-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-37855-6_7/.

Deligne:1973:QFQ

[Del73m] P. Deligne. Quadriques. (French) [Quadrics]. Lecture Notes in Mathematics, 340:62–81, 1973. CODEN LNMAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060508/.

Deligne:1973:VUN

[Del73n] Pierre Deligne. Variétés unirationnelles non rationnelles [d’après M. Artin et D. Mumford]. (French) []. Lecture Notes in Mathematics, 317:45–57, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069275/.

Dellacherie:1973:TFT

[Del73o] C. Dellacherie. Sur les théorèmes fondamentaux de la théorie générale des processus. (French) []. Lecture Notes in Mathematics, 321:38–47, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071393/.

Dellacherie:1973:TAT

[Del73p] C. Dellacherie. Temps d’arrêt totalement inaccessibles. (French) []. Lecture Notes in Mathematics, 321:36–37, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071392.

Dellacherie:1973:CGF

[Del73q] C. Dellacherie. Un crible généralisé. (French) []. Lecture Notes in Mathematics, 321:33–35, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071391/.

Dellacherie:1973:CES

[Del73r] C. Dellacherie. Une conjecture sur les ensembles semi-polaires. (French) []. Lecture Notes in Mathematics, 321:51–57, 1973. CO-DEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071395/.

Dellacherie:1973:DDT

[Del73s] C. Dellacherie. Une démonstration du théorème de Souslin– Lusin. (French) []. Lecture Notes in Mathematics, 321:48–50, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071394/.

Dellacherie:1974:ISP

[Del74a] C. Dellacherie. Intégrales stochastiques par rapport aux processus de Wiener et de Poisson. (French) []. Lecture Notes in Mathematics, 381:25–26, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057254.

Dellacherie:1974:EPM

[Del74b] C. Dellacherie. Un ensemble progressivement mesurable …. Lecture Notes in Mathematics, 381:22–24, 1974. CODEN LN-MAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057253/.

Demazure:1970:AGR

[Dem70a] M. Demazure. Automorphismes des groupes reductifs. (French) []. Lecture Notes in Mathematics, 153:323–409, 1970. CODEN LNMAA2. ISBN 3-540-05181-3 (print), 3-540-36339-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059033/.

Demazure:1970:DRF

[Dem70b] M. Demazure. Données radicielles. (French) []. Lecture Notes in Mathematics, 153:85–155, 1970. CODEN LNMAA2. ISBN 3-540-05181-3 (print), 3-540-36339-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059030/.

Demazure:1970:EIF

[Dem70c] M. Demazure. Extensions infinitesimales. (French) []. Lecture Notes in Mathematics, 151:83–159, 1970. CODEN LN-MAA2. ISBN 3-540-05179-1 (print), 3-540-36336-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058996/.

Demazure:1970:FTA

[Dem70d] M. Demazure. Fibrés tangents — algèbres de Lie. (French) []. Lecture Notes in Mathematics, 151:43–82, 1970. CODEN LN-MAA2. ISBN 3-540-05179-1 (print), 3-540-36336-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058995/.

Demazure:1970:GRG

[Dem70e] M. Demazure. Groupes reductifs — généralités. (French) []. Lecture Notes in Mathematics, 153:1–34, 1970. CODEN LN-MAA2. ISBN 3-540-05181-3 (print), 3-540-36339-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059028/.

Demazure:1970:GRR

[Dem70f] M. Demazure. Groupes réductifs de rang semi-simple. (French) []. Lecture Notes in Mathematics, 153:35–84, 1970. CODEN LNMAA2. ISBN 3-540-05181-3 (print), 3-540-36339-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059029/.

Demazure:1970:GRD

[Dem70g] M. Demazure. Groupes réductifs: Déploiements, sous-groupes, groupes-quotients. (French) []. Lecture Notes in Mathematics, 153:156–262, 1970. CODEN LNMAA2. ISBN 3-540-05181-3 (print), 3-540-36339-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059031/.

Demazure:1970:GRU

[Dem70h] M. Demazure. Groupes reductifs; Unicité des groupes Epinglés. (French) []. Lecture Notes in Mathematics, 153:263–322, 1970. CODEN LNMAA2. ISBN 3-540-05181-3 (print), 3-540-36339-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059032/.

Demazure:1970:TEF

[Dem70i] M. Demazure. Le théorème d’existence. (French) []. Lecture Notes in Mathematics, 153:410–425, 1970. CODEN LNMAA2. ISBN 3-540-05181-3 (print), 3-540-36339-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059034/.

Demazure:1970:SGP

[Dem70j] M. Demazure. Sous-groupes paraboliques des groupes reductifs. (French) []. Lecture Notes in Mathematics, 153:426–517, 1970. CODEN LNMAA2. ISBN 3-540-05181-3 (print), 3-540-36339-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059035/.

Demazure:1970:SAC

[Dem70k] M. Demazure. Structures algébriques — cohomologie des groupes. (French) []. Lecture Notes in Mathematics, 151:1–42, 1970. CODEN LNMAA2. ISBN 3-540-05179-1 (print), 3-540-36336-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058994/.

Demazure:1970:TFF

[Dem70l] M. Demazure. Topologies et faisceaux. (French) []. Lecture Notes in Mathematics, 151:160–250, 1970. CODEN LNMAA2. ISBN 3-540-05179-1 (print), 3-540-36336-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058997/.

Demazure:1971:MVA

[Dem71] Michel Demazure. Motifs des variétés algébriques. (French) []. Lecture Notes in Mathematics, 180:19–38, 1971. CODEN LN-MAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058822/.

Demazure:1972:CDG

[Dem72a] Michel Demazure. Classification of p-divisible groups. Lecture Notes in Mathematics, 302:73–93, 1972. CODEN LN-MAA2. ISBN 3-540-06092-8 (print), 3-540-38079-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060745/.

Demazure:1972:GSF

[Dem72b] Michel Demazure. Group-schemes and formal group-schemes. Lecture Notes in Mathematics, 302:21–49, 1972. CODEN LN-MAA2. ISBN 3-540-06092-8 (print), 3-540-38079-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060743/.

Demazure:1972:LDG

[Dem72c] Michel Demazure. Lectures on p-Divisible Groups, volume 302 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06092-8 (print), 3-540-38079-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 98 pp. LCCN QA3 .L28 no. 302. URL http:/

/link.springer.com/book/10.1007/BFb0060741;   http://
www.springerlink.com/content/978-3-540-38079-5.

Demazure:1972:ACA

[Dem72d] Michel Demazure. p-adic cohomology of abelian varieties. Lecture Notes in Mathematics, 302:94–98, 1972. CODEN LN-MAA2. ISBN 3-540-06092-8 (print), 3-540-38079-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060746/.

Demazure:1972:SFS

[Dem72e] Michel Demazure. Schemes and formal schemes. Lecture Notes in Mathematics, 302:1–20, 1972. CODEN LNMAA2. ISBN 3-540-06092-8 (print), 3-540-38079-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060742/.

Demazure:1972:WGD

[Dem72f] Michel Demazure. Witt groups and Dieudonné modules. Lecture Notes in Mathematics, 302:50–72, 1972. CODEN LN-MAA2. ISBN 3-540-06092-8 (print), 3-540-38079-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060744/.

Dennis:1973:S

[Den73a] R. Keith Dennis. Stability for K_{2}. Lecture Notes in Mathematics, 353:85–94, 1973. CODEN LNMAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059262/.

Deny:1973:DRT

[Den73b] Jacques Deny. Développements récents de la théorie du potentiel (travaux de Jacques Faraut et de Francis Hirsch). (French) []. Lecture Notes in Mathematics, 317:59–72, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069276/.

Deuring:1973:CFE

[Deu73a] Max Deuring. Constant field extensions. Lecture Notes in Mathematics, 314:92–149, 1973. CODEN LNMAA2. ISBN 3-540-06152-5 (print), 3-540-38368-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060948/.

Deuring:1973:FFV

[Deu73b] Max Deuring. Function fields and valuations. Lecture Notes in Mathematics, 314:1–19, 1973. CODEN LNMAA2. ISBN 3-540-06152-5 (print), 3-540-38368-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060945/.

Deuring:1973:LTA

[Deu73c] Max Deuring. Lectures on the Theory of Algebraic Functions of One Variable, volume 314 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06152-5 (print), 3-540-38368-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 151 pp. LCCN QA3 .L28 no. 314.

URL http://link.springer.com/book/10.1007/BFb0060944;
 http://www.springerlink.com/content/978-3-540-38368-
0.

Deuring:1973:RRT

[Deu73d] Max Deuring. The Riemann–Roch theorem. Lecture Notes in Mathematics, 314:20–54, 1973. CODEN LNMAA2. ISBN 3-540-06152-5 (print), 3-540-38368-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060946/.

Deuring:1973:ZFF

[Deu73e] Max Deuring. Zeta function and L-functions. Lecture Notes in Mathematics, 314:54–92, 1973. CODEN LNMAA2. ISBN 3-540-06152-5 (print), 3-540-38368-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060947/.

DeVore:1972:APO

[DeV72a] Ronald A. DeVore. Algebraic polynomial operators. Lecture Notes in Mathematics, 293:170–204, 1972. CODEN LN-MAA2. ISBN 3-540-06038-3 (print), 3-540-37995-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059499/.

DeVore:1972:ACFa

[DeV72b] Ronald A. DeVore. Approximation of classes of functions. Lecture Notes in Mathematics, 293:205–237, 1972. CODEN LN-MAA2. ISBN 3-540-06038-3 (print), 3-540-37995-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059500/.

DeVore:1972:ACFb

[DeV72c] Ronald A. DeVore. The Approximation of Continuous Functions by Positive Linear Operators, volume 293 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CO-DEN LNMAA2. ISBN 3-540-06038-3 (print), 3-540-37995-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 289 pp. LCCN QA3 .L28 no. 293. URL http:/

/link.springer.com/book/10.1007/BFb0059493;   http://
www.springerlink.com/content/978-3-540-37995-9.

DeVore:1972:FAP

[DeV72d] Ronald A. DeVore. Fundamental approximation properties of positive linear operators. Lecture Notes in Mathematics, 293: 22–47, 1972. CODEN LNMAA2. ISBN 3-540-06038-3 (print), 3-540-37995-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059495/.

DeVore:1972:IT

[DeV72e] Ronald A. DeVore. Inverse theorems. Lecture Notes in Mathematics, 293:238–277, 1972. CODEN LNMAA2. ISBN 3-540-06038-3 (print), 3-540-37995-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059501/.

DeVore:1972:P

[DeV72f] Ronald A. DeVore. Preliminaries. Lecture Notes in Mathematics, 293:1–21, 1972. CODEN LNMAA2. ISBN 3-540-06038-3 (print), 3-540-37995-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059494/.

DeVore:1972:SCB

[DeV72g] Ronald A. DeVore. Saturation in c[a,b]. Lecture Notes in Mathematics, 293:121–169, 1972. CODEN LNMAA2. ISBN 3-540-06038-3 (print), 3-540-37995-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059498/.

DeVore:1972:SCO

[DeV72h] Ronald A. DeVore. Saturation of convolution operators. Lecture Notes in Mathematics, 293:48–102, 1972. CODEN LN-MAA2. ISBN 3-540-06038-3 (print), 3-540-37995-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059496/.

DeVore:1972:TPO

[DeV72i] Ronald A. DeVore. Trigonometric polynomial operators. Lecture Notes in Mathematics, 293:103–120, 1972. CODEN LN-MAA2. ISBN 3-540-06038-3 (print), 3-540-37995-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059497/.

Devlin:1973:AC

[Dev73a] Keith J. Devlin. Aspects of Constructibility, volume 354 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 240 pp. LCCN QA3 .L28 no. 354. URL http:/

/link.springer.com/book/10.1007/BFb0059290;   http://
www.springerlink.com/content/978-3-540-37821-1.

Devlin:1973:ACC

[Dev73b] Keith J. Devlin. The axiom of constructibility, the condensation lemma, and the consistency of the generalised continuum hypothesis. Lecture Notes in Mathematics, 354:29–42, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059293/.

Devlin:1973:CPI

[Dev73c] Keith J. Devlin. The combinatorial principles ie104-1. Lecture Notes in Mathematics, 354:104–114, 1973. CODEN LN-MAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059299/.

Devlin:1973:CC

[Dev73d] Keith J. Devlin. Constructibility and the continuum. Lecture Notes in Mathematics, 354:59–72, 1973. CODEN LN-MAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059296/.

Devlin:1973:CU

[Dev73e] Keith J. Devlin. The constructible universe. Lecture Notes in Mathematics, 354:24–28, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059292/.

Devlin:1973:FSJ

[Dev73f] Keith J. Devlin. The fine structure of the Jensen hierarchy. Lecture Notes in Mathematics, 354:86–103, 1973. CODEN LN-MAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059298/.

Devlin:1973:GM

[Dev73g] Keith J. Devlin. Gap-1 morasses. Lecture Notes in Mathematics, 354:145–170, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059304/.

Devlin:1973:GTCa

[Dev73h] Keith J. Devlin. The gap-1 two cardinals theorem. Lecture Notes in Mathematics, 354:138–144, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059303/.

Devlin:1973:GTCb

[Dev73i] Keith J. Devlin. The gap-2 two cardinals theorem (Jensen). Lecture Notes in Mathematics, 354:171–182, 1973. CODEN LN-MAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059305/.

Devlin:1973:GSH

[Dev73j] Keith J. Devlin. The generalised Souslin hypothesis for successor cardinals. Lecture Notes in Mathematics, 354:115–118, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059300/.

Devlin:1973:ICG

[Dev73k] Keith J. Devlin. Ineffable cardinals and the generalised Kurepa hypothesis. Lecture Notes in Mathematics, 354:119–129, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059301/.

Devlin:1973:IM

[Dev73l] Keith J. Devlin. The inner model l[μ]. Lecture Notes in Mathematics, 354:214–228, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059310/.

Devlin:1973:JHC

[Dev73m] Keith J. Devlin. The Jensen hierarchy of constructible sets. Lecture Notes in Mathematics, 354:73–85, 1973. CODEN LN-MAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059297/.

Devlin:1973:KH

[Dev73n] Keith J. Devlin. The Kurepa hypothesis. Lecture Notes in Mathematics, 354:53–58, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059295/.

Devlin:1973:LLC

[Dev73o] Keith J. Devlin. Large large cardinals. Lecture Notes in Mathematics, 354:188–205, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059308/.

Devlin:1973:MLC

[Dev73p] Keith J. Devlin. Medium large cardinals. Lecture Notes in Mathematics, 354:186–187, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059307.

Devlin:1973:RC

[Dev73q] Keith J. Devlin. Relative constructibility. Lecture Notes in Mathematics, 354:206–213, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059309/.

Devlin:1973:SLC

[Dev73r] Keith J. Devlin. Small large cardinals. Lecture Notes in Mathematics, 354:183–185, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059306/.

Devlin:1973:SH

[Dev73s] Keith J. Devlin. The Souslin hypothesis. Lecture Notes in Mathematics, 354:43–52, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059294/.

Devlin:1973:WCC

[Dev73t] Keith J. Devlin. Weakly compact cardinals and the generalised Souslin hypothesis. Lecture Notes in Mathematics, 354:130– 137, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059302/.

Devlin:1973:ZFS

[Dev73u] Keith J. Devlin. Zermelo–Fraenkel set theory. Lecture Notes in Mathematics, 354:1–23, 1973. CODEN LNMAA2. ISBN 3-540-06522-9 (print), 3-540-37821-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059291/.

Devinatz:1974:DIT

[Dev74] A. Devinatz. On a deficiency index theorem of W. N. Everitt. Lecture Notes in Mathematics, 415:103–108, 1974. CODEN LN-MAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065515/.

Deza:1974:BCR

[Dez74] Mikhaïl Deza. Bref compte rendu des résultats lies au problème de la comparaison des bruits additifs avec un nombre donné de combinaisons de bruits. (French) []. Lecture Notes in Mathematics, 398:118–125, 1974. CODEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065773/.

deFeriet:1974:TGI

[dF74a] Joseph Kampé de Fériet. La théorie généralisée de l’information et la mesure subjective de l’information. (French) []. Lecture Notes in Mathematics, 398:1–35, 1974. CODEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065765/.

Dickman:1974:TCV

[DF74b] Raymond F. Dickman, Jr. and Peter Fletcher, editors. Topology Conference: Virginia Polytechnic Institute and State University, March 22–24, 1973, volume 375 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 375. URL http:/

/link.springer.com/book/10.1007/BFb0064004;   http://
www.springerlink.com/content/978-3-540-37948-5.

deFeriet:1974:TIA

[dFP74] J. Kampé de Fériet and C. F. Picard, editors. Théories de l’Information: Actes des Rencontres de Marseille–Luminy 5 au 7 Juin 1973, volume 398 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 398. URL http:/

/link.springer.com/book/10.1007/BFb0065764;   http://
www.springerlink.com/content/978-3-540-37798-6.

Demazure:1970:GTM

[DG70a] Michel Demazure and Alexander Grothendieck. Groupes de Type Multiplicatif, et Structure des Schémas en Groupes Généraux, volume 152 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05180-5 (print), 3-540-36337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 654 pp. LCCN QA3 .L28 no. 152. URL http:

//link.springer.com/book/10.1007/BFb0059005;   http:/
/www.springerlink.com/content/978-3-540-36337-8.

Demazure:1970:PGS

[DG70b] Michel Demazure and Alexander Grothendieck. Propriétés Générales des Schémas en Groupes. (French) [General Properties of Schemes in Groups], volume 151 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05179-1 (print), 3-540-36336-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xv + 564 pp. LCCN QA3 .L28 no. 151. URL http:/

/link.springer.com/book/10.1007/BFb0058993;   http://
www.springerlink.com/content/978-3-540-36336-1.

Demazure:1970:SSG

[DG70c] Michel Demazure and Alexander Grothendieck. Structure des Schémas en Groupes Réductifs, volume 153 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LN-MAA2. ISBN 3-540-05181-3 (print), 3-540-36337-8 (e-book), 3-540-36339-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 529 pp. LCCN QA3 .L28 no. 153. URL http:/

/link.springer.com/book/10.1007/BFb0059027;   http://
www.springerlink.com/content/978-3-540-36339-2.

Davis:1973:MIP

[DG73] Edward D. Davis and Anthony V. Geramita. Maximal ideals in polynomial rings. Lecture Notes in Mathematics, 311:57– 60, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068919/.

Dhombres:1974:LIL

[Dho74] J. G. Dhombres. Linear interpolation and linear extension of functions. Lecture Notes in Mathematics, 399:75–83, 1974. CO-DEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063567/.

DiPaola:1971:RGS

[Di 71] Jane W. Di Paola. On a result of Goethals and Seidel. Lecture Notes in Mathematics, 186:75–84, 1971. CODEN LN-MAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059426/.

DeMeyer:1971:AR

[DI71a] Frank DeMeyer and Edward Ingraham. Applications and remarks. Lecture Notes in Mathematics, 181:131–149, 1971. CO-DEN LNMAA2. ISBN 3-540-05371-9 (print), 3-540-36484-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061231/.

DeMeyer:1971:CSA

[DI71b] Frank DeMeyer and Edward Ingraham. Central separable algebras and the Brauer group. Lecture Notes in Mathematics, 181: 39–78, 1971. CODEN LNMAA2. ISBN 3-540-05371-9 (print), 3-540-36484-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061228/.

DeMeyer:1971:GT

[DI71c] Frank DeMeyer and Edward Ingraham. Galois theory. Lecture Notes in Mathematics, 181:79–114, 1971. CODEN LN-MAA2. ISBN 3-540-05371-9 (print), 3-540-36484-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061229/.

DeMeyer:1971:P

[DI71d] Frank DeMeyer and Edward Ingraham. Preliminaries. Lecture Notes in Mathematics, 181:1–38, 1971. CODEN LNMAA2. ISBN 3-540-05371-9 (print), 3-540-36484-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061227/.

DeMeyer:1971:SAC

[DI71e] Frank DeMeyer and Edward Ingraham. Separable Algebras Over Commutative Rings, volume 181 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05371-9 (print), 3-540-36484-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 157 pp. LCCN QA3 .L28 no. 181.

URL http://link.springer.com/book/10.1007/BFb0061226;
 http://www.springerlink.com/content/978-3-540-36484-
9.

DeMeyer:1971:STE

[DI71f] Frank DeMeyer and Edward Ingraham. The six-term exact sequence. Lecture Notes in Mathematics, 181:115–130, 1971. CO-DEN LNMAA2. ISBN 3-540-05371-9 (print), 3-540-36484-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061230/.

Diamond:1972:TOT

[Dia72] Harold G. Diamond. Two oscillation theorems. Lecture Notes in Mathematics, 251:113–118, 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058788/.

Dieudonne:1972:TIA

[Die72] Jean Dieudonné. La théorie des invariants au XIX^{e} siècle. (French) []. Lecture Notes in Mathematics, 244:257–274, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058706/.

Dinges:1971:ILP

[Din71a] Hermann Dinges. Inequalities leading to a proof of the classical martingale-convergence theorem. Lecture Notes in Mathematics, 190:9–12, 1971. CODEN LNMAA2. ISBN 3-540-05396-4 (print), 3-540-36515-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065885/.

Dinges:1971:MRM

[Din71b] Hermann Dinges, editor. Martingales: A Report on a Meeting at Oberwolfach May 17–23, 1970, volume 190 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05396-4 (print), 3-540-36515-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 190.

URL http://link.springer.com/book/10.1007/BFb0065883;
 http://www.springerlink.com/content/978-3-540-36515-
0.

Dinolt:1971:EPN

[Din71c] George W. Dinolt. An extremal problem for non-separable matroids. Lecture Notes in Mathematics, 211:31–49, 1971. CODEN LNMAA2. ISBN 3-540-05545-2 (print), 3-540-36877-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061073/.

Dineen:1972:CHD

[Din72] Sean Dineen. Convexité holomorphe en dimension infinie. (French) []. Lecture Notes in Mathematics, 275:177–181, 1972. CODEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058501/.

Dineen:1973:HCL

[Din73] Seán Dineen. Holomorphically complete locally convex topological vector spaces. Lecture Notes in Mathematics, 332:77– 111, 1973. CODEN LNMAA2. ISBN 3-540-06372-2 (print), 3-540-46980-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060909/.

Dineen:1974:HFS

[Din74a] Seán Dineen. Holomorphic functions and surjective limits. Lecture Notes in Mathematics, 364:1–12, 1974. CODEN LN-MAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069001/.

Dinges:1974:SS

[Din74b] H. Dinges. Stopping sequences. Lecture Notes in Mathematics, 381:27–36, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057255/.

Dionne:1974:WFS

[Dio74] Ph. A. Dionne. Wave front sets and hypoelliptic operators. Lecture Notes in Mathematics, 399:84–93, 1974. CODEN LN-MAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063568/.

Dixmier:1971:AHM

[Dix71] Jacques Dixmier. Les algèbres hilbertiennes modulaires de Tomita. (French) []. Lecture Notes in Mathematics, 180:129– 143, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058828/.

Dixon:1973:FSL

[Dix73] John D. Dixon. Free subgroups of linear groups. Lecture Notes in Mathematics, 319:45–56, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058928/.

Dixmier:1974:CRI

[Dix74] Jacques Dixmier. Certaines représentations infinies des algèbres de Lie semi-simples. (French) []. Lecture Notes in Mathematics, 383:141–156, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057306/.

deJager:1972:SEP

[dJ72] E. M. de Jager. Singular elliptic perturbations of vanishing firstorder differential operators. Lecture Notes in Mathematics, 280: 73–86, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066920/.

Devlin:1974:CZF

[DJ74a] Dr. Keith Devlin and Dr. Håvard Johnsbråten. Con(ZF) → Con(ZF C + GCH + SH). Lecture Notes in Mathematics, 405: 113–121, 1974. CODEN LNMAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065989/.

Devlin:1974:BM

[DJ74b] Dr. Keith Devlin and Dr. Håvard Johnsbråten. Back matter. Lecture Notes in Mathematics, 405:122–132, 1974. CODEN LN-MAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/content/pdf/bbm:978-3-540-37822-8/
1.pdf.

Devlin:1974:CP

[DJ74c] Dr. Keith Devlin and Dr. Håvard Johnsbråten. The combinatorial property ◇. Lecture Notes in Mathematics, 405:23–31, 1974.

CODEN LNMAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065982/.

Devlin:1974:FM

[DJ74d] Dr. Keith Devlin and Dr. Håvard Johnsbråten. Front matter. Lecture Notes in Mathematics, 405:i–viii, 1974. CODEN LNMAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/content/pdf/bfm:978-3-540-37822-8/1.pdf.

Devlin:1974:HST

[DJ74e] Dr. Keith Devlin and Dr. Håvard Johnsbråten. Homogeneous Souslin trees and lines. Lecture Notes in Mathematics, 405:32– 45, 1974. CODEN LNMAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065983/.

Devlin:1974:HJK

[DJ74f] Dr. Keith Devlin and Dr. Håvard Johnsbråten. How Jensen killed a Souslin tree. Lecture Notes in Mathematics, 405:97–112, 1974. CODEN LNMAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065988/.

Devlin:1974:IFJ

[DJ74g] Dr. Keith Devlin and Dr. Håvard Johnsbråten. Iterated forcing Jensen style. Lecture Notes in Mathematics, 405:74–96, 1974. CODEN LNMAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065987/.

Devlin:1974:MAC

[DJ74h] Dr. Keith Devlin and Dr. Håvard Johnsbråten. Martin’s axiom and the consistency of SH. Lecture Notes in Mathematics, 405: 54–67, 1974. CODEN LNMAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065985/.

Devlin:1974:P

[DJ74i] Dr. Keith Devlin and Dr. Håvard Johnsbråten. Preliminaries. Lecture Notes in Mathematics, 405:1–7, 1974. CODEN LN-MAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065980/.

Devlin:1974:RST

[DJ74j] Dr. Keith Devlin and Dr. Håvard Johnsbråten. Rigid Souslin trees and lines. Lecture Notes in Mathematics, 405:46–53, 1974. CODEN LNMAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065984/.

Devlin:1974:SP

[DJ74k] Dr. Keith Devlin and Dr. Håvard Johnsbråten. The Souslin Problem, volume 405 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 132 pp. LCCN QA3 .L28 no. 405. URL http:/

/link.springer.com/book/10.1007/BFb0065979;   http://
www.springerlink.com/content/978-3-540-37822-8.

Devlin:1974:SH

[DJ74l] Dr. Keith Devlin and Dr. Håvard Johnsbråten. Souslin’s hypothesis. Lecture Notes in Mathematics, 405:8–22, 1974. CODEN LNMAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065981/.

Devlin:1974:TCZ

[DJ74m] Dr. Keith Devlin and Dr. Håvard Johnsbråten. Towards con(ZFC+CH+SH): a false start. Lecture Notes in Mathematics, 405:68–73, 1974. CODEN LNMAA2. ISBN 3-540-06860-0 (print), 3-540-37822-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065986/.

Deligne:1973:GMG

[DK73a] Pierre Deligne and Nicholas M. Katz. Groupes de Monodromie en Géométrie Algébrique, volume 340 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 438 pp. LCCN QA3 .L28 no. 340.

URL http://link.springer.com/book/10.1007/BFb0060505;
 http://www.springerlink.com/content/978-3-540-37764-
1.

Deligne:1973:MFO

[DK73b] Pierre Deligne and Willem Kuijk, editors. Modular Functions of One Variable II: Proceedings International Summer School University of Antwerp, RUCA July 17–August 3, 1972, volume 349 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06558-X (print), 3-540-37855-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 349. URL http://link.springer.com/book/
10.1007/978-3-540-37855-6;   http://www.springerlink.
com/content/978-3-540-37855-6.

Dowling:1974:ESM

[DK74] T. A. Dowling and D. G. Kelly. Elementary strong maps and transversal geometries. Lecture Notes in Mathematics, 411:176– 186, 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066193/.

delaHarpe:1972:CBL

[dlH72a] Pierre de la Harpe. Classical Banach–Lie Algebras and Banach– Lie Groups of Operators in Hilbert Space, volume 285 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05984-9 (print), 3-540-37970-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 160 pp. LCCN QA3 .L28 no. 285. URL http:/

/link.springer.com/book/10.1007/BFb0071306;   http://
www.springerlink.com/content/978-3-540-37970-6.

delaHarpe:1972:CIB

[dlH72b] Pierre de la Harpe. Classical involutive Banach–Lie algebras and groups of bounded and compact operators. Lecture Notes in Mathematics, 285:72–114, 1972. CODEN LNMAA2. ISBN 3-540-05984-9 (print), 3-540-37970-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071311/.

delaHarpe:1972:CIL

[dlH72c] Pierre de la Harpe. Classical involutive Lie algebras of finite rank operators. Lecture Notes in Mathematics, 285:23–71, 1972. CODEN LNMAA2. ISBN 3-540-05984-9 (print), 3-540-37970-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071310/.

delaHarpe:1972:DTC

[dlH72d] Pierre de la Harpe. Detailed table of contents. Lecture Notes in Mathematics, 285:16–20, 1972. CODEN LNMAA2. ISBN 3-540-05984-9 (print), 3-540-37970-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071308/.

delaHarpe:1972:EID

[dlH72e] Pierre de la Harpe. Examples of infinite dimensional Hilbert symmetric spaces. Lecture Notes in Mathematics, 285:115–133, 1972. CODEN LNMAA2. ISBN 3-540-05984-9 (print), 3-540-37970-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071312/.

delaHarpe:1972:I

[dlH72f] Pierre de la Harpe. Introduction. Lecture Notes in Mathematics, 285:1–15, 1972. CODEN LNMAA2. ISBN 3-540-05984-9 (print), 3-540-37970-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071307/.

delaHarpe:1972:CCC

[dlH72g] Pierre de la Harpe. On the cohomology of the classical complex Lie algebras of compact operators. Lecture Notes in Mathematics, 285:134–146, 1972. CODEN LNMAA2. ISBN 3-540-05984-9 (print), 3-540-37970-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071313/.

delaHarpe:1972:SNC

[dlH72h] Pierre de la Harpe. Some notations and conventions. Lecture Notes in Mathematics, 285:21–22, 1972. CODEN LN-MAA2. ISBN 3-540-05984-9 (print), 3-540-37970-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071309.

deMontcheuil:1970:ACO

[dM70] J. Morreau de Montcheuil. Application de contrôle optimal aux trajectoires de Suées. (French) []. Lecture Notes in Mathematics, 132:264–280, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066686/.

deSamLazaro:1972:GPM

[dM72] J. de Sam Lazaro and P.-A. Meyer. Un gros processus de Markov application à certains flots. (French) []. Lecture Notes in Mathematics, 258:109–112, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059467/.

Desolneux-Moulis:1974:TFI

[DM74] Madame N. Desolneux-Moulis. Un théorème de fonctions implicites dans les espaces de Fréchet. Application à un problème de déformation de structures complexes. (French) []. Lecture Notes in Mathematics, 410:1–14, 1974. CODEN LNMAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066035/.

Dellacherie:1973:PFA

[DMW73a] C. Dellacherie, P.-A. Meyer, and M. Weil. Potentiels de fonctionnelles additives. Un contre-exemple de Knight (exposé de C. Dellacherie). (French) []. Lecture Notes in Mathematics, 321:58– 60, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071396/.

Dellacherie:1973:SMM

[DMW73b] C. Dellacherie, P.-A. Meyer, and M. Weil. A semi-Markovian model for the Brownian motion. Lecture Notes in Mathematics, 321:vii–viii, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/

10.1007/BFb0071388;   http://link.springer.com/content/
pdf/bfm:978-3-540-40023-3/1.pdf.

Dellacherie:1973:SPV

[DMW73c] C. Dellacherie, P.-A. Meyer, and M. Weil. Séminaire de Probabilités VII, volume 321 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 322 pp. LCCN QA3 .L28 no. 321.

URL http://link.springer.com/book/10.1007/BFb0071387;
 http://www.springerlink.com/content/978-3-540-40023-
3.

Dellacherie:1974:SPV

[DMW74] C. Dellacherie, P.-A. Meyer, and M. Weil. Séminaire de Probabilités VIII Université de Strasbourg, volume 381 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 354 pp. LCCN QA3 .L28 no. 381. URL http:/

/link.springer.com/book/10.1007/BFb0057249;   http://
www.springerlink.com/content/978-3-540-38384-0.

Dobbs:1970:CCD

[Dob70a] David E. Dobbs. Cech Cohomological Dimensions for Commutative Rings, volume 147 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04936-3 (print), 3-540-36310-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 176 pp. LCCN QA3 .L28 no. 147.

URL http://link.springer.com/book/10.1007/BFb0060274;
 http://www.springerlink.com/content/978-3-540-36310-
1.

Dobbs:1970:CDF

[Dob70b] David E. Dobbs. Cohomological dimension of fields. Lecture Notes in Mathematics, 147:1–68, 1970. CODEN LNMAA2. ISBN 3-540-04936-3 (print), 3-540-36310-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060275/.

Dobbs:1970:GCD

[Dob70c] David E. Dobbs. A generalization of cohomological dimension for rings. Lecture Notes in Mathematics, 147:109–140, 1970. CODEN LNMAA2. ISBN 3-540-04936-3 (print), 3-540-36310-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060277/.

Dobbs:1970:NTA

[Dob70d] David E. Dobbs. Number theoretic applications of a Cech dimension theory. Lecture Notes in Mathematics, 147:141–171, 1970. CODEN LNMAA2. ISBN 3-540-04936-3 (print), 3-540-36310-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060278/.

Dobbs:1970:CDT

[Dob70e] David E. Dobbs. On Cech dimension theories for rings. Lecture Notes in Mathematics, 147:69–108, 1970. CODEN LN-MAA2. ISBN 3-540-04936-3 (print), 3-540-36310-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060276/.

Dolbeault:1970:TRH

[Dol70a] Pierre Dolbeault. Theory of residues and homology. Lecture Notes in Mathematics, 116:152–163, 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059252/.

Doleans:1970:DCC

[Dol70b] Catherine Doléans. Diffusions à coefficients continus: le problème d’après un exposé de Catherine Doléans. (French) []. Lecture Notes in Mathematics, 124:241–251, 1970. CODEN LN-MAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059347/.

Dolezal:1970:RLC

[Dol70c] Vaclav Dolezal. On a representation of linear continuous operators defined on distributions. Lecture Notes in Mathematics, 144: 44–48, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059919/.

Dolbeault:1971:CRF

[Dol71a] Pierre Dolbeault. Courants résidus des formes semi-méromorphes. (French) []. Lecture Notes in Mathematics, 205:56–70, 1971. CO-DEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059833/.

Dolbeault:1971:TRF

[Dol71b] Pierre Dolbeault. Théorie des résidus. (French) []. Lecture Notes in Mathematics, 205:232–243, 1971. CODEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059843/.

Dolbeault:1972:VPR

[Dol72] Pierre Dolbeault. Valeurs principales et résidus sur les espaces analytiques complexes d’après M. Herrera et D. Lieberman [6]. (French) []. Lecture Notes in Mathematics, 275:14–26, 1972. CO-DEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058490/.

Domar:1973:SSC

[Dom73] Yngve Domar. On spectral synthesis in commutative Banach algebras using closed ideals of finite co-dimension. Lecture Notes in Mathematics, 336:63–78, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065788/.

Doob:1971:SSM

[Doo71] J. L. Doob. State space for Markov processes. Lecture Notes in Mathematics, 190:13–14, 1971. CODEN LNMAA2. ISBN 3-540-05396-4 (print), 3-540-36515-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065886.

Dostal:1971:EAU

[Dos71] M. A. Dostal. Espaces analytiquement uniformes. (French) []. Lecture Notes in Mathematics, 205:113–128, 1971. CODEN LN-MAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059836/.

Dostal:1974:SRR

[Dos74] M. A. Dostal. Some recent results on topological vector spaces. Lecture Notes in Mathematics, 384:20–91, 1974. CODEN LN-MAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057540/.

Douady:1971:PFA

[Dou71]  Adrien Douady.  Prolongement de faisceaux analytiques

cohérents. (French) []. Lecture Notes in Mathematics, 180:39– 54, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058823/.

Douady:1973:TID

[Dou73] Adrien Douady. Le théorème des images directes de Grauert [d’après Kiehl–Verdier]. (French) []. Lecture Notes in Mathematics, 317:73–87, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069277/.

Douglas:1973:ISN

[DP73] R. G. Douglas and Carl Pearcy. Invariant subspaces of nonquasitriangular operators. Lecture Notes in Mathematics, 345: 13–57, 1973. CODEN LNMAA2. ISBN 3-540-06496-6 (print), 3-540-37808-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058916/.

Dlab:1972:BR

[DR72] Vlastimil Dlab and Claus Michael Ringel. Balanced rings. Lecture Notes in Mathematics, 246:73–143, 1972. CODEN LN-MAA2. ISBN 3-540-05760-9 (print), 3-540-37181-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059564/.

Deligne:1973:SMC

[DR73] P. Deligne and M. Rapoport. Les schémas de modules de courbes elliptiques. (French) []. Lecture Notes in Mathematics, 349:143– 316, 1973. CODEN LNMAA2. ISBN 3-540-06558-X (print), 3-540-37855-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-37855-6_4/.

Dress:1973:CTI

[Dre73] Andreas W. M. Dress. Contributions to the theory of induced representations. Lecture Notes in Mathematics, 342:181–240, 1973. CODEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073727/.

Dror:1971:GWT

[Dro71] Emmanuel Dror. A generalization of the Whitehead theorem. Lecture Notes in Mathematics, 249:13–22, 1971. CODEN LN-MAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060891/.

Drury:1972:USI

[Dru72] S. W. Drury. Unions of sets of interpolation. Lecture Notes in Mathematics, 266:23–33, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059634/.

Dubuc:1970:CFE

[DS70a] Eduardo Dubuc and Ross Street. Categorically, the final examination for the summer institute at Bowdoin College (Maine) 1969. Lecture Notes in Mathematics, 137:138–139, 1970. CODEN LNMAA2. ISBN 3-540-04926-6 (print), 3-540-36292-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060444.

Dubuc:1970:DT

[DS70b] Eduardo Dubuc and Ross Street. Dinatural transformations. Lecture Notes in Mathematics, 137:126–137, 1970. CODEN LN-MAA2. ISBN 3-540-04926-6 (print), 3-540-36292-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060443/.

Dennis:1973:FSC

[DS73] R. Keith Dennis and Michael R. Stein. The functor K_{2}: a survey of computations and problems. Lecture Notes in Mathematics, 342:241–303, 1973. CODEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073728/.

Lazaro:1970:TSE

[dSL70] J. de Sam Lazaro. Théorème de Stone et espérances conditionnelles. (French) []. Lecture Notes in Mathematics, 124:132– 150, 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059339/.

Dubuc:1970:CCb

[Dub70a] Eduardo J. Dubuc. Complete categories. Lecture Notes in Mathematics, 145:111–148, 1970. CODEN LNMAA2. ISBN 3-540-04934-7 (print), 3-540-36307-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060489/.

Dubuc:1970:CCa

[Dub70b] Eduardo J. Dubuc. Completeness concepts. Lecture Notes in Mathematics, 145:7–59, 1970. CODEN LNMAA2. ISBN 3-540-04934-7 (print), 3-540-36307-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060487/.

Dubuc:1970:FC

[Dub70c] Eduardo J. Dubuc. Functor categories. Lecture Notes in Mathematics, 145:149–166, 1970. CODEN LNMAA2. ISBN 3-540-04934-7 (print), 3-540-36307-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060490/.

Dubuc:1970:KEE

[Dub70d] Eduardo J. Dubuc. Kan extensions in Enriched Category Theory, volume 145 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04934-7 (print), 3-540-36307-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 173 pp. LCCN QA3 .L28 no. 145. URL http:/

/link.springer.com/book/10.1007/BFb0060485;   http://
www.springerlink.com/content/978-3-540-36307-1.

Dubuc:1970:Mb

[Dub70e] Eduardo J. Dubuc. V -monads. Lecture Notes in Mathematics, 145:60–110, 1970. CODEN LNMAA2. ISBN 3-540-04934-7 (print), 3-540-36307-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060488/.

Dubuc:1970:Ma

[Dub70f] Eduardo J. Dubuc. V -monomorphisms. Lecture Notes in Mathematics, 145:1–6, 1970. CODEN LNMAA2. ISBN 3-540-04934-7 (print), 3-540-36307-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060486/.

Dudley:1970:RLF

[Dud70] R. M. Dudley. Random linear functionals: Some recent results. Lecture Notes in Mathematics, 170:62–70, 1970. CODEN LN-MAA2. ISBN 3-540-05284-4 (print), 3-540-36417-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079066/.

Dunkel:1970:FDE

[Dun70a] Gregory M. Dunkel. Function differential equations: Examples and problems. Lecture Notes in Mathematics, 144:49–63, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059920/.

Dunkel:1970:MSY

[Dun70b] Gregory M. Dunkel. Maximum sustainable yields as a control problem in integral equations. Lecture Notes in Mathematics, 144:64–69, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059921/.

Duncan:1972:SBV

[Dun72] T. E. Duncan. Some Banach-valued processes with applications. Lecture Notes in Mathematics, 294:85–96, 1972. CODEN LN-MAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064936/.

Dupont:1974:GMM

[Dup74a] Todd Dupont. Galerkin methods for modeling gas pipelines. Lecture Notes in Mathematics, 430:112–130, 1974. CODEN LN-MAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066267/.

Dupuis:1974:MHT

[Dup74b] Claire Dupuis. Mesure de Hausdorff de la trajectoire de certains processus à accroissements indépendants et stationnaires. (French) []. Lecture Notes in Mathematics, 381:37–77, 1974. CO-DEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057256/.

deVeubeke:1970:MMP

[dV70] B. Fraeijs de Veubeke. A maximum-minimum principle for Bang– Bang systems. Lecture Notes in Mathematics, 112:225–247, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060207/.

Dwork:1973:OAM

[Dwo73] B. Dwork. The U_{p} operator of Atkin on modular functions of level 2 with growth conditions. Lecture Notes in Mathematics, 350: 57–67, 1973. CODEN LNMAA2. ISBN 3-540-06483-4 (print), 3-540-37802-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-37802-0_2/.

Dwyer:1974:PDE

[Dwy74] Thomas A. W. Dwyer III. Partial differential equations in holomorphic Fock spaces. Lecture Notes in Mathematics, 384:252– 259, 1974. CODEN LNMAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057551/.

Dyer:1973:SPS

[Dye73] R. H. Dyer. Some properties of solutions of the Navier–Stokes equations. Lecture Notes in Mathematics, 316:67–84, 1973. CO-DEN LNMAA2. ISBN 3-540-06159-2 (print), 3-540-38370-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069624/.

Dzhrbashyan:1974:HAC

[Dzh74] M. M. Dzhrbashyan. Harmonic analysis in the complex domain and applications in the theory of analytical and infinitely differentiable functions. Lecture Notes in Mathematics, 399:94–118, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063569/.

Earle:1974:QMU

[Ear74] C. J. Earle. Quasiconformal mappings and uniformization. Lecture Notes in Mathematics, 400:15–23, 1974. CODEN LN-MAA2. ISBN 3-540-06840-6 (print), 3-540-37776-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065673/.

Eastham:1974:ESA

[Eas74] M. S. P. Eastham. The essential self-adjointness of Schrödinger operators. Lecture Notes in Mathematics, 415:109–117, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065516/.

Edelbaum:1970:OTL

[Ede70] T. N. Edelbaum. Optimum thrust-limited orbit transfer in strong gravity fields. Lecture Notes in Mathematics, 132:70–78, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066675/.

Edmunds:1974:NEE

[Edm74] D. E. Edmunds. Nonlinear elliptic equations. Lecture Notes in Mathematics, 415:118–129, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065517/.

Edwards:1974:EHT

[Edw74] David A. Edwards. Etale homotopy theory and shape. Lecture Notes in Mathematics, 428:58–107, 1974. CODEN LN-MAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070535/.

Eisenbud:1973:TCA

[EE73] David Eisenbud and E. Graham Evans, Jr. Three conjectures about modules over polynomial rings. Lecture Notes in Mathematics, 311:78–89, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068921/.

Edmunds:1973:GEF

[EG73] C. C. Edmunds and N. D. Gupta. On groups of exponent four IV. Lecture Notes in Mathematics, 319:57–70, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058929/.

Everitt:1974:IAC

[EG74] W. N. Everitt and M. Giertz. Inequalities associated with certain partial differential operators. Lecture Notes in Mathematics, 415: 133–136, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065519/.

Eakin:1973:CPR

[EH73] Paul Eakin and William Heinzer. A cancellation problem for rings. Lecture Notes in Mathematics, 311:61–77, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068920/.

Engelking:1970:TWO

[EHM70] R. Engelking, R. W. Heath, and E. Michael. Topological wellordering and continuous selections. Lecture Notes in Mathematics, 171:8–11, 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069712/.

Edwards:1972:LCG

[EHR72] Robert E. Edwards, Edwin Hewitt, and Kenneth A. Ross. Lacunarity for compact groups, III. Lecture Notes in Mathematics, 266:35–42, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059635/.

Erdos:1973:CNG

[EHR73] P. Erdős, A. Hajnal, and B. Rothchild. On chromatic number of graphs and set systems. Lecture Notes in Mathematics, 337: 531–538, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066788/.

Eichler:1970:AMT

[Eic70] M. Eichler. Algebraic methods in the theory of modular forms. Lecture Notes in Mathematics, 155:88–96, 1970. CODEN LN-MAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060316/.

Eichler:1971:AMF

[Eic71a] Martin Eichler. Applications to modular forms. Lecture Notes in Mathematics, 210:88–114, 1971. CODEN LNMAA2. ISBN 3-540-05519-3 (print), 3-540-36869-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071045/.

Eichler:1971:GM

[Eic71b] Martin Eichler. Graded modules. Lecture Notes in Mathematics, 210:4–48, 1971. CODEN LNMAA2. ISBN 3-540-05519-3 (print), 3-540-36869-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071043/.

Eichler:1971:GRI

[Eic71c] Martin Eichler. Graded rings and ideals. Lecture Notes in Mathematics, 210:49–87, 1971. CODEN LNMAA2. ISBN 3-540-05519-3 (print), 3-540-36869-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071044/.

Eichler:1971:I

[Eic71d] Martin Eichler. Introduction. Lecture Notes in Mathematics, 210:1–3, 1971. CODEN LNMAA2. ISBN 3-540-05519-3 (print), 3-540-36869-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071042/.

Eichler:1971:PVM

[Eic71e] Martin Eichler, editor. Projective Varieties and Modular Forms: Course Given at the University of Maryland, Spring 1970, volume 210 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05519-3 (print), 3-540-36869-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 210. URL http:/

/link.springer.com/book/10.1007/BFb0071041;   http://
www.springerlink.com/content/978-3-540-36869-4.

Eichler:1973:BPM

[Eic73] M. Eichler. The basis problem for modular forms and the traces of the Hecke operators. Lecture Notes in Mathematics, 320:75– 151, 1973. CODEN LNMAA2. ISBN 3-540-06219-X (print), 3-540-38509-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38509-7_4/.

Eisenberg:1971:GLB

[Eis71a] Bernard Eisenberg. Generalized lower bounds for the absolute values of the coefficients of chromatic polynomials. Lecture Notes in Mathematics, 186:85–94, 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059427/.

Eisenberg:1971:TEM

[Eis71b] Murray Eisenberg. A theorem on extensions of minimal sets. Lecture Notes in Mathematics, 235:61–64, 1971. CODEN LN-MAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059196/.

Ekeland:1970:SMS

[Eke70] Ivar Ekeland. Suites minimisantes et solutions de certains problèmes de contrôle optimal. (French) []. Lecture Notes in Mathematics, 132:79–84, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066676/.

Elman:1973:QSH

[EL73] Richard Elman and T. Y. Lam. On the quaternion symbol homomorphism g_{F}: k_{2}F → B(F). Lecture Notes in Mathematics, 342: 447–463, 1973. CODEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073734/.

Elgot:1971:ATP

[Elg71] Calvin C. Elgot. Algebraic theories and program schemes. Lecture Notes in Mathematics, 188:71–88, 1971. CODEN LN-MAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059694/.

Ellis:1973:GAE

[Ell73] Robert Ellis. A group associated with an extension. Lecture Notes in Mathematics, 318:86–94, 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061725/.

Endl:1974:FKD

[End74] Kurt Endl. Über die funktionentheoretische Konstruktion von dicht-vollständigen Systemen in allgemeinen normierten Räumen. (German) []. Lecture Notes in Mathematics, 419:80– 89, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064714/.

Engeler:1970:PTA

[Eng70] Erwin Engeler. Proof theory and the accuracy of computations. Lecture Notes in Mathematics, 125:62–72, 1970. CODEN LN-MAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060624/.

Engeler:1971:SME

[Eng71a] Erwin Engeler. Structure and meaning of elementary programs. Lecture Notes in Mathematics, 188:89–101, 1971. CODEN LN-MAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059695/.

Engeler:1971:SSA

[Eng71b] Erwin Engeler. Symposium on Semantics of Algorithmic Languages, volume 188 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 0-387-05377-8, 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 372 pp. LCCN QA3 .L28 no. 188.

URL http://link.springer.com/book/10.1007/BFb0059689;
 http://www.springerlink.com/content/978-3-540-36499-
3.

Engels:1973:SIG

[Eng73] Hermann Engels. Spezielle interpolationsquadraturen vom gauss’schen typ. (German) []. Lecture Notes in Mathematics, 333: 54–68, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060687/.

Erdos:1970:SAP

[ER70] P. Erdős and A. Rényi. On some applications of probability methods to additive number theoretic problems. Lecture Notes in Mathematics, 160:37–44, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060644/.

Erdos:1974:EPG

[Erd74] Paul Erdős. Extremal problems on graphs and hypergraphs. Lecture Notes in Mathematics, 411:75–84, 1974. CODEN LN-MAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066181/.

Ermoliev:1970:ROS

[Erm70] Ju. Ermoliev. Random optimization and stochastic programming. Lecture Notes in Mathematics, 112:104–115, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060200/.

Erdos:1972:PMH

[ES72a] P. Erdős and S. Shelah. On problems of Moser and Hanson. Lecture Notes in Mathematics, 303:75–79, 1972. CODEN LN-MAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067359/.

Erdos:1972:ISA

[ES72b] P. Erdős and M. V. Subbarao. On the iterates of some arithmetic functions. Lecture Notes in Mathematics, 251:119–125, 1972. CO-DEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058789/.

Everitt:1972:CTO

[ES72c] W. N. Everitt and B. D. Sleeman, editors. Conference on the Theory of Ordinary and Partial Differential Equations: Held in Dundee/Scotland, March 28–31, 1972, volume 280 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 280.

URL http://link.springer.com/book/10.1007/BFb0066914;
 http://www.springerlink.com/content/978-3-540-37618-
7.

Espinosa:1974:CSI

[Esp74] Ruben J. Espinosa. Collocation solutions of integro-differential equations. Lecture Notes in Mathematics, 430:154–183, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066269/.

Esser:1973:DEG

[Ess73] H. Esser. Zur Diskretisierung Von Extremalproblemen. (German) []. Lecture Notes in Mathematics, 333:69–88, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060688/.

Eisenstat:1974:ASM

[ESS74a] S. C. Eisenstat, M. H. Schultz, and A. H. Sherman. The application of sparse matrix methods to the numerical solution of nonlinear elliptic partial differential equations. Lecture Notes in Mathematics, 430:131–153, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066268/.

Essen:1974:DIA

[Ess74b] Matts Essén. A differential inequality with applications to subharmonic functions. Lecture Notes in Mathematics, 415:130– 132, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065518/.

Essen:1974:GAH

[Ess74c] Matts Essén. The generalized Ahlfors–Heins theorem in R^{3}. Lecture Notes in Mathematics, 419:90–98, 1974. CODEN LN-MAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064715/.

Evans:1971:NSS

[Eva71] D. J. Evans. Numerical solution of the Sturm Liouville problem with periodic boundary conditions. Lecture Notes in Mathematics, 228:277–289, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069462/.

Exbrayat:1970:FAE

[Exb70] Jean-Marie Exbrayat. Fonctions analytiques sur un espace de Banach. (French) []. Lecture Notes in Mathematics, 116:30–38, 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059243/.

Eymard:1971:ACF

[Eym71] Pierre Eymard. Algèbres A_{p} et convoluteurs de L^{p}. (French) []. Lecture Notes in Mathematics, 180:55–72, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058824/.

Eymard:1972:CDF

[Eym72a] Pierre Eymard. Compléments divers. (French) [Various supplements]. Lecture Notes in Mathematics, 300:83–109, 1972. CO-DEN LNMAA2. ISBN 3-540-06086-3 (print), 3-540-38071-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060756/.

Eymard:1972:EHMa

[Eym72b] Pierre Eymard. Espaces homogènes moyennables et propriétés de Reiter. (French) []. Lecture Notes in Mathematics, 300:21– 39, 1972. CODEN LNMAA2. ISBN 3-540-06086-3 (print), 3-540-38071-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060753/.

Eymard:1972:EHMb

[Eym72c] Pierre Eymard. Espaces homogènes moyennables et représentations quasi-regulières. (French) []. Lecture Notes in Mathematics, 300: 40–59, 1972. CODEN LNMAA2. ISBN 3-540-06086-3 (print), 3-540-38071-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060754/.

Eymard:1972:IFI

[Eym72d] Pierre Eymard. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 300:3–5, 1972. CODEN LNMAA2. ISBN 3-540-06086-3 (print), 3-540-38071-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060751/.

Eymard:1972:MIP

[Eym72e] Pierre Eymard. Moyennes invariantes et propriétés du point fixe. (French) []. Lecture Notes in Mathematics, 300:6–20, 1972. CO-DEN LNMAA2. ISBN 3-540-06086-3 (print), 3-540-38071-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060752/.

Eymard:1972:MIRa

[Eym72f] Pierre Eymard. Moyennes invariantes et représentations des produits semi-directs. (French) []. Lecture Notes in Mathematics, 300:60–82, 1972. CODEN LNMAA2. ISBN 3-540-06086-3 (print), 3-540-38071-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060755/.

Eymard:1972:MIRb

[Eym72g] Pierre Eymard. Moyennes Invariantes et Représentations Unitaires. (French) [], volume 300 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06086-3 (print), 3-540-38071-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 113 pp. LCCN QA3 .L28 no. 300.

URL http://link.springer.com/book/10.1007/BFb0060750;
 http://www.springerlink.com/content/978-3-540-38071-
9.

Faber:1974:HR

[Fab74] Vance Faber. Hypergraph reconstruction. Lecture Notes in Mathematics, 411:85–94, 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066182/.

Faith:1972:MFE

[Fai72] Carl Faith. Modules finite over endomorphism ring. Lecture Notes in Mathematics, 246:145–189, 1972. CODEN LNMAA2. ISBN 3-540-05760-9 (print), 3-540-37181-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059565/.

Faierman:1974:ETM

[Fai74] M. Faierman. The expansion theorem in multi-parameter Sturm– Liouville theory. Lecture Notes in Mathematics, 415:137–142, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065520/.

Fan:1970:ETF

[Fan70] Ky Fan. Extensions of two fixed point theorems of F. E. Browder. Lecture Notes in Mathematics, 171:12–16, 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069713/.

Fary:1972:GAB

[Fár72] I. Fáry. Group action and Betti sheaf. Lecture Notes in Mathematics, 299:311–322, 1972. CODEN LNMAA2. ISBN 3-540-06078-2 (print), 3-540-38066-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066767/.

Faraut:1973:FBV

[Far73] Jacques Faraut. Fonction brownienne sur une variété riemannienne. (French) []. Lecture Notes in Mathematics, 321:61–76, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071397/.

Faraut:1974:TDP

[Far74] Jacques Faraut. Théorie du Potentiel et Analyse Harmonique, volume 404 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 245 (est.) pp. LCCN QA3 .L28 no. 404. URL http:/

/link.springer.com/book/10.1007/BFb0060606;   http://
www.springerlink.com/content/978-3-540-37789-4.

Faurre:1970:IPM

[Fau70] P. Faurre. Identification par minimisation d’une représentation Markovienne de processus aléatoire. (French) []. Lecture Notes in Mathematics, 132:85–107, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066677/.

Fay:1973:BRS

[Fay73a] John D. Fay. Bordered Riemann surfaces. Lecture Notes in Mathematics, 352:108–133, 1973. CODEN LNMAA2. ISBN 3-540-06517-2 (print), 3-540-37815-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060096/.

Fay:1973:CUC

[Fay73b] John D. Fay. Cyclic unramified coverings. Lecture Notes in Mathematics, 352:61–84, 1973. CODEN LNMAA2. ISBN 3-540-06517-2 (print), 3-540-37815-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060094/.

Fay:1973:DRS

[Fay73c] John D. Fay. Degenerate Riemann surfaces. Lecture Notes in Mathematics, 352:37–60, 1973. CODEN LNMAA2. ISBN 3-540-06517-2 (print), 3-540-37815-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060093/.

Fay:1973:PF

[Fay73d] John D. Fay. The prime-form. Lecture Notes in Mathematics, 352: 16–36, 1973. CODEN LNMAA2. ISBN 3-540-06517-2 (print), 3-540-37815-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060092/.

Fay:1973:RDC

[Fay73e] John D. Fay. Ramified double coverings. Lecture Notes in Mathematics, 352:85–107, 1973. CODEN LNMAA2. ISBN 3-540-06517-2 (print), 3-540-37815-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060095/.

Fay:1973:RTF

[Fay73f] John D. Fay. Riemann’s theta function. Lecture Notes in Mathematics, 352:1–15, 1973. CODEN LNMAA2. ISBN 3-540-06517-2 (print), 3-540-37815-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060091/.

Fay:1973:TFR

[Fay73g] John D. Fay. Theta Functions on Riemann Surfaces, volume 352 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06517-2 (print), 3-540-37815-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 137 pp. LCCN QA3 .L28 no. 352. URL http:

//link.springer.com/book/10.1007/BFb0060090;   http:/
/www.springerlink.com/content/978-3-540-37815-0.

Feferman:1972:IPL

[Fef72] Solomon Feferman. Infinitary properties, local functors, and systems of ordinal functions. Lecture Notes in Mathematics, 255: 63–97, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059538/.

Feichtinger:1970:PT

[Fei70] Oskar Feichtinger. Properties of the λ topology. Lecture Notes in Mathematics, 171:17–23, 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069714/.

Feldman:1970:ACS

[Fel70] Professor Jacob Feldman. Absolute continuity of stochastic processes. Lecture Notes in Mathematics, 170:71–86, 1970. CODEN LNMAA2. ISBN 3-540-05284-4 (print), 3-540-36417-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079067/.

Felgner:1971:CEZ

[Fel71a] Ulrich Felgner. Cohen extensions of ZF-models. Lecture Notes in Mathematics, 223:76–166, 1971. CODEN LNMAA2. ISBN 3-540-05591-6 (print), 3-540-36908-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061164/.

Felgner:1971:CS

[Fel71b] Ulrich Felgner. Constructible sets. Lecture Notes in Mathematics, 223:23–45, 1971. CODEN LNMAA2. ISBN 3-540-05591-6 (print), 3-540-36908-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061162/.

Felgner:1971:FMS

[Fel71c] Ulrich Felgner. Fraenkel–Mostowski–Specker models. Lecture Notes in Mathematics, 223:46–75, 1971. CODEN LN-MAA2. ISBN 3-540-05591-6 (print), 3-540-36908-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061163/.

Felgner:1971:MZS

[Fel71d] Ulrich Felgner. Models of ZF-Set Theory, volume 223 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05591-6 (print), 3-540-36908-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 173 pp. LCCN QA3 .L28 no. 223. URL http:/

/link.springer.com/book/10.1007/BFb0061160;   http://
www.springerlink.com/content/978-3-540-36908-0.

Felgner:1971:P

[Fel71e] Ulrich Felgner. Prerequisites. Lecture Notes in Mathematics, 223: 1–22, 1971. CODEN LNMAA2. ISBN 3-540-05591-6 (print), 3-540-36908-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061161/.

Fell:1972:NLM

[Fel72] J. M. G. Fell. A new look at Mackey’s imprimitivity theorem. Lecture Notes in Mathematics, 266:43–58, 1972. CODEN LN-MAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059636/.

Fenyo:1974:DE

[Fen74] I. Fenyö. On the differential equation ∑_{r=0}^{n} c_{r}(pD_{1} + qD_{2})^{r}u = 0.

Lecture Notes in Mathematics, 415:143, 1974. CODEN LN-MAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065521.

Ferrier:1970:FTF

[Fer70a] J. P. Ferrier. Fonctions tempérées. (French) []. Lecture Notes in Mathematics, 164:11–16, 1970. CODEN LNMAA2. ISBN 3-540-05191-0 (print), 3-540-36376-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069599/.

Ferrier:1970:GSB

[Fer70b] J. P. Ferrier. Généralités sur les structures à bornes. (French) []. Lecture Notes in Mathematics, 164:5–10, 1970. CODEN LNMAA2. ISBN 3-540-05191-0 (print), 3-540-36376-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069598/.

Ferrier:1970:CSF

[Fer70c] J. P. Ferrier. Le calcul symbolique. (French) []. Lecture Notes in Mathematics, 164:23–34, 1970. CODEN LNMAA2. ISBN 3-540-05191-0 (print), 3-540-36376-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069601/.

Ferrier:1970:SWF

[Fer70d] J. P. Ferrier. Le spectre de Waelbroeck. (French) []. Lecture Notes in Mathematics, 164:17–22, 1970. CODEN LNMAA2. ISBN 3-540-05191-0 (print), 3-540-36376-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069600/.

Ferrier:1970:SAC

[Fer70e] J.-P. Ferrier. Séminaire sur les Algèbres Complètes. (French) [Seminar on Complete Algebras], volume 164 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05191-0 (print), 3-540-36376-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 69 pp. LCCN QA3 .L28 no. 164. URL http:/

/link.springer.com/book/10.1007/BFb0069597;   http://
www.springerlink.com/content/978-3-540-36376-7.

Ferrier:1970:ESS

[Fer70f] Jean-Pierre Ferrier. Ensembles spectraux d’une sous-algèbre; applications. (French) []. Lecture Notes in Mathematics, 164:42– 53, 1970. CODEN LNMAA2. ISBN 3-540-05191-0 (print), 3-540-36376-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069603/.

Ferrier:1970:FSP

[Fer70g] Jean-Pierre Ferrier. Fonctions spectrales pour z dans une sousalgèbre de θ(δ). (French) []. Lecture Notes in Mathematics, 164: 35–41, 1970. CODEN LNMAA2. ISBN 3-540-05191-0 (print), 3-540-36376-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069602/.

Ferrier:1970:PSA

[Fer70h] Jean-Pierre Ferrier. Un problème de spectre dans les algèbres complètes. (French) []. Lecture Notes in Mathematics, 116:101– 118, 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059248/.

Ferrier:1972:QCF

[Fer72] J. P. Ferrier. Questions de convexite. (French) []. Lecture Notes in Mathematics, 277:77–117, 1972. CODEN LNMAA2. ISBN 3-540-05934-2 (print), 3-540-37613-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058879/.

Ferguson:1974:MST

[Fer74a] Le Baron O. Ferguson. Müntz–Szasz theorem with integral coefficients I. Lecture Notes in Mathematics, 399:119–122, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063570/.

Fernique:1974:DSD

[Fer74b] X. Fernique. Une démonstration simple du théorème de R. M. Dudley et Marek Kanter sur les lois zéro–un pour les mesures stables. (French) []. Lecture Notes in Mathematics, 381:78–79, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057257.

Ferrier:1974:ILB

[Fer74c] J.-P. Ferrier. Inductive limits of Banach spaces and complex analysis. Lecture Notes in Mathematics, 399:123–128, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063571/.

Ferrier:1974:IFH

[Fer74d] Jean-Pierre Ferrier. Idéaux de fonctions holomorphes avec poids et calcul fonctionnel. (French) []. Lecture Notes in Mathematics, 410:86–96, 1974. CODEN LNMAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066042/.

Fesmire:1974:RNO

[Fes74] Steven H. Fesmire. Representation of nonlinear operators with the Hammerstein property. Lecture Notes in Mathematics, 399: 129–148, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063572/.

Frucht:1971:LNE

[FGQ71] Roberto Frucht, Allan Gewirtz, and Louis V. Quintas. The least number of edges for graphs having automorphism group of order three. Lecture Notes in Mathematics, 186:95–104, 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059428/.

Fichera:1972:AUP

[Fic72] G. Fichera. Abstract unilateral problems. Lecture Notes in Mathematics, 280:51–64, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066918/.

Fichera:1974:AAF

[Fic74] Gaetano Fichera. On the approximation of analytic functions by rational functions. Lecture Notes in Mathematics, 419:99– 109, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064716/.

Fife:1973:SPQ

[Fif73] Paul C. Fife. Singular perturbation by a quasilinear operator. Lecture Notes in Mathematics, 322:87–100, 1973. CODEN LN-MAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060562/.

Fillmore:1973:PCO

[Fil73] P. A. Fillmore, editor. Proceedings of a Conference on Operator Theory: Dalhousie University, Halifax, Nova Scotia April 13th and 14th, 1973, volume 345 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06496-6 (print), 3-540-37808-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 345. URL http:/

/link.springer.com/book/10.1007/BFb0058913;   http://
www.springerlink.com/content/978-3-540-37808-2.

Fine:1972:LCF

[Fin72] Kit Fine. Logics containing S4 without the finite model property. Lecture Notes in Mathematics, 255:98–102, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059539/.

Fink:1974:APD

[Fin74a] A. M. Fink. Almost Periodic Differential Equations, volume 377 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 285 (est.) pp. LCCN QA3 .L28 no. 377. URL http:/

/link.springer.com/book/10.1007/BFb0070324;   http://
www.springerlink.com/content/978-3-540-38307-9.

Fink:1974:APF

[Fin74b] A. M. Fink. Almost periodic functions. Lecture Notes in Mathematics, 377:1–15, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070325/.

Fink:1974:AAP

[Fin74c] A. M. Fink. Asymptotic almost periodic functions and other weaker conditions. Lecture Notes in Mathematics, 377:154– 169, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070333/.

Fink:1974:A

[Fin74d] A. M. Fink. Averaging. Lecture Notes in Mathematics, 377: 265–285, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070338/.

Fink:1974:EDK

[Fin74e] A. M. Fink. Exponential dichotomy and kinematic similarity. Lecture Notes in Mathematics, 377:121–141, 1974. CODEN LN-MAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070331/.

Fink:1974:FOE

[Fin74f] A. M. Fink. First order equations. Lecture Notes in Mathematics, 377:217–233, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070336/.

Fink:1974:FPM

[Fin74g] A. M. Fink. Fixed point methods. Lecture Notes in Mathematics, 377:142–153, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070332/.

Fink:1974:FST

[Fin74h] A. M. Fink. The Fourier series theory. Lecture Notes in Mathematics, 377:29–55, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070327/.

Fink:1974:LAP

[Fin74i] A. M. Fink. Linear almost periodic equations. Lecture Notes in Mathematics, 377:97–120, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070330/.

Fink:1974:LCC

[Fin74j] A. M. Fink. Linear constant coefficient equations. Lecture Notes in Mathematics, 377:77–96, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070329/.

Fink:1974:ME

[Fin74k] A. M. Fink. Modules and exponents. Lecture Notes in Mathematics, 377:56–76, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070328/.

Fink:1974:SOE

[Fin74l] A. M. Fink. Second order equations. Lecture Notes in Mathematics, 377:234–264, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070337/.

Fink:1974:SSa

[Fin74m] A. M. Fink. Separated solutions. Lecture Notes in Mathematics, 377:170–178, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070334/.

Fink:1974:SSb

[Fin74n] A. M. Fink. Stable solutions. Lecture Notes in Mathematics, 377: 179–216, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070335/.

Fink:1974:UAP

[Fin74o] A. M. Fink. Uniformly almost periodic families. Lecture Notes in Mathematics, 377:16–28, 1974. CODEN LNMAA2. ISBN 3-540-06729-9 (print), 3-540-38307-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070326/.

Fitting:1972:CSF

[Fit72] Melvin Fitting. An ϵ-calculus system for first-order S4. Lecture Notes in Mathematics, 255:103–110, 1972. CODEN LN-MAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059540/.

Flaschel:1972:RMG

[FK72a] P. Flaschel and W. Klingenberg. Die Riemannsche Mannigfaltigkeit H_{1}(I, M). (German) []. Lecture Notes in Mathematics, 282:71–111, 1972. CODEN LNMAA2. ISBN 3-540-05968-7 (print), 3-540-37961-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0080919/.

Flaschel:1972:PGK

[FK72b] P. Flaschel and W. Klingenberg. Periodische Geodätische auf Kompakten Riemannschen Mannigfaltigkeiten. (German) []. Lecture Notes in Mathematics, 282:112–199, 1972. CODEN LN-MAA2. ISBN 3-540-05968-7 (print), 3-540-37961-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0080920/.

Flaschel:1972:RHP

[FK72c] P. Flaschel and W. Klingenberg. Riemannsche Hilbertmannigfaltigkeiten. Periodische Geodätische: Mit einem Anhang von H. Karcher. (German) [Riemannian Hilbert manifolds. Periodic geodesic: With an appendix by H. Karcher], volume 282 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05968-7 (print), 3-540-37961-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 211 pp. LCCN QA3 .L28 no. 282. URL http:/

/link.springer.com/book/10.1007/BFb0080917;   http://
www.springerlink.com/content/978-3-540-37961-4.

Flaschel:1972:RGH

[FK72d] P. Flaschel and W. Klingenberg. Zur Riemannschen Geometrie auf Hilbertmannigfaltigkeiten. (German) []. Lecture Notes in Mathematics, 282:1–70, 1972. CODEN LNMAA2. ISBN 3-540-05968-7 (print), 3-540-37961-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0080918/.

Fuchs:1970:LKA

[FL70] A. Fuchs and G. Letta. L’inégalité de Kullback. Application à la théorie de l’estimation. (French) []. Lecture Notes in Mathematics, 124:108–131, 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059338/.

Fleischman:1970:SVM

[Fle70] W. M. Fleischman, editor. Set-Valued Mappings, Selections and Topological Properties of 2^{x}: Proceedings of the conference held at the State University of New York at Buffalo May 8–10, 1969, volume 171 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 171. URL http:/

/link.springer.com/book/10.1007/BFb0069710;   http://
www.springerlink.com/content/978-3-540-36419-1.

Frei:1971:LCC

[FM71] Armin Frei and John L. MacDonald. Limit-colimit commutation in abelian categories. Lecture Notes in Mathematics, 195: 199–204, 1971. CODEN LNMAA2. ISBN 3-540-05442-1 (print), 3-540-36548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072311/.

Fucik:1973:CLS

[FNSS73a] Svatopluk Fučík, Jindřich Nečas, Jiří Souček, and Vladimír Souček. The converse of Ljusternik–Schnirelmann theory. Lecture Notes in Mathematics, 346:152–166, 1973. CODEN LN-MAA2. ISBN 3-540-06484-2 (print), 3-540-37804-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059366/.

Fucik:1973:F

[FNSS73b] Svatopluk Fučík, Jindřich Nečas, Jiří Souček, and Vladimír Souček. Foreword. Lecture Notes in Mathematics, 346:2–7, 1973. CODEN LNMAA2. ISBN 3-540-06484-2 (print), 3-540-37804-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059361/.

Fucik:1973:FAN

[FNSS73c] Svatopluk Fučík, Jindřich Nečas, Jiří Souček, and Vladimír Souček. Fredholm alternative for nonlinear operators. Lecture Notes in Mathematics, 346:53–76, 1973. CODEN LN-MAA2. ISBN 3-540-06484-2 (print), 3-540-37804-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059363/.

Fucik:1973:LST

[FNSS73d] Svatopluk Fučík, Jindřich Nečas, Jiří Souček, and Vladimír Souček. Ljusternik–Schnirelmann theory. Lecture Notes in Mathematics, 346:77–115, 1973. CODEN LNMAA2. ISBN 3-540-06484-2 (print), 3-540-37804-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059364/.

Fucik:1973:MST

[FNSS73e] Svatopluk Fučík, Jindřich Nečas, Jiří Souček, and Vladimír Souček. Morse–Sard theorem. Lecture Notes in Mathematics, 346:116–151, 1973. CODEN LNMAA2. ISBN 3-540-06484-2 (print), 3-540-37804-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059365/.

Fucik:1973:P

[FNSS73f] Svatopluk Fučík, Jindřich Nečas, Jiří Souček, and Vladimír Souček. Preliminaries. Lecture Notes in Mathematics, 346:8– 52, 1973. CODEN LNMAA2. ISBN 3-540-06484-2 (print), 3-540-37804-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059362/.

Fucik:1973:SAN

[FNSS73g] Svatopluk Fučík, Jindřich Nečas, Jiří Souček, and Vladimír Souček. Spectral Analysis of Nonlinear Operators, volume 346 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06484-2 (print), 3-540-37804-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 166 (est.) pp. LCCN QA3 .L28 no. 346. URL http:/

/link.springer.com/book/10.1007/BFb0059360;   http://
www.springerlink.com/content/978-3-540-37804-4.

Forst:1974:CNS

[For74] Gunnar Forst. A characterization of non-symmetric translation invariant Dirichlet forms. Lecture Notes in Mathematics, 404: 113–125, 1974. CODEN LNMAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060611/.

Fournier:1971:RCM

[Fou71] J. C. Fournier. Représentation sur un Corps des matroïdes d’ordre ≤ 8. (French) []. Lecture Notes in Mathematics, 211: 50–61, 1971. CODEN LNMAA2. ISBN 3-540-05545-2 (print), 3-540-36877-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061074/.

Fournier:1974:CPQ

[Fou74] J. C. Fournier. Une condition pour qu’un hypergraphe, ou son complémentaire, soit fortement isomorphe à un hypergraphe complet. (French) []. Lecture Notes in Mathematics, 411:95– 98, 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066183/.

Fox:1974:STC

[Fox74] R. H. Fox. Shape theory and covering spaces. Lecture Notes in Mathematics, 375:71–90, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064013/.

Frazho:1974:NIA

[FPC74] D. B. Frazho, W. F. Powers, and R. P. Canale. Numerical integration aspects of a nutrient utilization ecological problem. Lecture Notes in Mathematics, 362:167–182, 1974. CODEN LN-MAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066591/.

Fraisse:1970:ADT

[Fra70a] Roland Fraïssé. Aspects du théorème de complétude selon Herbrand. (French) [Aspects of the completeness theorem according to Herbrand]. Lecture Notes in Mathematics, 125:73–86, 1970. CODEN LNMAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060625/.

Fraser:1970:SCS

[Fra70b] R. B. Fraser. Sequences of contractive set-valued maps. Lecture Notes in Mathematics, 171:24–26, 1970. CODEN LN-MAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069715/.

Frank:1974:CFE

[Fra74] Evelyn Frank. Continued fraction expansions for the ratios of hypergeometric functions evaluated on computers. Lecture Notes in Mathematics, 419:110–119, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064717/.

Freyd:1970:HC

[Fre70] Peter Freyd. Homotopy is not concrete. Lecture Notes in Mathematics, 168:25–34, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058516/.

Frederickson:1971:DSS

[Fre71a] Paul O. Frederickson. Dirichlet series solutions for certain functional differential equations. Lecture Notes in Mathematics, 243: 249–254, 1971. CODEN LNMAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058733/.

Frederickson:1971:PIF

[Fre71b] Paul O. Frederickson. On the Poincaré index for flows on 2manifolds. Lecture Notes in Mathematics, 235:65–68, 1971. CO-DEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059197/.

Freeman:1971:DVT

[Fre71c] Michael Freeman. A differential version of a theorem of Mergelyan. Lecture Notes in Mathematics, 184:37–49, 1971. CO-DEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059277/.

Frehse:1973:MSM

[Fre73] Jens Frehse. Morrey space methods in the theory of elliptic difference equations. Lecture Notes in Mathematics, 333:96–117, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060690/.

French:1974:CSL

[Fre74a] James Austin French. Coincidence of small and large inductive dimension. Lecture Notes in Mathematics, 378:132–139, 1974. CODEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068466/.

Freud:1974:PAR

[Fre74b] Géza Freud. On polynomial approximation with respect to general weights. Lecture Notes in Mathematics, 399:149–179, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063573/.

Friedman:1972:SAB

[Fri72] Avner Friedman. Stability and angular behavior of solutions of stochastic differential equations. Lecture Notes in Mathematics, 294:14–20, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064931/.

Friedman:1973:CMS

[Fri73a] Harvey Friedman. Countable models of set theories. Lecture Notes in Mathematics, 337:539–573, 1973. CODEN LN-MAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066789/.

Friedman:1973:E

[Fri73b] Harvey Friedman. Errata. Lecture Notes in Mathematics, 337: 573, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066790.

Friedman:1973:SAK

[Fri73c] Harvey Friedman. Some applications of Kleene’s methods for intuitionistic systems. Lecture Notes in Mathematics, 337:113– 170, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066773/.

Friedman:1974:FSD

[Fri74] Avner Friedman. Fundamental solutions for degenerate parabolic equations. Lecture Notes in Mathematics, 415:144–148, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065522/.

Frolik:1974:BRC

[Fro74] Zdeněk Frolík. Basic refinements of the category of uniform spaces. Lecture Notes in Mathematics, 378:140–158, 1974. CO-DEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068467/.

Foata:1970:FGP

[FS70a] Dominique Foata and Marcel-P. Schützenberger. Fonctions génératrices des polynômes eulériens. (French) []. Lecture Notes in Mathematics, 138:66–78, 1970. CODEN LNMAA2. ISBN 3-540-04927-4 (print), 3-540-36294-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060804/.

Foata:1970:IHN

[FS70b] Dominique Foata and Marcel-P. Schützenberger. Introduction et historique des nombres d’Euler. (French) []. Lecture Notes in Mathematics, 138:1–7, 1970. CODEN LNMAA2. ISBN 3-540-04927-4 (print), 3-540-36294-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060800/.

Foata:1970:FEF

[FS70c] Dominique Foata and Marcel-P. Schützenberger. La formule exponentielle. (French) []. Lecture Notes in Mathematics, 138:47– 65, 1970. CODEN LNMAA2. ISBN 3-540-04927-4 (print), 3-540-36294-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060803/.

Foata:1970:PEF

[FS70d] Dominique Foata and Marcel-P. Schützenberger. Les polynômes eulériens. (French) []. Lecture Notes in Mathematics, 138:27– 46, 1970. CODEN LNMAA2. ISBN 3-540-04927-4 (print), 3-540-36294-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060802/.

Foata:1970:SAF

[FS70e] Dominique Foata and Marcel-P. Schützenberger. Les sommes alternées A_{n}(−1) et B_{n}(−1). (French) []. Lecture Notes in Mathematics, 138:79–91, 1970. CODEN LNMAA2. ISBN 3-540-04927-4 (print), 3-540-36294-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060805/.

Foata:1970:PGS

[FS70f] Dominique Foata and Marcel-P. Schützenberger. Propriétés générales des systèmes d’excédances et de montées. (French) []. Lecture Notes in Mathematics, 138:8–26, 1970. CODEN LN-MAA2. ISBN 3-540-04927-4 (print), 3-540-36294-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060801/.

Foata:1970:TGP

[FS70g] Dominique Foata and Marcel-P. Schützenberger. Théorie Géométrique des Polynômes Eulériens. (French) [Geometric theory of Eulerian polynomials], volume 138 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CO-DEN LNMAA2. ISBN 3-540-04927-4 (print), 3-540-36294-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 94 pp. LCCN QA3 .L28 no. 138. URL http:/

/link.springer.com/book/10.1007/BFb0060799;   http://
www.springerlink.com/content/978-3-540-36294-4.

Fuchs:1972:CPM

[Fuc72] L. Fuchs. The cancellation property for modules. Lecture Notes in Mathematics, 246:191–212, 1972. CODEN LNMAA2. ISBN 3-540-05760-9 (print), 3-540-37181-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059566/.

Fuchs:1974:FRF

[Fuc74] W. H. J. Fuchs. A formula relevant to functions meromorphic in an angle and its applications. Lecture Notes in Mathematics, 419:120–124, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064718/.

Fuglede:1972:A

[Fug72a] Bent Fuglede. Applications. Lecture Notes in Mathematics, 289: 146–185, 1972. CODEN LNMAA2. ISBN 3-540-06005-7 (print), 3-540-37986-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068452/.

Fuglede:1972:CAP

[Fug72b] Bent Fuglede. Capacity in axiomatic potential theory. Lecture Notes in Mathematics, 289:32–66, 1972. CODEN LN-MAA2. ISBN 3-540-06005-7 (print), 3-540-37986-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068450/.

Fuglede:1972:FHFa

[Fug72c] Bent Fuglede. Finely harmonic and finely hyperharmonic functions. Lecture Notes in Mathematics, 289:67–145, 1972. CODEN LNMAA2. ISBN 3-540-06005-7 (print), 3-540-37986-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068451/.

Fuglede:1972:FHFb

[Fug72d] Bent Fuglede. Finely Harmonic Functions, volume 289 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06005-7 (print), 3-540-37986-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 188 pp. LCCN QA3 .L28 no. 289. URL http:/

/link.springer.com/book/10.1007/BFb0068447;   http://
www.springerlink.com/content/978-3-540-37986-7.

Fuglede:1972:I

[Fug72e] Bent Fuglede. Introduction. Lecture Notes in Mathematics, 289: 1–13, 1972. CODEN LNMAA2. ISBN 3-540-06005-7 (print), 3-540-37986-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068448/.

Fuglede:1972:P

[Fug72f] Bent Fuglede. Preliminaries. Lecture Notes in Mathematics, 289: 14–31, 1972. CODEN LNMAA2. ISBN 3-540-06005-7 (print), 3-540-37986-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068449/.

Fukushima:1973:GMP

[Fuk73] Masatoshi Fukushima. On the generation of Markov processes by symmetric forms. Lecture Notes in Mathematics, 330:46– 79, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061480/.

Furstenberg:1973:UEH

[Fur73] Harry Furstenberg. The unique ergodigity of the horocycle flow. Lecture Notes in Mathematics, 318:95–115, 1973. CODEN LN-MAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061726/.

Gabriel:1970:CPQ

[Gab70a] P. Gabriel. Construction de préschémas quotient. (French) []. Lecture Notes in Mathematics, 151:251–286, 1970. CODEN LN-MAA2. ISBN 3-540-05179-1 (print), 3-540-36336-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058998/.

Gabriel:1970:EISa

[Gab70b] P. Gabriel. Étude infinitesimale des schémas en groupes. (French) []. Lecture Notes in Mathematics, 151:411–475, 1970. CODEN LNMAA2. ISBN 3-540-05179-1 (print), 3-540-36336-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059001/.

Gabriel:1970:EISb

[Gab70c] P. Gabriel. Étude infinitesimale des schémas en groups. (French) []. Lecture Notes in Mathematics, 151:476–562, 1970. CODEN LNMAA2. ISBN 3-540-05179-1 (print), 3-540-36336-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059002/.

Gabriel:1970:GGA

[Gab70d] P. Gabriel. Généralités sur les groupes algébriques. (French) []. Lecture Notes in Mathematics, 151:287–317, 1970. CODEN LN-MAA2. ISBN 3-540-05179-1 (print), 3-540-36336-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058999/.

Gabriel:1971:RAL

[Gab71] Pierre Gabriel. Représentations des algèbres de Lie résolubles. (French) []. Lecture Notes in Mathematics, 179:1–22, 1971. CO-DEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058802/.

Gabbay:1972:CIT

[Gab72] Dov M. Gabbay. Craig’s interpolation theorem for modal logics. Lecture Notes in Mathematics, 255:111–127, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059541/.

Gaifman:1972:NMS

[Gai72] Haim Gaifman. A note on models and submodels of arithmetic. Lecture Notes in Mathematics, 255:128–144, 1972. CODEN LN-MAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059542/.

Gaier:1974:DCM

[Gai74] Dieter Gaier. Determination of conformal modules of ring domains and quadrilaterals. Lecture Notes in Mathematics, 399: 180–188, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0063574/.

Galambos:1972:DAM

[Gal72a] János Galambos. Distribution of additive and multiplicative functions. Lecture Notes in Mathematics, 251:127–139, 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058790/.

Galambos:1972:RGT

[Gal72b] János Galambos. The role of graph theory in some sieve arguments of probability theory. Lecture Notes in Mathematics, 303: 81–87, 1972. CODEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067360/.

Gamelin:1971:PAT

[Gam71] T. W. Gamelin. Polynomial approximation on thin sets. Lecture Notes in Mathematics, 184:50–78, 1971. CODEN LN-MAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059278/.

Ganea:1971:SPN

[Gan71a] Tudor Ganea. Some problems on numerical homotopy invariants. Lecture Notes in Mathematics, 249:23–30, 1971. CODEN LNMAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060892/.

Ganelius:1971:CTT

[Gan71b] Tord H. Ganelius. Complex Tauberian theorems. Lecture Notes in Mathematics, 232:57–66, 1971. CODEN LNMAA2. ISBN 3-540-05657-2 (print), 3-540-36979-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059955/.

Ganelius:1971:CEN

[Gan71c] Tord H. Ganelius. Convolution equations without non-trivial solutions. Lecture Notes in Mathematics, 232:53–56, 1971. CODEN LNMAA2. ISBN 3-540-05657-2 (print), 3-540-36979-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059954/.

Ganelius:1971:DAT

[Gan71d] Tord H. Ganelius. Different approaches to Tauberian theorems. Lecture Notes in Mathematics, 232:1–17, 1971. CODEN LN-MAA2. ISBN 3-540-05657-2 (print), 3-540-36979-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059949/.

Ganelius:1971:MR

[Gan71e] Tord H. Ganelius. Miscellaneous remarks. Lecture Notes in Mathematics, 232:67–71, 1971. CODEN LNMAA2. ISBN 3-540-05657-2 (print), 3-540-36979-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059956/.

Ganelius:1971:RDR

[Gan71f] Tord H. Ganelius. Rapidly decreasing remainders. Lecture Notes in Mathematics, 232:46–52, 1971. CODEN LNMAA2. ISBN 3-540-05657-2 (print), 3-540-36979-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059953/.

Ganelius:1971:RTW

[Gan71g] Tord H. Ganelius. Remainder theorems in Wiener’s form. Lecture Notes in Mathematics, 232:22–33, 1971. CODEN LN-MAA2. ISBN 3-540-05657-2 (print), 3-540-36979-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059951/.

Ganelius:1971:SPT

[Gan71h] Tord H. Ganelius. Some precise theorems in several dimensions. Lecture Notes in Mathematics, 232:34–45, 1971. CODEN LN-MAA2. ISBN 3-540-05657-2 (print), 3-540-36979-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059952/.

Ganelius:1971:TRT

[Gan71i] Tord H. Ganelius. Tauberian Remainder Theorems, volume 232 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05657-2 (print), 3-540-36979-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 75 pp. LCCN QA3 .L28 no. 232. URL http:/

/link.springer.com/book/10.1007/BFb0059948;   http://
www.springerlink.com/content/978-3-540-36979-0.

Ganelius:1971:TGR

[Gan71j] Tord H. Ganelius. A typical general remainder theorem. Lecture Notes in Mathematics, 232:18–21, 1971. CODEN LN-MAA2. ISBN 3-540-05657-2 (print), 3-540-36979-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059950/.

Garnett:1972:AC

[Gar72a] John Garnett. Analytic capacity. Lecture Notes in Mathematics, 297:6–36, 1972. CODEN LNMAA2. ISBN 3-540-06073-1 (print), 3-540-38003-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060914/.

Garnett:1972:ACM

[Gar72b] John Garnett. Analytic Capacity and Measure, volume 297 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06073-1 (print), 3-540-38003-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 138 pp. LCCN QA3 .L28 no. 297. URL http:/

/link.springer.com/book/10.1007/BFb0060912;   http://
www.springerlink.com/content/978-3-540-38003-0.

Garnett:1972:AA

[Gar72c] John Garnett. Applications to approximation. Lecture Notes in Mathematics, 297:107–130, 1972. CODEN LNMAA2. ISBN 3-540-06073-1 (print), 3-540-38003-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060918/.

Garnett:1972:CT

[Gar72d] John Garnett. The Cauchy transform. Lecture Notes in Mathematics, 297:37–57, 1972. CODEN LNMAA2. ISBN 3-540-06073-1 (print), 3-540-38003-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060915/.

Garnett:1972:HM

[Gar72e] John Garnett. Hausdorff measure. Lecture Notes in Mathematics, 297:58–82, 1972. CODEN LNMAA2. ISBN 3-540-06073-1 (print), 3-540-38003-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060916/.

Garnett:1972:I

[Gar72f] John Garnett. Introduction. Lecture Notes in Mathematics, 297: 1–5, 1972. CODEN LNMAA2. ISBN 3-540-06073-1 (print), 3-540-38003-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060913/.

Garnett:1972:SE

[Gar72g] John Garnett. Some examples. Lecture Notes in Mathematics, 297:83–106, 1972. CODEN LNMAA2. ISBN 3-540-06073-1 (print), 3-540-38003-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060917/.

Garling:1973:LBM

[Gar73] D. J. H. Garling. Lattice bounding mappings. Lecture Notes in Mathematics, 331:218–221, 1973. CODEN LNMAA2. ISBN 3-540-06367-6 (print), 3-540-46977-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068233/.

Gardiner:1974:AFE

[Gar74a] Frederick P. Gardiner. Automorphic forms and Eichler cohomology. Lecture Notes in Mathematics, 400:24–47, 1974. CODEN LNMAA2. ISBN 3-540-06840-6 (print), 3-540-37776-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065674/.

Garnir:1974:SAF

[Gar74b] H. G. Garnir. Solovay’s axiom and functional analysis. Lecture Notes in Mathematics, 399:189–204, 1974. CODEN LN-MAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063575/.

Gatesoupe:1973:SAF

[Gat73a] Michel Gatesoupe. Sur les sous-algèbres fermées d’algèbres de groupes abeliens compacts qui sont des algèbres de Beurling. (French) []. Lecture Notes in Mathematics, 336:28–39, 1973. CO-DEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065785/.

Gatterdam:1973:HTC

[Gat73b] R. W. Gatterdam. The Higman theorem for E^{n}(A) computable groups. Lecture Notes in Mathematics, 319:71–74, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058930/.

Gundy:1970:MIA

[GC70] R. F. Gundy and F. Chersi. Martingales et intégrabilité de X log^{+} X d’après. (French) []. Lecture Notes in Mathematics, 124: 37–46, 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059331/.

Guivarch:1974:PDF

[GC74] Y. Guivarc’h and J.-P. Conze. Propriété de droite fixe et fonctions harmoniques positives. (French) []. Lecture Notes in Mathematics, 404:126–132, 1974. CODEN LNMAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060612/.

Gebuhrer:1974:CPM

[Geb74] M. O. Gebuhrer. Une classe de processus de Markov en mécanique relativiste. Laplaciens généralisés sur les espaces symétriques de type non compact. (French) []. Lecture Notes in Mathematics, 381:80–133, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057258/.

Gehring:1974:ICH

[Geh74] F. W. Gehring. Inequalities for condensers, hyperbolic capacity, and extremal lengths. Lecture Notes in Mathematics, 419: 133–136, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064720/.

Gekeler:1972:RLS

[Gek72] Eckart Gekeler. Zur Randwerttechnik bei der Lösung schwach nichtlinearer parabolischer Differentialgleichungen. (German) []. Lecture Notes in Mathematics, 267:71–80, 1972. CODEN LN-MAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061614/.

Gekeler:1973:BPL

[Gek73] Eckart Gekeler. Zur Berechnung Periodischer Lösungen bei Parabolischen Randwertproblemen. (German) []. Lecture Notes in Mathematics, 333:118–127, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060691/.

Gelbaum:1974:UAO

[Gel74] Bernard R. Gelbaum. Q-uniform algebras and operator theory. Lecture Notes in Mathematics, 399:205–208, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063576/.

Gentzsch:1974:NBQ

[Gen74] Wolfgang Gentzsch. Zur numerischen Behandlung quasilinearer elliptischer Differentialgleichungen 4. Ordnung in 2 Dimensionen. (German) []. Lecture Notes in Mathematics, 395:169–175, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060670/.

Geoghegan:1974:ICD

[Geo74] Ross Geoghegan. On integral currents and the Dold–Thom construction. Lecture Notes in Mathematics, 428:241–281, 1974. CO-DEN LNMAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070542/.

Gersten:1973:HTR

[Ger73a] S. M. Gersten. Higher K-theory of rings. Lecture Notes in Mathematics, 341:3–42, 1973. CODEN LNMAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067049/.

Gersten:1973:PAH

[Ger73b] S. M. Gersten. Problems about higher K-functors. Lecture Notes in Mathematics, 341:43–56, 1973. CODEN LNMAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067050/.

Gersten:1973:SES

[Ger73c] S. M. Gersten. Some exact sequences in the higher K-theory of rings. Lecture Notes in Mathematics, 341:211–243, 1973. CODEN LNMAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067059/.

Gerard:1974:MIP

[Gér74] Raymond Gérard. Mesure de l’irregularité en un point singulier d’un système différentiel. (French) []. Lecture Notes in Mathematics, 409:418–429, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068113/.

Geyer:1974:IBF

[Gey74] W. D. Geyer. Invarianten binärer formen. (German) []. Lecture Notes in Mathematics, 412:36–69, 1974. CODEN LN-MAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066153/.

Gioia:1972:TAF

[GG72] Anthony A. Gioia and Donald L. Goldsmith, editors. The Theory of Arithmetic Functions: Proceedings of the Conference at Western Michigan University, April 29–May 1, 1971, volume 251 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 251. URL http:/

/link.springer.com/book/10.1007/BFb0058781;   http://
www.springerlink.com/content/978-3-540-37098-7.

Geller:1971:CD

[GH71] Dennis Geller and Frank Harary. Connectivity in digraphs. Lecture Notes in Mathematics, 186:105–115, 1971. CODEN LN-MAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059429/.

Grimm:1973:DIB

[GH73] L. J. Grimm and L. M. Hall. Differential inequalities and boundary problems for functional-differential equations. Lecture Notes in Mathematics, 312:41–53, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060044/.

Gilbert:1974:DPQ

[GH74a] Robert P. Gilbert and George C. Hsiao. On Dirichlet’s problem for quasi-linear elliptic equations. Lecture Notes in Mathematics, 430:184–236, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066270/.

Goodman:1974:HCP

[GH74b] S. Goodman and S. Hedetniemi. On the Hamiltonian completion problem. Lecture Notes in Mathematics, 406:262–272, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066448/.

Guthrie:1974:SSQ

[GH74c] J. A. Guthrie and Michael Henry. Sequential spaces and quasiperfect mappings. Lecture Notes in Mathematics, 378:159–163, 1974. CODEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068468/.

Giambalvo:1972:CLB

[Gia72] V. Giambalvo. Cobordism of line bundles with restricted characteristic class. Lecture Notes in Mathematics, 298:170–171, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070040.

Gibson:1972:PFG

[Gib72] P. M. Gibson. The Pfaffian and 1-factors of graphs II. Lecture Notes in Mathematics, 303:89–98, 1972. CODEN LN-MAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067361/.

Gilsinn:1970:DSI

[Gil70] David E. Gilsinn. Domains of stability for integral manifolds. Lecture Notes in Mathematics, 144:70–75, 1970. CODEN LN-MAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059922/.

Gilmer:1973:PLC

[Gil73] Robert Gilmer. Prüfer-like conditions on the set of overrings of an integral domain. Lecture Notes in Mathematics, 311:90– 102, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068922/.

Gilbert:1974:BVP

[Gil74a] Robert P. Gilbert. Boundary value problems associated with e[u] = 0. Lecture Notes in Mathematics, 365:34–61, 1974. CO-DEN LNMAA2. ISBN 3-540-06690-X (print), 3-540-37953-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064478/.

Gilbert:1974:CME

[Gil74b] Robert P. Gilbert. Constructive Methods for Elliptic Equations, volume 365 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06690-X (print), 3-540-37953-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 397 pp. LCCN QA3 .L28 no. 365. URL http:/

/link.springer.com/book/10.1007/BFb0064475;   http://
www.springerlink.com/content/978-3-540-37953-9.

Gilbert:1974:EED

[Gil74c] Robert P. Gilbert. Elliptic equations in dimension n ≥3. Lecture Notes in Mathematics, 365:116–176, 1974. CODEN LN-MAA2. ISBN 3-540-06690-X (print), 3-540-37953-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064481/.

Gilbert:1974:EPD

[Gil74d] Robert P. Gilbert. Elliptic partial differential equations in two variables with analytic coefficients. Lecture Notes in Mathematics, 365:14–33, 1974. CODEN LNMAA2. ISBN 3-540-06690-X (print), 3-540-37953-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064477/.

Gilbert:1974:ESFb

[Gil74e] Robert P. Gilbert. Elliptic systems of first order differential equations: Further results and constructive methods. Lecture Notes in Mathematics, 365:363–390, 1974. CODEN LNMAA2. ISBN 3-540-06690-X (print), 3-540-37953-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064486/.

Gilbert:1974:ESFa

[Gil74f] Robert P. Gilbert. Elliptic systems of first order differential equations: General theory. Lecture Notes in Mathematics, 365:302– 362, 1974. CODEN LNMAA2. ISBN 3-540-06690-X (print), 3-540-37953-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064485/.

Gilbert:1974:FRS

[Gil74g] Robert P. Gilbert. Further representations for solutions of higher order elliptic differential equations with analytic coefficients. Lecture Notes in Mathematics, 365:88–115, 1974. CODEN LN-MAA2. ISBN 3-540-06690-X (print), 3-540-37953-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064480/.

Gilbert:1974:HOH

[Gil74h] Robert P. Gilbert. Higher order and higher dimensional elliptic problems. Lecture Notes in Mathematics, 365:210–257, 1974. CODEN LNMAA2. ISBN 3-540-06690-X (print), 3-540-37953-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064483/.

Gilbert:1974:HOL

[Gil74i] Robert P. Gilbert. Higher order, linear elliptic equations of two independent variables. Lecture Notes in Mathematics, 365:62– 87, 1974. CODEN LNMAA2. ISBN 3-540-06690-X (print), 3-540-37953-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064479/.

Gilbert:1974:IVP

[Gil74j] Robert P. Gilbert. Initial value problems for semilinear elliptic equations. Lecture Notes in Mathematics, 365:258–301, 1974. CODEN LNMAA2. ISBN 3-540-06690-X (print), 3-540-37953-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064484/.

Gilbert:1974:I

[Gil74k] Robert P. Gilbert. Introduction. Lecture Notes in Mathematics, 365:1–13, 1974. CODEN LNMAA2. ISBN 3-540-06690-X (print), 3-540-37953-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064476/.

Gilbert:1974:SPS

[Gil74l] Robert P. Gilbert. Singularity properties of solutions to partial differential equations. Lecture Notes in Mathematics, 365: 177–209, 1974. CODEN LNMAA2. ISBN 3-540-06690-X (print), 3-540-37953-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064482/.

Giraud:1972:CT

[Gir72] Jean Giraud. Classifying topos. Lecture Notes in Mathematics, 274:43–56, 1972. CODEN LNMAA2. ISBN 3-540-05920-2 (print), 3-540-37609-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073964/.

Girard:1973:ORI

[Gir73a] J. Y. Girard. Ouelques résultats sur les interprétations fonctionnelles. (French) []. Lecture Notes in Mathematics, 337:232– 252, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066776/.

Giri:1973:DDF

[Gir73b] N. Giri. On discriminant decision function in complex Gaussian distributions. Lecture Notes in Mathematics, 296:139–148, 1973. CODEN LNMAA2. ISBN 3-540-06211-4 (print), 3-540-38485-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059822/.

Giroux:1974:PPZ

[Gir74] André Giroux. On polynomials with a prescribed zero. Lecture Notes in Mathematics, 399:209–211, 1974. CODEN LN-MAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063577/.

Gitler:1972:OCO

[Git72] Samuel Gitler. Operaciones cohomologicas de orden supérior. (Spanish) [Higher order cohomological operations]. Lecture Notes in Mathematics, 279:95–133, 1972. CODEN LNMAA2. ISBN 3-540-05944-X (print), 3-540-37616-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058510/.

Godement:1972:GT

[GJ72a] Roger Godement and Hervé Jacquet. Global theory. Lecture Notes in Mathematics, 260:136–184, 1972. CODEN LN-MAA2. ISBN 3-540-05797-8 (print), 3-540-37436-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070265/.

Godement:1972:LT

[GJ72b] Roger Godement and Hervé Jacquet. Local theory. Lecture Notes in Mathematics, 260:1–135, 1972. CODEN LNMAA2. ISBN 3-540-05797-8 (print), 3-540-37436-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070264/.

Godement:1972:ZFS

[GJ72c] Roger Godement and Hervé Jacquet. Zeta Functions of Simple Algebras, volume 260 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05797-8 (print), 3-540-37436-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 188 pp. LCCN QA3 .L28 no. 260. URL http:/

/link.springer.com/book/10.1007/BFb0070263;   http://
www.springerlink.com/content/978-3-540-37436-7.

Grodzovsky:1970:WLI

[GK70] G. L. Grodzovsky and B. N. Kiforenko. Weight limitation influence on optimum motion parameters of a body of variable mass. Lecture Notes in Mathematics, 112:248–253, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060208/.

Gross:1972:FBT

[GK72] Kenneth I. Gross and Ray A. Kunze. Fourier Bessel transforms and holomorphic discrete series. Lecture Notes in Mathematics, 266:79–122, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059639/.

Gitler:1970:TSM

[GL70] S. Gitler and Kee Yuen Lam. The K theory of Stiefel manifolds. Lecture Notes in Mathematics, 168:35–66, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058517/.

Gulick:1972:CHA

[GL72] Denny Gulick and Ronald L. Lipsman, editors. Conference on Harmonic Analysis: College Park, Maryland, 1971, volume 266 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 266. URL http:/

/link.springer.com/book/10.1007/BFb0059631;   http://
www.springerlink.com/content/978-3-540-37479-4.

Gilbert:1974:NSS

[GL74] R. P. Gilbert and Peter Linz. The numerical solution of some elliptic boundary value problems by integral operator methods. Lecture Notes in Mathematics, 430:237–252, 1974. CODEN LN-MAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066271/.

Glaeser:1971:LFD

[Gla71a] G. Glaeser. L’interpolation des fonctions Différentiables de plusieurs variables. (French) []. Lecture Notes in Mathematics, 209:1–33, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068890/.

Glaeser:1971:TPD

[Gla71b] G. Glaeser. Sur le Théorème de Préparation Différentiable. (French) []. Lecture Notes in Mathematics, 192:121–132, 1971. CODEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066817/.

Glaser:1971:SHS

[Gla71c] Leslie C. Glaser. On suspensions of homology spheres. Lecture Notes in Mathematics, 197:8–16, 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068608/.

Glimm:1970:MPF

[Gli70] James Glimm. Mathematical problems in the foundations of quantum field theory. Lecture Notes in Mathematics, 140:58– 67, 1970. CODEN LNMAA2. ISBN 3-540-04929-0 (print), 3-540-36298-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100066/.

Glicksberg:1972:FST

[Gli72] I. Glicksberg. Fourier–Stieltjes transforms with an isolated value. Lecture Notes in Mathematics, 266:59–72, 1972. CODEN LN-MAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059637/.

Glowinski:1970:RNP

[Glo70] R. Glowinski. Resolution numérique d’un problème non classique de calcul des variations par réduction à une équation intégrale non linéaire. (French) []. Lecture Notes in Mathematics, 132: 108–129, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066678/.

Gloede:1973:FCU

[Glo73] K. Gloede. Filters closed under Mahlo’s and Gaifman’s operation. Lecture Notes in Mathematics, 337:495–530, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066787/.

Gitler:1971:UDC

[GM71a] Samuel Gitler and R. James Milgram. Unstable divisibility of the Chern character. Lecture Notes in Mathematics, 249:31–33, 1971. CODEN LNMAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060893/.

Grothendieck:1971:AFG

[GM71b] Alexander Grothendieck and Jacob P. Murre. An application: the fundamental group of the spectrum of a complete local ring, of dimension two, minus a closed set. Lecture Notes in Mathematics, 208:122–128, 1971. CODEN LNMAA2. ISBN 3-540-05499-5 (print), 3-540-36857-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069617/.

Grothendieck:1971:CTC

[GM71c] Alexander Grothendieck and Jacob P. Murre. Comparison of two 2-cohomology classes. Lecture Notes in Mathematics, 208: 82–96, 1971. CODEN LNMAA2. ISBN 3-540-05499-5 (print), 3-540-36857-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069614/.

Grothendieck:1971:DTR

[GM71d] Alexander Grothendieck and Jacob P. Murre. Descent of tamely ramified coverings. Lecture Notes in Mathematics, 208:110– 121, 1971. CODEN LNMAA2. ISBN 3-540-05499-5 (print), 3-540-36857-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069616/.

Grothendieck:1971:ESN

[GM71e] Alexander Grothendieck and Jacob P. Murre. Extension of some notions from the theory of schemes to the theory of formal schemes. Lecture Notes in Mathematics, 208:44–51, 1971. CO-DEN LNMAA2. ISBN 3-540-05499-5 (print), 3-540-36857-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069611/.

Grothendieck:1971:KC

[GM71f] Alexander Grothendieck and Jacob P. Murre. Kummer coverings. Lecture Notes in Mathematics, 208:1–29, 1971. CODEN LNMAA2. ISBN 3-540-05499-5 (print), 3-540-36857-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069609/.

Grothendieck:1971:TFGc

[GM71g] Alexander Grothendieck and Jacob P. Murre. The Tame Fundamental Group of a Formal Neighbourhood of a Divisor with Normal Crossings on a Scheme, volume 208 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05499-5 (print), 3-540-36857-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 133 pp. LCCN QA3 .L28 no. 208. URL http:/

/link.springer.com/book/10.1007/BFb0069608;   http://
www.springerlink.com/content/978-3-540-36857-1.

Grothendieck:1971:TFGa

[GM71h] Alexander Grothendieck and Jacob P. Murre. The tame fundamental group of a formal neighbourhood of an irreducible divisor. Lecture Notes in Mathematics, 208:69–81, 1971. CODEN LNMAA2. ISBN 3-540-05499-5 (print), 3-540-36857-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069613/.

Grothendieck:1971:TFGb

[GM71i] Alexander Grothendieck and Jacob P. Murre. The tame fundamental group of a formal neighbourhood of an irreducible divisor (continued). Lecture Notes in Mathematics, 208:97–109, 1971. CODEN LNMAA2. ISBN 3-540-05499-5 (print), 3-540-36857-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069615/.

Grothendieck:1971:TRCb

[GM71j] Alexander Grothendieck and Jacob P. Murre. Tamely ramified coverings of formal schemes. Lecture Notes in Mathematics, 208: 52–68, 1971. CODEN LNMAA2. ISBN 3-540-05499-5 (print), 3-540-36857-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069612/.

Grothendieck:1971:TRCa

[GM71k] Alexander Grothendieck and Jacob P. Murre. Tamely ramified coverings of schemes. Lecture Notes in Mathematics, 208:30– 43, 1971. CODEN LNMAA2. ISBN 3-540-05499-5 (print), 3-540-36857-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069610/.

Goodman:1972:FBC

[GM72] Nicolas D. Goodman and John Myhill. The formalization of Bishop’s constructive mathematics. Lecture Notes in Mathematics, 274:83–96, 1972. CODEN LNMAA2. ISBN 3-540-05920-2 (print), 3-540-37609-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073966/.

Glover:1974:MEL

[GM74] Henry H. Glover and Guido Mislin. Metastable embedding and 2-localization. Lecture Notes in Mathematics, 418:48–57, 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070641/.

Gobert:1974:PIS

[Gob74] J. Gobert. A priori inequalities for systems of partial differential equations. Lecture Notes in Mathematics, 399:212–219, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063578/.

Godbillon:1971:FL

[God71a] C. Godbillon. Formalisme lagrangien. Lecture Notes in Mathematics, 206:9–11, 1971. CODEN LNMAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070122/.

Godbillon:1971:TDA

[God71b] Claude Godbillon. Travaux de D. Anosov et S. Smale sur les difféomorphismes. (French) []. Lecture Notes in Mathematics, 179:23–35, 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058803/.

Godement:1971:FAP

[God71c] R. Godement. Formes automorphes et produits euleriens. (French) []. Lecture Notes in Mathematics, 179:37–53, 1971. CO-DEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058804/.

Godbillon:1972:PEH

[God72] Claude Godbillon. Problèmes d’existence et d’homotopie dans les feuilletages. (French) []. Lecture Notes in Mathematics, 244: 167–181, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058701/.

Godbillon:1974:CAL

[God74a] Claude Godbillon. Cohomologies d’algèbres de Lie de champs de vecteurs formels. (French) []. Lecture Notes in Mathematics, 383:69–87, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057302/.

Godbillon:1974:FLF

[God74b] Claude Godbillon. Feuilletages de Lie. (French) []. Lecture Notes in Mathematics, 392:10–13, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064104/.

Gollwitzer:1971:SSH

[Gol71a] H. E. Gollwitzer. Small solutions of a high frequency linear oscillator. Lecture Notes in Mathematics, 183:28–34, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060408/.

Golub:1971:SME

[Gol71b] G. H. Golub. Some modified eigenvalue problems. Lecture Notes in Mathematics, 228:56, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069447.

Golub:1971:DMS

[Gol71c] Gene Golub. Direct methods for solving elliptic difference equations. Lecture Notes in Mathematics, 193:1–19, 1971. CODEN LNMAA2. ISBN 3-540-05422-7 (print), 3-540-36538-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060341/.

Golub:1971:MMM

[Gol71d] Gene Golub. Matrix methods in mathematical programming. Lecture Notes in Mathematics, 193:21–39, 1971. CODEN LN-MAA2. ISBN 3-540-05422-7 (print), 3-540-36538-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060342/.

Goldie:1972:SNR

[Gol72] Alfred W. Goldie. The structure of Noetherian rings. Lecture Notes in Mathematics, 246:213–321, 1972. CODEN LN-MAA2. ISBN 3-540-05760-9 (print), 3-540-37181-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059567/.

Gollwitzer:1973:AIE

[Gol73] H. E. Gollwitzer. Admissibility and the integral equations of asymptotic theory. Lecture Notes in Mathematics, 312:23–40, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060043/.

Goldberg:1974:RRS

[Gol74a] Richard R. Goldberg. Recent results on Segal algebras. Lecture Notes in Mathematics, 399:220–229, 1974. CODEN LN-MAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063579/.

Goldschmidt:1974:FHC

[Gol74b] Hubert Goldschmidt. Le formalisme de Hamilton–Cartan en calcul des variations. (French) []. Lecture Notes in Mathematics, 392: 76–89, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064114/.

Goldsmith:1974:HBA

[Gol74c] Deborah Louise Goldsmith. Homotopy of braids — in answer to a question of E. Artin. Lecture Notes in Mathematics, 375: 91–96, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064014/.

Gomes:1970:GPA

[Gom70] A. Pereira Gomes. Groupes pro-analytiques et fonctions analytiques. (French) []. Lecture Notes in Mathematics, 116:119– 127, 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059249/.

Gonshor:1974:ECV

[Gon74] Harry Gonshor. Enlargements contain various kinds of completions. Lecture Notes in Mathematics, 369:60–70, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066000/.

Goodwyn:1971:TEB

[Goo71] L. Wayne Goodwyn. Topological entropy bounds measuretheoretic entropy. Lecture Notes in Mathematics, 235:69–84, 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059198/.

Goodman:1973:ATC

[Goo73a] Nicolas D. Goodman. The arithmetic theory of constructions. Lecture Notes in Mathematics, 337:274–298, 1973. CODEN LN-MAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066778/.

Goodwyn:1973:HSW

[Goo73b] L. Wayne Goodwyn. How should we define topological entropy? Lecture Notes in Mathematics, 318:116–119, 1973. CODEN LN-MAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061727/.

Gordon:1970:VFS

[Gor70] R. Gordon. Vector fields on spheres. Lecture Notes in Mathematics, 136:254–264, 1970. CODEN LNMAA2. ISBN 3-540-04925-8 (print), 3-540-36289-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059025/.

Gorenflo:1971:DSP

[Gor71] Rudolf Gorenflo. On difference schemes for parabolic differential equations with derivative boundary conditions. Lecture Notes in Mathematics, 228:57–69, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069448/.

Gorenflo:1973:GMF

[Gor73] Rudolf Gorenflo. Über S. Gerschgorins Methode der Fehlerabschatzung bei Differenzenverfahren. (German) []. Lecture Notes in Mathematics, 333:128–143, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060692/.

Gottschalk:1971:AA

[Got71] Walter Gottschalk. Ambits (abstract). Lecture Notes in Mathematics, 235:85, 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059199.

Gottschalk:1973:SGD

[Got73] Walter Gottschalk. Some general dynamical notions. Lecture Notes in Mathematics, 318:120–125, 1973. CODEN LN-MAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061728/.

Goulaouic:1971:TSO

[Gou71] Charles Goulaouic. Sur la théorie spectrale des opérateurs elliptiques. (French) []. Lecture Notes in Mathematics, 179:231– 244, 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058815/.

Goulaouic:1972:RSD

[Gou72a] C. Goulaouic. Regularity for some degenerate differential operators. Lecture Notes in Mathematics, 280:65–72, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066919/.

Gourlay:1972:HCD

[Gou72b] A. R. Gourlay. The hopscotch class of difference methods for partial differential equations. Lecture Notes in Mathematics, 267: 81–98, 1972. CODEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061615/.

Goulaouic:1974:DEB

[Gou74a] C. Goulaouic. Degenerate elliptic boundary value problems and applications. Lecture Notes in Mathematics, 419:137–142, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064721/.

Gourlay:1974:RDH

[Gou74b] A. R. Gourlay. Recent developments of the hopscotch idea. Lecture Notes in Mathematics, 363:44–54, 1974. CODEN LN-MAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069124/.

Gowrisankaran:1974:MLO

[Gow74] Kohur Gowrisankaran. Measurability of lattice operations in a cone. Lecture Notes in Mathematics, 399:230–234, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063580/.

Graham:1972:EGS

[GP72] R. L. Graham and H. O. Pollak. On embedding graphs in squashed cubes. Lecture Notes in Mathematics, 303:99–110, 1972. CODEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067362/.

Gallaher:1974:UGF

[GP74] L. J. Gallaher and I. E. Perlin. Use of Green’s functions in the numerical solution of two-point boundary value problems. Lecture Notes in Mathematics, 362:374–407, 1974. CODEN LN-MAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066602/.

Gewirtz:1971:AFP

[GQ71] A. Gewirtz and Louis V. Quintas. Antitone fixed point maps. Lecture Notes in Mathematics, 186:117–136, 1971. CODEN LN-MAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059430/.

Grauert:1970:VAK

[GR70] Hans Grauert and Oswald Riemenschneider. Verschwindungssätze Für Analytische Kohomologiegruppen Auf Komplexen Räumen. (German) []. Lecture Notes in Mathematics, 155: 97–109, 1970. CODEN LNMAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060317/.

Grothendieck:1971:REG

[GR71] A. (Alexandre) Grothendieck and M. Raynaud. Rev{êtements étales et groupe fondamental: SGA 1, Séminaire de géometrie algébrique du Bois-Marie, 1960–61, volume 224 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 0-354-00361-4, 0-387-05614-9, 3-540-05614-9 (print), 3-540-36910-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xxii + 447 pp. LCCN QA3 .L28 no. 224. URL http:/

/link.springer.com/book/10.1007/BFb0058656;   http://
www.springerlink.com/content/978-3-540-36910-3.

Grothendieck:1972:MNM

[GR72] A. Grothendieck and M. Raynaud. Modèles de Néron et monodromie. (French) []. Lecture Notes in Mathematics, 288:313– 523, 1972. CODEN LNMAA2. ISBN 3-540-05987-3 (print), 3-540-37984-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068694/.

Gray:1971:RMC

[Gra71a] J. W. Gray, editor. Reports of the Midwest Category Seminar V, volume 195 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05442-1 (print), 3-540-36548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 195. URL http:/

/link.springer.com/book/10.1007/BFb0072304;   http://
www.springerlink.com/content/978-3-540-36548-8.

Gray:1971:MMC

[Gra71b] John W. Gray. The meeting of the Midwest Category Seminar in Zürich August 24–30, 1970. Lecture Notes in Mathematics, 195: 248–255, 1971. CODEN LNMAA2. ISBN 3-540-05442-1 (print), 3-540-36548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072315/.

Graham:1973:CSF

[Gra73] Jimmie Graham. Continuous symbols on fields of formal power series. Lecture Notes in Mathematics, 342:474–486, 1973. CO-DEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073736/.

Gramain:1974:GDE

[Gra74a] André Gramain. Groupe des difféomorphismes et espace de Teichmüller d’une surface. (French) []. Lecture Notes in Mathematics, 383:157–170, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057307/.

Grant:1974:SIG

[Gra74b] Douglas D. Grant. The stability index of graphs. Lecture Notes in Mathematics, 403:29–52, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057373/.

Grauert:1974:DKK

[Gra74c] Hans Grauert. Deformation kompakter komplexer Räume. (German) []. Lecture Notes in Mathematics, 412:70–74, 1974. CODEN LNMAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066154/.

Gray:1974:Cb

[Gra74d] John W. Gray. 2-categories. Lecture Notes in Mathematics, 391: 16–37, 1974. CODEN LNMAA2. ISBN 3-540-06830-9 (print), 3-540-37768-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061282/.

Gray:1974:AMC

[Gra74e] John W. Gray. Adjoint morphisms in 2-categories. Lecture Notes in Mathematics, 391:136–165, 1974. CODEN LNMAA2. ISBN 3-540-06830-9 (print), 3-540-37768-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061286/.

Gray:1974:B

[Gra74f] John W. Gray. Bicategories. Lecture Notes in Mathematics, 391: 38–54, 1974. CODEN LNMAA2. ISBN 3-540-06830-9 (print), 3-540-37768-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061283/.

Gray:1974:Ca

[Gra74g] John W. Gray. Categories. Lecture Notes in Mathematics, 391: 1–15, 1974. CODEN LNMAA2. ISBN 3-540-06830-9 (print), 3-540-37768-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061281/.

Gray:1974:FCT

[Gra74h] John W. Gray. Formal Category Theory: Adjointness for 2-Categories, volume 391 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06830-9 (print), 3-540-37768-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 282 pp. LCCN QA3 .L28 no. 391.

URL http://link.springer.com/book/10.1007/BFb0061280;
 http://www.springerlink.com/content/978-3-540-37768-

Gray:1974:P

[Gra74i] John W. Gray. Properties of Fun(A, B) and Pseud(A, B). Lecture Notes in Mathematics, 391:55–100, 1974. CODEN LN-MAA2. ISBN 3-540-06830-9 (print), 3-540-37768-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061284/.

Gray:1974:PCC

[Gra74j] John W. Gray. Properties of 2-comma categories. Lecture Notes in Mathematics, 391:101–135, 1974. CODEN LNMAA2. ISBN 3-540-06830-9 (print), 3-540-37768-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061285/.

Gray:1974:QA

[Gra74k] John W. Gray. Quasi-adjointness. Lecture Notes in Mathematics, 391:166–271, 1974. CODEN LNMAA2. ISBN 3-540-06830-9 (print), 3-540-37768-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061287/.

Green:1971:GF

[Gre71a] L. W. Green. Geodesic flows. Lecture Notes in Mathematics, 206: 25–28, 1971. CODEN LNMAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070132/.

Greenstadt:1971:CD

[Gre71b] J. Greenstadt. Cell discretization. Lecture Notes in Mathematics, 228:70–85, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069449/.

Green:1973:GLD

[Gre73a] Leon W. Green. Group-like decompositions of Riemannian bundles. Lecture Notes in Mathematics, 318:126–139, 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061729/.

Greenlee:1973:RBM

[Gre73b] W. M. Greenlee. Remarks on branching from multiple eigenvalues. Lecture Notes in Mathematics, 322:101–121, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060563/.

Greenfield:1974:AHP

[Gre74] Stephen J. Greenfield. Approximation and homotopy properties for holomorphic extension in ℓ^{2}. Lecture Notes in Mathematics, 364:150–156, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069014/.

Griffiths:1971:STM

[Gri71a] Phillip A. Griffiths. Some transcendental methods in the study of algebraic cycles. Lecture Notes in Mathematics, 185:1–46, 1971. CODEN LNMAA2. ISBN 3-540-05372-7 (print), 3-540-36493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058762/.

Grillet:1971:DCS

[Gri71b] Pierre-Antoine Grillet. Directed colimits and sheaves in some non-abelian categories. Lecture Notes in Mathematics, 195:36– 69, 1971. CODEN LNMAA2. ISBN 3-540-05442-1 (print), 3-540-36548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0072307/.

Grillet:1971:RC

[Gri71c] Pierre-Antoine Grillet. Regular categories. Lecture Notes in Mathematics, 236:121–222, 1971. CODEN LNMAA2. ISBN 3-540-05678-5 (print), 3-540-36999-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058581/.

Grisvard:1972:RLE

[Gri72] Pierre Grisvard. Résolution locale d’une équation différentielle selon Nirenberg et Trèves. (French) []. Lecture Notes in Mathematics, 244:183–196, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058702/.

Grigelionis:1973:NFT

[Gri73] B. Grigelionis. On nonlinear filtering theory and absolute continuity of measures, corresponding to stochastic processes. Lecture Notes in Mathematics, 330:80–94, 1973. CODEN LN-MAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061481/.

Gross:1970:AWM

[Gro70a] Leonard Gross. Abstract Wiener measure and infinite dimensional potential theory. Lecture Notes in Mathematics, 140:84– 116, 1970. CODEN LNMAA2. ISBN 3-540-04929-0 (print), 3-540-36298-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100068/.

Grossman:1970:SDD

[Gro70b] Steven E. Grossman. Stability in n-dimensional differential-delay equations. Lecture Notes in Mathematics, 144:76–78, 1970. CO-DEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059923/.

Grothendieck:1970:CCG

[Gro70c] A. Grothendieck. Caractérisation et classification des groupes de type multiplicatif. (French) []. Lecture Notes in Mathematics, 152:77–126, 1970. CODEN LNMAA2. ISBN 3-540-05180-5 (print), 3-540-36337-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059008/.

Grothendieck:1970:CRA

[Gro70d] A. Grothendieck. Critères de representabilité. Applications aux sous-groupes de type multiplicatif des schémas en groupes affines. (French) []. Lecture Notes in Mathematics, 152:127–179, 1970. CODEN LNMAA2. ISBN 3-540-05180-5 (print), 3-540-36337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059009/.

Grothendieck:1970:GTM

[Gro70e] A. Grothendieck. Groupes de type multiplicatif: Homomorphismes dans un schéma en groupes. (French) []. Lecture Notes in Mathematics, 152:37–76, 1970. CODEN LNMAA2. ISBN 3-540-05180-5 (print), 3-540-36337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059007/.

Grothendieck:1970:GDF

[Gro70f] A. Grothendieck. Groupes diagonalisables. (French) []. Lecture Notes in Mathematics, 152:1–36, 1970. CODEN LNMAA2. ISBN 3-540-05180-5 (print), 3-540-36337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059006/.

Grothendieck:1970:TMG

[Gro70g] A. Grothendieck. Tores maximaux, groupe de Weyl, sous-groupes de Cartan, centre reductif des schémas en groupes lisses et affines. (French) []. Lecture Notes in Mathematics, 152:180–248, 1970. CODEN LNMAA2. ISBN 3-540-05180-5 (print), 3-540-36337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059010/.

Grothendieck:1970:ERG

[Gro70h] Alexander Grothendieck. Éléments réguliers des groupes algébriques et des algèbres de Lie. (French) []. Lecture Notes in Mathematics, 152:249–295, 1970. CODEN LNMAA2. ISBN 3-540-05180-5 (print), 3-540-36337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059011/.

Grothendieck:1971:EPP

[Gro71a] A. Grothendieck. Esquisse d’un programme pour une theôrie des intersections sur les schémas généraux. (French) []. Lecture Notes in Mathematics, 225:1–77, 1971. CODEN LNMAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066284/.

Grothendieck:1971:POT

[Gro71b] A. Grothendieck. Problèmes ouverts en théorie des intersections. (French) []. Lecture Notes in Mathematics, 225:667–689, 1971.

CODEN LNMAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066297/.

Grosswald:1972:OT

[Gro72a] Emil Grosswald. Oscillation theorems. Lecture Notes in Mathematics, 251:141–168, 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058791/.

Grothendieck:1972:BFG

[Gro72b] A. Grothendieck. Biextensions de faisceaux de groupes. (French) []. Lecture Notes in Mathematics, 288:133–217, 1972. CODEN LNMAA2. ISBN 3-540-05987-3 (print), 3-540-37984-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068692/.

Grothendieck:1972:CBP

[Gro72c] A. Grothendieck. Complements sur les biextensions. Propriétés générales des biextensions des schémas en groupes. (French) []. Lecture Notes in Mathematics, 288:218–312, 1972. CODEN LN-MAA2. ISBN 3-540-05987-3 (print), 3-540-37984-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068693/.

Grothendieck:1972:FFS

[Gro72d] A. Grothendieck. Foncteurs fibrés, supports, étude cohomologique des morphismes finis. (French) []. Lecture Notes in Mathematics, 270:366–412, 1972. CODEN LNMAA2. ISBN 3-540-06012-X (print), 3-540-37987-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061324/.

Grothendieck:1972:STE

[Gro72e] A. Grothendieck. Site et topos étales d’un schéma. (French) []. Lecture Notes in Mathematics, 270:341–365, 1972. CODEN LN-MAA2. ISBN 3-540-06012-X (print), 3-540-37987-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061323/.

Grothendieck:1972:GMG

[Gro72f] A. (Alexandre) Grothendieck. Groupes de Monodromie en Géométrie Algébrique, volume 288 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05987-3 (print), 3-540-37984-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 523 pp. LCCN QA3 .L28 no. 288.

URL http://link.springer.com/book/10.1007/BFb0068688;
 http://www.springerlink.com/content/978-3-540-37984-
3.

Gruenberg:1970:CET

[Gru70a] Karl W. Gruenberg. Classical extension theory. Lecture Notes in Mathematics, 143:65–84, 1970. CODEN LNMAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059167/.

Gruenberg:1970:CD

[Gru70b] Karl W. Gruenberg. Cohomological dimension. Lecture Notes in Mathematics, 143:119–184, 1970. CODEN LNMAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059170/.

Gruenberg:1970:CTG

[Gru70c] Karl W. Gruenberg. Cohomological Topics in Group Theory, volume 143 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 275 pp. LCCN QA3 .L28 no. 143. URL http:/

/link.springer.com/book/10.1007/BFb0059162;   http://
www.springerlink.com/content/978-3-540-36303-3.

Gruenberg:1970:CHG

[Gru70d] Karl W. Gruenberg. The cohomology and homology groups. Lecture Notes in Mathematics, 143:15–29, 1970. CODEN LN-MAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059164/.

Gruenberg:1970:CC

[Gru70e] Karl W. Gruenberg. Comments and corrections. Lecture Notes in Mathematics, 143:A1–A4, 1970. CODEN LNMAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0120208/.

Gruenberg:1970:ECF

[Gru70f] Karl W. Gruenberg. Extension categories: Finite groups. Lecture Notes in Mathematics, 143:249–275, 1970. CODEN LN-MAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059173/.

Gruenberg:1970:ECG

[Gru70g] Karl W. Gruenberg. Extension categories: General theory. Lecture Notes in Mathematics, 143:185–220, 1970. CODEN LN-MAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059171/.

Gruenberg:1970:FGb

[Gru70h] Karl W. Gruenberg. Finite p-groups. Lecture Notes in Mathematics, 143:97–117, 1970. CODEN LNMAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059169/.

Gruenberg:1970:FPF

[Gru70i] Karl W. Gruenberg. Fixed point free action. Lecture Notes in Mathematics, 143:1–13, 1970. CODEN LNMAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059163/.

Gruenberg:1970:FGa

[Gru70j] Karl W. Gruenberg. Free groups. Lecture Notes in Mathematics, 143:51–64, 1970. CODEN LNMAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059166/.

Gruenberg:1970:MCM

[Gru70k] Karl W. Gruenberg. More cohomological machinery. Lecture Notes in Mathematics, 143:85–95, 1970. CODEN LN-MAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059168/.

Gruenberg:1970:MMT

[Gru70l] Karl W. Gruenberg. More module theory. Lecture Notes in Mathematics, 143:221–247, 1970. CODEN LNMAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059172/.

Gruenberg:1970:PR

[Gru70m] Karl W. Gruenberg. Presentations and resolutions. Lecture Notes in Mathematics, 143:31–50, 1970. CODEN LNMAA2. ISBN 3-540-04932-0 (print), 3-540-36303-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059165/.

Gruson:1972:DET

[Gru72] L. Gruson. La dualité dans les espaces de type convexe. (French) []. Lecture Notes in Mathematics, 277:118–165, 1972. CODEN LNMAA2. ISBN 3-540-05934-2 (print), 3-540-37613-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058880/.

Gruenberg:1973:RTM

[Gru73a] K. W. Gruenberg. Ring theoretic methods and finiteness conditions in infinite soluble group theory. Lecture Notes in Mathematics, 319:75–84, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058931/.

Gruman:1973:ODO

[Gru73b] Lawrence Gruman. Opérateurs différentiels d’ordre infini dans des espaces de fonctions entières. (French) []. Lecture Notes in Mathematics, 332:69–76, 1973. CODEN LNMAA2. ISBN 3-540-06372-2 (print), 3-540-46980-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060908/.

Gruman:1974:MTP

[Gru74a] L. Gruman. Majorations de type L^{p} pour les fonctions plurisousharmoniques d’après Hörmander. (French) []. Lecture Notes in Mathematics, 410:77–85, 1974. CODEN LN-MAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066041/.

Grunsky:1974:AFO

[Gru74b] Helmut Grunsky. Analytische Fortsetzung über offene Randkomponenten einer berandeten Riemannschen Fläche. (German) []. Lecture Notes in Mathematics, 419:143–155, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064722/.

Grzegorczyk:1970:DPT

[Grz70] Andrzej Grzegorczyk. Decision procedure for theories categorical in Alef_{0}. Lecture Notes in Mathematics, 125:87–100, 1970. CODEN LNMAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060626/.

Grothendieck:1970:ERS

[GS70] A. Grothendieck and J.-P. Serre. Éléments réguliers: Suite, application aux groupes algébriques. (French) []. Lecture Notes in Mathematics, 152:296–348, 1970. CODEN LNMAA2. ISBN 3-540-05180-5 (print), 3-540-36337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059012/.

Graham:1971:SGF

[GS71] R. L. Graham and J. H. Spencer. On small graphs with forced monochromatic triangles. Lecture Notes in Mathematics, 186: 137–141, 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059431/.

Gorenflo:1974:MEI

[GS74a] R. Gorenflo and H. J. Schaum. Monoton einschliessende Iterationsverfahren for invers-isotone Diskretisierung nicht-linearer Zwei-Punkt-Randwertaufgaben zweiter Ordnung. (German) []. Lecture Notes in Mathematics, 395:177–198, 1974. CODEN LN-MAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060671/.

Guthrie:1974:EM

[GS74b] J. A. Guthrie and H. E. Stone. Expansions of mappings. Lecture Notes in Mathematics, 378:164–171, 1974. CODEN LN-MAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068469/.

Gabriel:1971:AEP

[GU71a] Peter Gabriel and Friedrich Ulmer. Abschätzung von Erzeugungsund Präsentierungszahlen. (German) []. Lecture Notes in Mathematics, 221:163–169, 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059410/.

Gabriel:1971:AKG

[GU71b] Peter Gabriel and Friedrich Ulmer. Algebraische Kategorien. (German) [Algebraic categories]. Lecture Notes in Mathematics, 221:131–142, 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059408/.

Gabriel:1971:APO

[GU71c] Peter Gabriel and Friedrich Ulmer. Anhang, Δ-präsentierbare Objekte. (German) []. Lecture Notes in Mathematics, 221:185– 192, 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059412/.

Gabriel:1971:CSK

[GU71d] Peter Gabriel and Friedrich Ulmer. Charakterisierung und Stetigkeitseigenschaften der Kanschen Erweiterungen. (German) []. Lecture Notes in Mathematics, 221:170–184, 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059411/.

Gabriel:1971:DFG

[GU71e] Peter Gabriel and Friedrich Ulmer. Dichte Funktoren. (German) [Density functors]. Lecture Notes in Mathematics, 221: 38–44, 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059400/.

Gabriel:1971:ERE

[GU71f] Peter Gabriel and Friedrich Ulmer. Echte und reguläre Epimorphismen. (German) []. Lecture Notes in Mathematics, 221:16– 23, 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059398/.

Gabriel:1971:GGS

[GU71g] Peter Gabriel and Friedrich Ulmer. Garben. (German) [Sheaves]. Lecture Notes in Mathematics, 221:143–162, 1971. CODEN LN-MAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059409/.

Gabriel:1971:KEB

[GU71h] Peter Gabriel and Friedrich Ulmer. Kan’sche Erweiterungen, Beispiele. (German) []. Lecture Notes in Mathematics, 221:24– 37, 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059399/.

Gabriel:1971:KSFa

[GU71i] Peter Gabriel and Friedrich Ulmer. Kategorien α-stetiger Funktoren. (German) [Categories of α-continuous functors]. Lecture Notes in Mathematics, 221:55–62, 1971. CODEN LN-MAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059402/.

Gabriel:1971:KSFb

[GU71j] Peter Gabriel and Friedrich Ulmer. Kategorien Σ-stetiger Funktoren. (German) [Categories of Σ-continuous functors]. Lecture Notes in Mathematics, 221:85–104, 1971. CODEN LN-MAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059405/.

Gabriel:1971:LEK

[GU71k] Peter Gabriel and Friedrich Ulmer. Lokal α-erzeugbare Kategorien. (German) [Local α-producible categories]. Lecture Notes in Mathematics, 221:105–120, 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059406/.

Gabriel:1971:LPKa

[GU71l] Peter Gabriel and Friedrich Ulmer. Lokal α-präsentierbare Kategorien. (German) []. Lecture Notes in Mathematics, 221:71– 84, 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059404/.

Gabriel:1971:LPKb

[GU71m] Peter Gabriel and Friedrich Ulmer. Lokal präsentierbare Kategorien. (German) [Local presentable categories], volume 221 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 200 pp. LCCN QA3 .L28 no. 221. URL http:/

/link.springer.com/book/10.1007/BFb0059396;   http://
www.springerlink.com/content/978-3-540-36886-1.

Gabriel:1971:PBE

[GU71n] Peter Gabriel and Friedrich Ulmer. Präsentierbare bzw. erzeugbare Objekte und Generatoren, Beispiele. (German) []. Lecture Notes in Mathematics, 221:63–70, 1971. CODEN LN-MAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059403/.

Gabriel:1971:RBG

[GU71o] Peter Gabriel and Friedrich Ulmer. Retrake, Beispiele. (German) []. Lecture Notes in Mathematics, 221:45–54, 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059401/.

Gabriel:1971:TNG

[GU71p] Peter Gabriel and Friedrich Ulmer. Terminologie und Notation. (German) []. Lecture Notes in Mathematics, 221:13–15, 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059397/.

Gabriel:1971:TLP

[GU71q] Peter Gabriel and Friedrich Ulmer. Tripel in lokal präsentierbaren Kategorien. (German) []. Lecture Notes in Mathematics, 221: 121–130, 1971. CODEN LNMAA2. ISBN 3-540-05578-9 (print), 3-540-36886-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059407/.

Guichardet:1972:BAT

[Gui72a] Alain Guichardet. Boolean algebras of tensor decompositions of a Hilbert space. Lecture Notes in Mathematics, 261:62–79, 1972. CODEN LNMAA2. ISBN 3-540-05803-6 (print), 3-540-37455-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070312/.

Guichardet:1972:CPD

[Gui72b] Alain Guichardet. Conditionally positive definite functions and infinitely divisible positive definite functions. Lecture Notes in Mathematics, 261:47–61, 1972. CODEN LNMAA2. ISBN 3-540-05803-6 (print), 3-540-37455-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070311/.

Guichardet:1972:CP

[Gui72c] Alain Guichardet. Continuous products. Lecture Notes in Mathematics, 261:120–147, 1972. CODEN LNMAA2. ISBN 3-540-05803-6 (print), 3-540-37455-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070315/.

Guichardet:1972:FPD

[Gui72d] Alain Guichardet. Factorizable positive definite functions on current groups. Lecture Notes in Mathematics, 261:80–92, 1972. CODEN LNMAA2. ISBN 3-540-05803-6 (print), 3-540-37455-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070313/.

Guichardet:1972:GMT

[Gui72e] Alain Guichardet. Gaussian measures on topological vector spaces. Lecture Notes in Mathematics, 261:93–119, 1972. CO-DEN LNMAA2. ISBN 3-540-05803-6 (print), 3-540-37455-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070314/.

Guichardet:1972:I

[Gui72f] Alain Guichardet. Introduction. Lecture Notes in Mathematics, 261:1–2, 1972. CODEN LNMAA2. ISBN 3-540-05803-6 (print), 3-540-37455-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070307.

Guichardet:1972:PDF

[Gui72g] Alain Guichardet. Positive definite functions of type (S). Lecture Notes in Mathematics, 261:30–46, 1972. CODEN LN-MAA2. ISBN 3-540-05803-6 (print), 3-540-37455-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070310/.

Guichardet:1972:SHSa

[Gui72h] Alain Guichardet. The symmetric Hilbert space of a Hilbert space. Lecture Notes in Mathematics, 261:18–29, 1972. CODEN LNMAA2. ISBN 3-540-05803-6 (print), 3-540-37455-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070309/.

Guichardet:1972:SHSb

[Gui72i] Alain Guichardet. Symmetric Hilbert Spaces and Related Topics: Infinitely Divisible Positive Definite Functions Continuous Products and Tensor Products Gaussian and Poissonian Stochastic Processes, volume 261 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05803-6 (print), 3-540-37455-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 197 pp. LCCN QA3 .L28 no. 261. URL http:/

/link.springer.com/book/10.1007/BFb0070306;   http://
www.springerlink.com/content/978-3-540-37455-8.

Guichardet:1972:SMS

[Gui72j] Alain Guichardet. The symmetric measure space of a measure space. Lecture Notes in Mathematics, 261:3–17, 1972. CODEN LNMAA2. ISBN 3-540-05803-6 (print), 3-540-37455-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070308/.

Gunning:1971:ASS

[Gun71] R. C. Gunning. Analytic structures on the space of flat vector bundles over a compact Riemann surface. Lecture Notes in Mathematics, 185:47–62, 1971. CODEN LNMAA2. ISBN 3-540-05372-7 (print), 3-540-36493-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058763/.

Gupta:1974:NIE

[Gup74] Chaitan P. Gupta. On nonlinear integral equations of Hammerstein type with unbounded linear mapping. Lecture Notes in Mathematics, 384:184–238, 1974. CODEN LNMAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057549/.

Gusak:1973:CPT

[Gus73a] D. V. Gusak. On the continuous passage through a fixed level of a homogeneous process with independent increments on a Markov chain. Lecture Notes in Mathematics, 330:95–103, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061482/.

Gustafson:1973:TRG

[Gus73b] William H. Gustafson. The theory of relative Grothendieck rings. Lecture Notes in Mathematics, 353:95–112, 1973. CODEN LN-MAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059263/.

Garnir:1974:FAA

[GUW74] H. G. Garnir, K. R. Unni, and J. H. Williamson, editors. Functional Analysis and its Applications: International Conference, Madras, 1973, volume 399 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 399. URL http:/

/link.springer.com/book/10.1007/BFb0063562;   http://
www.springerlink.com/content/978-3-540-37827-3.

Guy:1971:LRC

[Guy71] Richard K. Guy. Latest results on crossing numbers. Lecture Notes in Mathematics, 186:143–156, 1971. CODEN LN-MAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059432/.

Guy:1972:CNG

[Guy72] Richard K. Guy. Crossing numbers of graphs. Lecture Notes in Mathematics, 303:111–124, 1972. CODEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067363/.

Grothendieck:1972:CFT

[GV72a] A. Grothendieck and J.-L. Verdier. Conditions de finitude. Topos et sites fibrés. Applications aux questions de passage à la limite. (French) []. Lecture Notes in Mathematics, 270:163–340, 1972. CODEN LNMAA2. ISBN 3-540-06012-X (print), 3-540-37987-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061322/.

Grothendieck:1972:PFP

[GV72b] A. Grothendieck and J. L. Verdier. Préfaisceaux. (French) [Presheaves]. Lecture Notes in Mathematics, 269:1–217, 1972. CODEN LNMAA2. ISBN 3-540-05896-6 (print), 3-540-37549-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081552/.

Grothendieck:1972:T

[GV72c] A. Grothendieck and J. L. Verdier. Topos. Lecture Notes in Mathematics, 269:299–518, 1972. CODEN LNMAA2. ISBN 3-540-05896-6 (print), 3-540-37549-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081555/.

Glicksberg:1972:RFS

[GW72] I. Glicksberg and I. Wik. The range of Fourier–Stieltjes transforms of parts of measures. Lecture Notes in Mathematics, 266: 73–77, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059638/.

Gatterdam:1973:CGT

[GW73] R. W. Gatterdam and K. W. Weston, editors. Conference on Group Theory: University of Wisconsin–Parkside 1972, volume 319 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 319. URL http:/

/link.springer.com/book/10.1007/BFb0058922;   http://
www.springerlink.com/content/978-3-540-38479-3.

Ganelius:1974:DAM

[GW74a] Tord Ganelius and Stig Westlund. The degree of approximation in Müntz’s theorem. Lecture Notes in Mathematics, 419: 125–132, 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064719/.

Gilbert:1974:ISE

[GW74b] Robert P. Gilbert and Richard J. Weinacht. Iterative schemes for elliptic systems. Lecture Notes in Mathematics, 430:253– 260, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066272/.

Haack:1974:SZL

[Haa74] Wolfgang Haack. Systeme von zwei linearen partiellen Differentialgleichungen erster Ordnung in zwei Variablen vom hyperbolischen Typus mit parabolischer Anfangskurve. (German) [].

Lecture Notes in Mathematics, 419:156–165, 1974. CODEN LN-MAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064723/.

Habets:1974:SPV

[Hab74] P. Habets. Singular perturbations of a vector boundary value problem. Lecture Notes in Mathematics, 415:149–154, 1974. CO-DEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065523/.

Haefliger:1971:LTG

[Hae71a] A. Haefliger. Lectures on the theorem of Gromov. Lecture Notes in Mathematics, 209:128–141, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068900/.

Haefliger:1971:HI

[Hae71b] André Haefliger. Homotopy and integrability. Lecture Notes in Mathematics, 197:133–163, 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068615/.

Haefliger:1973:CCF

[Hae73] André Haefliger. Sur les classes caractéristiques des feuilletages. (French) []. Lecture Notes in Mathematics, 317:239–260, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069285/.

Hager:1974:VLU

[Hag74] Anthony W. Hager. Vector lattices of uniformly continuous functions and some categorical methods in uniform spaces. Lecture Notes in Mathematics, 378:172–187, 1974. CODEN LN-MAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068470/.

Haimo:1973:IAA

[Hai73] Franklin Haimo. The index of an algebra automorphism group. Lecture Notes in Mathematics, 319:85–90, 1973. CODEN LN-MAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058932/.

Hajek:1970:PTD

[Háj70] Otomar Hájek. Prolongations in topological dynamics. Lecture Notes in Mathematics, 144:79–89, 1970. CODEN LN-MAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059924/.

Hajek:1973:NMM

[Háj73] Otomar Hájek. Non-minimality of 3-manifolds. Lecture Notes in Mathematics, 318:140–142, 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061730/.

Haken:1974:SSP

[Hak74] Wolfgang Haken. Some special presentations of homotopy 3spheres. Lecture Notes in Mathematics, 375:97–107, 1974. CO-DEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064015/.

Halanay:1970:HOO

[Hal70a] A. Halanay. Hyperstability, oscillations and optimal stabilization. Lecture Notes in Mathematics, 144:90–114, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059925/.

Halpern:1970:ATS

[Hal70b] Benjamin Halpern. Algebraic topology and set-valued maps. Lecture Notes in Mathematics, 171:27–33, 1970. CODEN LN-MAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069716/.

Hale:1971:FDE

[Hal71] Jack K. Hale. Functional differential equations. Lecture Notes in Mathematics, 183:9–22, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060406/.

Hall:1973:NGE

[Hal73] Marshall Hall, Jr. Notes on groups of exponent four. Lecture Notes in Mathematics, 319:91–118, 1973. CODEN LN-MAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058933/.

Hamm:1971:TIS

[Ham71] H. A. Hamm. Topology of isolated singularities of complex spaces. Lecture Notes in Mathematics, 209:213–217, 1971. CO-DEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068906/.

Hanson:1970:SRC

[Han70] D. L. Hanson. Some results on convergence rates for weighted averages. Lecture Notes in Mathematics, 160:53–63, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060646/.

Hansen:1971:HTA

[Han71a] W. Hansen. Hunt’s theorem and axiomatic potential theory. Lecture Notes in Mathematics, 190:15–18, 1971. CODEN LN-MAA2. ISBN 3-540-05396-4 (print), 3-540-36515-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065887/.

Hansen:1971:CHS

[Han71b] Wolfhard Hansen. Cohomology in harmonic spaces. Lecture Notes in Mathematics, 226:63–101, 1971. CODEN LNMAA2. ISBN 3-540-05638-6 (print), 3-540-36912-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060481/.

Hannula:1973:MCC

[Han73] T. A. Hannula. The Morita context and the construction of QF rings. Lecture Notes in Mathematics, 353:113–130, 1973. CODEN LNMAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059264/.

Hartshorne:1970:AOS

[Har70a] Robin Hartshorne. Affine open subsets. Lecture Notes in Mathematics, 156:59–80, 1970. CODEN LNMAA2. ISBN 3-540-05184-8 (print), 3-540-36345-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067841/.

Hartshorne:1970:AGA

[Har70b] Robin Hartshorne. Algebraic geometry and analytic geometry. Lecture Notes in Mathematics, 156:210–236, 1970. CODEN LN-MAA2. ISBN 3-540-05184-8 (print), 3-540-36345-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067845/.

Hartshorne:1970:AD

[Har70c] Robin Hartshorne. Ample divisors. Lecture Notes in Mathematics, 156:1–58, 1970. CODEN LNMAA2. ISBN 3-540-05184-8 (print), 3-540-36345-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067840/.

Hartshorne:1970:ASA

[Har70d] Robin Hartshorne. Ample Subvarieties of Algebraic Varieties: Notes written in Collaboration with C. Musili, volume 156 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05184-8 (print), 3-540-36345-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 256 pp. LCCN QA3 .L28 no. 156. URL http:/

/link.springer.com/book/10.1007/BFb0067839;   http://
www.springerlink.com/content/978-3-540-36345-3.

Hartshorne:1970:FRF

[Har70e] Robin Hartshorne. Formal-rational functions along a subvariety. Lecture Notes in Mathematics, 156:186–209, 1970. CODEN LNMAA2. ISBN 3-540-05184-8 (print), 3-540-36345-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067844/.

Hartshorne:1970:GHC

[Har70f] Robin Hartshorne. Generalization to higher codimensions. Lecture Notes in Mathematics, 156:81–161, 1970. CODEN LN-MAA2. ISBN 3-540-05184-8 (print), 3-540-36345-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067842/.

Hartshorne:1970:GLT

[Har70g] Robin Hartshorne. The Grothendieck–Lefschetz theorems. Lecture Notes in Mathematics, 156:162–185, 1970. CODEN LN-MAA2. ISBN 3-540-05184-8 (print), 3-540-36345-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067843/.

Harris:1971:ATD

[Har71a] W. A. Harris, Jr. Analytic theory of difference equations. Lecture Notes in Mathematics, 183:45–58, 1971. CODEN LN-MAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060411/.

Harris:1971:ATL

[Har71b] W. A. Harris, Jr. Analytic theory of linear differential systems. Lecture Notes in Mathematics, 243:229–237, 1971. CODEN LN-MAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058731/.

Hartman:1971:DW

[Har71c] Philip Hartman. Disconjugacy and Wronskians. Lecture Notes in Mathematics, 243:208–218, 1971. CODEN LNMAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058729/.

Harary:1972:RRG

[Har72a] Frank Harary. Recent results on generalized Ramsey theory for graphs. Lecture Notes in Mathematics, 303:125–138, 1972. CO-DEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067364/.

Harris:1972:FPK

[Har72b] Christopher J. Harris. The Fokker–Planck–Kolmogorov equation in the analysis of nonlinear feedback stochastic systems. Lecture Notes in Mathematics, 294:230–238, 1972. CODEN LN-MAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064944/.

Harris:1973:SPB

[Har73] W. A. Harris, Jr. Singularly perturbed boundary value problems revisited. Lecture Notes in Mathematics, 312:54–64, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060045/.

Harary:1974:RRG

[Har74a] Frank Harary. Recent results on graphical enumeration. Lecture Notes in Mathematics, 406:29–36, 1974. CODEN LN-MAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066432/.

Harary:1974:RRT

[Har74b] Frank Harary. Recent results on trees. Lecture Notes in Mathematics, 406:1–9, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic). URL http://link.springer.com/
chapter/10.1007/BFb0066429/;   http://link.springer.
com/content/pdf/bfm:978-3-540-37809-9/1.pdf.

Harary:1974:SGR

[Har74c] Frank Harary. A survey of generalized Ramsey theory. Lecture Notes in Mathematics, 406:10–17, 1974. CODEN LN-MAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066430/.

Harary:1974:SRC

[Har74d] Frank Harary. A survey of the reconstruction conjecture. Lecture Notes in Mathematics, 406:18–28, 1974. CODEN LN-MAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066431/.

Harley:1974:MCI

[Har74e] Peter W. Harley III. Metrization of closed images of metric spaces. Lecture Notes in Mathematics, 378:188–191, 1974. CO-DEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068471/.

Harper:1974:HT

[Har74f] John R. Harper. The mod3 homotopy type of F_{4}. Lecture Notes in Mathematics, 418:58–67, 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070642/.

Harper:1974:DFS

[Har74g] John R. Harper. modp decompositions of finite H-spaces. Lecture Notes in Mathematics, 428:44–51, 1974. CODEN LN-MAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070533/.

Harris:1974:RS

[Har74h] Douglas Harris. Representation of spaces. Lecture Notes in Mathematics, 378:192–200, 1974. CODEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068472/.

Harris:1974:BSH

[Har74i] Lawrence A. Harris. Bounded symmetric homogeneous domains in infinite dimensional spaces. Lecture Notes in Mathematics, 364: 13–40, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069002/.

Harzallah:1974:DHI

[Har74j] K. Harzallah. Distances hilbertiennes invariantes sur un espace homogène. (French) []. Lecture Notes in Mathematics, 404:133– 137, 1974. CODEN LNMAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060613/.

Hass:1972:KDH

[Has72] R. Hass. Konvergenz von Differenzenverfahren für halblineare Anfangswertaufgaben. (German) []. Lecture Notes in Mathematics, 267:99–116, 1972. CODEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061616/.

Hass:1974:BSD

[Has74] R. Hass. Bemerkungen zur Stabilität von Differenzenverfahren für lineare und halblineare Anfangswertaufgaben. (German) []. Lecture Notes in Mathematics, 395:19–31, 1974. CODEN LN-MAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060662/.

Hatcher:1973:PI

[Hat73] Allen E. Hatcher. Pseudo-isotopy and K_{2}. Lecture Notes in Mathematics, 342:328–336, 1973. CODEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073731/.

Haussmann:1972:SLS

[Hau72] U. G. Haussmann. Stabilization of linear systems with multiplicative noise. Lecture Notes in Mathematics, 294:125–130, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064938/.

Hausen:1973:PAG

[Hau73] Jutta Hausen. Primary abelian groups and the normal structure of their automorphism groups. Lecture Notes in Mathematics, 319:119–122, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058934/.

Haver:1974:CPM

[Hav74] William E. Haver. A covering property for metric spaces. Lecture Notes in Mathematics, 375:108–113, 1974. CODEN LN-MAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064016/.

Hayli:1974:MDI

[Hay74] Avram Hayli. The method of the doubly individual step for N-body computations. Lecture Notes in Mathematics, 362:304– 312, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066598/.

Heine:1970:SMI

[HB70] Walter Heine and John V. Breakwell. Symmetric minimumimpulse rendezvous between certain non-coplanar orbits. Lecture Notes in Mathematics, 132:130–150, 1970. CODEN LN-MAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066679/.

Harish-Chandra:1970:BCS

[HC70a] Harish-Chandra. The behavior of the characters of the supercuspidal representations on the regular set. Lecture Notes in Mathematics, 162:43–62, 1970. CODEN LNMAA2. ISBN 3-540-05189-9 (print), 3-540-36372-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061275/.

Harish-Chandra:1970:ECD

[HC70b] Harish-Chandra. Existence of characters for the discrete series. Lecture Notes in Mathematics, 162:4–11, 1970. CODEN LNMAA2. ISBN 3-540-05189-9 (print), 3-540-36372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061271/.

Harish-Chandra:1970:ECG

[HC70c] Harish-Chandra. Existence of characters in the general case. Lecture Notes in Mathematics, 162:12–19, 1970. CODEN LN-MAA2. ISBN 3-540-05189-9 (print), 3-540-36372-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061272/.

Harish-Chandra:1970:HAR

[HC70d] Harish-Chandra. Harmonic Analysis on Reductive p-adic Groups, volume 162 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05189-9 (print), 3-540-36372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 125 pp. LCCN QA3 .L28 no. 162. URL http:

//link.springer.com/book/10.1007/BFb0061269;   http:/
/www.springerlink.com/content/978-3-540-36372-9.

Harish-Chandra:1970:I

[HC70e] Harish-Chandra. Introduction. Lecture Notes in Mathematics, 162:1–3, 1970. CODEN LNMAA2. ISBN 3-540-05189-9 (print), 3-540-36372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061270/.

Harish-Chandra:1970:LSCa

[HC70f] Harish-Chandra. The local summability of |D|^{− \frac{1}{2} −ε} (char θ = 0). Lecture Notes in Mathematics, 162:86–91, 1970. CODEN LNMAA2. ISBN 3-540-05189-9 (print), 3-540-36372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061277/.

Harish-Chandra:1970:LSCb

[HC70g] Harish-Chandra. The local summability of the characters of the supercuspidal representations (char θ = 0). Lecture Notes in Mathematics, 162:92–125, 1970. CODEN LNMAA2. ISBN 3-540-05189-9 (print), 3-540-36372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061278/.

Harish-Chandra:1970:MC

[HC70h] Harish-Chandra. The mapping ‘F_{f}’ (char θ = 0). Lecture Notes in Mathematics, 162:63–85, 1970. CODEN LNMAA2. ISBN 3-540-05189-9 (print), 3-540-36372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061276/.

Harish-Chandra:1970:SAS

[HC70i] Harish-Chandra. Some applications of the Schwartz space of a semisimple Lie group. Lecture Notes in Mathematics, 140:1– 7, 1970. CODEN LNMAA2. ISBN 3-540-04929-0 (print), 3-540-36298-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0100063/.

Harish-Chandra:1970:S

[HC70j] Harish-Chandra. The space A(G, τ). Lecture Notes in Mathematics, 162:30–42, 1970. CODEN LNMAA2. ISBN 3-540-05189-9 (print), 3-540-36372-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061274/.

Harish-Chandra:1970:SFS

[HC70k] Harish-Chandra. Supercusp forms and supercuspidal representations. Lecture Notes in Mathematics, 162:20–29, 1970. CODEN LNMAA2. ISBN 3-540-05189-9 (print), 3-540-36372-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061273/.

Harish-Chandra:1972:TEI

[HC72] Harish-Chandra. On the theory of the Eisenstein integral. Lecture Notes in Mathematics, 266:123–149, 1972. CODEN LN-MAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059640/.

Heath:1974:SSP

[Hea74a] David Heath. Skorokhod stopping via potential theory. Lecture Notes in Mathematics, 381:150–154, 1974. CODEN LN-MAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057260/.

Heath:1974:SCS

[Hea74b] R. W. Heath. Some comments on simultaneous extensions of mappings on closed subsets. Lecture Notes in Mathematics, 375: 114–119, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064017/.

Heath:1974:SRE

[Hea74c] R. W. Heath. Some remarks on extension theorems in monotonically normal spaces. Lecture Notes in Mathematics, 378:201– 206, 1974. CODEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068473/.

Hechler:1974:DSU

[Hec74a] Stephen H. Hechler. A dozen small uncountable cardinals. Lecture Notes in Mathematics, 378:207–218, 1974. CODEN LN-MAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068474/.

Hector:1974:AGD

[Hec74b] Gilbert Hector. Actions de groupes de diffeomorphismes de {0, 1}. (French) []. Lecture Notes in Mathematics, 392:14–22, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064105/.

Hedetniemi:1971:GM

[Hed71] Stephen T. Hedetniemi. Graphs of (0, 1)-matrices. Lecture Notes in Mathematics, 186:157–171, 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059433/.

Hedstrom:1972:SNE

[Hed72] G. W. Hedstrom. Some numerical experiments with Dafermos’s method for nonlinear hyperbolic equations. Lecture Notes in Mathematics, 267:117–138, 1972. CODEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061617/.

Heinkel:1973:CSP

[Hei73a] Bernard Heinkel. Une condition suffisante pour la continuité presque sûre des trajectoires de certains processus gaussiens.

(French) []. Lecture Notes in Mathematics, 321:77–94, 1973. CO-DEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071398/.

Heinzer:1973:NII

[Hei73b] William Heinzer. Noetherian intersections of integral domains II. Lecture Notes in Mathematics, 311:107–119, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068924/.

Heinkel:1974:TDD

[Hei74a] B. Heinkel. Thèorémes de dérivation du type de celui de Lebesgue et continuité presque sure des trajectoires de certains processus gaussiens. (French) []. Lecture Notes in Mathematics, 381:155– 171, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057261/.

Heins:1974:GRP

[Hei74b] Albert E. Heins. The generalized radiation problem. Lecture Notes in Mathematics, 419:166–177, 1974. CODEN LN-MAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064724/.

Helgason:1972:HAN

[Hel72] Sigurdur Helgason. Harmonic analysis in the non-Euclidean disk. Lecture Notes in Mathematics, 266:151–156, 1972. CODEN LN-MAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059641/.

Helson:1973:CDS

[Hel73] Henry Helson. Convergence of Dirichlet series. Lecture Notes in Mathematics, 336:153–160, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065797/.

Hemminger:1972:LD

[Hem72] Robert L. Hemminger. Line digraphs. Lecture Notes in Mathematics, 303:149–163, 1972. CODEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067366/.

Henrici:1971:CAD

[Hen71] Peter Henrici. Circular arithmetic and the determination of polynomial zeros. Lecture Notes in Mathematics, 228:86–92, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069450/.

Henry:1973:ASF

[Hen73] Myron S. Henry. Approximate solutions of functional differential equations. Lecture Notes in Mathematics, 333:144–152, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060693/.

Herrera:1970:RCS

[Her70] M. Herrera. Residues on complex spaces. Lecture Notes in Mathematics, 155:110–114, 1970. CODEN LNMAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060318/.

Herve:1971:APF

[Her71a] Michel Hervé, editor. Analytic and Plurisubharmonic Functions: in Finite and Infinite Dimensional Spaces Course Given at the University of Maryland, Spring 1970, volume 198 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05472-3 (print), 3-540-36654-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 198.

URL http://link.springer.com/book/10.1007/BFb0070695;
 http://www.springerlink.com/content/978-3-540-36654-
6.

Herve:1971:ACB

[Her71b] Michel Hervé. Analytic continuation on Banach spaces. Lecture Notes in Mathematics, 185:63–75, 1971. CODEN LN-MAA2. ISBN 3-540-05372-7 (print), 3-540-36493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058764/.

Herve:1971:CPT

[Her71c] Michel Hervé. Classical potential theory. Lecture Notes in Mathematics, 198:1–30, 1971. CODEN LNMAA2. ISBN 3-540-05472-3 (print), 3-540-36654-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070696/.

Herve:1971:CAI

[Her71d] Michel Hervé. Complex analysis in infinite dimensional vector spaces. Lecture Notes in Mathematics, 198:56–90, 1971. CODEN LNMAA2. ISBN 3-540-05472-3 (print), 3-540-36654-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070698/.

Herve:1971:PSA

[Her71e] Michel Hervé. Plurisuperharmonicity and separate analyticity with respect to a finite number of variables. Lecture Notes in Mathematics, 198:31–55, 1971. CODEN LNMAA2. ISBN 3-540-05472-3 (print), 3-540-36654-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070697/.

Hertling:1972:NTH

[Her72a] Jörg Hertling. Numerical treatment of Hammerstein-equations by variational methods. Lecture Notes in Mathematics, 267:265– 288, 1972. CODEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061622/.

Herz:1972:PES

[Her72b] C. Herz. Problems of extrapolation and spectral synthesis on groups. Lecture Notes in Mathematics, 266:157–166, 1972. CO-DEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059642/.

Hertling:1973:NCP

[Her73] Jörg Hertling. Numerical calculation of primary bifurcation points of the Hammerstein operator. Lecture Notes in Mathematics, 333:153–163, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060694/.

Hertling:1974:NLE

[Her74a] J. Hertling. Zur numerischen Lösung einiger singulärer Differential- und Integralgleichungen. (German) []. Lecture Notes in Mathematics, 395:263–274, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060677/.

Herve:1974:LPI

[Her74b] Michel Hervé. Lindelöf’s principle in infinite dimensions. Lecture Notes in Mathematics, 364:41–57, 1974. CODEN LN-MAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069003/.

Hervier:1974:WPB

[Her74c] Yves Hervier. On the Weierstrass problem in Banach spaces. Lecture Notes in Mathematics, 364:157–167, 1974. CODEN LN-MAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069015/.

Hervier:1974:PLB

[Her74d] Yves Hervier. Sur le problème de Levi Banachique. (French) []. Lecture Notes in Mathematics, 410:18–27, 1974. CODEN LNMAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066037/.

Herz:1974:GNC

[Her74e] Carl Herz. Généralisations de la notion des classes H_{p} de Hardy. (French) []. Lecture Notes in Mathematics, 404:138–148, 1974. CODEN LNMAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060614/.

Hess:1974:NPL

[Hes74a] Peter Hess. Nonlinear perturbations of linear elliptic operators. Lecture Notes in Mathematics, 415:167–172, 1974. CODEN LN-MAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065525/.

Hess:1974:SNE

[Hes74b] Peter Hess. On some nonlinear elliptic boundary value problems. Lecture Notes in Mathematics, 399:235–247, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063581/.

Heyer:1970:CDG

[Hey70a] Herbert Heyer. Chu-Dualität. (German) [Chu duality]. Lecture Notes in Mathematics, 150:329–363, 1970. CODEN LN-MAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069790/.

Heyer:1970:DKG

[Hey70b] Herbert Heyer. Darstellungen kompakter Gruppen. (German) []. Lecture Notes in Mathematics, 150:75–96, 1970. CODEN LNMAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069782/.

Heyer:1970:DLG

[Hey70c] Herbert Heyer. Dualität Lokalkompakter Gruppen. (German) [Duality of Locally Compact Groups], volume 150 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CO-DEN LNMAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiii + 372 pp. LCCN QA3 .L28 no. 150. URL http:/

/link.springer.com/book/10.1007/BFb0069778;   http://
www.springerlink.com/content/978-3-540-36313-2.

Heyer:1970:DPG

[Hey70d] Herbert Heyer. Dualitätssatz von Pontrjagin. (German) []. Lecture Notes in Mathematics, 150:148–163, 1970. CODEN LN-MAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069784/.

Heyer:1970:DTGa

[Hey70e] Herbert Heyer. Dualitätssatz von Takahashi. (German) []. Lecture Notes in Mathematics, 150:266–289, 1970. CODEN LN-MAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069788/.

Heyer:1970:DTGb

[Hey70f] Herbert Heyer. Dualitätssatz von Tatsuuma. (German) []. Lecture Notes in Mathematics, 150:290–328, 1970. CODEN LN-MAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069789/.

Heyer:1970:DTK

[Hey70g] Herbert Heyer. Dualitätssätze von Tannaka und Krein. (German) []. Lecture Notes in Mathematics, 150:97–147, 1970. CODEN LNMAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069783/.

Heyer:1970:IDG

[Hey70h] Herbert Heyer. Irreduzible Darstellungen. (German) []. Lecture Notes in Mathematics, 150:57–74, 1970. CODEN LN-MAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069781/.

Heyer:1970:MFGa

[Hey70i] Herbert Heyer. Maximal fastperiodische gruppen: Elementare theorie. Lecture Notes in Mathematics, 150:164–191, 1970. CO-DEN LNMAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069785/.

Heyer:1970:MFGb

[Hey70j] Herbert Heyer. Maximal fastperiodische Gruppen: Struktursätze. (German) []. Lecture Notes in Mathematics, 150:192– 226, 1970. CODEN LNMAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069786/.

Heyer:1970:VG

[Hey70k] Herbert Heyer. Vorbereitungen. (German) []. Lecture Notes in Mathematics, 150:1–24, 1970. CODEN LNMAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069779/.

Heyer:1970:ZGG

[Hey70l] Herbert Heyer. Zentrale Gruppen. (German) [Central groups]. Lecture Notes in Mathematics, 150:227–265, 1970. CODEN LN-MAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069787/.

Heyer:1970:ZDG

[Hey70m] Herbert Heyer. Zyklische Darstellungen. (German) []. Lecture Notes in Mathematics, 150:25–56, 1970. CODEN LN-MAA2. ISBN 3-540-04939-8 (print), 3-540-36313-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069780/.

Heyer:1972:IDP

[Hey72] Herbert Heyer. Infinitely divisible probability measures on compact groups. Lecture Notes in Mathematics, 247:55–249, 1972. CODEN LNMAA2. ISBN 3-540-05729-3 (print), 3-540-37117-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058552/.

Heyer:1973:GCD

[Hey73] Herbert Heyer. Groups with chu duality. Lecture Notes in Mathematics, 296:181–215, 1973. CODEN LNMAA2. ISBN 3-540-06211-4 (print), 3-540-38485-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059825/.

Helton:1973:IOT

[HH73] J. William Helton and Roger E. Howe. Integral operators: Traces, index, and homology. Lecture Notes in Mathematics, 345:141– 209, 1973. CODEN LNMAA2. ISBN 3-540-06496-6 (print), 3-540-37808-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058919/.

Hausmann:1971:EHT

[HHP71] J.-Cl. Hausmann, A. Haefliger, and V. Poénaru. Extension d’une homotopie transverse à un feuilletage topologique. (French) []. Lecture Notes in Mathematics, 197:164–175, 1971. CODEN LN-MAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068616/.

Hickin:1973:CTL

[Hic73] Kenneth K. Hickin. Countable type local theorems in algebra. Lecture Notes in Mathematics, 319:123–125, 1973. CODEN LN-MAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058935/.

Higman:1974:CCG

[Hig74] D. G. Higman. Coherent configurations and generalized polygons. Lecture Notes in Mathematics, 403:1–5, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057370/.

Hilton:1971:SAT

[Hil71] Peter J. Hilton. Symposium on Algebraic Topology, volume 249 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 111 pp. LCCN QA3 .L28 no. 249. URL http:/

/link.springer.com/book/10.1007/BFb0060889;   http://
www.springerlink.com/content/978-3-540-37082-6.

Hilton:1974:LNS

[Hil74a] Peter Hilton. Localization of nilpotent spaces. Lecture Notes in Mathematics, 428:18–43, 1974. CODEN LNMAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070532/.

Hilton:1974:DLN

[Hil74b] Peter Hilton. On direct limits of nilpotent groups. Lecture Notes in Mathematics, 418:68–77, 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070643/.

Hilton:1974:LGT

[Hil74c] Prof. Dr. Peter J. Hilton. Localization in Group Theory and Homotopy Theory: and Related Topics, volume 418 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 171 (est.) pp. LCCN QA3 .L28 no. 418. URL http:/

/link.springer.com/book/10.1007/BFb0070635;   http://
www.springerlink.com/content/978-3-540-37268-4.

Hirschowitz:1971:BEF

[Hir71a] André Hirschowitz. Bornologie des espaces de fonctions analytiques en dimension infinie. (French) []. Lecture Notes in Mathematics, 205:21–33, 1971. CODEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059831/.

Hirschowitz:1971:DNO

[Hir71b] André Hirschowitz. Diverses notions d’ouverts d’analyticité en dimension infinie. (French) []. Lecture Notes in Mathematics, 205: 11–20, 1971. CODEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059830/.

Hirzebruch:1971:PCR

[Hir71c] F. Hirzebruch. Pontrjagin classes of rational homology manifolds and the signature of some affine hypersurfaces. Lecture Notes in Mathematics, 209:207–212, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068905/.

Hirai:1972:SIR

[Hir72a] Takeshi Hirai. Structure of induced representations and characters of irreducible representations of complex semisimple Lie groups. Lecture Notes in Mathematics, 266:167–188, 1972. CO-DEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059643/.

Hirschelmann:1972:AUP

[Hir72b] Arnulf Hirschelmann. An application of ultra-products to prime rings with polynomial identities. Lecture Notes in Mathematics, 255:145–148, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059543/.

Hirschowitz:1972:CAE

[Hir72c] André Hirschowitz. Convexité abstraite et espaces normés. (French) []. Lecture Notes in Mathematics, 275:1–13, 1972. CO-DEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058489/.

Hirzebruch:1972:HMG

[Hir72d] F. Hirzebruch. The Hilbert modular group, resolution of the singularities at the cusps and related problems. Lecture Notes in Mathematics, 244:275–288, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058707/.

Hirschowitz:1973:GCA

[Hir73] André Hirschowitz. Le groupe de Cremona d’après Demazure. (French) []. Lecture Notes in Mathematics, 317:261–276, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069286/.

Hirsch:1974:TSF

[Hir74a] Francis Hirsch. Transformation de Stieltjes et fonctions opérant sur les potentiels abstraits. (French) []. Lecture Notes in Mathematics, 404:149–163, 1974. CODEN LNMAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060615/.

Hirschfeld:1974:MAS

[Hir74b] Joram Hirschfeld. Models of arithmetic and the semi-ring of recursive functions. Lecture Notes in Mathematics, 369:99–105, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066003/.

Hirzebruch:1974:KHM

[Hir74c] F. Hirzebruch. Kurven auf den Hilbertschen Modulflächen und Klassenzahlrelationen. (German) []. Lecture Notes in Mathematics, 412:75–93, 1974. CODEN LNMAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066155/.

Hajian:1970:EEM

[HK70] Arshag B. Hajian and Shizuo Kakutani. Example of an ergodic measure preserving transformation on an infinite measure space. Lecture Notes in Mathematics, 160:45–52, 1970. CODEN LN-MAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060645/.

Herzog:1971:KMC

[HK71a] Jürgen Herzog and Ernst Kunz. Der kanonische Modul eines Cohen–Macaulay-Rings. (German) [The canonical module of a Cohen–Macaulay ring], volume 238 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05683-1 (print), 3-540-37000-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 103 pp. LCCN QA3 .L28 no. 238. URL http:/

/link.springer.com/book/10.1007/BFb0059377;   http://
www.springerlink.com/content/978-3-540-37000-0.

Herzog:1971:VKGa

[HK71b] Jürgen Herzog and Ernst Kunz. Vortrag: E. Kunz. (German) [Lecture: E. Kunz]. Lecture Notes in Mathematics, 238:17– 24, 1971. CODEN LNMAA2. ISBN 3-540-05683-1 (print), 3-540-37000-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059379/.

Herzog:1971:VKGb

[HK71c] Jürgen Herzog and Ernst Kunz. Vortrag: E. Kunz. (German) [Lecture: E. Kunz]. Lecture Notes in Mathematics, 238:85– 102, 1971. CODEN LNMAA2. ISBN 3-540-05683-1 (print), 3-540-37000-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059384/.

Herzog:1971:VHS

[HK71d] Jürgen Herzog and Ernst Kunz. Vortrag: Heinz Schmitt. (German) [Lecture: Heinz Schmitt]. Lecture Notes in Mathematics, 238:33–44, 1971. CODEN LNMAA2. ISBN 3-540-05683-1 (print), 3-540-37000-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059381/.

Herzog:1971:VJHa

[HK71e] Jürgen Herzog and Ernst Kunz. Vortrag: J. Herzog. (German) [Lecture: J. Herzog]. Lecture Notes in Mathematics, 238:25– 32, 1971. CODEN LNMAA2. ISBN 3-540-05683-1 (print), 3-540-37000-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059380/.

Herzog:1971:VJHb

[HK71f] Jürgen Herzog and Ernst Kunz. Vortrag: J. Herzog. (German) [Lecture: J. Herzog]. Lecture Notes in Mathematics, 238:59– 84, 1971. CODEN LNMAA2. ISBN 3-540-05683-1 (print), 3-540-37000-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059383/.

Herzog:1971:VME

[HK71g] Jürgen Herzog and Ernst Kunz. Vortrag: Margret Engelken. (German) []. Lecture Notes in Mathematics, 238:1–16, 1971. CO-DEN LNMAA2. ISBN 3-540-05683-1 (print), 3-540-37000-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059378/.

Herzog:1971:VRW

[HK71h] Jürgen Herzog and Ernst Kunz. Vortrag: R. Waldi. (German) [Lecture: R. Waldi]. Lecture Notes in Mathematics, 238:45– 58, 1971. CODEN LNMAA2. ISBN 3-540-05683-1 (print), 3-540-37000-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059382/.

Huss:1971:CMS

[HK71i] R. Huss and R. Kalaba. Computation of the moments of solutions of certain random two point boundary value problems. Lecture Notes in Mathematics, 228:93–102, 1971. CODEN LN-MAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069451/.

Hansen:1974:PHN

[HL74a] P. Hansen and M. Las Vergnas, C. N. R. S. On a property of hypergraphs with no cycles of length greater than two. Lecture Notes in Mathematics, 411:99–101, 1974. CODEN LN-MAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066184/.

Harris:1974:AIL

[HL74b] W. A. Harris, Jr. and D. S. Lutz. Asymptotic integration of linear systems. Lecture Notes in Mathematics, 415:155–166, 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065524/.

Hurd:1974:VSN

[HL74c] Albert Hurd and Peter Loeb, editors. Victoria Symposium on Nonstandard Analysis: University of Victoria 1972, volume 369 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 369. URL http:/

/link.springer.com/book/10.1007/BFb0065992;   http://
www.springerlink.com/content/978-3-540-37928-7.

Hajek:1971:GDD

[HLM71] O. Hájek, A. J. Lohwater, and R. McCann, editors. Global Differentiable Dynamics: Proceedings of the Conference held at Case Western Reserve University, Cleveland, Ohio, June 2–6, 1969, volume 235 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 235. URL http:/

/link.springer.com/book/10.1007/BFb0059188;   http://
www.springerlink.com/content/978-3-540-36996-7.

Henson:1974:INH

[HM74a] C. Ward Henson and L. C. Moore, Jr. Invariance of the nonstandard hulls of a uniform space. Lecture Notes in Mathematics, 369:85–98, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066002/.

Henson:1974:SRN

[HM74b] C. Ward Henson and L. C. Moore, Jr. Semi-reflexivity of the nonstandard hulls of a locally convex space. Lecture Notes in Mathematics, 369:71–84, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066001/.

Hilton:1971:SBS

[HMR71] Peter Hilton, Guido Mislin, and Joseph Roitberg. Sphere bundles over spheres and non-cancellation phenomena. Lecture Notes in Mathematics, 249:34–46, 1971. CODEN LNMAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060894/.

Hsia:1973:PCO

[HMR73] John S. Hsia, Manohar L. Madan, and Thomas G. Ralley, editors. Proceedings of the Conference on Orders, Group Rings and Related Topics, volume 353 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 353. URL http:/

/link.springer.com/book/10.1007/BFb0059257;   http://
www.springerlink.com/content/978-3-540-37818-1.

Hofmann:1974:ADL

[HMS74a] Karl Heinrich Hofmann, Michael Mislove, and Albert Stralka. Application of duality to lattice theory. Lecture Notes in Mathematics, 396:53–91, 1974. CODEN LNMAA2. ISBN 3-540-06807-4 (print), 3-540-37741-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065933/.

Hofmann:1974:ADS

[HMS74b] Karl Heinrich Hofmann, Michael Mislove, and Albert Stralka. Applications of duality to the structure theory of compact zero dimensional semilattices. Lecture Notes in Mathematics, 396: 92–108, 1974. CODEN LNMAA2. ISBN 3-540-06807-4 (print), 3-540-37741-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065934/.

Hofmann:1974:CDS

[HMS74c] Karl Heinrich Hofmann, Michael Mislove, and Albert Stralka. The category of discrete semilattices and the category of compact zero-dimensional semilattices. Lecture Notes in Mathematics, 396:5–27, 1974. CODEN LNMAA2. ISBN 3-540-06807-4 (print), 3-540-37741-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065931/.

Hofmann:1974:CTC

[HMS74d] Karl Heinrich Hofmann, Michael Mislove, and Albert Stralka. The character theory of compact and discrete semilattices. Lecture Notes in Mathematics, 396:28–52, 1974. CODEN LN-MAA2. ISBN 3-540-06807-4 (print), 3-540-37741-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065932/.

Hofmann:1974:PDC

[HMS74e] Karl Heinrich Hofmann, Michael Mislove, and Albert Stralka. The Pontryagin Duality of Compact 0-Dimensional Semilattices and its Applications, volume 396 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06807-4 (print), 3-540-37741-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvi + 122 pp. LCCN QA3 .L28 no. 396.

URL http://link.springer.com/book/10.1007/BFb0065929;
 http://www.springerlink.com/content/978-3-540-37741-
2.

Hofmann:1974:P

[HMS74f] Karl Heinrich Hofmann, Michael Mislove, and Albert Stralka. Preliminaries. Lecture Notes in Mathematics, 396:1–4, 1974. CO-DEN LNMAA2. ISBN 3-540-06807-4 (print), 3-540-37741-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065930/.

Hogbe-Nlend:1971:ADE

[HN71a] H. Hogbe-Nlend. Applications différentiables entre espaces bornologiques. (French) []. Lecture Notes in Mathematics, 213:137– 158, 1971. CODEN LNMAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069429/.

Hogbe-Nlend:1971:BBD

[HN71b] H. Hogbe-Nlend. Bornologies à base denombrable dans les espaces vectories topologiques. (French) []. Lecture Notes in Mathematics, 213:90–100, 1971. CODEN LNMAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069425/.

Hogbe-Nlend:1971:BEB

[HN71c] H. Hogbe-Nlend. Bornologies Équicontinues — bornologies de Schwartz. (French) []. Lecture Notes in Mathematics, 213:56– 66, 1971. CODEN LNMAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069423/.

Hogbe-Nlend:1971:BNF

[HN71d] H. Hogbe-Nlend. Bornologies Nucléaires. (French) [Nuclear bornologies]. Lecture Notes in Mathematics, 213:67–89, 1971. CODEN LNMAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069424/.

Hogbe-Nlend:1971:CEF

[HN71e] H. Hogbe-Nlend. Contre-exemples. (French) [Counterexamples]. Lecture Notes in Mathematics, 213:104–111, 1971. CODEN LN-MAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069427/.

Hogbe-Nlend:1971:DEB

[HN71f] H. Hogbe-Nlend. Distributions dans les espaces bornologiques et distributions opérationnelles. (French) []. Lecture Notes in Mathematics, 213:112–136, 1971. CODEN LNMAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069428/.

Hogbe-Nlend:1971:DBF

[HN71g] H. Hogbe-Nlend. Dualité bornologique. (French) [Bornological duality]. Lecture Notes in Mathematics, 213:46–55, 1971. CO-DEN LNMAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069422/.

Hogbe-Nlend:1971:EBC

[HN71h] H. Hogbe-Nlend. Espaces bornologiques complets. (French) []. Lecture Notes in Mathematics, 213:28–40, 1971. CODEN LN-MAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069420/.

Hogbe-Nlend:1971:MCF

[HN71i] H. Hogbe-Nlend. La Mackey-convergence. (French) []. Lecture Notes in Mathematics, 213:7–21, 1971. CODEN LNMAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069418/.

Hogbe-Nlend:1971:TDG

[HN71j] H. Hogbe-Nlend. Le théorème du graphe Mackey–Ferme. (French) []. Lecture Notes in Mathematics, 213:41–45, 1971. CO-DEN LNMAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069421/.

Check title: Ferme ou Fermé??

Hogbe-Nlend:1971:NFF

[HN71k] H. Hogbe-Nlend. Notions fondamentales. (French) [Fundamental concepts]. Lecture Notes in Mathematics, 213:1–6, 1971. CODEN LNMAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069417/.

Hogbe-Nlend:1971:SEM

[HN71l] H. Hogbe-Nlend. Séparation — espaces Mackey-complets. (French) []. Lecture Notes in Mathematics, 213:22–27, 1971. CO-DEN LNMAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069419/.

Hogbe-Nlend:1971:TBA

[HN71m] H. Hogbe-Nlend. Théorie des Bornologies et Applications. (French) [Theory of Bornologies and Applications], volume 213 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 168 pp. LCCN QA3 .L28 no. 213. URL http:/

/link.springer.com/book/10.1007/BFb0069416;   http://
www.springerlink.com/content/978-3-540-36878-6.

Hogbe-Nlend:1971:TBP

[HN71n] H. Hogbe-Nlend. Topologies et bornologies pseudo-convexes. (French) []. Lecture Notes in Mathematics, 213:101–103, 1971. CODEN LNMAA2. ISBN 3-540-05546-0 (print), 3-540-36878-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069426/.

Hogbe-Nlend:1973:TBT

[HN73] H. Hogbe-Nlend. Techniques de bornologie en théorie des espaces vectoriels topologiques. (French) []. Lecture Notes in Mathematics, 331:84–162, 1973. CODEN LNMAA2. ISBN 3-540-06367-6 (print), 3-540-46977-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068228/.

Hunsaker:1974:EHS

[HN74] W. N. Hunsaker and S. A. Naimpally. C-embedding of Hausdorff spaces in certain extensions. Lecture Notes in Mathematics, 375:

137–142, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064019/.

Hamrick:1972:UBM

[HO72] Gary Hamrick and Erich Ossa. Unitary bordism of monogenic groups and isometries. Lecture Notes in Mathematics, 298:172– 182, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070041/.

Handelsman:1973:AAC

[HO73] Richard A. Handelsman and W. E. Olmstead. Asymptotic analysis of a class of nonlinear integral equations. Lecture Notes in Mathematics, 322:122–129, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060564/.

Hoare:1971:PPA

[Hoa71] C. A. R. Hoare. Procedures and parameters: an axiomatic approach. Lecture Notes in Mathematics, 188:102–116, 1971. CO-DEN LNMAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059696/.

Hochster:1973:CMM

[Hoc73] Melvin Hochster. Cohen–Macaulay modules. Lecture Notes in Mathematics, 311:120–152, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068925/.

Hodges:1972:CML

[Hod72] Wilfrid Hodges, editor. Conference in Mathematical Logic — London ’70, volume 255 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 255. URL http:/

/link.springer.com/book/10.1007/BFb0059533;   http://
www.springerlink.com/content/978-3-540-37162-5.

Hodel:1974:SRM

[Hod74] R. E. Hodel. Some results in metrization theory, 1950–1972. Lecture Notes in Mathematics, 375:120–136, 1974. CODEN LN-MAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064018/.

Hofmann:1970:A

[Hof70a] K. H. Hofmann. C∗-algebras. Lecture Notes in Mathematics, 129: 28–34, 1970. CODEN LNMAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071216/.

Hofmann:1970:C

[Hof70b] K. H. Hofmann. C∗-cogebras. Lecture Notes in Mathematics, 129: 48–67, 1970. CODEN LNMAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071220/.

Hofmann:1970:ACM

[Hof70c] K. H. Hofmann. Algebras over a category with multiplication. Lecture Notes in Mathematics, 129:11–18, 1970. CODEN LN-MAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071213/.

Hofmann:1970:B

[Hof70d] K. H. Hofmann. Bigebras. Lecture Notes in Mathematics, 129: 23–27, 1970. CODEN LNMAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071215/.

Hofmann:1970:CCM

[Hof70e] K. H. Hofmann. Cogebras over a category with multiplication. Lecture Notes in Mathematics, 129:18–22, 1970. CODEN LN-MAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071214/.

Hofmann:1970:CSC

[Hof70f] K. H. Hofmann. The confunctor spec for commutative C∗- algebras. Lecture Notes in Mathematics, 129:44–46, 1970. CO-DEN LNMAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071218/.

Hofmann:1970:DD

[Hof70g] K. H. Hofmann. Definition of duality. Lecture Notes in Mathematics, 129:1–2, 1970. CODEN LNMAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071211.

Hofmann:1970:DIS

[Hof70h] K. H. Hofmann. Discrete involutive semigroups and duality. Lecture Notes in Mathematics, 129:126–139, 1970. CODEN LN-MAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071225/.

Hofmann:1970:DCG

[Hof70i] K. H. Hofmann. Dualities for compact groups. Lecture Notes in Mathematics, 129:112–122, 1970. CODEN LNMAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071223/.

Hofmann:1970:DTD

[Hof70j] K. H. Hofmann. Dualities for totally disconnected compact semigroups. Lecture Notes in Mathematics, 129:122–126, 1970. CO-DEN LNMAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071224/.

Hofmann:1970:DCSa

[Hof70k] K. H. Hofmann. The duality of compact semigroups. Lecture Notes in Mathematics, 129:67–78, 1970. CODEN LN-MAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071221/.

Hofmann:1970:DCSb

[Hof70l] K. H. Hofmann. The Duality of Compact Semigroups and C∗- Bigebras, volume 129 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 142 pp. LCCN QA3 .L28 no. 129. URL http:/

/link.springer.com/book/10.1007/BFb0071210;   http://
www.springerlink.com/content/978-3-540-36269-2.

Hofmann:1970:FCC

[Hof70m] K. H. Hofmann. Finite coproducts of commutative C∗-algebras with identity. Lecture Notes in Mathematics, 129:46–48, 1970. CODEN LNMAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071219/.

Hofmann:1970:IBT

[Hof70n] K. H. Hofmann. Involutive bigebras and their relation to C∗- bigebras. Lecture Notes in Mathematics, 129:78–112, 1970. CO-DEN LNMAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071222/.

Hofmann:1970:MCC

[Hof70o] K. H. Hofmann. A multiplication for the category C∗. Lecture Notes in Mathematics, 129:34–44, 1970. CODEN LN-MAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071217/.

Hofmann:1970:MC

[Hof70p] K. H. Hofmann. Multiplicative categories. Lecture Notes in Mathematics, 129:3–10, 1970. CODEN LNMAA2. ISBN 3-540-04918-5 (print), 3-540-36269-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071212/.

Hoffman:1971:CAG

[Hof71] A. J. Hoffman. Combinatorial aspects of Gerschgorin’s theorem. Lecture Notes in Mathematics, 186:173–179, 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059434/.

Hoffman:1972:LPS

[Hof72a] Alan J. Hoffman. On limit points of spectral radii of non-negative symmetric integral matrices. Lecture Notes in Mathematics, 303: 165–172, 1972. CODEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067367/.

Hofmann:1972:ROT

[Hof72b] Karl Heinrich Hofmann, editor. Ring and Operator Theory. Los Angeles, USA 1970–1971. Lectures on Operator Algebras, volume 247 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05729-3 (print), 3-540-37117-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 786 pp. LCCN QA3 .L28 no. 247. URL http:

//link.springer.com/book/10.1007/BFb0058550;   http:/
/www.springerlink.com/content/978-3-540-37117-5.

Hofmann:1972:TUR

[Hof72c] Karl Heinrich Hofmann, editor. Tulane University ring and operator theory year, 1970/71. Lectures on Rings and Modules, volume 246 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05760-9 (print), 3-540-37181-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 661 pp. LCCN QA3 .L28 no. 246. URL http:

//link.springer.com/book/10.1007/BFb0059562;   http:/
/www.springerlink.com/content/978-3-540-37181-6.

Holmes:1972:CP

[Hol72a] Richard B. Holmes. Comments on the problems. Lecture Notes in Mathematics, 257:128–144, 1972. CODEN LNMAA2. ISBN 3-540-05764-1 (print), 3-540-37182-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059454/.

Holmes:1972:COB

[Hol72b] Richard B. Holmes. A Course on Optimization and Best Approximation, volume 257 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05764-1 (print), 3-540-37182-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 233 pp. LCCN QA3 .L28 no. 257. URL http:/

/link.springer.com/book/10.1007/BFb0059450;   http://
www.springerlink.com/content/978-3-540-37182-3.

Holmes:1972:P

[Hol72c] Richard B. Holmes. Preliminaries. Lecture Notes in Mathematics, 257:1–13, 1972. CODEN LNMAA2. ISBN 3-540-05764-1 (print), 3-540-37182-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059451/.

Holmes:1972:SST

[Hol72d] Richard B. Holmes. Selected special topics. Lecture Notes in Mathematics, 257:145–233, 1972. CODEN LNMAA2. ISBN 3-540-05764-1 (print), 3-540-37182-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059455/.

Holmes:1972:TBA

[Hol72e] Richard B. Holmes. Theory of best approximation. Lecture Notes in Mathematics, 257:76–127, 1972. CODEN LNMAA2. ISBN 3-540-05764-1 (print), 3-540-37182-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059453/.

Holmes:1972:TO

[Hol72f] Richard B. Holmes. Theory of optimization. Lecture Notes in Mathematics, 257:14–75, 1972. CODEN LNMAA2. ISBN 3-540-05764-1 (print), 3-540-37182-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059452/.

Holevo:1973:SPQ

[Hol73] A. S. Holevo. Statistical problems in quantum physics. Lecture Notes in Mathematics, 330:104–119, 1973. CODEN LN-MAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061483/.

Holton:1974:S

[Hol74a] D. A. Holton. Stability. Lecture Notes in Mathematics, 403: 53–55, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057374/.

Holton:1974:CMP

[Hol74b] Derek A. Holton, editor. Combinatorial Mathematics: Proceedings of the Second Australian Conference [University of Melbourne, 1973], volume 403 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 403. URL http:/

/link.springer.com/book/10.1007/BFb0057369;   http://
www.springerlink.com/content/978-3-540-37837-2.

Hong:1974:LOR

[Hon74a] Sung Sa Hong. Limit-operators and reflective subcategories. Lecture Notes in Mathematics, 378:219–227, 1974. CODEN LN-MAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068475/.

Honig:1974:GFL

[Hön74b] Chaim Samuel Hönig. The Green function of a linear differential equation with lateral condition. Lecture Notes in Mathematics, 384:147–167, 1974. CODEN LNMAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057546/.

Hoppensteadt:1971:GAB

[Hop71] Frank C. Hoppensteadt. A geometric approach to boundary value problems for nonlinear ordinary differential equations with a small parameter. Lecture Notes in Mathematics, 183:35–41, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060409/.

Horowitz:1970:NFI

[Hor70a] S. Horowitz. A note on σ-finite invariant measures. Lecture Notes in Mathematics, 160:64–70, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060647/.

Horvath:1970:SCV

[Hor70b] John Horváth, editor. Several Complex Variables I Maryland 1970: Proceedings of the International Mathematical Conference, held at College Park, April 6–17, 1970, volume 155 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 155.

URL http://link.springer.com/book/10.1007/BFb0060312;
 http://www.springerlink.com/content/978-3-540-36344-
6.

Horvath:1971:SCV

[Hor71] John Horváth, editor. Several Complex Variables II Maryland 1970: Proceedings of the International Mathematical Conference, held at College Park, April 6–17, 1970, volume 185 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05372-7 (print), 3-540-36493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 185.

URL http://link.springer.com/book/10.1007/BFb0058761;
 http://www.springerlink.com/content/978-3-540-36493-
1.

Horvath:1973:LCS

[Hor73] John Horváth. Locally convex spaces. Lecture Notes in Mathematics, 331:41–83, 1973. CODEN LNMAA2. ISBN 3-540-06367-6 (print), 3-540-46977-X (e-book). ISSN 0075-8434 (print),

1617-9692 (electronic). URL http://link.springer.com/
chapter/10.1007/BFb0068226/;   http://link.springer.
com/chapter/10.1007/BFb0068227/.

Houzel:1972:PTF

[Hou72a] C. Houzel. Produits tensoriels. (French) [Tensor products]. Lecture Notes in Mathematics, 277:166–229, 1972. CODEN LN-MAA2. ISBN 3-540-05934-2 (print), 3-540-37613-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058881/.

Houzel:1972:SBF

[Hou72b] C. Houzel. Séminaire Banach. (French) [Banach seminar], volume 277 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05934-2 (print), 3-540-37613-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 229 (est.) pp. LCCN QA3 .L28 no. 277. URL http:/

/link.springer.com/book/10.1007/BFb0058876;   http://
www.springerlink.com/content/978-3-540-37613-2.

Howes:1971:SBG

[How71] Leonard Howes. On subdominantly bounded graphs — summary of results. Lecture Notes in Mathematics, 186:181–183, 1971. CODEN LNMAA2. ISBN 3-540-05386-7 (print), 3-540-36508-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059435/.

Howard:1974:PSA

[How74] Bernard E. Howard. Phase space analysis in numerical integration of ordinary differential equations. Lecture Notes in Mathematics, 362:107–127, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066588/.

Hilton:1970:CTF

[HR70] Peter Hilton and Joseph Roitberg. On the classification of torsion-free H-spaces of rank 2. Lecture Notes in Mathematics, 168:67–74, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058518/.

Hilton:1971:QFO

[HR71] Peter Hilton and Joseph Roitberg. On quasi-fibrations and orthogonal bundles. Lecture Notes in Mathematics, 196:100– 106, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067752/.

Harary:1974:NGP

[HR74a] Frank Harary and Ronald C. Read. Is the null-graph a pointless concept? Lecture Notes in Mathematics, 406:37–44, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066433/.

Hirsch:1974:ODC

[HR74b] Francis Hirsch and Jean-Pierre Roth. Opérateurs dissipatifs et codissipatifs invariants sur un espace homogene. (French) []. Lecture Notes in Mathematics, 404:229–245, 1974. CODEN LN-MAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060618/.

Hoobler:1974:AT

[HR74c] R. Hoobler and D. L. Rector. Arithmetic K-theory. Lecture Notes in Mathematics, 418:78–95, 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070644/.

Huck:1973:BDI

[HRS^{+}73a] H. Huck, R. Roitzsch, U. Simon, W. Vortisch, R. Walden, B. Wegner, and W. Wendland. Beweismethoden der Differentialgeometrie im Großen. (German) [], volume 335 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06385-4 (print), 3-540-46990-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 160 (est.) pp. LCCN QA3 .L28 no. 335. URL http:/

/link.springer.com/book/10.1007/BFb0061770;   http://
www.springerlink.com/content/978-3-540-46990-2.

Huck:1973:IG

[HRS^{+}73b] H. Huck, R. Roitzsch, U. Simon, W. Vortisch, R. Walden, B. Wegner, and W. Wendland. Die Indexmethode. (German) []. Lecture Notes in Mathematics, 335:40–69, 1973. CODEN LNMAA2. ISBN 3-540-06385-4 (print), 3-540-46990-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061772/.

Huck:1973:IIG

[HRS^{+}73c] H. Huck, R. Roitzsch, U. Simon, W. Vortisch, R. Walden, B. Wegner, and W. Wendland. Die Integralformelmethode Integralformeln. (German) []. Lecture Notes in Mathematics, 335: 70–108, 1973. CODEN LNMAA2. ISBN 3-540-06385-4 (print), 3-540-46990-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061773/.

Huck:1973:MG

[HRS^{+}73d] H. Huck, R. Roitzsch, U. Simon, W. Vortisch, R. Walden, B. Wegner, and W. Wendland. Die Maximummethode. (German) []. Lecture Notes in Mathematics, 335:109–160, 1973. CODEN LNMAA2. ISBN 3-540-06385-4 (print), 3-540-46990-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061774/.

Huck:1973:EG

[HRS^{+}73e] H. Huck, R. Roitzsch, U. Simon, W. Vortisch, R. Walden, B. Wegner, and W. Wendland. Einführung. (German) []. Lecture Notes in Mathematics, 335:10–39, 1973. CODEN LN-MAA2. ISBN 3-540-06385-4 (print), 3-540-46990-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061771/.

Holm:1970:CPD

[HS70] P. Holm and E. Spanier. Compact perturbations and degree theory. Lecture Notes in Mathematics, 168:75–89, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058519/.

Hsieh:1971:ATD

[HS71] P. F. Hsieh and A. W. J. Stoddart, editors. Analytic Theory of Differential Equations: The Proceedings of the Conference at Western Michigan University, Kalamazoo, from 30 April to 2 May 1970, volume 183 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 183. URL http:/

/link.springer.com/book/10.1007/BFb0060404;   http://
www.springerlink.com/content/978-3-540-36454-2.

Hedetniemi:1972:LGT

[HS72] S. T. Hedetniemi and P. J. Slater. Line graphs of triangleless graphs and iterated clique graphs. Lecture Notes in Mathematics, 303:139–147, 1972. CODEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067365/.

Harris:1973:SAD

[HS73a] B. Harris and J. Stasheff. Suspension, automorphisms, and division algebras. Lecture Notes in Mathematics, 342:337–346, 1973. CODEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073732/.

Harris:1973:SOD

[HS73b] William A. Harris, Jr. and Yasutaka Sibuya, editors. Symposium on Ordinary Differential Equations: Minneapolis, Minnesota, May 29–30, 1972, volume 312 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 312. URL http:/

/link.springer.com/book/10.1007/BFb0060041;   http://
www.springerlink.com/content/978-3-540-38353-6.

Harary:1974:WGI

[HS74a] Frank Harary and Allen J. Schwenk. Which graphs have integral spectra? Lecture Notes in Mathematics, 406:45–51, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066434/.

Hayden:1974:PID

[HS74b] T. L. Hayden and T. J. Suffridge, editors. Proceedings on Infinite Dimensional Holomorphy: University of Kentucky 1973, volume 364 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 364. URL http:/

/link.springer.com/book/10.1007/BFb0069000;   http://
www.springerlink.com/content/978-3-540-37915-7.

Hsiang:1971:M

[Hsi71] Wu-Chung Hsiang. Manifolds with π_{1} = Z^{k}. Lecture Notes in Mathematics, 197:36–43, 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068610/.

Hsieh:1973:BSN

[Hsi73] P. F. Hsieh. Bounded solutions of nonlinear equations at an irregular type singularity. Lecture Notes in Mathematics, 312: 65–73, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060046/.

Hengartner:1974:CI

[HT74] W. Hengartner and R. Theodorescu. Concentration and information. Lecture Notes in Mathematics, 398:99–110, 1974. CODEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065771/.

Hull:1974:DSS

[Hul74] T. E. Hull. The development of software for solving ordinary differential equations. Lecture Notes in Mathematics, 363:55– 63, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069125/.

Humphreys:1973:VMC

[Hum73] J. E. Humphreys. Variations on Milnor’s computation of K_{2}Z. Lecture Notes in Mathematics, 342:304–307, 1973. CODEN LN-MAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073729/.

Hung:1974:MIT

[Hun74] Nguyen-Trung Hung. Sur les mesures d’information de type inf. (French) []. Lecture Notes in Mathematics, 398:62–75, 1974. CO-DEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065768/.

Hurd:1974:NPB

[Hur74] A. E. Hurd. Near periods and Bohr compactifications. Lecture Notes in Mathematics, 369:106–112, 1974. CODEN LN-MAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066004/.

Husseini:1970:SF

[Hus70] S. Y. Husseini. Spherical fibrations. Lecture Notes in Mathematics, 168:107–124, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058521/.

Husemoller:1973:HF

[Hus73] Dale Husemoller. On the homology of the fibre of ψ^{q} − 1. Lecture Notes in Mathematics, 341:199–204, 1973. CODEN LN-MAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067057/.

Husain:1974:QBA

[Hus74a] T. Husain. Quasicomplemented Banach algebras. Lecture Notes in Mathematics, 399:248–253, 1974. CODEN LNMAA2. ISBN 3-540-06869-4 (print), 3-540-37827-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063582/.

Husch:1974:SRA

[Hus74b] L. S. Husch. Semi-regular actions. Lecture Notes in Mathematics, 375:143–148, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064020/.

Isaev:1970:IOR

[IDS70] V. K. Isaev, B. Kh. Davidson, and V. V. Sonin. An investigation of optimal rocket flight in the vicinity of a planet. Lecture Notes in Mathematics, 112:266–285, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060210/.

Igarashi:1971:SAL

[Iga71] Shigeru Igarashi. Semantics of Algol-like statements. Lecture Notes in Mathematics, 188:117–177, 1971. CODEN LN-MAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059697/.

Ihara:1973:OCW

[Iha73] Shunsuke Ihara. Optimal coding in white Gaussian channel with feedback. Lecture Notes in Mathematics, 330:120–123, 1973. CO-DEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061484/.

Igusa:1971:CRS

[iI71] Jun ichi Igusa. On certain representations of semi-simple algebraic groups. Lecture Notes in Mathematics, 185:76–86, 1971. CODEN LNMAA2. ISBN 3-540-05372-7 (print), 3-540-36493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058765/.

Illusie:1971:CFR

[Ill71a] L. Illusie. Conditions de finitude relatives. (French) []. Lecture Notes in Mathematics, 225:222–273, 1971. CODEN LN-MAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066287/.

Illusie:1971:ERG

[Ill71b] L. Illusie. Existence de Résolutions globales. (French) []. Lecture Notes in Mathematics, 225:160–221, 1971. CODEN LN-MAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066286/.

Illusie:1971:GCF

[Ill71c] L. Illusie. Généralités sur les conditions de finitude dans les catégories derivées. (French) []. Lecture Notes in Mathematics, 225:78–159, 1971. CODEN LNMAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066285/.

Illusie:1971:GGT

[Ill71d] L. Illusie. Groupes de Grothendieck des topos anneles. (French) []. Lecture Notes in Mathematics, 225:274–296, 1971. CODEN LNMAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066288/.

Illusie:1971:AHR

[Ill71e] Luc Illusie. Algébre homotopique relative. (French) []. Lecture Notes in Mathematics, 239:1–113, 1971. CODEN LN-MAA2. ISBN 3-540-05686-6 (print), 3-540-37001-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059053/.

Illusie:1971:CCC

[Ill71f] Luc Illusie. Classes de Chern des complexes parfaits. (French) []. Lecture Notes in Mathematics, 239:263–342, 1971. CODEN LNMAA2. ISBN 3-540-05686-6 (print), 3-540-37001-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059057/.

Illusie:1971:CCDa

[Ill71g] Luc Illusie. Complexe cotangent: Definition et premières propriétés. (French) []. Lecture Notes in Mathematics, 239:114– 150, 1971. CODEN LNMAA2. ISBN 3-540-05686-6 (print), 3-540-37001-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059054/.

Illusie:1971:CCDb

[Ill71h] Luc Illusie. Complexe Cotangent et Déformations I. (French) [Complex cotangent and deformations I], volume 239 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05686-6 (print), 3-540-37001-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xv + 355 pp. LCCN QA3 .L28 no. 239. URL http:/

/link.springer.com/book/10.1007/BFb0059052;   http://
www.springerlink.com/content/978-3-540-37001-7.

Illusie:1971:CCE

[Ill71i] Luc Illusie. Complexe cotangent et extensions infinitesimales. (French) []. Lecture Notes in Mathematics, 239:151–224, 1971.

CODEN LNMAA2. ISBN 3-540-05686-6 (print), 3-540-37001-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059055/.

Illusie:1971:DMF

[Ill71j] Luc Illusie. Deformations de modules. (French) []. Lecture Notes in Mathematics, 239:225–262, 1971. CODEN LNMAA2. ISBN 3-540-05686-6 (print), 3-540-37001-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059056/.

Illusie:1971:EFE

[Ill71k] Luc Illusie. Errata. (French) [Errata]. Lecture Notes in Mathematics, 239:e1–e2, 1971. CODEN LNMAA2. ISBN 3-540-05686-6 (print), 3-540-37001-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059058.

Illusie:1972:BM

[Ill72a] Luc Illusie. Back matter. Lecture Notes in Mathematics, 274: 191–192, 1972. CODEN LNMAA2. ISBN 3-540-05920-2 (print), 3-540-37609-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-37609-5/1.pdf.

Illusie:1972:CFC

[Ill72b] Luc Illusie. Catégories formelles, complexes de de Rham et cohomologie cristalline. (French) []. Lecture Notes in Mathematics, 283:245–292, 1972. CODEN LNMAA2. ISBN 3-540-05976-8 (print), 3-540-37962-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059576/.

Illusie:1972:CDF

[Ill72c] Luc Illusie. Cohomologie de diagrammes. (French) []. Lecture Notes in Mathematics, 283:1–179, 1972. CODEN LN-MAA2. ISBN 3-540-05976-8 (print), 3-540-37962-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059574/.

Illusie:1972:CCDb

[Ill72d] Luc Illusie. Complexe Cotangent et Déformations II. (French) [Complex cotangent and deformations II], volume 283 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05976-8 (print), 3-540-37962-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 304 pp. LCCN QA3 .L28 no. 283. URL http:/

/link.springer.com/book/10.1007/BFb0059573;   http://
www.springerlink.com/content/978-3-540-37962-1.

Illusie:1972:CCDa

[Ill72e] Luc Illusie. Cotangent complex and deformations of torsors and group schemes. Lecture Notes in Mathematics, 274:159– 189, 1972. CODEN LNMAA2. ISBN 3-540-05920-2 (print), 3-540-37609-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073969/.

Illusie:1972:DES

[Ill72f] Luc Illusie. Déformations Équivariantes de G-schémas et déformations de schémas en groupes. (French) []. Lecture Notes in Mathematics, 283:180–244, 1972. CODEN LNMAA2. ISBN 3-540-05976-8 (print), 3-540-37962-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059575/.

Illusie:1972:EFE

[Ill72g] Luc Illusie. Errata. (French) [Errata]. Lecture Notes in Mathematics, 283:i, 1972. CODEN LNMAA2. ISBN 3-540-05976-8 (print), 3-540-37962-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059577.

Illusie:1973:TQC

[Ill73] Luc Illusie. Travaux de Quillen sur la cohomologie des groupes. (French) []. Lecture Notes in Mathematics, 317:89–105, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069278/.

Infante:1971:SRA

[Inf71] E. F. Infante. Some results and applications of generalized dynamical systems. Lecture Notes in Mathematics, 243:91–105, 1971. CODEN LNMAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058721/.

Ingleton:1971:CRT

[Ing71] A. W. Ingleton. Conditions for representability and transversality of matroids. Lecture Notes in Mathematics, 211:62–66, 1971. CODEN LNMAA2. ISBN 3-540-05545-2 (print), 3-540-36877-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061075/.

Ivanilov:1970:POD

[IP70] Yu. P. Ivanilov and A. A. Petrov. The problems of optimalization in dynamic multi-branch industrial models. Lecture Notes in Mathematics, 112:286–293, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060211/.

Ito:1972:SDC

[Itô72] Kiyosi Itô. Stochastic differentials of continuous local quasimartingales. Lecture Notes in Mathematics, 294:1–7, 1972. CO-DEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064929/.

Iversen:1970:CPS

[Ive70a] Birger Iversen. Construction of the Picard scheme of a family of curves. Lecture Notes in Mathematics, 174:52–68, 1970. CODEN LNMAA2. ISBN 3-540-05301-8 (print), 3-540-36439-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069478/.

Iversen:1970:ISR

[Ive70b] Birger Iversen. Invertible sheaves and rational maps into C^{(g)}. Lecture Notes in Mathematics, 174:32–51, 1970. CODEN LN-MAA2. ISBN 3-540-05301-8 (print), 3-540-36439-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069477/.

Iversen:1970:LD

[Ive70c] Birger Iversen. The linear determinant. Lecture Notes in Mathematics, 174:1–19, 1970. CODEN LNMAA2. ISBN 3-540-05301-8 (print), 3-540-36439-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069475/.

Iversen:1970:LDA

[Ive70d] Birger Iversen. Linear Determinants with Applications to the Picard Scheme of a Family of Algebraic Curves, volume 174 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05301-8 (print), 3-540-36439-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 69 pp. LCCN QA3 .L28 no. 174. URL http:/

/link.springer.com/book/10.1007/BFb0069474;   http://
www.springerlink.com/content/978-3-540-36439-9.

Iversen:1970:RFS

[Ive70e] Birger Iversen. Representation of n-fold sections by symmetric products. Lecture Notes in Mathematics, 174:20–31, 1970. CODEN LNMAA2. ISBN 3-540-05301-8 (print), 3-540-36439-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069476/.

Iversen:1973:EM

[Ive73a] Birger Iversen. Etale morphisms. Lecture Notes in Mathematics, 310:63–88, 1973. CODEN LNMAA2. ISBN 3-540-06137-1 (print), 3-540-38338-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060795/.

Iversen:1973:GLS

[Ive73b] Birger Iversen. Generic Local Structure of the Morphisms in Commutative Algebra, volume 310 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06137-1 (print), 3-540-38338-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 108 pp. LCCN QA3 .L28 no. 310.

URL http://link.springer.com/book/10.1007/BFb0060790;
 http://www.springerlink.com/content/978-3-540-38338-
3.

Iversen:1973:I

[Ive73c] Birger Iversen. Introduction. Lecture Notes in Mathematics, 310: 1–2, 1973. CODEN LNMAA2. ISBN 3-540-06137-1 (print), 3-540-38338-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060791.

Iversen:1973:LGG

[Ive73d] Birger Iversen. Leitfaden. (German) [Guide]. Lecture Notes in Mathematics, 310:3, 1973. CODEN LNMAA2. ISBN 3-540-06137-1 (print), 3-540-38338-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060792.

Iversen:1973:SM

[Ive73e] Birger Iversen. Smooth morphisms. Lecture Notes in Mathematics, 310:33–62, 1973. CODEN LNMAA2. ISBN 3-540-06137-1 (print), 3-540-38338-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060794/.

Iversen:1973:SFT

[Ive73f] Birger Iversen. Some fundamental theorems. Lecture Notes in Mathematics, 310:89–102, 1973. CODEN LNMAA2. ISBN 3-540-06137-1 (print), 3-540-38338-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060796/.

Iversen:1973:UM

[Ive73g] Birger Iversen. Unramified morphisms. Lecture Notes in Mathematics, 310:4–32, 1973. CODEN LNMAA2. ISBN 3-540-06137-1 (print), 3-540-38338-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060793/.

Ikeda:1973:LSC

[IW73] Nobuyuki Ikeda and Shinzo Watanabe. The local structure of a class of diffusions and related problems. Lecture Notes in Mathematics, 330:124–169, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061485/.

Iwahori:1970:CIF

[Iwa70] Nagayoshi Iwahori. Centralizers of involutions in finite Chevalley groups. Lecture Notes in Mathematics, 131:267–295, 1970.

CODEN LNMAA2. ISBN 3-540-04920-7 (print), 3-540-36272-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0081547/.

Iwano:1971:AES

[Iwa71a] Masahiro Iwano. Analytic expressions of solutions with the order of positive powers of x near an irregular type singularity. Lecture Notes in Mathematics, 243:134–147, 1971. CODEN LN-MAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058724/.

Iwano:1971:BSS

[Iwa71b] Masahiro Iwano. Bounded solutions and stable domains of nonlinear ordinary differential equations. Lecture Notes in Mathematics, 183:59–127, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060412/.

Jacobs:1970:LFP

[Jac70] Konrad Jacobs. Lipschitz functions and the prevalence of strict ergodicity for continuous-time flows. Lecture Notes in Mathematics, 160:87–124, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060650/.

Jacquet:1972:AFP

[Jac72a] Hervé Jacquet. Automorphic Forms on GL(2) Part II, volume 278 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05931-8 (print), 3-540-37612-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 140 (est.) pp. LCCN QA3 .L28 no. 278. URL http:/

/link.springer.com/book/10.1007/BFb0058503;   http://
www.springerlink.com/content/978-3-540-37612-5.

Jacquet:1972:EVT

[Jac72b] Hervé Jacquet. Espaces vectoriels topologiques et bornologiques généralités. (French) []. Lecture Notes in Mathematics, 277:34– 76, 1972. CODEN LNMAA2. ISBN 3-540-05934-2 (print), 3-540-37613-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058878/.

Jacquet:1972:LTa

[Jac72c] Hervé Jacquet. Local theory for GL(2) × GL(2). Lecture Notes in Mathematics, 278:1–116, 1972. CODEN LNMAA2. ISBN 3-540-05931-8 (print), 3-540-37612-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058504/.

Jacquet:1972:LTb

[Jac72d] Hervé Jacquet. Local theory for GL(2) × GL(2). Lecture Notes in Mathematics, 278:117–140, 1972. CODEN LNMAA2. ISBN 3-540-05931-8 (print), 3-540-37612-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058505/.

James:1970:DFS

[Jam70a] I. M. James. On the decomposability of fibre spaces. Lecture Notes in Mathematics, 168:125–134, 1970. CODEN LN-MAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058522/.

Jameson:1970:LLR

[Jam70b] Graham Jameson. Linear lattices and Riesz spaces. Lecture Notes in Mathematics, 141:40–77, 1970. CODEN LNMAA2. ISBN 3-540-04930-4 (print), 3-540-36300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059133/.

Jameson:1970:LLT

[Jam70c] Graham Jameson. Linear lattices with a topology. Lecture Notes in Mathematics, 141:148–169, 1970. CODEN LNMAA2. ISBN 3-540-04930-4 (print), 3-540-36300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059135/.

Jameson:1970:OA

[Jam70d] Graham Jameson. Ordered algebras. Lecture Notes in Mathematics, 141:170–184, 1970. CODEN LNMAA2. ISBN 3-540-04930-4 (print), 3-540-36300-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059136/.

Jameson:1970:OLSa

[Jam70e] Graham Jameson. Ordered linear spaces. Lecture Notes in Mathematics, 141:1–39, 1970. CODEN LNMAA2. ISBN 3-540-04930-4 (print), 3-540-36300-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059132/.

Jameson:1970:OLSb

[Jam70f] Graham Jameson. Ordered linear spaces, volume 141 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04930-4 (print), 3-540-36300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xv + 194 pp. LCCN QA3 .L28 no. 141. URL http:/

/link.springer.com/book/10.1007/BFb0059130;   http://
www.springerlink.com/content/978-3-540-36300-2.

Jameson:1970:OTL

[Jam70g] Graham Jameson. Ordered topological linear spaces. Lecture Notes in Mathematics, 141:78–147, 1970. CODEN LN-MAA2. ISBN 3-540-04930-4 (print), 3-540-36300-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059134/.

Jameson:1970:P

[Jam70h] Graham Jameson. Preliminaries. Lecture Notes in Mathematics, 141:viii–xv, 1970. CODEN LNMAA2. ISBN 3-540-04930-4 (print), 3-540-36300-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059131.

James:1971:BS

[Jam71] I. M. James. Bibliography on H-spaces. Lecture Notes in Mathematics, 196:137–156, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067756/.

James:1972:SMR

[Jam72a] D. J. G. James. Stability of model-reference systems with random inputs. Lecture Notes in Mathematics, 294:147–159, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064940/.

James:1972:TPS

[Jam72b] Ioan M. James. Two problems studied by Heinz Hopf. Lecture Notes in Mathematics, 279:134–174, 1972. CODEN LN-MAA2. ISBN 3-540-05944-X (print), 3-540-37616-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058511/.

Janssen:1971:MBT

[Jan71] Klaus Janßen. Martin boundary and H^{√}-theory of harmonic spaces. Lecture Notes in Mathematics, 226:102–151, 1971. CO-DEN LNMAA2. ISBN 3-540-05638-6 (print), 3-540-36912-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060482/.

Janin:1974:VOM

[Jan74a] Guy Janin. A variable order method for the numerical integration of the gravitational N-body problem. Lecture Notes in Mathematics, 362:291–303, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066597/.

Janos:1974:ACT

[Jan74b] Ludvik Janos. An application of combinatorial techniques to a topological problem. Lecture Notes in Mathematics, 403:56– 60, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057375/.

Jategaonkar:1970:ITF

[Jat70a] A. V. Jategaonkar. Ideal theory of fully left Goldie Ipli-rings. Lecture Notes in Mathematics, 123:93–122, 1970. CODEN LN-MAA2. ISBN 3-540-04912-6 (print), 3-540-36258-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060433/.

Jategaonkar:1970:LGR

[Jat70b] A. V. Jategaonkar. Left Goldie rings. Lecture Notes in Mathematics, 123:1–33, 1970. CODEN LNMAA2. ISBN 3-540-04912-6 (print), 3-540-36258-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060431/.

Jategaonkar:1970:LPI

[Jat70c] A. V. Jategaonkar. Left principal ideal rings, volume 123 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04912-6 (print), 3-540-36258-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 145 pp. LCCN QA3 .L28 no. 123. URL http:/

/link.springer.com/book/10.1007/BFb0060430;   http://
www.springerlink.com/content/978-3-540-36258-6.

Jategaonkar:1970:PD

[Jat70d] A. V. Jategaonkar. Pli-domains. Lecture Notes in Mathematics, 123:123–136, 1970. CODEN LNMAA2. ISBN 3-540-04912-6 (print), 3-540-36258-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060434/.

Jategaonkar:1970:SPR

[Jat70e] A. V. Jategaonkar. Structure of pli-rings. Lecture Notes in Mathematics, 123:34–92, 1970. CODEN LNMAA2. ISBN 3-540-04912-6 (print), 3-540-36258-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060432/.

Jaworowski:1970:SVF

[Jaw70] Jan W. Jaworowski. Set-valued fixed points theorems for approximative retracts. Lecture Notes in Mathematics, 171:34–39, 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069717/.

Jech:1971:AC

[Jec71a] Thomas J. Jech. Axiom of choice. Lecture Notes in Mathematics, 217:15–16, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061138.

Jech:1971:AR

[Jec71b] Thomas J. Jech. Axiom of regularity. Lecture Notes in Mathematics, 217:19–20, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061140.

Jech:1971:AZF

[Jec71c] Thomas J. Jech. Axioms of Zermelo–Fraenkel. Lecture Notes in Mathematics, 217:2–6, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061133/.

Jech:1971:CA

[Jec71d] Thomas J. Jech. Cardinal arithmetic. Lecture Notes in Mathematics, 217:16–18, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061139/.

Jech:1971:CN

[Jec71e] Thomas J. Jech. Cardinal numbers. Lecture Notes in Mathematics, 217:9–12, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061135/.

Jech:1971:CAG

[Jec71f] Thomas J. Jech. Consistency of AC and GCH. Lecture Notes in Mathematics, 217:35–36, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061143.

Jech:1971:CS

[Jec71g] Thomas J. Jech. Constructible sets. Lecture Notes in Mathematics, 217:32–35, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061142/.

Jech:1971:FS

[Jec71h] Thomas J. Jech. Finite sets. Lecture Notes in Mathematics, 217: 12–13, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061136.

Jech:1971:FC

[Jec71i] Thomas J. Jech. Formulas and classes. Lecture Notes in Mathematics, 217:1–2, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061132.

Jech:1971:ICH

[Jec71j] Thomas J. Jech. Independence of the continuum hypothesis and collapsing of cardinals. Lecture Notes in Mathematics, 217:62– 74, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061148/.

Jech:1971:LM

[Jec71k] Thomas J. Jech. Lebesgue measurability. Lecture Notes in Mathematics, 217:78–90, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061150/.

Jech:1971:LST

[Jec71l] Thomas J. Jech. Lectures in Set Theory with Particular Emphasis on the Method of Forcing, volume 217 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CO-DEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 137 pp. LCCN QA3 .L28 no. 217. URL http:/

/link.springer.com/book/10.1007/BFb0061131;   http://
www.springerlink.com/content/978-3-540-36882-3.

Jech:1971:MFB

[Jec71m] Thomas J. Jech. The method of forcing and Boolean-valued models. Lecture Notes in Mathematics, 217:51–62, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061147/.

Jech:1971:MTM

[Jec71n] Thomas J. Jech. More on transitive models. Lecture Notes in Mathematics, 217:36–41, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061144/.

Jech:1971:OD

[Jec71o] Thomas J. Jech. Ordinal definability. Lecture Notes in Mathematics, 217:41–47, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061145/.

Jech:1971:ON

[Jec71p] Thomas J. Jech. Ordinal numbers. Lecture Notes in Mathematics, 217:6–9, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061134/.

Jech:1971:RN

[Jec71q] Thomas J. Jech. Real numbers. Lecture Notes in Mathematics, 217:13–15, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061137/.

Jech:1971:RCB

[Jec71r] Thomas J. Jech. Remarks on complete Boolean algebras. Lecture Notes in Mathematics, 217:47–51, 1971. CODEN LN-MAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061146/.

Jech:1971:SP

[Jec71s] Thomas J. Jech. Suslin’s problem. Lecture Notes in Mathematics, 217:90–99, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061151/.

Jech:1971:TM

[Jec71t] Thomas J. Jech. Transitive models. Lecture Notes in Mathematics, 217:20–32, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061141/.

Jech:1971:TAB

[Jec71u] Thomas J. Jech. Two applications of Boolean-valued models in the theory of Boolean algebras. Lecture Notes in Mathematics, 217:74–78, 1971. CODEN LNMAA2. ISBN 3-540-05564-9 (print), 3-540-36882-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061149/.

Jensen:1972:ATD

[Jen72a] C. U. Jensen. Applications à la théorie des dimensions homologiques des anneaux et des modules. (French) []. Lecture Notes in Mathematics, 254:38–50, 1972. CODEN LN-MAA2. ISBN 3-540-05737-4 (print), 3-540-37122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058400/.

Jensen:1972:AAM

[Jen72b] C. U. Jensen. Applications aux anneaux et modules complets. (French) []. Lecture Notes in Mathematics, 254:68–77, 1972. CO-DEN LNMAA2. ISBN 3-540-05737-4 (print), 3-540-37122-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058403/.

Jensen:1972:CTP

[Jen72c] C. U. Jensen. Des conditions topologiques pour l’annulation de. (French) []. Lecture Notes in Mathematics, 254:56–67, 1972. CO-DEN LNMAA2. ISBN 3-540-05737-4 (print), 3-540-37122-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058402/.

Jensen:1972:RCD

[Jen72d] C. U. Jensen. Des résultats complémentaires sur la dimension cohomologique d’un ensemble ordonné. (French) []. Lecture Notes in Mathematics, 254:51–55, 1972. CODEN LNMAA2. ISBN 3-540-05737-4 (print), 3-540-37122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058401/.

Jensen:1972:IFF

[Jen72e] C. U. Jensen. Introduction des foncteurs. (French) [Introduction to functors]. Lecture Notes in Mathematics, 254:1–11, 1972. CODEN LNMAA2. ISBN 3-540-05737-4 (print), 3-540-37122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058396/.

Jensen:1972:FDL

[Jen72f] C. U. Jensen. Les Foncteurs Dérivés de \underline{lim} et leurs Applications en Théorie des Modules. (French) [Derivative functors of \underline{lim} and their applications in the theory of modules], volume 254 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05737-4 (print), 3-540-37122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 97 (est.) pp. LCCN QA3 .L28 no. 254. URL http:

//link.springer.com/book/10.1007/BFb0058395;   http:/
/www.springerlink.com/content/978-3-540-37122-9.

Jensen:1972:SSP

[Jen72g] C. U. Jensen. Les suites spectrales pour ????. (French) [Spectral sequences for ????]. Lecture Notes in Mathematics, 254:31– 37, 1972. CODEN LNMAA2. ISBN 3-540-05737-4 (print), 3-540-37122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058399/.

Jensen:1972:REP

[Jen72h] C. U. Jensen. Résultats explicites pour un ensemble d’indices dénombrable. (French) []. Lecture Notes in Mathematics, 254: 12–23, 1972. CODEN LNMAA2. ISBN 3-540-05737-4 (print), 3-540-37122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058397/.

Jensen:1972:DCE

[Jen72i] C. U. Jensen. Sur la dimension cohomologique d’un ensemble de puissance x_{K}. (French) []. Lecture Notes in Mathematics, 254: 24–30, 1972. CODEN LNMAA2. ISBN 3-540-05737-4 (print), 3-540-37122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058398/.

Jensen:1972:APS

[Jen72j] C. U. Jensen. Sur l’annulation de pour des systèmes projectifs de modules de type fini sur un anneau noethérien. (French) []. Lecture Notes in Mathematics, 254:78–97, 1972. CODEN LNMAA2. ISBN 3-540-05737-4 (print), 3-540-37122-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058404/.

Jussila:1971:FIS

[JG71] O. Jussila and A. Grothendieck. Formalisme des intersections sur les schémas algébriques propres. (French) []. Lecture Notes in Mathematics, 225:519–594, 1971. CODEN LNMAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066294/.

Jacquet:1970:AF

[JL70a] H. Jacquet and R. P. Langlands. Automorphic Forms on GL(2), volume 114 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04903-7 (print), 3-540-36234-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 548 pp. LCCN QA3 .L28 no. 114. URL http:/

/link.springer.com/book/10.1007/BFb0058988;   http://
www.springerlink.com/content/978-3-540-36234-0.

Jacquet:1970:GT

[JL70b] H. Jacquet and R. P. Langlands. Global theory. Lecture Notes in Mathematics, 114:301–416, 1970. CODEN LNMAA2. ISBN 3-540-04903-7 (print), 3-540-36234-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058990/.

Jacquet:1970:LT

[JL70c] H. Jacquet and R. P. Langlands. Local theory. Lecture Notes in Mathematics, 114:1–300, 1970. CODEN LNMAA2. ISBN 3-540-04903-7 (print), 3-540-36234-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058989/.

Jacquet:1970:QA

[JL70d] H. Jacquet and R. P. Langlands. Quaternion algebras. Lecture Notes in Mathematics, 114:417–548, 1970. CODEN LN-MAA2. ISBN 3-540-04903-7 (print), 3-540-36234-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058991/.

Jones:1971:PCI

[JL71] C. B. Jones and P. Lucas. Proving correctness of implementation techniques. Lecture Notes in Mathematics, 188:178–211, 1971. CODEN LNMAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059698/.

Jurkat:1973:ESM

[JLP73] W. Jurkat, D. Lutz, and A. Peyerimhoff. Effective solutions for meromorphic second order differential equations. Lecture Notes in Mathematics, 312:100–107, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060049/.

Jonsson:1972:A

[Jón72a] Bjarni Jónsson. Addenda. Lecture Notes in Mathematics, 250: 213–214, 1972. CODEN LNMAA2. ISBN 3-540-05722-6 (print), 3-540-37096-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058653.

Jonsson:1972:ARS

[Jón72b] Bjarni Jónsson. Algebras and relational structures. Lecture Notes in Mathematics, 250:16–65, 1972. CODEN LNMAA2. ISBN 3-540-05722-6 (print), 3-540-37096-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058650/.

Jonsson:1972:CR

[Jón72c] Bjarni Jónsson. Congruence relations. Lecture Notes in Mathematics, 250:148–212, 1972. CODEN LNMAA2. ISBN 3-540-05722-6 (print), 3-540-37096-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058652/.

Jonsson:1972:P

[Jón72d] Bjarni Jónsson. Preliminaries. Lecture Notes in Mathematics, 250:1–15, 1972. CODEN LNMAA2. ISBN 3-540-05722-6 (print), 3-540-37096-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058649/.

Jonsson:1972:S

[Jón72e] Bjarni Jónsson. Substructures. Lecture Notes in Mathematics, 250:66–147, 1972. CODEN LNMAA2. ISBN 3-540-05722-6 (print), 3-540-37096-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058651/.

Jonsson:1972:TUA

[Jón72f] Bjarni Jónsson. Topics in Universal Algebra, volume 250 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05722-6 (print), 3-540-37096-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 220 pp. LCCN QA3 .L28 no. 250. URL http:/

/link.springer.com/book/10.1007/BFb0058648;   http://
www.springerlink.com/content/978-3-540-37096-3.

Jones:1974:HSN

[Jon74] F. Burton Jones. Hereditarily separable, non-completely regular spaces. Lecture Notes in Mathematics, 375:149–152, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064021/.

Jorgens:1972:PDO

[Jör72] K. Jörgens. Perturbations of the Dirac operator. Lecture Notes in Mathematics, 280:87–102, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066921/.

Joseph:1973:RAB

[Jos73] Daniel D. Joseph. Remarks about bifurcation and stability of quasi-periodic solutions which bifurcate from periodic solutions of the Navier Stokes equations. Lecture Notes in Mathematics, 322: 130–158, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060565/.

Josefson:1974:CLP

[Jos74] Bengt Josefson. A counterexample in the Levi problem. Lecture Notes in Mathematics, 364:168–177, 1974. CODEN LN-MAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069016/.

Jouanolou:1973:QCT

[Jou73a] J. P. Jouanolou. Quelques calculs en K-théorie des schémas. (French) []. Lecture Notes in Mathematics, 341:317–335, 1973. CODEN LNMAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067064/.

Jouanolou:1973:SEM

[Jou73b] J. P. Jouanolou. Une suite exacte de Mayer–Vietoris en K-théorie algébrique. (French) []. Lecture Notes in Mathematics, 341:293– 316, 1973. CODEN LNMAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067063/.

Jorgens:1973:ESH

[JW73a] Konrad Jörgens and Joachim Weidmann. The essential spectrum of the Hamiltonian of an N-particle system with external forces. Lecture Notes in Mathematics, 313:87–102, 1973. CODEN LNMAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060832/.

Jorgens:1973:ESI

[JW73b] Konrad Jörgens and Joachim Weidmann. The essential spectrum of the internal Hamiltonian of a free system. Lecture Notes in Mathematics, 313:103–118, 1973. CODEN LNMAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060833/.

Jorgens:1973:E

[JW73c] Konrad Jörgens and Joachim Weidmann. Examples. Lecture Notes in Mathematics, 313:32–38, 1973. CODEN LN-MAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060826/.

Jorgens:1973:I

[JW73d] Konrad Jörgens and Joachim Weidmann. Introduction. Lecture Notes in Mathematics, 313:1–6, 1973. CODEN LNMAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060822/.

Jorgens:1973:LBE

[JW73e] Konrad Jörgens and Joachim Weidmann. A lower bound of the essential spectrum. Lecture Notes in Mathematics, 313:78– 86, 1973. CODEN LNMAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060831/.

Jorgens:1973:PH

[JW73f] Konrad Jörgens and Joachim Weidmann. N-particle Hamiltonians. Lecture Notes in Mathematics, 313:45–54, 1973. CODEN LNMAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060828/.

Jorgens:1973:OAO

[JW73g] Konrad Jörgens and Joachim Weidmann. Operators acting only on part of the variables. Lecture Notes in Mathematics, 313: 39–44, 1973. CODEN LNMAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060827/.

Jorgens:1973:PSI

[JW73h] Konrad Jörgens and Joachim Weidmann. Perturbations small at infinity. Lecture Notes in Mathematics, 313:19–31, 1973. CODEN LNMAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060825/.

Jorgens:1973:PT

[JW73i] Konrad Jörgens and Joachim Weidmann. Proof of theorem 11.16. Lecture Notes in Mathematics, 313:119–132, 1973. CODEN LN-MAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060834/.

Jorgens:1973:SO

[JW73j] Konrad Jörgens and Joachim Weidmann. Schrödinger operators. Lecture Notes in Mathematics, 313:12–18, 1973. CODEN LNMAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060824/.

Jorgens:1973:SES

[JW73k] Konrad Jörgens and Joachim Weidmann. Spectra and essential spectra of selfadjoint and essentially selfadjoint operators. Lecture Notes in Mathematics, 313:7–11, 1973. CODEN LNMAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060823/.

Jorgens:1973:SPH

[JW73l] Konrad Jörgens and Joachim Weidmann. Spectral Properties of Hamiltonian Operators, volume 313 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 140 pp. LCCN QA3 .L28 no. 313. URL http:/

/link.springer.com/book/10.1007/BFb0060821;   http://
www.springerlink.com/content/978-3-540-38354-3.

Jorgens:1973:SHF

[JW73m] Konrad Jörgens and Joachim Weidmann. The spectrum of the Hamiltonian of a free system. Lecture Notes in Mathematics, 313: 66–77, 1973. CODEN LNMAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060830/.

Jorgens:1973:SH

[JW73n] Konrad Jörgens and Joachim Weidmann. Symmetries of the Hamiltonian. Lecture Notes in Mathematics, 313:55–65, 1973. CODEN LNMAA2. ISBN 3-540-06151-7 (print), 3-540-38354-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060829/.

Kadison:1970:SAM

[Kad70] Richard V. Kadison. Some analytic methods in the theory of operator algebras. Lecture Notes in Mathematics, 140:8–29, 1970. CODEN LNMAA2. ISBN 3-540-04929-0 (print), 3-540-36298-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0100064/.

Kahane:1973:IPF

[Kah73] Jean-Pierre Kahane. Idéaux primaires fermés dans certaines algèbres de Banach de fonctions analytiques. (French) []. Lecture Notes in Mathematics, 336:5–14, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065782/.

Kainen:1974:SRR

[Kai74] Paul C. Kainen. Some recent results in topological graph theory. Lecture Notes in Mathematics, 406:76–108, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066436/.

Kakutani:1973:EEM

[Kak73] Shizuo Kakutani. Examples of ergodic measure preserving transformations which are weakly mixing but not strongly mixing. Lecture Notes in Mathematics, 318:143–149, 1973. CODEN LN-MAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061731/.

Kantor:1970:HCF

[Kan70] J.-M. Kantor. Hyperfonctions cohérentes. (French) []. Lecture Notes in Mathematics, 116:75–100, 1970. CODEN LN-MAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059247/.

Kaneyuki:1971:AAG

[Kan71a] Soji Kaneyuki. The affine automorphism groups of Siegel domains. Lecture Notes in Mathematics, 241:1–11, 1971. CODEN LNMAA2. ISBN 3-540-05702-1 (print), 3-540-37028-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060968/.

Kaneyuki:1971:CTU

[Kan71b] Soji Kaneyuki. Cayley transforms of universal domains. Lecture Notes in Mathematics, 241:52–63, 1971. CODEN LN-MAA2. ISBN 3-540-05702-1 (print), 3-540-37028-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060974/.

Kaneyuki:1971:GBI

[Kan71c] Soji Kaneyuki. A generalization of Borel imbeddings. Lecture Notes in Mathematics, 241:42–51, 1971. CODEN LN-MAA2. ISBN 3-540-05702-1 (print), 3-540-37028-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060973/.

Kaneyuki:1971:HBD

[Kan71d] Soji Kaneyuki. Homogeneous Bounded Domains and Siegel Domains, volume 241 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05702-1 (print), 3-540-37028-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 89 pp. LCCN QA3 .L28 no. 241. URL http:/

/link.springer.com/book/10.1007/BFb0060967;   http://
www.springerlink.com/content/978-3-540-37028-4.

Kaneyuki:1971:IS

[Kan71e] Soji Kaneyuki. The Iwasawa subgroups. Lecture Notes in Mathematics, 241:12–17, 1971. CODEN LNMAA2. ISBN 3-540-05702-1 (print), 3-540-37028-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060969/.

Kaneyuki:1971:A

[Kan71f] Soji Kaneyuki. j-algebras. Lecture Notes in Mathematics, 241: 18–25, 1971. CODEN LNMAA2. ISBN 3-540-05702-1 (print), 3-540-37028-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060970/.

Kaneyuki:1971:RSD

[Kan71g] Soji Kaneyuki. Realizations as Siegel domains of type III. Lecture Notes in Mathematics, 241:71–82, 1971. CODEN LN-MAA2. ISBN 3-540-05702-1 (print), 3-540-37028-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060976/.

Kaneyuki:1971:RHC

[Kan71h] Soji Kaneyuki. Remarks on Harish-Chandra imbeddings. Lecture Notes in Mathematics, 241:64–70, 1971. CODEN LN-MAA2. ISBN 3-540-05702-1 (print), 3-540-37028-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060975/.

Kaneyuki:1971:UD

[Kan71i] Soji Kaneyuki. Universal domains. Lecture Notes in Mathematics, 241:35–41, 1971. CODEN LNMAA2. ISBN 3-540-05702-1 (print), 3-540-37028-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060972/.

Kaneyuki:1971:UA

[Kan71j] Soji Kaneyuki. Universal j-algebras. Lecture Notes in Mathematics, 241:26–34, 1971. CODEN LNMAA2. ISBN 3-540-05702-1 (print), 3-540-37028-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060971/.

Kantor:1971:ITR

[Kan71k] J.-M. Kantor. Introduction à la théorie des résidus. (French) []. Lecture Notes in Mathematics, 205:1–10, 1971. CODEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059829/.

Kantor:1972:ODE

[Kan72] J.-M. Kantor. Opérateurs différentiels sur les espaces analytiques. (French) []. Lecture Notes in Mathematics, 275:147–157, 1972. CODEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058498/.

Kanda:1973:RMP

[Kan73a] Mamoru Kanda. Remarks on Markov processes having Green functions with isotropic singularity. Lecture Notes in Mathematics, 330:170–175, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061486/.

Kaneko:1973:FPE

[Kan73b] Akira Kaneko. Fundamental principle and extension of solutions of linear differential equations with constant coefficients. Lecture Notes in Mathematics, 287:122–134, 1973. CODEN LN-MAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068148/.

Kantor:1974:ODS

[Kan74] Jean-Marie Kantor. Opérateurs différentiels sur les singularités quotients. (French) []. Lecture Notes in Mathematics, 409:335– 340, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068108/.

Kaplan:1970:ASD

[Kap70] James L. Kaplan. Absolutely stable differential equations which are not integrally stable. Lecture Notes in Mathematics, 144: 115–120, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059926/.

Kaplansky:1973:CR

[Kap73] Irving Kaplansky. Commutative rings. Lecture Notes in Mathematics, 311:153–166, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068926/.

Karoubi:1970:ACO

[Kar70a] Max Karoubi. Algèbres de Clifford et opérateurs de Fredholm. (French) []. Lecture Notes in Mathematics, 136:66–106, 1970. CODEN LNMAA2. ISBN 3-540-04925-8 (print), 3-540-36289-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059022/.

Karoubi:1970:FDT

[Kar70b] Max Karoubi. Foncteurs dérivés et K-théorie. (French) []. Lecture Notes in Mathematics, 136:107–186, 1970. CODEN LN-MAA2. ISBN 3-540-04925-8 (print), 3-540-36289-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059023/.

Karoubi:1970:TEF

[Kar70c] Max Karoubi. Sur la K-théorie équivariante. (French) [On equivariant K-theory]. Lecture Notes in Mathematics, 136:187–253, 1970. CODEN LNMAA2. ISBN 3-540-04925-8 (print), 3-540-36289-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059024/.

Karoubi:1973:CGF

[Kar73a] Max Karoubi. Cobordisme et groupes formels. (French) []. Lecture Notes in Mathematics, 317:141–165, 1973. CODEN LN-MAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069281/.

Karoubi:1973:PTH

[Kar73b] Max Karoubi. Périodicité de la K-théorie hermitienne. (French) []. Lecture Notes in Mathematics, 343:301–411, 1973. CODEN LNMAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061372/.

Karoubi:1973:SPC

[Kar73c] Max Karoubi. Some problems and conjectures in algebraic Ktheory. Lecture Notes in Mathematics, 343:52–56, 1973. CODEN LNMAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061369/.

Kato:1970:ASC

[Kat70a] Junji Kato. Absolute stability of control systems with multiple feedback. Lecture Notes in Mathematics, 144:128–133, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059929/.

Kato:1970:LFD

[Kat70b] Junji Kato. On linear functional differential equations. Lecture Notes in Mathematics, 144:121–123, 1970. CODEN LN-MAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059927/.

Kato:1970:UAS

[Kat70c] Junji Kato. Uniformly asymptotic stability and total stability. Lecture Notes in Mathematics, 144:124–127, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059928/.

Kato:1971:LRT

[Kat71] Junji Kato. On Liapunov–Razumikhin type theorems. Lecture Notes in Mathematics, 243:54–65, 1971. CODEN LN-MAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058718/.

Katz:1973:ECP

[Kat73a] N. Katz. Étude cohomologique des pinceaux de Lefschetz. (French) []. Lecture Notes in Mathematics, 340:254–327, 1973.

CODEN LNMAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060514/.

Katz:1973:NCI

[Kat73b] N. Katz. Le niveau de la cohomologie des intersections complétes. (French) []. Lecture Notes in Mathematics, 340:363–400, 1973. CODEN LNMAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060517/.

Katz:1973:TGF

[Kat73c] N. Katz. Le théorème de Griffiths. (French) []. Lecture Notes in Mathematics, 340:341–362, 1973. CODEN LNMAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060516/.

Katz:1973:FCP

[Kat73d] N. Katz. Une formule de congruence pour la fonction ζ. (French) []. Lecture Notes in Mathematics, 340:401–438, 1973. CODEN LNMAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060518/.

Katz:1973:PLT

[Kat73e] N. M. Katz. Pinceaux de Lefschetz: Théorème d’existence. (French) []. Lecture Notes in Mathematics, 340:212–253, 1973. CODEN LNMAA2. ISBN 3-540-06433-8 (print), 3-540-37764-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060513/.

Katz:1973:TDF

[Kat73f] Nicholas Katz. Travaux de Dwork. (French) []. Lecture Notes in Mathematics, 317:167–200, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069282/.

Katz:1973:APM

[Kat73g] Nicholas M. Katz. p-Adic properties of modular schemes and modular forms. Lecture Notes in Mathematics, 350:69–190, 1973. CODEN LNMAA2. ISBN 3-540-06483-4 (print), 3-540-37802-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-37802-0_3/.

Katznelson:1973:SAE

[Kat73h] Yitzhak Katznelson. Suites aléatoires d’entiers. (French) []. Lecture Notes in Mathematics, 336:148–152, 1973. CODEN LN-MAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065796/.

Kaufman:1972:HDL

[Kau72] Robert Kaufman. Hausdorff dimension, lacunary series, and some problems on exceptional sets. Lecture Notes in Mathematics, 266: 189–195, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059644/.

Kauffman:1974:ILC

[Kau74] Louis H. Kauffman. An invariant of link concordance. Lecture Notes in Mathematics, 375:153–157, 1974. CODEN LN-MAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064022/.

Kawakubo:1972:TAS

[Kaw72a] K. Kawakubo. Topological s^{1} and Z_{2k} actions on spheres. Lecture Notes in Mathematics, 298:14–18, 1972. CODEN LN-MAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070031/.

Kawakubo:1972:ISF

[Kaw72b] Katsuo Kawakubo. Invariants for semi-free s^{1}-actions. Lecture Notes in Mathematics, 298:1–13, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070030/.

Kawai:1973:GER

[Kaw73a] Takahiro Kawai. On the global existence of real analytic solutions of linear differential equations. Lecture Notes in Mathematics, 287:99–121, 1973. CODEN LNMAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068147/.

Kawata:1973:CLP

[Kaw73b] Tatsuo Kawata. On a class of linear processes. Lecture Notes in Mathematics, 330:193–212, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061488/.

Kazarinoff:1971:IBV

[Kaz71] N. D. Kazarinoff. Initial-boundary value problems and moving boundaries. Lecture Notes in Mathematics, 183:195–199, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060419/.

Kazamaki:1972:ELM

[Kaz72a] N. Kazamaki. Examples on local martingales. Lecture Notes in Mathematics, 258:98–100, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059464/.

Kazamaki:1972:KDL

[Kaz72b] N. Kazamaki. Krickeberg’s decomposition for local martingales. Lecture Notes in Mathematics, 258:101–104, 1972. CODEN LN-MAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059465/.

Kazamaki:1972:NSI

[Kaz72c] N. Kazamaki. Note on a stochastic integral equation. Lecture Notes in Mathematics, 258:105–108, 1972. CODEN LN-MAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059466/.

Kazamaki:1973:NMF

[Kaz73] N. Kazamaki. Une note sur les martingales faibles. (French) []. Lecture Notes in Mathematics, 321:118–121, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071400/.

Keesling:1970:CRP

[Kee70] James Keesling. Compactness related properties in hyperspaces. Lecture Notes in Mathematics, 171:40–44, 1970. CODEN LN-MAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069718/.

Keesling:1974:MLC

[Kee74] James Keesling. On movability and local connectivity. Lecture Notes in Mathematics, 375:158–167, 1974. CODEN LN-MAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064023/.

Keimel:1971:RLO

[Kei71] Klaus Keimel. The representation of lattice-ordered groups and rings by sections in sheaves. Lecture Notes in Mathematics, 248: 1–98, 1971. CODEN LNMAA2. ISBN 3-540-05714-5 (print), 3-540-37081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058563/.

Keisler:1974:MCF

[Kei74] H. Jerome Keisler. Monotone complete fields. Lecture Notes in Mathematics, 369:113–115, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066005/.

Kelly:1972:AAC

[Kel72a] G. M. Kelly. An abstract approach to coherence. Lecture Notes in Mathematics, 281:106–147, 1972. CODEN LNMAA2. ISBN 3-540-05963-6 (print), 3-540-37958-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059557/.

Kelly:1972:CET

[Kel72b] G. M. Kelly. A cut-elimination theorem. Lecture Notes in Mathematics, 281:196–213, 1972. CODEN LNMAA2. ISBN 3-540-05963-6 (print), 3-540-37958-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059559/.

Kelly:1972:MVF

[Kel72c] G. M. Kelly. Many-variable functorial calculus. I. Lecture Notes in Mathematics, 281:66–105, 1972. CODEN LNMAA2. ISBN 3-540-05963-6 (print), 3-540-37958-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059556/.

Kelemen:1974:QMQ

[Kel74a] Peter J. Kelemen. Quantum mechanics, quantum field theory and hyper-quantum mechanics. Lecture Notes in Mathematics, 369:116–121, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066006/.

Keller:1974:CDF

[Kel74b] Prof. Dr. Hans Heinrich Keller. Continuously differentiable functions. Lecture Notes in Mathematics, 417:59–81, 1974. CODEN LNMAA2. ISBN 3-540-06962-3 (print), 3-540-37267-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070566/.

Keller:1974:DCL

[Kel74c] Prof. Dr. Hans Heinrich Keller. Differential Calculus in Locally Convex Spaces, volume 417 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06962-3 (print), 3-540-37267-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 143 pp. LCCN QA3 .L28 no. 417.

URL http://link.springer.com/book/10.1007/BFb0070564;
 http://www.springerlink.com/content/978-3-540-37267-
7.

Keller:1974:FC

[Kel74d] Prof. Dr. Hans Heinrich Keller. Functions of class C^{p}. Lecture Notes in Mathematics, 417:83–113, 1974. CODEN LN-MAA2. ISBN 3-540-06962-3 (print), 3-540-37267-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070567/.

Keller:1974:SMM

[Kel74e] Prof. Dr. Hans Heinrich Keller. Spaces of multilinear mappings. Lecture Notes in Mathematics, 417:13–58, 1974. CODEN LN-MAA2. ISBN 3-540-06962-3 (print), 3-540-37267-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070565/.

Kelly:1974:CTL

[Kel74f] G. M. Kelly. Coherence theorems for Lax algebras and for distributive laws. Lecture Notes in Mathematics, 420:281–375, 1974. CODEN LNMAA2. ISBN 3-540-06966-6 (print), 3-540-37270-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063106/.

Kelly:1974:DA

[Kel74g] G. M. Kelly. Doctrinal adjunction. Lecture Notes in Mathematics, 420:257–280, 1974. CODEN LNMAA2. ISBN 3-540-06966-6 (print), 3-540-37270-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063105/.

Kelly:1974:CD

[Kel74h] G. M. Kelly. On clubs and doctrines. Lecture Notes in Mathematics, 420:181–256, 1974. CODEN LNMAA2. ISBN 3-540-06966-6 (print), 3-540-37270-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063104/.

Kelly:1974:CSP

[Kel74i] Prof. Dr. Gregory M. Kelly, editor. Category Seminar: Proceedings Sydney Category Theory Seminar 1972/1973, volume 420 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06966-6 (print), 3-540-37270-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 420. URL http:/

/link.springer.com/book/10.1007/BFb0063096;   http://
www.springerlink.com/content/978-3-540-37270-7.

Kent:1973:LCA

[Ken73] Joseph F. Kent. Locally connected almost periodic minimal sets. Lecture Notes in Mathematics, 318:150–166, 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061732/.

Kerber:1971:ART

[Ker71a] Adalbert Kerber. Application to the representation theory of symmetric and alternating groups. Lecture Notes in Mathematics, 240:114–181, 1971. CODEN LNMAA2. ISBN 3-540-05693-9 (print), 3-540-37004-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067947/.

Kerber:1971:I

[Ker71b] Adalbert Kerber. Introduction. Lecture Notes in Mathematics, 240:1–3, 1971. CODEN LNMAA2. ISBN 3-540-05693-9 (print), 3-540-37004-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067944/.

Kerber:1971:RPG

[Ker71c] Adalbert Kerber. Representations of Permutation Groups I, volume 240 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05693-9 (print), 3-540-37004-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 192 pp. LCCN QA3 .L28 no. 240. URL http:

//link.springer.com/book/10.1007/BFb0067943;   http:/
/www.springerlink.com/content/978-3-540-37004-8.

Kerber:1971:RWP

[Ker71d] Adalbert Kerber. Representations of wreath products. Lecture Notes in Mathematics, 240:59–113, 1971. CODEN LN-MAA2. ISBN 3-540-05693-9 (print), 3-540-37004-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067946/.

Kerber:1971:WPG

[Ker71e] Adalbert Kerber. Wreath products of groups. Lecture Notes in Mathematics, 240:4–58, 1971. CODEN LNMAA2. ISBN 3-540-05693-9 (print), 3-540-37004-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067945/.

Kervaire:1971:KCC

[Ker71f] Michel A. Kervaire. Knot Cobordism in codimension two. Lecture Notes in Mathematics, 197:83–105, 1971. CODEN LN-MAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068613/.

Kerzman:1973:REE

[Ker73] Norberto Kerzman. Remarks on estimates for the ¯∂ equation. Lecture Notes in Mathematics, 336:111–124, 1973. CODEN LN-MAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065792/.

Keune:1973:DFA

[Keu73] Frans Keune. Derived functors and algebraic K-theory. Lecture Notes in Mathematics, 341:166–176, 1973. CODEN LN-MAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067055/.

Khalili-Francon:1973:PGW

[KF73] E. Khalili-Françon. Processus de Galton–Watson. (French) []. Lecture Notes in Mathematics, 321:122–135, 1973. CODEN LN-MAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071401/.

Karoubi:1970:SHS

[KGLZ70] M. Karoubi, R. Gordon, P. Löffler, and M. Zisman. Séminaire Heidelberg–Saarbrücken–Strasbourg sur la K-Théorie, volume 136 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04925-8 (print), 3-540-36289-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 264 (est.) pp. LCCN QA3 .L28 no. 136. URL http:/

/link.springer.com/book/10.1007/BFb0059019;   http://
www.springerlink.com/content/978-3-540-36289-0.

Khabbaz:1971:GAS

[Kha71] S. A. Khabbaz. Geometric aspects of the singular solutions of certain differential equations. Lecture Notes in Mathematics, 209: 58–76, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068892/.

Kikuchi:1971:CE

[Kik71] Norio Kikuchi. On contingent equations. Lecture Notes in Mathematics, 243:169–181, 1971. CODEN LNMAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058726/.

Kimura:1971:RPR

[Kim71] Tosihusa Kimura. On the Riemann problem on Riemann surfaces. Lecture Notes in Mathematics, 243:219–228, 1971. CODEN LNMAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058730/.

Kimura:1973:MSD

[Kim73] Tosihusa Kimura. On meromorphic solutions of the difference equation y(x + 1) = y(x) + 1 + λ/y(x). Lecture Notes in Mathematics, 312:74–86, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060047/.

Kinoshita:1972:ICA

[Kin72] Shin’ichi Kinoshita. On infinite cyclic actions on contractible open 3-manifolds and strong irreducibility. Lecture Notes in Mathematics, 299:323–327, 1972. CODEN LNMAA2. ISBN 3-540-06078-2 (print), 3-540-38066-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066768/.

King:1974:EFE

[Kin74] J. Thomas King. Extrapolation in the finite element method with penalty. Lecture Notes in Mathematics, 430:261–270, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066273/.

Ku:1972:CIF

[KK72] Hsu-Tung Ku and Mei-Chin Ku. Characteristic invariants of free differentiable actions of S^{1} and S^{3} on homotopy spheres. Lecture Notes in Mathematics, 298:19–40, 1972. CODEN LN-MAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070032/.

Kawada:1973:VGP

[KK73] Takayuki Kawada and Norio Kôno. On the variation of Gaussian processes. Lecture Notes in Mathematics, 330:176–192, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061487/.

Kempf:1973:EET

[KKMSD73a] G. Kempf, F. Knudsen, D. Mumford, and B. Saint-Donat. Equivariant embeddings of tori. Lecture Notes in Mathematics, 339: 1–52, 1973. CODEN LNMAA2. ISBN 3-540-06432-X (print), 3-540-37755-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070319/.

Kempf:1973:SSR

[KKMSD73b] G. Kempf, F. Knudsen, D. Mumford, and B. Saint-Donat. Semistable reduction. Lecture Notes in Mathematics, 339:53–108, 1973. CODEN LNMAA2. ISBN 3-540-06432-X (print), 3-540-37755-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070320/.

Kempf:1973:TE

[KKMSD73c] G. Kempf, Finn Faye Knudsen, D. Mumford, and B. Saint-Donat. Toroidal Embeddings I, volume 339 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06432-X (print), 3-540-37755-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 209 pp. LCCN QA3 .L28 no. 339. URL http:/

/link.springer.com/book/10.1007/BFb0070318;   http://
www.springerlink.com/content/978-3-540-37755-9.

Klasa:1973:SG

[Kla73] Stan Klasa. On Steinberg groups. Lecture Notes in Mathematics, 353:131–138, 1973. CODEN LNMAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059265/.

Kleiman:1971:TFP

[Kle71] S. Kleiman. Les théorèmes de finitude pour le foncteur de Picard. (French) []. Lecture Notes in Mathematics, 225:616–666, 1971. CODEN LNMAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066296/.

Klee:1972:WGP

[Kle72] Victor Klee. Which generalized prisms admit H-circuits? Lecture Notes in Mathematics, 303:173–178, 1972. CODEN LN-MAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067368/.

Kleene:1973:RRS

[Kle73a] S. C. Kleene. Realizability: a retrospective survey. Lecture Notes in Mathematics, 337:95–112, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066772/.

Kleinberg:1973:IC

[Kle73b] E. M. Kleinberg. Infinitary combinatorics. Lecture Notes in Mathematics, 337:361–418, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066782/.

Klimko:1970:WRC

[Kli70] Eugene M. Klimko. Weak ratio convergence of measures in infinite measure spaces. Lecture Notes in Mathematics, 160:125– 132, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060651/.

Kelly:1972:CC

[KLLM72] G. M. Kelly, M. Laplaza, G. Lewis, and Saunders Mac Lane. Coherence in Categories, volume 281 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05963-6 (print), 3-540-37958-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 235 (est.) pp. LCCN QA3 .L28 no. 281.

URL http://link.springer.com/book/10.1007/BFb0059553;
 http://www.springerlink.com/content/978-3-540-37958-
4.

Kelly:1972:CCN

[KM72] G. M. Kelly and Saunders Mac Lane. Closed coherence for a natural transformation. Lecture Notes in Mathematics, 281:1– 28, 1972. CODEN LNMAA2. ISBN 3-540-05963-6 (print), 3-540-37958-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059554/.

Kock:1974:TTF

[KM74] Anders Kock and Chr. Juul Mikkelsen. Topos-theoretic factorization of non-standard extensions. Lecture Notes in Mathematics, 369:122–143, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066007/.

Ku:1972:PSCa

[KMSS72a] H. T. Ku, L. N. Mann, J. L. Sicks, and J. C. Su, editors. Proceedings of the Second Conference on Compact Transformation Groups: University of Massachusetts, Amherst, 1971, volume 298 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 298. URL http:/

/link.springer.com/book/10.1007/BFb0070029;   http://
www.springerlink.com/content/978-3-540-38063-4.

Ku:1972:PSCb

[KMSS72b] H. T. Ku, L. N. Mann, J. L. Sicks, and J. C. Su, editors. Proceedings of the Second Conference on Compact Transformation Groups: University of Massachusetts, Amherst, 1971, volume 299 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06078-2 (print), 3-540-38066-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 299. URL http:/

/link.springer.com/book/10.1007/BFb0066759;   http://
www.springerlink.com/content/978-3-540-38066-5.

Kambayashi:1974:UAG

[KMT74a] Prof. Dr. Tatsuji Kambayashi, Prof. Dr. Masayoshi Miyanishi, and Prof. Dr. Mitsuhiro Takeuchi. Unipotent Algebraic Groups, volume 414 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06960-7 (print), 3-540-37265-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 165 pp. LCCN QA3 .L28 no. 414. URL http:

//link.springer.com/book/10.1007/BFb0070517;   http:/
/www.springerlink.com/content/978-3-540-37265-3.

Kambayashi:1974:AUG

[KMT74b] Tatsuji Kambayashi, Masayoshi Miyanishi, and Mitsuhiro Takeuchi. Actions of unipotent group schemes. Lecture Notes in Mathematics, 414:108–130, 1974. CODEN LNMAA2. ISBN 3-540-06960-7 (print), 3-540-37265-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070525/.

Kambayashi:1974:DTC

[KMT74c] Tatsuji Kambayashi, Masayoshi Miyanishi, and Mitsuhiro Takeuchi. Decomposition theorems for central extensions of com̄ utative group schemes; application to the two-dimensional unipotent groups. Lecture Notes in Mathematics, 414:29–39, 1974. CODEN LNMAA2. ISBN 3-540-06960-7 (print), 3-540-37265-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070521/.

Kambayashi:1974:FAL

[KMT74d] Tatsuji Kambayashi, Masayoshi Miyanishi, and Mitsuhiro Takeuchi. Forms of the affine line and geometry of the groups of Russell type. Lecture Notes in Mathematics, 414:58–107, 1974. CODEN LNMAA2. ISBN 3-540-06960-7 (print), 3-540-37265-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070524/.

Kambayashi:1974:FVG

[KMT74e] Tatsuji Kambayashi, Masayoshi Miyanishi, and Mitsuhiro Takeuchi. Forms of vector groups; groups of Russell type. Lecture Notes in Mathematics, 414:17–28, 1974. CODEN LN-MAA2. ISBN 3-540-06960-7 (print), 3-540-37265-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070520/.

Kambayashi:1974:HUG

[KMT74f] Tatsuji Kambayashi, Masayoshi Miyanishi, and Mitsuhiro Takeuchi. The hyperalgebra of a unipotent group scheme. Lecture Notes in Mathematics, 414:141–145, 1974. CODEN LN-MAA2. ISBN 3-540-06960-7 (print), 3-540-37265-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070527/.

Kambayashi:1974:NCS

[KMT74g] Tatsuji Kambayashi, Masayoshi Miyanishi, and Mitsuhiro Takeuchi. Notations, conventions and some basic preliminary facts. Lecture Notes in Mathematics, 414:5–16, 1974. CODEN LNMAA2. ISBN 3-540-06960-7 (print), 3-540-37265-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070519/.

Kambayashi:1974:TUA

[KMT74h] Tatsuji Kambayashi, Masayoshi Miyanishi, and Mitsuhiro Takeuchi. On the theory of unipotent algebraic groups over an arbitrary ground field. Lecture Notes in Mathematics, 414: 1–4, 1974. CODEN LNMAA2. ISBN 3-540-06960-7 (print), 3-540-37265-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070518/.

Kambayashi:1974:QCT

[KMT74i] Tatsuji Kambayashi, Masayoshi Miyanishi, and Mitsuhiro Takeuchi. The question of commutativity for two-dimensional unipotent groups. Lecture Notes in Mathematics, 414:46–57, 1974. CODEN LNMAA2. ISBN 3-540-06960-7 (print), 3-540-37265-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070523/.

Kambayashi:1974:USU

[KMT74j] Tatsuji Kambayashi, Masayoshi Miyanishi, and Mitsuhiro Takeuchi. The underlying scheme of a unipotent algebraic group. Lecture Notes in Mathematics, 414:131–140, 1974. CODEN LN-MAA2. ISBN 3-540-06960-7 (print), 3-540-37265-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070526/.

Kambayashi:1974:WUG

[KMT74k] Tatsuji Kambayashi, Masayoshi Miyanishi, and Mitsuhiro Takeuchi. Wound unipotent groups. Lecture Notes in Mathematics, 414:40–45, 1974. CODEN LNMAA2. ISBN 3-540-06960-7 (print), 3-540-37265-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070522/.

Keynes:1973:CTE

[KN73] Harvey B. Keynes and Dan Newton. Choquet theory and ergodic measures for compact group extensions. Lecture Notes in Mathematics, 318:167–173, 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061733/.

Knight:1971:RCS

[Kni71] Frank B. Knight. A reduction of continuous square-integrable martingales to Brownian motion. Lecture Notes in Mathematics, 190:19–31, 1971. CODEN LNMAA2. ISBN 3-540-05396-4 (print), 3-540-36515-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065888/.

Knight:1974:ESO

[Kni74] Frank B. Knight. Existence of small oscillations at zeros of Brownian motion. Lecture Notes in Mathematics, 381:134–149, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057259/.

Knobloch:1972:OCP

[Kno72] H. W. Knobloch. On optimal control problems with bounded state variables. Lecture Notes in Mathematics, 280:103–119, 1972.

CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066922/.

Knops:1973:LCO

[Kno73a] R. J. Knops. Logarithmic convexity and other techniques applied to problems in continuum mechanics. Lecture Notes in Mathematics, 316:31–54, 1973. CODEN LNMAA2. ISBN 3-540-06159-2 (print), 3-540-38370-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069622/.

Knops:1973:SNW

[Kno73b] R. J. Knops, editor. Symposium on Non-Well-Posed Problems and Logarithmic Convexity: Held in Heriot–Watt University, Edinburgh/Scotland March 22–24, 1972, volume 316 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06159-2 (print), 3-540-38370-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 316.

URL http://link.springer.com/book/10.1007/BFb0069620;
 http://www.springerlink.com/content/978-3-540-38370-
3.

Knuth:1971:EFS

[Knu71a] Donald E. Knuth. Examples of formal semantics. Lecture Notes in Mathematics, 188:212–235, 1971. CODEN LNMAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059699/.

Knutson:1971:ASa

[Knu71b] Donald Knutson. Algebraic spaces. Lecture Notes in Mathematics, 203:91–152, 1971. CODEN LNMAA2. ISBN 3-540-05496-0 (print), 3-540-36663-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059753/.

Knutson:1971:ASb

[Knu71c] Donald Knutson. Algebraic Spaces, volume 203 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CO-DEN LNMAA2. ISBN 3-540-05496-0 (print), 3-540-36663-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 261 pp. LCCN QA3 .L28 no. 203. URL http:/

/link.springer.com/book/10.1007/BFb0059750;   http://
www.springerlink.com/content/978-3-540-36663-8.

Knutson:1971:ETS

[Knu71d] Donald Knutson. The etale topology of schemes. Lecture Notes in Mathematics, 203:29–90, 1971. CODEN LNMAA2. ISBN 3-540-05496-0 (print), 3-540-36663-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059752/.

Knutson:1971:FT

[Knu71e] Donald Knutson. The finiteness theorem. Lecture Notes in Mathematics, 203:176–203, 1971. CODEN LNMAA2. ISBN 3-540-05496-0 (print), 3-540-36663-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059755/.

Knutson:1971:FAS

[Knu71f] Donald Knutson. Formal algebraic spaces. Lecture Notes in Mathematics, 203:204–251, 1971. CODEN LNMAA2. ISBN 3-540-05496-0 (print), 3-540-36663-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059756/.

Knutson:1971:I

[Knu71g] Donald Knutson. Introduction. Lecture Notes in Mathematics, 203:1–28, 1971. CODEN LNMAA2. ISBN 3-540-05496-0 (print), 3-540-36663-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059751/.

Knutson:1971:QSN

[Knu71h] Donald Knutson. Quasicoherent sheaves on Noetherian locally separated algebraic spaces. Lecture Notes in Mathematics, 203: 153–175, 1971. CODEN LNMAA2. ISBN 3-540-05496-0 (print), 3-540-36663-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059754/.

Knudsen:1973:CNP

[Knu73a] Finn F. Knudsen. Construction of nice polyhedral subdivisions. Lecture Notes in Mathematics, 339:109–164, 1973. CODEN LN-MAA2. ISBN 3-540-06432-X (print), 3-540-37755-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070321/.

Knudsen:1973:FA

[Knu73b] Finn F. Knudsen. Further applications. Lecture Notes in Mathematics, 339:165–209, 1973. CODEN LNMAA2. ISBN 3-540-06432-X (print), 3-540-37755-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070322/.

Knutson:1973:R

[Knu73c] D. Knutson. λ-rings. Lecture Notes in Mathematics, 308:5– 58, 1973. CODEN LNMAA2. ISBN 3-540-06184-3 (print), 3-540-38408-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069219/.

Knutson:1973:FT

[Knu73d] D. Knutson. The fundamental theorem. Lecture Notes in Mathematics, 308:124–193, 1973. CODEN LNMAA2. ISBN 3-540-06184-3 (print), 3-540-38408-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069221/.

Knutson:1973:I

[Knu73e] D. Knutson. Introduction. Lecture Notes in Mathematics, 308: 1–4, 1973. CODEN LNMAA2. ISBN 3-540-06184-3 (print), 3-540-38408-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069218/.

Knutson:1973:RTG

[Knu73f] D. Knutson. The representation theory of groups. Lecture Notes in Mathematics, 308:59–123, 1973. CODEN LNMAA2. ISBN 3-540-06184-3 (print), 3-540-38408-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069220/.

Knutson:1973:RRT

[Knu73g] Donald Knutson. λ-Rings and the Representation Theory of the Symmetric Group, volume 308 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06184-3 (print), 3-540-38408-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 203 pp. LCCN QA3 .L28 no. 308.

URL http://link.springer.com/book/10.1007/BFb0069217;
 http://www.springerlink.com/content/978-3-540-38408-
3.

Kondo:1973:CPK

[KŌ73a] Ryōji Kondō and Yōichi Ōshima. A characterization of potential kernels for recurrent Markov chains with strong Feller transition function. Lecture Notes in Mathematics, 330:213–238, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061489/.

Kotani:1973:MPSa

[KO73b] Shinichi Kotani and Yasunori Okabe. On a Markovian property of stationary Gaussian processes with a multi-dimensional parameter. Lecture Notes in Mathematics, 287:153–163, 1973. CODEN LNMAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068150/.

Knus:1974:PFP

[KO74a] Max-Albert Knus and Manuel Ojanguren. I. Préliminaires. (French) [I. Preliminaries]. Lecture Notes in Mathematics, 389: 3–27, 1974. CODEN LNMAA2. ISBN 3-540-06791-4 (print), 3-540-38406-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057800/.

Knus:1974:ITD

[KO74b] Max-Albert Knus and Manuel Ojanguren. II. Théories de la déscente. (French) []. Lecture Notes in Mathematics, 389:28– 70, 1974. CODEN LNMAA2. ISBN 3-540-06791-4 (print), 3-540-38406-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057801/.

Knus:1974:IAS

[KO74c] Max-Albert Knus and Manuel Ojanguren. III. Algébres separables. Lecture Notes in Mathematics, 389:71–105, 1974. CODEN LNMAA2. ISBN 3-540-06791-4 (print), 3-540-38406-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057802/.

Knus:1974:IAA

[KO74d] Max-Albert Knus and Manuel Ojanguren. IV. Automorphismes des algèbres d’Azumaya. (French) []. Lecture Notes in Mathematics, 389:106–118, 1974. CODEN LNMAA2. ISBN 3-540-06791-4 (print), 3-540-38406-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057803/.

Knus:1974:TDA

[KO74e] Max-Albert Knus and Manuel Ojanguren. Théorie de la Descente et Algèbres d’Azumaya. (French) [], volume 389 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06791-4 (print), 3-540-38406-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 153 (est.) pp. LCCN QA3 .L28 no. 389. URL http:/

/link.springer.com/book/10.1007/BFb0057799;   http://
www.springerlink.com/content/978-3-540-38406-9.

Knus:1974:VCA

[KO74f] Max-Albert Knus and Manuel Ojanguren. V. La cohomologie d’Amitsur. (French) []. Lecture Notes in Mathematics, 389:119– 150, 1974. CODEN LNMAA2. ISBN 3-540-06791-4 (print), 3-540-38406-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057804/.

Knus:1974:VRS

[KO74g] Max-Albert Knus and Manuel Ojanguren. VI. Remarques et sources. (French) []. Lecture Notes in Mathematics, 389:151– 153, 1974. CODEN LNMAA2. ISBN 3-540-06791-4 (print), 3-540-38406-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057805/.

Koh:1972:QMO

[Koh72] Kwangil Koh. Quasisimple modules and other topics in ring theory. Lecture Notes in Mathematics, 246:323–428, 1972. CODEN LNMAA2. ISBN 3-540-05760-9 (print), 3-540-37181-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059568/.

Kolar:1972:DMA

[Kol72] W. Kolar. Über Differenzenverfahren von monotoner Art für nichtlineare parabolische Randwertprobleme. (German) []. Lecture Notes in Mathematics, 267:139–157, 1972. CODEN LN-MAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061618/.

Komkov:1972:CBC

[Kom72a] Vadim Komkov. Classification of the boundary conditions in optimal control theory of beams and thin plates. Lecture Notes in Mathematics, 253:210–235, 1972. CODEN LNMAA2. ISBN 3-540-05734-X (print), 3-540-37121-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058910/.

Komkov:1972:IR

[Kom72b] Vadim Komkov. Introductory remarks. Lecture Notes in Mathematics, 253:1–3, 1972. CODEN LNMAA2. ISBN 3-540-05734-X (print), 3-540-37121-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058906/.

Komkov:1972:OCV

[Kom72c] Vadim Komkov. The optimal control of vibrating beams. Lecture Notes in Mathematics, 253:61–118, 1972. CODEN LN-MAA2. ISBN 3-540-05734-X (print), 3-540-37121-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058908/.

Komkov:1972:OCTb

[Kom72d] Vadim Komkov. Optimal Control Theory for the Damping of Vibrations of Simple Elastic Systems, volume 253 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CO-DEN LNMAA2. ISBN 3-540-05734-X (print), 3-540-37121-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 240 pp. LCCN QA3 .L28 no. 253. URL http:/

/link.springer.com/book/10.1007/BFb0058905;   http://
www.springerlink.com/content/978-3-540-37121-2.

Komkov:1972:OCTa

[Kom72e] Vadim Komkov. Optimal control theory for thin plates. Lecture Notes in Mathematics, 253:119–209, 1972. CODEN LN-MAA2. ISBN 3-540-05734-X (print), 3-540-37121-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058909/.

Komkov:1972:SSR

[Kom72f] Vadim Komkov. A summary of some results on controls of hyperbolic partial differential equations. Lecture Notes in Mathematics, 253:4–60, 1972. CODEN LNMAA2. ISBN 3-540-05734-X (print), 3-540-37121-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058907/.

Komatsu:1973:FTP

[Kom73a] Hikosaburo Komatsu. Foundation of the theory of pseudodifferential equations. Lecture Notes in Mathematics, 287:315– 456, 1973. CODEN LNMAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068155/.

Komatsu:1973:HLP

[Kom73b] Hikosaburo Komatsu. Hyperfunctions and linear partial differential equations. Lecture Notes in Mathematics, 287:180–191, 1973. CODEN LNMAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068152/.

Komatsu:1973:HPD

[Kom73c] Hikosaburo Komatsu, editor. Hyperfunctions and Pseudo-Differential Equations: Proceedings of a Conference at Katata, 1971, volume 287 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 287. URL http:/

/link.springer.com/book/10.1007/BFb0068143;   http://
www.springerlink.com/content/978-3-540-38506-6.

Komatsu:1973:ITH

[Kom73d] Hikosaburo Komatsu. An introduction to the theory of hyperfunctions. Lecture Notes in Mathematics, 287:3–40, 1973. CO-DEN LNMAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068144/.

Komatsu:1973:RCS

[Kom73e] Hikosaburo Komatsu. Relative cohomology of sheaves of solutions of differential equations. Lecture Notes in Mathematics, 287: 192–261, 1973. CODEN LNMAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068153/.

Komatsu:1973:SSP

[Kom73f] Hikosaburo Komatsu. Structure of systems of pseudo-differential equations. Lecture Notes in Mathematics, 287:457–523, 1973. CODEN LNMAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068156/.

Komatsu:1973:TM

[Kom73g] Hikosaburo Komatsu. Theory of microfunctions. Lecture Notes in Mathematics, 287:265–314, 1973. CODEN LNMAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068154/.

Komatsu:1973:UH

[Kom73h] Hikosaburo Komatsu. Ultradistributions and hyperfunctions. Lecture Notes in Mathematics, 287:164–179, 1973. CODEN LN-MAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068151/.

Koranyi:1970:RBV

[Kor70] Adam Korányi. A remark on boundary values of functions of several complex variables. Lecture Notes in Mathematics, 155: 115–121, 1970. CODEN LNMAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060319/.

Korn:1974:TDT

[Kor74] D. G. Korn. Transonic design in two dimensions. Lecture Notes in Mathematics, 430:271–288, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066274/.

Kostant:1970:LMAa

[Kos70a] Professor Bertram Kostant. Lectures in modern analysis and applications I. Lecture Notes in Mathematics, 170:209–211, 1970. CODEN LNMAA2. ISBN 3-540-05284-4 (print), 3-540-36417-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079069/.

Kostant:1970:LMAb

[Kos70b] Professor Bertram Kostant. Lectures in modern analysis and applications II. Lecture Notes in Mathematics, 170:212–213, 1970. CODEN LNMAA2. ISBN 3-540-05284-4 (print), 3-540-36417-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079070.

Kostant:1970:QUR

[Kos70c] Professor Bertram Kostant. Quantization and unitary representations. Lecture Notes in Mathematics, 170:87–208, 1970. CO-DEN LNMAA2. ISBN 3-540-05284-4 (print), 3-540-36417-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0079068/.

Koszul:1971:TJS

[Kos71] Jean-Louis Koszul. Travaux de J. Stallings sur la décomposition des groupes en produits libres. (French) []. Lecture Notes in Mathematics, 179:173–185, 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058811/.

Kotani:1973:MPSb

[Kot73] Shinichi Kotani. On a Markov property for stationary Gaussian processes with a multidimensional parameter. Lecture Notes in Mathematics, 330:239–250, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061490/.

Kowalski:1970:CUE

[Kow70] Robert Kowalski. The case for using equality axioms in automatic demonstration. Lecture Notes in Mathematics, 125:112– 127, 1970. CODEN LNMAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060628/.

Kowada:1973:CRE

[Kow73] Masasi Kowada. Convergence rate in the ergodic theorem for an analytic flow on the torus. Lecture Notes in Mathematics, 330: 251–254, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061491/.

Kozin:1972:SLS

[Koz72] F. Kozin. Stability of the linear stochastic system. Lecture Notes in Mathematics, 294:186–229, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064943/.

Khabbaz:1971:NCP

[KP71] Samir Khabbaz and Everett Pitcher. Nondegenerate critical points of deficiency one. Lecture Notes in Mathematics, 192: 271–281, 1971. CODEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066827/.

Koch:1972:CRI

[KP72] R. J. Koch and G. Pall. Centralizers of rootless integral matrices. Lecture Notes in Mathematics, 299:76–80, 1972. CODEN LNMAA2. ISBN 3-540-06078-2 (print), 3-540-38066-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066761/.

Kawakubo:1972:IMT

[KR72] Katsuo Kawakubo and Frank Raymond. The index of manifolds with toral actions and geometric interpretations of the σ(∞, (S^{1}, M^{n})) invariant of Atiyah and Singer. Lecture Notes in Mathematics, 298:228–233, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070045/.

Karoui:1973:PDF

[KR73] N. Karoui and H. Reinhard. Processus de diffusion dans R^{n}. (French) []. Lecture Notes in Mathematics, 321:95–117, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0071399/.

Kra:1974:DS

[Kra74] Irwin Kra. Deformation spaces. Lecture Notes in Mathematics, 400:48–70, 1974. CODEN LNMAA2. ISBN 3-540-06840-6 (print), 3-540-37776-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065675/.

Kreisel:1970:HPS

[Kre70a] G. Kreisel. Hilbert’s programme and the search for automatic proof procedures. Lecture Notes in Mathematics, 125:128–146, 1970. CODEN LNMAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060629/.

Krengel:1970:TFI

[Kre70b] Ulrich Krengel. Transformations without finite invariant measure have finite strong generators. Lecture Notes in Mathematics, 160:133–157, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060652/.

Kreith:1973:AOT

[Kre73a] Kurt Kreith. Abstract oscillation theory. Lecture Notes in Mathematics, 324:42–56, 1973. CODEN LNMAA2. ISBN 3-540-06258-0 (print), 3-540-40005-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067543/.

Kreith:1973:CT

[Kre73b] Kurt Kreith. The classical theory. Lecture Notes in Mathematics, 324:3–10, 1973. CODEN LNMAA2. ISBN 3-540-06258-0 (print), 3-540-40005-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067539/.

Kreith:1973:COT

[Kre73c] Kurt Kreith. Complex oscillation theory. Lecture Notes in Mathematics, 324:57–65, 1973. CODEN LNMAA2. ISBN 3-540-06258-0 (print), 3-540-40005-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067544/.

Kreith:1973:DEE

[Kre73d] Kurt Kreith. Differential equations of even order. Lecture Notes in Mathematics, 324:66–95, 1973. CODEN LNMAA2. ISBN 3-540-06258-0 (print), 3-540-40005-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067545/.

Kreith:1973:FOD

[Kre73e] Kurt Kreith. Fourth order differential equations. Lecture Notes in Mathematics, 324:96–101, 1973. CODEN LNMAA2. ISBN 3-540-06258-0 (print), 3-540-40005-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067546/.

Kreith:1973:I

[Kre73f] Kurt Kreith. Introduction. Lecture Notes in Mathematics, 324: 1–2, 1973. CODEN LNMAA2. ISBN 3-540-06258-0 (print), 3-540-40005-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067538.

Kreith:1973:NE

[Kre73g] Kurt Kreith. Nonselfadjoint equations. Lecture Notes in Mathematics, 324:11–16, 1973. CODEN LNMAA2. ISBN 3-540-06258-0 (print), 3-540-40005-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067540/.

Kreith:1973:OT

[Kre73h] Kurt Kreith. Oscillation Theory, volume 324 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CO-DEN LNMAA2. ISBN 3-540-06258-0 (print), 3-540-40005-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 101 (est.) pp. LCCN QA3 .L28 no. 324. URL http:/

/link.springer.com/book/10.1007/BFb0067537;   http://
www.springerlink.com/content/978-3-540-40005-9.

Kreith:1973:PDE

[Kre73i] Kurt Kreith. Partial differential equations. Lecture Notes in Mathematics, 324:17–29, 1973. CODEN LNMAA2. ISBN 3-540-06258-0 (print), 3-540-40005-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067541/.

Kreith:1973:RTA

[Kre73j] Kurt Kreith. Related topics in analysis. Lecture Notes in Mathematics, 324:30–41, 1973. CODEN LNMAA2. ISBN 3-540-06258-0 (print), 3-540-40005-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067542/.

Kree:1974:SFE

[Kré74a] Paul Krée. Solutions faibles d’équations aux dérivées fonctionnelles I. (French) []. Lecture Notes in Mathematics, 410:142– 181, 1974. CODEN LNMAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066046/.

Kreiss:1974:BCH

[Kre74b] Heinz-Otto Kreiss. Boundary conditions for hyperbolic differential equations. Lecture Notes in Mathematics, 363:64–74, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069126/.

Kreth:1974:ANL

[Kre74c] H. Kreth. Ein Äquivalenzsatz bei der numerischen Lösung quasilinearer Anfangswertaufgaben. (German) []. Lecture Notes in Mathematics, 395:33–55, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060663/.

Krieger:1970:AWA

[Kri70] Wolfgang Krieger. On the Araki–Woods asymptotic ratio set and non-singular transformations of a measure space. Lecture Notes in Mathematics, 160:158–177, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060653/.

Krivine:1971:TCT

[Kri71] Jean-Louis Krivine. Théorèmes de consistance en théorie de la mesure de R. Solovay. (French) []. Lecture Notes in Mathematics, 179:187–197, 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058812/.

Krickeberg:1973:MPP

[Kri73] Klaus Krickeberg. Moments of point processes. Lecture Notes in Mathematics, 296:70–101, 1973. CODEN LNMAA2. ISBN 3-540-06211-4 (print), 3-540-38485-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059820/.

Kronk:1972:CNT

[Kro72] Hudson V. Kronk. The chromatic number of triangle-free graphs. Lecture Notes in Mathematics, 303:179–181, 1972. CODEN LN-MAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067369/.

Krogh:1974:CSI

[Kro74] Fred T. Krogh. Changing stepsize in the integration of differential equations using modified divided differences. Lecture Notes in Mathematics, 362:22–71, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066584/.

Kruglov:1973:CNC

[Kru73] V. M. Kruglov. Convergence of numeric characteristics of sums of independent random variables and global theorems. Lecture Notes in Mathematics, 330:255–286, 1973. CODEN LN-MAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061492/.

Kelley:1970:AGP

[KS70] Henry J. Kelley and Jason L. Speyer. Accelerated gradient projection. Lecture Notes in Mathematics, 132:151–158, 1970. CO-DEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066680/.

Khabbaz:1971:NFA

[KS71a] Samir Khabbaz and Gilbert Stengle. Normal forms for analytic matrix valued functions. Lecture Notes in Mathematics, 209:34– 57, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068891/.

Kirby:1971:STT

[KS71b] R. C. Kirby and L. C. Siebenmann. Some theorems on topological manifolds. Lecture Notes in Mathematics, 197:1–7, 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068607/.

Knapp:1972:ITP

[KS72] A. W. Knapp and E. M. Stein. Irreducibility theorems for the principal series. Lecture Notes in Mathematics, 266:197– 214, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059645/.

Kaminker:1973:HPV

[KS73a] Jerome Kaminker and Claude Schochet. Ext(X) from a homological point of view. Lecture Notes in Mathematics, 345:129– 140, 1973. CODEN LNMAA2. ISBN 3-540-06496-6 (print), 3-540-37808-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058918/.

Kuijk:1973:MFOb

[KS73b] Willem Kuijk and Jean-Pierre Serre, editors. Modular Functions of One Variable III: Proceedings International Summer School University of Antwerp, RUCA July 17–August 3, 1972, volume 350 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06483-4 (print), 3-540-37802-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN

QA3 .L28 no. 350. URL http://link.springer.com/book/
10.1007/978-3-540-37802-0;   http://www.springerlink.
com/content/978-3-540-37802-0.

Kegel:1974:EPP

[KS74a] O. H. Kegel and A. Schleiermacher. Embeddings of projective planes. Lecture Notes in Mathematics, 403:61–62, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057376.

Kelly:1974:REC

[KS74b] G. M. Kelly and Ross Street. Review of the elements of 2categories. Lecture Notes in Mathematics, 420:75–103, 1974. CO-DEN LNMAA2. ISBN 3-540-06966-6 (print), 3-540-37270-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063101/.

Keimel:1971:LAS

[KT71] Klaus Keimel and Silviu Teleman. Lectures on the Applications of Sheaves to Ring Theory, volume 248 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LN-MAA2. ISBN 0-387-05714-5, 3-540-05714-5 (print), 3-540-37081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 315 pp. LCCN QA3 .L28 no. 248. URL http:/

/link.springer.com/book/10.1007/BFb0058562;   http://
www.springerlink.com/content/978-3-540-37081-9.

Kubota:1971:SNT

[Kub71] Tomio Kubota. Some number-theoretical results on real analytic automorphic forms. Lecture Notes in Mathematics, 185:

87–96, 1971. CODEN LNMAA2. ISBN 3-540-05372-7 (print), 3-540-36493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058766/.

Kubo:1973:EDS

[Kub73] Izumi Kubo. Ergodicity of the dynamical system of a particle on a domain with irregular walls. Lecture Notes in Mathematics, 330:287–295, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061493/.

Kuhlmann:1974:CID

[Kuh74] Norbert Kuhlmann. Sur la coherence des images directes. (French) []. Lecture Notes in Mathematics, 409:347–356, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068110/.

Kuiper:1971:MAP

[Kui71a] Nicolaas H. Kuiper, editor. Manifolds — Amsterdam 1970: Proceedings of the Nuffic Summer School on Manifolds Amsterdam, August 17–29, 1970, volume 197 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 197. URL http:/

/link.springer.com/book/10.1007/BFb0068606;   http://
www.springerlink.com/content/978-3-540-36653-9.

Kuiper:1971:MRC

[Kui71b] Nicolaas H. Kuiper. Morse relations for curvature and tightness. Lecture Notes in Mathematics, 209:77–89, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068893/.

Kuijk:1973:MFOa

[Kui73a] Willem Kuijk, editor. Modular Functions of One Variable I: Proceedings International Summer School University of Antwerp, RUCA July 17–August 3, 1972, volume 320 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06219-X (print), 3-540-38509-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 320.

URL http://link.springer.com/book/10.1007/978-3-540-
38509-7;   http://www.springerlink.com/content/978-3-
540-38509-7.

Kuiper:1973:VRT

[Kui73b] Nicolaas H. Kuiper. Sur les variétés riemanniennes très pincées. (French) []. Lecture Notes in Mathematics, 317:201–218, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069283/.

Kujala:1971:AD

[Kuj71a] Robert O. Kujala. On algebraic divisors in C^{K}. Lecture Notes in Mathematics, 184:223–230, 1971. CODEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059287/.

Kujala:1971:P

[Kuj71b] Robert O. Kujala. Problems. Lecture Notes in Mathematics, 184: 231–234, 1971. CODEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059288/.

Kumpera:1974:IDP

[Kum74] A. Kumpera. Invariants différentiels d’un pseudogroupe de Lie. (French) []. Lecture Notes in Mathematics, 392:121–162, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064119/.

Kunen:1973:XMN

[Kun73] Kenneth Kunen. XVI. A model for the negation of the Axiom of Choice. Lecture Notes in Mathematics, 337:489–494, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066786/.

Kuo:1971:RTA

[Kuo71] T.-C. Kuo. The ratio test for analytic Whitney stratifications. Lecture Notes in Mathematics, 192:141–149, 1971. CODEN LN-MAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066819/.

Kuratowski:1970:SPC

[Kur70a] K. Kuratowski. Some problems concerning semi-continuous setvalued mappings. Lecture Notes in Mathematics, 171:45–48, 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069719/.

Kurzweil:1970:GSF

[Kur70b] Jaroslav Kurzweil. Global solutions of functional differential equations. Lecture Notes in Mathematics, 144:134–139, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059930/.

Kurzweil:1971:CIC

[Kur71] J. Kurzweil. Conditions for integrability of certain equations. Lecture Notes in Mathematics, 206:9, 1971. CODEN LN-MAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070121.

Kushner:1972:SS

[Kus72] H. J. Kushner. Stochastic stability. Lecture Notes in Mathematics, 294:97–124, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064937/.

LopezdeMedrano:1972:CDC

[Lan72] Santiago López de Medrano. Cobordism of diffeomorphisms of (k − 1)-connected 2k-manifolds. Lecture Notes in Mathematics, 298:217–227, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070044/.

Lachlan:1972:ENL

[Lac72] A. H. Lachlan. Embedding nondistributive lattices in the recursively enumerable degrees. Lecture Notes in Mathematics, 255: 149–177, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059544/.

Lachlan:1973:PMC

[Lac73] A. H. Lachlan. The priority method for the construction of recursively enumerable sets. Lecture Notes in Mathematics, 337: 299–310, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066779/.

Lambek:1971:AS

[Lam71a] Joachim Lambek. Additive semantics. Lecture Notes in Mathematics, 177:15–31, 1971. CODEN LNMAA2. ISBN 3-540-05340-9 (print), 3-540-36448-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061031/.

Lambek:1971:ET

[Lam71b] Joachim Lambek. Embedding theorems. Lecture Notes in Mathematics, 177:64–78, 1971. CODEN LNMAA2. ISBN 3-540-05340-9 (print), 3-540-36448-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061034/.

Lambek:1971:ETT

[Lam71c] Joachim Lambek. An exposition of torsion theories. Lecture Notes in Mathematics, 177:1–14, 1971. CODEN LNMAA2. ISBN 3-540-05340-9 (print), 3-540-36448-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061030/.

Lambek:1971:PM

[Lam71d] Joachim Lambek. Protorsion modules. Lecture Notes in Mathematics, 177:54–63, 1971. CODEN LNMAA2. ISBN 3-540-05340-9 (print), 3-540-36448-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061033/.

Lambek:1971:TIR

[Lam71e] Joachim Lambek. Torsion ideals and rings of quotients. Lecture Notes in Mathematics, 177:32–53, 1971. CODEN LN-MAA2. ISBN 3-540-05340-9 (print), 3-540-36448-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061032/.

Lambek:1971:TTA

[Lam71f] Joachim Lambek. Torsion Theories, Additive Semantics, and Rings of Quotients, volume 177 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05340-9 (print), 3-540-36448-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 94 pp. LCCN QA3 .L28 no. 177.

URL http://link.springer.com/book/10.1007/BFb0061029;
 http://www.springerlink.com/content/978-3-540-36448-
1.

Lambek:1972:DSC

[Lam72] Joachim Lambek. Deductive systems and categories III. Cartesian closed categories, intuitionist propositional calculus, and combinatory logic. Lecture Notes in Mathematics, 274:57–82, 1972. CODEN LNMAA2. ISBN 3-540-05920-2 (print), 3-540-37609-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073965/.

Lam:1973:ITG

[Lam73] Ping-Fun Lam. The intermediate transformation groups. Lecture Notes in Mathematics, 318:174–181, 1973. CODEN LN-MAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061734/.

Lambert:1974:NMS

[Lam74] J. D. Lambert. Nonlinear methods for stiff systems of ordinary differential equations. Lecture Notes in Mathematics, 363:75– 88, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069127/.

Langlands:1970:PTA

[Lan70] Professor R. P. Langlands. Problems in the theory of automorphic forms to Salomon Bochner in gratitude. Lecture Notes in Mathematics, 170:18–61, 1970. CODEN LNMAA2. ISBN 3-540-05284-4 (print), 3-540-36417-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079065/.

Lanford:1973:BPS

[Lan73a] Oscar E. Lanford III. Bifurcation of periodic solutions into invariant tori: The work of Ruelle and Takens. Lecture Notes in Mathematics, 322:159–192, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060566/.

Langlands:1973:MFA

[Lan73b] R. P. Langlands. Modular forms and ℓ-adic representations. Lecture Notes in Mathematics, 349:361–500, 1973. CODEN LN-MAA2. ISBN 3-540-06558-X (print), 3-540-37855-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-37855-6_
6/.

Langrand:1974:PFM

[Lan74] Claude Langrand. Précapacités fortes et mesures d’information. (French) [Strong precapacities and information measures]. Lecture Notes in Mathematics, 398:36–48, 1974. CODEN LN-MAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065766/.

Laplaza:1972:CD

[Lap72a] Miguel L. Laplaza. Coherence for distributivity. Lecture Notes in Mathematics, 281:29–65, 1972. CODEN LNMAA2. ISBN 3-540-05963-6 (print), 3-540-37958-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059555/.

Laplaza:1972:NRC

[Lap72b] Miguel L. Laplaza. A new result of coherence for distributivity. Lecture Notes in Mathematics, 281:214–235, 1972. CODEN LNMAA2. ISBN 3-540-05963-6 (print), 3-540-37958-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059560/.

Lasota:1970:BVP

[Las70] Andrzej Lasota. Boundary value problems for second order differential equations. Lecture Notes in Mathematics, 144:140– 152, 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059931/.

LasVergnas:1971:DTM

[Las71] Michel Las Vergnas. Sur la dualité en théorie des matroïdes. (French) [On the duality in matroid theory]. Lecture Notes in Mathematics, 211:67–85, 1971. CODEN LNMAA2. ISBN 3-540-05545-2 (print), 3-540-36877-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061076/.

LasVergnas:1974:HBF

[Las74] M. Las Vergnas, C. N. R. S. Sur les hypergraphes bichromatiques. (French) []. Lecture Notes in Mathematics, 411:102–110, 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066185/.

Latour:1972:CNS

[Lat72] François Latour. Chirurgie non simplement connexe (d’après C. T. C. Wall). (French) []. Lecture Notes in Mathematics, 244: 289–322, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058708/.

Laudet:1970:AOF

[Lau70] M. Laudet. Allocution d’ouverture. (French) [Opening remarks]. Lecture Notes in Mathematics, 125:1–5, 1970. CODEN LN-MAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060621/.

Lawvere:1972:TAG

[Law72a] F. W. Lawvere, editor. Toposes, Algebraic Geometry and Logic: Dalhousie University, Halifax, January 16–19, 1971, volume 274 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05920-2 (print), 3-540-37609-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 274. URL http:/

/link.springer.com/book/10.1007/BFb0073961;   http://
www.springerlink.com/content/978-3-540-37609-5.

Lawvere:1972:I

[Law72b] F. William Lawvere. Introduction. Lecture Notes in Mathematics, 274:1–12, 1972. CODEN LNMAA2. ISBN 3-540-05920-2 (print), 3-540-37609-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073962/.

Lawton:1973:SCC

[Law73] Wayne Lawton. The structure of compact connected groups which admit an expansive automorphism. Lecture Notes in Mathematics, 318:182–196, 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061735/.

Lazarov:1972:QTM

[Laz72] Connor Lazarov. Quillen’s theorem for MU. Lecture Notes in Mathematics, 298:183–190, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070042/.

Lazet:1973:AAE

[Laz73] D. Lazet. Applications analytiques dans les espaces bornologiques. (French) []. Lecture Notes in Mathematics, 332:1–47, 1973. CODEN LNMAA2. ISBN 3-540-06372-2 (print), 3-540-46980-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060905/.

Lascoux:1970:FDF

[LB70a] Alain Lascoux and Marcel Berger. C^{ω} fibres en droites. (French) []. Lecture Notes in Mathematics, 154:43–49, 1970. CODEN LNMAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069339/.

Lascoux:1970:CCF

[LB70b] Alain Lascoux and Marcel Berger. Classes de Chern. (French) []. Lecture Notes in Mathematics, 154:70–76, 1970. CODEN LNMAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069343/.

Lascoux:1970:CVK

[LB70c] Alain Lascoux and Marcel Berger. Cohomologie des variétés Kähleriennes. (French) []. Lecture Notes in Mathematics, 154: 31–38, 1970. CODEN LNMAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069337/.

Lascoux:1970:CFH

[LB70d] Alain Lascoux and Marcel Berger. Cohomologie et formes harmoniques. (French) []. Lecture Notes in Mathematics, 154:23– 30, 1970. CODEN LNMAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069336/.

Lascoux:1970:C

[LB70e] Alain Lascoux and Marcel Berger. Connexions. Lecture Notes in Mathematics, 154:62–69, 1970. CODEN LNMAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069342/.

Lascoux:1970:EF

[LB70f] Alain Lascoux and Marcel Berger. Éclatements. (French) []. Lecture Notes in Mathematics, 154:20–22, 1970. CODEN LN-MAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069335/.

Lascoux:1970:EFV

[LB70g] Alain Lascoux and Marcel Berger. Espaces fibres vectoriels. (French) []. Lecture Notes in Mathematics, 154:39–42, 1970. CO-DEN LNMAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069338/.

Lascoux:1970:SRF

[LB70h] Alain Lascoux and Marcel Berger. Surfaces de Riemann. (French) []. Lecture Notes in Mathematics, 154:50–55, 1970. CODEN LNMAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069340/.

Lascoux:1970:TKF

[LB70i] Alain Lascoux and Marcel Berger. Théorème de Kodaira. (French) []. Lecture Notes in Mathematics, 154:56–61, 1970. CO-DEN LNMAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069341/.

Lascoux:1970:VVR

[LB70j] Alain Lascoux and Marcel Berger. Variétés C^{∞} — variétés Riemanniennes. (French) []. Lecture Notes in Mathematics, 154:1– 7, 1970. CODEN LNMAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069332/.

Lascoux:1970:VF

[LB70k] Alain Lascoux and Marcel Berger. Variétés C^{ω}. (French) []. Lecture Notes in Mathematics, 154:8–15, 1970. CODEN LN-MAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069333/.

Lascoux:1970:VKC

[LB70l] Alain Lascoux and Marcel Berger. Variétés Kähleriennes Compactes. (French) [Compact Kahlerian varieties], volume 154 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970.

CODEN LNMAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 83 pp. LCCN QA3 .L28 no. 154. URL http:/

/link.springer.com/book/10.1007/BFb0069331;   http://
www.springerlink.com/content/978-3-540-36341-5.

Lascoux:1970:VKF

[LB70m] Alain Lascoux and Marcel Berger. Variétés Kähleriennes. (French) []. Lecture Notes in Mathematics, 154:16–19, 1970. CO-DEN LNMAA2. ISBN 3-540-05182-1 (print), 3-540-36341-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069334/.

Leblanc:1973:ABA

[Leb73a] Noël Leblanc. Algèbres de Banach associées à un opérateur différentiel de Sturm–Liouville. (French) []. Lecture Notes in Mathematics, 336:40–50, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065786/.

Lebowitz:1973:ETS

[Leb73b] Joel L. Lebowitz. Ergodic theory and statistical mechanics of non-equilibrium processes. Lecture Notes in Mathematics, 322: 193–209, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060567/.

Lehner:1973:BM

[Leh73a] Joseph Lehner. Back matter. Lecture Notes in Mathematics, 320: 195, 1973. CODEN LNMAA2. ISBN 3-540-06219-X (print), 3-540-38509-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-38509-7/1.pdf.

Lehner:1973:C

[Leh73b] Joseph Lehner. On the A_{q}(Γ) ⊂ B_{q}(Γ) conjecture. Lecture Notes in Mathematics, 320:187–194, 1973. CODEN LNMAA2. ISBN 3-540-06219-X (print), 3-540-38509-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38509-7_7/.

Lelong:1970:FPE

[Lel70a] P. Lelong. Fonctions plurisousharmoniques et ensembles polaires sur une algèbre de fonctions holomorphes. (French) []. Lecture Notes in Mathematics, 116:1–20, 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059241/.

Lelong:1970:SPL

[Lel70b] Pierre Lelong, editor. Séminaire Pierre Lelong (Analyse) Année 1969, volume 116 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 116; QA300 .S46 1970; QA3 .L28 no. 116. URL http:/

/link.springer.com/book/10.1007/BFb0059240;   http://
www.springerlink.com/content/978-3-540-36240-1.

Lelong:1971:TBS

[Lel71a] P. Lelong. Théorème de Banach–Steinhaus pour les polynômes; applications entières d’espaces vectoriels complexes. (French) []. Lecture Notes in Mathematics, 205:87–112, 1971. CODEN LN-MAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059835/.

Lelong:1971:RRA

[Lel71b] Pierre Lelong. Recent results on analytic mappings and plurisubharmonic functions in topological linear spaces. Lecture Notes in Mathematics, 185:97–124, 1971. CODEN LNMAA2. ISBN 3-540-05372-7 (print), 3-540-36493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058767/.

Lelong:1971:SPL

[Lel71c] Pierre Lelong, editor. Séminaire Pierre Lelong (Analyse) Année 1970, volume 205 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 205. URL http:/

/link.springer.com/book/10.1007/BFb0059828;   http://
www.springerlink.com/content/978-3-540-36867-0.

Lelong:1972:REP

[Lel72a] Pierre Lelong. Remarques sur l’exposé précédent et le “lemme polynômial” de Leja. (French) []. Lecture Notes in Mathematics, 275:143–146, 1972. CODEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058497/.

Lelong:1972:SPL

[Lel72b] Pierre Lelong, editor. Séminaire Pierre Lelong (Analyse) Année 1970–1971, volume 275 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 275. URL http:/

/link.springer.com/book/10.1007/BFb0058488;   http://
www.springerlink.com/content/978-3-540-37610-1.

Lelong:1972:TCP

[Lel72c] Pierre Lelong. Topologies sur les courants positifs fermés et les ensembles analytiques sur une variété analytique complexe dénombrable à l’infini. (French) []. Lecture Notes in Mathematics, 275:27–70, 1972. CODEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058491/.

Lelong:1972:VAA

[Lel72d] Pierre Lelong. Valeurs algébriques d’une application méromorphe (d’après E. Bombieri). (French) []. Lecture Notes in Mathematics, 244:29–45, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058695/.

Lelong:1973:EEC

[Lel73a] Pierre Lelong. Éléments extrémaux sur le cône des courants positifs fermés. (French) []. Lecture Notes in Mathematics, 332:112– 131, 1973. CODEN LNMAA2. ISBN 3-540-06372-2 (print), 3-540-46980-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060910/.

Anonymous:1973:SPL

[Lel73b] Pierre Lelong, editor. Séminaire Pierre Lelong (Analyse) Année 1971–1972, volume 332 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06372-2 (print), 3-540-46980-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 332. URL http:/

/link.springer.com/book/10.1007/BFb0060904;   http://
www.springerlink.com/content/978-3-540-46980-3.

Lelek:1974:RWC

[Lel74a] A. Lelek. Report on weakly confluent mappings. Lecture Notes in Mathematics, 375:168–170, 1974. CODEN LNMAA2. ISBN 3-540-06684-5 (print), 3-540-37948-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064024/.

Lelong:1974:RTS

[Lel74b] P. Lelong. Remarques sur un théorème de support pour une classe de courants. (French) []. Lecture Notes in Mathematics, 410:97– 106, 1974. CODEN LNMAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066043/.

Lelong:1974:PFT

[Lel74c] Pierre Lelong. Plurisubharmonic functions in topological vector spaces: Polar sets and problems of measure. Lecture Notes in Mathematics, 364:58–68, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069004/.

Lelong:1974:SPL

[Lel74d] Prof. Dr. Pierre Lelong, editor. Séminaire Pierre Lelong (Analyse) Année 1972–1973, volume 410 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 410.

URL http://link.springer.com/book/10.1007/BFb0066034;
 http://www.springerlink.com/content/978-3-540-37816-
7.

Lemaire:1970:JED

[Lem70] Bernard Lemaire. Jeux dans les Équations aux derivées partielles. (French) []. Lecture Notes in Mathematics, 132:181–195, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066682/.

Lemaire:1971:FCW

[Lem71] Jean-Michel Lemaire. A finite complex whose rational homotopy is not finitely generated. Lecture Notes in Mathematics, 196: 114–121, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067754/.

Lemaire:1974:ACH

[Lem74a] Dr. Jean-Michel Lemaire. Algèbres Connexes et Homologie des Espaces de Lacets, volume 422 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06968-2 (print), 3-540-37278-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 134 pp. LCCN QA3 .L28 no. 422.

URL http://link.springer.com/book/10.1007/BFb0070509;
 http://www.springerlink.com/content/978-3-540-37278-
3.

Lemaire:1974:ADC

[Lem74b] Dr. Jean-Michel Lemaire. Application à la determination de certaines algèbres de Pontryagin d’espaces de lacets. (French) []. Lecture Notes in Mathematics, 422:54–74, 1974. CODEN LNMAA2. ISBN 3-540-06968-2 (print), 3-540-37278-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070512/.

Lemaire:1974:HAD

[Lem74c] Dr. Jean-Michel Lemaire. Homologie des algèbres différentielles libres. (French) []. Lecture Notes in Mathematics, 422:18–53, 1974. CODEN LNMAA2. ISBN 3-540-06968-2 (print), 3-540-37278-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070511/.

Lemaire:1974:PF

[Lem74d] Dr. Jean-Michel Lemaire. La propriété P_{□} . (French) []. Lecture Notes in Mathematics, 422:103–130, 1974. CODEN LN-MAA2. ISBN 3-540-06968-2 (print), 3-540-37278-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070514/.

Lemaire:1974:PAC

[Lem74e] Dr. Jean-Michel Lemaire. Présentations d’algèbres connexes. (French) []. Lecture Notes in Mathematics, 422:1–17, 1974. CO-DEN LNMAA2. ISBN 3-540-06968-2 (print), 3-540-37278-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070510/.

Lemaire:1974:CAD

[Lem74f] Dr. Jean-Michel Lemaire. Sur certaines algèbres différentielles libres. (French) []. Lecture Notes in Mathematics, 422:75–102, 1974. CODEN LNMAA2. ISBN 3-540-06968-2 (print), 3-540-37278-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070513/.

Leptin:1972:DVA

[Lep72] Horst Leptin. Darstellung Verallgemeinerter L^{1}-Algebren II. (German) []. Lecture Notes in Mathematics, 247:251–307, 1972. CODEN LNMAA2. ISBN 3-540-05729-3 (print), 3-540-37117-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058553/.

Leroux:1971:SSA

[Ler71] Pierre Leroux. Structure et sémantique abstraites: Extension à des catégories de morphismes d’une paire de foncteurs adjoints. (French) []. Lecture Notes in Mathematics, 195:154–198, 1971. CODEN LNMAA2. ISBN 3-540-05442-1 (print), 3-540-36548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072310/.

Lerman:1973:AOP

[Ler73] Manuel Lerman. Admissible ordinals and priority arguments. Lecture Notes in Mathematics, 337:311–344, 1973. CODEN LN-MAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066780/.

Levin:1971:NVE

[Lev71a] J. J. Levin. On a nonlinear Volterra equation. Lecture Notes in Mathematics, 243:66–75, 1971. CODEN LNMAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058719/.

Levine:1971:ISM

[Lev71b] H. I. Levine. II singularities of mappings. Lecture Notes in Mathematics, 192:22–57, 1971. CODEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066811/.

Levine:1971:IES

[Lev71c] H. I. Levine. III equivalence and stability. Lecture Notes in Mathematics, 192:58–89, 1971. CODEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066812/.

Levine:1971:SDM

[Lev71d] H. I. Levine. Singularities of differentiable mappings. Lecture Notes in Mathematics, 192:1–21, 1971. CODEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066810/.

Levine:1971:BS

[Lev71e] Harold I. Levine. Blowing up singularities. Lecture Notes in Mathematics, 209:90–103, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068894/.

Levine:1973:LCC

[Lev73] Howard A. Levine. Logarithmic convexity and the Cauchy problem for P (t)u_{tt} + M(t)u_{t} + N(t)u = C in Hilbert space. Lecture Notes in Mathematics, 316:102–160, 1973. CODEN LN-MAA2. ISBN 3-540-06159-2 (print), 3-540-38370-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069626/.

Lewis:1972:CAM

[Lew72a] D. J. Lewis. Curves with abnormally many integral points. Lecture Notes in Mathematics, 251:169–180, 1972. CODEN LN-MAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058792/.

Lewis:1972:CCF

[Lew72b] Geoffrey Lewis. Coherence for a closed functor. Lecture Notes in Mathematics, 281:148–195, 1972. CODEN LNMAA2. ISBN 3-540-05963-6 (print), 3-540-37958-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059558/.

Lichtenbaum:1973:VZF

[Lic73] Stephen Lichtenbaum. Values of zeta-functions, étale cohomology, and algebraic K-theory. Lecture Notes in Mathematics, 342: 487–501, 1973. CODEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0073737/.

Lichnerowicz:1974:LLA

[Lic74] André Lichnerowicz. L’algèbre de Lie des automorphismes symplectiques. (French) []. Lecture Notes in Mathematics, 392:163– 174, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064120/.

Lieb:1974:SBE

[Lie74] Ingo Lieb. Solutions bornées des équations de Cauchy–Riemann. (French) []. Lecture Notes in Mathematics, 409:310–326, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068106/.

Linton:1971:MYL

[Lin71] F. E. J. Linton. The multilinear Yoneda lemmas: Toccata, fugue, and fantasia on themes by Eilenberg–Kelly and Yoneda. Lecture Notes in Mathematics, 195:209–229, 1971. CODEN LN-MAA2. ISBN 3-540-05442-1 (print), 3-540-36548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072313/.

Lindner:1974:SFE

[Lin74] Charles C. Lindner. A survey of finite embedding theorems for partial latin squares and quasigroups. Lecture Notes in Mathematics, 406:109–152, 1974. CODEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066437/.

Lions:1970:RMP

[Lio70] J.-L. Lions. Remarques sur la méthode de pénalisation et applications. (French) []. Lecture Notes in Mathematics, 112:21–32, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060195/.

Lions:1971:PUF

[Lio71] Jacques-Louis Lions. Sur les problèmes unilatéraux. (French) []. Lecture Notes in Mathematics, 179:55–77, 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058805/.

Lions:1973:COSb

[Lio73a] J.-L. Lions. Contrôle optimal de systèmes à fonction Coût singulière. (French) []. Lecture Notes in Mathematics, 323:579– 613, 1973. CODEN LNMAA2. ISBN 3-540-06264-5 (print), 3-540-40009-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060536/.

Lions:1973:COSa

[Lio73b] J.-L. Lions. Contrôle optimal de systèmes distribués singuliers. (French) []. Lecture Notes in Mathematics, 323:513–578, 1973. CODEN LNMAA2. ISBN 3-540-06264-5 (print), 3-540-40009-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060535/.

Lions:1973:EES

[Lio73c] J.-L. Lions. Équations d’évolution sans changement de type. (French) []. Lecture Notes in Mathematics, 323:245–357, 1973. CODEN LNMAA2. ISBN 3-540-06264-5 (print), 3-540-40009-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060532/.

Lions:1973:PSP

[Lio73d] J.-L. Lions. Perturbations Singulières dans les Problèmes aux Limites et en Contrôle Optimal. (French) [Singular perturbations in boundary value problems and in optimal control], volume 323 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06264-5 (print), 3-540-40009-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 645 pp. LCCN QA3 .L28 no. 323. URL http:/

/link.springer.com/book/10.1007/BFb0060528;   http://
www.springerlink.com/content/978-3-540-40009-7.

Lions:1973:PEA

[Lio73e] J.-L. Lions. Problèmes d’évolution avec changement de type. (French) []. Lecture Notes in Mathematics, 323:406–512, 1973. CODEN LNMAA2. ISBN 3-540-06264-5 (print), 3-540-40009-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060534/.

Lions:1973:PEE

[Lio73f] J.-L. Lions. Problèmes elliptiques-elliptiques de couches limites. (French) []. Lecture Notes in Mathematics, 323:94–215, 1973. CODEN LNMAA2. ISBN 3-540-06264-5 (print), 3-540-40009-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060530/.

Lions:1973:PREa

[Lio73g] J.-L. Lions. Problèmes raides elliptiques–elliptiques. (French) [Straight elliptic–elliptic problems]. Lecture Notes in Mathematics, 323:1–93, 1973. CODEN LNMAA2. ISBN 3-540-06264-5 (print), 3-540-40009-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060529/.

Lions:1973:PREb

[Lio73h] J.-L. Lions. Problèmes raides elliptiques-elliptiques avec couches limites. (French) []. Lecture Notes in Mathematics, 323:216– 244, 1973. CODEN LNMAA2. ISBN 3-540-06264-5 (print), 3-540-40009-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060531/.

Lions:1973:PSA

[Lio73i] J.-L. Lions. Problèmes stationnaires avec changement de type. (French) []. Lecture Notes in Mathematics, 323:358–405, 1973. CODEN LNMAA2. ISBN 3-540-06264-5 (print), 3-540-40009-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060533/.

Lipschutz:1973:PCC

[Lip73] Seymour Lipschutz. On powers, conjugacy classes and smallcancellation groups. Lecture Notes in Mathematics, 319:126– 132, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058936/.

Lipman:1974:PSF

[Lip74a] Joseph Lipman. Picard schemes of formal schemes; application to rings with discrete divisor class group. Lecture Notes in Mathematics, 412:94–132, 1974. CODEN LNMAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066156/.

Lipsman:1974:CIS

[Lip74b] Ronald L. Lipsman. Chapter I. Introduction to semisimple theory. Lecture Notes in Mathematics, 388:1–44, 1974. CODEN LNMAA2. ISBN 3-540-06790-6 (print), 3-540-38402-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057146/.

Lipsman:1974:CIRa

[Lip74c] Ronald L. Lipsman. Chapter II. Results on induced representations. Lecture Notes in Mathematics, 388:45–66, 1974. CODEN LNMAA2. ISBN 3-540-06790-6 (print), 3-540-38402-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057147/.

Lipsman:1974:CIRb

[Lip74d] Ronald L. Lipsman. Chapter III. Representations of group extensions. Lecture Notes in Mathematics, 388:67–86, 1974. CODEN LNMAA2. ISBN 3-540-06790-6 (print), 3-540-38402-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057148/.

Lipsman:1974:CIN

[Lip74e] Ronald L. Lipsman. Chapter IV. Nilpotent groups. Lecture Notes in Mathematics, 388:87–105, 1974. CODEN LNMAA2. ISBN 3-540-06790-6 (print), 3-540-38402-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057149/.

Lipsman:1974:CVR

[Lip74f] Ronald L. Lipsman. Chapter V. Representations of algebraic groups. Lecture Notes in Mathematics, 388:106–140, 1974. CO-DEN LNMAA2. ISBN 3-540-06790-6 (print), 3-540-38402-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057150/.

Lipsman:1974:CVS

[Lip74g] Ronald L. Lipsman. Chapter VI. Solvable groups. Lecture Notes in Mathematics, 388:141–149, 1974. CODEN LNMAA2. ISBN 3-540-06790-6 (print), 3-540-38402-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057151/.

Lipsman:1974:GRS

[Lip74h] Ronald L. Lipsman. Group Representations: A Survey of Some Current Topics, volume 388 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06790-6 (print), 3-540-38402-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 166 pp. LCCN QA3 .L28 no. 388.

URL http://link.springer.com/book/10.1007/BFb0057145;
 http://www.springerlink.com/content/978-3-540-38402-

Little:1971:GS

[Lit71] J. A. Little. Geometric singularities. Lecture Notes in Mathematics, 209:118–120, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068897/.

Little:1974:EKM

[Lit74a] Charles H. C. Little. An extension of Kasteleyn’s method of enumerating the 1-factors of planar graphs. Lecture Notes in Mathematics, 403:63–72, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057377/.

Litwin:1974:CDV

[Lit74b] J. Litwin. Complexe de Dolbeault sur les variétés de Stein. (French) []. Lecture Notes in Mathematics, 410:67–76, 1974. CO-DEN LNMAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066040/.

Litwin:1974:PLA

[Lit74c] J. Litwin. Problèmes linéaires d’analyse complexe d’après Mitiaguine et Henkine. (French) []. Lecture Notes in Mathematics, 410:28–43, 1974. CODEN LNMAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066038/.

Liulevicius:1971:A

[Liu71] Arunas Liulevicius. On the algebra BP_{∗}(BP). Lecture Notes in Mathematics, 249:47–53, 1971. CODEN LNMAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060895/.

Linnik:1973:ADT

[LKR73] Yu. V. Linnik, L. B. Klebanov, and A. L. Rukhin. On ‘attraction domains’ in the theory of sequential estimation. Lecture Notes in Mathematics, 330:296–303, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061494/.

Langmann:1974:AEA

[LL74a] Klaus Langmann and Werner Lütkebohmert. Algebraische Eigenschaften analytischer Moduln. (German) []. Lecture Notes in Mathematics, 367:79–108, 1974. CODEN LNMAA2. ISBN 3-540-06683-7 (print), 3-540-37945-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068372/.

Langmann:1974:CRH

[LL74b] Klaus Langmann and Werner Lütkebohmert. Cousinverteilungen in Ringen holomorpher Funktionen. (German) []. Lecture Notes in Mathematics, 367:46–78, 1974. CODEN LNMAA2. ISBN 3-540-06683-7 (print), 3-540-37945-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068371/.

Langmann:1974:CF

[LL74c] Klaus Langmann and Werner Lütkebohmert. Cousinverteilungen und Fortsetzungssätze, volume 367 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06683-7 (print), 3-540-37945-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 151 pp. LCCN QA3 .L28 no. 367.

URL http://link.springer.com/book/10.1007/BFb0068368;
 http://www.springerlink.com/content/978-3-540-37945-
4.

Langmann:1974:FUG

[LL74d] Klaus Langmann and Werner Lütkebohmert. Fortsetzungssätze für Untergarben. (German) []. Lecture Notes in Mathematics, 367:109–149, 1974. CODEN LNMAA2. ISBN 3-540-06683-7 (print), 3-540-37945-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068373/.

Langmann:1974:TUG

[LL74e] Klaus Langmann and Werner Lütkebohmert. Theorem A für Unterringe. (German) []. Lecture Notes in Mathematics, 367: 23–45, 1974. CODEN LNMAA2. ISBN 3-540-06683-7 (print), 3-540-37945-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068370/.

Langmann:1974:VGP

[LL74f] Klaus Langmann and Werner Lütkebohmert. Vorbereitungen. (German) [Preliminaries]. Lecture Notes in Mathematics, 367: 1–22, 1974. CODEN LNMAA2. ISBN 3-540-06683-7 (print), 3-540-37945-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068369/.

Lehto:1974:TAC

[LLN74] Prof. Olli Lehto, Prof. Ilppo Simo Louhivaara, and Prof. Rolf Nevanlinna, editors. Topics in Analysis: Colloquium on Mathematical Analysis Jyväskylä 1970, volume 419 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06965-8 (print), 3-540-37907-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 419.

URL http://link.springer.com/book/10.1007/BFb0064704;
 http://www.springerlink.com/content/978-3-540-37907-
2.

Laudet:1970:SAD

[LLNS70] M. Laudet, D. Lacombe, L. Nolin, and M. Schützenberger, editors. Symposium on Automatic Demonstration: Held at Versailles/France, December 1968, volume 125 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 125.

URL http://link.springer.com/book/10.1007/BFb0060620;
 http://www.springerlink.com/content/978-3-540-36262-
3.

Laurent:1970:MDP

[LM70] P.-J. Laurent and B. Martinet. Méthodes duales pour le calcul du minimum d’une fonction convexe sur une intersection de convexes. (French) []. Lecture Notes in Mathematics, 132:159– 180, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066681/.

Lobry:1972:QPG

[Lob72] C. Lobry. Quelques propriétés ‘génériques’ des systèmes à commande. (French) []. Lecture Notes in Mathematics, 280:120– 130, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066923/.

Loday:1973:AAD

[Lod73] Jean-Louis Loday. Applications algébriques du tore dans la sphere et de S^{p} × S^{q} dans S^{p+q}. (French) []. Lecture Notes in Mathematics, 342:79–91, 1973. CODEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073720/.

Loeb:1974:NRB

[Loe74a] Peter A. Loeb. A nonstandard representation of Borel measures and σ-finite measures. Lecture Notes in Mathematics, 369:144– 152, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066008/.

Loeb:1974:NCR

[Loe74b] Peter A. Loeb. A note on continuity for Robinson’s predistributions. Lecture Notes in Mathematics, 369:153–154, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066009.

Loffler:1970:CA

[Lof70] Peter Loffler. The Clifford algebras C^{p,q}. Lecture Notes in Mathematics, 136:1–10, 1970. CODEN LNMAA2. ISBN 3-540-04925-8 (print), 3-540-36289-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059020/.

Lojasiewicz:1971:WFM

[Łoj71] S. Łojasiewicz. Whitney fields and the Malgrange–Mather preparation theorem. Lecture Notes in Mathematics, 192:106–115, 1971. CODEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066815/.

London:1971:EIA

[Lon71] Ralph L. London. Experience with inductive assertions for proving programs correct. Lecture Notes in Mathematics, 188:236– 251, 1971. CODEN LNMAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059700/.

Lorenz:1970:BEG

[Lor70a] Falko Lorenz. Bemerkungen über Erweiterung des Grundkörpers. (German) []. Lecture Notes in Mathematics, 130:52–56, 1970. CODEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069640/.

Lorenz:1970:SWG

[Lor70b] Falko Lorenz. Der ‘Satz 7’ von Witt. (German) []. Lecture Notes in Mathematics, 130:41–42, 1970. CODEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069636.

Lorenz:1970:PGb

[Lor70c] Falko Lorenz. Die Primideale in B(k). (German) []. Lecture Notes in Mathematics, 130:57–64, 1970. CODEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069641/.

Lorenz:1970:HPQ

[Lor70d] Falko Lorenz. Ein ‘Hasse-Prinzip’ für quadratische Formen über reellen Körpern. (German) []. Lecture Notes in Mathematics, 130: 39–40, 1970. CODEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069635.

Lorenz:1970:SCG

[Lor70e] Falko Lorenz. Ein Satz von Cassels. (German) []. Lecture Notes in Mathematics, 130:18–22, 1970. CODEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069631/.

Lorenz:1970:PGa

[Lor70f] Falko Lorenz. Präliminarien. (German) []. Lecture Notes in Mathematics, 130:5–18, 1970. CODEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069630/.

Lorenz:1970:QFF

[Lor70g] Falko Lorenz. Quadratische Formen über Funktionenkörpern mit reell abgeschlossenem Grundkörper. (German) []. Lecture Notes in Mathematics, 130:65–74, 1970. CODEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069642/.

Lorenz:1970:QFK

[Lor70h] Falko Lorenz. Quadratische Formen über Körpern. (German) [Quadratic forms over fields], volume 130 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CO-DEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 77 pp. LCCN QA3 .L28 no. 130. URL http:/

/link.springer.com/book/10.1007/BFb0069629;   http://
www.springerlink.com/content/978-3-540-36270-8.

Lorenz:1970:QFL

[Lor70i] Falko Lorenz. Quadratische Formen über lokalen Körpern. (German) []. Lecture Notes in Mathematics, 130:43–45, 1970. CODEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069637/.

Lorenz:1970:QFRa

[Lor70j] Falko Lorenz. Quadratische Formen über nicht reellen Körpern. (German) []. Lecture Notes in Mathematics, 130:45–50, 1970.

CODEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069638/.

Lorenz:1970:QFRb

[Lor70k] Falko Lorenz. Quadratische Formen über reellen Körpern. (German) []. Lecture Notes in Mathematics, 130:50–52, 1970. CODEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069639/.

Lorenz:1970:QSK

[Lor70l] Falko Lorenz. Quadratsummen und die Stufe eines Körpers. (German) []. Lecture Notes in Mathematics, 130:27–33, 1970. CODEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069633/.

Lorenz:1970:RMF

[Lor70m] Falko Lorenz. Runde und multiplikative Formen. (German) []. Lecture Notes in Mathematics, 130:22–27, 1970. CODEN LN-MAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069632/.

Lorenz:1970:TNN

[Lor70n] Falko Lorenz. Torsionselemente, Nullteiler und nilpotente Elemente im Wittschen Ring. (German) []. Lecture Notes in Mathematics, 130:33–38, 1970. CODEN LNMAA2. ISBN 3-540-04919-3 (print), 3-540-36270-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069634/.

Lorimer:1974:CBD

[Lor74] Peter Lorimer. A class of block designs having the same parameters as the design of points and lines in a projective 3space. Lecture Notes in Mathematics, 403:73–78, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057378/.

Losfeld:1974:IGR

[Los74] Joseph Losfeld. Information généralisée et relation d’ordre. (French) []. Lecture Notes in Mathematics, 398:49–61, 1974. CO-DEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065767/.

Loud:1973:BPS

[Lou73] W. S. Loud. Branching of periodic solutions. Lecture Notes in Mathematics, 312:87–99, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060048/.

Loveland:1970:LFR

[Lov70] D. W. Loveland. A linear format for resolution. Lecture Notes in Mathematics, 125:147–162, 1970. CODEN LNMAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060630/.

Lovasz:1974:MTH

[Lov74] L. Lovász. Minimax theorems for hypergraphs. Lecture Notes in Mathematics, 411:111–126, 1974. CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066186/.

Lowenthal:1973:OG

[Low73] F. Lowenthal. The orders of generation of SL(2, R). Lecture Notes in Mathematics, 319:133–134, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058937.

Lascaux:1971:SPS

[LR71] P. Lascaux and P.-A. Raviart. Stabilité et précision des schémas DSN pour l’Équation de transport en géométrie sphérique. (French) []. Lecture Notes in Mathematics, 228:103–136, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069452/.

Levinson:1972:PCD

[LR72] Henry W. Levinson and Elvira Strasser Rapaport. Planarity of Cayley diagrams. Lecture Notes in Mathematics, 303:183– 188, 1972. CODEN LNMAA2. ISBN 3-540-06096-0 (print), 3-540-38114-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067370/.

Lavrentiev:1970:IHC

[LRV70a] Lavrentiev, Romanov, and Vasiliev. Inverse heat conduction problems with continuously active sources. Lecture Notes in Mathematics, 167:39–50, 1970. CODEN LNMAA2. ISBN 3-540-05282-8 (print), 3-540-36404-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067432/.

Lavrentiev:1970:IPS

[LRV70b] Lavrentiev, Romanov, and Vasiliev. Inverse problems for secondorder elliptic equations. Lecture Notes in Mathematics, 167:51– 56, 1970. CODEN LNMAA2. ISBN 3-540-05282-8 (print), 3-540-36404-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067433/.

Lavrentiev:1970:LID

[LRV70c] Lavrentiev, Romanov, and Vasiliev. Linearized inverse dynamic problem for the telegraph equation. Lecture Notes in Mathematics, 167:22–32, 1970. CODEN LNMAA2. ISBN 3-540-05282-8 (print), 3-540-36404-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067430/.

Lavrentiev:1970:LIK

[LRV70d] Lavrentiev, Romanov, and Vasiliev. Linearized inverse kinematic problem for the wave equation. Lecture Notes in Mathematics, 167:33–38, 1970. CODEN LNMAA2. ISBN 3-540-05282-8 (print), 3-540-36404-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067431/.

Lavrentiev:1970:MIP

[LRV70e] Lavrentiev, Romanov, and Vasiliev. Multidimensional Inverse Problems for Differential Equations, volume 167 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05282-8 (print), 3-540-36404-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 56 (est.) pp. LCCN QA3 .L28 no. 167. URL http:/

/link.springer.com/book/10.1007/BFb0067428;   http://
www.springerlink.com/content/978-3-540-36404-7.

Lavrentiev:1970:SPI

[LRV70f] Lavrentiev, Romanov, and Vasiliev. Some problems of integral geometry. Lecture Notes in Mathematics, 167:1–21, 1970. CO-DEN LNMAA2. ISBN 3-540-05282-8 (print), 3-540-36404-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067429/.

Lepore:1972:SLC

[LS72] John A. Lepore and Robert A. Stoltz. Stability of linear cylindrical shells subjected to stochastic excitations. Lecture Notes in Mathematics, 294:239–251, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064945/.

Lions:1970:EDC

[LT70] J.-L. Lions and R. Temam. Éclatement et décentralisation en calcul des variations. (French) []. Lecture Notes in Mathematics, 132:196–217, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066683/.

Lindenstrauss:1973:CBS

[LT73] Joram Lindenstrauss and Lior Tzafriri. Classical Banach spaces, volume 338 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 0-387-06408-7, 3-540-06408-7. ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 243 pp. LCCN QA3 .L28 no. 338; QA3 .L28 no. 338;

QA322.2. URL http://www.springerlink.com/content/978-

3-540-06408-4. Reprinted in [LT96].

Lindenstrauss:1996:CBS

[LT96] Joram Lindenstrauss and Lior Tzafriri, editors. Classical Banach Spaces I: Sequence Spaces, volume 338. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CO-DEN LNMAA2. ISBN 3-540-60628-9 (print), 3-540-37732-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xiii + 188

pp. LCCN QA3 .L28 no. 338. URL http://www.springerlink.
com/content/978-3-540-37732-0. Reprint of [LT73].

Luckham:1970:RTR

[Luc70] David Luckham. Refinement theorems in resolution theory. Lecture Notes in Mathematics, 125:163–190, 1970. CODEN LN-MAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060631/.

Lucien:1971:CS

[Luc71a] Waelbroeck Lucien. Compactological spaces. Lecture Notes in Mathematics, 230:44–58, 1971. CODEN LNMAA2. ISBN 3-540-05650-5 (print), 3-540-36938-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061237/.

Lucien:1971:CBS

[Luc71b] Waelbroeck Lucien. Completant bounded structures. Lecture Notes in Mathematics, 230:23–43, 1971. CODEN LN-MAA2. ISBN 3-540-05650-5 (print), 3-540-36938-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061236/.

Lucien:1971:DVV

[Luc71c] Waelbroeck Lucien. Differentiable vector-valued functions. Lecture Notes in Mathematics, 230:59–77, 1971. CODEN LN-MAA2. ISBN 3-540-05650-5 (print), 3-540-36938-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061238/.

Lucien:1971:FAF

[Luc71d] Waelbroeck Lucien. Fréchet algebras. (French) []. Lecture Notes in Mathematics, 230:112–117, 1971. CODEN LNMAA2. ISBN 3-540-05650-5 (print), 3-540-36938-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061241/.

Lucien:1971:GMT

[Luc71e] Waelbroeck Lucien. The Gelfand–Mazùr theorem. Lecture Notes in Mathematics, 230:78–92, 1971. CODEN LNMAA2. ISBN 3-540-05650-5 (print), 3-540-36938-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061239/.

Lucien:1971:HFC

[Luc71f] Waelbroeck Lucien. The holomorphic functional calculus. Lecture Notes in Mathematics, 230:93–111, 1971. CODEN LN-MAA2. ISBN 3-540-05650-5 (print), 3-540-36938-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061240/.

Lucien:1971:PF

[Luc71g] Waelbroeck Lucien. A polarization formula. Lecture Notes in Mathematics, 230:118–133, 1971. CODEN LNMAA2. ISBN 3-540-05650-5 (print), 3-540-36938-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061242/.

Lucien:1971:TEC

[Luc71h] Waelbroeck Lucien. Topological extensions of the complex field. Lecture Notes in Mathematics, 230:134–149, 1971. CODEN LN-MAA2. ISBN 3-540-05650-5 (print), 3-540-36938-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061243/.

Luckhardt:1973:CBR

[Luc73a] Horst Luckhardt. The calculus T ∪ BR of the bar recursive functionals. Lecture Notes in Mathematics, 306:76–87, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060879/.

Luckhardt:1973:CPR

[Luc73b] Horst Luckhardt. The calculus T of the primitive recursive functionals. Lecture Notes in Mathematics, 306:59–67, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060877/.

Luckhardt:1973:CPC

[Luc73c] Horst Luckhardt. Consistency proof by computation. Computation of T ∪ BR^{o...o} ∪ μ. Lecture Notes in Mathematics, 306: 98–119, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060882/.

Luckhardt:1973:EE

[Luc73d] Horst Luckhardt. Elimination of extensionality. Lecture Notes in Mathematics, 306:28–40, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060874/.

Luckhardt:1973:EGF

[Luc73e] Horst Luckhardt. Extensional Gödel Functional Interpretation: A Consistency Proof of Classical Analysis, volume 306 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 161 pp. LCCN QA3 .L28 no. 306. URL http:/

/link.springer.com/book/10.1007/BFb0060871;   http://
www.springerlink.com/content/978-3-540-38283-6.

Luckhardt:1973:FSC

[Luc73f] Horst Luckhardt. A formal system of classical analysis. Lecture Notes in Mathematics, 306:13–27, 1973. CODEN LN-MAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060873/.

Luckhardt:1973:FICb

[Luc73g] Horst Luckhardt. Functional interpretation of classical (AC)^{o}- , (ωAC)-analysis with (ER)-qf and functional interpretation in the narrower sense of Heyting-analysis plus (ER)-qf, (MP), (\underline{V}), (Λ)^{o}, (ω AC) in T ∪ BR. Lecture Notes in Mathematics, 306: 88–92, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060880/.

Luckhardt:1973:FICa

[Luc73h] Horst Luckhardt. Functional interpretation of classical arithmetic plus (ER)-qf, (AC)-qf and functional interpretation in the narrower sense of Heyting-analysis plus (ER)-qf, (MP), (\underline{V}) in T. Lecture Notes in Mathematics, 306:68–75, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060878/.

Luckhardt:1973:FCF

[Luc73i] Horst Luckhardt. Further consequences from the functional interpretation of classical analysis. Lecture Notes in Mathematics, 306:93–97, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060881/.

Luckhardt:1973:GBI

[Luc73j] Horst Luckhardt. Generalization of bar induction BI_{D} and the inductive generation processes to trees over species. Lecture Notes in Mathematics, 306:129–133, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060884/.

Luckhardt:1973:GID

[Luc73k] Horst Luckhardt. Generalized inductive definitions. Lecture Notes in Mathematics, 306:120–128, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/BFb0060883/.

Luckhardt:1973:GFI

[Luc73l] Horst Luckhardt. Gödel’s functional interpretation in the narrower sense. Lecture Notes in Mathematics, 306:50–58, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060876/.

Luckhardt:1973:IS

[Luc73m] Horst Luckhardt. Introduction and survey. Lecture Notes in Mathematics, 306:1–12, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060872/.

Luckhardt:1973:M

[Luc73n] Horst Luckhardt. A model for T ∪ BR. Lecture Notes in Mathematics, 306:134–136, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060885/.

Luckhardt:1973:BRM

[Luc73o] Horst Luckhardt. On the bar recursive model of classical analysis and the general bar induction over species. Lecture Notes in Mathematics, 306:137–156, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060886/.

Luckhardt:1973:TCI

[Luc73p] Horst Luckhardt. Translation of classical into intuitionistic approximated theories. Lecture Notes in Mathematics, 306:41– 49, 1973. CODEN LNMAA2. ISBN 3-540-06119-3 (print), 3-540-38283-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060875/.

Lutz:1974:SCC

[Lut74] R. Lutz. Structures de contact en codimension quelconque. (French) []. Lecture Notes in Mathematics, 392:23–29, 1974. CO-DEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064106/.

Lee:1972:ECN

[LW72] Chung N. Lee and Arthur G. Wasserman. Equivariant characteristic numbers. Lecture Notes in Mathematics, 298:191–216, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070043/.

Locher:1973:MVN

[LZ73] F. Locher and K. Zeller. MCC-Verfahren (Numerische Čebyšev-Entwicklung Einer Stammfunktion). (German) []. Lecture Notes in Mathematics, 333:175–182, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060696/.

Maass:1970:SRS

[Maa70] H. Maass. Some remarks on Selberg’s zeta functions. Lecture Notes in Mathematics, 155:122–131, 1970. CODEN LN-MAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060320/.

Maass:1971:AGB

[Maa71a] Hans Maaß. Automorphism groups of bilinear forms. Lecture Notes in Mathematics, 216:8–21, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058627/.

Maass:1971:DO

[Maa71b] Hans Maaß. The differential operator M_{α}. Lecture Notes in Mathematics, 216:309–318, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058644/.

Maass:1971:DSC

[Maa71c] Hans Maaß. Dirichlet series corresponding to modular forms. Lecture Notes in Mathematics, 216:202–219, 1971. CODEN LN-MAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058640/.

Maass:1971:FA

[Maa71d] Hans Maaß. Final aspects. Lecture Notes in Mathematics, 216: 319–325, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058645/.

Maass:1971:FDM

[Maa71e] Hans Maaß. The fundamental domain of the modular group. Lecture Notes in Mathematics, 216:165–180, 1971. CODEN LN-MAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058637/.

Maass:1971:G

[Maa71f] Hans Maaß. A generalization of J_{S}(X). Lecture Notes in Mathematics, 216:84–102, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058632/.

Maass:1971:GRS

[Maa71g] Hans Maaß. Geometry in the representation space. Lecture Notes in Mathematics, 216:22–30, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058628/.

Maass:1971:GCQ

[Maa71h] Hans Maaß. Größen-characters of quadratic forms. Lecture Notes in Mathematics, 216:140–154, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058635/.

Maass:1971:MFD

[Maa71i] Hans Maaß. Modular forms of degree n. Lecture Notes in Mathematics, 216:181–196, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058638/.

Maass:1971:MGD

[Maa71j] Hans Maaß. The modular group of degree n. Lecture Notes in Mathematics, 216:155–164, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058636/.

Maass:1971:NAE

[Maa71k] Hans Maaß. Non-analytic Eisenstein series. Lecture Notes in Mathematics, 216:300–308, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058643/.

Maass:1971:PRT

[Maa71l] Hans Maaß. Preliminary remarks on topological groups. Lecture Notes in Mathematics, 216:1–7, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058626/.

Maass:1971:RTP

[Maa71m] Hans Maaß. The reduction theory of positive quadratic forms. Lecture Notes in Mathematics, 216:122–139, 1971. CODEN LN-MAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058634/.

Maass:1971:RES

[Maa71n] Hans Maaß. Report on Eisenstein series of the modular group. Lecture Notes in Mathematics, 216:197–201, 1971. CODEN LN-MAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058639/.

Maass:1971:RS

[Maa71o] Hans Maaß. The Riemannian space. Lecture Notes in Mathematics, 216:103–121, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058633/.

Maass:1971:RSA

[Maa71p] Hans Maaß. The Riemannian space of all positive matrices. Lecture Notes in Mathematics, 216:63–83, 1971. CODEN LN-MAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058631/.

Maass:1971:SZF

[Maa71q] Hans Maaß. Selberg’s zeta functions. Lecture Notes in Mathematics, 216:253–299, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058642/.

Maass:1971:SMF

[Maa71r] Hans Maaß, editor. Siegel’s Modular Forms and Dirichlet Series: Course Given at the University of Maryland, 1969– 1970, volume 216 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 216. URL http:/

/link.springer.com/book/10.1007/BFb0058625;   http://
www.springerlink.com/content/978-3-540-36881-6.

Maass:1971:SG

[Maa71s] Hans Maaß. Symplectic geometry. Lecture Notes in Mathematics, 216:31–50, 1971. CODEN LNMAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058629/.

Maass:1971:WSR

[Maa71t] Hans Maaß. Weakly symmetric Riemannian spaces. Lecture Notes in Mathematics, 216:51–62, 1971. CODEN LN-MAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058630/.

Maass:1971:ZFA

[Maa71u] Hans Maaß. Zeta functions attached to quadratic forms. Lecture Notes in Mathematics, 216:220–252, 1971. CODEN LN-MAA2. ISBN 3-540-05563-0 (print), 3-540-36881-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058641/.

MacLane:1970:RMC

[MAB^{+}70] S. Mac Lane, H. Applegate, M. Barr, B. Day, E. Dubuc, Phreilambud, A. Pultr, R. Street, M. Tierney, and S. Swierczkowski, editors. Reports of the Midwest Category Seminar IV, volume 137 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04926-6 (print), 3-540-36292-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 137. URL http:/

/link.springer.com/book/10.1007/BFb0060437;   http://
www.springerlink.com/content/978-3-540-36292-0.

MacLane:1970:MBC

[Mac70] Saunders Mac Lane. The Milgram bar construction as a tensor product of functors. Lecture Notes in Mathematics, 168:135– 152, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058523/.

Macintyre:1972:DPD

[Mac72] Angus Macintyre. Direct powers with distinguished diagonal. Lecture Notes in Mathematics, 255:178–203, 1972. CODEN LN-MAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059545/.

MacRae:1973:ERA

[Mac73] R. E. MacRae. On Euclidean rings of algebraic functions. Lecture Notes in Mathematics, 311:167–170, 1973. CODEN LN-MAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068927/.

Mahony:1974:SMM

[Mah74] Louis M. Mahony. The structure of maps from manifolds to spheres. Lecture Notes in Mathematics, 428:144–199, 1974. CO-DEN LNMAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070538/.

Maisonneuve:1971:ERT

[Mai71] B. Maisonneuve. Ensembles régénératifs, temps locaux et subordinateurs. (French) []. Lecture Notes in Mathematics, 191:147– 169, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058856/.

Maisonneuve:1972:TDT

[Mai72] B. Maisonneuve. Topologies du type de Skorohod. (French) []. Lecture Notes in Mathematics, 258:113–117, 1972. CODEN LN-MAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059468/.

Malgrange:1973:OFD

[Mal73] Bernard Malgrange. Opérateurs de Fourier [d’après Hörmander et Maslov]. (French) [Fourier operators [according to Hörmander et Maslov]]. Lecture Notes in Mathematics, 317:219–238, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069284/.

Malraison:1974:HAC

[Mal74] Pierre J. Malraison, Jr. Homotopy associative categories. Lecture Notes in Mathematics, 428:108–131, 1974. CODEN LN-MAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070536/.

Manna:1971:MTP

[Man71] Zohar Manna. Mathematical theory of partial correctness. Lecture Notes in Mathematics, 188:252–269, 1971. CODEN LN-MAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059701/.

Mandl:1972:AIF

[Man72] Petr Mandl. An application of Itô’s formula to stochastic control systems. Lecture Notes in Mathematics, 294:8–13, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064930/.

Mandelbrojt:1973:APF

[Man73a] S. Mandelbrojt. Avant-propos. (French) [Preface]. Lecture Notes in Mathematics, 336:1–4, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065781/.

Mandelbrojt:1973:FPP

[Man73b] S. Mandelbrojt. Fonctions presque périodiques séries adhérentes. (French) []. Lecture Notes in Mathematics, 336:134–143, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065794/.

Marchal:1970:EDS

[Mar70a] Christian Marchal. Étude du sens optimal des commutations dans la théorie de Contensou–Pontryagin. (French) []. Lecture Notes in Mathematics, 112:198–222, 1970. CODEN LN-MAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060205/.

Marec:1970:CER

[Mar70b] Jean-Pierre Marec. Contribution à l’étude des rendez-vous multiimpulsionnels, optimaux, de durée moyenne, entre orbites quasicirculaires, proches, non coplanaires. (French) []. Lecture Notes in Mathematics, 112:159–197, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060204/.

Marec:1970:UEV

[Mar70c] Jean-Pierre Marec. Utilisation des Équations aux variations dans l’Étude des trajectoires spatiales optimales; application aux solutions singulières réticentes. (French) []. Lecture Notes in Mathematics, 132:218–251, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066684/.

Markley:1970:NTM

[Mar70d] Nelson G. Markley. Non-trivial minimal sets — a survey. Lecture Notes in Mathematics, 144:153–159, 1970. CODEN LN-MAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059932/.

Martineau:1970:SFA

[Mar70e] André Martineau. Les supports des fonctionnelles analytiques. (French) []. Lecture Notes in Mathematics, 116:175–195, 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059255/.

Markoe:1971:CAF

[Mar71a] Andrew Markoe. Cohomology of analytic families of differential complexes. Lecture Notes in Mathematics, 184:175–181, 1971. CODEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059284/.

Mars:1971:NTG

[Mar71b] J. G. M. Mars. Les nombres de Tamagawa de groupes semisimples. (French) []. Lecture Notes in Mathematics, 179:79– 94, 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058806/.

Martineau:1971:FAN

[Mar71c] André Martineau. Fonctionnelles analytiques non linéaires et représentation de Polya pour une fonction entière de n variables de type exponentiel. (French) []. Lecture Notes in Mathematics, 205:129–165, 1971. CODEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059837/.

Martinet:1971:FCV

[Mar71d] J. Martinet. Formes de contact sur les Variétés de dimension 3. (French) []. Lecture Notes in Mathematics, 209:142–163, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068901/.

Martinet:1971:CEC

[Mar71e] Jacques Martinet. Un contre-exemple à une conjecture d’E. Noether. (French) []. Lecture Notes in Mathematics, 180:145– 154, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058829/.

Markley:1973:CS

[Mar73a] Nelson G. Markley. Characteristic sequences. Lecture Notes in Mathematics, 318:197–200, 1973. CODEN LNMAA2. ISBN 3-540-06187-8 (print), 3-540-38414-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061736/.

Markus:1973:OCL

[Mar73b] Lawrence Markus. Optimal control of limit cycles or what control theory can do to cure a heart attack or to cause one. Lecture Notes in Mathematics, 312:108–134, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060050/.

Maruyama:1973:AOT

[Mar73c] G. Maruyama. Applications of Ornstein’s theory to stationary processes. Lecture Notes in Mathematics, 330:304–309, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061495/.

Marden:1974:KGD

[Mar74a] Albert Marden. Kleinian groups and 3-dimensional topology a survey. Lecture Notes in Mathematics, 400:108–121, 1974. CO-DEN LNMAA2. ISBN 3-540-06840-6 (print), 3-540-37776-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065679/.

Martinet:1974:CCS

[Mar74b] Jean Martinet. Classes caractéristiques des systèmes de Pfaff. (French) []. Lecture Notes in Mathematics, 392:30–36, 1974. CO-DEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064107/.

Mason:1971:CTI

[Mas71] J. H. Mason. A characterization of transversal independence spaces. Lecture Notes in Mathematics, 211:86–94, 1971. CODEN LNMAA2. ISBN 3-540-05545-2 (print), 3-540-36877-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061077/.

Maskit:1974:GBK

[Mas74a] Bernard Maskit. Good and bad Kleinian groups. Lecture Notes in Mathematics, 400:94–107, 1974. CODEN LNMAA2. ISBN 3-540-06840-6 (print), 3-540-37776-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065678/.

Masur:1974:CTS

[Mas74b] Howard Masur. The curvature of Teichmüller space. Lecture Notes in Mathematics, 400:122–123, 1974. CODEN LN-MAA2. ISBN 3-540-06840-6 (print), 3-540-37776-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065680.

Matsushima:1970:AES

[Mat70] Yozo Matsushima. On automorphisms and equivalences of Siegel domains. Lecture Notes in Mathematics, 155:132–140, 1970. CO-DEN LNMAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060321/.

Mather:1971:NPM

[Mat71a] J. N. Mather. On Nirenberg’s proof of Malgrange’s preparation theorem. Lecture Notes in Mathematics, 192:116–120, 1971. CO-DEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066816/.

Mather:1971:SMV

[Mat71b] J. N. Mather. Stability of C^{∞} mappings: VI the nice dimensions. Lecture Notes in Mathematics, 192:207–253, 1971. CODEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066824/.

Mather:1971:SMG

[Mat71c] John N. Mather. Stable map-germs and algebraic geometry. Lecture Notes in Mathematics, 197:176–193, 1971. CODEN LN-MAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068617/.

Mathias:1972:SPC

[Mat72] A. R. D. Mathias. Solution of problems of Choquet and Puritz. Lecture Notes in Mathematics, 255:204–210, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059546/.

Matlis:1973:DCMb

[Mat73a] Eben Matlis. 1-Dimensional Cohen–Macaulay Rings, volume 327 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 157 pp. LCCN QA3 .L28 no. 327. URL http:/

/link.springer.com/book/10.1007/BFb0061666;   http://
www.springerlink.com/content/978-3-540-46923-0.

Matlis:1973:ADM

[Mat73b] Eben Matlis. Artinian divisible modules. Lecture Notes in Mathematics, 327:42–55, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061671/.

Matlis:1973:CIR

[Mat73c] Eben Matlis. The canonical ideal of R. Lecture Notes in Mathematics, 327:141–152, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061681/.

Matlis:1973:CCR

[Mat73d] Eben Matlis. The closed components of R. Lecture Notes in Mathematics, 327:68–78, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061673/.

Matlis:1973:CE

[Mat73e] Eben Matlis. Compatible extensions. Lecture Notes in Mathematics, 327:28–33, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061669/.

Matlis:1973:C

[Mat73f] Eben Matlis. Completions. Lecture Notes in Mathematics, 327: 11–27, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061668/.

Matlis:1973:FNR

[Mat73g] Eben Matlis. The first neighborhood ring. Lecture Notes in Mathematics, 327:103–115, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061678/.

Matlis:1973:GR

[Mat73h] Eben Matlis. Gorenstein rings. Lecture Notes in Mathematics, 327:116–125, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061679/.

Matlis:1973:DCMa

[Mat73i] Eben Matlis. h-divisible and cotorsion modules. Lecture Notes in Mathematics, 327:1–10, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061667/.

Matlis:1973:IC

[Mat73j] Eben Matlis. The integral closure. Lecture Notes in Mathematics, 327:90–95, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061676/.

Matlis:1973:L

[Mat73k] Eben Matlis. Localizations. Lecture Notes in Mathematics, 327: 34–41, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061670/.

Matlis:1973:M

[Mat73l] Eben Matlis. Multiplicities. Lecture Notes in Mathematics, 327: 126–140, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061680/.

Matlis:1973:PD

[Mat73m] Eben Matlis. The primary decomposition. Lecture Notes in Mathematics, 327:96–102, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061677/.

Matlis:1973:SSU

[Mat73n] Eben Matlis. Semi-simple and uniserial divisible modules. Lecture Notes in Mathematics, 327:85–89, 1973. CODEN LN-MAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061675/.

Matlis:1973:SDM

[Mat73o] Eben Matlis. Simple divisible modules. Lecture Notes in Mathematics, 327:79–84, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061674/.

Matlis:1973:SUR

[Mat73p] Eben Matlis. Strongly unramified ring extensions. Lecture Notes in Mathematics, 327:56–67, 1973. CODEN LNMAA2. ISBN 3-540-06327-7 (print), 3-540-46923-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061672/.

Maumary:1973:PSG

[Mau73] Serge Maumary. Proper surgery groups and wall-Novikov groups. Lecture Notes in Mathematics, 343:526–539, 1973. CODEN LN-MAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061378/.

May:1970:GAA

[May70] J. Peter May. A general algebraic approach to Steenrod operations. Lecture Notes in Mathematics, 168:153–231, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058524/.

May:1972:O

[May72a] J. P. May. A_{∞} and E_{∞} operads. Lecture Notes in Mathematics, 271:19–29, 1972. CODEN LNMAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067494/.

May:1972:SIL

[May72b] J. P. May. E_{∞} spaces and infinite loop sequences. Lecture Notes in Mathematics, 271:137–152, 1972. CODEN LNMAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067505/.

May:1972:AT

[May72c] J. P. May. The approximation theorem. Lecture Notes in Mathematics, 271:50–59, 1972. CODEN LNMAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067497/.

May:1972:CC

[May72d] J. P. May. A categorical construction. Lecture Notes in Mathematics, 271:82–93, 1972. CODEN LNMAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067500/.

May:1972:CQF

[May72e] J. P. May. Cofibrations and quasi-fibrations. Lecture Notes in Mathematics, 271:60–69, 1972. CODEN LNMAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067498/.

May:1972:GR

[May72f] J. P. May. Geometric realization and S_{∗}, C_{∗}, and Ω_{∗}. Lecture Notes in Mathematics, 271:113–125, 1972. CODEN LN-MAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067503/.

May:1972:GRS

[May72g] J. P. May. Geometric realization of simplicial spaces. Lecture Notes in Mathematics, 271:100–112, 1972. CODEN LN-MAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067502/.

May:1972:GIL

[May72h] J. P. May. The Geometry of Iterated Loop Spaces, volume 271 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 161 (est.) pp. LCCN QA3 .L28 no. 271. URL http:/

/link.springer.com/book/10.1007/BFb0067491;   http://
www.springerlink.com/content/978-3-540-37603-3.

May:1972:ILS

[May72i] J. P. May. Iterated loop spaces and the. Lecture Notes in Mathematics, 271:39–49, 1972. CODEN LNMAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067496/.

May:1972:LCO

[May72j] J. P. May. The little cubes operads. Lecture Notes in Mathematics, 271:30–38, 1972. CODEN LNMAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067495/.

May:1972:MC

[May72k] J. P. May. Monoidal categories. Lecture Notes in Mathematics, 271:94–99, 1972. CODEN LNMAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067501/.

May:1972:OM

[May72l] J. P. May. Operads and monads. Lecture Notes in Mathematics, 271:10–18, 1972. CODEN LNMAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067493/.

May:1972:OS

[May72m] J. P. May. Operads and-spaces. Lecture Notes in Mathematics, 271:1–9, 1972. CODEN LNMAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067492/.

May:1972:RPS

[May72n] J. P. May. The recognition principle and A_{∞} spaces. Lecture Notes in Mathematics, 271:126–136, 1972. CODEN LN-MAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067504/.

May:1972:RCR

[May72o] J. P. May. Remarks concerning the recognition principle. Lecture Notes in Mathematics, 271:153–161, 1972. CODEN LN-MAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067506/.

May:1972:SCP

[May72p] J. P. May. The smash and composition products. Lecture Notes in Mathematics, 271:70–81, 1972. CODEN LNMAA2. ISBN 3-540-05904-0 (print), 3-540-37603-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067499/.

Mayne:1972:PIL

[May72q] D. Q. Mayne. Parametrization and identification of linear multivariable systems. Lecture Notes in Mathematics, 294:56–61, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064934/.

Mazet:1971:NGA

[Maz71] P. Mazet. Nullstellensatz en géométrie analytique banachique. (French) []. Lecture Notes in Mathematics, 205:71–86, 1971. CO-DEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059834/.

Mazet:1972:PVV

[Maz72] P. Mazet. Propriétés de θ (e) vis-a-vis des décompositions primaires. (French) []. Lecture Notes in Mathematics, 275:71– 81, 1972. CODEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058492/.

Mazur:1973:CES

[Maz73] Barry Mazur. Courbes elliptiques et symboles modulaires. (French) []. Lecture Notes in Mathematics, 317:277–294, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069287/.

Mazet:1974:TID

[Maz74] P. Mazet. Un théorème d’image directe propre. (French) []. Lecture Notes in Mathematics, 410:107–116, 1974. CODEN LN-MAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066044/.

Miele:1970:MGM

[MC70] A. Miele and J. W. Cantrell. Memory gradient method for the minimization of functions. Lecture Notes in Mathematics, 132: 252–263, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066685/.

McAuley:1970:COS

[McA70] Louis F. McAuley. Concerning open selections. Lecture Notes in Mathematics, 171:49–53, 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069720/.

McAuley:1971:COO

[McA71] Louis F. McAuley. Concerning one-to-one continuous images of the reals. Lecture Notes in Mathematics, 235:86–92, 1971. CO-DEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059200/.

McAuley:1974:AGM

[McA74a] Dr. Louis F. McAuley, editor. Algebraic and Geometrical Methods in Topology: Conference on Topological Methods in Algebraic Topology SUNY Binghamton, October 3–7, 1973, volume 428 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 428. URL http:/

/link.springer.com/book/10.1007/BFb0070530;   http://
www.springerlink.com/content/978-3-540-37300-1.

McAvaney:1974:CST

[McA74b] K. L. McAvaney. Counting stable trees. Lecture Notes in Mathematics, 403:79–85, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057379/.

McCann:1971:LSS

[McC71] Roger C. McCann. Local sections for simultaneous local dynamical systems. Lecture Notes in Mathematics, 235:93–102, 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059201/.

McDowell:1974:WJG

[McD74] R. H. McDowell. The works of J. de Groot. Lecture Notes in Mathematics, 378:1–15, 1974. CODEN LNMAA2. ISBN 3-540-06741-8 (print), 3-540-38323-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068456/.

McGehee:1972:HS

[McG72] O. Carruth McGehee. Helson sets in T^{n}. Lecture Notes in Mathematics, 266:229–237, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059647/.

McGehee:1973:SMT

[McG73] Richard McGehee. The stable manifold theorem via an isolating block. Lecture Notes in Mathematics, 312:135–144, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060051/.

Meghea:1971:AHF

[Meg71a] Constantin Meghea. Applications harmoniques. (French) [Harmonic applications]. Lecture Notes in Mathematics, 222:95– 105, 1971. CODEN LNMAA2. ISBN 3-540-05579-7 (print), 3-540-36902-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061005/.

Meghea:1971:CEHb

[Meg71b] Constantin Meghea. Compactification des Espaces Harmoniques, volume 222 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05579-7 (print), 3-540-36902-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 108 pp. LCCN QA3 .L28 no. 222. URL http:

//link.springer.com/book/10.1007/BFb0060998;   http:/
/www.springerlink.com/content/978-3-540-36902-8.

Meghea:1971:CEHa

[Meg71c] Constantin Meghea. Compactification des espaces harmoniques. (French) []. Lecture Notes in Mathematics, 222:53–81, 1971. CO-DEN LNMAA2. ISBN 3-540-05579-7 (print), 3-540-36902-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061003/.

Meghea:1971:FHF

[Meg71d] Constantin Meghea. Fonctions harmonisables. (French) []. Lecture Notes in Mathematics, 222:26–52, 1971. CODEN LN-MAA2. ISBN 3-540-05579-7 (print), 3-540-36902-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061002/.

Meghea:1971:IF

[Meg71e] Constantin Meghea. Introduction. (French) []. Lecture Notes in Mathematics, 222:1–2, 1971. CODEN LNMAA2. ISBN 3-540-05579-7 (print), 3-540-36902-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060999.

Meghea:1971:CWF

[Meg71f] Constantin Meghea. Le compactifié de Wiener. (French) []. Lecture Notes in Mathematics, 222:82–94, 1971. CODEN LN-MAA2. ISBN 3-540-05579-7 (print), 3-540-36902-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061004/.

Meghea:1971:LG

[Meg71g] Constantin Meghea. Leitfaden. (German) []. Lecture Notes in Mathematics, 222:3–4, 1971. CODEN LNMAA2. ISBN 3-540-05579-7 (print), 3-540-36902-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061000.

Meghea:1971:PDF

[Meg71h] Constantin Meghea. Problème de Dirichlet. (French) []. Lecture Notes in Mathematics, 222:5–25, 1971. CODEN LNMAA2. ISBN 3-540-05579-7 (print), 3-540-36902-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061001/.

Meis:1972:VHO

[Mei72] Theodor Meis. Ein Verfahren höherer Ordnung zur Diskretisierung von nichtlinearen, selbstadjungierten Randwertproblemen.

(German) [A higher-order method for the discretization of nonlinear, self-adjoint boundary-value problems]. Lecture Notes in Mathematics, 267:289–304, 1972. CODEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061623/.

Meisen:1974:RRC

[Mei74] Jeanne Meisen. Relations in regular categories. Lecture Notes in Mathematics, 418:96–102, 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070645/.

Mesarovic:1971:MTG

[Mes71a] M. D. Mesarović. Mathematical theory of general systems. Lecture Notes in Mathematics, 206:14–15, 1971. CODEN LN-MAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070125.

Mesarovic:1971:MTM

[Mes71b] M. D. Mesarović. Mathematical theory of multi-level systems. Lecture Notes in Mathematics, 206:15, 1971. CODEN LN-MAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070126.

Messing:1972:C

[Mes72a] William Messing. Conventions. Lecture Notes in Mathematics, 264:10, 1972. CODEN LNMAA2. ISBN 3-540-05840-0 (print), 3-540-37475-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058303.

Messing:1972:CABa

[Mes72b] William Messing. The crystals associated to Barsotti–Tate groups. Lecture Notes in Mathematics, 264:112–149, 1972. CO-DEN LNMAA2. ISBN 3-540-05840-0 (print), 3-540-37475-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058307/.

Messing:1972:CABb

[Mes72c] William Messing. The Crystals Associated to Barsotti–Tate Groups: with Applications to Abelian Schemes, volume 264 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05840-0 (print), 3-540-37475-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 190 pp. LCCN QA3 .L28 no. 264. URL http:/

/link.springer.com/book/10.1007/BFb0058301;   http://
www.springerlink.com/content/978-3-540-37475-6.

Messing:1972:DE

[Mes72d] William Messing. Definitions and examples. Lecture Notes in Mathematics, 264:11–22, 1972. CODEN LNMAA2. ISBN 3-540-05840-0 (print), 3-540-37475-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058304/.

Messing:1972:DTA

[Mes72e] William Messing. The deformation theory and applications. Lecture Notes in Mathematics, 264:150–174, 1972. CODEN LN-MAA2. ISBN 3-540-05840-0 (print), 3-540-37475-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058308/.

Messing:1972:DPE

[Mes72f] William Messing. Divided powers, exponentials and crystals. Lecture Notes in Mathematics, 264:77–111, 1972. CODEN LN-MAA2. ISBN 3-540-05840-0 (print), 3-540-37475-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058306/.

Messing:1972:I

[Mes72g] William Messing. Introduction. Lecture Notes in Mathematics, 264:1–9, 1972. CODEN LNMAA2. ISBN 3-540-05840-0 (print), 3-540-37475-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058302/.

Messing:1972:RBB

[Mes72h] William Messing. The relation between Barsotti–Tate and formal Lie groups. Lecture Notes in Mathematics, 264:23–76, 1972.

CODEN LNMAA2. ISBN 3-540-05840-0 (print), 3-540-37475-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058305/.

Meuer:1973:VZL

[Meu73] Hans-Werner Meuer. Eine Variante Des Zweischritt-Lax– Wendroff-Verfahrens. (German) []. Lecture Notes in Mathematics, 333:183–201, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060697/.

Meyer:1970:QIM

[Mey70a] P.-A. Meyer. Quelques inégalités sur les martingales, d’après Dubins et Freedman. (French) []. Lecture Notes in Mathematics, 124:162–169, 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059341/.

Meyer:1970:SPI

[Mey70b] Paul-André Meyer. Séminaire de Probabilités IV Université de Strasbourg, volume 124 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 282 pp. LCCN QA3 .L28 no. 124.

URL http://link.springer.com/book/10.1007/BFb0059328;
 http://www.springerlink.com/content/978-3-540-36260-
9.

Meyer:1970:EEH

[Mey70c] Yves Meyer. Étude des ensembles harmonieux dans le cas réel. (French) []. Lecture Notes in Mathematics, 117:14–23, 1970. CO-DEN LNMAA2. ISBN 3-540-04906-1 (print), 3-540-36243-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069683/.

Meyer:1970:IFI

[Mey70d] Yves Meyer. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 117:5–6, 1970. CODEN LNMAA2. ISBN 3-540-04906-1 (print), 3-540-36243-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069681.

Meyer:1970:RMF

[Mey70e] Yves Meyer. La répartition modulo 1. (French) []. Lecture Notes in Mathematics, 117:24–38, 1970. CODEN LNMAA2. ISBN 3-540-04906-1 (print), 3-540-36243-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069684/.

Meyer:1970:NPN

[Mey70f] Yves Meyer. Nombres de Pisot, Nombres de Salem et Analyse Harmonique, volume 117 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04906-1 (print), 3-540-36243-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 62 (est.) pp. LCCN QA3 .L28 no. 117.

URL http://link.springer.com/book/10.1007/BFb0069680;
 http://www.springerlink.com/content/978-3-540-36243-
2.

Meyer:1970:RAP

[Mey70g] Yves Meyer. Retour au problème de la synthèse. (French) []. Lecture Notes in Mathematics, 117:54–62, 1970. CODEN LN-MAA2. ISBN 3-540-04906-1 (print), 3-540-36243-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069686/.

Meyer:1970:TGE

[Mey70h] Yves Meyer. Théorie générale des ensembles harmonieux. (French) []. Lecture Notes in Mathematics, 117:6–13, 1970. CO-DEN LNMAA2. ISBN 3-540-04906-1 (print), 3-540-36243-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069682/.

Meyer:1970:EFC

[Mey70i] Yves Meyer. Un espace de fonctions continues. (French) []. Lecture Notes in Mathematics, 117:39–53, 1970. CODEN LN-MAA2. ISBN 3-540-04906-1 (print), 3-540-36243-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069685/.

Meyer:1971:DST

[Mey71a] P.-A. Meyer. Démonstration simplifiée d’un théorème de Knight. (French) []. Lecture Notes in Mathematics, 191:191–195, 1971.

CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058859/.

Meyer:1971:PPP

[Mey71b] P.-A. Meyer. Processus de Poisson ponctuels, d’après K. Itô. (French) []. Lecture Notes in Mathematics, 191:177–190, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058858/.

Meyer:1971:RIF

[Mey71c] P.-A. Meyer. Représentation intégrale des fonctions excessives résultats de Mokobodzki. (French) []. Lecture Notes in Mathematics, 191:196–208, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058860/.

Meyer:1971:ADF

[Mey71d] P.-A. Meyer. Sur un article de Dubins. (French) []. Lecture Notes in Mathematics, 191:170–176, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058857/.

Meyer:1971:TRT

[Mey71e] P.-A. Meyer. Un théorème sur la répartition des temps locaux. (French) []. Lecture Notes in Mathematics, 191:209–210, 1971. CODEN LNMAA2. ISBN 3-540-05397-2 (print), 3-540-36517-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058861.

Meyer:1971:NSI

[Mey71f] Paul A. Meyer. Non-square integrable martingales etc. Lecture Notes in Mathematics, 190:38–43, 1971. CODEN LN-MAA2. ISBN 3-540-05396-4 (print), 3-540-36515-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065890/.

Meyer:1971:SIM

[Mey71g] Paul A. Meyer. Square integrable martingales, a survey. Lecture Notes in Mathematics, 190:32–37, 1971. CODEN LN-MAA2. ISBN 3-540-05396-4 (print), 3-540-36515-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065889/.

Meyer:1971:DPI

[Mey71h] Paul-André Meyer. Démonstration probabiliste d’une identité de convolution. (French) []. Lecture Notes in Mathematics, 179: 245–259, 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058816/.

Meyer:1972:MHF

[Mey72a] P.-A. Meyer. La mesure de H. Föllmer en théorie des surmartingales. (French) []. Lecture Notes in Mathematics, 258: 118–129, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059469/.

Meyer:1972:RRB

[Mey72b] P.-A. Meyer. Les résultats récents de Burkholder, Davis et Gundy. (French) []. Lecture Notes in Mathematics, 258:151– 158, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059471/.

Meyer:1972:TAA

[Mey72c] P.-A. Meyer. Temps d’arrêt algébriquement prévisibles. (French) []. Lecture Notes in Mathematics, 258:159–163, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059472/.

Meyer:1972:NTH

[Mey72d] P.-A. Meyer. Une note sur le théorème de Hunt. (French) []. Lecture Notes in Mathematics, 258:164–167, 1972. CODEN LN-MAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059473/.

Meyer:1972:BD

[Mey72e] Paul-André Meyer. Basic definitions. Lecture Notes in Mathematics, 284:1–23, 1972. CODEN LNMAA2. ISBN 3-540-05983-0 (print), 3-540-37968-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069774/.

Meyer:1972:DMT

[Mey72f] Paul-André Meyer. Discrete martingale theory. Lecture Notes in Mathematics, 284:24–73, 1972. CODEN LNMAA2. ISBN 3-540-05983-0 (print), 3-540-37968-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069775/.

Meyer:1972:MSI

[Mey72g] Paul-André Meyer. Martingales and Stochastic Integrals I, volume 284 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05983-0 (print), 3-540-37968-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 73 (est.) pp. LCCN QA3 .L28 no. 284. URL http:

//link.springer.com/book/10.1007/BFb0069773;   http:/
/www.springerlink.com/content/978-3-540-37968-3.

Meyer:1972:RAS

[Mey72h] Yves Meyer. Recent advances in spectral synthesis. Lecture Notes in Mathematics, 266:239–253, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059648/.

Meyer:1973:SLT

[Mey73a] K. R. Meyer. Stability of a Lurie type equation. Lecture Notes in Mathematics, 312:145–150, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060052/.

Meyer:1973:AEP

[Mey73b] P.-A. Meyer. Applications de l’expose precedent aux processus de Markov. (French) []. Lecture Notes in Mathematics, 321:172– 179, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071405/.

Meyer:1973:CPM

[Mey73c] P.-A. Meyer. Chirurgie sur un processus de Markov d’après Knight et Pittenger. (French) []. Lecture Notes in Mathematics, 321:146–154, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0071403/.

Meyer:1973:DEB

[Mey73d] P.-A. Meyer. Le dual de ‘H’ est ‘BMO’ (cas continu). (French) []. Lecture Notes in Mathematics, 321:136–145, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071402/.

Meyer:1973:LMA

[Mey73e] P.-A. Meyer. Limites médiales, d’après Mokobodzki. (French) [Medial limits, according to Mokobodzki]. Lecture Notes in Mathematics, 321:198–204, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071407/.

Meyer:1973:PPM

[Mey73f] P.-A. Meyer. Présentation des processus de Markov. (French) []. Lecture Notes in Mathematics, 307:165–198, 1973. CODEN LNMAA2. ISBN 3-540-06126-6 (print), 3-540-38321-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059709/.

Meyer:1973:RDT

[Mey73g] P.-A. Meyer. Résultats d’Azéma en ‘théorie générale des processus’. (French) []. Lecture Notes in Mathematics, 321:180– 197, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0071406/.

Meyer:1973:MAP

[Mey73h] P.-A. Meyer. Une mise au point sur les systèmes de Lévy. Remarques sur l’expose de A. Benveniste. (French) []. Lecture Notes in Mathematics, 321:25–32, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0071390/.

Meyer:1973:PMM

[Mey73i] Yves Meyer. Sur un problème de M. Mandelbrojt. (French) []. Lecture Notes in Mathematics, 336:161–167, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065798/.

Meyer:1974:QPC

[Mey74a] J. C. Meyer. Quelques problémes concernant les cliques des hypergraphes h-complets et q-parti h-complets. (French) []. Lecture Notes in Mathematics, 411:127–139, 1974. CODEN LN-MAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066187/.

Meyer:1974:EAMa

[Mey74b] P.-A. Meyer. Ensembles aléatoires markoviens homogènes I. (French) []. Lecture Notes in Mathematics, 381:176–190, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057263/.

Meyer:1974:EAMb

[Mey74c] P.-A. Meyer. Ensembles aléatoires markoviens homogènes (II). (French) []. Lecture Notes in Mathematics, 381:191–211, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057264/.

Meyer:1974:EAMc

[Mey74d] P.-A. Meyer. Ensembles aléatoires markoviens homogènes (V). (French) []. Lecture Notes in Mathematics, 381:242–261, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057267/.

Meyer:1974:TAR

[Mey74e] P.-A. Meyer. Les travaux d’Azéma sur le retournement du temps. (French) []. Lecture Notes in Mathematics, 381:262–288, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057268/.

Meyer:1974:TPM

[Mey74f] P.-A. Meyer. Transformation des processus de Markov. (French) []. Lecture Notes in Mathematics, 390:1–36, 1974. CODEN LNMAA2. ISBN 3-540-06816-3 (print), 3-540-37765-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0082433/.

Meyer:1974:NCR

[Mey74g] P.-A. Meyer. Une note sur la compactification de Ray. (French) []. Lecture Notes in Mathematics, 381:289, 1974. CODEN LN-MAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057269.

Meyer:1974:TDL

[Mey74h] Paul-André Meyer. Le théorème de dérivation de Lebesgue par rapport à une resolvante. (French) []. Lecture Notes in Mathematics, 383:88–97, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057303/.

Michael:1970:SCS

[Mic70] E. Michael. A survey of continuous selections. Lecture Notes in Mathematics, 171:54–58, 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069721/.

Michler:1972:BCG

[Mic72] G. O. Michler. Blocks and centers of group algebras. Lecture Notes in Mathematics, 246:429–563, 1972. CODEN LN-MAA2. ISBN 3-540-05760-9 (print), 3-540-37181-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059569/.

Micula:1974:NLN

[Mic74] Gh. Micula. Die numerische Lösung nichtlinearer Differentialgleichungen unter Verwendung von Spline-Funktionen. (German) []. Lecture Notes in Mathematics, 395:57–83, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060664/.

Milgram:1971:SPB

[Mil71] R. James Milgram. Surgery, B_{P L}, B_{T OP} and the PL bordism rings. Lecture Notes in Mathematics, 249:54–74, 1971. CODEN LNMAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060896/.

Miller:1973:NUC

[Mil73a] K. Miller. Non-unique continuation for certain ODE’s in Hilbert space and for uniformly parabolic and elliptic equations in selfadjoint divergence form. Lecture Notes in Mathematics, 316:85– 101, 1973. CODEN LNMAA2. ISBN 3-540-06159-2 (print), 3-540-38370-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069625/.

Miller:1973:SQR

[Mil73b] K. Miller. Stabilized quasi-reversibility and other nearly-best-possible methods for non-well-posed problems. Lecture Notes in Mathematics, 316:161–176, 1973. CODEN LNMAA2. ISBN 3-540-06159-2 (print), 3-540-38370-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069627/.

Milgram:1974:CG

[Mil74a] R. James Milgram. Calculating the groups. Lecture Notes in Mathematics, 368:70–74, 1974. CODEN LNMAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070465/.

Milgram:1974:CSH

[Mil74b] R. James Milgram. Calculations of the stable homotopy of K(π, n)’s. Lecture Notes in Mathematics, 368:75–80, 1974. CO-DEN LNMAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070466/.

Milgram:1974:C

[Mil74c] R. James Milgram. The cohomology of the F_{n}. Lecture Notes in Mathematics, 368:37–43, 1974. CODEN LNMAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070460/.

Milgram:1974:EME

[Mil74d] R. James Milgram. An example for the metastable exact sequence. Lecture Notes in Mathematics, 368:91–97, 1974. CODEN LNMAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070468/.

Milgram:1974:FCS

[Mil74e] R. James Milgram. Further calculations for some truncated projective spaces. Lecture Notes in Mathematics, 368:98–106, 1974. CODEN LNMAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070469/.

Milgram:1974:Ib

[Mil74f] R. James Milgram. The inclusion X → Ω^{n}Σ^{n}X. Lecture Notes in Mathematics, 368:26–34, 1974. CODEN LNMAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070458/.

Milgram:1974:Ia

[Mil74g] R. James Milgram. Introduction. Lecture Notes in Mathematics, 368:1–10, 1974. CODEN LNMAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070456/.

Milgram:1974:ILS

[Mil74h] R. James Milgram. Iterated loop spaces. Lecture Notes in Mathematics, 368:11–25, 1974. CODEN LNMAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070457/.

Milgram:1974:LSF

[Mil74i] R. James Milgram. The loop space functor for resolutions. Lecture Notes in Mathematics, 368:56–61, 1974. CODEN LN-MAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070463/.

Milgram:1974:M

[Mil74j] R. James Milgram. The map F_{n}(X) → ΩF_{n−1}(ΣX). Lecture Notes in Mathematics, 368:34–37, 1974. CODEN LN-MAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070459/.

Milgram:1974:MES

[Mil74k] R. James Milgram. The metastable exact sequence. Lecture Notes in Mathematics, 368:61–70, 1974. CODEN LNMAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070464/.

Milgram:1974:SCS

[Mil74l] R. James Milgram. Some calculations of the stable homotopy groups for the K(Z, n). Lecture Notes in Mathematics, 368:81– 90, 1974. CODEN LNMAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070467/.

Milgram:1974:SIL

[Mil74m] R. James Milgram. The structure of iterated loop spaces in the metastable range. Lecture Notes in Mathematics, 368:44– 50, 1974. CODEN LNMAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070461/.

Milgram:1974:UAS

[Mil74n] R. James Milgram. An unstable Adams spectral sequence. Lecture Notes in Mathematics, 368:51–56, 1974. CODEN LN-MAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070462/.

Milgram:1974:UHS

[Mil74o] R. James Milgram. Unstable Homotopy from the Stable Point of View, volume 368 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06655-1 (print), 3-540-37925-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 109 pp. LCCN QA3 .L28 no. 368. URL http:/

/link.springer.com/book/10.1007/BFb0070455;   http://
www.springerlink.com/content/978-3-540-37925-6.

Miller:1974:NDG

[Mil74p] R. H. Miller. Numerical difficulties with the gravitational nbody problem. Lecture Notes in Mathematics, 362:260–275, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066595/.

Mislin:1971:GS

[Mis71a] Guido Mislin. The genus of an H-space. Lecture Notes in Mathematics, 249:75–83, 1971. CODEN LNMAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060897/.

Mislin:1971:S

[Mis71b] Guido Mislin. H-spaces modp (I). Lecture Notes in Mathematics, 196:5–10, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067739/.

Mislin:1974:NGF

[Mis74] Guido Mislin. Nilpotent groups with finite commutator subgroups. Lecture Notes in Mathematics, 418:103–120, 1974. CO-DEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070646/.

Mitteau:1974:EAH

[Mit74a] Jean-Claude Mitteau. Existence des applications harmoniques. (French) [Existence of harmonic applications]. Lecture Notes in Mathematics, 392:90–96, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064115/.

Mittelmann:1974:FEV

[Mit74b] H. D. Mittelmann. Finite-Element Verfahren bei quasi-linearen elliptischen randwertproblemen. (German) []. Lecture Notes in Mathematics, 395:199–214, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060672/.

Moyer:1970:CPE

[MK70] H. Gardner Moyer and Henry J. Kelley. Conjugate points on extremal rocket paths. Lecture Notes in Mathematics, 112:223– 224, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060206.

Maisonneuve:1970:TLP

[MM70] B. Maisonneuve and Ph. Morando. Temps locaux pour les ensembles régénératifs. (French) []. Lecture Notes in Mathematics, 124: 151–161, 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059340/.

Malliavin:1972:CSA

[MM72] Marie-Paule Malliavin and Paul Malliavin. Calcul symbolique sur l’algèbre de Wiener centrale de su(3). (French) []. Lecture Notes in Mathematics, 266:215–228, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059646/.

Maisonneuve:1974:EAMa

[MM74a] B. Maisonneuve and P.-A. Meyer. Ensembles aléatoires markoviens homogènes (III). (French) []. Lecture Notes in Mathematics, 381:212–226, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057265/.

Maisonneuve:1974:EAM

[MM74b] B. Maisonneuve and P.-A. Meyer. Ensembles aléatoires markoviens homogènes. introduction et bibliographie. (French) [Homogeneous Markov random sets. Introduction and bibliography]. Lecture Notes in Mathematics, 381:172–175, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057262/.

Maisonneuve:1974:EAMb

[MM74c] B. Maisonneuve and P.-A. Meyer. Ensembles aléatoires markoviens homogènes (IV). (French) []. Lecture Notes in Mathematics, 381:227–241, 1974. CODEN LNMAA2. ISBN 3-540-06783-3 (print), 3-540-38384-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057266/.

Malliavin:1974:FLL

[MM74d] Marie-Paule Malliavin and Paul Malliavin. Factorisations et lois limites de la diffusion horizontale au-dessus d’un espace riemannien symétrique. (French) []. Lecture Notes in Mathematics, 404: 164–217, 1974. CODEN LNMAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060616/.

Mazur:1974:ECU

[MM74e] Barry Mazur and William Messing. Explicit constructions of universal extensions. Lecture Notes in Mathematics, 370:1–59, 1974. CODEN LNMAA2. ISBN 3-540-06659-4 (print), 3-540-37933-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061629/.

Mazur:1974:UEC

[MM74f] Barry Mazur and William Messing. Universal extensions and crystals. Lecture Notes in Mathematics, 370:60–132, 1974. CO-DEN LNMAA2. ISBN 3-540-06659-4 (print), 3-540-37933-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061630/.

Mazur:1974:UEO

[MM74g] Barry Mazur and William Messing. Universal Extensions and One Dimensional Crystalline Cohomology, volume 370 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06659-4 (print), 3-540-37933-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 134 pp. LCCN QA3 .L28 no. 370. URL http:/

/link.springer.com/book/10.1007/BFb0061628;   http://
www.springerlink.com/content/978-3-540-37933-1.

Mitchell:1974:CEF

[MM74h] A. R. Mitchell and R. McLeod. Curved elements in the finite element method. Lecture Notes in Mathematics, 363:89–104, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069128/.

Moiseev:1970:CMO

[Moi70] N. N. Moiseev, editor. Colloquium on Methods of Optimization: Held in Novosibirsk/USSR, June 1968, volume 112 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 112.

URL http://link.springer.com/book/10.1007/BFb0060193;
 http://www.springerlink.com/content/978-3-540-36204-
3.

Moishezon:1974:MCV

[Moi74] Boris Moishezon. Modifications of complex varieties and the Chow lemma. Lecture Notes in Mathematics, 412:133–139, 1974. CODEN LNMAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066157/.

Mokobodzki:1970:DRD

[Mok70a] Gabriel Mokobodzki. Densité relative de deux potentiels comparables. (French) []. Lecture Notes in Mathematics, 124:170– 194, 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059342/.

Mokobodzki:1970:QPR

[Mok70b] Gabriel Mokobodzki. Quelques propriétés remarquables des opérateurs presque positifs. (French) []. Lecture Notes in Mathematics, 124:195–207, 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059343/.

Mokobodzki:1971:SCP

[Mok71] Gabriel Mokobodzki. Structure des cônes de potentiels. (French) []. Lecture Notes in Mathematics, 180:239–252, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058834/.

Mokobodzki:1972:PQD

[Mok72] Gabriel Mokobodzki. Pseudo-quotient de deux mesures par rapport à un cône de potentiels application à la dualité. (French) []. Lecture Notes in Mathematics, 258:173–176, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059475/.

Montgomery:1971:AFL

[Mon71a] Hugh L. Montgomery. Arithmetic formulations of the large sieve. Lecture Notes in Mathematics, 227:23–30, 1971. CODEN LN-MAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060854/.

Montgomery:1971:BVT

[Mon71b] Hugh L. Montgomery. The Bombieri–Vinogradov theorem. Lecture Notes in Mathematics, 227:133–140, 1971. CODEN LN-MAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060866/.

Montgomery:1971:CMV

[Mon71c] Hugh L. Montgomery. Classical mean value theorems. Lecture Notes in Mathematics, 227:50–57, 1971. CODEN LN-MAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060857/.

Montgomery:1971:DZF

[Mon71d] Hugh L. Montgomery. Distribution of zeros of L-functions. Lecture Notes in Mathematics, 227:95–116, 1971. CODEN LN-MAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060863/.

Montgomery:1971:FRC

[Mon71e] Hugh L. Montgomery. Further results and conjectures concerning mean and large moduli. Lecture Notes in Mathematics, 227: 69–73, 1971. CODEN LNMAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060860/.

Montgomery:1971:LMT

[Mon71f] Hugh L. Montgomery. Large moduli theorems. Lecture Notes in Mathematics, 227:64–68, 1971. CODEN LNMAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060859/.

Montgomery:1971:LS

[Mon71g] Hugh L. Montgomery. The large sieve. Lecture Notes in Mathematics, 227:11–22, 1971. CODEN LNMAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060853/.

Montgomery:1971:LCN

[Mon71h] Hugh L. Montgomery. Least character non-residues and argL(12 + it, x). Lecture Notes in Mathematics, 227:117–129, 1971. CODEN LNMAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060864/.

Montgomery:1971:LAP

[Mon71i] Hugh L. Montgomery. A lemma in additive prime number theory. Lecture Notes in Mathematics, 227:141–144, 1971. CODEN LNMAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060867/.

Montgomery:1971:LBR

[Mon71j] Hugh L. Montgomery. A lower bound of Roth. Lecture Notes in Mathematics, 227:45–49, 1971. CODEN LNMAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060856/.

Montgomery:1971:MMF

[Mon71k] Hugh L. Montgomery. Mean moduli of L-functions. Lecture Notes in Mathematics, 227:74–84, 1971. CODEN LNMAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060861/.

Montgomery:1971:MVT

[Mon71l] Hugh L. Montgomery. The mean value theorem of Barban. Lecture Notes in Mathematics, 227:145–154, 1971. CODEN LN-MAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060868/.

Montgomery:1971:NMV

[Mon71m] Hugh L. Montgomery. New mean value theorems. Lecture Notes in Mathematics, 227:58–63, 1971. CODEN LNMAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060858/.

Montgomery:1971:PNT

[Mon71n] Hugh L. Montgomery. The prime number theorems of Hoheisel and Selberg. Lecture Notes in Mathematics, 227:130–132, 1971. CODEN LNMAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060865/.

Montgomery:1971:TBP

[Mon71o] Hugh L. Montgomery. Three basic principles. Lecture Notes in Mathematics, 227:1–10, 1971. CODEN LNMAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060852/.

Montgomery:1971:TMN

[Mon71p] Hugh L. Montgomery. Topics in Multiplicative Number Theory, volume 227 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 178 pp. LCCN QA3 .L28 no. 227. URL http:

//link.springer.com/book/10.1007/BFb0060851;   http:/
/www.springerlink.com/content/978-3-540-36935-6.

Montgomery:1971:WSA

[Mon71q] Hugh L. Montgomery. A weighted sieve and its application. Lecture Notes in Mathematics, 227:31–44, 1971. CODEN LN-MAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060855/.

Montgomery:1971:ZFR

[Mon71r] Hugh L. Montgomery. Zero-free regions and the proliferation of zeros. Lecture Notes in Mathematics, 227:85–94, 1971. CODEN LNMAA2. ISBN 3-540-05641-6 (print), 3-540-36935-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060862/.

Moore:1970:SPL

[Moo70] John C. Moore. Some properties of the loop homology of commutative coalgebras. Lecture Notes in Mathematics, 168:232– 245, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058525/.

Moore:1971:BSS

[Moo71] John C. Moore. Bockstein spectral sequences, modified Bockstein spectral sequences, and Hopf algebras over certain Frobenius rings. Lecture Notes in Mathematics, 249:84–93, 1971. CO-DEN LNMAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060898/.

Moore:1974:CNI

[Moo74] Hays Moore. Comparison of numerical integration techniques for orbital applications. Lecture Notes in Mathematics, 362:149– 166, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066590/.

Morel:1970:EEH

[Mor70] D. Morel. Une extension des espaces de Hardy Liée À l’approximation. (French) []. Lecture Notes in Mathematics, 164: 54–61, 1970. CODEN LNMAA2. ISBN 3-540-05191-0 (print), 3-540-36376-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069604/.

Morlet:1971:HTV

[Mor71a] Claude Morlet. Hauptvermutung et triangulation des variétés (d’après Kirby, Siebenmann et aussi Lées, Wall, etc …). (French) []. Lecture Notes in Mathematics, 179:261–278, 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058817/.

Morris:1971:CAN

[Mor71b] John Ll. Morris, editor. Conference on Applications of Numerical Analysis: Held in Dundee/Scotland, March 23–26, 1971, volume 228 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 228. URL http:/

/link.springer.com/book/10.1007/BFb0069442;   http://
www.springerlink.com/content/978-3-540-36976-9.

Morris:1971:STN

[Mor71c] John Ll. Morris, editor. Symposium on the Theory of Numerical Analysis: Held in Dundee/Scotland, September 15–23, 1970, volume 193 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05422-7 (print), 3-540-36538-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 193; QA3 .L35 v.193. URL http:

//link.springer.com/book/10.1007/BFb0060340;   http:/
/www.springerlink.com/content/978-3-540-36538-9.

Morrow:1971:NSD

[Mor71d] J. Morrow. Non-singular deformations of singular spaces with normal crossings. Lecture Notes in Mathematics, 209:234–243, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068908/.

Morozan:1972:BPS

[Mor72] Toader Morozan. Boundedness properties for stochastic systems. Lecture Notes in Mathematics, 294:21–34, 1972. CODEN LN-MAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064932/.

Morimoto:1973:EWT

[Mor73] Mitsuo Morimoto. Edge of the wedge theorem and hyperfunction. Lecture Notes in Mathematics, 287:41–81, 1973. CODEN LNMAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068145/.

Morton:1974:DDS

[Mor74] K. W. Morton. The design of difference schemes for studying physical instabilities. Lecture Notes in Mathematics, 363:105– 117, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069129/.

Moskowitz:1970:ANC

[Mos70] Samuel E. Moskowitz. Approximate non-coplanar orbital transfers with minimum expenditure of fuel. Lecture Notes in Mathematics, 132:281–291, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066687/.

Moss:1972:SBR

[Mos72] J. M. B. Moss. Some B. Russell’s sprouts (1903–1908). Lecture Notes in Mathematics, 255:211–250, 1972. CODEN LN-MAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059547/.

Mott:1973:GDA

[Mot73] Joe L. Mott. The group of divisibility and its applications. Lecture Notes in Mathematics, 311:194–208, 1973. CODEN LN-MAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068929/.

Moulis:1971:VDI

[Mou71] Nicole Moulis. Variétés de dimension infinie. (French) []. Lecture Notes in Mathematics, 180:253–267, 1971. CODEN LN-MAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058835/.

Moulis:1972:CDG

[Mou72a] Nicole Moulis. Contractibilité du groupe linéaire d’un espace de Hilbert. (French) []. Lecture Notes in Mathematics, 259:19– 27, 1972. CODEN LNMAA2. ISBN 3-540-05789-7 (print), 3-540-37186-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058450/.

Moulis:1972:DPFb

[Mou72b] Nicole Moulis. Definition et propriétés fondamentales des structures de Fredholm et des structures etalées sur une variété Hilbertienne. (French) []. Lecture Notes in Mathematics, 259:28–40, 1972. CODEN LNMAA2. ISBN 3-540-05789-7 (print), 3-540-37186-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058451/.

Moulis:1972:DPFa

[Mou72c] Nicole Moulis. Definition et propriétés fondamentales des structures de Fredholm et etalées sur un fibre vectoriel. (French) []. Lecture Notes in Mathematics, 259:1–18, 1972. CODEN LN-MAA2. ISBN 3-540-05789-7 (print), 3-540-37186-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058449/.

Moulis:1972:DEF

[Mou72d] Nicole Moulis. Diffeomorphismes étalés fondamentaux. (French) []. Lecture Notes in Mathematics, 259:41–52, 1972. CODEN LNMAA2. ISBN 3-540-05789-7 (print), 3-540-37186-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058452/.

Moulis:1972:EVT

[Mou72e] Nicole Moulis. Existence de voisinages tubulaires. (French) []. Lecture Notes in Mathematics, 259:53–60, 1972. CODEN LN-MAA2. ISBN 3-540-05789-7 (print), 3-540-37186-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058453/.

Moulis:1972:PVE

[Mou72f] Nicole Moulis. Plongement de variétés étalées (existence, isotopie). (French) []. Lecture Notes in Mathematics, 259:73–92, 1972. CODEN LNMAA2. ISBN 3-540-05789-7 (print), 3-540-37186-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058455/.

Moulis:1972:SFV

[Mou72g] Nicole Moulis. Structures de Fredholm sur les Variétés Hilbertiennes, volume 259 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05789-7 (print), 3-540-37186-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 123 pp. LCCN QA3 .L28 no. 259. URL http:/

/link.springer.com/book/10.1007/BFb0058448;   http://
www.springerlink.com/content/978-3-540-37186-1.

Moulis:1972:TMC

[Mou72h] Nicole Moulis. Théorème de Mazur et classification des structures de Fredholm sur une variété Hilbertienne. (French) []. Lecture Notes in Mathematics, 259:114–120, 1972. CODEN LN-MAA2. ISBN 3-540-05789-7 (print), 3-540-37186-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058457/.

Moulis:1972:TSF

[Mou72i] Nicole Moulis. Théorème de stabilité. (French) []. Lecture Notes in Mathematics, 259:93–113, 1972. CODEN LNMAA2. ISBN 3-540-05789-7 (print), 3-540-37186-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058456/.

Moulis:1972:TIA

[Mou72j] Nicole Moulis. Théorème d’isotopie ambiante étalée des voisinages tubulaires. (French) []. Lecture Notes in Mathematics, 259: 61–72, 1972. CODEN LNMAA2. ISBN 3-540-05789-7 (print), 3-540-37186-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058454/.

Moussu:1974:CEF

[Mou74] Robert Moussu. Sur les classes exotiques des feuilletages. (French) []. Lecture Notes in Mathematics, 392:37–42, 1974. CO-DEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064108/.

Mozzochi:1971:MT

[Moz71a] Charles J. Mozzochi. The main theorem. Lecture Notes in Mathematics, 199:8–10, 1971. CODEN LNMAA2. ISBN 3-540-05475-8 (print), 3-540-36656-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061169/.

Mozzochi:1971:PCF

[Moz71b] Charles J. Mozzochi. On the Pointwise Convergence of Fourier Series, volume 199 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05475-8 (print), 3-540-36656-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 87 pp. LCCN QA3 .L28 no. 199. URL http:/

/link.springer.com/book/10.1007/BFb0061167;   http://
www.springerlink.com/content/978-3-540-36656-0.

Mozzochi:1971:PTa

[Moz71c] Charles J. Mozzochi. A proof of theorem (2.2). Lecture Notes in Mathematics, 199:11–18, 1971. CODEN LNMAA2. ISBN 3-540-05475-8 (print), 3-540-36656-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061170/.

Mozzochi:1971:PTb

[Moz71d] Charles J. Mozzochi. A proof of theorem (3.6). Lecture Notes in Mathematics, 199:19, 1971. CODEN LNMAA2. ISBN 3-540-05475-8 (print), 3-540-36656-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061171.

Mozzochi:1971:PTc

[Moz71e] Charles J. Mozzochi. A proof of theorem (4.2). Lecture Notes in Mathematics, 199:20–23, 1971. CODEN LNMAA2. ISBN 3-540-05475-8 (print), 3-540-36656-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061172/.

Mozzochi:1971:PTd

[Moz71f] Charles J. Mozzochi. A proof of theorem (5.2). Lecture Notes in Mathematics, 199:24–43, 1971. CODEN LNMAA2. ISBN 3-540-05475-8 (print), 3-540-36656-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061173/.

Mozzochi:1971:TSW

[Moz71g] Charles J. Mozzochi. A theorem of Stein and Weiss. Lecture Notes in Mathematics, 199:1–7, 1971. CODEN LNMAA2. ISBN 3-540-05475-8 (print), 3-540-36656-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061168/.

Moore:1971:NFA

[MP71] John C. Moore and Franklin P. Peterson. Nearly Frobenius algebras and their module categories. Lecture Notes in Mathematics, 249:94–98, 1971. CODEN LNMAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060899/.

Maruyama:1973:PSJ

[MP73a] G. Maruyama and Yu. V. Prokhorov, editors. Proceedings of the Second Japan–USSR Symposium on Probability Theory [Kyoto, 1972], volume 330 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 0-387-06358-7, 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 330; QA3 .L28 no. 330 QA273.A1. URL http:/

/link.springer.com/book/10.1007/BFb0061476;   http://
www.springerlink.com/content/978-3-540-46956-8.

Murthy:1973:PR

[MP73b] M. Pavaman Murthy and Claudio Pedrini. K_{0} and K_{1} of polynomial rings. Lecture Notes in Mathematics, 342:109–121, 1973. CODEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073722/.

Markoe:1971:FSP

[MR71] A. Markoe and H. Rossi. Families of strongly pseudoconvex manifolds. Lecture Notes in Mathematics, 184:182–207, 1971. CODEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059285/.

Meyer:1972:SRT

[MR72] P.-A. Meyer and H. Rost. Le schéma de remplissage en temps continu. (French) []. Lecture Notes in Mathematics, 258:130– 150, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059470/.

Mathias:1973:CSS

[MR73] A. R. D. Mathias and H. Rogers, editors. Cambridge Summer School in Mathematical Logic: Held in Cambridge/England, August 1–21, 1971, volume 337 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 337. URL http:/

/link.springer.com/book/10.1007/BFb0066770;   http://
www.springerlink.com/content/978-3-540-36884-7.

Mrowka:1970:SCS

[Mró70] S. Mrówka. Some comments on the space of subsets. Lecture Notes in Mathematics, 171:59–63, 1970. CODEN LN-MAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069722/.

Muller:1970:BM

[MS70a] Gert H. Müller and Dirk Siefkes. Back matter. Lecture Notes in Mathematics, 120:125–133, 1970. CODEN LNMAA2. ISBN 3-540-04909-6 (print), 3-540-36252-5 (e-book), 3-662-35848-4 (print), 3-662-36678-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-662-36678-3/1.pdf.

Muller:1970:BDP

[MS70b] Gert H. Müller and Dirk Siefkes. Benefits of the decision procedure. Lecture Notes in Mathematics, 120:86–124, 1970. CO-DEN LNMAA2. ISBN 3-662-35848-4 (print), 3-662-36678-9 (e-book), 3-540-04909-6 (print), 3-540-36252-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-36678-3_2/;
http://link.springer.com/chapter/10.1007/BFb0061049/.

Muller:1970:DCS

[MS70c] Gert H. Müller and Dirk Siefkes. Decidability and completeness of SC. Lecture Notes in Mathematics, 120:1–85, 1970. CODEN LNMAA2. ISBN 3-540-04909-6 (print), 3-540-36252-5 (e-book), 3-662-35848-4 (print), 3-662-36678-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-36678-3_1/;
http://link.springer.com/chapter/10.1007/BFb0061048/.

Muller:1970:DTB

[MS70d] Gert H. Müller and Dirk Siefkes. Decidable Theories. Büchi’s Monadic Second Order Successor Arithmetic, volume 120 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04909-6 (print), 3-540-36252-5 (e-book), 3-662-35848-4 (print), 3-662-36678-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). xii + 133 pp. LCCN QA3 .L28 no. 120. URL http://link.

springer.com/book/10.1007/978-3-662-36678-3;   http:/
/link.springer.com/book/10.1007/BFb0061047;   http://
www.springerlink.com/content/978-3-540-36252-4;   http:
//www.springerlink.com/content/978-3-662-36678-3.

Muller:1970:FM

[MS70e] Gert H. Müller and Dirk Siefkes. Front matter. Lecture Notes in Mathematics, 120:n2–xii, 1970. CODEN LNMAA2. ISBN 3-540-04909-6 (print), 3-540-36252-5 (e-book), 3-662-35848-4 (print), 3-662-36678-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-662-36678-3/1.pdf.

MacRae:1973:SIE

[MS73a] R. E. MacRae and Pierre Samuel. Subfields of index 2 of elliptic function fields. Lecture Notes in Mathematics, 311:171–193, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068928/.

Mathai:1973:AFS

[MS73b] A. M. Mathai and R. K. Saxena. Application of G-function in statistics. Lecture Notes in Mathematics, 348:189–230, 1973. CO-DEN LNMAA2. ISBN 3-540-06482-6 (print), 3-540-37800-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060474/.

Mathai:1973:CRF

[MS73c] A. M. Mathai and R. K. Saxena. Computable representations of a G-function in the logarithmic case. Lecture Notes in Mathematics, 348:157–188, 1973. CODEN LNMAA2. ISBN 3-540-06482-6 (print), 3-540-37800-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060473/.

Mathai:1973:FIS

[MS73d] A. M. Mathai and R. K. Saxena. Finite and infinite series of G-functions. Lecture Notes in Mathematics, 348:108–156, 1973. CODEN LNMAA2. ISBN 3-540-06482-6 (print), 3-540-37800-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060472/.

Mathai:1973:GHF

[MS73e] A. M. Mathai and R. K. Saxena. Generalized Hypergeometric Functions with Applications in Statistics and Physical Sciences, volume 348 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06482-6 (print), 3-540-37800-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 314 pp. LCCN QA3 .L28 no. 348. URL http:/

/link.springer.com/book/10.1007/BFb0060468;   http://
www.springerlink.com/content/978-3-540-37800-6.

Mathai:1973:IF

[MS73f] A. M. Mathai and R. K. Saxena. Integrals of G-functions. Lecture Notes in Mathematics, 348:69–107, 1973. CODEN LN-MAA2. ISBN 3-540-06482-6 (print), 3-540-37800-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060471/.

Mathai:1973:MF

[MS73g] A. M. Mathai and R. K. Saxena. Meijer’s G-function. Lecture Notes in Mathematics, 348:1–40, 1973. CODEN LNMAA2. ISBN 3-540-06482-6 (print), 3-540-37800-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060469/.

Mathai:1973:OAF

[MS73h] A. M. Mathai and R. K. Saxena. Other applications of the G-function. Lecture Notes in Mathematics, 348:231–260, 1973. CODEN LNMAA2. ISBN 3-540-06482-6 (print), 3-540-37800-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060475/.

Mathai:1973:PCM

[MS73i] A. M. Mathai and R. K. Saxena. Particular cases of Meijer’s Gfunction. Lecture Notes in Mathematics, 348:41–68, 1973. CO-DEN LNMAA2. ISBN 3-540-06482-6 (print), 3-540-37800-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060470/.

Muller:1973:DTI

[MS73j] G. H. Müller and D. Siefkes. Decidable Theories II: The Monadic Second Order Theory of All Countable Ordinals, volume 328 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06345-5 (print), 3-540-46946-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 217 pp. LCCN QA3 .L28 no. 328. URL http:

//link.springer.com/book/10.1007/BFb0082720;   http:/
/www.springerlink.com/content/978-3-540-46946-9.

Meyer:1973:RJH

[MT73] P.-A. Meyer and M. Traki. Réduites et jeux de hasard. (French) []. Lecture Notes in Mathematics, 321:155–171, 1973. CODEN LNMAA2. ISBN 3-540-06287-4 (print), 3-540-40023-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0071404/.

Murakami:1972:AAG

[Mur72a] Shingo Murakami. Affine automorphism groups of Siegel domains. Lecture Notes in Mathematics, 286:9–23, 1972. CODEN LNMAA2. ISBN 3-540-05985-7 (print), 3-540-37982-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058570/.

Murakami:1972:AHS

[Mur72b] Shingo Murakami. Affinely homogeneous Siegel domains. Lecture Notes in Mathematics, 286:23–29, 1972. CODEN LN-MAA2. ISBN 3-540-05985-7 (print), 3-540-37982-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058571/.

Murakami:1972:E

[Mur72c] Shingo Murakami. Examples. Lecture Notes in Mathematics, 286: 76–95, 1972. CODEN LNMAA2. ISBN 3-540-05985-7 (print), 3-540-37982-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058577/.

Murakami:1972:GSD

[Mur72d] Shingo Murakami. Generalised Siegel domains. Lecture Notes in Mathematics, 286:29–34, 1972. CODEN LNMAA2. ISBN 3-540-05985-7 (print), 3-540-37982-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058572/.

Murakami:1972:I

[Mur72e] Shingo Murakami. Introduction. Lecture Notes in Mathematics, 286:1–3, 1972. CODEN LNMAA2. ISBN 3-540-05985-7 (print), 3-540-37982-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058568/.

Murakami:1972:LAH

[Mur72f] Shingo Murakami. Lie algebra (D) for homogeneous Siegel domains. Lecture Notes in Mathematics, 286:62–75, 1972. CODEN LNMAA2. ISBN 3-540-05985-7 (print), 3-540-37982-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058576/.

Murakami:1972:LAS

[Mur72g] Shingo Murakami. Lie algebras (D) for Siegel domains. Lecture Notes in Mathematics, 286:52–62, 1972. CODEN LN-MAA2. ISBN 3-540-05985-7 (print), 3-540-37982-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058575/.

Murakami:1972:ASD

[Mur72h] Shingo Murakami. On Automorphisms of Siegel Domains, volume 286 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05985-7 (print), 3-540-37982-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 95 pp. LCCN QA3 .L28 no. 286. URL http:/

/link.springer.com/book/10.1007/BFb0058567;   http://
www.springerlink.com/content/978-3-540-37982-9.

Murakami:1972:SD

[Mur72i] Shingo Murakami. Siegel domains. Lecture Notes in Mathematics, 286:3–9, 1972. CODEN LNMAA2. ISBN 3-540-05985-7 (print), 3-540-37982-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058569/.

Murakami:1972:S

[Mur72j] Shingo Murakami. Structure of (D) for D with c = 1/2. Lecture Notes in Mathematics, 286:38–52, 1972. CODEN LN-MAA2. ISBN 3-540-05985-7 (print), 3-540-37982-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058574/.

Murakami:1972:VFB

[Mur72k] Shingo Murakami. Vector fields belonging to (D). Lecture Notes in Mathematics, 286:34–38, 1972. CODEN LNMAA2. ISBN 3-540-05985-7 (print), 3-540-37982-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058573/.

Murre:1974:SRC

[Mur74] J. P. Murre. Some results on cubic threefolds. Lecture Notes in Mathematics, 412:140–160, 1974. CODEN LNMAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066158/.

Mordeson:1970:G

[MV70a] John N. Mordeson and Bernard Vinograde. Generators. Lecture Notes in Mathematics, 173:1–73, 1970. CODEN LNMAA2. ISBN 3-540-05295-X (print), 3-540-36431-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060993/.

Mordeson:1970:IF

[MV70b] John N. Mordeson and Bernard Vinograde. Intermediate fields. Lecture Notes in Mathematics, 173:74–92, 1970. CODEN LN-MAA2. ISBN 3-540-05295-X (print), 3-540-36431-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060994/.

Mordeson:1970:SA

[MV70c] John N. Mordeson and Bernard Vinograde. Some applications. Lecture Notes in Mathematics, 173:93–134, 1970. CODEN LN-MAA2. ISBN 3-540-05295-X (print), 3-540-36431-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060995/.

Mordeson:1970:SAP

[MV70d] John N. Mordeson and Bernard Vinograde. Structure of Arbitrary Purely Inseparable Extension Fields, volume 173 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CO-DEN LNMAA2. ISBN 3-540-05295-X (print), 3-540-36431-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 138 pp. LCCN QA3 .L28 no. 173. URL http:/

/link.springer.com/book/10.1007/BFb0060992;   http://
www.springerlink.com/content/978-3-540-36431-3.

Manna:1971:TAP

[MW71] Zohar Manna and Richard J. Waldinger. Towards automatic program synthesis. Lecture Notes in Mathematics, 188:270– 310, 1971. CODEN LNMAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059702/.

Meyer:1972:RRR

[MW72] P.-A. Meyer and J. B. Walsh. Un résultat sur les résolvantes de Ray. (French) []. Lecture Notes in Mathematics, 258:168– 172, 1972. CODEN LNMAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059474/.

Montgomery:1972:DPF

[MY72] Deane Montgomery and C. T. Yang. Differentiable pseudofree circle actions on homotopy seven spheres. Lecture Notes in Mathematics, 298:41–101, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070033/.

Myhill:1973:SPI

[Myh73] John Myhill. Some properties of intuitionistic Zermelo–Frankel set theory. Lecture Notes in Mathematics, 337:206–231, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066775/.

Nachbin:1971:UHT

[Nac71] Leopoldo Nachbin. Uniformité d’holomorphie et type exponentiel. (French) []. Lecture Notes in Mathematics, 205:216– 224, 1971. CODEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059841/.

Nachbin:1974:FAA

[Nac74a] Leopoldo Nachbin, editor. Functional Analysis and Applications: Proceedings of the Symposium of Analysis Universidade Federal de Pernambuco Recife, Pernambuco, Brasil, July 9–29, 1972, volume 384 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 384. URL http:/

/link.springer.com/book/10.1007/BFb0057538;   http://
www.springerlink.com/content/978-3-540-38342-0.

Nachbin:1974:GID

[Nac74b] Leopoldo Nachbin. A glimpse at infinite dimensional holomorphy. Lecture Notes in Mathematics, 364:69–79, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069005/.

Nadler:1970:SRM

[Nad70] Sam B. Nadler, Jr. Some results on multi-valued contraction mappings. Lecture Notes in Mathematics, 171:64–69, 1970. CO-DEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069723/.

Nagasawa:1972:BPM

[Nag72] Masao Nagasawa. Branching property of Markov processes. Lecture Notes in Mathematics, 258:177–197, 1972. CODEN LN-MAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059476/.

Nakao:1973:CTS

[Nak73] Shintaro Nakao. Comparison theorems for solutions of onedimensional stochastic differential equations. Lecture Notes in Mathematics, 330:310–315, 1973. CODEN LNMAA2. ISBN 3-540-06358-7 (print), 3-540-46956-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061496/.

Namikawa:1974:SD

[Nam74] Yukihiko Namikawa. Studies on degeneration. Lecture Notes in Mathematics, 412:165–210, 1974. CODEN LNMAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066159/.

Narasimhan:1970:CBC

[Nar70] Raghavan Narasimhan. Cohomology with bounds on complex spaces. Lecture Notes in Mathematics, 155:141–150, 1970. CO-DEN LNMAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060322/.

Narens:1974:FEG

[Nar74a] Louis Narens. Field embeddings of generalized metric spaces. Lecture Notes in Mathematics, 369:155–170, 1974. CODEN LN-MAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066010/.

Narens:1974:HTG

[Nar74b] Louis Narens. Homeomorphism types of generalized metric spaces. Lecture Notes in Mathematics, 369:171–179, 1974. CO-DEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066011/.

Nashed:1974:ARS

[Nas74] M. Zuhair Nashed. Approximate regularized solutions to improperly posed linear integral and operator equations. Lecture Notes in Mathematics, 430:289–332, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066275/.

Neri:1971:C

[Ner71a] Umberto Neri. Convolutions. Lecture Notes in Mathematics, 200: 6–18, 1971. CODEN LNMAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079052/.

Neri:1971:DFT

[Ner71b] Umberto Neri. Distributions and Fourier transforms. Lecture Notes in Mathematics, 200:124–150, 1971. CODEN LN-MAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079058/.

Neri:1971:FT

[Ner71c] Umberto Neri. Fourier transforms. Lecture Notes in Mathematics, 200:19–54, 1971. CODEN LNMAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079053/.

Neri:1971:HT

[Ner71d] Umberto Neri. The Hilbert transform. Lecture Notes in Mathematics, 200:55–81, 1971. CODEN LNMAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079054/.

Neri:1971:I

[Ner71e] Umberto Neri. Introduction. Lecture Notes in Mathematics, 200: 1, 1971. CODEN LNMAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079050.

Neri:1971:N

[Ner71f] Umberto Neri. Notation. Lecture Notes in Mathematics, 200: 122–123, 1971. CODEN LNMAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079057.

Neri:1971:Pb

[Ner71g] Umberto Neri. Preface. Lecture Notes in Mathematics, 200: 121, 1971. CODEN LNMAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079056.

Neri:1971:Pa

[Ner71h] Umberto Neri. Preliminaries. Lecture Notes in Mathematics, 200: 2–5, 1971. CODEN LNMAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0079051/.

Neri:1971:SIO

[Ner71i] Umberto Neri. Singular integral operators. Lecture Notes in Mathematics, 200:238–272, 1971. CODEN LNMAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0079061/.

Neri:1971:SIb

[Ner71j] Umberto Neri. Singular Integrals, volume 200 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CO-DEN LNMAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 272 pp. LCCN QA3 .L28 no. 200. URL http:/

/link.springer.com/book/10.1007/BFb0079049;   http://
www.springerlink.com/content/978-3-540-36864-9.

Neri:1971:SIS

[Ner71k] Umberto Neri. Singular integrals and Sobolev spaces. Lecture Notes in Mathematics, 200:151–211, 1971. CODEN LN-MAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079059/.

Neri:1971:SIa

[Ner71l] Umberto Neri. Singular integrals in E^{n}. Lecture Notes in Mathematics, 200:82–117, 1971. CODEN LNMAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079055/.

Neri:1971:SH

[Ner71m] Umberto Neri. Spherical harmonics. Lecture Notes in Mathematics, 200:212–237, 1971. CODEN LNMAA2. ISBN 3-540-05502-9 (print), 3-540-36864-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0079060/.

Neto:1974:FSH

[Net74] J. Barros Neto. Fundamental solutions of hypoelliptic boundary value problems. Lecture Notes in Mathematics, 384:139– 146, 1974. CODEN LNMAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057545/.

Neustadt:1970:OCP

[Neu70] Lucien W. Neustadt. Optimal control problems with operator equation restrictions. Lecture Notes in Mathematics, 132:292– 306, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066688/.

Neumann:1974:I

[Neu74] B. H. Neumann. Introduction. Lecture Notes in Mathematics, 372:1–7, 1974. CODEN LNMAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21571-5_1/;
http://link.springer.com/chapter/10.1007/BFb0065155/;
 http://link.springer.com/content/pdf/bfm:978-3-662-
21571-5/1.pdf.

Newberger:1974:MTH

[New74a] Edward Newberger. A majorization technique for hyperbolic equations. Lecture Notes in Mathematics, 430:333–362, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066276/.

Newman:1974:PSI

[New74b] M. F. Newman, editor. Proceedings of the Second International Conference on the Theory of Groups: Australian National University, August 13–24, 1973, volume 372 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LN-MAA2. ISBN 3-540-06845-7 (print), 3-540-37801-4 (e-book), 3-662-21571-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 372. URL http://link.

springer.com/book/10.1007/978-3-662-21571-5;   http:/
/link.springer.com/book/10.1007/BFb0065154;   http://
www.springerlink.com/content/978-3-540-37801-3;   http:
//www.springerlink.com/content/978-3-662-21571-5.

Naveira:1974:ZOC

[NF74] A. M. Naveira and C. Fuertes. Les zéros des opérateurs courbure holomorphes non-negatifs. (French) []. Lecture Notes in Mathematics, 392:61–63, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064112/.

Nicolis:1973:MPT

[Nic73] Grégoire Nicolis. Mathematical problems in theoretical biology. Lecture Notes in Mathematics, 322:210–230, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060568/.

Nirenberg:1971:PMP

[Nir71] L. Nirenberg. A proof of the Malgrange preparation theorem. Lecture Notes in Mathematics, 192:97–105, 1971. CODEN LN-MAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066814/.

Noble:1971:SAN

[Nob71] Ben Noble. Some applications of the numerical solution of integral equations to boundary value problems. Lecture Notes in Mathematics, 228:137–154, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069453/.

Nohel:1971:PVE

[Noh71] John A. Nohel. Perturbations of Volterra equations and admissibility. Lecture Notes in Mathematics, 243:40–53, 1971. CODEN LNMAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058717/.

Norguet:1971:ITC

[Nor71] François Norguet. Introduction à la théorie cohomologique des résidus. (French) []. Lecture Notes in Mathematics, 205:34– 55, 1971. CODEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059832/.

Norguet:1974:FPV

[Nor74a] François Norguet, editor. Fonctions de Plusieurs Variables Complexes: Séminaire François Norguet Octobre 1970–Décembre 1973, volume 409 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 409. URL http:/

/link.springer.com/book/10.1007/BFb0068093;   http://
www.springerlink.com/content/978-3-540-37814-3.

Norguet:1974:FHD

[Nor74b] François Norguet. Fonctions holomorphes definies par des intégrales. (French) []. Lecture Notes in Mathematics, 409:161, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068097.

Norguet:1974:IFP

[Nor74c] François Norguet. Introduction aux fonctions de plusieurs variables complexes représentations intégrales. (French) []. Lecture Notes in Mathematics, 409:1–97, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068094/.

Norguet:1974:RCV

[Nor74d] François Norguet. Remarques sur la cohomologie des variétés analytiques complexes. (French) []. Lecture Notes in Mathematics, 409:162–175, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068098/.

Noverraz:1970:FAT

[Nov70] Philippe Noverraz. Fonctions analytiques et théorèmes de prolongement dans les espaces vectoriels topologiques complexes. (French) []. Lecture Notes in Mathematics, 116:21–29, 1970. CO-DEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059242/.

Noverraz:1972:CFE

[Nov72] Ph. Noverraz. Convexité fonctionnelle dans les espaces de Banach à base. (French) []. Lecture Notes in Mathematics, 275: 166–176, 1972. CODEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058500/.

Noverraz:1973:TCT

[Nov73] Philippe Noverraz. Sur le théorème de Cartan–Thullen–Oka en dimension infinie. (French) []. Lecture Notes in Mathematics, 332: 59–68, 1973. CODEN LNMAA2. ISBN 3-540-06372-2 (print), 3-540-46980-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060907/.

Noverraz:1974:AHP

[Nov74a] Philippe Noverraz. Approximation of holomorphic or plurisubharmonic functions in certain Banach spaces. Lecture Notes in Mathematics, 364:178–185, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069017/.

Noverraz:1974:CPC

[Nov74b] Philippe Noverraz. Completion pseudo-convexe en dimension infinie. (French) []. Lecture Notes in Mathematics, 410:15–17, 1974.

CODEN LNMAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066036/.

Narayan:1971:UAS

[NS71] Jack Narayan and Gilbert Stengle. Uniform asymptotic splitting of linear differential equations. Lecture Notes in Mathematics, 183:170–175, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060417/.

Nakamizo:1972:ASO

[NS72] Takayoshi Nakamizo and Yoshikazu Sawaragi. Analytical study on n-th order linear system with stochastic coefficients. Lecture Notes in Mathematics, 294:173–185, 1972. CODEN LN-MAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064942/.

Neunzert:1972:TNE

[NW72] H. Neunzert and J. Wick. Theoretische und numerische Ergebnisse zur nichtlinearen Vlasov-Gleichung. (German) []. Lecture Notes in Mathematics, 267:159–185, 1972. CODEN LN-MAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061619/.

Neunzert:1974:ALI

[NW74] H. Neunzert and J. Wick. Die Approximation der Lösung von Integro-Differentialgleichungen durch endliche Punktmengen. (German) []. Lecture Notes in Mathematics, 395:275–290, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060678/.

Osteyee:1974:CBG

[OG74a] David Bridston Osteyee and Irving John Good. Comparison between I(R_{T}: S_{T}), W_{T} (H_{N}/H_{S+N}), and J_{T} (N, S + N) for Gaussian signals and noise. Lecture Notes in Mathematics, 376:96– 97, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064139.

Osteyee:1974:CAA

[OG74b] David Bridston Osteyee and Irving John Good. Conclusions and areas for additional research. Lecture Notes in Mathematics, 376: 130, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064144.

Osteyee:1974:D

[OG74c] David Bridston Osteyee and Irving John Good. Divergence. Lecture Notes in Mathematics, 376:44–46, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064134/.

Osteyee:1974:E

[OG74d] David Bridston Osteyee and Irving John Good. Entropy. Lecture Notes in Mathematics, 376:13–17, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064130/.

Osteyee:1974:EMIa

[OG74e] David Bridston Osteyee and Irving John Good. Expected mutual information. Lecture Notes in Mathematics, 376:26–38, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064132/.

Osteyee:1974:EMIb

[OG74f] David Bridston Osteyee and Irving John Good. Expected mutual information, expected weight of evidence, and divergence for random processes. Lecture Notes in Mathematics, 376:47–53, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064135/.

Osteyee:1974:EMIc

[OG74g] David Bridston Osteyee and Irving John Good. Expected mutual information rate. Lecture Notes in Mathematics, 376:98– 111, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064140/.

Osteyee:1974:EWE

[OG74h] David Bridston Osteyee and Irving John Good. Expected weight of evidence. Lecture Notes in Mathematics, 376:39–43, 1974. CO-DEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064133/.

Osteyee:1974:EEW

[OG74i] David Bridston Osteyee and Irving John Good. Expressing expected weight of evidence for Gaussian generalized processes in terms of integral operators. Lecture Notes in Mathematics, 376: 92–95, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064138/.

Osteyee:1974:GPE

[OG74j] David Bridston Osteyee and Irving John Good. Gaussian processes with equal covariance functions including nonrandom signals in Gaussian noise. Lecture Notes in Mathematics, 376:118– 122, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064142/.

Osteyee:1974:IEW

[OG74k] David Bridston Osteyee and Irving John Good. Information in events and weight of evidence. Lecture Notes in Mathematics, 376:8–12, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064129/.

Osteyee:1974:IWE

[OG74l] David Bridston Osteyee and Irving John Good. Information, Weight of Evidence, the Singularity between Probability Measures and Signal Detection, volume 376 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 156 pp. LCCN QA3 .L28 no. 376.

URL http://link.springer.com/book/10.1007/BFb0064126;
 http://www.springerlink.com/content/978-3-540-38294-
2.

Osteyee:1974:I

[OG74m] David Bridston Osteyee and Irving John Good. Introduction. Lecture Notes in Mathematics, 376:1–2, 1974. CODEN LN-MAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064127.

Osteyee:1974:LS

[OG74n] David Bridston Osteyee and Irving John Good. List of symbols. Lecture Notes in Mathematics, 376:3–7, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064128/.

Osteyee:1974:OEE

[OG74o] David Bridston Osteyee and Irving John Good. Other expressions for expected mutual information. Lecture Notes in Mathematics, 376:81–91, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064137/.

Osteyee:1974:REW

[OG74p] David Bridston Osteyee and Irving John Good. Rate of expected weight of evidence. Lecture Notes in Mathematics, 376: 112–117, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064141/.

Osteyee:1974:RBC

[OG74q] David Bridston Osteyee and Irving John Good. Relationships between certain random processes and the singularity between probability measures. Lecture Notes in Mathematics, 376:54–80, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064136/.

Osteyee:1974:SBT

[OG74r] David Bridston Osteyee and Irving John Good. Singularity between two probability measures. Lecture Notes in Mathematics, 376:18–25, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064131/.

Osteyee:1974:SMR

[OG74s] David Bridston Osteyee and Irving John Good. Summary of the major results in this survey for Gaussian processes, including Gaussian signals and noise. Lecture Notes in Mathematics, 376: 123–129, 1974. CODEN LNMAA2. ISBN 3-540-06726-4 (print), 3-540-38294-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064143/.

Ogg:1973:SMF

[Ogg73] Andrew Ogg. Survey of modular functions of one variable. Lecture Notes in Mathematics, 320:1–35, 1973. CODEN LN-MAA2. ISBN 3-540-06219-X (print), 3-540-38509-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-38509-7_
1/;   http://link.springer.com/content/pdf/bfm:978-3-
540-38509-7/1.pdf.

Ohm:1973:HDE

[Ohm73] Jack Ohm. Homological dimension and Euler maps. Lecture Notes in Mathematics, 311:209–221, 1973. CODEN LN-MAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068930/.

Okamoto:1972:HAH

[Oka72] K. Okamoto. Harmonic analysis on homogeneous vector bundles. Lecture Notes in Mathematics, 266:255–271, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059649/.

Okubo:1971:CPS

[Oku71] Kenjiro Okubo. Connection problems for systems of linear differential equations. Lecture Notes in Mathematics, 243:238– 248, 1971. CODEN LNMAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058732/.

Olive:1971:SRQ

[Oli71] D. Olive. Singularities in relativistic quantum mechanics. Lecture Notes in Mathematics, 209:244–269, 1971. CODEN LN-MAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068909/.

OMalley:1974:BLM

[O’M74] R. E. O’Malley, Jr. Boundary layer methods for ordinary differential equations with small coefficients multiplying the highest derivatives. Lecture Notes in Mathematics, 430:363–389, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066277/.

Omori:1974:BT

[Omo74a] Prof. Hideki Omori. Basic theorems I. Lecture Notes in Mathematics, 427:41–52, 1974. CODEN LNMAA2. ISBN 3-540-07013-3 (print), 3-540-37295-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063403/.

Omori:1974:BTI

[Omo74b] Prof. Hideki Omori. Basic theorems II (Frobenius theorem). Lecture Notes in Mathematics, 427:75–82, 1974. CODEN LN-MAA2. ISBN 3-540-07013-3 (print), 3-540-37295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063406/.

Omori:1974:FTS

[Omo74c] Prof. Hideki Omori. Frobenius theorem on strong ILB–Lie groups. Lecture Notes in Mathematics, 427:83–94, 1974. CODEN LNMAA2. ISBN 3-540-07013-3 (print), 3-540-37295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063407/.

Omori:1974:GTS

[Omo74d] Prof. Hideki Omori. General theory of strong ILB–Lie groups and subgroups. Lecture Notes in Mathematics, 427:1–26, 1974. CODEN LNMAA2. ISBN 3-540-07013-3 (print), 3-540-37295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063401/.

Omori:1974:GD

[Omo74e] Prof. Hideki Omori. Groups of diffeomorphisms. Lecture Notes in Mathematics, 427:27–40, 1974. CODEN LNMAA2. ISBN 3-540-07013-3 (print), 3-540-37295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063402/.

Omori:1974:IDL

[Omo74f] Prof. Hideki Omori. Infinite Dimensional Lie Transformations Groups, volume 427 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-07013-3 (print), 3-540-37295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 149 pp. LCCN QA3 .L28 no. 427. URL http:/

/link.springer.com/book/10.1007/BFb0063400;   http://
www.springerlink.com/content/978-3-540-37295-0.

Omori:1974:LAV

[Omo74g] Prof. Hideki Omori. Lie algebras of vector fields. Lecture Notes in Mathematics, 427:123–139, 1974. CODEN LNMAA2. ISBN 3-540-07013-3 (print), 3-540-37295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063410/.

Omori:1974:LGG

[Omo74h] Prof. Hideki Omori. Linear groups and groups of diffeomorphisms. Lecture Notes in Mathematics, 427:140–149, 1974. CO-DEN LNMAA2. ISBN 3-540-07013-3 (print), 3-540-37295-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0063411/.

Omori:1974:ME

[Omo74i] Prof. Hideki Omori. Miscellaneous examples. Lecture Notes in Mathematics, 427:95–113, 1974. CODEN LNMAA2. ISBN 3-540-07013-3 (print), 3-540-37295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0063408/.

Omori:1974:PTG

[Omo74j] Prof. Hideki Omori. Primitive transformation groups. Lecture Notes in Mathematics, 427:114–122, 1974. CODEN LN-MAA2. ISBN 3-540-07013-3 (print), 3-540-37295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063409/.

Omori:1974:RSE

[Omo74k] Prof. Hideki Omori. Review of the smooth extension theorem and a remark on elliptic operators. Lecture Notes in Mathematics, 427:68–74, 1974. CODEN LNMAA2. ISBN 3-540-07013-3 (print), 3-540-37295-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063405/.

Omori:1974:VBS

[Omo74l] Prof. Hideki Omori. Vector bundle over strong ILB–Lie groups. Lecture Notes in Mathematics, 427:53–67, 1974. CODEN LN-MAA2. ISBN 3-540-07013-3 (print), 3-540-37295-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063404/.

Oort:1974:HCN

[Oor74] Frans Oort. Hyperelliptic curves over number fields. Lecture Notes in Mathematics, 412:211–218, 1974. CODEN LN-MAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066160/.

Opfer:1974:BGD

[Opf74] Gerhard Opfer. Eine Bemerkung zur Gewinnung diskretharmonischer Funktionen mit funktionentheoretischen Mitteln. (German) []. Lecture Notes in Mathematics, 395:215–222, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060673/.

Orlik:1972:CAM

[Orl72a] Peter Orlik. Circle actions on 3-manifolds. Lecture Notes in Mathematics, 291:1–21, 1972. CODEN LNMAA2. ISBN 3-540-06014-6 (print), 3-540-37989-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060330/.

Orlik:1972:ECI

[Orl72b] Peter Orlik. Equivariant Cobordism and the α-invariant. Lecture Notes in Mathematics, 291:66–81, 1972. CODEN LN-MAA2. ISBN 3-540-06014-6 (print), 3-540-37989-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060333/.

Orlik:1972:EP

[Orl72c] Peter Orlik. Equivariant plumbing. Lecture Notes in Mathematics, 291:22–31, 1972. CODEN LNMAA2. ISBN 3-540-06014-6 (print), 3-540-37989-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060331/.

Orlik:1972:F

[Orl72d] Peter Orlik. Fibering over S^{1}. Lecture Notes in Mathematics, 291: 115–127, 1972. CODEN LNMAA2. ISBN 3-540-06014-6 (print), 3-540-37989-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060336/.

Orlik:1972:FAF

[Orl72e] Peter Orlik. Free actions of finite groups on S^{3}. Lecture Notes in Mathematics, 291:103–114, 1972. CODEN LNMAA2. ISBN 3-540-06014-6 (print), 3-540-37989-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060335/.

Orlik:1972:FG

[Orl72f] Peter Orlik. Fundamental groups. Lecture Notes in Mathematics, 291:82–102, 1972. CODEN LNMAA2. ISBN 3-540-06014-6 (print), 3-540-37989-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060334/.

Orlik:1972:FT

[Orl72g] Peter Orlik. Further topics. Lecture Notes in Mathematics, 291: 128–150, 1972. CODEN LNMAA2. ISBN 3-540-06014-6 (print), 3-540-37989-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060337/.

Orlik:1972:HWH

[Orl72h] Peter Orlik. On the homology of weighted homogeneous manifolds. Lecture Notes in Mathematics, 298:260–269, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070047/.

Orlik:1972:RS

[Orl72i] Peter Orlik. Resolution of singularities. Lecture Notes in Mathematics, 291:32–65, 1972. CODEN LNMAA2. ISBN 3-540-06014-6 (print), 3-540-37989-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060332/.

Orlik:1972:SM

[Orl72j] Peter Orlik. Seifert Manifolds, volume 291 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06014-6 (print), 3-540-37989-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 155 pp. LCCN QA3 .L28 no. 291. URL http:/

/link.springer.com/book/10.1007/BFb0060329;   http://
www.springerlink.com/content/978-3-540-37989-8.

Ornstein:1970:IBS

[Orn70] D. S. Ornstein. Imbedding Bernoulli shifts in flows. Lecture Notes in Mathematics, 160:178–218, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060654/.

Ornstein:1970:EFI

[OS70] Donald S. Ornstein and Louis Sucheston. On the existence of a σ-finite invariant measure under a generalized Harris condition. Lecture Notes in Mathematics, 160:219–233, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060655/.

Osborne:1971:IEP

[Osb71] M. R. Osborne. On the inverse eigenvalue problem for matrices and related problems for difference and differential equations. Lecture Notes in Mathematics, 228:155–168, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069454/.

Ostrom:1970:FTP

[Ost70a] T. G. Ostrom. Finite Translation Planes, volume 158 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05186-4 (print), 3-540-36369-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 112 pp. LCCN QA3 .L28 no. 158. URL http:/

/link.springer.com/book/10.1007/BFb0070676;   http://
www.springerlink.com/content/978-3-540-36369-9.

Ostrom:1970:LGC

[Ost70b] T. G. Ostrom. Linear groups and collineation groups. Lecture Notes in Mathematics, 158:49–102, 1970. CODEN LN-MAA2. ISBN 3-540-05186-4 (print), 3-540-36369-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070678/.

Ostrom:1970:STP

[Ost70c] T. G. Ostrom. Some translation planes that are not well known. Lecture Notes in Mathematics, 158:4–48, 1970. CODEN LN-MAA2. ISBN 3-540-05186-4 (print), 3-540-36369-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070677/.

Takahashi:1972:DST

[oT72] Moto o Takahashi. Δ̃_{1}-definability in set theory. Lecture Notes in Mathematics, 255:281–304, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059549/.

Ovsyannikov:1970:POD

[OTF70] R. N. Ovsyannikov, V. V. Tokarev, and Yu. M. Fatkin. Parameters optimization of the dynamic system, universal for the given class of maneuvers. Lecture Notes in Mathematics, 112: 254–265, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060209/.

Ouchi:1973:ACP

[Ōuc73] Sunao Ōuchi. On abstract Cauchy problems in the sense of hyperfunction. Lecture Notes in Mathematics, 287:135–152, 1973. CODEN LNMAA2. ISBN 3-540-06218-1 (print), 3-540-38506-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068149/.

Orlik:1972:ERS

[OW72] Peter Orlik and Philip Wagreich. Equivariant resolution of singularities with C∗ action. Lecture Notes in Mathematics, 298: 270–290, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070048/.

Palmquist:1971:DCA

[Pal71] Paul H. Palmquist. The double category of adjoint squares. Lecture Notes in Mathematics, 195:123–153, 1971. CODEN LN-MAA2. ISBN 3-540-05442-1 (print), 3-540-36548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072309/.

Papangelou:1970:AKT

[Pap70] Fredos Papangelou. The ambrose-Kakutani theorem and the Poisson process. Lecture Notes in Mathematics, 160:234–240, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060656/.

Parry:1971:DSN

[Par71a] W. Parry. Dynamical systems on nilmanifolds. Lecture Notes in Mathematics, 206:36–38, 1971. CODEN LNMAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070137/.

Parry:1971:GED

[Par71b] W. Parry. Group extensions of discrete dynamical systems. Lecture Notes in Mathematics, 206:6–8, 1971. CODEN LN-MAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070120/.

Paris:1972:MA

[Par72a] J. B. Paris. On models of arithmetic. Lecture Notes in Mathematics, 255:251–280, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059548/.

Parthasarathy:1972:CS

[Par72b] T. Parthasarathy. Continuous selections. Lecture Notes in Mathematics, 263:3–13, 1972. CODEN LNMAA2. ISBN 3-540-05818-4 (print), 3-540-37464-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058348/.

Parthasarathy:1972:CSS

[Par72c] T. Parthasarathy. Continuous selections and solutions of generalised differential equations. Lecture Notes in Mathematics, 263: 24–34, 1972. CODEN LNMAA2. ISBN 3-540-05818-4 (print), 3-540-37464-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058350/.

Parthasarathy:1972:CSM

[Par72d] T. Parthasarathy. Continuous selections on metric continuum. Lecture Notes in Mathematics, 263:14–23, 1972. CODEN LN-MAA2. ISBN 3-540-05818-4 (print), 3-540-37464-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058349/.

Parthasarathy:1972:GTS

[Par72e] T. Parthasarathy. General theorems on selectors. Lecture Notes in Mathematics, 263:49–57, 1972. CODEN LNMAA2. ISBN 3-540-05818-4 (print), 3-540-37464-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058352/.

Parthasarathy:1972:I

[Par72f] T. Parthasarathy. Introduction. Lecture Notes in Mathematics, 263:1–2, 1972. CODEN LNMAA2. ISBN 3-540-05818-4 (print), 3-540-37464-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058347.

Parthasarathy:1972:MSA

[Par72g] T. Parthasarathy. Measurable selections and an application to stochastic games. Lecture Notes in Mathematics, 263:35– 48, 1972. CODEN LNMAA2. ISBN 3-540-05818-4 (print), 3-540-37464-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058351/.

Parthasarathy:1972:UST

[Par72h] T. Parthasarathy. On the uniformisation of sets in topological spaces. Lecture Notes in Mathematics, 263:76–84, 1972. CODEN LNMAA2. ISBN 3-540-05818-4 (print), 3-540-37464-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058355/.

Parthasarathy:1972:STT

[Par72i] T. Parthasarathy. Selection Theorems and their Applications, volume 263 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05818-4 (print), 3-540-37464-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 101 pp. LCCN QA3 .L28 no. 263. URL http:

//link.springer.com/book/10.1007/BFb0058346;   http:/
/www.springerlink.com/content/978-3-540-37464-0.

Parthasarathy:1972:SRS

[Par72j] T. Parthasarathy. Supplementary remarks on selection theorems. Lecture Notes in Mathematics, 263:85–101, 1972. CODEN LNMAA2. ISBN 3-540-05818-4 (print), 3-540-37464-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058356/.

Parthasarathy:1972:TAM

[Par72k] T. Parthasarathy. Two applications of measurable selections. Lecture Notes in Mathematics, 263:58–64, 1972. CODEN LN-MAA2. ISBN 3-540-05818-4 (print), 3-540-37464-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058353/.

Parthasarathy:1972:NMC

[Par72l] T. Parthasarathy. Von Neumann’s measurable choice theorem. Lecture Notes in Mathematics, 263:65–75, 1972. CODEN LN-MAA2. ISBN 3-540-05818-4 (print), 3-540-37464-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058354/.

Patnaik:1970:KFP

[Pat70] Surendra-Nath Patnaik. On the Kakutani fixed point theorem and its relationship with the selection problem I. Lecture Notes in Mathematics, 171:70–73, 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069724/.

Patterson3rd:1974:IMS

[Pat74] Walter Mead Patterson 3rd. Iterative Methods for the Solution of a Linear Operator Equation in Hilbert Space — a Survey, volume 394 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06805-8 (print), 3-540-38455-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 170 (est.) pp. LCCN QA3 .L28 no. 394. URL http:/

/link.springer.com/book/10.1007/BFb0061519;   http://
www.springerlink.com/content/978-3-540-38455-7.

Pawlak:1970:DAA

[Paw70] Z. Pawlak. Definitional approach to automatic demonstration. Lecture Notes in Mathematics, 125:191–193, 1970. CODEN LN-MAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060632/.

Payne:1973:SGR

[Pay73] L. E. Payne. Some general remarks on improperly posed problems for partial differential equations. Lecture Notes in Mathematics, 316:1–30, 1973. CODEN LNMAA2. ISBN 3-540-06159-2 (print), 3-540-38370-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069621/.

Powell:1971:DCA

[PBR71] M. J. D. Powell, I. Barrodale, and F. D. K. Roberts. The differential correction algorithm for rational L_{∞} approximation. Lecture Notes in Mathematics, 228:169, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069455.

Pour-El:1973:ACV

[PE73] Marian Boykan Pour-El. Abstract computability versus analoggenerability (a survey). Lecture Notes in Mathematics, 337:345– 360, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066781/.

Peckham:1974:HPG

[Pec74] I. A. Peckham. The Hamiltonian product of graphs. Lecture Notes in Mathematics, 403:86–95, 1974. CODEN LN-MAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057380/.

Pedrini:1973:CPE

[Ped73] Claudio Pedrini. On the K_{0} of certain polynomial extensions. Lecture Notes in Mathematics, 342:92–108, 1973. CODEN LN-MAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073721/.

Pereyra:1974:VOV

[Per74] Victor Pereyra. Variable order variable step finite difference methods for nonlinear boundary value problems. Lecture Notes in Mathematics, 363:118–133, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069130/.

Peterson:1970:SAA

[Pet70a] F. P. Peterson, editor. The Steenrod Algebra and Its Applications: a Conference to Celebrate N. E. Steenrod’s Sixtieth Birthday: Proceedings of the Conference held at the Battelle Memorial Institute, Columbus, Ohio March 30th–April 4th, 1970, volume 168 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 168. URL http:/

/link.springer.com/book/10.1007/BFb0058513;   http://
www.springerlink.com/content/978-3-540-36437-5.

Peterson:1970:CHA

[Pet70b] Franklin P. Peterson. Coalgebras over Hopf algebras. Lecture Notes in Mathematics, 168:246–249, 1970. CODEN LN-MAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058526/.

Petrie:1970:RTS

[Pet70c] Ted Petrie. Representation theory, surgery and free actions of finite groups on varieties and homotopy spheres. Lecture Notes in Mathematics, 168:250–266, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058527/.

Petrie:1972:IHC

[Pet72] Ted Petrie. Involutions on homotopy complex projective spaces and related topics. Lecture Notes in Mathematics, 298:234– 259, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070046/.

Petrie:1973:IET

[Pet73] Ted Petrie. Induction in equivariant K-theory and geometric applications. Lecture Notes in Mathematics, 343:540–572, 1973. CODEN LNMAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061379/.

Petolla:1974:QCM

[Pet74a] G. Petolla. Questionnaires de coût moyen minimal lorsque les coûts sont lies à la base des questions. (French) []. Lecture Notes in Mathematics, 398:165–185, 1974. CODEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065776/.

Petrich:1974:LTV

[Pet74b] Mario Petrich. Linearly topologized vector spaces and rings. Lecture Notes in Mathematics, 380:112–151, 1974. CODEN LN-MAA2. ISBN 3-540-06730-2 (print), 3-540-38314-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061473/.

Petrich:1974:RS

[Pet74c] Mario Petrich. Rings and Semigroups, volume 380 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06730-2 (print), 3-540-38314-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 182 pp. LCCN QA3 .L28 no. 380. URL http:/

/link.springer.com/book/10.1007/BFb0061470;   http://
www.springerlink.com/content/978-3-540-38314-7.

Petrich:1974:SRL

[Pet74d] Mario Petrich. Semigroups and rings of linear transformations. Lecture Notes in Mathematics, 380:1–64, 1974. CODEN LN-MAA2. ISBN 3-540-06730-2 (print), 3-540-38314-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061471/.

Petrich:1974:SRM

[Pet74e] Mario Petrich. Semiprime rings with minimal one-sided ideals. Lecture Notes in Mathematics, 380:65–111, 1974. CODEN LN-MAA2. ISBN 3-540-06730-2 (print), 3-540-38314-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061472/.

Pfeiffer:1970:OSO

[Pfe70] C. G. Pfeiffer. On optimal stochastic orbital transfer strategy. Lecture Notes in Mathematics, 112:128–138, 1970. CODEN LN-MAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060202/.

Pham:1971:DEI

[Pha71] F. Pham. Deformations Équisingulières des idéaux jacobiens de courbes planes. (French) []. Lecture Notes in Mathematics, 209: 218–233, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068907/.

Phillips:1973:CSE

[Phi73] Richard E. Phillips. Centrally separated elements in groups. Lecture Notes in Mathematics, 319:135–139, 1973. CODEN LN-MAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058938/.

Pierce:1972:CSA

[Pie72] R. S. Pierce. Closure spaces with applications to ring theory. Lecture Notes in Mathematics, 246:565–616, 1972. CODEN LN-MAA2. ISBN 3-540-05760-9 (print), 3-540-37181-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059570/.

Pimbley:1973:PPE

[Pim73] George H. Pimbley. On predator-prey equations simulating an immune response. Lecture Notes in Mathematics, 322:231– 243, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060569/.

Pisanelli:1971:DPF

[Pis71] Domingos Pisanelli. Développement de Pincherle. (French) []. Lecture Notes in Mathematics, 205:225–231, 1971. CODEN LN-MAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059842/.

Pisanelli:1974:ETM

[Pis74] Domingos Pisanelli. Un ’extension des théorémes de Malgrange et Martineau. (French) []. Lecture Notes in Mathematics, 384: 239–251, 1974. CODEN LNMAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057550/.

Pitrat:1970:HIU

[Pit70] Jacques Pitrat. Heuristic interest of using metatheorems. Lecture Notes in Mathematics, 125:194–206, 1970. CODEN LN-MAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060633/.

Pittnauer:1972:AES

[Pit72a] Franz Pittnauer. Asymptotische Entwicklungen spezieller Funktionen. (German) []. Lecture Notes in Mathematics, 301:114– 153, 1972. CODEN LNMAA2. ISBN 3-540-06090-1 (print), 3-540-38077-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059529/.

Pittnauer:1972:EGb

[Pit72b] Franz Pittnauer. Eindeutigkeitssätze. (German) []. Lecture Notes in Mathematics, 301:80–92, 1972. CODEN LNMAA2. ISBN 3-540-06090-1 (print), 3-540-38077-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059527/.

Pittnauer:1972:EGa

[Pit72c] Franz Pittnauer. Existenzsätze. (German) []. Lecture Notes in Mathematics, 301:49–79, 1972. CODEN LNMAA2. ISBN 3-540-06090-1 (print), 3-540-38077-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059526/.

Pittnauer:1972:GGB

[Pit72d] Franz Pittnauer. Grundeigenschaften. (German) [Basic properties]. Lecture Notes in Mathematics, 301:1–48, 1972. CODEN LNMAA2. ISBN 3-540-06090-1 (print), 3-540-38077-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059525/.

Pittnauer:1972:NAG

[Pit72e] Franz Pittnauer. Numerische Anwendungen. (German) []. Lecture Notes in Mathematics, 301:154–173, 1972. CODEN LN-MAA2. ISBN 3-540-06090-1 (print), 3-540-38077-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059530/.

Pittnauer:1972:SFS

[Pit72f] Franz Pittnauer. Struktureigenschaften. (French) [Structural properties]. Lecture Notes in Mathematics, 301:93–113, 1972. CODEN LNMAA2. ISBN 3-540-06090-1 (print), 3-540-38077-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059528/.

Pittnauer:1972:VAR

[Pit72g] Franz Pittnauer. Vorlesungen über asymptotische Reihen. (German) [Lectures on asymptotic series], volume 301 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CO-DEN LNMAA2. ISBN 3-540-06090-1 (print), 3-540-38077-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 186 pp. LCCN QA3 .L28 no. 301. URL http:/

/link.springer.com/book/10.1007/BFb0059524;   http://
www.springerlink.com/content/978-3-540-38077-1.

Pittnauer:1973:ALF

[Pit73] Franz Pittnauer. Asymptotische Lösungen Von Funktionalgleichungen. (German) []. Lecture Notes in Mathematics, 333:202– 210, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060698/.

Pleijel:1972:GFP

[Ple72] Å. Pleijel. Green’s functions for pairs of formally selfadjoint ordinary differential operators. Lecture Notes in Mathematics, 280: 131–146, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066924/.

Pontryagin:1970:PMC

[PMN70] L. S. Pontryagin, E. F. Mishenko, and P. Nepomiastchy. Le problème d’un mobile contrôlable poursuivi par un autre. (French) []. Lecture Notes in Mathematics, 132:307–310, 1970. CODEN LNMAA2. ISBN 3-540-04921-5 (print), 3-540-36275-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066689/.

Poenaru:1971:HTD

[Poé71a] Valentin Poénaru. Homotopy theory and differentiable singularities. Lecture Notes in Mathematics, 197:106–132, 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068614/.

Poenaru:1971:TJC

[Poe71b] Valentin Poenaru. Travaux de J. Cerf (isotopie et pseudoisotopie). (French) []. Lecture Notes in Mathematics, 180:155– 176, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058830/.

Poenaru:1972:TCF

[Poe72] Valentin Poenaru. Le théorème de s-cobordisme. (French) []. Lecture Notes in Mathematics, 244:197–219, 1972. CODEN LN-MAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058703/.

Poenaru:1974:CLT

[Poé74a] Prof. Dr. Valentin Poénaru. Calcul de l’Algèbre E(G); le théorème de Stallings sur les groupes à deux bouts. (French) []. Lecture Notes in Mathematics, 421:118–144, 1974. CODEN LN-MAA2. ISBN 3-540-06967-4 (print), 3-540-37272-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068171/.

Poenaru:1974:GDF

[Poé74b] Prof. Dr. Valentin Poénaru. Groupes Discrets. (French) [Discrete groups], volume 421 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06967-4 (print), 3-540-37272-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 216 pp. LCCN QA3 .L28 no. 421. URL http:/

/link.springer.com/book/10.1007/BFb0068167;   http://
www.springerlink.com/content/978-3-540-37272-1.

Poenaru:1974:TBF

[Poé74c] Prof. Dr. Valentin Poénaru. La théorie des bouts. (French) []. Lecture Notes in Mathematics, 421:44–80, 1974. CODEN LN-MAA2. ISBN 3-540-06967-4 (print), 3-540-37272-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068169/.

Poenaru:1974:TSG

[Poé74d] Prof. Dr. Valentin Poénaru. Le théorème de Stallings sur les groupes à une infinité de bouts. (French) []. Lecture Notes in Mathematics, 421:145–159, 1974. CODEN LNMAA2. ISBN 3-540-06967-4 (print), 3-540-37272-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068172/.

Poenaru:1974:PGS

[Poé74e] Prof. Dr. Valentin Poénaru. Pre-groupes et structures bipolaires. (French) []. Lecture Notes in Mathematics, 421:81–117, 1974. CODEN LNMAA2. ISBN 3-540-06967-4 (print), 3-540-37272-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068170/.

Poenaru:1974:PTG

[Poé74f] Prof. Dr. Valentin Poénaru. Préliminaires de théorie des groupes, e.a.d.s. (French) []. Lecture Notes in Mathematics, 421:1–43, 1974. CODEN LNMAA2. ISBN 3-540-06967-4 (print), 3-540-37272-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068168/.

Poenaru:1974:VDT

[Poé74g] Prof. Dr. Valentin Poénaru. Variétés de dimension 3: Théorème de Kneser–Grushko–Stallings, les trois théorèmes de Papakyriakopoulos, Applications. (French) []. Lecture Notes in Mathematics, 421:160–216, 1974. CODEN LNMAA2. ISBN 3-540-06967-4 (print), 3-540-37272-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068173/.

Poenaru:1974:ADF

[Poé74h] Valentin Poénaru. Analyse Différentielle. (French) [Differential analysis], volume 371 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 228 pp. LCCN QA3 .L28 no. 371. URL http:/

/link.springer.com/book/10.1007/BFb0060715;   http://
www.springerlink.com/content/978-3-540-37934-8.

Poenaru:1974:AAC

[Poe74i] Valentin Poenaru. Appendice au chapitre II. (French) [Appendix to Chapter II]. Lecture Notes in Mathematics, 371:64–65, 1974. CODEN LNMAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060719.

Poenaru:1974:CAG

[Poe74j] Valentin Poenaru. Caractérisation des applications (et des germes d’applications) stables. (French) []. Lecture Notes in Mathematics, 371:210–227, 1974. CODEN LNMAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060727/.

Poenaru:1974:GAS

[Poe74k] Valentin Poenaru. Germes d’applications C^{∞}. (Stabilité, jets k-suffisants, e.a.d.s.). (French) []. Lecture Notes in Mathematics, 371:166–209, 1974. CODEN LNMAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060726/.

Poenaru:1974:I

[Poe74l] Valentin Poenaru. Introduction. Lecture Notes in Mathematics, 371:103, 1974. CODEN LNMAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060723.

Poenaru:1974:IFI

[Poe74m] Valentin Poenaru. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 371:1–10, 1974. CODEN LNMAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060716/.

Poenaru:1974:DDP

[Poe74n] Valentin Poenaru. La division des distributions à partir de la résolution des singularités de Hironaka. (French) []. Lecture Notes in Mathematics, 371:104–123, 1974. CODEN LNMAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060724/.

Poenaru:1974:TDM

[Poe74o] Valentin Poenaru. Le théorème de division de Mather. (French) []. Lecture Notes in Mathematics, 371:11–36, 1974. CODEN LNMAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060717/.

Poenaru:1974:TPW

[Poe74p] Valentin Poenaru. Le théorème de preparation de Weierstrass– Malgrange–Mather. (French) []. Lecture Notes in Mathematics, 371:37–63, 1974. CODEN LNMAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060718/.

Poenaru:1974:TRL

[Poe74q] Valentin Poenaru. Le théorème ‘de récollement’ de Łojasiewicz. (French) []. Lecture Notes in Mathematics, 371:80–89, 1974. CO-DEN LNMAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060721/.

Poenaru:1974:TSS

[Poe74r] Valentin Poenaru. Le théorème de synthèse spectrale de Whitney. (French) [The Whitney theorem on spectral synthesis]. Lecture Notes in Mathematics, 371:90–102, 1974. CODEN LN-MAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060722/.

Poenaru:1974:TEW

[Poe74s] Valentin Poenaru. Le théorème d’extension de Whitney. (French) []. Lecture Notes in Mathematics, 371:66–79, 1974. CODEN LNMAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060720/.

Poenaru:1974:SAD

[Poe74t] Valentin Poenaru. Sur la stabilité des applications différentiables (stabilité infinitésimale ⇒ stabilité). (French) []. Lecture Notes in Mathematics, 371:124–165, 1974. CODEN LNMAA2. ISBN 3-540-06665-9 (print), 3-540-37934-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060725/.

Pohl:1971:SDG

[Poh71] W. F. Pohl. Singularities in the differential geometry of submanifolds. Lecture Notes in Mathematics, 209:104–113, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068895/.

Popp:1970:ASB

[Pop70a] Herbert Popp. Anwendungen der Sätze (11.4) und (12.1). Besonderheiten bei wilder Verzweigung. Beispiele und abschliessende Bemerkungen über Überlagerungen von Kurven. (German) []. Lecture Notes in Mathematics, 176:116–126, 1970. CODEN LN-MAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068088/.

Popp:1970:ABG

[Pop70b] Herbert Popp. Anwendungen und Beispiele. (German) []. Lecture Notes in Mathematics, 176:57–66, 1970. CODEN LN-MAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068082/.

Popp:1970:VZV

[Pop70c] Herbert Popp. Das Verhalten zahm verzweigter Überlageruncen in Lokalen. (German) []. Lecture Notes in Mathematics, 176: 30–37, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068079/.

Popp:1970:FNS

[Pop70d] Herbert Popp. Die Fundamentalgruppe normaler Schemata. Der klassische Fall. (German) []. Lecture Notes in Mathematics, 176: 1–15, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068076/.

Popp:1970:SFE

[Pop70e] Herbert Popp. Die Struktur der Faktorkommutatorgruppe von Π_{1}^{(z)}(P^{n} −C). Erzeugende und Relationen für Π_{1}^{(z)}(p^{n} −C), wenn C nur normale Schnitte als Singularitäten hat. (German) [The structure of the factor commutator group of Π_{1}^{(z)}(P^{n} − C). generators and relations for Π_{1}^{(z)}(p^{n} −C) if C has only standard cuts as singularities]. Lecture Notes in Mathematics, 176:38–48, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068080/.

Popp:1970:SPTb

[Pop70f] Herbert Popp. Die Struktur des p-primen Teils der Fundamentalgruppe einer in n Punkten Punktierten, projektiven Kurve in Charakteristik p > 0. (German) []. Lecture Notes in Mathematics, 176:106–115, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068087/.

Popp:1970:SPTa

[Pop70g] Herbert Popp. Die Struktur des p-primen Teils der Fundamentalgruppe einer irreduziblen, projektiven und regulären Kurve vom Geschlecht g in Charakteristik p > 0. (German) []. Lecture Notes in Mathematics, 176:98–105, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068086/.

Popp:1970:SFI

[Pop70h] Herbert Popp. Die Struktur von Π_{1}^{(z)}(X − C), falls X eine irreduzible, reguläre, projektive Fläche ist und die Kurve C “nicht zu grosse” Singularitäten hat. (German) []. Lecture Notes in Mathematics, 176:49–56, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068081/.

Popp:1970:EWR

[Pop70i] Herbert Popp. Ein Erzeugendensystem für Π_{1}^{(z)}(X − C), wenn X eine reguläre, einfach zusammenhängende, projektive Mannigfaltigkeit der Dimension ≥ 2 ist und die hyperfläche C nur normale Schnitte als Singularitäten hat. (German) []. Lecture Notes in Mathematics, 176:24–29, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068078/.

Popp:1970:EOF

[Pop70j] Herbert Popp. Einige offene Fragen. (German) []. Lecture Notes in Mathematics, 176:146–152, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068090/.

Popp:1970:EUK

[Pop70k] Herbert Popp. Einiges über Überlagerungen von Kurven. (German) []. Lecture Notes in Mathematics, 176:67–79, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068083/.

Popp:1970:FAM

[Pop70l] Herbert Popp. Fundamentalgruppen algebraischer Mannigfaltigkeiten, volume 176 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 156 pp. LCCN QA3 .L28 no. 176.

URL http://link.springer.com/book/10.1007/BFb0068075;
 http://www.springerlink.com/content/978-3-540-36446-
7.

Popp:1970:UPU

[Pop70m] Herbert Popp. Überlagerungen von Produkten. Unabhängigkeit von Π_{1} (X) bei Konstantenerweiterung. Deformation und Hochheben Etaler Überlagerungen. (German) []. Lecture Notes in Mathematics, 176:80–89, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068084/.

Popp:1970:VSA

[Pop70n] Herbert Popp. Vier Sätze der algebraischen Geometrie. (German) []. Lecture Notes in Mathematics, 176:16–23, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068077/.

Popp:1970:ZFV

[Pop70o] Herbert Popp. Zurück Zu Flächen. Das Verhalten von Π_{1}(X − C), wenn die Kurve C in einer algebraischen Familie auf der Fläche X variiert. Anwendungen. (German) []. Lecture Notes in Mathematics, 176:127–145, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068089/.

Popp:1970:ZKH

[Pop70p] Herbert Popp. Zurück zu Kurven. Hochheben von Kurven nach Charakteristik O unter Erhaltung des Geschlechts. (German) [Back to curves. Lifting of curves of characteristic O with preservation of gender]. Lecture Notes in Mathematics, 176:90–97, 1970. CODEN LNMAA2. ISBN 3-540-05324-7 (print), 3-540-36446-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068085/.

Popp:1974:MAM

[Pop74a] Herbert Popp. Modulräume algebraischer Mannigfaltigkeiten. (German) []. Lecture Notes in Mathematics, 412:219–242, 1974. CODEN LNMAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066161/.

Popp:1974:CAV

[Pop74b] Prof. Dr. Herbert Popp. Classification of Algebraic Varieties and Compact Complex Manifolds, volume 412 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974.

CODEN LNMAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 333 pp. LCCN QA3 .L28 no. 412. URL http:/

/link.springer.com/book/10.1007/BFb0066151;   http://
www.springerlink.com/content/978-3-540-37877-8.

Porteous:1971:GDT

[Por71a] I. R. Porteous. Geometric differentiation — a Thomist view of differential geometry. Lecture Notes in Mathematics, 209:122– 127, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068899/.

Porteous:1971:SSM

[Por71b] I. R. Porteous. Simple singularities of maps. Lecture Notes in Mathematics, 192:286–307, 1971. CODEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066829/.

Porter:1971:USM

[Por71c] Gerald J. Porter. The universal H-space of a monoid. Lecture Notes in Mathematics, 196:63–67, 1971. CODEN LN-MAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067748/.

Powers:1970:LFP

[Pow70] M. J. Powers. Lefschetz fixed point theorems for multi-valued maps. Lecture Notes in Mathematics, 171:74–81, 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069725/.

Parikh:1974:CPU

[PP74a] Rohit Parikh and Milton Parnes. Conditional probabilities and uniform sets. Lecture Notes in Mathematics, 369:180–194, 1974. CODEN LNMAA2. ISBN 3-540-06656-X (print), 3-540-37928-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066012/.

Petolla:1974:ASQ

[PP74b] Sylvette Petolla and Claude-François Picard. Axiomatique et systématique des questionnaires. (French) []. Lecture Notes in Mathematics, 398:126–151, 1974. CODEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065774/.

Prawitz:1970:PPM

[Pra70] Dag Prawitz. A proof procedure with matrix reduction. Lecture Notes in Mathematics, 125:207–214, 1970. CODEN LN-MAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060634/.

Priddy:1973:TSG

[Pri73a] Stewart B. Priddy. Transfer, symmetric groups, and stable homotopy theory. Lecture Notes in Mathematics, 341:244–255, 1973. CODEN LNMAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067060/.

Prikry:1973:DCU

[Pri73b] Karel Prikry. On descendingly complete ultrafilters. Lecture Notes in Mathematics, 337:459–488, 1973. CODEN LN-MAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066785/.

Priouret:1974:PDE

[Pri74] P. Priouret. Processus de diffusion et Équations différentielles stochastiques. (French) []. Lecture Notes in Mathematics, 390: 37–113, 1974. CODEN LNMAA2. ISBN 3-540-06816-3 (print), 3-540-37765-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0082434/.

Propoy:1970:TDG

[Pro70] A. Propoy. On the theory of dynamic games. Lecture Notes in Mathematics, 112:80–94, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060198/.

Prolla:1974:MCF

[Pro74] João B. Prolla. Modules of continuous functions. Lecture Notes in Mathematics, 384:123–128, 1974. CODEN LNMAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057543/.

Pjateckii-Sapiro:1970:RFM

[PŠ70] I. I. Pjateckii-Šapiro. Reduction of the fields of modular functions and the rings of functions on p-adic manifolds. Lecture Notes in Mathematics, 155:151–164, 1970. CODEN LNMAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060323/.

Parthasarathy:1972:ACF

[PS72a] K. R. Parthasarathy and K. Schmidt. The analysis of cocycles of first order on some special groups. Lecture Notes in Mathematics, 272:83–104, 1972. CODEN LNMAA2. ISBN 3-540-05908-3 (print), 3-540-37607-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058343/.

Parthasarathy:1972:LTU

[PS72b] K. R. Parthasarathy and K. Schmidt. Limit theorems for uniformly infinitesimal families of positive definite kernels. Lecture Notes in Mathematics, 272:48–83, 1972. CODEN LN-MAA2. ISBN 3-540-05908-3 (print), 3-540-37607-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058342/.

Parthasarathy:1972:PDKa

[PS72c] K. R. Parthasarathy and K. Schmidt. Positive definite kernels and continuous tensor products. Lecture Notes in Mathematics, 272:1–48, 1972. CODEN LNMAA2. ISBN 3-540-05908-3 (print), 3-540-37607-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058341/.

Parthasarathy:1972:PDKb

[PS72d] K. R. Parthasarathy and Klaus Schmidt. Positive Definite Kernels, Continuous Tensor Products, and Central Limit Theorems of Probability Theory, volume 272 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LN-MAA2. ISBN 0-387-05908-3, 3-540-05908-3 (print), 3-540-37607-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 107 pp. LCCN QA3 .L28 no. 272. URL http:

//link.springer.com/book/10.1007/BFb0058340;   http:/
/www.springerlink.com/content/978-3-540-37607-1.

Piateckii-Shapiro:1973:ZFM

[PS73] I. I. Piateckii-Shapiro. Zeta-functions of modular curves. Lecture Notes in Mathematics, 349:317–360, 1973. CODEN LNMAA2. ISBN 3-540-06558-X (print), 3-540-37855-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-37855-6_5/.

Padgett:1973:GI

[PT73a] W. J. Padgett and R. L. Taylor. General introduction. Lecture Notes in Mathematics, 360:1–3, 1973. CODEN LNMAA2. ISBN 3-540-06585-7 (print), 3-540-37904-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069732/.

Padgett:1973:LLNb

[PT73b] W. J. Padgett and R. L. Taylor. Laws of large numbers for Fréchet spaces. Lecture Notes in Mathematics, 360:84–91, 1973. CODEN LNMAA2. ISBN 3-540-06585-7 (print), 3-540-37904-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069738/.

Padgett:1973:LLNa

[PT73c] W. J. Padgett and R. L. Taylor. Laws of large numbers for random variables and separable Hilbert spaces. Lecture Notes in Mathematics, 360:31–40, 1973. CODEN LNMAA2. ISBN 3-540-06585-7 (print), 3-540-37904-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069735/.

Padgett:1973:MP

[PT73d] W. J. Padgett and R. L. Taylor. Mathematical preliminaries. Lecture Notes in Mathematics, 360:4–14, 1973. CODEN LN-MAA2. ISBN 3-540-06585-7 (print), 3-540-37904-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069733/.

Padgett:1973:RES

[PT73e] W. J. Padgett and R. L. Taylor. Random elements in separable metric spaces. Lecture Notes in Mathematics, 360:15–30, 1973. CODEN LNMAA2. ISBN 3-540-06585-7 (print), 3-540-37904-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069734/.

Padgett:1973:SA

[PT73f] W. J. Padgett and R. L. Taylor. Some applications. Lecture Notes in Mathematics, 360:92–103, 1973. CODEN LNMAA2. ISBN 3-540-06585-7 (print), 3-540-37904-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069739/.

Padgett:1973:SLL

[PT73g] W. J. Padgett and R. L. Taylor. Strong laws of large numbers for normed linear spaces. Lecture Notes in Mathematics, 360: 41–65, 1973. CODEN LNMAA2. ISBN 3-540-06585-7 (print), 3-540-37904-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069736/.

Padgett:1973:WLL

[PT73h] W. J. Padgett and R. L. Taylor. Weak laws of large numbers for normed linear spaces. Lecture Notes in Mathematics, 360: 66–83, 1973. CODEN LNMAA2. ISBN 3-540-06585-7 (print), 3-540-37904-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069737/.

Padgett:1973:LLNc

[PT73i] W. J. Padgett and Robert Lee Taylor. Laws of Large Numbers for Normed Linear Spaces and Certain Fréchet Spaces, volume 360 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 0-387-06585-7, 3-540-06585-7 (print), 3-540-37904-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 111 pp. LCCN QA3 .L28 no. 360; QA3 .L28 no. 360 QA273.43. URL http:/

/link.springer.com/book/10.1007/BFb0069731;   http://
www.springerlink.com/content/978-3-540-37904-1.

Pukanszky:1972:AAG

[Puk72] L. Pukanszky. Action of algebraic groups of automorphisms on the dual of a class of type I groups. Lecture Notes in Mathematics, 266:273–279, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059650/.

Pultr:1970:RAC

[Pul70] A. Pultr. The right adjoints into the categories of relational systems. Lecture Notes in Mathematics, 137:100–113, 1970. CODEN LNMAA2. ISBN 3-540-04926-6 (print), 3-540-36292-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060441/.

Quillen:1973:FGG

[Qui73a] Daniel Quillen. Finite generation of the groups K_{i} of rings of algebraic integers. Lecture Notes in Mathematics, 341:179–198, 1973. CODEN LNMAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067056/.

Quillen:1973:HAT

[Qui73b] Daniel Quillen. Higher algebraic K-theory: I. Lecture Notes in Mathematics, 341:85–147, 1973. CODEN LNMAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067053/.

Quintana:1973:ARB

[Qui73c] Ricardo B. Quintana, Jr. An attack on the restricted Burnside problem for groups of exponent 8 on 2 generators. Lecture Notes in Mathematics, 319:140–147, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058939/.

Rabinowitz:1971:NSL

[Rab71] Paul H. Rabinowitz. A nonlinear Sturm–Liouville theorem. Lecture Notes in Mathematics, 243:182–192, 1971. CODEN LN-MAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058727/.

Rado:1974:RTI

[Rad74] Richard Rado. Reconstruction theorems for infinite hypergraphs. Lecture Notes in Mathematics, 411:140–146, 1974. CODEN LN-MAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066188/.

Rahilly:1974:DCC

[Rah74] Alan Rahilly. Derivable chains containing generalized Hall planes. Lecture Notes in Mathematics, 403:96–100, 1974. CODEN LN-MAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057381/.

Ramalho:1974:NEC

[Ram74a] R. Ramalho. A note on the eigenvalues of compact operators. Lecture Notes in Mathematics, 384:260–270, 1974. CODEN LN-MAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057552/.

Ramis:1974:SRR

[Ram74b] J.-P. Ramis. Some recent results and open problems in infinite dimensional analytic geometry. Lecture Notes in Mathematics, 364: 186–195, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069018/.

Ramis:1974:TFP

[Ram74c] Jean-Pierre Ramis. Théorèmes de finitude pour les espaces pconvexes, q-concaves et (p, q)-convexes-concaves. (French) []. Lecture Notes in Mathematics, 409:357–393, 1974. CODEN LN-MAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068111/.

Rang:1973:NIE

[Ran73a] E. R. Rang. A nonlinear integral equation relating distillation processes. Lecture Notes in Mathematics, 312:151–163, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060053/.

Ranicki:1973:ATI

[Ran73b] A. A. Ranicki. Algebraic L-theory III. twisted Laurent extensions. Lecture Notes in Mathematics, 343:412–463, 1973. CODEN LNMAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061373/.

Rao:1970:GM

[Rao70] M. M. Rao. Generalized martingales. Lecture Notes in Mathematics, 160:241–261, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060657/.

Rao:1972:CAS

[Rao72a] K. Nageswara Rao. On certain arithmetical sums. Lecture Notes in Mathematics, 251:181–192, 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058793/.

Rao:1972:DDB

[Rao72b] Murali Rao. Doob decomposition and Burkholder inequalities. Lecture Notes in Mathematics, 258:198–201, 1972. CODEN LN-MAA2. ISBN 3-540-05773-0 (print), 3-540-37183-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059477/.

Ratliff:1973:CCD

[Rat73] Louis J. Ratliff, Jr. Chain conjectures and H-domains. Lecture Notes in Mathematics, 311:222–238, 1973. CODEN LN-MAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068931/.

Rautmann:1972:NSP

[Rau72] Reimund Rautmann. Ein Näherungsverfahren für spezielle parabolische Anfangswertaufgaben mit Operatoren. (German) []. Lecture Notes in Mathematics, 267:187–231, 1972. CODEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061620/.

Rautmann:1974:LAS

[Rau74] Reimund Rautmann. Lösung einer Anfangswertaufgabe für substantielle Differentialgleichungen. (German) []. Lecture Notes in Mathematics, 395:85–108, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060665/.

Raviart:1971:RNC

[Rav71] P.-A. Raviart. Resolution numérique de certains problèmes hyperboliques non linéaires. Méthode de pseudo–viscosité. (French) []. Lecture Notes in Mathematics, 228:170–206, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069456/.

Raynaud:1970:CSTb

[Ray70a] M. Raynaud. Complements sur les sous-Tores d’un préschéma en groupes. Application aux groupes lisses. (French) []. Lecture Notes in Mathematics, 152:349–483, 1970. CODEN LN-MAA2. ISBN 3-540-05180-5 (print), 3-540-36337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059013/.

Raynaud:1970:GAU

[Ray70b] M. Raynaud. Groupes algébriques unipotents extensions entre groupes unipotents et groupes de type multiplicatif. (French) []. Lecture Notes in Mathematics, 152:532–631, 1970. CODEN LNMAA2. ISBN 3-540-05180-5 (print), 3-540-36337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059015/.

Raynaud:1970:GRU

[Ray70c] M. Raynaud. Groupes de rang unipotent nul. (French) []. Lecture Notes in Mathematics, 152:484–531, 1970. CODEN LN-MAA2. ISBN 3-540-05180-5 (print), 3-540-36337-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059014/.

Raynaud:1970:AGF

[Ray70d] Michel Raynaud. Action d’un groupe fini sur un anneau. (French) []. Lecture Notes in Mathematics, 169:103–111, 1970. CODEN LNMAA2. ISBN 3-540-05283-6 (print), 3-540-36406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069581/.

Raynaud:1970:ALH

[Ray70e] Michel Raynaud. Anneaux locaux henséliens. (French) [Henselian local rings], volume 169 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05283-6 (print), 3-540-36406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 129 pp. LCCN QA3 .L28 no. 169.

URL http://link.springer.com/book/10.1007/BFb0069571;
 http://www.springerlink.com/content/978-3-540-36406-
1.

Raynaud:1970:AUG

[Ray70f] Michel Raynaud. Anneaux unibranches et géométriquement unibranches. (French) []. Lecture Notes in Mathematics, 169:97– 102, 1970. CODEN LNMAA2. ISBN 3-540-05283-6 (print), 3-540-36406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069580/.

Raynaud:1970:AAF

[Ray70g] Michel Raynaud. Applications au foncteur de Picard et aux correspondances divisorielles. (French) []. Lecture Notes in Mathematics, 119:37–55, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059509/.

Raynaud:1970:CGE

[Ray70h] Michel Raynaud. Cas des groupes et des espaces homogènes. (French) []. Lecture Notes in Mathematics, 119:95–102, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059512/.

Raynaud:1970:CSG

[Ray70i] Michel Raynaud. Cas des schémas en groupes lisses à fibres affines et des espaces homogènes sous iceux. (French) []. Lecture Notes in Mathematics, 119:103–117, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059513/.

Raynaud:1970:COB

[Ray70j] Michel Raynaud. Cas où la base est un schéma régulier de dimension 1. (French) []. Lecture Notes in Mathematics, 119:118– 125, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059514/.

Raynaud:1970:CSTa

[Ray70k] Michel Raynaud. Compléments sur les sous-tores d’un schéma en groupes. (French) []. Lecture Notes in Mathematics, 119:126– 144, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059515/.

Raynaud:1970:CEQ

[Ray70l] Michel Raynaud. Contre-exemples et questions ouvertes. (French) []. Lecture Notes in Mathematics, 119:213–216, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059521/.

Raynaud:1970:CHF

[Ray70m] Michel Raynaud. Couples henséliens. (French) []. Lecture Notes in Mathematics, 169:112–128, 1970. CODEN LNMAA2. ISBN 3-540-05283-6 (print), 3-540-36406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069582/.

Raynaud:1970:EHC

[Ray70n] Michel Raynaud. Elimination des hypothèses de connexité. (French) []. Lecture Notes in Mathematics, 119:145–158, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059516/.

Raynaud:1970:EAN

[Ray70o] Michel Raynaud. Étude des algèbres nettes. (French) []. Lecture Notes in Mathematics, 169:21–38, 1970. CODEN LN-MAA2. ISBN 3-540-05283-6 (print), 3-540-36406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069574/.

Raynaud:1970:EAE

[Ray70p] Michel Raynaud. Exemples d’algèbres étales. (French) []. Lecture Notes in Mathematics, 169:66–71, 1970. CODEN LN-MAA2. ISBN 3-540-05283-6 (print), 3-540-36406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069577/.

Raynaud:1970:ESA

[Ray70q] Michel Raynaud. Exemples de schémas abéliens non localement projectifs. (French) []. Lecture Notes in Mathematics, 119:185– 191, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059518/.

Raynaud:1970:FASa

[Ray70r] Michel Raynaud. Faisceaux amples sur les schémas à groupe d’opérateur. (French) []. Lecture Notes in Mathematics, 119:74– 94, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059511/.

Raynaud:1970:FASb

[Ray70s] Michel Raynaud. Faisceaux amples sur les schémas en groupes et les espaces homogènes, volume 119 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 218 pp. LCCN QA3 .L28 no. 119.

URL http://link.springer.com/book/10.1007/BFb0059504;
 http://www.springerlink.com/content/978-3-540-36249-

Raynaud:1970:FMN

[Ray70t] Michel Raynaud. Formalisme des morphismes nets et étales. Exemples. (French) []. Lecture Notes in Mathematics, 169:13– 20, 1970. CODEN LNMAA2. ISBN 3-540-05283-6 (print), 3-540-36406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069573/.

Raynaud:1970:HF

[Ray70u] Michel Raynaud. Hensélisation. (French) []. Lecture Notes in Mathematics, 169:80–96, 1970. CODEN LNMAA2. ISBN 3-540-05283-6 (print), 3-540-36406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069579/.

Raynaud:1970:IFI

[Ray70v] Michel Raynaud. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 119:1–3, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059505/.

Raynaud:1970:TDC

[Ray70w] Michel Raynaud. Le théorème du cube. (French) []. Lecture Notes in Mathematics, 119:56–73, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059510/.

Raynaud:1970:MQF

[Ray70x] Michel Raynaud. Morphismes quasi-finis. Main theorem de Zariski. (French) []. Lecture Notes in Mathematics, 169:39– 50, 1970. CODEN LNMAA2. ISBN 3-540-05283-6 (print), 3-540-36406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069575/.

Raynaud:1970:NAL

[Ray70y] Michel Raynaud. Notion d’anneau local hensélien. (French) []. Lecture Notes in Mathematics, 169:1–12, 1970. CODEN LN-MAA2. ISBN 3-540-05283-6 (print), 3-540-36406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069572/.

Raynaud:1970:PPE

[Ray70z] Michel Raynaud. Propriétés de permanence. Exemples d’anneaux henséliens. (French) []. Lecture Notes in Mathematics, 169:72– 79, 1970. CODEN LNMAA2. ISBN 3-540-05283-6 (print), 3-540-36406-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069578/.

Raynaud:1970:QPG

[Ray70-27] Michel Raynaud. Quasi-projectivité globale de certains schémas en groupes. Application à la descente. (French) []. Lecture Notes in Mathematics, 119:159–184, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059517/.

Raynaud:1970:QNF

[Ray70-28] Michel Raynaud. Quelques notations. (French) []. Lecture Notes in Mathematics, 119:4–6, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059506/.

Raynaud:1970:RDF

[Ray70-29] Michel Raynaud. Rappels sur les diviseurs et les faisceaux inversibles. (French) []. Lecture Notes in Mathematics, 119:7– 13, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059507/.

Raynaud:1970:SLA

[Ray70-30] Michel Raynaud. Structure locale des algèbres nettes et étalés. Critère jacobien. (French) [Local structure of the sharp and spread algebras. Jacobian criterion]. Lecture Notes in Mathematics, 169:51–65, 1970. CODEN LNMAA2. ISBN 3-540-05283-6 (print), 3-540-36406-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069576/.

Raynaud:1970:FID

[Ray70-31] Michel Raynaud. Sur le foncteur image directe. (French) []. Lecture Notes in Mathematics, 119:14–36, 1970. CODEN LN-MAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059508/.

Raynaud:1970:TLI

[Ray70-32] Michel Raynaud. Torseurs localement isotriviaux sous un groupe algébrique. (French) []. Lecture Notes in Mathematics, 119:204– 212, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059520/.

Raynaud:1970:TSS

[Ray70-33] Michel Raynaud. Torseurs sous un schéma abélien. (French) []. Lecture Notes in Mathematics, 119:192–203, 1970. CODEN LNMAA2. ISBN 3-540-04908-8 (print), 3-540-36249-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059519/.

Raynaud:1971:TRM

[Ray71a] Michel Raynaud. Travaux récents de M. Artin. (French) []. Lecture Notes in Mathematics, 179:279–295, 1971. CODEN LN-MAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058818/.

Raynaud:1971:GAG

[Ray71b] Mme M. Raynaud. Géométrie Algébrique et Géométrie analytique. (French) []. Lecture Notes in Mathematics, 224:311–343, 1971. CODEN LNMAA2. ISBN 0-354-00361-4, 0-387-05614-9, 3-540-05614-9 (print), 3-540-36910-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058667/.

Raynaud:1971:PCF

[Ray71c] Mme M. Raynaud. Propriété cohomologique des faisceaux d’ensembles et des faisceaux de groupes non commutatifs. (French) []. Lecture Notes in Mathematics, 224:344–439, 1971. CODEN LNMAA2. ISBN 0-354-00361-4, 0-387-05614-9, 3-540-05614-9 (print), 3-540-36910-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058668/.

Rayner:1971:TCD

[Ray71d] C. B. Rayner. Thom’s cell decomposition as a stratified set. Lecture Notes in Mathematics, 192:150–152, 1971. CODEN LN-MAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066820/.

Raynaud:1972:CDM

[Ray72a] Michel Raynaud. Compactification du module des courbes. (French) []. Lecture Notes in Mathematics, 244:47–61, 1972. CO-DEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058696/.

Raynaud:1972:PFD

[Ray72b] Michèle Raynaud. Propriétés de finitude du groupe fondamental. (French) []. Lecture Notes in Mathematics, 288:25–31, 1972. CODEN LNMAA2. ISBN 3-540-05987-3 (print), 3-540-37984-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068690/.

Raynaud:1974:CAC

[Ray74] Michel Raynaud. Construction analytique de courbes en géométrie non archimédienne. (French) []. Lecture Notes in Mathematics, 383:171–185, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057308/.

rd:1974:GM

[rd74a] Walter Mead Patterson rd. Gradient methods. Lecture Notes in Mathematics, 394:118–170, 1974. CODEN LNMAA2. ISBN 3-540-06805-8 (print), 3-540-38455-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061524/.

rd:1974:I

[rd74b] Walter Mead Patterson rd. Introduction. Lecture Notes in Mathematics, 394:1–7, 1974. CODEN LNMAA2. ISBN 3-540-06805-8 (print), 3-540-38455-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061520/.

rd:1974:IMC

[rd74c] Walter Mead Patterson rd. Iterative methods in complex Hilbert spaces. Lecture Notes in Mathematics, 394:31–73, 1974. CODEN LNMAA2. ISBN 3-540-06805-8 (print), 3-540-38455-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061522/.

rd:1974:IMR

[rd74d] Walter Mead Patterson rd. Iterative methods in real Hilbert spaces. Lecture Notes in Mathematics, 394:8–30, 1974. CODEN LNMAA2. ISBN 3-540-06805-8 (print), 3-540-38455-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061521/.

rd:1974:SAM

[rd74e] Walter Mead Patterson rd. Successive approximation methods. Lecture Notes in Mathematics, 394:74–117, 1974. CODEN LN-MAA2. ISBN 3-540-06805-8 (print), 3-540-38455-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061523/.

Rector:1971:LSH

[Rec71] David L. Rector. Loop structures on the homotopy type of S^{3}. Lecture Notes in Mathematics, 249:99–105, 1971. CODEN LN-MAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060900/.

Reeken:1973:BTG

[Ree73] Michael Reeken. Bifurcation theory for gradient systems. Lecture Notes in Mathematics, 322:244–260, 1973. CODEN LN-MAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060570/.

Reeb:1974:PEP

[Ree74] G. Reeb. À propos de l’équation (de Painlevé) y′′ = 6y^{2} + x. (French) []. Lecture Notes in Mathematics, 392:43–46, 1974. CO-DEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0064109/.

Reiter:1971:ASA

[Rei71a] Hans Reiter. L^{1}-Algebras and Segal Algebras, volume 231 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 108 (est.) pp. LCCN QA3 .L28 no. 231. URL http:/

/link.springer.com/book/10.1007/BFb0060759;   http://
www.springerlink.com/content/978-3-540-36939-4.

Reiter:1971:ACA

[Rei71b] Hans Reiter. L^{1}-algebras of compact and of abelian groups. Lecture Notes in Mathematics, 231:95–108, 1971. CODEN LN-MAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060776/.

Reiter:1971:AUBa

[Rei71c] Hans Reiter. Approximate units in Banach algebras: definitions. Lecture Notes in Mathematics, 231:27–29, 1971. CODEN LNMAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060765/.

Reiter:1971:AUBb

[Rei71d] Hans Reiter. Approximate units in Banach algebras: lemmas. Lecture Notes in Mathematics, 231:30–33, 1971. CODEN LN-MAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060766/.

Reiter:1971:AUS

[Rei71e] Hans Reiter. Approximate units in Segal algebras. Lecture Notes in Mathematics, 231:34–38, 1971. CODEN LNMAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060767/.

Reiter:1971:CRI

[Rei71f] Hans Reiter. Canonical representation of the ideals. Lecture Notes in Mathematics, 231:51–55, 1971. CODEN LN-MAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060770/.

Reiter:1971:CIAa

[Rei71g] Hans Reiter. Certain ideals in L^{1}-algebras. Lecture Notes in Mathematics, 231:1–7, 1971. CODEN LNMAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060760/.

Reiter:1971:CIS

[Rei71h] Hans Reiter. Certain ideals in Segal algebras. Lecture Notes in Mathematics, 231:43–50, 1971. CODEN LNMAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060769/.

Reiter:1971:CIAb

[Rei71i] Hans Reiter. Closed ideals with approximate units in Segal algebras. Lecture Notes in Mathematics, 231:69–76, 1971. CODEN LNMAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060773/.

Reiter:1971:E

[Rei71j] Hans Reiter. Examples. Lecture Notes in Mathematics, 231: 24–26, 1971. CODEN LNMAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060764/.

Reiter:1971:PA

[Rei71k] Hans Reiter. The property P_{1}, applications. Lecture Notes in Mathematics, 231:8–11, 1971. CODEN LNMAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060761/.

Reiter:1971:SA

[Rei71l] Hans Reiter. Segal algebras. Lecture Notes in Mathematics, 231: 16–23, 1971. CODEN LNMAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060763/.

Reiter:1971:SAN

[Rei71m] Hans Reiter. Segal algebras and normal subgroups. Lecture Notes in Mathematics, 231:56–61, 1971. CODEN LNMAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060771/.

Reiter:1971:SAA

[Rei71n] Hans Reiter. Segal algebras on abelian groups. Lecture Notes in Mathematics, 231:85–94, 1971. CODEN LNMAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060775/.

Reiter:1971:SAQ

[Rei71o] Hans Reiter. Segal algebras on quotient groups. Lecture Notes in Mathematics, 231:62–68, 1971. CODEN LNMAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060772/.

Reiter:1971:SGR

[Rei71p] Hans Reiter. Some general results on Segal algebras. Lecture Notes in Mathematics, 231:39–42, 1971. CODEN LN-MAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060768/.

Reiter:1971:SPE

[Rei71q] Hans Reiter. Some properties equivalent to P_{1}. relativization. Lecture Notes in Mathematics, 231:12–15, 1971. CODEN LN-MAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060762/.

Reiter:1971:SSA

[Rei71r] Hans Reiter. Symmetric Segal algebras on compact groups. Lecture Notes in Mathematics, 231:77–84, 1971. CODEN LN-MAA2. ISBN 3-540-05651-3 (print), 3-540-36939-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060774/.

Reiter:1972:ASA

[Rei72] Hans Reiter. L^{1}-algebras and Segal algebras. Lecture Notes in Mathematics, 266:281–285, 1972. CODEN LNMAA2. ISBN 3-540-05856-7 (print), 3-540-37479-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059651/.

Reiji:1974:FSG

[Rei74a] Takahashi Reiji. Fonctions sphériques dans les groupes Sp(n, l). (French) []. Lecture Notes in Mathematics, 404:218–228, 1974. CODEN LNMAA2. ISBN 3-540-06842-2 (print), 3-540-37789-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060617/.

Reinhart:1974:HIF

[Rei74b] Bruce L. Reinhart. Holonomy invariants for framed foliations. Lecture Notes in Mathematics, 392:47–52, 1974. CODEN LN-MAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064110/.

Renyi:1973:UFC

[Rén73] Alfred Rényi. Uniform flows in cascade graphs. Lecture Notes in Mathematics, 296:1–18, 1973. CODEN LNMAA2. ISBN 3-540-06211-4 (print), 3-540-38485-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059818/.

Revuz:1970:ATM

[Rev70] Daniel Revuz. Application d’un théorème de Mokobodzky aux opérateurs potentiels dans le cas recurrent. (French) []. Lecture Notes in Mathematics, 124:208–215, 1970. CODEN LN-MAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059344/.

Roggenkamp:1970:HIE

[RHD70a] Klaus W. Roggenkamp and Verena Huber-Dyson. The Higman ideal and extensions of lattices. Lecture Notes in Mathematics, 115:243–271, 1970. CODEN LNMAA2. ISBN 3-540-04904-5 (print), 3-540-36237-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068801/.

Roggenkamp:1970:HA

[RHD70b] Klaus W. Roggenkamp and Verena Huber-Dyson. Homological algebra. Lecture Notes in Mathematics, 115:65–123, 1970. CO-DEN LNMAA2. ISBN 3-540-04904-5 (print), 3-540-36237-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068798/.

Roggenkamp:1970:LO

[RHD70c] Klaus W. Roggenkamp and Verena Huber-Dyson. Lattices over Orders I, volume 115 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04904-5 (print), 3-540-36237-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xix + 290 pp. LCCN QA3 .L28 no. 115. URL http:/

/link.springer.com/book/10.1007/BFb0068796;   http://
www.springerlink.com/content/978-3-540-36237-1.

Roggenkamp:1970:MO

[RHD70d] Klaus W. Roggenkamp and Verena Huber-Dyson. Maximal orders. Lecture Notes in Mathematics, 115:178–242, 1970. CODEN LNMAA2. ISBN 3-540-04904-5 (print), 3-540-36237-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068800/.

Roggenkamp:1970:MTS

[RHD70e] Klaus W. Roggenkamp and Verena Huber-Dyson. Morita theorems and separable algebras. Lecture Notes in Mathematics, 115: 124–177, 1970. CODEN LNMAA2. ISBN 3-540-04904-5 (print), 3-540-36237-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068799/.

Roggenkamp:1970:PRM

[RHD70f] Klaus W. Roggenkamp and Verena Huber-Dyson. Preliminaries on rings and modules. Lecture Notes in Mathematics, 115: 1–64, 1970. CODEN LNMAA2. ISBN 3-540-04904-5 (print), 3-540-36237-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068797/.

Rhoades:1974:FPI

[Rho74] B. E. Rhoades. Fixed point iterations using infinite matrices, II. Lecture Notes in Mathematics, 430:390–394, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066278/.

Riccia:1971:SSF

[Ric71] Giacomo Della Riccia. Stable semi-flows (one-parameter semigroups) on locally compact or complete metric spaces. Lecture Notes in Mathematics, 235:59–60, 1971. CODEN LN-MAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059195.

Rickart:1974:HED

[Ric74] C. E. Rickart. Holomorphic extensions and domains of holomorphy for general function algebras. Lecture Notes in Mathematics, 364:80–91, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069006/.

Rim:1972:FDT

[Rim72] D. S. Rim. Formal deformation theory. Lecture Notes in Mathematics, 288:32–132, 1972. CODEN LNMAA2. ISBN 3-540-05987-3 (print), 3-540-37984-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068691/.

Ringrose:1972:COA

[Rin72a] J. R. Ringrose. Cohomology of operator algebras. Lecture Notes in Mathematics, 247:355–434, 1972. CODEN LNMAA2. ISBN 3-540-05729-3 (print), 3-540-37117-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058555/.

Ringrose:1972:LTF

[Rin72b] J. R. Ringrose. Lectures on the trace in a finite von Neumann algebra. Lecture Notes in Mathematics, 247:309–354, 1972. CO-DEN LNMAA2. ISBN 3-540-05729-3 (print), 3-540-37117-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058554/.

Rivano:1972:C

[Riv72a] Neantro Saavedra Rivano. ⊗-categories. Lecture Notes in Mathematics, 265:11–86, 1972. CODEN LNMAA2. ISBN 3-540-05844-3 (print), 3-540-37477-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059110/.

Rivano:1972:CT

[Riv72b] Neantro Saavedra Rivano. Catégories Tannakiennes, volume 265 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05844-3 (print), 3-540-37477-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 379 (est.) pp. LCCN QA3 .L28 no. 265. URL http:/

/link.springer.com/book/10.1007/BFb0059108;   http://
www.springerlink.com/content/978-3-540-37477-0.

Rivano:1972:DTF

[Riv72c] Neantro Saavedra Rivano. Dualité de Tannaka. (French) []. Lecture Notes in Mathematics, 265:161–205, 1972. CODEN LN-MAA2. ISBN 3-540-05844-3 (print), 3-540-37477-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059112/.

Rivano:1972:ETG

[Riv72d] Neantro Saavedra Rivano. Exemples tires de la géométrie algébrique. (French) []. Lecture Notes in Mathematics, 265:305– 379, 1972. CODEN LNMAA2. ISBN 3-540-05844-3 (print), 3-540-37477-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059115/.

Rivano:1972:GFF

[Riv72e] Neantro Saavedra Rivano. Graduations et filtrations. (French) []. Lecture Notes in Mathematics, 265:207–237, 1972. CODEN LNMAA2. ISBN 3-540-05844-3 (print), 3-540-37477-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059113/.

Rivano:1972:I

[Riv72f] Neantro Saavedra Rivano. Introduction. Lecture Notes in Mathematics, 265:5–9, 1972. CODEN LNMAA2. ISBN 3-540-05844-3 (print), 3-540-37477-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059109/.

Rivano:1972:PCT

[Riv72g] Neantro Saavedra Rivano. Polarisations des catégories tannakiennes sur IR. (French) []. Lecture Notes in Mathematics, 265: 239–304, 1972. CODEN LNMAA2. ISBN 3-540-05844-3 (print), 3-540-37477-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059114/.

Rivano:1972:RLS

[Riv72h] Neantro Saavedra Rivano. Représentations linéaires de schémas en groupes affines. (French) []. Lecture Notes in Mathematics, 265:87–159, 1972. CODEN LNMAA2. ISBN 3-540-05844-3 (print), 3-540-37477-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059111/.

Raynaud:1971:TRR

[RK71] M. Raynaud and S. Kleiman. Un théorème de représentabilité relative sur le foncteur de Picard. (French) []. Lecture Notes in Mathematics, 225:595–615, 1971. CODEN LNMAA2. ISBN 3-540-05647-5 (print), 3-540-36936-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066295/.

Robin:1971:CDC

[Rob71a] G. Robin. Sur la cohomologie du complémentaire d’une hypersurface. (French) []. Lecture Notes in Mathematics, 205:204– 215, 1971. CODEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059840/.

Robinson:1971:GPC

[Rob71b] R. C. Robinson. Generic properties of conservative systems. Lecture Notes in Mathematics, 206:35–36, 1971. CODEN LN-MAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070136.

Robert:1973:BM

[Rob73a] Alain Robert. Back matter. Lecture Notes in Mathematics, 326: 254–264, 1973. CODEN LNMAA2. ISBN 3-540-06309-9 (print), 3-540-46916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-46916-2/1.pdf.

Robert:1973:CFCa

[Rob73b] Alain Robert. Compléments. (French) [Complements]. Lecture Notes in Mathematics, 326:225–253, 1973. CODEN LNMAA2. ISBN 3-540-06309-9 (print), 3-540-46916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-46916-2_4/.

Robert:1973:CFCb

[Rob73c] Alain Robert. Compléments. (French) [Complements]. Lecture Notes in Mathematics, 326:268, 1973. CODEN LNMAA2. ISBN 3-540-06309-9 (print), 3-540-46916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-46916-2_8.

Robert:1973:CECa

[Rob73d] Alain Robert. Complex elliptic curves. Lecture Notes in Mathematics, 326:1–74, 1973. CODEN LNMAA2. ISBN 3-540-06309-9 (print), 3-540-46916-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-46916-2_1/.

Robert:1973:CECb

[Rob73e] Alain Robert. Complex elliptic curves. Lecture Notes in Mathematics, 326:265, 1973. CODEN LNMAA2. ISBN 3-540-06309-9 (print), 3-540-46916-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-46916-2_5.

Robert:1973:DPa

[Rob73f] Alain Robert. Division points. Lecture Notes in Mathematics, 326:173–224, 1973. CODEN LNMAA2. ISBN 3-540-06309-9 (print), 3-540-46916-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-46916-2_3/.

Robert:1973:DPb

[Rob73g] Alain Robert. Division points. Lecture Notes in Mathematics, 326:267, 1973. CODEN LNMAA2. ISBN 3-540-06309-9 (print), 3-540-46916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-46916-2_7.

Robert:1973:ECCa

[Rob73h] Alain Robert. Elliptic curves in characteristic zero. Lecture Notes in Mathematics, 326:75–172, 1973. CODEN LNMAA2. ISBN 3-540-06309-9 (print), 3-540-46916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-46916-2_2/.

Robert:1973:ECCb

[Rob73i] Alain Robert. Elliptic curves in characteristic zero. Lecture Notes in Mathematics, 326:266, 1973. CODEN LNMAA2. ISBN 3-540-06309-9 (print), 3-540-46916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-46916-2_6.

Robert:1973:ECN

[Rob73j] Alain Robert, editor. Elliptic Curves: Notes from Postgraduate Lectures Given in Lausanne 1971/72, volume 326 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06309-9 (print), 3-540-46916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 326.

URL http://link.springer.com/book/10.1007/978-3-540-
46916-2;   http://www.springerlink.com/content/978-3-
540-46916-2.

Robert:1973:FAF

[Rob73k] Alain Robert. Formes automorphes sur Gl_{2}. (French) []. Lecture Notes in Mathematics, 317:295–318, 1973. CODEN LN-MAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069288/.

Robert:1973:FM

[Rob73l] Alain Robert. Front matter. Lecture Notes in Mathematics, 326: i–viii, 1973. CODEN LNMAA2. ISBN 3-540-06309-9 (print), 3-540-46916-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-46916-2/1.pdf.

Roberts:1973:CAT

[Rob73m] L. Roberts. Comparison of algebraic and topological K-theory. Lecture Notes in Mathematics, 342:74–78, 1973. CODEN LN-MAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073719/.

Roberts:1973:BCA

[Rob73n] Leslie Roberts. Base change for K_{0} of algebraic varieties. Lecture Notes in Mathematics, 342:122–134, 1973. CODEN LN-MAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073723/.

Robinson:1973:GWI

[Rob73o] Derek J. S. Robinson. Groups whose images have a transitive normality relation. Lecture Notes in Mathematics, 319:148– 155, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058940/.

Roberts:1974:CPF

[Rob74] Leigh Roberts. Characterisation of a pregeometry by its flats. Lecture Notes in Mathematics, 403:101–104, 1974. CODEN LN-MAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057382/.

Roggenkamp:1970:GL

[Rog70a] Klaus W. Roggenkamp. Genera of lattices. Lecture Notes in Mathematics, 142:70–104, 1970. CODEN LNMAA2. ISBN 3-540-04931-2 (print), 3-540-36301-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058957/.

Roggenkamp:1970:GG

[Rog70b] Klaus W. Roggenkamp. Grothendieck groups. Lecture Notes in Mathematics, 142:105–175, 1970. CODEN LNMAA2. ISBN 3-540-04931-2 (print), 3-540-36301-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058958/.

Roggenkamp:1970:LOI

[Rog70c] Klaus W. Roggenkamp. Lattices over Orders II, volume 142 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04931-2 (print), 3-540-36301-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 387 pp. LCCN QA3 .L28 no. 142. URL http:/

/link.springer.com/book/10.1007/BFb0058955;   http://
www.springerlink.com/content/978-3-540-36301-9.

Roggenkamp:1970:MOO

[Rog70d] Klaus W. Roggenkamp. Modules over orders, one-sided ideals over maximal orders. (German) []. Lecture Notes in Mathematics, 142:1–69, 1970. CODEN LNMAA2. ISBN 3-540-04931-2 (print), 3-540-36301-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058956/.

Roggenkamp:1970:NIL

[Rog70e] Klaus W. Roggenkamp. The number of indecomposable lattices over orders. Lecture Notes in Mathematics, 142:287–382, 1970. CODEN LNMAA2. ISBN 3-540-04931-2 (print), 3-540-36301-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058960/.

Roggenkamp:1970:STO

[Rog70f] Klaus W. Roggenkamp. Special types of orders. Lecture Notes in Mathematics, 142:176–286, 1970. CODEN LNMAA2. ISBN 3-540-04931-2 (print), 3-540-36301-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058959/.

Roitberg:1971:HCR

[Roi71] Joseph Roitberg. On the homotopy classification of rank 3-H-spaces. Lecture Notes in Mathematics, 196:34–38, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067743/.

Roitberg:1974:NGH

[Roi74] Joseph Roitberg. Nilpotent groups, homotopy types and rational Lie algebras. Lecture Notes in Mathematics, 418:132–138, 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070648/.

Roos:1974:LCF

[Roo74] Guy Roos. L’intégrale de Cauchy dans C^{n}. (French) []. Lecture Notes in Mathematics, 409:176–195, 1974. CODEN LN-MAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068099/.

Rost:1971:EES

[Ros71] Hermann Rost. Einige ergebnisse Über die stoppverteilungen eines Markoff-prozesses. (German) []. Lecture Notes in Mathematics, 190:44–48, 1971. CODEN LNMAA2. ISBN 3-540-05396-4 (print), 3-540-36515-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065891/.

Rosenberg:1972:FSD

[Ros72] Harold Rosenberg. Feuilletages sur des sphères [d’après H. Blaine Lawson]. (French) []. Lecture Notes in Mathematics, 244:221– 232, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058704/.

Roseblade:1973:PGR

[Ros73] J. E. Roseblade. Polycyclic group rings and the Nullstellensatz. Lecture Notes in Mathematics, 319:156–167, 1973. CODEN LN-MAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058941/.

Rosenberg:1974:CEC

[Ros74] Harold Rosenberg. Contre-exemple à la conjecture de Seifert. (French) []. Lecture Notes in Mathematics, 383:294–306, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057315/.

Roussarie:1974:PSI

[Rou74] R. Roussarie. Phenômes de stabilité et d’instabilité dans les feuilletages. (French) []. Lecture Notes in Mathematics, 392:53– 60, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064111/.

Roxin:1971:SGP

[Rox71] Emilio O. Roxin. Some global problems in differential games. Lecture Notes in Mathematics, 235:103–116, 1971. CODEN LN-MAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059202/.

Royden:1971:RKM

[Roy71] H. L. Royden. Remarks on the Kobayashi metric. Lecture Notes in Mathematics, 185:125–137, 1971. CODEN LNMAA2. ISBN 3-540-05372-7 (print), 3-540-36493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058768/.

Royden:1974:MTS

[Roy74] H. L. Royden. Metrics on Teichmüller space. Lecture Notes in Mathematics, 400:71–78, 1974. CODEN LNMAA2. ISBN 3-540-06840-6 (print), 3-540-37776-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065676/.

Rector:1974:LGH

[RS74] David Rector and James Stasheff. Lie groups from a homotopy point of view. Lecture Notes in Mathematics, 418:121–131, 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070647/.

Reiner:1973:CGO

[RU73] I. Reiner and S. Ullom. Class groups of orders and a Mayer– Vietoris sequence. Lecture Notes in Mathematics, 353:139–151, 1973. CODEN LNMAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059266/.

Rubel:1973:SFS

[Rub73] L. A. Rubel. A survey of a Fourier series method for meromorphic functions. Lecture Notes in Mathematics, 336:51–62, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065787/.

Rudin:1973:FFH

[Rud73] W. Rudin. Factorisation des fonctions holomorphes bornées. (French) []. Lecture Notes in Mathematics, 336:168–169, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065799.

Ruelle:1970:NMP

[Rue70] D. Ruelle. New methods and problems in statistical mechanics. Lecture Notes in Mathematics, 140:68–83, 1970. CODEN LNMAA2. ISBN 3-540-04929-0 (print), 3-540-36298-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100067/.

Ruget:1970:CAD

[Rug70] Gabriel Ruget. Cycles analytiques de dimension infinie. (French) []. Lecture Notes in Mathematics, 116:56–64, 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059245/.

Robinson:1970:ASA

[RW70] George Robinson and Lawrence Wos. Axiom systems in automatic theorem proving. Lecture Notes in Mathematics, 125:215– 236, 1970. CODEN LNMAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060635/.

Ryavec:1972:UMT

[Rya72] C. Ryavec. The use of measure theoretic methods in the study of additive arithmetic functions. Lecture Notes in Mathematics, 251:193–203, 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058794/.

Sachs:1972:IEN

[Sac72] A. Sachs. Iterationsverfahren für elliptische (nichtlineare) Differenzenoperatoren in Divergenzform. (German) [Iterative methods for elliptic (nonlinear) difference operators in divergence form]. Lecture Notes in Mathematics, 267:305–322, 1972. CO-DEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061624/.

Sacerdote:1973:SUR

[Sac73] G. S. Sacerdote. SQ-universal 1-relator groups. Lecture Notes in Mathematics, 319:168, 1973. CODEN LNMAA2. ISBN 3-540-06205-X (print), 3-540-38479-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058942.

Sachs:1974:SAL

[Sac74a] A. Sachs. Zur Struktur eines Algorithmus zur Lösung freier Randwertprobleme parabolischer Differentialoperatoren. (German) []. Lecture Notes in Mathematics, 395:223–233, 1974. CO-DEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060674/.

Sacksteder:1974:MIU

[Sac74b] Richard Sacksteder. The measures invariant under an expanding map. Lecture Notes in Mathematics, 392:179–194, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064122/.

Sadler:1974:CVH

[Sad74] Charles Sadler. Champs de vecteurs holomorphes et trajectoires sur les espaces projectifs complexes. (French) []. Lecture Notes in Mathematics, 409:223–241, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068103/.

Saito:1971:DST

[Sai71a] T. Saito. Dynamical systems on an n-torus. Lecture Notes in Mathematics, 206:18–19, 1971. CODEN LNMAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070128.

Saito:1971:GBT

[Sai71b] T. Saito. Generalisation of Bendixson’s theory. Lecture Notes in Mathematics, 206:23–25, 1971. CODEN LNMAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070131/.

Saito:1971:MPF

[Sai71c] T. Saito. On the measure-preserving flow on the torus. Lecture Notes in Mathematics, 206:1–2, 1971. CODEN LNMAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070118.

Saito:1972:GNA

[Sai72a] Teishirô Saitô. Generations of von Neumann algebras. Lecture Notes in Mathematics, 247:435–531, 1972. CODEN LN-MAA2. ISBN 3-540-05729-3 (print), 3-540-37117-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058556/.

Saito:1972:HOR

[Sai72b] Teishirô Saitô. Hyponormal operators and related topics. Lecture Notes in Mathematics, 247:533–664, 1972. CODEN LN-MAA2. ISBN 3-540-05729-3 (print), 3-540-37117-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058557/.

Sakawa:1971:PEP

[Sak71] Yoshiyuki Sakawa. On pursuit and evasion problems. Lecture Notes in Mathematics, 243:193–207, 1971. CODEN LN-MAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058728/.

Sally:1973:NGI

[Sal73] Judith Sally. On the number of generators of ideals of dimension zero. Lecture Notes in Mathematics, 311:239–242, 1973. CODEN LNMAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068932/.

Sallantin:1974:ITS

[Sal74] Jean Sallantin. Informations et trajectoires sur un système de propositions. (French) []. Lecture Notes in Mathematics, 398: 76–89, 1974. CODEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065769/.

Sanchez:1971:CMR

[Sán71] David A. Sánchez. Characteristic multipliers and the Riccati equation. Lecture Notes in Mathematics, 183:42–44, 1971. CO-DEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060410/.

SanRoman:1974:GAC

[San74] J. Sancho San Román. Sur des groupes associés à certaines varieties tridimensionelles. (French) []. Lecture Notes in Mathematics, 392:175–178, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0064121/.

Satake:1971:CSA

[Sat71] Ichirô Satake. ¯∂-cohomology spaces associated with a certain group extension. Lecture Notes in Mathematics, 185:138–168, 1971. CODEN LNMAA2. ISBN 3-540-05372-7 (print), 3-540-36493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058769/.

Sattinger:1973:SLT

[Sat73a] D. H. Sattinger. Six lectures on the transition to instability. Lecture Notes in Mathematics, 322:261–287, 1973. CODEN LN-MAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060571/.

Sattinger:1973:BSE

[Sat73b] David H. Sattinger. Bifurcation at a simple eigenvalue. Lecture Notes in Mathematics, 309:77–102, 1973. CODEN LN-MAA2. ISBN 3-540-06133-9 (print), 3-540-38333-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060083/.

Sattinger:1973:BPS

[Sat73c] David H. Sattinger. Bifurcation of periodic solutions. Lecture Notes in Mathematics, 309:103–124, 1973. CODEN LN-MAA2. ISBN 3-540-06133-9 (print), 3-540-38333-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060084/.

Sattinger:1973:FA

[Sat73d] David H. Sattinger. Functional analysis. Lecture Notes in Mathematics, 309:48–76, 1973. CODEN LNMAA2. ISBN 3-540-06133-9 (print), 3-540-38333-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060082/.

Sattinger:1973:I

[Sat73e] David H. Sattinger. Introduction. Lecture Notes in Mathematics, 309:1–13, 1973. CODEN LNMAA2. ISBN 3-540-06133-9 (print), 3-540-38333-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060080/.

Sattinger:1973:MPH

[Sat73f] David H. Sattinger. The mathematical problems of hydrodynamic stability. Lecture Notes in Mathematics, 309:125–140, 1973. CODEN LNMAA2. ISBN 3-540-06133-9 (print), 3-540-38333-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060085/.

Sattinger:1973:NEB

[Sat73g] David H. Sattinger. Nonlinear elliptic boundary value problems of second order. Lecture Notes in Mathematics, 309:14–47, 1973. CODEN LNMAA2. ISBN 3-540-06133-9 (print), 3-540-38333-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060081/.

Sattinger:1973:RW

[Sat73h] David H. Sattinger. The real world. Lecture Notes in Mathematics, 309:173–184, 1973. CODEN LNMAA2. ISBN 3-540-06133-9 (print), 3-540-38333-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060087/.

Sattinger:1973:TSB

[Sat73i] David H. Sattinger. Topics in Stability and Bifurcation Theory, volume 309 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06133-9 (print), 3-540-38333-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 190 pp. LCCN QA3 .L28 no. 309. URL http:/

/link.springer.com/book/10.1007/BFb0060079;   http://
www.springerlink.com/content/978-3-540-38333-8.

Sattinger:1973:TDT

[Sat73j] David H. Sattinger. Topological degree theory and applications. Lecture Notes in Mathematics, 309:141–172, 1973. CODEN LN-MAA2. ISBN 3-540-06133-9 (print), 3-540-38333-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060086/.

Sibuya:1973:ESF

[SC73] Yasutaka Sibuya and Robert H. Cameron. An entire solution of the functional equation f(λ) + f(ωλ)f(ω^{−1}λ) = 1, (ω^{5} = 1). Lecture Notes in Mathematics, 312:194–202, 1973. CODEN LN-MAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060056/.

Scarpellini:1971:FIS

[Sca71a] Bruno Scarpellini. A formally intuitionistic system as strong as classical analysis. Lecture Notes in Mathematics, 212:79– 140, 1971. CODEN LNMAA2. ISBN 3-540-05541-X (print), 3-540-36875-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068787/.

Scarpellini:1971:FIT

[Sca71b] Bruno Scarpellini. A formally intuitonistic theory equivalent to classical transfinite induction with respect to recursive wellfounded trees with function parameters. Lecture Notes in Mathematics, 212:161–192, 1971. CODEN LNMAA2. ISBN 3-540-05541-X (print), 3-540-36875-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068789/.

Scarpellini:1971:HF

[Sca71c] Bruno Scarpellini. Harrop formulas. Lecture Notes in Mathematics, 212:217–247, 1971. CODEN LNMAA2. ISBN 3-540-05541-X (print), 3-540-36875-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068791/.

Scarpellini:1971:IP

[Sca71d] Bruno Scarpellini. Introduction and preliminaries. Lecture Notes in Mathematics, 212:1–33, 1971. CODEN LNMAA2. ISBN 3-540-05541-X (print), 3-540-36875-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068784/.

Scarpellini:1971:ISN

[Sca71e] Bruno Scarpellini. The intuitionistic system of number theory. Lecture Notes in Mathematics, 212:70–78, 1971. CODEN LN-MAA2. ISBN 3-540-05541-X (print), 3-540-36875-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068786/.

Scarpellini:1971:MP

[Sca71f] Bruno Scarpellini. The Markov principle. Lecture Notes in Mathematics, 212:248–255, 1971. CODEN LNMAA2. ISBN 3-540-05541-X (print), 3-540-36875-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068792/.

Scarpellini:1971:PTI

[Sca71g] Bruno Scarpellini. Proof Theory and Intuitionistic Systems, volume 212 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05541-X (print), 3-540-36875-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 291 pp. LCCN QA3 .L28 no. 212. URL http:/

/link.springer.com/book/10.1007/BFb0068783;   http://
www.springerlink.com/content/978-3-540-36875-5.

Scarpellini:1971:RCP

[Sca71h] Bruno Scarpellini. Relative consistency proof of ZTN with respect to ZTi / I_{N}∗. Lecture Notes in Mathematics, 212:256–290, 1971. CODEN LNMAA2. ISBN 3-540-05541-X (print), 3-540-36875-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068793/.

Scarpellini:1971:RGS

[Sca71i] Bruno Scarpellini. A review of Gentzen’s second consistency proof. Lecture Notes in Mathematics, 212:34–69, 1971. CODEN LNMAA2. ISBN 3-540-05541-X (print), 3-540-36875-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068785/.

Scarpellini:1971:SCB

[Sca71j] Bruno Scarpellini. A system containing barinduction with respect to decidable predicates. Lecture Notes in Mathematics, 212:193–216, 1971. CODEN LNMAA2. ISBN 3-540-05541-X (print), 3-540-36875-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068790/.

Scarpellini:1971:TIR

[Sca71k] Bruno Scarpellini. Transfinite induction with respect to recursive wellorderings without function parameters. Lecture Notes in Mathematics, 212:141–160, 1971. CODEN LNMAA2. ISBN 3-540-05541-X (print), 3-540-36875-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068788/.

Schapira:1970:CFF

[Sch70a] Pierre Schapira. Cohomologie des faisceaux. (French) [Sheaf cohomology]. Lecture Notes in Mathematics, 126:8–43, 1970. CO-DEN LNMAA2. ISBN 3-540-04915-0 (print), 3-540-36263-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059486/.

Schapira:1970:HFH

[Sch70b] Pierre Schapira. Hyperfonctions. (French) [Hyperfunctions]. Lecture Notes in Mathematics, 126:44–76, 1970. CODEN LN-MAA2. ISBN 3-540-04915-0 (print), 3-540-36263-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059487/.

Schapira:1970:OEF

[Sch70c] Pierre Schapira. Opérateurs elliptiques. (French) []. Lecture Notes in Mathematics, 126:77–97, 1970. CODEN LNMAA2. ISBN 3-540-04915-0 (print), 3-540-36263-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059488/.

Schapira:1970:RF

[Sch70d] Pierre Schapira. Rappels. (French) []. Lecture Notes in Mathematics, 126:1–7, 1970. CODEN LNMAA2. ISBN 3-540-04915-0 (print), 3-540-36263-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059485/.

Schapira:1970:RDF

[Sch70e] Pierre Schapira. Résultats divers. (French) []. Lecture Notes in Mathematics, 126:98–117, 1970. CODEN LNMAA2. ISBN 3-540-04915-0 (print), 3-540-36263-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059489/.

Schapira:1970:THF

[Sch70f] Pierre Schapira. Théorie des Hyperfonctions. (French) [Theory of hyperfunctions], volume 126 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04915-0 (print), 3-540-36263-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 152 (est.) pp. LCCN QA3 .L28 no. 126.

URL http://link.springer.com/book/10.1007/BFb0059484;
 http://www.springerlink.com/content/978-3-540-36263-
0.

Schapira:1970:VAB

[Sch70g] Pierre Schapira. Valeurs au bord des fonctions holomorphes. (French) []. Lecture Notes in Mathematics, 126:118–152, 1970. CODEN LNMAA2. ISBN 3-540-04915-0 (print), 3-540-36263-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059490/.

Schweitzer:1970:JCI

[Sch70h] Paul A. Schweitzer. Joint Cobordism of immersions. Lecture Notes in Mathematics, 168:267–282, 1970. CODEN LN-MAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058528/.

Schapira:1971:UHT

[Sch71a] P. Schapira. Utilisation des hyperfonctions dans les théorèmes de dualité de la géométrie analytique. (French) []. Lecture Notes in Mathematics, 205:166–182, 1971. CODEN LNMAA2. ISBN 3-540-05505-3 (print), 3-540-36867-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059838/.

Schnorr:1971:CZDc

[Sch71b] C. P. Schnorr. Charakterisierung der Zufallsfolgen durch Invarianzeigenschaften. (German) []. Lecture Notes in Mathematics, 218:89–91, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112471/.

Schnorr:1971:CZDa

[Sch71c] C. P. Schnorr. Charakterisierung der Zufallsfolgen durch konstruktive Nullmengen nach L. E. J. Brouwer. (German) []. Lecture Notes in Mathematics, 218:60–69, 1971. CODEN LN-MAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112467/.

Schnorr:1971:CHF

[Sch71d] C. P. Schnorr. Charakterisierung hyperzufälliger Folgen durch Invarianzeigenschaften. (German) []. Lecture Notes in Mathematics, 218:45–51, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112465/.

Schnorr:1971:CZDb

[Sch71e] C. P. Schnorr. Charakterisierung von Zufallsfolgen durch das Prinzip vom ausgeschlossenen Spielsystem. (German) []. Lecture Notes in Mathematics, 218:70–77, 1971. CODEN LN-MAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112468/.

Schnorr:1971:DSG

[Sch71f] C. P. Schnorr. Darstellung des starken Gesetzes der großen Zahlen durch Martingale. (German) []. Lecture Notes in Mathematics, 218:78–82, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112469/.

Schnorr:1971:NBK

[Sch71g] C. P. Schnorr. Der naive Begriff des Kollektivs nach von Mises. (German) []. Lecture Notes in Mathematics, 218:10–20, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112461/.

Schnorr:1971:OZG

[Sch71h] C. P. Schnorr. Die Ordnung eines Zufallsgesetzes. (German) []. Lecture Notes in Mathematics, 218:120–128, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112475/.

Schnorr:1971:PNK

[Sch71i] C. P. Schnorr. Die Programmkomplexität nach Kolmogoroff. (German) []. Lecture Notes in Mathematics, 218:107–119, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112474/.

Schnorr:1971:DEA

[Sch71j] C. P. Schnorr. Durch endliche Automaten darstellbare Zufallsgesetze. (German) [From finite automata to representable laws of chance]. Lecture Notes in Mathematics, 218:146–149, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112478/.

Schnorr:1971:EMS

[Sch71k] C. P. Schnorr. Einige modifizierte Spielsysteme. (German) []. Lecture Notes in Mathematics, 218:92–97, 1971. CODEN LN-MAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112472/.

Schnorr:1971:EAW

[Sch71l] C. P. Schnorr. Erste Ansätze zur widerspruchsfreien Definition der Kollektive und ihre Kritik durch Ville. (German) []. Lecture Notes in Mathematics, 218:21–28, 1971. CODEN LN-MAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112462/.

Schnorr:1971:HFG

[Sch71m] C. P. Schnorr. Hyperzufällige Folgen. (German) []. Lecture Notes in Mathematics, 218:29–37, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0112463/.

Schnorr:1971:HFP

[Sch71n] C. P. Schnorr. Hyperzufällige Folgen und das Prinzip vom ausgeschlossenen Spielsystem. (German) []. Lecture Notes in Mathematics, 218:38–44, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0112464/.

Schnorr:1971:IZG

[Sch71o] C. P. Schnorr. Invarianzeigenschaften von Zufallsfolgen. (German) []. Lecture Notes in Mathematics, 218:83–88, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112470/.

Schnorr:1971:KMW

[Sch71p] C. P. Schnorr. Kritik der Maß-Wahrscheinlichkeitstheorie. (German) []. Lecture Notes in Mathematics, 218:4–9, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0112460/.

Schnorr:1971:VQR

[Sch71q] C. P. Schnorr. Voraussagbare und quasi-rekursive Folgen. (German) []. Lecture Notes in Mathematics, 218:140–145, 1971. CO-DEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112477/.

Schnorr:1971:VEG

[Sch71r] C. P. Schnorr. Vorwort und Einleitung. (German) []. Lecture Notes in Mathematics, 218:1–3, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0112459/.

Schnorr:1971:WEG

[Sch71s] C. P. Schnorr. Weitere Einwände gegen den Begriff der Zufallsfolge im Sinne von Martin-Löf. (German) []. Lecture Notes in Mathematics, 218:52–59, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0112466/.

Schnorr:1971:ZOF

[Sch71t] C. P. Schnorr. Zufallsfolgen als optimale Folgen für die Bank. (German) []. Lecture Notes in Mathematics, 218:98–106, 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112473/.

Schnorr:1971:ZEO

[Sch71u] C. P. Schnorr. Zufallsgesetze von exponentieller Ordnung. (German) []. Lecture Notes in Mathematics, 218:129–139, 1971. CO-DEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0112476/.

Schnorr:1971:ZWA

[Sch71v] Claus-Peter Schnorr. Zufälligkeit und Wahrscheinlichkeit: Eine algorithmische Begründung der Wahrscheinlichkeitstheorie. (German) [Randomness and probability. An algorithmic foundation of probability theory], volume 218 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05566-5 (print), 3-540-36883-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 212 pp. LCCN QA3 .L28 no. 218; QA3 .L28 no. 218 QA273.4. URL http:/

/link.springer.com/book/10.1007/BFb0112458;   http://
www.springerlink.com/content/978-3-540-36883-0.

Schreiber:1971:NPT

[Sch71w] Jean-Pierre Schreiber. Nombres de Pisot et travaux d’Yves Meyer. (French) []. Lecture Notes in Mathematics, 180:269– 279, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058836/.

Schubert:1971:OS

[Sch71x] C. F. Schubert. Overdetermined systems on L^{2}(R_{n}^{+}). Lecture Notes in Mathematics, 183:221–225, 1971. CODEN LN-MAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060423/.

Schiffmann:1972:TEC

[Sch72a] Gérard Schiffmann. Théorie Élémentaire des catégories. (French) []. Lecture Notes in Mathematics, 277:1–33, 1972. CODEN LNMAA2. ISBN 3-540-05934-2 (print), 3-540-37613-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058877/.

Schiffmann:1972:ADT

[Sch72b] Gérard Schiffmann. Un analogue du théorème de Borel–Weil– Bott dans le cas non compact. (French) []. Lecture Notes in Mathematics, 244:323–336, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058709/.

Schroder:1972:EDG

[Sch72c] Johann Schröder. Einschliessungsaussagen bei Differentialgleichungen. (German) []. Lecture Notes in Mathematics, 267:23– 49, 1972. CODEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061612/.

Schultz:1972:TAH

[Sch72d] Reinhard Schultz. Z_{2}-torus actions on homotopy spheres. Lecture Notes in Mathematics, 298:117–118, 1972. CODEN LN-MAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070035.

Schultz:1972:SCA

[Sch72e] Reinhard Schultz. Semifree circle actions with twisted fixed point sets. Lecture Notes in Mathematics, 298:102–116, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070034/.

Schwartz:1972:PTA

[Sch72f] Laurent Schwartz. Produits tensoriels g_{p} et d_{p}, applications p-sommantes, applications p-radonifiantes. (French) []. Lecture Notes in Mathematics, 244:63–88, 1972. CODEN LN-MAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058697/.

Schober:1973:EFE

[Sch73a] Glenn Schober. Estimates for Fredholm eigenvalues based on quasiconformal mapping. Lecture Notes in Mathematics, 333: 211–217, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060699/.

Schutz:1973:CP

[Sch73b] John W. Schutz. Collinear particles. Lecture Notes in Mathematics, 361:103–146, 1973. CODEN LNMAA2. ISBN 3-540-06591-1 (print), 3-540-37906-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066802/.

Schutz:1973:CSS

[Sch73c] John W. Schutz. Collinear sub-SPRAYS after coincidence. Lecture Notes in Mathematics, 361:71–102, 1973. CODEN LN-MAA2. ISBN 3-540-06591-1 (print), 3-540-37906-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066801/.

Schutz:1973:CR

[Sch73d] John W. Schutz. Concluding remarks. Lecture Notes in Mathematics, 361:300–301, 1973. CODEN LNMAA2. ISBN 3-540-06591-1 (print), 3-540-37906-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066806.

Schutz:1973:CCP

[Sch73e] John W. Schutz. Conditionally complete particles. Lecture Notes in Mathematics, 361:42–58, 1973. CODEN LNMAA2. ISBN 3-540-06591-1 (print), 3-540-37906-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066799/.

Schutz:1973:FSR

[Sch73f] John W. Schutz. Foundations of Special Relativity: Kinematic Axioms for Minkowski Space–Time, volume 361 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06591-1 (print), 3-540-37906-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xx + 314 pp. LCCN QA3 .L28 no. 361. URL http:/

/link.springer.com/book/10.1007/BFb0066796;   http://
www.springerlink.com/content/978-3-540-37906-5.

Schutz:1973:IC

[Sch73g] John W. Schutz. Implications of collinearity. Lecture Notes in Mathematics, 361:59–70, 1973. CODEN LNMAA2. ISBN 3-540-06591-1 (print), 3-540-37906-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066800/.

Schutz:1973:I

[Sch73h] John W. Schutz. Introduction. Lecture Notes in Mathematics, 361:1–6, 1973. CODEN LNMAA2. ISBN 3-540-06591-1 (print), 3-540-37906-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066797/.

Schutz:1973:KAM

[Sch73i] John W. Schutz. Kinematic axioms for Minkowski space-time. Lecture Notes in Mathematics, 361:7–41, 1973. CODEN LN-MAA2. ISBN 3-540-06591-1 (print), 3-540-37906-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066798/.

Schutz:1973:ODK

[Sch73j] John W. Schutz. One-dimensional kinematics. Lecture Notes in Mathematics, 361:233–249, 1973. CODEN LNMAA2. ISBN 3-540-06591-1 (print), 3-540-37906-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066804/.

Schutz:1973:TP

[Sch73k] John W. Schutz. Theory of parallels. Lecture Notes in Mathematics, 361:147–232, 1973. CODEN LNMAA2. ISBN 3-540-06591-1 (print), 3-540-37906-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066803/.

Schutz:1973:TDK

[Sch73l] John W. Schutz. Three-dimensional kinematics. Lecture Notes in Mathematics, 361:250–299, 1973. CODEN LNMAA2. ISBN 3-540-06591-1 (print), 3-540-37906-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066805/.

Schweiger:1973:BDa

[Sch73m] Fritz Schweiger. Basic definitions. Lecture Notes in Mathematics, 334:1–8, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059846/.

Schweiger:1973:BDb

[Sch73n] Fritz Schweiger. Billingsley dimension. Lecture Notes in Mathematics, 334:65–68, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059860/.

Schweiger:1973:BCL

[Sch73o] Fritz Schweiger. The Borel–Cantelli lemma of Schmidt–Philipp. Lecture Notes in Mathematics, 334:49–54, 1973. CODEN LN-MAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059855/.

Schweiger:1973:CT

[Sch73p] Fritz Schweiger. Comparison theorems. Lecture Notes in Mathematics, 334:68–75, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059861/.

Schweiger:1973:CE

[Sch73q] Fritz Schweiger. Conditional expectations. Lecture Notes in Mathematics, 334:17–21, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059849/.

Schweiger:1973:CR

[Sch73r] Fritz Schweiger. Convergence results. Lecture Notes in Mathematics, 334:42–49, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059854/.

Schweiger:1973:C

[Sch73s] Fritz Schweiger. Cylinders. Lecture Notes in Mathematics, 334: 8–15, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059847/.

Schweiger:1973:EIM

[Sch73t] Fritz Schweiger. Ergodic invariant measures. Lecture Notes in Mathematics, 334:90–94, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059863/.

Schweiger:1973:ETb

[Sch73u] Fritz Schweiger. The ergodic theorem. Lecture Notes in Mathematics, 334:28–34, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059852/.

Schweiger:1973:ETa

[Sch73v] Fritz Schweiger. Ergodicity of the transformation. Lecture Notes in Mathematics, 334:22–23, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059850.

Schweiger:1973:EEI

[Sch73w] Fritz Schweiger. Existence of an equivalent invariant measure. Lecture Notes in Mathematics, 334:23–28, 1973. CODEN LN-MAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059851/.

Schweiger:1973:HD

[Sch73x] Fritz Schweiger. Hausdorff dimension. Lecture Notes in Mathematics, 334:63–65, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059859/.

Schweiger:1973:HM

[Sch73y] Fritz Schweiger. Hausdorff measures. Lecture Notes in Mathematics, 334:60–63, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059858/.

Schweiger:1973:IF

[Sch73z] Fritz Schweiger. Increasing σ-fields. Lecture Notes in Mathematics, 334:15–17, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059848/.

Schweiger:1973:KT

[Sch73-27] Fritz Schweiger. Kuzmin’s theorem. Lecture Notes in Mathematics, 334:34–42, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059853/.

Schweiger:1973:MTD

[Sch73-28] Fritz Schweiger. The main theorem of dimension theory of Jacobi algorithm. Lecture Notes in Mathematics, 334:75–90, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059862/.

Schweiger:1973:MTJ

[Sch73-29] Fritz Schweiger. The Metrical Theory of Jacobi–Perron Algorithm, volume 334 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 111 pp. LCCN QA3 .L28 no. 334. URL http:/

/link.springer.com/book/10.1007/BFb0059845;   http://
www.springerlink.com/content/978-3-540-47010-6.

Schweiger:1973:OM

[Sch73-30] Fritz Schweiger. Outer measures. Lecture Notes in Mathematics, 334:56–60, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059857/.

Schweiger:1973:PC

[Sch73-31] Fritz Schweiger. Proof of the conjecture for n = 1 and n = 2. Lecture Notes in Mathematics, 334:99–104, 1973. CODEN LNMAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059865/.

Schweiger:1973:SEK

[Sch73-32] Fritz Schweiger. Some extensions of Kuzmin’s theorem. Lecture Notes in Mathematics, 334:54–56, 1973. CODEN LN-MAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059856/.

Schweiger:1973:VAM

[Sch73-33] Fritz Schweiger. Volume as approximation measure. Lecture Notes in Mathematics, 334:94–98, 1973. CODEN LN-MAA2. ISBN 3-540-06388-9 (print), 3-540-47010-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059864/.

Schaback:1974:KMS

[Sch74a] Robert Schaback. Kollokation mit mehrdimensionalen Spline-Funktionen. (German) [Collocation with multi-dimensional spline functions]. Lecture Notes in Mathematics, 395:291–300, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060679/.

Schapira:1974:FMS

[Sch74b] Pierre Schapira. Le faisceau C de M. Sato. (French) []. Lecture Notes in Mathematics, 409:196–203, 1974. CODEN LN-MAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068100/.

Schmid:1974:AAQ

[Sch74c] Wilfried Schmid. Abbildungen in arithmetische Quotienten hermitesch symmetrischer Räume. (German) []. Lecture Notes in Mathematics, 412:243–258, 1974. CODEN LNMAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066162/.

Schottenloher:1974:BDB

[Sch74d] Martin Schottenloher. Biemann domains: Basic results and open problems. Lecture Notes in Mathematics, 364:196–212, 1974. CO-DEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069019/.

Schottenloher:1974:BSB

[Sch74e] Martin Schottenloher. Bounding sets in Banach spaces and regular classes of analytic functions. Lecture Notes in Mathematics, 384:109–122, 1974. CODEN LNMAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057542/.

Schwenk:1974:CCP

[Sch74f] Allen J. Schwenk. Computing the characteristic polynomial of a graph. Lecture Notes in Mathematics, 406:153–172, 1974. CO-DEN LNMAA2. ISBN 3-540-06854-6 (print), 3-540-37809-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066438/.

Scott:1970:CV

[Sco70] Dana Scott. Constructive validity. Lecture Notes in Mathematics, 125:237–275, 1970. CODEN LNMAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060636/.

Scott:1971:LFD

[Sco71] Dana Scott. The lattice of flow diagrams. Lecture Notes in Mathematics, 188:311–366, 1971. CODEN LNMAA2. ISBN 3-540-05377-8 (print), 3-540-36499-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059703/.

Scott:1972:CL

[Sco72] Dana Scott. Continuous lattices. Lecture Notes in Mathematics, 274:97–136, 1972. CODEN LNMAA2. ISBN 3-540-05920-2 (print), 3-540-37609-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0073967/.

Saint-Donat:1972:TDC

[SD72] B. Saint-Donat. Techniques de descente cohomologique. (French) []. Lecture Notes in Mathematics, 270:83–162, 1972. CODEN LNMAA2. ISBN 3-540-06012-X (print), 3-540-37987-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061321/.

Swinnerton-Dyer:1973:ARC

[SD73] H. P. F. Swinnerton-Dyer. On ℓ-adic representations and congruences for coefficients of modular forms. Lecture Notes in Mathematics, 350:1–55, 1973. CODEN LNMAA2. ISBN 3-540-06483-4 (print), 3-540-37802-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-37802-0_1/;   http://link.springer.
com/content/pdf/bfm:978-3-540-37802-0/1.pdf.

Swan:1970:MOb

[SE70a] Richard G. Swan and E. Graham Evans. K_{0} of a maximal order. Lecture Notes in Mathematics, 149:126–140, 1970. CODEN LNMAA2. ISBN 3-540-04938-X (print), 3-540-36312-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059157/.

Swan:1970:KG

[SE70b] Richard G. Swan and E. Graham Evans. K_{1} and G_{1}. Lecture Notes in Mathematics, 149:140–167, 1970. CODEN LN-MAA2. ISBN 3-540-04938-X (print), 3-540-36312-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059158/.

Swan:1970:CT

[SE70c] Richard G. Swan and E. Graham Evans. Cancellation theorems. Lecture Notes in Mathematics, 149:167–204, 1970. CODEN LN-MAA2. ISBN 3-540-04938-X (print), 3-540-36312-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059159/.

Swan:1970:FT

[SE70d] Richard G. Swan and E. Graham Evans. Finiteness theorems. Lecture Notes in Mathematics, 149:37–54, 1970. CODEN LN-MAA2. ISBN 3-540-04938-X (print), 3-540-36312-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059153/.

Swan:1970:FF

[SE70e] Richard G. Swan and E. Graham Evans. Frobenius functors. Lecture Notes in Mathematics, 149:13–36, 1970. CODEN LN-MAA2. ISBN 3-540-04938-X (print), 3-540-36312-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059152/.

Swan:1970:I

[SE70f] Richard G. Swan and E. Graham Evans. Introduction. Lecture Notes in Mathematics, 149:1–13, 1970. CODEN LNMAA2. ISBN 3-540-04938-X (print), 3-540-36312-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059151/.

Swan:1970:MOa

[SE70g] Richard G. Swan and E. Graham Evans. Maximal orders. Lecture Notes in Mathematics, 149:83–103, 1970. CODEN LN-MAA2. ISBN 3-540-04938-X (print), 3-540-36312-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059155/.

Swan:1970:O

[SE70h] Richard G. Swan and E. Graham Evans. Orders. Lecture Notes in Mathematics, 149:104–126, 1970. CODEN LNMAA2. ISBN 3-540-04938-X (print), 3-540-36312-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059156/.

Swan:1970:R

[SE70i] Richard G. Swan and E. Graham Evans. Results on K_{0} and G_{0}. Lecture Notes in Mathematics, 149:54–83, 1970. CODEN LNMAA2. ISBN 3-540-04938-X (print), 3-540-36312-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059154/.

Segal:1970:MTQ

[Seg70] Irving E. Segal. The mathematical theory of quantum fields. Lecture Notes in Mathematics, 140:30–57, 1970. CODEN LN-MAA2. ISBN 3-540-04929-0 (print), 3-540-36298-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0100065/.

Sehgal:1973:LPM

[Seh73] S. K. Sehgal. Lie properties in modular group algebras. Lecture Notes in Mathematics, 353:152–160, 1973. CODEN LN-MAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059267/.

Seifert:1971:APS

[Sei71] George Seifert. Almost periodic solutions by the method of averaging. Lecture Notes in Mathematics, 243:123–133, 1971. CO-DEN LNMAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058723/.

Seip:1972:DGD

[Sei72a] Ulrich Seip. Differenzierbarkeit. (German) [Differentiability]. Lecture Notes in Mathematics, 273:49–57, 1972. CODEN LN-MAA2. ISBN 3-540-05909-1 (print), 3-540-37608-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058381/.

Seip:1972:FAG

[Sei72b] Ulrich Seip. Fixpunktsätze und Anwendungen bei gewöhnlichen Differentialgleichungen. (German) []. Lecture Notes in Mathematics, 273:107–114, 1972. CODEN LNMAA2. ISBN 3-540-05909-1 (print), 3-540-37608-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058386/.

Seip:1972:IVD

[Sei72c] Ulrich Seip. Integration vs differentiation. Lecture Notes in Mathematics, 273:95–106, 1972. CODEN LNMAA2. ISBN 3-540-05909-1 (print), 3-540-37608-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058385/.

Seip:1972:KKEa

[Sei72d] Ulrich Seip. Kategorien kompakt erzeugter Räume. (German) []. Lecture Notes in Mathematics, 273:11–22, 1972. CODEN LNMAA2. ISBN 3-540-05909-1 (print), 3-540-37608-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058379/.

Seip:1972:KKEb

[Sei72e] Ulrich Seip. Kategorien kompakt erzeugter reeller Vektorräume. (German) []. Lecture Notes in Mathematics, 273:23–48, 1972. CODEN LNMAA2. ISBN 3-540-05909-1 (print), 3-540-37608-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058380/.

Seip:1972:KVG

[Sei72f] Ulrich Seip. Kategorische Vorbemerkungen. (German) []. Lecture Notes in Mathematics, 273:1–10, 1972. CODEN LNMAA2. ISBN 3-540-05909-1 (print), 3-540-37608-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058378/.

Seip:1972:KEV

[Sei72g] Ulrich Seip. Kompakt erzeugte Vektorräume und Analysis, volume 273 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05909-1 (print), 3-540-37608-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 119 pp. LCCN QA3 .L28 no. 273. URL http:

//link.springer.com/book/10.1007/BFb0058377;   http:/
/www.springerlink.com/content/978-3-540-37608-8.

Seip:1972:MSD

[Sei72h] Ulrich Seip. Mehrfache stetige Differenzierbarkeit. (German) [Multiple continuous differentiability]. Lecture Notes in Mathematics, 273:75–83, 1972. CODEN LNMAA2. ISBN 3-540-05909-1 (print), 3-540-37608-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058383/.

Seip:1972:SDG

[Sei72i] Ulrich Seip. Stetige Differenzierbarkeit. (German) [Continuous differentiability]. Lecture Notes in Mathematics, 273:59–74, 1972. CODEN LNMAA2. ISBN 3-540-05909-1 (print), 3-540-37608-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058382/.

Seip:1972:UDG

[Sei72j] Ulrich Seip. Unendliche Differenzierbarkeit. (German) [Infinite differentiability]. Lecture Notes in Mathematics, 273:85–93, 1972. CODEN LNMAA2. ISBN 3-540-05909-1 (print), 3-540-37608-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058384/.

Sell:1973:DDI

[Sel73] George R. Sell. Dichotomies for differential and integral equations. Lecture Notes in Mathematics, 312:188–193, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060055/.

Serre:1971:TCE

[Ser71a] Jean-Pierre Serre. p-torsion des courbes elliptiques. (French) []. Lecture Notes in Mathematics, 180:281–294, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058837/.

Serre:1971:TBF

[Ser71b] Jean-Pierre Serre. Travaux de Baker. (French) []. Lecture Notes in Mathematics, 180:73–86, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058825/.

Serre:1972:CGD

[Ser72] Jean-Pierre Serre. Cohomologie des groupes discrets. (French) []. Lecture Notes in Mathematics, 244:337–350, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058710/.

Serre:1973:BM

[Ser73a] Jean-Pierre Serre. Back matter. Lecture Notes in Mathematics, 5:207–217, 1973. CODEN LNMAA2. ISBN 3-540-06084-7 (print), 3-662-21553-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-662-21553-1/1.pdf.

Serre:1973:CGP

[Ser73b] Jean-Pierre Serre. Cohomologie des groupes profinis. (French) []. Lecture Notes in Mathematics, 5:1–86, 1973. CODEN LNMAA2. ISBN 3-540-06084-7 (print), 3-662-21553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-662-21553-1_1/.

Serre:1973:CGCa

[Ser73c] Jean-Pierre Serre. Cohomologie galoisienne — cas commutatif. Lecture Notes in Mathematics, 5:87–136, 1973. CODEN LN-MAA2. ISBN 3-540-06084-7 (print), 3-662-21553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-662-21553-1_

2/. See erratum [Ser73h].

Serre:1973:CGCb

[Ser73d] Jean-Pierre Serre, editor. Cohomologie Galoisienne: Cours au Collège de France, 1962–1963, volume 5 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06084-7 (print), 3-662-21553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 5.

URL http://link.springer.com/book/10.1007/978-3-662-
21553-1;   http://www.springerlink.com/content/978-3-
662-21553-1.

Serre:1973:CGN

[Ser73e] Jean-Pierre Serre. Cohomologie galoisienne non commutative. (French) [Noncommutative Galois cohomology]. Lecture Notes in Mathematics, 5:137–181, 1973. CODEN LNMAA2. ISBN 3-540-06084-7 (print), 3-662-21553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-21553-1_3/. See erratum

[Ser73i].

Serre:1973:CFM

[Ser73f] Jean-Pierre Serre. Congruences et formes modulaires. (French) []. Lecture Notes in Mathematics, 317:319–338, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069289/.

Serre:1973:E

[Ser73g] Jean-Pierre Serre. Erratum. Lecture Notes in Mathematics, 5: 213, 1973. CODEN LNMAA2. ISBN 3-540-06084-7 (print), 3-662-21553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21553-1_8.

Serre:1973:ECGa

[Ser73h] Jean-Pierre Serre. Erratum to: Cohomologie Galoisienne — Cas Commutatif. Lecture Notes in Mathematics, 5:213, 1973. CODEN LNMAA2. ISBN 3-540-06084-7 (print), 3-662-21553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21553-1_5. See [Ser73c].

Serre:1973:ECGb

[Ser73i] Jean-Pierre Serre. Erratum to: Cohomologie Galoisienne non Commutative. Lecture Notes in Mathematics, 5:213, 1973. CODEN LNMAA2. ISBN 3-540-06084-7 (print), 3-662-21553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21553-1_6. See [Ser73e].

Serre:1973:EDC

[Ser73j] Jean-Pierre Serre. Erratum to: Dualité dans la cohomologie des groupes profinis. (French) []. Lecture Notes in Mathematics, 5: 213, 1973. CODEN LNMAA2. ISBN 3-540-06084-7 (print), 3-662-21553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21553-1_7. See [Ver73].

Serre:1973:FMF

[Ser73k] Jean-Pierre Serre. Formes modulaires et fonctions zêta padiques. (French) []. Lecture Notes in Mathematics, 350:191– 268, 1973. CODEN LNMAA2. ISBN 3-540-06483-4 (print), 3-540-37802-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-37802-0_4/.

Serre:1973:FM

[Ser73l] Jean-Pierre Serre. Front matter. Lecture Notes in Mathematics, 5:n2–vii, 1973. CODEN LNMAA2. ISBN 3-540-06084-7 (print), 3-662-21553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-662-21553-1/1.pdf.

Sethna:1972:UBC

[Set72] P. R. Sethna. Ultimate behaviour of a class of stochastic differential systems dependent on a parameter. Lecture Notes in Mathematics, 294:273–282, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064947/.

Shapiro:1971:AHB

[Sha71a] Harold S. Shapiro. Applications of the Hahn–Banach theorem and dual extremal problems. Lecture Notes in Mathematics, 187: 68–80, 1971. CODEN LNMAA2. ISBN 3-540-05376-X (print), 3-540-36497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058981/.

Shapiro:1971:ATE

[Sha71b] Harold S. Shapiro. Approximation theory and extremal problems in Hilbert spaces. Lecture Notes in Mathematics, 187:81– 116, 1971. CODEN LNMAA2. ISBN 3-540-05376-X (print), 3-540-36497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058982/.

Shapiro:1971:ATH

[Sha71c] Harold S. Shapiro. Approximation theory in homogeneous Banach spaces. Lecture Notes in Mathematics, 187:200–233, 1971. CODEN LNMAA2. ISBN 3-540-05376-X (print), 3-540-36497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058985/.

Shapiro:1971:BAE

[Sha71d] Harold S. Shapiro. Best approximation and extremal problems in other norms. Lecture Notes in Mathematics, 187:52–67, 1971. CODEN LNMAA2. ISBN 3-540-05376-X (print), 3-540-36497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058980/.

Shapiro:1971:BUA

[Sha71e] Harold S. Shapiro. Best uniform approximation. Lecture Notes in Mathematics, 187:5–37, 1971. CODEN LNMAA2. ISBN 3-540-05376-X (print), 3-540-36497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058978/.

Shapiro:1971:GAD

[Sha71f] Harold S. Shapiro. General aspects of ‘degree of approximation’. Lecture Notes in Mathematics, 187:165–199, 1971. CODEN LN-MAA2. ISBN 3-540-05376-X (print), 3-540-36497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058984/.

Shapiro:1971:IFG

[Sha71g] Harold S. Shapiro. The interpolation formula and Gaussian quadrature. Lecture Notes in Mathematics, 187:38–51, 1971. CO-DEN LNMAA2. ISBN 3-540-05376-X (print), 3-540-36497-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058979/.

Shapiro:1971:I

[Sha71h] Harold S. Shapiro. Introduction. Lecture Notes in Mathematics, 187:1–4, 1971. CODEN LNMAA2. ISBN 3-540-05376-X (print), 3-540-36497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058977/.

Shapiro:1971:MEF

[Sha71i] Harold S. Shapiro. Minimal extrapolation of Fourier transforms. Lecture Notes in Mathematics, 187:117–164, 1971. CODEN LN-MAA2. ISBN 3-540-05376-X (print), 3-540-36497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058983/.

Shapiro:1971:TAT

[Sha71j] Harold S. Shapiro. Topics in Approximation Theory, volume 187 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05376-X (print), 3-540-36497-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 233 (est.) pp. LCCN QA3 .L28 no. 187. URL http:/

/link.springer.com/book/10.1007/BFb0058976;   http://
www.springerlink.com/content/978-3-540-36497-9.

Shapiro:1971:SMF

[Sha71k] Leonard Shapiro. On the structure of minimal flows. Lecture Notes in Mathematics, 235:123–128, 1971. CODEN LN-MAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059204/.

Shapiro:1971:SC

[Sha71l] Leonard Shapiro. The structure of H-cascades. Lecture Notes in Mathematics, 235:117–122, 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059203/.

Shay:1971:DAB

[Sha71m] P. Brian Shay. Discoherently associative bifunctors on groups. Lecture Notes in Mathematics, 195:29–35, 1971. CODEN LN-MAA2. ISBN 3-540-05442-1 (print), 3-540-36548-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0072306/.

Shaneson:1973:HTT

[Sha73a] Julius L. Shaneson. Hermitian K-theory in topology. Lecture Notes in Mathematics, 343:1–40, 1973. CODEN LNMAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061367/.

Shaneson:1973:SPH

[Sha73b] Julius L. Shaneson. Some problems in Hermitian K-theory. Lecture Notes in Mathematics, 343:41–51, 1973. CODEN LN-MAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061368/.

Sharpe:1973:SU

[Sha73c] Richard Sharpe. Surgery and unitary K_{2}. Lecture Notes in Mathematics, 343:464–470, 1973. CODEN LNMAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061374/.

Shiffman:1971:EAS

[Shi71a] Bernard Shiffman. Extending analytic subvarieties. Lecture Notes in Mathematics, 184:208–222, 1971. CODEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059286/.

Shimura:1971:CFR

[Shi71b] Goro Shimura. Class fields over real quadratic fields in the theory of modular functions. Lecture Notes in Mathematics, 185: 169–188, 1971. CODEN LNMAA2. ISBN 3-540-05372-7 (print), 3-540-36493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058770/.

Shimura:1973:CM

[Shi73a] Goro Shimura. Complex multiplication. Lecture Notes in Mathematics, 320:37–56, 1973. CODEN LNMAA2. ISBN 3-540-06219-X (print), 3-540-38509-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-38509-7_2/.

Shimura:1973:MFH

[Shi73b] Goro Shimura. Modular forms of half integral weight. Lecture Notes in Mathematics, 320:57–74, 1973. CODEN LNMAA2. ISBN 3-540-06219-X (print), 3-540-38509-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38509-7_3/.

Shub:1971:ASP

[Shu71a] M. Shub. Appendix to Smale’s paper: Diagonals and relative equilibria. Lecture Notes in Mathematics, 197:199–201, 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068619/.

Shub:1971:I

[Shu71b] M. Shub. Instability. Lecture Notes in Mathematics, 206:28– 29, 1971. CODEN LNMAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070133.

Sibuya:1971:BVP

[Sib71a] Yasutaka Sibuya. A boundary value problem in the complex plane. Lecture Notes in Mathematics, 183:128–144, 1971. CO-DEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060413/.

Sibuya:1971:PIS

[Sib71b] Yasutaka Sibuya. Perturbation at an irregular singular point. Lecture Notes in Mathematics, 243:148–168, 1971. CODEN LN-MAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058725/.

Sibony:1974:PAF

[Sib74] Nessim Sibony. Prolongement analytique des fonctions holomorphes bornées. (French) []. Lecture Notes in Mathematics, 410: 44–66, 1974. CODEN LNMAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066039/.

Siciak:1972:PLA

[Sic72] J. Siciak. A polynomial lemma and analytic mappings in topological vector spaces. Lecture Notes in Mathematics, 275:131– 142, 1972. CODEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058496/.

Siciak:1974:WAC

[Sic74] Józef Siciak. Weak analytic continuation from compact subsets of C^{n}. Lecture Notes in Mathematics, 364:92–96, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069007/.

Siebenmann:1974:LTD

[Sie74] Laurent Siebenmann. L’invariance topologique du type simple d’homotopie. (French) []. Lecture Notes in Mathematics, 383: 186–209, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057309/.

Sigrist:1971:MBS

[Sig71a] François Sigrist. H-maps between spheres. Lecture Notes in Mathematics, 196:39–41, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067744/.

Sigrist:1971:SAR

[Sig71b] François Sigrist, editor. H-Spaces: Actes de la réunion de Neuchâtel (Suisse), août 1970, volume 196 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 196.

URL http://link.springer.com/book/10.1007/BFb0067737;
 http://www.springerlink.com/content/978-3-540-36621-

Silberger:1970:ARS

[Sil70a] Allan J. Silberger. PGL_{2} over the p-adics: its Representations, Spherical Functions, and Fourier Analysis, volume 166 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05193-7 (print), 3-540-36394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 204 pp. LCCN QA3 .L28 no. 166. URL http:/

/link.springer.com/book/10.1007/BFb0059369;   http://
www.springerlink.com/content/978-3-540-36394-1.

Silberger:1970:CCG

[Sil70b] Allan J. Silberger. Computation of the characters of G and of K. Lecture Notes in Mathematics, 166:108–138, 1970. CODEN LNMAA2. ISBN 3-540-05193-7 (print), 3-540-36394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059373/.

Silberger:1970:DAI

[Sil70c] Allan J. Silberger. A description of all irreducible representations of K = PGL(2, 0). Lecture Notes in Mathematics, 166: 52–107, 1970. CODEN LNMAA2. ISBN 3-540-05193-7 (print), 3-540-36394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059372/.

Silberger:1970:FTI

[Sil70d] Allan J. Silberger. The Fourier transformation and inversion formulas for S_{u}^{0} for all u. Lecture Notes in Mathematics, 166:139– 201, 1970. CODEN LNMAA2. ISBN 3-540-05193-7 (print), 3-540-36394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059374/.

Silberger:1970:IG

[Sil70e] Allan J. Silberger. Introduction and generalities. Lecture Notes in Mathematics, 166:1–35, 1970. CODEN LNMAA2. ISBN 3-540-05193-7 (print), 3-540-36394-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059370/.

Silberger:1970:IUR

[Sil70f] Allan J. Silberger. The irreducible unitary representations of PGL(2, Ω) and their characters. Lecture Notes in Mathematics, 166:36–51, 1970. CODEN LNMAA2. ISBN 3-540-05193-7 (print), 3-540-36394-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059371/.

Silverstein:1974:DDF

[Sil74a] Prof. Martin L. Silverstein. Decomposition of the Dirichlet form. Lecture Notes in Mathematics, 426:97–129, 1974. CODEN LN-MAA2. ISBN 3-540-07012-5 (print), 3-540-37292-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073685/.

Silverstein:1974:E

[Sil74b] Prof. Martin L. Silverstein. Examples. Lecture Notes in Mathematics, 426:220–270, 1974. CODEN LNMAA2. ISBN 3-540-07012-5 (print), 3-540-37292-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073687/.

Silverstein:1974:GT

[Sil74c] Prof. Martin L. Silverstein. General theory. Lecture Notes in Mathematics, 426:1–96, 1974. CODEN LNMAA2. ISBN 3-540-07012-5 (print), 3-540-37292-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073684/.

Silverstein:1974:ST

[Sil74d] Prof. Martin L. Silverstein. Structure theory. Lecture Notes in Mathematics, 426:130–219, 1974. CODEN LNMAA2. ISBN 3-540-07012-5 (print), 3-540-37292-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0073686/.

Silverstein:1974:SMP

[Sil74e] Prof. Martin L. Silverstein. Symmetric Markov Processes, volume 426 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-07012-5 (print), 3-540-37292-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 287 pp. LCCN QA3 .L28 no. 426. URL http:

//link.springer.com/book/10.1007/BFb0073683;   http:/
/www.springerlink.com/content/978-3-540-37292-9.

Simader:1972:DBV

[Sim72a] Christian G. Simader. On Dirichlet’s Boundary Value Problem: An L_{p}-Theory Based on a Generalization of Gårding’s Inequality, volume 268 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05903-2 (print), 3-540-37589-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 199 (est.) pp. LCCN QA3 .L28 no. 268. URL http:/

/link.springer.com/book/10.1007/BFb0058774;   http://
www.springerlink.com/content/978-3-540-37589-0.

Simader:1972:O

[Sim72b] Christian G. Simader. Outline. Lecture Notes in Mathematics, 268:1–12, 1972. CODEN LNMAA2. ISBN 3-540-05903-2 (print), 3-540-37589-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058775/.

Simader:1972:PES

[Sim72c] Christian G. Simader. A priori estimates for solutions of linear elliptic functional equations with constant coefficients. Lecture Notes in Mathematics, 268:13–83, 1972. CODEN LN-MAA2. ISBN 3-540-05903-2 (print), 3-540-37589-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058776/.

Simader:1972:RET

[Sim72d] Christian G. Simader. Regularity and existence theorems for uniformly elliptic functional equations. Lecture Notes in Mathematics, 268:133–199, 1972. CODEN LNMAA2. ISBN 3-540-05903-2 (print), 3-540-37589-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058778/.

Simader:1972:RCL

[Sim72e] Christian G. Simader. A representation for continuous linear functionals on W_{o}^{m,p}(G)(1 < p < ∞) and its applications: a generalization of Gårding’s inequality and existence theorems.

Lecture Notes in Mathematics, 268:84–132, 1972. CODEN LN-MAA2. ISBN 3-540-05903-2 (print), 3-540-37589-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058777/.

Simpson:1973:GFS

[Sim73] R. B. Simpson. Groundwater flow as a singular perturbation problem and remarks about numerical methods. Lecture Notes in Mathematics, 322:288–297, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060572/.

Simonovits:1974:NHE

[Sim74] M. Simonovits. Note on a hypergraph extremal problem. Lecture Notes in Mathematics, 411:147–151, 1974. CODEN LN-MAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066189/.

Singh:1973:CMS

[Sin73] J. M. Singh. A complete metric space of sub-algebras. Lecture Notes in Mathematics, 296:165–180, 1973. CODEN LN-MAA2. ISBN 3-540-06211-4 (print), 3-540-38485-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059824/.

Sion:1973:D

[Sio73a] Maurice Sion. Differentiation. Lecture Notes in Mathematics, 355:82–124, 1973. CODEN LNMAA2. ISBN 3-540-06542-3 (print), 3-540-37839-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060136/.

Sion:1973:MIF

[Sio73b] Maurice Sion. Measurable and integrable functions. Lecture Notes in Mathematics, 355:38–81, 1973. CODEN LN-MAA2. ISBN 3-540-06542-3 (print), 3-540-37839-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060135/.

Sion:1973:SVO

[Sio73c] Maurice Sion. Semigroup valued outer measures. Lecture Notes in Mathematics, 355:1–37, 1973. CODEN LNMAA2. ISBN 3-540-06542-3 (print), 3-540-37839-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060134/.

Sion:1973:TSV

[Sio73d] Maurice Sion. A Theory of Semigroup Valued Measures, volume 355 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06542-3 (print), 3-540-37839-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 140 pp. LCCN QA3 .L28 no. 355. URL http:

//link.springer.com/book/10.1007/BFb0060133;   http:/
/www.springerlink.com/content/978-3-540-37839-6.

Siu:1971:GGD

[Siu71a] Yum-Tong Siu. Generalizations of Grauert’s direct image theorem. Lecture Notes in Mathematics, 184:151–174, 1971. CODEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059283/.

Siu:1971:OTE

[Siu71b] Yum-Tong Siu. An Osgood type extension theorem for coherent analytic sheaves. Lecture Notes in Mathematics, 185:189– 241, 1971. CODEN LNMAA2. ISBN 3-540-05372-7 (print), 3-540-36493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058771/.

Stakgold:1973:NPP

[SJS73] Ivar Stakgold, Daniel D. Joseph, and David H. Sattinger, editors. Nonlinear Problems in the Physical Sciences and Biology: Proceedings of a Battelle Summer Institute, Seattle, July 3–28, 1972, volume 322 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 322. URL http:/

/link.springer.com/book/10.1007/BFb0060558;   http://
www.springerlink.com/content/978-3-540-38558-5.

Skoda:1972:CFE

[Sko72] H. Skoda. Croissance des fonctions entières s’annulant sur une hypersurface donnée de C^{n}. (French) []. Lecture Notes in Mathematics, 275:82–105, 1972. CODEN LNMAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058493/.

Skoda:1973:FHD

[Sko73] Henri Skoda. Formulation hilbertienne du Nullstellensatz dans les algèbres de fonctions holomorphes. (French) []. Lecture Notes in Mathematics, 336:95–103, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065790/.

Skoda:1974:NMP

[Sko74] H. Skoda. Nouvelle méthode pour l’étude des potentiels associés aux ensembles analytiques. (French) []. Lecture Notes in Mathematics, 410:117–141, 1974. CODEN LNMAA2. ISBN 3-540-06858-9 (print), 3-540-37816-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066045/.

Shioda:1974:SAS

[SM74a] Tetsuji Shioda and Naoki Mitani. Singular abelian surfaces and binary quadratic forms. Lecture Notes in Mathematics, 412:259– 287, 1974. CODEN LNMAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066163/.

Sleeman:1974:OPD

[SM74b] Dr. B. D. Sleeman and Dr. I. M. Michael, editors. Ordinary and Partial Differential Equations: Proceedings of the Conference held at Dundee, Scotland, 26–29 March, 1974, volume 415 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06959-3 (print), 3-540-37264-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 415. URL http:/

/link.springer.com/book/10.1007/BFb0065505;   http://
www.springerlink.com/content/978-3-540-37264-6.

Smale:1971:PNR

[Sma71a] S. Smale. Problems on the nature of relative equilibria in celestial mechanics. Lecture Notes in Mathematics, 197:194–198, 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0068618/.

Smale:1971:TM

[Sma71b] S. Smale. Topology and mechanics. Lecture Notes in Mathematics, 206:33–34, 1971. CODEN LNMAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070135.

Smale:1971:SGD

[Sma71c] Stephen Smale. Stability and genericity in dynamical systems. Lecture Notes in Mathematics, 180:177–185, 1971. CODEN LN-MAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058831/.

Smith:1970:CCA

[Smi70a] Larry Smith. The cobar construction: Applications. Lecture Notes in Mathematics, 134:118–138, 1970. CODEN LN-MAA2. ISBN 3-540-04923-1 (print), 3-540-36286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060350/.

Smith:1970:EMS

[Smi70b] Larry Smith. The Eilenberg–Moore spectral sequence as a Kunneth spectral sequence. Lecture Notes in Mathematics, 134:1– 53, 1970. CODEN LNMAA2. ISBN 3-540-04923-1 (print), 3-540-36286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060348/.

Smith:1970:LEM

[Smi70c] Larry Smith. Lectures on the Eilenberg–Moore Spectral Sequence, volume 134 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04923-1 (print), 3-540-36286-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 142 pp. LCCN QA3 .L28 no. 134. URL http:/

/link.springer.com/book/10.1007/BFb0060347;   http://
www.springerlink.com/content/978-3-540-36286-9.

Smith:1970:MF

[Smi70d] Larry Smith. Multiplicative fibrations. Lecture Notes in Mathematics, 134:54–117, 1970. CODEN LNMAA2. ISBN 3-540-04923-1 (print), 3-540-36286-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060349/.

Smith:1972:GAF

[Smi72] David A. Smith. Generalized arithmetic function algebras. Lecture Notes in Mathematics, 251:205–245, 1972. CODEN LN-MAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058795/.

Smith:1974:IET

[Smi74] Haywood Smith, Jr. Integration errors and their effects on macroscopic properties of calculated N-body systems. Lecture Notes in Mathematics, 362:360–373, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066601/.

Smorodinsky:1971:CPE

[Smo71a] Meir Smorodinsky. Conditional probabilities and expectations. Doob’s martingale theorem. Lecture Notes in Mathematics, 214: 13–18, 1971. CODEN LNMAA2. ISBN 3-540-05556-8 (print), 3-540-36879-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066090/.

Smorodinsky:1971:D

[Smo71b] Meir Smorodinsky. Discussion. Lecture Notes in Mathematics, 214:58–60, 1971. CODEN LNMAA2. ISBN 3-540-05556-8 (print), 3-540-36879-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066097/.

Smorodinsky:1971:ECE

[Smo71c] Meir Smorodinsky. Entropy and conditional entropy. Lecture Notes in Mathematics, 214:19–28, 1971. CODEN LN-MAA2. ISBN 3-540-05556-8 (print), 3-540-36879-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066091/.

Smorodinsky:1971:EGK

[Smo71d] Meir Smorodinsky. Entropy and generators. Krieger’s theorem. Lecture Notes in Mathematics, 214:54–57, 1971. CODEN LN-MAA2. ISBN 3-540-05556-8 (print), 3-540-36879-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066096/.

Smorodinsky:1971:ETb

[Smo71e] Meir Smorodinsky. Entropy of a transformation. Lecture Notes in Mathematics, 214:29–33, 1971. CODEN LNMAA2. ISBN 3-540-05556-8 (print), 3-540-36879-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066092/.

Smorodinsky:1971:ETa

[Smo71f] Meir Smorodinsky. Ergodic theorem. Lecture Notes in Mathematics, 214:8–12, 1971. CODEN LNMAA2. ISBN 3-540-05556-8 (print), 3-540-36879-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066089/.

Smorodinsky:1971:ETE

[Smo71g] Meir Smorodinsky. Ergodic Theory, Entropy, volume 214 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05556-8 (print), 3-540-36879-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 64 pp. LCCN QA3 .L28 no. 214. URL http:/

/link.springer.com/book/10.1007/BFb0066086;   http://
www.springerlink.com/content/978-3-540-36879-3.

Smorodinsky:1971:E

[Smo71h] Meir Smorodinsky. Examples. Lecture Notes in Mathematics, 214:2–7, 1971. CODEN LNMAA2. ISBN 3-540-05556-8 (print), 3-540-36879-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066088/.

Smorodinsky:1971:IBS

[Smo71i] Meir Smorodinsky. Isomorphism of Bernoulli shifts Sinai and Ornstein’s theorems. Lecture Notes in Mathematics, 214:43– 53, 1971. CODEN LNMAA2. ISBN 3-540-05556-8 (print), 3-540-36879-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066095/.

Smorodinsky:1971:MKA

[Smo71j] Meir Smorodinsky. Mixing, Kolmogorov-automorphisms. Lecture Notes in Mathematics, 214:36–42, 1971. CODEN LN-MAA2. ISBN 3-540-05556-8 (print), 3-540-36879-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066094/.

Smorodinsky:1971:NS

[Smo71k] Meir Smorodinsky. Notation, symbols. Lecture Notes in Mathematics, 214:1, 1971. CODEN LNMAA2. ISBN 3-540-05556-8 (print), 3-540-36879-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066087.

Smorodinsky:1971:SMB

[Smo71l] Meir Smorodinsky. The Shannon–McMillan–Breiman theorem. Lecture Notes in Mathematics, 214:34–35, 1971. CODEN LN-MAA2. ISBN 3-540-05556-8 (print), 3-540-36879-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066093.

Smorynski:1973:AKM

[Smo73] C. A. Smoryński. Applications of Kripke models. Lecture Notes in Mathematics, 344:324–391, 1973. CODEN LNMAA2. ISBN 3-540-06491-5 (print), 3-540-37806-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066744/;
 http://link.springer.com/content/pdf/bfm:978-3-540-
37806-8/1.pdf.

Snell:1971:PTT

[Sne71] L. J. Snell. On the proofs of two theorems in probabilistic potential theory. Lecture Notes in Mathematics, 190:49–59, 1971. CODEN LNMAA2. ISBN 3-540-05396-4 (print), 3-540-36515-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065892/.

Stakgold:1973:NPN

[SP73] I. Stakgold and L. E. Payne. Nonlinear problems in nuclear reactor analysis. Lecture Notes in Mathematics, 322:298–307, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060573/.

Spencer:1971:FDO

[Spe71] D. C. Spencer. Flat differential operators. Lecture Notes in Mathematics, 184:85–108, 1971. CODEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059280/.

Spencer:1974:CAD

[Spe74] D. C. Spencer. Curvatures associated with differential operators. Lecture Notes in Mathematics, 392:64–75, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064113/.

Spijker:1972:ETN

[Spi72] M. N. Spijker. Equivalence theorems for nonlinear finitedifference methods. Lecture Notes in Mathematics, 267:233– 264, 1972. CODEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061621/.

Spijker:1974:TSE

[Spi74a] M. N. Spijker. Two-sided error estimates in the numerical solution of initial value problems. Lecture Notes in Mathematics, 395: 109–122, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060666/.

Spitzer:1974:IPM

[Spi74b] Frank L. Spitzer. Introduction aux processus de Markov à Paramètre dans Z_{ν}. (French) []. Lecture Notes in Mathematics, 390:114–189, 1974. CODEN LNMAA2. ISBN 3-540-06816-3 (print), 3-540-37765-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0082435/.

Springer:1970:CSG

[Spr70a] T. A. Springer. Characters of special groups. Lecture Notes in Mathematics, 131:121–166, 1970. CODEN LNMAA2. ISBN 3-540-04920-7 (print), 3-540-36272-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081545/.

Springer:1970:CFF

[Spr70b] T. A. Springer. Cusp forms for finite groups. Lecture Notes in Mathematics, 131:97–120, 1970. CODEN LNMAA2. ISBN 3-540-04920-7 (print), 3-540-36272-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081544/.

Springer:1974:CGC

[Spr74] T. A. Springer. Caractères de groupes de Chevalley finis. (French) []. Lecture Notes in Mathematics, 383:210–233, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057310/.

Stummel:1973:DCC

[SR73] Friedrich Stummel and Jürgen Reinhardt. Discrete convergence of continuous mappings in metric spaces. Lecture Notes in Mathematics, 333:218–242, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060700/.

Springer:1970:CC

[SS70] T. A. Springer and R. Steinberg. Conjugacy classes. Lecture Notes in Mathematics, 131:167–266, 1970. CODEN LN-MAA2. ISBN 3-540-04920-7 (print), 3-540-36272-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081546/.

Siu:1971:C

[ST71a] Yum-Tong Siu and Günther Trautmann. ρ-convexity. Lecture Notes in Mathematics, 172:113–120, 1971. CODEN LN-MAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060584/.

Siu:1971:CCM

[ST71b] Yum-Tong Siu and Günther Trautmann. Closedness of coboundary modules. Lecture Notes in Mathematics, 172:80–97, 1971. CODEN LNMAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060582/.

Siu:1971:D

[ST71c] Yum-Tong Siu and Günther Trautmann. Duality. Lecture Notes in Mathematics, 172:98–112, 1971. CODEN LNMAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060583/.

Siu:1971:EAC

[ST71d] Yum-Tong Siu and Günther Trautmann. Extension of analytic covers. Lecture Notes in Mathematics, 172:121–127, 1971. CO-DEN LNMAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060585/.

Siu:1971:GSE

[ST71e] Yum-Tong Siu and Günther Trautmann. Gap-Sheaves and Estension of Coherent Analytic Subsheaves, volume 172 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 172 pp. LCCN QA3 .L28 no. 172. URL http:/

/link.springer.com/book/10.1007/BFb0060577;   http://
www.springerlink.com/content/978-3-540-36429-0.

Siu:1971:G

[ST71f] Yum-Tong Siu and Günther Trautmann. Globalization. Lecture Notes in Mathematics, 172:153–165, 1971. CODEN LN-MAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060588/.

Siu:1971:HN

[ST71g] Yum-Tong Siu and Günther Trautmann. Historical notes. Lecture Notes in Mathematics, 172:166–168, 1971. CODEN LN-MAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060589/.

Siu:1971:P

[ST71h] Yum-Tong Siu and Günther Trautmann. Preliminaries. Lecture Notes in Mathematics, 172:1–21, 1971. CODEN LNMAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060578/.

Siu:1971:PDR

[ST71i] Yum-Tong Siu and Günther Trautmann. Primary decomposition and relative gap-sheaves. Lecture Notes in Mathematics, 172: 45–61, 1971. CODEN LNMAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060580/.

Siu:1971:SLC

[ST71j] Yum-Tong Siu and Günther Trautmann. Sheaves of local cohomology and absolute gap-sheaves. Lecture Notes in Mathematics, 172:62–79, 1971. CODEN LNMAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060581/.

Siu:1971:SCS

[ST71k] Yum-Tong Siu and Günther Trautmann. Singularities of coherent sheaves and their cohomology classes. Lecture Notes in Mathematics, 172:22–44, 1971. CODEN LNMAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060579/.

Siu:1971:SE

[ST71l] Yum-Tong Siu and Günther Trautmann. Subsheaf extension. Lecture Notes in Mathematics, 172:144–152, 1971. CODEN LN-MAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060587/.

Siu:1971:SEP

[ST71m] Yum-Tong Siu and Günther Trautmann. Subvariety extension. Projection lemma. Lecture Notes in Mathematics, 172:128– 143, 1971. CODEN LNMAA2. ISBN 3-540-05294-1 (print), 3-540-36429-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060586/.

Stasheff:1970:Sa

[Sta70a] James Stasheff. A_{n}-spaces. Lecture Notes in Mathematics, 161: 48–57, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065908/.

Stasheff:1970:ALS

[Sta70b] James Stasheff. Associativity: Loop spaces and topological groups. Lecture Notes in Mathematics, 161:14–19, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065901/.

Stasheff:1970:BCS

[Sta70c] James Stasheff. The bar construction spectral sequence. Lecture Notes in Mathematics, 161:23–26, 1970. CODEN LN-MAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065903/.

Stasheff:1970:DBC

[Sta70d] James Stasheff. Differentials in the bar construction. Lecture Notes in Mathematics, 161:44–47, 1970. CODEN LN-MAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065907/.

Stasheff:1970:SHPa

[Sta70e] James Stasheff. H-spaces from a homotopy point of view. Lecture Notes in Mathematics, 161:1–2, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065897.

Stasheff:1970:SHPb

[Sta70f] James Stasheff. H-spaces from a Homotopy Point of View, volume 161 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 95 pp. LCCN QA3 .L28 no. 161. URL http:/

/link.springer.com/book/10.1007/BFb0065896;   http://
www.springerlink.com/content/978-3-540-36314-9.

Stasheff:1970:SWF

[Sta70g] James Stasheff. H-spaces which are finite complexes. Lecture Notes in Mathematics, 161:20–22, 1970. CODEN LN-MAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065902/.

Stasheff:1970:HA

[Sta70h] James Stasheff. Homotopy associativity. Lecture Notes in Mathematics, 161:27–30, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065904/.

Stasheff:1970:HCb

[Sta70i] James Stasheff. Homotopy commutativity. Lecture Notes in Mathematics, 161:65–70, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065910/.

Stasheff:1970:HCa

[Sta70j] James Stasheff. The Hopf construction. Lecture Notes in Mathematics, 161:3–6, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065898/.

Stasheff:1970:ILS

[Sta70k] James Stasheff. Infinite loop spaces. Lecture Notes in Mathematics, 161:75–80, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065912/.

Stasheff:1970:MSA

[Sta70l] James Stasheff. Maps into an H-space: Algebraic structure, inverses, other multiplications, etc. Lecture Notes in Mathematics, 161:10–13, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065900/.

Stasheff:1970:MS

[Sta70m] James Stasheff. Maps of H-spaces. Lecture Notes in Mathematics, 161:31–37, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065905/.

Stasheff:1970:MPG

[Sta70n] James Stasheff. Massey products and generalized bar construction. Lecture Notes in Mathematics, 161:59–64, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065909/.

Stasheff:1970:OIL

[Sta70o] James Stasheff. Operations in iterated loop spaces. Lecture Notes in Mathematics, 161:81–88, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065913/.

Stasheff:1970:PP

[Sta70p] James Stasheff. The projective plane. Lecture Notes in Mathematics, 161:7–9, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065899/.

Stasheff:1970:SIM

[Sta70q] James Stasheff. Spaces induced by H-maps. Lecture Notes in Mathematics, 161:38–43, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065906/.

Stasheff:1970:Sb

[Sta70r] James Stasheff. Structure on B_{X}. Lecture Notes in Mathematics, 161:71–74, 1970. CODEN LNMAA2. ISBN 3-540-04940-1 (print), 3-540-36314-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065911/.

Stasheff:1971:FFC

[Sta71a] James D. Stasheff. Families of finite H-complexes — revisited. Lecture Notes in Mathematics, 196:1–4, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067738/.

Stasheff:1971:SP

[Sta71b] James D. Stasheff. H-space problems. Lecture Notes in Mathematics, 196:122–136, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067755/.

Stasheff:1971:ILS

[Sta71c] James D. Stasheff. Infinite loop spaces — an historical survey. Lecture Notes in Mathematics, 196:43–53, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067746/.

Stasheff:1971:SBS

[Sta71d] James Dillon Stasheff. Sphere bundles over spheres as H-spaces modp > 2. Lecture Notes in Mathematics, 249:106–110, 1971. CODEN LNMAA2. ISBN 3-540-05715-3 (print), 3-540-37082-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060901/.

Stampacchia:1972:RRT

[Sta72] G. Stampacchia. Recent results in the theory of variational inequalities. Lecture Notes in Mathematics, 280:147–154, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066925/.

Stammbach:1973:CE

[Sta73a] Urs Stammbach. Central extensions. Lecture Notes in Mathematics, 359:95–145, 1973. CODEN LNMAA2. ISBN 3-540-06569-5 (print), 3-540-37870-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067182/.

Stammbach:1973:EHT

[Sta73b] Urs Stammbach. Elements of homology theory. Lecture Notes in Mathematics, 359:11–35, 1973. CODEN LNMAA2. ISBN 3-540-06569-5 (print), 3-540-37870-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067179/.

Stammbach:1973:EAK

[Sta73c] Urs Stammbach. Extensions in V with Abelian kernel. Lecture Notes in Mathematics, 359:37–61, 1973. CODEN LN-MAA2. ISBN 3-540-06569-5 (print), 3-540-37870-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067180/.

Stammbach:1973:HGT

[Sta73d] Urs Stammbach. Homology in Group Theory, volume 359 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06569-5 (print), 3-540-37870-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 183 pp. LCCN QA3 .L28 no. 359. URL http:/

/link.springer.com/book/10.1007/BFb0067177;   http://
www.springerlink.com/content/978-3-540-37870-9.

Stammbach:1973:LNG

[Sta73e] Urs Stammbach. Localization of nilpotent groups. Lecture Notes in Mathematics, 359:147–178, 1973. CODEN LNMAA2. ISBN 3-540-06569-5 (print), 3-540-37870-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067183/.

Stammbach:1973:LCS

[Sta73f] Urs Stammbach. The lower central series. Lecture Notes in Mathematics, 359:63–93, 1973. CODEN LNMAA2. ISBN 3-540-06569-5 (print), 3-540-37870-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067181/.

Stammbach:1973:VG

[Sta73g] Urs Stammbach. Varieties of groups. Lecture Notes in Mathematics, 359:1–9, 1973. CODEN LNMAA2. ISBN 3-540-06569-5 (print), 3-540-37870-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067178/.

Staples:1973:CRC

[Sta73h] John Staples. Combinator realizability of constructive finite type analysis. Lecture Notes in Mathematics, 337:253–273, 1973. CO-DEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066777/.

Stark:1973:CNC

[Sta73i] H. M. Stark. Class-numbers of complex quadratic fields. Lecture Notes in Mathematics, 320:153–174, 1973. CODEN LNMAA2. ISBN 3-540-06219-X (print), 3-540-38509-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38509-7_5/.

Stasheff:1974:SNM

[Sta74a] James Stasheff. H-space newsletter — May, 1974. Lecture Notes in Mathematics, 418:139–141, 1974. CODEN LNMAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070649/.

Stasheff:1974:PTC

[Sta74b] James D. Stasheff. Parallel transport and classification of fibrations. Lecture Notes in Mathematics, 428:1–17, 1974. CODEN LNMAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070531/.

Stasheff:1974:DLG

[Sta74c] James Dillon Stasheff. The modp decomposition of Lie groups. Lecture Notes in Mathematics, 418:142–149, 1974. CODEN LN-MAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070650/.

Stein:1970:VLP

[Ste70a] Professor E. M. Stein. Variations on the Littlewood–Paley theme. Lecture Notes in Mathematics, 170:1–17, 1970. CODEN LN-MAA2. ISBN 3-540-05284-4 (print), 3-540-36417-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0079064/.

Stewart:1970:AGS

[Ste70b] Ian Stewart. Algebras with a given star. Lecture Notes in Mathematics, 127:48–57, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061090/.

Stewart:1970:BF

[Ste70c] Ian Stewart. Baer ≠ fitting. Lecture Notes in Mathematics, 127: 74–79, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061094/.

Stewart:1970:BD

[Ste70d] Ian Stewart. Basic definitions. Lecture Notes in Mathematics, 127:1–8, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061082/.

Stewart:1970:BR

[Ste70e] Ian Stewart. The bear radical. Lecture Notes in Mathematics, 127:65–70, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061092/.

Stewart:1970:CD

[Ste70f] Ian Stewart. The Cartan decomposition. Lecture Notes in Mathematics, 127:24–27, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061086/.

Stewart:1970:CS

[Ste70g] Ian Stewart. Cartan subalgebras. Lecture Notes in Mathematics, 127:15–18, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061084/.

Stewart:1970:DD

[Ste70h] Ian Stewart. Dynkin diagrams. Lecture Notes in Mathematics, 127:34–43, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061088/.

Stewart:1970:KF

[Ste70i] Ian Stewart. The Killing form. Lecture Notes in Mathematics, 127:19–23, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061085/.

Stewart:1970:LA

[Ste70j] Ian Stewart. Lie Algebras, volume 127 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 97 pp. LCCN QA3 .L28 no. 127. URL http:/

/link.springer.com/book/10.1007/BFb0061081;   http://
www.springerlink.com/content/978-3-540-36264-7.

Stewart:1970:LAW

[Ste70k] Ian Stewart. Lie algebras in which every subalgebra is a subideal. Lecture Notes in Mathematics, 127:80–89, 1970. CODEN LN-MAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061095/.

Stewart:1970:MCS

[Ste70l] Ian Stewart. The minimal condition for subideals. Lecture Notes in Mathematics, 127:90–95, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061096/.

Stewart:1970:OR

[Ste70m] Ian Stewart. Other radicals. Lecture Notes in Mathematics, 127: 71–73, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061093/.

Stewart:1970:RNA

[Ste70n] Ian Stewart. Representations of nilpotent algebras. Lecture Notes in Mathematics, 127:9–14, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061083/.

Stewart:1970:SAO

[Ste70o] Ian Stewart. Some astronomical observations. Lecture Notes in Mathematics, 127:44–47, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061089/.

Stewart:1970:SDA

[Ste70p] Ian Stewart. Subideals, derivations, automorphisms. Lecture Notes in Mathematics, 127:58–64, 1970. CODEN LN-MAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061091/.

Stewart:1970:SFR

[Ste70q] Ian Stewart. Systems of fundamental roots. Lecture Notes in Mathematics, 127:28–33, 1970. CODEN LNMAA2. ISBN 3-540-04916-9 (print), 3-540-36264-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061087/.

Steenrod:1971:PAA

[Ste71a] Norman Steenrod. Polynomial algebras over the algebra of cohomology operations. Lecture Notes in Mathematics, 196:85– 99, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067751/.

Stenstrom:1971:CMQ

[Ste71b] Bo Stenström. Categories of modules of quotients. Lecture Notes in Mathematics, 237:33–57, 1971. CODEN LNMAA2. ISBN 3-540-05690-4 (print), 3-540-37002-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059906/.

Stenstrom:1971:GPR

[Ste71c] Bo Stenström. General properties of rings of quotients. Lecture Notes in Mathematics, 237:58–92, 1971. CODEN LN-MAA2. ISBN 3-540-05690-4 (print), 3-540-37002-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059907/.

Stenstrom:1971:MCR

[Ste71d] Bo Stenström. Maximal and classical rings of quotients. Lecture Notes in Mathematics, 237:110–129, 1971. CODEN LN-MAA2. ISBN 3-540-05690-4 (print), 3-540-37002-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059909/.

Stenstrom:1971:RMQ

[Ste71e] Bo Stenström. Rings and Modules of Quotients, volume 237 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05690-4 (print), 3-540-37002-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 136 pp. LCCN QA3 .L28 no. 237. URL http:/

/link.springer.com/book/10.1007/BFb0059904;   http://
www.springerlink.com/content/978-3-540-37002-4.

Stenstrom:1971:SIR

[Ste71f] Bo Stenström. Self-injective rings. Lecture Notes in Mathematics, 237:93–109, 1971. CODEN LNMAA2. ISBN 3-540-05690-4 (print), 3-540-37002-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059908/.

Stenstrom:1971:TT

[Ste71g] Bo Stenström. Torsion theory. Lecture Notes in Mathematics, 237:1–32, 1971. CODEN LNMAA2. ISBN 3-540-05690-4 (print), 3-540-37002-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059905/.

Stetter:1971:SDI

[Ste71h] Hans J. Stetter. Stability of discretizations on infinite intervals. Lecture Notes in Mathematics, 228:207–222, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069457/.

Stehle:1972:PDD

[Ste72] Jean-Luc Stehlé. Plongements du disque dans C^{2}. (French) []. Lecture Notes in Mathematics, 275:119–130, 1972. CODEN LN-MAA2. ISBN 3-540-05921-0 (print), 3-540-37610-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058495/.

Steinberg:1974:AHS

[Ste74a] Robert Steinberg. Abstract homomorphisms of simple algebraic groups. Lecture Notes in Mathematics, 383:307–326, 1974. CO-DEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057316/.

Steinberg:1974:AAV

[Ste74b] Robert Steinberg. Affine algebraic varieties, affine algebraic groups and their orbits. Lecture Notes in Mathematics, 366:1– 21, 1974. CODEN LNMAA2. ISBN 3-540-06657-8 (print), 3-540-37931-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067855/.

Steinberg:1974:CCA

[Ste74c] Robert Steinberg. Conjugacy Classes in Algebraic Groups, volume 366 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06657-8 (print), 3-540-37931-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 159 pp. LCCN QA3 .L28 no. 366. URL http:

//link.springer.com/book/10.1007/BFb0067854;   http:/
/www.springerlink.com/content/978-3-540-37931-7.

Steinberg:1974:FPJ

[Ste74d] Robert Steinberg. First part: Jordan decompositions, unipotent and diagonalizable groups. Lecture Notes in Mathematics, 366: 22–45, 1974. CODEN LNMAA2. ISBN 3-540-06657-8 (print), 3-540-37931-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067856/.

Steinberg:1974:RSA

[Ste74e] Robert Steinberg. Reductive and semisimple algebraic groups, regular and subregular elements. Lecture Notes in Mathematics, 366:76–155, 1974. CODEN LNMAA2. ISBN 3-540-06657-8 (print), 3-540-37931-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067858/.

Steinberg:1974:SPQ

[Ste74f] Robert Steinberg. Second part: Quotients and solvable groups. Lecture Notes in Mathematics, 366:46–75, 1974. CODEN LN-MAA2. ISBN 3-540-06657-8 (print), 3-540-37931-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067857/.

Stenger:1974:CEB

[Ste74g] Frank Stenger. On the convergence and error of the Bubnov– Galerkin method. Lecture Notes in Mathematics, 362:434–450, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic).

Sterboul:1974:CND

[Ste74h] F. Sterboul. On the chromatic number of the direct product of hypergraphs. Lecture Notes in Mathematics, 411:165–174, 1974.

CODEN LNMAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066191/.

Sterboul:1974:CVC

[Ste74i] F. Sterboul. Sur une conjecture de V. Chvatal. (French) []. Lecture Notes in Mathematics, 411:152–164, 1974. CODEN LN-MAA2. ISBN 3-540-06846-5 (print), 3-540-37803-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066190/.

Stetter:1974:CFD

[Ste74j] Hans J. Stetter. Cyclic finite-difference methods for ordinary differential equations. Lecture Notes in Mathematics, 363:134– 143, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069131/.

Stoll:1970:AF

[Sto70a] Wilhelm Stoll. Admissible families. Lecture Notes in Mathematics, 135:17–28, 1970. CODEN LNMAA2. ISBN 3-540-04924-X (print), 3-540-36287-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059120/.

Stoll:1970:CSP

[Sto70b] Wilhelm Stoll. The construction of singular potentials. Lecture Notes in Mathematics, 135:52–63, 1970. CODEN LN-MAA2. ISBN 3-540-04924-X (print), 3-540-36287-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059123/.

Stoll:1970:CPF

[Sto70c] Wilhelm Stoll. The construction of the proximity form. Lecture Notes in Mathematics, 135:81–100, 1970. CODEN LN-MAA2. ISBN 3-540-04924-X (print), 3-540-36287-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059125/.

Stoll:1970:DPF

[Sto70d] Wilhelm Stoll. The definition of the proximity form. Lecture Notes in Mathematics, 135:29–42, 1970. CODEN LN-MAA2. ISBN 3-540-04924-X (print), 3-540-36287-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059121/.

Stoll:1970:E

[Sto70e] Wilhelm Stoll. Equidistribution. Lecture Notes in Mathematics, 135:114–134, 1970. CODEN LNMAA2. ISBN 3-540-04924-X (print), 3-540-36287-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059127/.

Stoll:1970:FMT

[Sto70f] Wilhelm Stoll. The first main theorem. Lecture Notes in Mathematics, 135:43–51, 1970. CODEN LNMAA2. ISBN 3-540-04924-X (print), 3-540-36287-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059122/.

Stoll:1970:I

[Sto70g] Wilhelm Stoll. Introduction. Lecture Notes in Mathematics, 135: 5–16, 1970. CODEN LNMAA2. ISBN 3-540-04924-X (print), 3-540-36287-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059119/.

Stoll:1970:LB

[Sto70h] Wilhelm Stoll. Level bumps. Lecture Notes in Mathematics, 135:101–113, 1970. CODEN LNMAA2. ISBN 3-540-04924-X (print), 3-540-36287-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059126/.

Stoll:1970:PSP

[Sto70i] Wilhelm Stoll. Properties of singular potentials. Lecture Notes in Mathematics, 135:64–80, 1970. CODEN LNMAA2. ISBN 3-540-04924-X (print), 3-540-36287-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059124/.

Stoll:1970:VDHa

[Sto70j] Wilhelm Stoll. Value distribution of holomorphic maps. Lecture Notes in Mathematics, 155:165–190, 1970. CODEN LN-MAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060324/.

Stoll:1970:VDHb

[Sto70k] Wilhelm Stoll. Value Distribution of Holomorphic Maps into Compact Complex Manifolds, volume 135 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04924-X (print), 3-540-36287-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). i + 267 pp. LCCN QA3 .L28 no. 135. URL http:/

/link.springer.com/book/10.1007/BFb0059118;   http://
www.springerlink.com/content/978-3-540-36287-6.

Stoll:1971:FIS

[Sto71] Wilhelm Stoll. Fiber integration and some applications. Lecture Notes in Mathematics, 184:109–120, 1971. CODEN LN-MAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059281/.

Stone:1972:CSC

[Sto72a] David A. Stone. Classifying stratifications and Cobordism theories. Lecture Notes in Mathematics, 252:98–132, 1972. CODEN LNMAA2. ISBN 3-540-05726-9 (print), 3-540-37112-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058589/.

Stone:1972:D

[Sto72b] David A. Stone. Definitions. Lecture Notes in Mathematics, 252: 1–42, 1972. CODEN LNMAA2. ISBN 3-540-05726-9 (print), 3-540-37112-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058585/.

Stone:1972:E

[Sto72c] David A. Stone. Existence. Lecture Notes in Mathematics, 252: 43–55, 1972. CODEN LNMAA2. ISBN 3-540-05726-9 (print), 3-540-37112-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058586/.

Stone:1972:OBB

[Sto72d] David A. Stone. Obstructions to block bundles. Lecture Notes in Mathematics, 252:133–151, 1972. CODEN LNMAA2. ISBN 3-540-05726-9 (print), 3-540-37112-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058590/.

Stone:1972:OQ

[Sto72e] David A. Stone. Open questions. Lecture Notes in Mathematics, 252:152–158, 1972. CODEN LNMAA2. ISBN 3-540-05726-9 (print), 3-540-37112-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058591/.

Stone:1972:SP

[Sto72f] David A. Stone. Stratified Polyhedra, volume 252 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-05726-9 (print), 3-540-37112-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 193 pp. LCCN QA3 .L28 no. 252. URL http:/

/link.springer.com/book/10.1007/BFb0058584;   http://
www.springerlink.com/content/978-3-540-37112-0.

Stone:1972:ST

[Sto72g] David A. Stone. Symmetry of transversality. Lecture Notes in Mathematics, 252:159–178, 1972. CODEN LNMAA2. ISBN 3-540-05726-9 (print), 3-540-37112-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058592/.

Stone:1972:T

[Sto72h] David A. Stone. Transversality. Lecture Notes in Mathematics, 252:79–97, 1972. CODEN LNMAA2. ISBN 3-540-05726-9 (print), 3-540-37112-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058588/.

Stone:1972:U

[Sto72i] David A. Stone. Uniqueness. Lecture Notes in Mathematics, 252: 56–78, 1972. CODEN LNMAA2. ISBN 3-540-05726-9 (print), 3-540-37112-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058587/.

Storrer:1972:GPD

[Sto72j] Hans H. Storrer. On Goldman’s primary decomposition. Lecture Notes in Mathematics, 246:617–661, 1972. CODEN LN-MAA2. ISBN 3-540-05760-9 (print), 3-540-37181-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059571/.

Stoer:1974:EMS

[Sto74] Josef Stoer. Extrapolation methods for the solution of initial value problems and their practical realization. Lecture Notes in Mathematics, 362:1–21, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066583/.

Strauss:1971:GGR

[Str71] Aaron Strauss. Global growth rates for solutions of certain perturbed differential systems. Lecture Notes in Mathematics, 235: 129–133, 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059205/.

Street:1972:SFS

[Str72] Anne Penfold Street. Sum-free sets. Lecture Notes in Mathematics, 292:124–271, 1972. CODEN LNMAA2. ISBN 3-540-06035-9 (print), 3-540-37994-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069910/.

Strang:1974:DPP

[Str74a] Gilbert Strang. The dimension of piecewise polynomial spaces, and one-sided approximation. Lecture Notes in Mathematics, 363: 144–152, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069132/.

Street:1974:EWM

[Str74b] Anne Penfold Street. Eulerian washing machines. Lecture Notes in Mathematics, 403:105–108, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057383/.

Street:1974:EC

[Str74c] Ross Street. Elementary cosmoi I. Lecture Notes in Mathematics, 420:134–180, 1974. CODEN LNMAA2. ISBN 3-540-06966-6 (print), 3-540-37270-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0063103/.

Street:1974:FYL

[Str74d] Ross Street. Fibrations and Yoneda’s lemma in a 2-category. Lecture Notes in Mathematics, 420:104–133, 1974. CODEN LN-MAA2. ISBN 3-540-06966-6 (print), 3-540-37270-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0063102/.

Stummel:1972:SPE

[Stu72] F. Stummel. Singular perturbations of elliptic sesquilinear forms. Lecture Notes in Mathematics, 280:155–180, 1972. CODEN LN-MAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066926/.

Stummel:1974:DML

[Stu74] F. Stummel. Difference methods for linear initial value problems. Lecture Notes in Mathematics, 395:123–135, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060667/.

Stys:1974:MPS

[Sty74] T. Styś. The maximum principle for the system of nonlinear differential equations of the Monge–Ampère type. Lecture Notes in Mathematics, 395:235–243, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060675/.

Subbarao:1972:SAC

[Sub72] M. V. Subbarao. On some arithmetic convolutions. Lecture Notes in Mathematics, 251:247–271, 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058796/.

Sucheston:1970:CET

[Suc70] Louis Sucheston, editor. Contributions to Ergodic Theory and Probability: Proceedings of the First Midwestern Conference on Ergodic Theory held at the Ohio State University, March 27–30, 1970, volume 160 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 160. URL http:/

/link.springer.com/book/10.1007/BFb0060639;   http://
www.springerlink.com/content/978-3-540-36371-2.

Sugiyama:1971:DIT

[Sug71] Shohei Sugiyama. Difference inequalities and their applications to stability problems. Lecture Notes in Mathematics, 243:1– 15, 1971. CODEN LNMAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058714/.

Sullivan:1971:CIA

[Sul71a] D. Sullivan. Combinatorial invariants of analytic spaces. Lecture Notes in Mathematics, 192:165–177, 1971. CODEN LN-MAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066822/.

Sullivan:1971:SS

[Sul71b] D. Sullivan. Singularities in spaces. Lecture Notes in Mathematics, 209:196–206, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068904/.

Sullivan:1971:GPI

[Sul71c] Dennis Sullivan. Geometric periodicity and the invariants of manifolds. Lecture Notes in Mathematics, 197:44–75, 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068611/.

Summers:1974:BCW

[Sum74] W. H. Summers. The bounded case of the weighted approximation problem. Lecture Notes in Mathematics, 384:177–183, 1974. CODEN LNMAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0057548/.

Sundaresan:1970:EPC

[Sun70a] Kondagunta Sundaresan. Extreme points of convex sets and selection theorems. Lecture Notes in Mathematics, 171:82–87, 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069726/.

Sunyach:1970:CTD

[Sun70b] Ch. Sunyach. Complétude et théorème du graphe fermé. (French) []. Lecture Notes in Mathematics, 116:168–174, 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059254/.

Suryanarayana:1972:NBU

[Sur72] D. Suryanarayana. The number of bi-unitary divisors of an integer. Lecture Notes in Mathematics, 251:273–282, 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058797/.

Stroock:1971:DPMa

[SV71a] Daniel W. Stroock and S. R. S. Varadhan. Diffusion processes and martingales I. Lecture Notes in Mathematics, 190:60–66, 1971. CODEN LNMAA2. ISBN 3-540-05396-4 (print), 3-540-36515-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065893/.

Stroock:1971:DPMb

[SV71b] Daniel W. Stroock and S. R. S. Varadhan. Diffusion processes and martingales II. Lecture Notes in Mathematics, 190:67–75, 1971. CODEN LNMAA2. ISBN 3-540-05396-4 (print), 3-540-36515-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065894/.

Street:1974:SFS

[SW74a] Anne Penfold Street and Earl Glen Whitehead, Jr. Sum-free sets, difference sets and cyclotomy. Lecture Notes in Mathematics, 403:109–124, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057384/.

Swartz:1974:CEC

[SW74b] Blair Swartz and Burton Wendroff. The comparative efficiency of certain finite element and finite difference methods for a hyperbolic problem. Lecture Notes in Mathematics, 363:153–163, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069133/.

Swan:1970:TFG

[Swa70] Richard G. Swan. K-Theory of Finite Groups and Orders, volume 149 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04938-X (print), 3-540-36312-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 237 pp. LCCN QA3 .L28 no. 149. URL http:

//link.springer.com/book/10.1007/BFb0059150;   http:/
/www.springerlink.com/content/978-3-540-36312-5.

Swan:1973:TAC

[Swa73] Richard G. Swan. K-theory and algebraic correspondences. Lecture Notes in Mathematics, 353:161–179, 1973. CODEN LN-MAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059268/.

Swierczkowski:1970:THC

[Świ70] S. Świerczkowski. Triads in the homology of categories. Lecture Notes in Mathematics, 137:114–125, 1970. CODEN LN-MAA2. ISBN 3-540-04926-6 (print), 3-540-36292-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060442/.

Szekeres:1974:PDT

[Sze74] G. Szekeres. Polyhedral decomposition of trivalent graphs. Lecture Notes in Mathematics, 403:125–127, 1974. CODEN LN-MAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057385/.

Szpiro:1973:TKK

[Szp73] Lucien Szpiro. Travaux de Kempf, Kleiman, Laksov sur les diviseurs exceptionnels. (French) []. Lecture Notes in Mathematics, 317:339–353, 1973. CODEN LNMAA2. ISBN 3-540-06179-7 (print), 3-540-38403-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069290/.

Taam:1970:LMAa

[Taa70a] C. T. Taam. Lectures in modern analysis and applications II, volume 140 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04929-0 (print), 3-540-36298-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 119 pp. LCCN QA3 .L28 no. 140. URL http:

//link.springer.com/book/10.1007/BFb0100062;   http:/
/www.springerlink.com/content/978-3-540-36298-2.

Taam:1970:LMAb

[Taa70b] C. T. Taam. Lectures in Modern Analysis and Applications III, volume 170 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05284-4 (print), 3-540-36417-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 213 pp. LCCN QA3 .L28 no. 170. URL http:

//link.springer.com/book/10.1007/BFb0079063;   http:/
/www.springerlink.com/content/978-3-540-36417-7.

Tachikawa:1973:NLQ

[Tac73a] Hiroyuki Tachikawa. Are Noetherian, left QF-3 rings artinian, so QF-3? Lecture Notes in Mathematics, 351:72–90, 1973. CODEN LNMAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060003/.

Tachikawa:1973:CAN

[Tac73b] Hiroyuki Tachikawa. Conjectures arising from Nakayama’s conjecture. Lecture Notes in Mathematics, 351:105–119, 1973. CO-DEN LNMAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060005/.

Tachikawa:1973:DD

[Tac73c] Hiroyuki Tachikawa. Dominant dimension. Lecture Notes in Mathematics, 351:91–104, 1973. CODEN LNMAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060004/.

Tachikawa:1973:DCC

[Tac73d] Hiroyuki Tachikawa. The double centralizer condition. Lecture Notes in Mathematics, 351:131–144, 1973. CODEN LN-MAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060007/.

Tachikawa:1973:LQR

[Tac73e] Hiroyuki Tachikawa. Left QF-3 rings. Lecture Notes in Mathematics, 351:40–53, 1973. CODEN LNMAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060001/.

Tachikawa:1973:MD

[Tac73f] Hiroyuki Tachikawa. Morita duality. Lecture Notes in Mathematics, 351:22–40, 1973. CODEN LNMAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060000/.

Tachikawa:1973:P

[Tac73g] Hiroyuki Tachikawa. Preliminaries. Lecture Notes in Mathematics, 351:1–5, 1973. CODEN LNMAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059998/.

Tachikawa:1973:QR

[Tac73h] Hiroyuki Tachikawa. QF-13 rings. Lecture Notes in Mathematics, 351:144–155, 1973. CODEN LNMAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060008/.

Tachikawa:1973:QRC

[Tac73i] Hiroyuki Tachikawa. QF-3 rings and categories of projective modules. Lecture Notes in Mathematics, 351:155–167, 1973. CODEN LNMAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060009/.

Tachikawa:1973:QRR

[Tac73j] Hiroyuki Tachikawa. QF-3 rings and rings of finite representation type. Lecture Notes in Mathematics, 351:120–130, 1973. CODEN LNMAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060006/.

Tachikawa:1973:QFR

[Tac73k] Hiroyuki Tachikawa. Quasi-Frobenius Rings and Generalizations QF-3 and QF-1 Rings, volume 351 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xi + 172 pp. LCCN QA3 .L28 no. 351.

URL http://link.springer.com/book/10.1007/BFb0059997;
 http://www.springerlink.com/content/978-3-540-37812-

9. Notes by Claus Michael Ringel.

Tachikawa:1973:SPR

[Tac73l] Hiroyuki Tachikawa. Semi-perfect rings. Lecture Notes in Mathematics, 351:5–22, 1973. CODEN LNMAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059999/.

Tachikawa:1973:SQR

[Tac73m] Hiroyuki Tachikawa. The structure of QF-3 rings. Lecture Notes in Mathematics, 351:53–71, 1973. CODEN LNMAA2. ISBN 3-540-06501-6 (print), 3-540-37812-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060002/.

Takesaki:1970:CTM

[Tak70a] M. Takesaki. The commutation theorem for modular Hilbert algebras. Lecture Notes in Mathematics, 128:20–23, 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065837/.

Takesaki:1970:FMH

[Tak70b] M. Takesaki. Formulation of the modular Hilbert algebra. Lecture Notes in Mathematics, 128:52–57, 1970. CODEN LN-MAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065843/.

Takesaki:1970:GHA

[Tak70c] M. Takesaki. Generalized Hilbert algebras. Lecture Notes in Mathematics, 128:8–19, 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065836/.

Takesaki:1970:I

[Tak70d] M. Takesaki. Introduction. Lecture Notes in Mathematics, 128: 1–4, 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065833/.

Takesaki:1970:MAG

[Tak70e] M. Takesaki. The modular automorphism group and the Kubo– Martin–Schwinger boundary condition. Lecture Notes in Mathematics, 128:63–72, 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065846/.

Takesaki:1970:MHA

[Tak70f] M. Takesaki. Modular Hilbert algebras. Lecture Notes in Mathematics, 128:5–8, 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065835/.

Takesaki:1970:MO

[Tak70g] M. Takesaki. The modular operator Δ. Lecture Notes in Mathematics, 128:31–35, 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065840/.

Takesaki:1970:OPA

[Tak70h] M. Takesaki. The one-parameter automorphisms defined by the modular operator Δ. Lecture Notes in Mathematics, 128:44– 52, 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065842/.

Takesaki:1970:P

[Tak70i] M. Takesaki. Preliminaries. Lecture Notes in Mathematics, 128: 4–5, 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065834.

Takesaki:1970:RMO

[Tak70j] M. Takesaki. The resolvent of the modular operator Δ. Lecture Notes in Mathematics, 128:35–44, 1970. CODEN LN-MAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065841/.

Takesaki:1970:SAS

[Tak70k] M. Takesaki. Self-adjoint subalgebras of generalized Hilbert algebras. Lecture Notes in Mathematics, 128:23–25, 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065838/.

Takesaki:1970:SFM

[Tak70l] M. Takesaki. Semi-finiteness and the modular automorphism group. Lecture Notes in Mathematics, 128:73–120, 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065847/.

Takesaki:1970:SA

[Tak70m] M. Takesaki. The spectral algebra. Lecture Notes in Mathematics, 128:25–31, 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065839/.

Takesaki:1970:SRN

[Tak70n] M. Takesaki. The standard representation of von Neumann algebras. Lecture Notes in Mathematics, 128:60–63, 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065845/.

Takesaki:1970:TPD

[Tak70o] M. Takesaki. Tensor product and direct sum of modular Hilbert algebras. Lecture Notes in Mathematics, 128:57–60, 1970. CO-DEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0065844/.

Takesaki:1970:TTM

[Tak70p] M. Takesaki. Tomita’s Theory of Modular Hilbert Algebras and its Applications, volume 128 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04917-7 (print), 3-540-36267-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ii + 123 pp. LCCN QA3 .L28 no. 128.

URL http://link.springer.com/book/10.1007/BFb0065832;
 http://www.springerlink.com/content/978-3-540-36267-
8.

Takens:1971:S

[Tak71a] F. Takens. A solution. Lecture Notes in Mathematics, 197: 231, 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068623.

Takens:1971:ZTS

[Tak71b] Floris Takens. On Zeeman’s tolerance stability conjecture. Lecture Notes in Mathematics, 197:209–219, 1971. CODEN LN-MAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068621/.

Takens:1971:PCM

[Tak71c] Floris Takens. Problems concerning manifolds. Lecture Notes in Mathematics, 197:220–230, 1971. CODEN LNMAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068622/.

Takesaki:1972:DNA

[Tak72] Masamichi Takesaki. Duality and von Neumann algebras. Lecture Notes in Mathematics, 247:665–786, 1972. CODEN LN-MAA2. ISBN 3-540-05729-3 (print), 3-540-37117-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058558/.

Tate:1971:CIV

[Tat71] John Tate. Classes d’isogénie des variétés abéliennes sur un corps fini (d’après T. Honda). (French) []. Lecture Notes in Mathematics, 179:95–110, 1971. CODEN LNMAA2. ISBN 3-540-05356-5 (print), 3-540-36451-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058807/.

Taubert:1974:DGA

[Tau74] Klaus Taubert. Differenzenverfahren für gewöhnliche Anfangswertaufgaben mit unstetiger rechter seite. (German) [Difference methods for ordinary initial value problems with discontinuous right-hand sides]. Lecture Notes in Mathematics, 395: 137–148, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060668/.

Taylor:1973:SGI

[Tay73] Lawrence R. Taylor. Surgery groups and inner automorphisms. Lecture Notes in Mathematics, 343:471–477, 1973. CODEN LN-MAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061375/.

Taylor:1974:GBD

[Tay74a] D. E. Taylor. Graphs and block designs associated with the threedimensional unitary groups. Lecture Notes in Mathematics, 403: 128–131, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057386/.

Taylor:1974:EHO

[Tay74b] Michael Taylor. Elliptic and hypoelliptic operators. Lecture Notes in Mathematics, 416:45–57, 1974. CODEN LNMAA2. ISBN 3-540-06961-5 (print), 3-540-37266-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101250/.

Taylor:1974:EBV

[Tay74c] Michael Taylor. Elliptic boundary value problems; Petrowsky parabolic operators. Lecture Notes in Mathematics, 416:84– 119, 1974. CODEN LNMAA2. ISBN 3-540-06961-5 (print), 3-540-37266-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0101252/.

Taylor:1974:IVP

[Tay74d] Michael Taylor. The initial value problem. Hyperbolic operators. Lecture Notes in Mathematics, 416:58–83, 1974. CODEN LNMAA2. ISBN 3-540-06961-5 (print), 3-540-37266-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101251/.

Taylor:1974:LPD

[Tay74e] Michael Taylor. Lectures on pseudo differential operators. Lecture Notes in Mathematics, 416:1–3, 1974. CODEN LNMAA2. ISBN 3-540-06961-5 (print), 3-540-37266-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101247/.

Taylor:1974:PSW

[Tay74f] Michael Taylor. Propagation of singularities; wave front sets. Lecture Notes in Mathematics, 416:120–138, 1974. CODEN LN-MAA2. ISBN 3-540-06961-5 (print), 3-540-37266-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0101253/.

Taylor:1974:PDOa

[Tay74g] Michael Taylor. Pseudo differential operators. Lecture Notes in Mathematics, 416:19–44, 1974. CODEN LNMAA2. ISBN 3-540-06961-5 (print), 3-540-37266-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101249/.

Taylor:1974:PDOb

[Tay74h] Michael Taylor. Pseudo Differential Operators, volume 416 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06961-5 (print), 3-540-37266-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 155 pp. LCCN QA3 .L28 no. 416. URL http:/

/link.springer.com/book/10.1007/BFb0101246;   http://
www.springerlink.com/content/978-3-540-37266-0.

Taylor:1974:SGI

[Tay74i] Michael Taylor. The sharp Gårding inequality. Lecture Notes in Mathematics, 416:139–149, 1974. CODEN LNMAA2. ISBN 3-540-06961-5 (print), 3-540-37266-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101254/.

Taylor:1974:SIO

[Tay74j] Michael Taylor. Singular integral operators on the circle. Lecture Notes in Mathematics, 416:4–18, 1974. CODEN LNMAA2. ISBN 3-540-06961-5 (print), 3-540-37266-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0101248/.

tomDieck:1972:BAE

[tD72a] Tammo tom Dieck. Bemerkungen Über Äquivariante Eulerklassen. (German) []. Lecture Notes in Mathematics, 298:152– 162, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070038/.

tomDieck:1972:EFP

[tD72b] Tammo tom Dieck. Existence of fixed points. Lecture Notes in Mathematics, 298:163–169, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070039/.

tomDieck:1970:CGC

[tDKP70a] Tammo tom Dieck, Klaus Heiner Kamps, and Dieter Puppe. Cofaserungen. (German) [Cofibrations]. Lecture Notes in Mathematics, 157:20–85, 1970. CODEN LNMAA2. ISBN 3-540-05185-6 (print), 3-540-36346-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059723/.

tomDieck:1970:FGF

[tDKP70b] Tammo tom Dieck, Klaus Heiner Kamps, and Dieter Puppe. Faserungen. (German) [Fibrations]. Lecture Notes in Mathematics, 157:86–158, 1970. CODEN LNMAA2. ISBN 3-540-05185-6 (print), 3-540-36346-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059724/.

tomDieck:1970:HHG

[tDKP70c] Tammo tom Dieck, Klaus Heiner Kamps, and Dieter Puppe. Homotopiemengen und Homotopiegruppen. (German) []. Lecture Notes in Mathematics, 157:159–254, 1970. CODEN LN-MAA2. ISBN 3-540-05185-6 (print), 3-540-36346-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059725/.

tomDieck:1970:HGH

[tDKP70d] Tammo tom Dieck, Klaus Heiner Kamps, and Dieter Puppe. Homotopietheorie. (German) [Homotopy theory], volume 157 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-05185-6 (print), 3-540-36346-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 265 pp. LCCN QA3 .L28 no. 157. URL http:/

/link.springer.com/book/10.1007/BFb0059721;   http://
www.springerlink.com/content/978-3-540-36346-0.

tomDieck:1970:KGG

[tDKP70e] Tammo tom Dieck, Klaus Heiner Kamps, and Dieter Puppe. Kategorientheoretische Grundlagen. Grundlagen der Homotopietheorie. (German) []. Lecture Notes in Mathematics, 157:1–19, 1970. CODEN LNMAA2. ISBN 3-540-05185-6 (print), 3-540-36346-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059722/.

Teleman:1971:THA

[Tel71] Silviu Teleman. Theory of harmonic algebras with applications to von Neumann algebras and cohomology of locally compact spaces (de Rham’s theorem). Lecture Notes in Mathematics, 248: 99–315, 1971. CODEN LNMAA2. ISBN 3-540-05714-5 (print), 3-540-37081-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058564/.

Temam:1971:AED

[Tem71a] Roger Temam. Approximation d’équations aux dérivées partielles par des méthodes de décomposition. (French) []. Lecture Notes in Mathematics, 180:295–303, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058838/.

Temam:1971:TPN

[Tem71b] Roger Temam. Table par noms d’auteurs [Séminaire Bourbaki, 1967/68 à 1969/70, Exposés 331 à 381]. (French) []. Lecture Notes in Mathematics, 180:305–310, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058839/.

Terrell:1970:LET

[Ter70] Thomas R. Terrell. Local ergodic theorems for N-parameter semigroups of operators. Lecture Notes in Mathematics, 160: 262–278, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060658/.

Thimm:1970:ECA

[Thi70] Walter Thimm. Extension of coherent analytic subsheaves. Lecture Notes in Mathematics, 155:191–202, 1970. CODEN LN-MAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060325/.

Thomas:1970:TMA

[Tho70] Emery Thomas. The Thom–Massey approach to embeddings. Lecture Notes in Mathematics, 168:283–307, 1970. CODEN LN-MAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058529/.

Thom:1971:BSS

[Tho71a] R. Thom. The bifurcation subset of a space of maps. Lecture Notes in Mathematics, 197:202–208, 1971. CODEN LN-MAA2. ISBN 3-540-05467-7 (print), 3-540-36653-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068620/.

Thom:1971:SSM

[Tho71b] R. Thom. Stratified sets and morphisms: Local models. Lecture Notes in Mathematics, 192:153–164, 1971. CODEN LN-MAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066821/.

Thomas:1971:SRV

[Tho71c] Emery Thomas. Some remarks on vector fields. Lecture Notes in Mathematics, 196:107–113, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067753/.

Thomee:1971:TST

[Tho71d] Vidar Thomée. Topics in stability theory for partial difference operators. Lecture Notes in Mathematics, 193:41–91, 1971. CODEN LNMAA2. ISBN 3-540-05422-7 (print), 3-540-36538-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060343/.

Thomee:1973:CES

[Tho73a] Vidar Thomée. Convergence estimates for semi-discrete Galerkin methods for initial-value problems. Lecture Notes in Mathematics, 333:243–262, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060701/.

Thompson:1973:SNL

[Tho73b] Colin J. Thompson. Some non-linear problems in statistical mechanics and biology. Lecture Notes in Mathematics, 322:308– 342, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060574/.

Thompson:1973:DAS

[Tho73c] Robert J. Thompson. Difference approximations for some functional differential equations. Lecture Notes in Mathematics, 333: 263–273, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060702/.

Thom:1974:CLV

[Tho74a] R. Thom. Sur le cut locus d’une variété plongée (résumé). (French) []. Lecture Notes in Mathematics, 392:97, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064116.

Thomee:1974:CAF

[Tho74b] V. Thomée. Convergence analysis of a finite difference scheme for a simple semi-linear hyperbolic equation. Lecture Notes in Mathematics, 395:149–166, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060669/.

Thomee:1974:SGM

[Tho74c] Vidar Thomée. Spline-Galerkin methods for initial-value problems with constant coefficients. Lecture Notes in Mathematics, 363:164–175, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069134/.

Tierney:1972:STC

[Tie72] Myles Tierney. Sheaf theory and the continuum hypothesis. Lecture Notes in Mathematics, 274:13–42, 1972. CODEN LN-MAA2. ISBN 3-540-05920-2 (print), 3-540-37609-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0073963/.

Timourian:1971:MDC

[Tim71] J. G. Timourian. Maps with 0-dimensional critical set. Lecture Notes in Mathematics, 192:282–285, 1971. CODEN LN-MAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066828/.

Tits:1971:GFS

[Tit71a] Jacques Tits. Groupes finis simples sporadiques. (French) []. Lecture Notes in Mathematics, 180:187–211, 1971. CODEN LN-MAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058832/.

Titus:1971:PCL

[Tit71b] C. J. Titus. A proof of a conjecture of Loewner and of the Carathéodory conjecture concerning umbilic points. Lecture Notes in Mathematics, 209:114–117, 1971. CODEN LN-MAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068896/.

Tits:1974:AS

[Tit74a] Jacques Tits. Appendix 1. Shadows. Lecture Notes in Mathematics, 386:232–248, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-38349-9_12/.

Tits:1974:AGR

[Tit74b] Jacques Tits. Appendix 2. Generators and relations. Lecture Notes in Mathematics, 386:249–273, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38349-9_13/.

Tits:1974:BM

[Tit74c] Jacques Tits. Back matter. Lecture Notes in Mathematics, 386: 274–302, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-38349-9/1.pdf.

Tits:1974:BSS

[Tit74d] Jacques Tits. The building of a semi-simple algebraic group. Lecture Notes in Mathematics, 386:75–91, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.

springer.com/chapter/10.1007/978-3-540-38349-9_5/.

Tits:1974:B

[Tit74e] Jacques Tits. Buildings. Lecture Notes in Mathematics, 386: 38–59, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38349-9_3/.

Tits:1974:BST

[Tit74f] Jacques Tits. Buildings of Spherical Type and Finite BN-Pairs, volume 386 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). x + 299 pp. LCCN QA3 .L28 no. 386. URL http://link.

springer.com/book/10.1007/978-3-540-38349-9;   http:/
/www.springerlink.com/content/978-3-540-38349-9.

Tits:1974:BTa

[Tit74g] Jacques Tits. Buildings of type A_{n}, D_{n}, E_{n}. Lecture Notes in Mathematics, 386:92–101, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38349-9_6/.

Tits:1974:BTP

[Tit74h] Jacques Tits. Buildings of type C_{n}. I. Polar spaces. Lecture Notes in Mathematics, 386:102–116, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-38349-9_7/.

Tits:1974:BTIa

[Tit74i] Jacques Tits. Buildings of type C_{n}. II. Projective embeddings of polar spaces. Lecture Notes in Mathematics, 386:117–176, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38349-9_8/.

Tits:1974:BTIb

[Tit74j] Jacques Tits. Buildings of type C_{n}. III. Non-embeddable polar spaces. Lecture Notes in Mathematics, 386:177–199, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38349-9_9/.

Tits:1974:BTb

[Tit74k] Jacques Tits. Buildings of type F_{4}. Lecture Notes in Mathematics, 386:200–217, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-38349-9_10/.

Tits:1974:C

[Tit74l] Jacques Tits. Complexes. Lecture Notes in Mathematics, 386: 1–12, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38349-9_1/.

Tits:1974:CC

[Tit74m] Jacques Tits. Coxeter complexes. Lecture Notes in Mathematics, 386:13–37, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-38349-9_2/.

Tits:1974:FBP

[Tit74n] Jacques Tits. Finite BN-pairs of irreducible type and rank ≥ 3. Lecture Notes in Mathematics, 386:218–231, 1974. CODEN LN-MAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http:/

/link.springer.com/chapter/10.1007/978-3-540-38349-9_
11/.

Tits:1974:FM

[Tit74o] Jacques Tits. Front matter. Lecture Notes in Mathematics, 386: i–x, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bfm:
978-3-540-38349-9/1.pdf.

Tits:1974:R

[Tit74p] Jacques Tits. Reduction. Lecture Notes in Mathematics, 386: 60–74, 1974. CODEN LNMAA2. ISBN 3-540-06757-4 (print), 3-540-38349-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-38349-9_4/.

Toda:1971:DLG

[Tod71] Hirosi Toda. modp decomposition of Lie groups. Lecture Notes in Mathematics, 196:20–24, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067741/.

Topsoe:1970:BRCa

[Top70a] Flemming Topsøe. Basic result on construction of a measure. Lecture Notes in Mathematics, 133:3–6, 1970. CODEN LN-MAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069483/.

Topsoe:1970:BRCb

[Top70b] Flemming Topsøe. Basic result on construction of an integral. Lecture Notes in Mathematics, 133:6–15, 1970. CODEN LN-MAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069484/.

Topsoe:1970:CWT

[Top70c] Flemming Topsøe. Compactness in the weak topology. Lecture Notes in Mathematics, 133:42–45, 1970. CODEN LN-MAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069490/.

Topsoe:1970:CMA

[Top70d] Flemming Topsøe. Construction of measures by approximation from outside and by approximation from inside. Lecture Notes in Mathematics, 133:26–31, 1970. CODEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069487/.

Topsoe:1970:CWC

[Top70e] Flemming Topsøe. Criteria for weak convergence. Lecture Notes in Mathematics, 133:45–47, 1970. CODEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069491/.

Topsoe:1970:DBP

[Top70f] Flemming Topsøe. Definition and basic properties of the weak topology. Lecture Notes in Mathematics, 133:40–42, 1970. CO-DEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069489/.

Topsoe:1970:FAT

[Top70g] Flemming Topsøe. Finitely additive theory. Lecture Notes in Mathematics, 133:15–21, 1970. CODEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069485/.

Topsoe:1970:FSP

[Top70h] Flemming Topsøe. First solution of the ξ-problem. Lecture Notes in Mathematics, 133:54–60, 1970. CODEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069494/.

Topsoe:1970:BMB

[Top70i] Flemming Topsøe. From “Baire” measures to “Borel” measures, an abstract approach. Lecture Notes in Mathematics, 133:21– 26, 1970. CODEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069486/.

Topsoe:1970:JC

[Top70j] Flemming Topsøe. Joint continuity. Lecture Notes in Mathematics, 133:66–68, 1970. CODEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069497/.

Topsoe:1970:MID

[Top70k] Flemming Topsøe. Measure and integral, definitions. Lecture Notes in Mathematics, 133:1–3, 1970. CODEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069482/.

Topsoe:1970:PPS

[Top70l] Flemming Topsøe. On the possibility of providing a space of measures with a vague topology. Lecture Notes in Mathematics, 133:31–39, 1970. CODEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069488/.

Topsoe:1970:S

[Top70m] Flemming Topsøe. On the structure of M_{+}(X). Lecture Notes in Mathematics, 133:47–51, 1970. CODEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069492/.

Topsoe:1970:PWC

[Top70n] Flemming Topsøe. Preservation of weak convergence. Lecture Notes in Mathematics, 133:68–71, 1970. CODEN LN-MAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069498/.

Topsoe:1970:PRQ

[Top70o] Flemming Topsøe. A problem related to questions of uniformity. Lecture Notes in Mathematics, 133:51–54, 1970. CODEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069493/.

Topsoe:1970:SSP

[Top70p] Flemming Topsøe. Second solution of the ξ-problem. Lecture Notes in Mathematics, 133:60–64, 1970. CODEN LN-MAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069495/.

Topsoe:1970:TM

[Top70q] Flemming Topsøe. Topology and Measure, volume 133 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 79 pp. LCCN QA3 .L28 no. 133. URL http:/

/link.springer.com/book/10.1007/BFb0069481;   http://
www.springerlink.com/content/978-3-540-36284-5.

Topsoe:1970:UC

[Top70r] Flemming Topsøe. Uniformity classes. Lecture Notes in Mathematics, 133:64–66, 1970. CODEN LNMAA2. ISBN 3-540-04922-3 (print), 3-540-36284-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069496/.

Tornig:1973:MDD

[Tör73] W. Törnig. Monotonieeigenschaften Von Diskretisierungen Des Dirichletproblems Quasilinearer Elliptischer Differentialgleichungen. (German) []. Lecture Notes in Mathematics, 333:274–291, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060703/.

Tounissoux:1974:APR

[Tou74] Daniel Tounissoux. Application of properties with regard to the order of questions. Lecture Notes in Mathematics, 398:194– 201, 1974. CODEN LNMAA2. ISBN 3-540-06844-9 (print), 3-540-37798-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065778/.

Tsokos:1971:ASR

[TP71a] Chris P. Tsokos and W. J. Padgett. Approximate solutions of the random Volterra integral equation. Lecture Notes in Mathematics, 233:48–75, 1971. CODEN LNMAA2. ISBN 3-540-05660-2 (print), 3-540-36992-9 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0059963/.

Tsokos:1971:GI

[TP71b] Chris P. Tsokos and W. J. Padgett. General introduction. Lecture Notes in Mathematics, 233:1–7, 1971. CODEN LNMAA2. ISBN 3-540-05660-2 (print), 3-540-36992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059960/.

Tsokos:1971:P

[TP71c] Chris P. Tsokos and W. J. Padgett. Preliminaries. Lecture Notes in Mathematics, 233:8–26, 1971. CODEN LNMAA2. ISBN 3-540-05660-2 (print), 3-540-36992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059961/.

Tsokos:1971:RDF

[TP71d] Chris P. Tsokos and W. J. Padgett. Random discrete Fredholm and Volterra equations. Lecture Notes in Mathematics, 233:106– 129, 1971. CODEN LNMAA2. ISBN 3-540-05660-2 (print), 3-540-36992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059965/.

Tsokos:1971:RIEa

[TP71e] Chris P. Tsokos and W. J. Padgett. A random integral equation of the Volterra type. Lecture Notes in Mathematics, 233:

27–47, 1971. CODEN LNMAA2. ISBN 3-540-05660-2 (print), 3-540-36992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059962/.

Tsokos:1971:RIEb

[TP71f] Chris P. Tsokos and W. J. Padgett. Random Integral Equations with Applications to Stochastic Systems, volume 233 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05660-2 (print), 3-540-36992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 174 pp. LCCN QA3 .L28 no. 233. URL http:/

/link.springer.com/book/10.1007/BFb0059959;   http://
www.springerlink.com/content/978-3-540-36992-9.

Tsokos:1971:SDSa

[TP71g] Chris P. Tsokos and W. J. Padgett. The stochastic differential systems. Lecture Notes in Mathematics, 233:130–143, 1971. CODEN LNMAA2. ISBN 3-540-05660-2 (print), 3-540-36992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059966/.

Tsokos:1971:SDSb

[TP71h] Chris P. Tsokos and W. J. Padgett. The stochastic differential systems. Lecture Notes in Mathematics, 233:144–155, 1971. CODEN LNMAA2. ISBN 3-540-05660-2 (print), 3-540-36992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059967/.

Tsokos:1971:SDSc

[TP71i] Chris P. Tsokos and W. J. Padgett. The stochastic differential systems with lag time. Lecture Notes in Mathematics, 233: 156–165, 1971. CODEN LNMAA2. ISBN 3-540-05660-2 (print), 3-540-36992-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059968/.

Tsokos:1971:SIE

[TP71j] Chris P. Tsokos and W. J. Padgett. A stochastic integral equation of the Fredholm type with application to systems theory. Lecture Notes in Mathematics, 233:76–105, 1971. CODEN LN-MAA2. ISBN 3-540-05660-2 (print), 3-540-36992-9 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059964/.

Trang:1974:FML

[Trá74a] Lê Dûng Tráng. Finitude de la monodromie locale des courbes planes. (French) []. Lecture Notes in Mathematics, 409:341– 346, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068109/.

Trang:1974:CEF

[Trá74b] Lê Dûng Tráng. Sur un critère d’équisingularité. (French) [On a criterium of equisingularity]. Lecture Notes in Mathematics, 409: 124–160, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068096/.

Trang:1974:TDT

[Trá74c] Lê Dûng Tráng. Un théorème du type de Lefschetz. (French) []. Lecture Notes in Mathematics, 409:242–278, 1974. CODEN LNMAA2. ISBN 3-540-06856-2 (print), 3-540-37814-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068104/.

Trebels:1973:APE

[Tre73a] Walter Trebels. Applications to particular expansions. Lecture Notes in Mathematics, 329:74–93, 1973. CODEN LN-MAA2. ISBN 3-540-06357-9 (print), 3-540-46951-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060964/.

Trebels:1973:GT

[Tre73b] Walter Trebels. General theory. Lecture Notes in Mathematics, 329:9–17, 1973. CODEN LNMAA2. ISBN 3-540-06357-9 (print), 3-540-46951-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060961/.

Trebels:1973:I

[Tre73c] Walter Trebels. Introduction. Lecture Notes in Mathematics, 329:1–8, 1973. CODEN LNMAA2. ISBN 3-540-06357-9 (print), 3-540-46951-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060960/.

Trebels:1973:MCB

[Tre73d] Walter Trebels. Multiplier criteria for (C, α)-bounded expansions. Lecture Notes in Mathematics, 329:18–50, 1973. CODEN LNMAA2. ISBN 3-540-06357-9 (print), 3-540-46951-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060962/.

Trebels:1973:MBF

[Tre73e] Walter Trebels. Multipliers for (C, α)-Bounded Fourier Expansions in Banach Spaces and Approximation Theory, volume 329 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06357-9 (print), 3-540-46951-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 103 pp. LCCN QA3 .L28 no. 329. URL http:

//link.springer.com/book/10.1007/BFb0060959;   http:/
/www.springerlink.com/content/978-3-540-46951-3.

Trebels:1973:PSM

[Tre73f] Walter Trebels. Particular summation methods. Lecture Notes in Mathematics, 329:51–73, 1973. CODEN LNMAA2. ISBN 3-540-06357-9 (print), 3-540-46951-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060963/.

Treves:1974:UPF

[Trè74] F. Trèves. Usefulness of pseudodifferential and Fourier integral operators in the study of the local solvability of linear partial differential equations. Lecture Notes in Mathematics, 384:92– 108, 1974. CODEN LNMAA2. ISBN 3-540-06752-3 (print), 3-540-38342-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057541/.

Troelstra:1973:MII

[Tro73a] A. S. Troelstra. Metamathematical Investigation of Intuitionistic Arithmetic and Analysis, volume 344 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CO-DEN LNMAA2. ISBN 3-540-06491-5 (print), 3-540-37806-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xvii + 485 pp. LCCN QA3 .L28 no. 344. URL http:/

/link.springer.com/book/10.1007/BFb0066739;   http://
www.springerlink.com/content/978-3-540-37806-8.

Troelstra:1973:NIS

[Tro73b] A. S. Troelstra. Notes on intuitionistic second order arithmetic. Lecture Notes in Mathematics, 337:171–205, 1973. CODEN LN-MAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066774/.

Tulcea:1970:SMV

[Tul70a] A. Ionescu Tulcea. Super-mean-valued functions and semipolar sets. Lecture Notes in Mathematics, 160:71–74, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060648/.

Tulcea:1970:LDB

[Tul70b] C. Ionescu Tulcea. Liftings and derivation bases. Lecture Notes in Mathematics, 160:75–86, 1970. CODEN LNMAA2. ISBN 3-540-05188-0 (print), 3-540-36371-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060649/.

Turrittin:1971:SME

[Tur71] H. L. Turrittin. Stokes multipliers for the equation ie145-1. Lecture Notes in Mathematics, 183:145–157, 1971. CODEN LN-MAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060414/.

Turpin:1973:LOB

[Tur73a] Philippe Turpin. Linear operators between Orlicz spaces. Lecture Notes in Mathematics, 331:222–226, 1973. CODEN LN-MAA2. ISBN 3-540-06367-6 (print), 3-540-46977-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068234/.

Turrittin:1973:MME

[Tur73b] H. L. Turrittin. My mathematical expectations. Lecture Notes in Mathematics, 312:1–22, 1973. CODEN LNMAA2. ISBN 3-540-06146-0 (print), 3-540-38353-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060042/.

Turner:1974:SDG

[Tur74] Edward C. Turner. A survey of diffeomorphism groups. Lecture Notes in Mathematics, 428:200–218, 1974. CODEN LN-MAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070539/.

Tutte:1971:WW

[Tut71] W. T. Tutte. Wheels and whirls. Lecture Notes in Mathematics, 211:1–4, 1971. CODEN LNMAA2. ISBN 3-540-05545-2 (print), 3-540-36877-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061071/.

Ueno:1974:ICT

[Uen74] Kenji Ueno. Introduction to classification theory of algebraic varieties and compact complex spaces. Lecture Notes in Mathematics, 412:288–332, 1974. CODEN LNMAA2. ISBN 3-540-06951-8 (print), 3-540-37877-4 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066164/.

Ulmer:1971:LPL

[Ulm71] Friedrich Ulmer. Locally α-presentable and locally α-generated categories. Lecture Notes in Mathematics, 195:230–247, 1971. CODEN LNMAA2. ISBN 3-540-05442-1 (print), 3-540-36548-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0072314/.

Ura:1971:LDS

[Ura71a] Taro Ura. Local dynamical systems and their isomorphisms. Lecture Notes in Mathematics, 243:76–90, 1971. CODEN LN-MAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058720/.

Urabe:1971:GFP

[Ura71b] Minoru Urabe. Green functions of pseudoperiodic differential operators. Lecture Notes in Mathematics, 243:106–122, 1971. CODEN LNMAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058722/.

Urabe:1971:JUS

[Ura71c] Minoru Urabe, editor. Japan–United States Seminar on Ordinary Differential and Functional Equations: Held in Kyoto/Japan, September 6–11, 1971, volume 243 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 243. URL http:/

/link.springer.com/book/10.1007/BFb0058713;   http://
www.springerlink.com/content/978-3-540-37080-2.

Vaisala:1971:APQ

[Väi71a] Jussi Väisälä. The analytic properties of quasiconformal mappings. Lecture Notes in Mathematics, 229:106–126, 1971. CO-DEN LNMAA2. ISBN 3-540-05648-3 (print), 3-540-36937-6 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0061220/.

Vaisala:1971:BRA

[Väi71b] Jussi Väisälä. Background in real analysis. Lecture Notes in Mathematics, 229:81–105, 1971. CODEN LNMAA2. ISBN 3-540-05648-3 (print), 3-540-36937-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061219/.

Vaisala:1971:LDQ

[Väi71c] Jussi Väisälä. Lectures on n-Dimensional Quasiconformal Mappings, volume 229 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05648-3 (print), 3-540-36937-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). xiv + 144 pp. LCCN QA3 .L28 no. 229. URL http:/

/link.springer.com/book/10.1007/BFb0061216;   http://
www.springerlink.com/content/978-3-540-36937-0.

Vaisala:1971:MP

[Väi71d] Jussi Väisälä. Mapping problems. Lecture Notes in Mathematics, 229:127–139, 1971. CODEN LNMAA2. ISBN 3-540-05648-3 (print), 3-540-36937-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061221/.

Vaisala:1971:MCF

[Väi71e] Jussi Väisälä. The modulus of a curve family. Lecture Notes in Mathematics, 229:1–40, 1971. CODEN LNMAA2. ISBN 3-540-05648-3 (print), 3-540-36937-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061217/.

Vaisala:1971:QM

[Väi71f] Jussi Väisälä. Quasiconformal mappings. Lecture Notes in Mathematics, 229:41–80, 1971. CODEN LNMAA2. ISBN 3-540-05648-3 (print), 3-540-36937-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061218/.

VanOsdol:1971:SRC

[Van71] Donovan H. Van Osdol. Sheaves in regular categories. Lecture Notes in Mathematics, 236:223–239, 1971. CODEN LN-MAA2. ISBN 3-540-05678-5 (print), 3-540-36999-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058582/.

VanDijk:1972:HAR

[Van72] G. Van Dijk. Harmonic analysis on reductive p-adic groups (after Harish-Chandra [4 (c)]). Lecture Notes in Mathematics, 244: 89–106, 1972. CODEN LNMAA2. ISBN 3-540-05720-X (print), 3-540-37094-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058698/.

Vasconcelos:1973:RGD

[Vas73] Wolmer V. Vasconcelos. Rings of global dimension two. Lecture Notes in Mathematics, 311:243–251, 1973. CODEN LN-MAA2. ISBN 3-540-06140-1 (print), 3-540-38340-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068933/.

Vauthier:1973:CAT

[Vau73] J. Vauthier. Comportement asymptotique des transformées de Fourier de distributions à support compact. (French) []. Lecture Notes in Mathematics, 336:125–133, 1973. CODEN LN-MAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065793/.

Velez:1974:CPS

[VCLN74] C. E. Velez, P. J. Cefola, A. C. Long, and K. S. Nimitz. Calculation of precision satellite orbits with nonsingular elements (VOP formulation). Lecture Notes in Mathematics, 362:183–206, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066592/.

vanDalen:1973:LI

[vD73] D. van Dalen. Lectures on intuitionism. Lecture Notes in Mathematics, 337:1–94, 1973. CODEN LNMAA2. ISBN 3-540-05569-X (print), 3-540-36884-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0066771/.

vandeCraats:1973:CSA

[vdC73] Jan van de Craats. Compound successive approximations for nonlinear two point boundary value problems. Lecture Notes in Mathematics, 333:44–53, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060686/.

vanderHouwen:1973:OSM

[vdH73] P. J. van der Houwen. One-step methods with adaptive stability functions for the integration of differential equations. Lecture Notes in Mathematics, 333:164–174, 1973. CODEN LN-MAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060695/.

vanderKallen:1973:AS

[vdK73a] W. L. J. van der Kallen. The action Âd. Structure of g_{Z}^{∗}, g_{R}^{∗}. Lecture Notes in Mathematics, 356:22–38, 1973. CODEN LN-MAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060179/.

vanderKallen:1973:ALC

[vdK73b] W. L. J. van der Kallen. Admissible lattices and the category L_{V}. Σ-connected components. Lecture Notes in Mathematics, 356: 39–46, 1973. CODEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060180/.

vanderKallen:1973:C

[vdK73c] W. L. J. van der Kallen. Conventions. Lecture Notes in Mathematics, 356:1, 1973. CODEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060176.

vanderKallen:1973:DSE

[vdK73d] W. L. J. van der Kallen. Degenerate sums. The extension r: g′Z → gZ. Lecture Notes in Mathematics, 356:5–22, 1973. CODEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060178/.

vanderKallen:1973:EGGb

[vdK73e] W. L. J. van der Kallen. The existence of ϕ: G∗ → G. Lecture Notes in Mathematics, 356:65–97, 1973. CODEN LN-MAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060186/.

vanderKallen:1973:EGGa

[vdK73f] W. L. J. van der Kallen. The extension ϕ: G∗ → G. Lecture Notes in Mathematics, 356:54–59, 1973. CODEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060183/.

vanderKallen:1973:EM

[vdK73g] W. L. J. van der Kallen. Extensions of G by a G-module. Lecture Notes in Mathematics, 356:59–61, 1973. CODEN LN-MAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060184/.

vanderKallen:1973:IS

[vdK73h] W. L. J. van der Kallen. G-invariant [p]-structures. Lecture Notes in Mathematics, 356:51–54, 1973. CODEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060182/.

vanderKallen:1973:M

[vdK73i] W. L. J. van der Kallen. The G-module kerπ. Lecture Notes in Mathematics, 356:46–51, 1973. CODEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060181/.

vanderKallen:1973:GFG

[vdK73j] W. L. J. van der Kallen. The group functor G∗. Lecture Notes in Mathematics, 356:138–140, 1973. CODEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060190/.

vanderKallen:1973:HG

[vdK73k] W. L. J. van der Kallen. The Hochschild groups. Lecture Notes in Mathematics, 356:61–65, 1973. CODEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060185/.

vanderKallen:1973:ICE

[vdK73l] W. L. J. van der Kallen. Infinitesimally Central Extensions of Chevalley Groups, volume 356 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 140 (est.) pp. LCCN QA3 .L28 no. 356.

URL http://link.springer.com/book/10.1007/BFb0060175;
 http://www.springerlink.com/content/978-3-540-37857-

vanderKallen:1973:ROC

[vdK73m] W. L. J. van der Kallen. Relations in the open cell. Lecture Notes in Mathematics, 356:97–118, 1973. CODEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060187/.

vanderKallen:1973:RWG

[vdK73n] W. L. J. van der Kallen. Representatives in G∗ of the Weyl group. Lecture Notes in Mathematics, 356:119–124, 1973. CO-DEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060188/.

vanderKallen:1973:TGR

[vdK73o] W. L. J. van der Kallen. The theorem of generators and relations and its consequences. Lecture Notes in Mathematics, 356: 124–138, 1973. CODEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060189/.

vanderKallen:1973:UCE

[vdK73p] W. L. J. van der Kallen. Universal central extensions. Central trick. Lecture Notes in Mathematics, 356:2–5, 1973. CODEN LNMAA2. ISBN 3-540-06559-8 (print), 3-540-37857-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060177/.

Velte:1972:KEN

[Vel72] W. Velte. Über komplementäre Extremalprobleme bei nichtlinearen Randwertaufgaben. (German) [On complementary extremal problems in nonlinear boundary-value problems]. Lecture Notes in Mathematics, 267:323–335, 1972. CODEN LN-MAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061625/.

Vergne:1971:IER

[Ver71] Michèle Vergne. Sur les intégrales d’entrelacement de R. A. Kunze et E. M. Stein. (French) []. Lecture Notes in Mathematics, 180:87–106, 1971. CODEN LNMAA2. ISBN 3-540-05357-3 (print), 3-540-36452-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058826/.

Verdier:1972:CTF

[Ver72a] J. L. Verdier. Cohomologie dans les topos. (French) []. Lecture Notes in Mathematics, 270:1–82, 1972. CODEN LNMAA2. ISBN 3-540-06012-X (print), 3-540-37987-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061320/.

Verdier:1972:FCF

[Ver72b] J. L. Verdier. Fonctorialité des catégories de faisceaux. (French) []. Lecture Notes in Mathematics, 269:265–297, 1972. CODEN LNMAA2. ISBN 3-540-05896-6 (print), 3-540-37549-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0081554/.

Verdier:1972:TFF

[Ver72c] J. L. Verdier. Topologies et faisceaux. (French) []. Lecture Notes in Mathematics, 269:219–263, 1972. CODEN LNMAA2. ISBN 3-540-05896-6 (print), 3-540-37549-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0081553/.

Verdier:1973:DCG

[Ver73] Jean-Louis Verdier. Dualité dans la cohomologie des groupes profinis. (French) []. Lecture Notes in Mathematics, 5:183–206, 1973. CODEN LNMAA2. ISBN 3-540-06084-7 (print), 3-662-21553-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-662-21553-1_4/. See erratum [Ser73j].

Verdier:1974:IPR

[Ver74] Jean-Louis Verdier. Indépendance par rapport à ℓ des polynômes caractéristiques des endomorphismes de Frobenius de la cohomologie ℓ-adique. (French) []. Lecture Notes in Mathematics, 383: 98–115, 1974. CODEN LNMAA2. ISBN 3-540-06796-5 (print), 3-540-38450-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057304/.

Vesentini:1970:PEF

[Ves70] Edoardo Vesentini. Periodic entire functions and reducible Banach algebras. Lecture Notes in Mathematics, 155:203–211, 1970. CODEN LNMAA2. ISBN 3-540-05183-X (print), 3-540-36344-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060326/.

vonFinckenstein:1973:DLQ

[vF73] Karl Graf Finck von Finckenstein. Differenzenverfahren Zur Lösung Quasilinearer Diffusionsgleichungen in Zylindersymmetrie. (German) []. Lecture Notes in Mathematics, 333:89–95, 1973. CODEN LNMAA2. ISBN 3-540-06378-1 (print), 3-540-46986-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060689/.

vonFinckenstein:1974:DDP

[vFD74] K. Graf Finck von Finckenstein and D. Düchs. Differenzenverfahren für Diffusionsgleichungen der Plasmaphysik. (German) []. Lecture Notes in Mathematics, 395:3–17, 1974. CODEN LN-MAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060661/.

Vidal:1974:GDC

[Vid74] Enrique Vidal, editor. Géométrie Différentielle: Colloque, Santiago de Compostela Espagne 1972, volume 392 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 392.

URL http://link.springer.com/book/10.1007/BFb0064102;
 http://www.springerlink.com/content/978-3-540-38452-
6.

vanLint:1971:BM

[vL71a] Jacobus H. van Lint. Back matter. Lecture Notes in Mathematics, 201:130–140, 1971. CODEN LNMAA2. ISBN 3-540-05476-6 (print), 3-662-20712-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-662-20712-3/1.pdf.

vanLint:1971:CT

[vL71b] Jacobus H. van Lint. Coding Theory, volume 201 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05476-6 (print), 3-662-20712-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 140 (est.) pp. LCCN QA3 .L28

no. 201. URL http://link.springer.com/book/10.1007/978-
3-662-20712-3;   http://www.springerlink.com/content/
978-3-662-20712-3.

vanLint:1971:CC

[vL71c] Jacobus H. van Lint. Cyclic codes. Lecture Notes in Mathematics, 201:42–60, 1971. CODEN LNMAA2. ISBN 3-540-05476-6 (print), 3-662-20712-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-20712-3_3/.

vanLint:1971:FM

[vL71d] Jacobus H. van Lint. Front matter. Lecture Notes in Mathematics, 201:i–vii, 1971. CODEN LNMAA2. ISBN 3-540-05476-6 (print), 3-662-20712-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-662-20712-3/1.pdf.

vanLint:1971:ICC

[vL71e] Jacobus H. van Lint. Important cyclic codes. Lecture Notes in Mathematics, 201:61–85, 1971. CODEN LNMAA2. ISBN 3-540-05476-6 (print), 3-662-20712-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-20712-3_4/.

vanLint:1971:I

[vL71f] Jacobus H. van Lint. Introduction. Lecture Notes in Mathematics, 201:1–14, 1971. CODEN LNMAA2. ISBN 3-540-05476-6 (print), 3-662-20712-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-20712-3_1/.

vanLint:1971:LC

[vL71g] Jacobus H. van Lint. Linear codes. Lecture Notes in Mathematics, 201:15–41, 1971. CODEN LNMAA2. ISBN 3-540-05476-6 (print), 3-662-20712-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-20712-3_2/.

vanLint:1971:PC

[vL71h] Jacobus H. van Lint. Perfect codes. Lecture Notes in Mathematics, 201:86–119, 1971. CODEN LNMAA2. ISBN 3-540-05476-6 (print), 3-662-20712-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-662-20712-3_5/.

vanLint:1971:WE

[vL71i] Jacobus H. van Lint. Weight enumeration. Lecture Notes in Mathematics, 201:120–129, 1971. CODEN LNMAA2. ISBN 3-540-05476-6 (print), 3-662-20712-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-662-20712-3_6/.

vanLint:1973:BM

[vL73a] Jacobus H. van Lint. Back matter. Lecture Notes in Mathematics, 201:130–139, 1973. CODEN LNMAA2. ISBN 3-540-06363-3 (print), 3-540-36657-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bbm:978-3-540-36657-7/1.pdf.

vanLint:1973:CT

[vL73b] Jacobus H. van Lint. Coding Theory, volume 201 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06363-3 (print), 3-540-36657-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 139 (est.) pp. LCCN QA3 .L28

no. 201. URL http://link.springer.com/book/10.1007/978-
3-540-36657-7;   http://www.springerlink.com/content/
978-3-540-36657-7.

vanLint:1973:CC

[vL73c] Jacobus H. van Lint. Cyclic codes. Lecture Notes in Mathematics, 201:42–60, 1973. CODEN LNMAA2. ISBN 3-540-06363-3 (print), 3-540-36657-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36657-7_3/.

vanLint:1973:FM

[vL73d] Jacobus H. van Lint. Front matter. Lecture Notes in Mathematics, 201:i–vii, 1973. CODEN LNMAA2. ISBN 3-540-06363-3 (print), 3-540-36657-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/content/
pdf/bfm:978-3-540-36657-7/1.pdf.

vanLint:1973:ICC

[vL73e] Jacobus H. van Lint. Important cyclic codes. Lecture Notes in Mathematics, 201:61–85, 1973. CODEN LNMAA2. ISBN 3-540-06363-3 (print), 3-540-36657-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36657-7_4/.

vanLint:1973:I

[vL73f] Jacobus H. van Lint. Introduction. Lecture Notes in Mathematics, 201:1–14, 1973. CODEN LNMAA2. ISBN 3-540-06363-3 (print), 3-540-36657-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36657-7_1/.

vanLint:1973:LC

[vL73g] Jacobus H. van Lint. Linear codes. Lecture Notes in Mathematics, 201:15–41, 1973. CODEN LNMAA2. ISBN 3-540-06363-3 (print), 3-540-36657-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36657-7_2/.

vanLint:1973:PC

[vL73h] Jacobus H. van Lint. Perfect codes. Lecture Notes in Mathematics, 201:86–119, 1973. CODEN LNMAA2. ISBN 3-540-06363-3 (print), 3-540-36657-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/978-3-540-36657-7_5/.

vanLint:1973:WE

[vL73i] Jacobus H. van Lint. Weight enumeration. Lecture Notes in Mathematics, 201:120–129, 1973. CODEN LNMAA2. ISBN 3-540-06363-3 (print), 3-540-36657-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/978-3-540-36657-7_6/.

vanLint:1974:CTS

[vL74a] Jacobus H. van Lint, editor. Combinatorial Theory Seminar Eindhoven University of Technology, volume 382 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 382.

URL http://link.springer.com/book/10.1007/BFb0057319;
 http://www.springerlink.com/content/978-3-540-38316-
1.

vanLint:1974:PC

[vL74b] Jacobus H. van Lint. I. Permutations and combinations. Lecture Notes in Mathematics, 382:1–3, 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057320/.

vanLint:1974:IIF

[vL74c] Jacobus H. van Lint. II. Inversion formulae. Lecture Notes in Mathematics, 382:4–20, 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057321/.

vanLint:1974:IGF

[vL74d] Jacobus H. van Lint. III. Generating functions and recursions. Lecture Notes in Mathematics, 382:21–32, 1974. CODEN LN-MAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057322/.

vanLint:1974:IP

[vL74e] Jacobus H. van Lint. IV. Partitions. Lecture Notes in Mathematics, 382:33–39, 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057323/.

vanLint:1974:IBS

[vL74f] Jacobus H. van Lint. IX. de Bruign sequences. Lecture Notes in Mathematics, 382:82–92, 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057328/.

vanLint:1974:VDR

[vL74g] Jacobus H. van Lint. V. Distinct representatives. Lecture Notes in Mathematics, 382:40–68, 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057324/.

vanLint:1974:VRT

[vL74h] Jacobus H. van Lint. VI. Ramsey’s theorem. Lecture Notes in Mathematics, 382:69–76, 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057325/.

vanLint:1974:VSE

[vL74i] Jacobus H. van Lint. VII. Some extremal problems. Lecture Notes in Mathematics, 382:77–79, 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057326/.

vanLint:1974:VCS

[vL74j] Jacobus H. van Lint. VIII. Convex spaces. Lecture Notes in Mathematics, 382:80–81, 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057327.

vanLint:1974:XBD

[vL74k] Jacobus H. van Lint. X. Block designs. Lecture Notes in Mathematics, 382:93–102, 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0057329/.

vanLint:1974:XDS

[vL74l] Jacobus H. van Lint. XI. Difference sets. Lecture Notes in Mathematics, 382:103–104, 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057330.

vanLint:1974:XFG

[vL74m] Jacobus H. van Lint. XII. Finite geometries. Lecture Notes in Mathematics, 382:105–110, 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057331/.

vanLint:1974:XOL

[vL74n] Jacobus H. van Lint. XIII. Orthogonal latin squares. Lecture Notes in Mathematics, 382:111–119, 1974. CODEN LN-MAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057332/.

vanLint:1974:XHM

[vL74o] Jacobus H. van Lint. XIV. Hadamard matrices. Lecture Notes in Mathematics, 382:120–124, 1974. CODEN LNMAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0057333/.

vanLint:1974:XCB

[vL74p] Jacobus H. van Lint. XV. Constructions of block designs. Lecture Notes in Mathematics, 382:125–129, 1974. CODEN LN-MAA2. ISBN 3-540-06735-3 (print), 3-540-38316-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057334/.

vonWaldenfels:1973:ATP

[vW73] Wilhelm von Waldenfels. An approach to the theory of pressure broadening of spectral lines. Lecture Notes in Mathematics, 296: 19–69, 1973. CODEN LNMAA2. ISBN 3-540-06211-4 (print), 3-540-38485-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059819/.

Wachspress:1971:IPN

[Wac71a] Eugene L. Wachspress. Iteration parameters in the numerical solution of elliptic problems. Lecture Notes in Mathematics, 193: 93–109, 1971. CODEN LNMAA2. ISBN 3-540-05422-7 (print), 3-540-36538-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060344/.

Wachspress:1971:RBF

[Wac71b] Eugene L. Wachspress. A rational basis for function approximation. Lecture Notes in Mathematics, 228:223–252, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069458/.

Wacker:1972:NHK

[Wac72] Hansjörg Wacker. Nichtlineare Homotopien zur Konstruktion von Startlösungen für Iterationsverfahren. (German) []. Lecture Notes in Mathematics, 267:51–67, 1972. CODEN LNMAA2. ISBN 3-540-05895-8 (print), 3-540-37540-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061613/.

Wachspress:1974:AGF

[Wac74a] Eugene L. Wachspress. Algebraic-geometry foundations for finiteelement computation. Lecture Notes in Mathematics, 363:177– 188, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069136/.

Wacker:1974:GLV

[Wac74b] Hansjörg Wacker. Globalisierung lokaler Verfahren. (German) []. Lecture Notes in Mathematics, 395:301–307, 1974. CODEN LNMAA2. ISBN 3-540-06832-5 (print), 3-540-37771-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060680/.

Waelbroeck:1971:TVSa

[Wae71a] Lucien Waelbroeck. Topological vector spaces. Lecture Notes in Mathematics, 230:1–22, 1971. CODEN LNMAA2. ISBN 3-540-05650-5 (print), 3-540-36938-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061235/.

Waelbroeck:1971:TVSb

[Wae71b] Lucien Waelbroeck. Topological Vector Spaces and Algebras, volume 230 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05650-5 (print), 3-540-36938-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vii + 158 pp. LCCN QA3 .L28 no. 230. URL http:/

/link.springer.com/book/10.1007/BFb0061234;   http://
www.springerlink.com/content/978-3-540-36938-7.

Waelbroeck:1973:SST

[Wae73] Lucien Waelbroeck, editor. Summer School on Topological Vector Spaces, volume 331 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06367-6 (print), 3-540-46977-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 331. URL http:/

/link.springer.com/book/10.1007/BFb0068225;   http://
www.springerlink.com/content/978-3-540-46977-3.

Waelbroeck:1974:HFC

[Wae74a] Lucien Waelbroeck. The holomorphic functional calculus and infinite dimensional holomorphy. Lecture Notes in Mathematics, 364:101–108, 1974. CODEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069009/.

Waelbroeck:1974:WAF

[Wae74b] Lucien Waelbroeck. Weak analytic functions and the closed graph theorem. Lecture Notes in Mathematics, 364:97–100, 1974. CO-DEN LNMAA2. ISBN 3-540-06619-5 (print), 3-540-37915-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069008/.

Wagner:1970:FSA

[Wag70] J. Wagner. Faisceau structural associé à une algèbre de Banach. (French) []. Lecture Notes in Mathematics, 116:164–167, 1970. CODEN LNMAA2. ISBN 3-540-04905-3 (print), 3-540-36240-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059253/.

Wagoner:1971:AIP

[Wag71] J. B. Wagoner. Algebraic invariants for pseudo-isotopies. Lecture Notes in Mathematics, 209:164–190, 1971. CODEN LN-MAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068902/.

Wagoner:1973:BSH

[Wag73] J. Wagoner. Buildings, stratifications, and higher K-theory. Lecture Notes in Mathematics, 341:148–165, 1973. CODEN LN-MAA2. ISBN 3-540-06434-6 (print), 3-540-37767-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067054/.

Wall:1971:IPT

[Wal71a] C. T. C. Wall. Introduction to the preparation theorem. Lecture Notes in Mathematics, 192:90–96, 1971. CODEN LN-MAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066813/.

Wall:1971:LSC

[Wal71b] C. T. C. Wall. Lectures on C^{∞}-stability and classification. Lecture Notes in Mathematics, 192:178–206, 1971. CODEN LN-MAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066823/.

Wall:1971:PLSa

[Wal71c] C. T. C. Wall, editor. Proceedings of Liverpool Singularities — Symposium I, volume 192 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 192. URL http:/

/link.springer.com/book/10.1007/BFb0066809;   http://
www.springerlink.com/content/978-3-540-36531-0.

Wall:1971:PLSb

[Wal71d] C. T. C. Wall, editor. Proceedings of Liverpool Singularities Symposium II, volume 209 of Lecture Notes in Mathematics.

Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 209. URL http:/

/link.springer.com/book/10.1007/BFb0068889;   http://
www.springerlink.com/content/978-3-540-36868-7.

Wall:1971:RGV

[Wal71e] C. T. C. Wall. Reflections on gradient vector fields. Lecture Notes in Mathematics, 209:191–195, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068903/.

Wall:1971:RGS

[Wal71f] C. T. C. Wall. Remark on geometric singularities. Lecture Notes in Mathematics, 209:121, 1971. CODEN LNMAA2. ISBN 3-540-05511-8 (print), 3-540-36868-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068898.

Wall:1971:SSS

[Wal71g] C. T. C. Wall. Stratified sets: a survey. Lecture Notes in Mathematics, 192:133–140, 1971. CODEN LNMAA2. ISBN 3-540-05402-2 (print), 3-540-36531-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0066818/.

Wall:1972:DBS

[Wal72a] Charles R. Wall. Density bounds for the sum of divisors function. Lecture Notes in Mathematics, 251:283–287, 1972. CODEN LNMAA2. ISBN 3-540-05723-4 (print), 3-540-37098-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058798/.

Wallis:1972:A

[Wal72b] Jennifer Seberry Wallis. Aftermath. Lecture Notes in Mathematics, 292:492–498, 1972. CODEN LNMAA2. ISBN 3-540-06035-9 (print), 3-540-37994-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069912/.

Wallis:1972:HM

[Wal72c] Jennifer Seberry Wallis. Hadamard matrices. Lecture Notes in Mathematics, 292:274–489, 1972. CODEN LNMAA2. ISBN 3-540-06035-9 (print), 3-540-37994-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069911/.

Wallis:1972:RS

[Wal72d] W. D. Wallis. Room squares. Lecture Notes in Mathematics, 292: 30–121, 1972. CODEN LNMAA2. ISBN 3-540-06035-9 (print), 3-540-37994-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069909/.

Walter:1972:SNA

[Wal72e] W. Walter. Some new aspects of the line method for parabolic differential equations. Lecture Notes in Mathematics, 280:181– 189, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066927/.

Waldhausen:1973:WGG

[Wal73a] Friedhelm Waldhausen. Whitehead groups of generalized free products. Lecture Notes in Mathematics, 342:155–179, 1973. CO-DEN LNMAA2. ISBN 3-540-06435-4 (print), 3-540-37770-0 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0073726/.

Wall:1973:FAT

[Wal73b] C. T. C. Wall. Foundations of algebraic L-theory. Lecture Notes in Mathematics, 343:266–300, 1973. CODEN LNMAA2. ISBN 3-540-06436-2 (print), 3-540-37773-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061371/.

Walter:1973:SPP

[Wal73c] Charles Walter. Stability properties and periodic behavior of controlled biochemical systems. Lecture Notes in Mathematics, 322: 343–357, 1973. CODEN LNMAA2. ISBN 3-540-06251-3 (print), 3-540-38558-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060575/.

Waldschmidt:1974:MBF

[Wal74a] Michel Waldschmidt. La méthode de Baker. (French) []. Lecture Notes in Mathematics, 402:227–247, 1974. CODEN LN-MAA2. ISBN 3-540-06874-0 (print), 3-540-37836-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065328/.

Waldschmidt:1974:MGF

[Wal74b] Michel Waldschmidt. La méthode de Gel’fond. (French) []. Lecture Notes in Mathematics, 402:71–99, 1974. CODEN LN-MAA2. ISBN 3-540-06874-0 (print), 3-540-37836-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065323/.

Waldschmidt:1974:MSF

[Wal74c] Michel Waldschmidt. La méthode de Schneider. (French) []. Lecture Notes in Mathematics, 402:42–70, 1974. CODEN LN-MAA2. ISBN 3-540-06874-0 (print), 3-540-37836-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065322/.

Waldschmidt:1974:NTF

[Wal74d] Michel Waldschmidt. Nombres Transcendants. (French) [Transcendental numbers], volume 402 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06874-0 (print), 3-540-37836-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 277 pp. LCCN QA3 .L28 no. 402.

URL http://link.springer.com/book/10.1007/BFb0065320;
 http://www.springerlink.com/content/978-3-540-37836-
5.

Waldschmidt:1974:PF

[Wal74e] Michel Waldschmidt. Préliminaires. (French) []. Lecture Notes in Mathematics, 402:1–41, 1974. CODEN LNMAA2. ISBN 3-540-06874-0 (print), 3-540-37836-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0065321/.

Waldschmidt:1974:PIA

[Wal74f] Michel Waldschmidt. Propriétés d’indépendance algébrique de la fonction exponentielle. (French) []. Lecture Notes in Mathematics, 402:183–226, 1974. CODEN LNMAA2. ISBN 3-540-06874-0 (print), 3-540-37836-7 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0065327/.

Waldschmidt:1974:TTF

[Wal74g] Michel Waldschmidt. Type de transcendance. (French) []. Lecture Notes in Mathematics, 402:100–142, 1974. CODEN LN-MAA2. ISBN 3-540-06874-0 (print), 3-540-37836-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065324/.

Waldschmidt:1974:CTF

[Wal74h] Michel Waldschmidt. Un critère de transcendance. (French) []. Lecture Notes in Mathematics, 402:143–162, 1974. CODEN LN-MAA2. ISBN 3-540-06874-0 (print), 3-540-37836-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065325/.

Waldschmidt:1974:ZPE

[Wal74i] Michel Waldschmidt. Zéros de polynômes exponentiels. (French) []. Lecture Notes in Mathematics, 402:163–182, 1974. CODEN LNMAA2. ISBN 3-540-06874-0 (print), 3-540-37836-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0065326/.

Walker:1974:CCS

[Wal74j] A. G. Walker. A certain conformal structure. Lecture Notes in Mathematics, 392:98–110, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064117/.

Wallis:1974:WME

[Wal74k] Jennifer Seberry Wallis. Williamson matrices of even order. Lecture Notes in Mathematics, 403:132–142, 1974. CODEN LN-MAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0057387/.

Wallis:1974:S

[Wal74l] W. D. Wallis. Supersquares. Lecture Notes in Mathematics, 403: 143–148, 1974. CODEN LNMAA2. ISBN 3-540-06903-8 (print), 3-540-37837-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0057388/.

Walter:1974:LMP

[Wal74m] Professor Dr. Wolfgang Walter. The line method for parabolic differential equations problems in boundary layer theory and existence of periodic solutions. Lecture Notes in Mathematics, 430: 395–413, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066279/.

Wang:1970:LRP

[Wan70] Hao Wang. On the long-range prospects of automatic theoremproving. Lecture Notes in Mathematics, 125:101–111, 1970. CO-DEN LNMAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060627/.

Wang:1972:FSF

[Wan72] Kai Wang. Free and semi-free smooth actions of S^{1} and S^{3} on homotopy spheres. Lecture Notes in Mathematics, 298:119– 130, 1972. CODEN LNMAA2. ISBN 3-540-06077-4 (print), 3-540-38063-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070036/.

Wang:1974:SAH

[Wan74] Kai Wang. Semifree actions on homotopy spheres. Lecture Notes in Mathematics, 428:219–229, 1974. CODEN LNMAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070540/.

Ward:1970:SVM

[War70] L. E. Ward, Jr. Set-valued mappings on partially ordered spaces. Lecture Notes in Mathematics, 171:88–99, 1970. CODEN LN-MAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069727/.

Ware:1973:SWR

[War73] Roger Ware. The structure of the Witt ring and quotients of abelian group rings. Lecture Notes in Mathematics, 353:180– 186, 1973. CODEN LNMAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059269/.

Wasow:1971:CCP

[Was71] Wolfgang Wasow. The central connection problem at turning points of linear differential equations. Lecture Notes in Mathematics, 183:158–164, 1971. CODEN LNMAA2. ISBN 3-540-05369-7 (print), 3-540-36454-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060415/.

Wassermann:1974:FDG

[Was74a] Gordon Wassermann. Finitely determined germs. Lecture Notes in Mathematics, 393:35–53, 1974. CODEN LNMAA2. ISBN 3-540-06794-9 (print), 3-540-38423-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061660/.

Wassermann:1974:P

[Was74b] Gordon Wassermann. Preliminaries. Lecture Notes in Mathematics, 393:1–34, 1974. CODEN LNMAA2. ISBN 3-540-06794-9 (print), 3-540-38423-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061659/.

Wassermann:1974:SEC

[Was74c] Gordon Wassermann. The seven elementary catastrophes: a classification theorem. Lecture Notes in Mathematics, 393:116– 152, 1974. CODEN LNMAA2. ISBN 3-540-06794-9 (print), 3-540-38423-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061663/.

Wassermann:1974:SUb

[Was74d] Gordon Wassermann. Stability of Unfoldings, volume 393 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974.

CODEN LNMAA2. ISBN 3-540-06794-9 (print), 3-540-38423-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). ix + 164 pp. LCCN QA3 .L28 no. 393. URL http:/

/link.springer.com/book/10.1007/BFb0061658;   http://
www.springerlink.com/content/978-3-540-38423-6.

Wassermann:1974:SUa

[Was74e] Gordon Wassermann. Stable unfoldings. Lecture Notes in Mathematics, 393:84–115, 1974. CODEN LNMAA2. ISBN 3-540-06794-9 (print), 3-540-38423-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061662/.

Wassermann:1974:UU

[Was74f] Gordon Wassermann. Universal unfoldings. Lecture Notes in Mathematics, 393:54–83, 1974. CODEN LNMAA2. ISBN 3-540-06794-9 (print), 3-540-38423-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061661/.

Watson:1974:CNS

[Wat74] G. A. Watson, editor. Conference on the Numerical Solution of Differential Equations: Dundee 1973, volume 363 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 0-387-06617-9, 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 363; QA3.L28 no.363. URL http:

//link.springer.com/book/10.1007/BFb0069119;   http:/
/www.springerlink.com/content/978-3-540-37914-0.

Wedig:1972:RIL

[Wed72] W. Wedig. Regions of instability for a linear system with random parametric excitation. Lecture Notes in Mathematics, 294: 160–172, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0064941/.

Weil:1970:QPF

[Wei70] Michel Weil. Quasi-processus. (French) []. Lecture Notes in Mathematics, 124:216–239, 1970. CODEN LNMAA2. ISBN 3-540-04913-4 (print), 3-540-36260-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059345/.

Weil:1971:BC

[Wei71a] André Weil. Basic concepts. Lecture Notes in Mathematics, 189: 17–21, 1971. CODEN LNMAA2. ISBN 3-540-05382-4 (print), 3-540-36502-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061204/.

Weil:1971:CC

[Wei71b] André Weil. The classical case. Lecture Notes in Mathematics, 189:1–8, 1971. CODEN LNMAA2. ISBN 3-540-05382-4 (print), 3-540-36502-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061202/.

Weil:1971:CL

[Wei71c] André Weil. The convergence lemmas. Lecture Notes in Mathematics, 189:35–39, 1971. CODEN LNMAA2. ISBN 3-540-05382-4 (print), 3-540-36502-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061206/.

Weil:1971:DS

[Wei71d] André Weil. Dirichlet series. Lecture Notes in Mathematics, 189: 9–15, 1971. CODEN LNMAA2. ISBN 3-540-05382-4 (print), 3-540-36502-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061203/.

Weil:1971:DSA

[Wei71e] André Weil. Dirichlet Series and Automorphic Forms: Lezioni Fermiane, volume 189 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05382-4 (print), 3-540-36502-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 164 (est.) pp. LCCN QA3 .L28 no. 189. URL http:/

/link.springer.com/book/10.1007/BFb0061201;   http://
www.springerlink.com/content/978-3-540-36502-0.

Weil:1971:E

[Wei71f] André Weil. Examples. Lecture Notes in Mathematics, 189:141– 164, 1971. CODEN LNMAA2. ISBN 3-540-05382-4 (print), 3-540-36502-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061212/.

Weil:1971:EP

[Wei71g] André Weil. The extension problem. Lecture Notes in Mathematics, 189:23–34, 1971. CODEN LNMAA2. ISBN 3-540-05382-4 (print), 3-540-36502-8 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061205/.

Weil:1971:FF

[Wei71h] André Weil. Function-fields. Lecture Notes in Mathematics, 189: 47–61, 1971. CODEN LNMAA2. ISBN 3-540-05382-4 (print), 3-540-36502-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061208/.

Weil:1971:HIP

[Wei71i] André Weil. Harmonicity at an infinite place. Lecture Notes in Mathematics, 189:63–104, 1971. CODEN LNMAA2. ISBN 3-540-05382-4 (print), 3-540-36502-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061209/.

Weil:1971:HSC

[Wei71j] André Weil. Harmonicity (special case). Lecture Notes in Mathematics, 189:105–111, 1971. CODEN LNMAA2. ISBN 3-540-05382-4 (print), 3-540-36502-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061210/.

Weil:1971:HO

[Wei71k] André Weil. Hecke operators. Lecture Notes in Mathematics, 189: 41–45, 1971. CODEN LNMAA2. ISBN 3-540-05382-4 (print), 3-540-36502-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061207/.

Weil:1971:NF

[Wei71l] André Weil. Number-fields. Lecture Notes in Mathematics, 189: 113–140, 1971. CODEN LNMAA2. ISBN 3-540-05382-4 (print), 3-540-36502-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0061211/.

Weiss:1973:WDA

[Wei73] Guido Weiss. Weak^{∗}-Dirichlet algebras induced by the ergodic Hilbert transform. Lecture Notes in Mathematics, 336:20–27, 1973. CODEN LNMAA2. ISBN 3-540-06392-7 (print), 3-540-47170-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0065784/.

Wells:1971:PCS

[Wel71a] R. O. Wells, Jr. Parametrizing the compact submanifolds of a period matrix domain by a Stein manifold. Lecture Notes in Mathematics, 184:121–150, 1971. CODEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059282/.

Welsh:1971:MBD

[Wel71b] D. J. A. Welsh. Matroids and block designs. Lecture Notes in Mathematics, 211:95–106, 1971. CODEN LNMAA2. ISBN 3-540-05545-2 (print), 3-540-36877-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061078/.

Wendland:1972:BVP

[Wen72] W. Wendland. On boundary value problems of generalized analytic functions. Lecture Notes in Mathematics, 280:190–201, 1972. CODEN LNMAA2. ISBN 3-540-05962-8 (print), 3-540-37618-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066928/.

Wendland:1974:IEM

[Wen74a] Wolfgang L. Wendland. An integral equation method for generalized analytic functions. Lecture Notes in Mathematics, 430: 414–452, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066280/.

Wendroff:1974:SGM

[Wen74b] Burton Wendroff. Spline-Galerkin methods for initial-value problems with variable coefficients. Lecture Notes in Mathematics, 363:189–195, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0069137/.

Wermer:1971:ES

[Wer71] John Wermer. On an example of Stolzenberg. Lecture Notes in Mathematics, 184:79–84, 1971. CODEN LNMAA2. ISBN 3-540-05370-0 (print), 3-540-36455-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059279/.

Whitehead:1970:WNS

[Whi70] George W. Whitehead. The work of Norman E. Steenrod in algebraic topology: an appreciation. Lecture Notes in Mathematics, 168:1–10, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058514/.

Widlund:1971:SRB

[Wid71a] O. B. Widlund. Some results on best possible error bounds for finite element methods and approximation with piecewise polynomial functions. Lecture Notes in Mathematics, 228:253–263, 1971. CODEN LNMAA2. ISBN 3-540-05656-4 (print), 3-540-36976-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069459/.

Widlund:1971:IFD

[Wid71b] Olof Widlund. Introduction to finite difference approximations to initial value problems for partial differential equations. Lecture Notes in Mathematics, 193:111–152, 1971. CODEN LN-MAA2. ISBN 3-540-05422-7 (print), 3-540-36538-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060345/.

Wielen:1974:NIB

[Wie74] Roland Wielen. On the numerical integration of the N-body problem for star clusters. Lecture Notes in Mathematics, 362:

276–290, 1974. CODEN LNMAA2. ISBN 3-540-06602-0 (print), 3-540-37911-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0066596/.

Wille:1970:AKG

[Wil70a] Rudolf Wille. Affine Koordinatisierung. (German) []. Lecture Notes in Mathematics, 113:26–35, 1970. CODEN LN-MAA2. ISBN 3-540-04902-9 (print), 3-540-36221-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060217/.

Wille:1970:AGG

[Wil70b] Rudolf Wille. Algebraische Grundbegriffe. (German) []. Lecture Notes in Mathematics, 113:5–11, 1970. CODEN LNMAA2. ISBN 3-540-04902-9 (print), 3-540-36221-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0060215/.

Wille:1970:EG

[Wil70c] Rudolf Wille. Einleitung. (German) []. Lecture Notes in Mathematics, 113:1–4, 1970. CODEN LNMAA2. ISBN 3-540-04902-9 (print), 3-540-36221-5 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0060214/.

Wille:1970:GEV

[Wil70d] Rudolf Wille. Geometrien mit eindeutigen Verbindungsgeraden. (German) []. Lecture Notes in Mathematics, 113:36–53, 1970. CODEN LNMAA2. ISBN 3-540-04902-9 (print), 3-540-36221-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060218/.

Wille:1970:GGG

[Wil70e] Rudolf Wille. Geometrische Grundbegriffe. (German) []. Lecture Notes in Mathematics, 113:12–25, 1970. CODEN LN-MAA2. ISBN 3-540-04902-9 (print), 3-540-36221-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060216/.

Wille:1970:KGC

[Wil70f] Rudolf Wille. Kongruenzklassengeometrien. (German) [Congruence class geometries], volume 113 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04902-9 (print), 3-540-36221-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iii + 99 pp. LCCN QA3 .L28 no. 113.

URL http://link.springer.com/book/10.1007/BFb0060213;
 http://www.springerlink.com/content/978-3-540-36221-
0.

Wille:1970:PGG

[Wil70g] Rudolf Wille. Pseudoaffine geometrien. (German) [pseudoaffine geometries]. Lecture Notes in Mathematics, 113:54–62, 1970. CO-DEN LNMAA2. ISBN 3-540-04902-9 (print), 3-540-36221-5 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060219/.

Wille:1970:RPK

[Wil70h] Rudolf Wille. Rahmenaussagen und primitive Klassen. (German) []. Lecture Notes in Mathematics, 113:63–82, 1970. CODEN LNMAA2. ISBN 3-540-04902-9 (print), 3-540-36221-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060220/.

Wille:1970:SSP

[Wil70i] Rudolf Wille. Schliessungsaussagen und starke primitive Klassen. (German) []. Lecture Notes in Mathematics, 113:83–95, 1970. CODEN LNMAA2. ISBN 3-540-04902-9 (print), 3-540-36221-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060221/.

Willems:1972:LFG

[Wil72] Jacques L. Willems. Lyapunov functions and global frequency domain stability criteria for a class of stochastic feedback systems. Lecture Notes in Mathematics, 294:131–146, 1972. CODEN LNMAA2. ISBN 3-540-06050-2 (print), 3-540-38000-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0064939/.

Williams:1973:I

[Wil73a] Floyd L. Williams. Introduction. Lecture Notes in Mathematics, 358:1–6, 1973. CODEN LNMAA2. ISBN 3-540-06567-9 (print), 3-540-37865-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0067485/.

Williams:1973:PIR

[Wil73b] Floyd L. Williams. Preliminaries on induced representations. Lecture Notes in Mathematics, 358:7–22, 1973. CODEN LN-MAA2. ISBN 3-540-06567-9 (print), 3-540-37865-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0067486/.

Williams:1973:RIC

[Wil73c] Floyd L. Williams. Representations induced by characters of a subgroup of the Cartan subgroup. Lecture Notes in Mathematics, 358:23–73, 1973. CODEN LNMAA2. ISBN 3-540-06567-9 (print), 3-540-37865-0 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0067487/.

Williams:1973:TPPa

[Wil73d] Floyd L. Williams. The tensor product of principal series representations of a complex semi-simple Lie group. Lecture Notes in Mathematics, 358:74–120, 1973. CODEN LNMAA2. ISBN 3-540-06567-9 (print), 3-540-37865-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0067488/.

Williams:1973:TPPb

[Wil73e] Floyd L. Williams. Tensor Products of Principal Series Representations: Reduction of Tensor Products of Principal Series Representations of Complex Semisimple Lie Groups, volume 358 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1973. CODEN LNMAA2. ISBN 3-540-06567-9 (print), 3-540-37865-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). vi + 132 pp. LCCN QA3 .L28 no. 358. URL http:

//link.springer.com/book/10.1007/BFb0067484;   http:/
/www.springerlink.com/content/978-3-540-37865-5.

Wilhelmson:1974:SPD

[Wil74a] Robert Wilhelmson. Solving partial differential equations using ILLIAC IV. Lecture Notes in Mathematics, 430:453–476, 1974. CODEN LNMAA2. ISBN 3-540-07021-4 (print), 3-540-37302-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0066281/.

Wilkerson:1974:DS

[Wil74b] Clarence Wilkerson. modp Decompositions of modp H-spaces. Lecture Notes in Mathematics, 428:52–57, 1974. CODEN LN-MAA2. ISBN 3-540-07019-2 (print), 3-540-37300-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070534/.

Wilkerson:1974:GCS

[Wil74c] Clarence Wilkerson. Genus and cancellation for h-spaces. Lecture Notes in Mathematics, 418:158–160, 1974. CODEN LN-MAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070652/.

Wilkerson:1974:SMC

[Wil74d] Clarence Wilkerson. Self-maps of classifying spaces. Lecture Notes in Mathematics, 418:150–157, 1974. CODEN LN-MAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070651/.

Willmore:1974:MCI

[Wil74e] T. J. Willmore. Minimal conformal immersions. Lecture Notes in Mathematics, 392:111–120, 1974. CODEN LNMAA2. ISBN 3-540-06797-3 (print), 3-540-38452-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0064118/.

Wloka:1973:GTS

[Wlo73] Joseph Wloka. Gelfand triplets and spectral theory. Lecture Notes in Mathematics, 331:163–182, 1973. CODEN LNMAA2. ISBN 3-540-06367-6 (print), 3-540-46977-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068229/.

Wolf:1971:CMU

[Wol71] Joseph A. Wolf. Complex manifolds and unitary representations. Lecture Notes in Mathematics, 185:242–287, 1971. CODEN LNMAA2. ISBN 3-540-05372-7 (print), 3-540-36493-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058772/.

Wong:1970:FS

[Won70] Raymond Y. T. Wong. Factors of subclasses of 2^{(Im)}. Lecture Notes in Mathematics, 171:100–101, 1970. CODEN LN-MAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069728.

Woodcock:1974:SRB

[Woo74a] A. E. R. Woodcock. Stereographic reconstructions of the behaviour space of the cuspoids catastrophes. Lecture Notes in Mathematics, 373:132–231, 1974. CODEN LNMAA2. ISBN 3-540-06681-0 (print), 3-540-37941-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068971/.

Woodcock:1974:SRCa

[Woo74b] A. E. R. Woodcock. Stereographic reconstructions of the catastrophe manifolds of the cuspoids: (2) the swallowtall. Lecture Notes in Mathematics, 373:149–165, 1974. CODEN LN-MAA2. ISBN 3-540-06681-0 (print), 3-540-37941-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068972/.

Woodcock:1974:SRCb

[Woo74c] A. E. R. Woodcock. Stereographic reconstructions of the catastrophe manifolds of the cuspoids: (3) the butterfly. Lecture Notes in Mathematics, 373:166–192, 1974. CODEN LNMAA2. ISBN 3-540-06681-0 (print), 3-540-37941-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068973/.

Woodcock:1974:SRCd

[Woo74d] A. E. R. Woodcock. Stereographic reconstructions of the catastrophe manifolds of the cuspoids: (5) the star. Lecture Notes in Mathematics, 373:212–231, 1974. CODEN LNMAA2. ISBN 3-540-06681-0 (print), 3-540-37941-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068975/.

Woodcock:1974:SRCc

[Woo74e] A. E. R. Woodcock. Stereographic reconstuctions of the catastrophe manifolds of the cuspoids: (4) the wigwam. Lecture Notes in Mathematics, 373:193–211, 1974. CODEN LNMAA2. ISBN 3-540-06681-0 (print), 3-540-37941-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068974/.

Woodcock:1974:GPR

[WP74a] A. E. R. Woodcock and T. Poston. The geometrical properties of the reduced double cusp. Lecture Notes in Mathematics, 373: 232–254, 1974. CODEN LNMAA2. ISBN 3-540-06681-0 (print), 3-540-37941-X (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0068976/.

Woodcock:1974:GSE

[WP74b] A. E. R. Woodcock and T. Poston. A Geometrical Study of the Elementary Catastrophes, volume 373 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06681-0 (print), 3-540-37941-X (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). 254 (est.) pp. LCCN QA3 .L28 no. 373. URL http://

link.springer.com/book/10.1007/BFb0068967;   http://
www.springerlink.com/content/978-3-540-37941-6.

Woodcock:1974:GECa

[WP74c] A. E. R. Woodcock and T. Poston. The geometry of the elementary catastrophes: (1). The cuspoids. Lecture Notes in Mathematics, 373:1–131, 1974. CODEN LNMAA2. ISBN 3-540-06681-0 (print), 3-540-37941-X (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0068968/.

Woodcock:1974:GECb

[WP74d] A. E. R. Woodcock and T. Poston. The geometry of the elementary catastrophes: (2). The hyperbolic and elliptic umbilics. Lecture Notes in Mathematics, 373:64–90, 1974. CODEN LNMAA2. ISBN 3-540-06681-0 (print), 3-540-37941-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0068969/.

Woodcock:1974:GECc

[WP74e] A. E. R. Woodcock and T. Poston. The geometry of the elementary catastrophes: (3). The parabolic umbilic. Lecture Notes in Mathematics, 373:91–131, 1974. CODEN LNMAA2. ISBN 3-540-06681-0 (print), 3-540-37941-X (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0068970/.

Wos:1970:PSS

[WR70] Lawrence Wos and George Robinson. Paramodulation and set of support. Lecture Notes in Mathematics, 125:276–310, 1970. CODEN LNMAA2. ISBN 3-540-04914-2 (print), 3-540-36262-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0060637/.

Wu:1971:DMS

[Wu71] Ta-Sun Wu. Disjointness of minimal sets. Lecture Notes in Mathematics, 235:134–140, 1971. CODEN LNMAA2. ISBN 3-540-05674-2 (print), 3-540-36996-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059206/.

Weiss:1974:ABL

[WW74a] Prof. Dr. Lionel Weiss and Prof. Dr. Jacob Wolfowitz. Asymptotic behavior of the likelihood function. Asymptotically sufficient statistics. Lecture Notes in Mathematics, 424:40–50, 1974. CODEN LNMAA2. ISBN 3-540-06970-4 (print), 3-540-37279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070231/.

Weiss:1974:EML

[WW74b] Prof. Dr. Lionel Weiss and Prof. Dr. Jacob Wolfowitz. Efficiency of maximum likelihood estimators. Lecture Notes in Mathematics, 424:51–60, 1974. CODEN LNMAA2. ISBN 3-540-06970-4 (print), 3-540-37279-2 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070232/.

Weiss:1974:MLE

[WW74c] Prof. Dr. Lionel Weiss and Prof. Dr. Jacob Wolfowitz. The maximum likelihood estimator. Lecture Notes in Mathematics, 424: 5–11, 1974. CODEN LNMAA2. ISBN 3-540-06970-4 (print), 3-540-37279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070228/.

Weiss:1974:MPEa

[WW74d] Prof. Dr. Lionel Weiss and Prof. Dr. Jacob Wolfowitz. The maximum probability estimator. Lecture Notes in Mathematics, 424: 12–29, 1974. CODEN LNMAA2. ISBN 3-540-06970-4 (print), 3-540-37279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070229/.

Weiss:1974:MPEc

[WW74e] Prof. Dr. Lionel Weiss and Prof. Dr. Jacob Wolfowitz. Maximum Probability Estimators and Related Topics, volume 424 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LN-MAA2. ISBN 0-387-06970-4, 3-540-06970-4 (print), 3-540-37279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 106 pp. LCCN QA3 .L28 no. 424; QA3 .L28 no. 424 QA273.

URL http://link.springer.com/book/10.1007/BFb0070226;
 http://www.springerlink.com/content/978-3-540-37279-
0.

Weiss:1974:MPEb

[WW74f] Prof. Dr. Lionel Weiss and Prof. Dr. Jacob Wolfowitz. Maximum probability estimators with a general loss function. Lecture Notes in Mathematics, 424:30–39, 1974. CODEN LNMAA2. ISBN 3-540-06970-4 (print), 3-540-37279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070230/.

Weiss:1974:PM

[WW74g] Prof. Dr. Lionel Weiss and Prof. Dr. Jacob Wolfowitz. Purpose of this monograph. Lecture Notes in Mathematics, 424:1–4, 1974. CODEN LNMAA2. ISBN 3-540-06970-4 (print), 3-540-37279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070227/.

Weiss:1974:TH

[WW74h] Prof. Dr. Lionel Weiss and Prof. Dr. Jacob Wolfowitz. Testing hypotheses. Lecture Notes in Mathematics, 424:61–92, 1974. CODEN LNMAA2. ISBN 3-540-06970-4 (print), 3-540-37279-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070233/.

Wallis:1972:CRS

[WWS72a] W. D. Wallis, Jennifer Seberry Wallis, and Anne Penfold Street. Combinatorics: Room Squares, Sum-Free Sets, Hadamard Matrices, volume 292 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06035-9 (print), 3-540-37994-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 498 (est.) pp. LCCN QA3 .L28 no. 292. URL http:/

/link.springer.com/book/10.1007/BFb0069907;   http://
www.springerlink.com/content/978-3-540-37994-2.

Wallis:1972:P

[WWS72b] W. D. Wallis, Jennifer Seberry Wallis, and Anne Penfold Street. Preliminaries. Lecture Notes in Mathematics, 292:2–28, 1972. CODEN LNMAA2. ISBN 3-540-06035-9 (print), 3-540-37994-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0069908/.

Yamada:1973:SSB

[Yam73] Toshihiko Yamada. The Schur subgroup of the Brauer group. Lecture Notes in Mathematics, 353:187–203, 1973. CODEN LN-MAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059270/.

Yamada:1974:BWT

[Yam74a] Toshihiko Yamada. The Brauer–Witt theorem. Lecture Notes in Mathematics, 397:20–33, 1974. CODEN LNMAA2. ISBN 3-540-06806-6 (print), 3-540-37733-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061707/.

Yamada:1974:CI

[Yam74b] Toshihiko Yamada. Chapter introduction. Lecture Notes in Mathematics, 397:1–3, 1974. CODEN LNMAA2. ISBN 3-540-06806-6 (print), 3-540-37733-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061704/.

Yamada:1974:CA

[Yam74c] Toshihiko Yamada. Cyclotomic algebras. Lecture Notes in Mathematics, 397:14–19, 1974. CODEN LNMAA2. ISBN 3-540-06806-6 (print), 3-540-37733-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061706/.

Yamada:1974:PSA

[Yam74d] Toshihiko Yamada. Properties of a Schur algebra. Lecture Notes in Mathematics, 397:89–94, 1974. CODEN LNMAA2. ISBN 3-540-06806-6 (print), 3-540-37733-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061710/.

Yamada:1974:SA

[Yam74e] Toshihiko Yamada. Schur algebras. Lecture Notes in Mathematics, 397:4–13, 1974. CODEN LNMAA2. ISBN 3-540-06806-6 (print), 3-540-37733-6 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061705/.

Yamada:1974:SSAb

[Yam74f] Toshihiko Yamada. The Schur subgroup of a 2-adic field. Lecture Notes in Mathematics, 397:54–88, 1974. CODEN LN-MAA2. ISBN 3-540-06806-6 (print), 3-540-37733-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061709/.

Yamada:1974:SSAa

[Yam74g] Toshihiko Yamada. The Schur subgroup of a p-adic field, p ≠ 2. Lecture Notes in Mathematics, 397:34–53, 1974. CODEN LNMAA2. ISBN 3-540-06806-6 (print), 3-540-37733-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061708/.

Yamada:1974:SSR

[Yam74h] Toshihiko Yamada. The Schur subgroup of a real field. Lecture Notes in Mathematics, 397:95–131, 1974. CODEN LN-MAA2. ISBN 3-540-06806-6 (print), 3-540-37733-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061711/.

Yamada:1974:SSI

[Yam74i] Toshihiko Yamada. The Schur subgroup of an imaginary field. Lecture Notes in Mathematics, 397:132–142, 1974. CODEN LN-MAA2. ISBN 3-540-06806-6 (print), 3-540-37733-6 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061712/.

Yamada:1974:SSB

[Yam74j] Toshihiko Yamada. The Schur Subgroup of the Brauer Group, volume 397 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06806-6 (print), 3-540-37733-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). v + 159 pp. LCCN QA3 .L28 no. 397. URL http:

//link.springer.com/book/10.1007/BFb0061703;   http:/
/www.springerlink.com/content/978-3-540-37733-7.

Yamada:1974:STS

[Yam74k] Toshihiko Yamada. Some theorems for a Schur algebra. Lecture Notes in Mathematics, 397:143–151, 1974. CODEN LN-MAA2. ISBN 3-540-06806-6 (print), 3-540-37733-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061713/.

Yamamuro:1974:CM

[Yam74l] Sadayuki Yamamuro. Compact mappings. Lecture Notes in Mathematics, 374:43–49, 1974. CODEN LNMAA2. ISBN 3-540-06709-4 (print), 3-540-37971-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061583/.

Yamamuro:1974:DFP

[Yam74m] Sadayuki Yamamuro. Definitions and fundamental properties. Lecture Notes in Mathematics, 374:6–42, 1974. CODEN LN-MAA2. ISBN 3-540-06709-4 (print), 3-540-37971-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061582/.

Yamamuro:1974:DMR

[Yam74n] Sadayuki Yamamuro. Differentiability of mappings of a real variable. Lecture Notes in Mathematics, 374:111–123, 1974. CODEN LNMAA2. ISBN 3-540-06709-4 (print), 3-540-37971-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061587/.

Yamamuro:1974:DSN

[Yam74o] Sadayuki Yamamuro. Differentiability of semi-norms. Lecture Notes in Mathematics, 374:76–97, 1974. CODEN LN-MAA2. ISBN 3-540-06709-4 (print), 3-540-37971-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061585/.

Yamamuro:1974:DCT

[Yam74p] Sadayuki Yamamuro. Differential Calculus in Topological Linear Spaces, volume 374 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1974. CODEN LNMAA2. ISBN 3-540-06709-4 (print), 3-540-37971-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 179 pp. LCCN QA3 .L28 no. 374. URL http:/

/link.springer.com/book/10.1007/BFb0061580;   http://
www.springerlink.com/content/978-3-540-37971-3.

Yamamuro:1974:I

[Yam74q] Sadayuki Yamamuro. Introduction. Lecture Notes in Mathematics, 374:1–5, 1974. CODEN LNMAA2. ISBN 3-540-06709-4 (print), 3-540-37971-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061581/.

Yamamuro:1974:IMT

[Yam74r] Sadayuki Yamamuro. Inverse mapping theorems. Lecture Notes in Mathematics, 374:50–75, 1974. CODEN LNMAA2. ISBN 3-540-06709-4 (print), 3-540-37971-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0061584/.

Yamamuro:1974:SDM

[Yam74s] Sadayuki Yamamuro. Sets of differentiable mappings. Lecture Notes in Mathematics, 374:124–139, 1974. CODEN LN-MAA2. ISBN 3-540-06709-4 (print), 3-540-37971-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0061588/.

Yamamuro:1974:S

[Yam74t] Sadayuki Yamamuro. Smoothness. Lecture Notes in Mathematics, 374:98–110, 1974. CODEN LNMAA2. ISBN 3-540-06709-4 (print), 3-540-37971-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0061586/.

Yates:1972:ACP

[Yat72a] C. E. M. Yates. Abstracts of contributed papers. Lecture Notes in Mathematics, 255:336–351, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059551/.

Yates:1972:ISI

[Yat72b] C. E. M. Yates. Initial segments and implications for the structure of degrees. Lecture Notes in Mathematics, 255:305–335, 1972. CODEN LNMAA2. ISBN 3-540-05744-7 (print), 3-540-37162-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059550/.

Yevtushenko:1970:ACO

[Yev70] Yu. G. Yevtushenko. Approximate calculation of optimal control by averaging method. Lecture Notes in Mathematics, 112: 116–127, 1970. CODEN LNMAA2. ISBN 3-540-04901-0 (print), 3-540-36204-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060201/.

Yorke:1970:SDE

[Yor70] J. A. Yorke, editor. Seminar on Differential Equations and Dynamical Systems, II: Seminar lectures at the University of Maryland 1969, volume 144 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04933-9 (print), 3-540-36306-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). LCCN QA3 .L28 no. 144. URL http:/

/link.springer.com/book/10.1007/BFb0059912;   http://
www.springerlink.com/content/978-3-540-36306-4.

Yorke:1971:STD

[Yor71] James A. Yorke. Selected topics in differential delay equations. Lecture Notes in Mathematics, 243:16–28, 1971. CODEN LN-MAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058715/.

Yoshizawa:1971:SAP

[Yos71] Taro Yoshizawa. Stability for almost periodic systems. Lecture Notes in Mathematics, 243:29–39, 1971. CODEN LN-MAA2. ISBN 3-540-05708-0 (print), 3-540-37080-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058716/.

Young:1970:CSL

[You70] Gail S. Young. Continuous selections on locally compact separable metric spaces. Lecture Notes in Mathematics, 171:102– 110, 1970. CODEN LNMAA2. ISBN 3-540-05293-3 (print), 3-540-36419-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069729/.

Young:1974:ASM

[You74] David M. Young. On the accelerated SSOR method for solving elliptic boundary value problems. Lecture Notes in Mathematics, 363:176, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0069135.

Zabrodsky:1970:COH

[Zab70] Alexander Zabrodsky. Cohomology operations and homotopy commutative H-spaces. Lecture Notes in Mathematics, 168:308– 317, 1970. CODEN LNMAA2. ISBN 3-540-05300-X (print), 3-540-36437-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058530/.

Zabrodsky:1971:SCC

[Zab71] Alexander Zabrodsky. On spherical classes in the cohomology of H-spaces. Lecture Notes in Mathematics, 196:25–33, 1971. CODEN LNMAA2. ISBN 3-540-05461-8 (print), 3-540-36621-0 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0067742/.

Zabrodsky:1974:EHT

[Zab74] Alexander Zabrodsky. p Equivalences and homotopy type. Lecture Notes in Mathematics, 418:161–171, 1974. CODEN LN-MAA2. ISBN 3-540-06963-1 (print), 3-540-37268-7 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070653/.

Zagier:1972:EPC

[Zag72a] Don Bernard Zagier. Equivariant Pontrjagin Classes and Applications to Orbit Spaces: Applications of the G-signature Theorem to Transformation Groups, Symmetric Products and Number Theory, volume 290 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CODEN LNMAA2. ISBN 3-540-06013-8 (print), 3-540-37988-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). 127 (est.) pp. LCCN QA3 .L28 no. 290. URL http:/

/link.springer.com/book/10.1007/BFb0060730;   http://
www.springerlink.com/content/978-3-540-37988-1.

Zagier:1972:STS

[Zag72b] Don Bernard Zagier. The G-signature theorem and some elementary number theory. Lecture Notes in Mathematics, 290: 96–127, 1972. CODEN LNMAA2. ISBN 3-540-06013-8 (print), 3-540-37988-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0060733/.

Zagier:1972:CRH

[Zag72c] Don Bernard Zagier. L-classes of rational homology manifolds. Lecture Notes in Mathematics, 290:1–31, 1972. CODEN LN-MAA2. ISBN 3-540-06013-8 (print), 3-540-37988-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060731/.

Zagier:1972:CSP

[Zag72d] Don Bernard Zagier. L-classes of symmetric products. Lecture Notes in Mathematics, 290:32–95, 1972. CODEN LN-MAA2. ISBN 3-540-06013-8 (print), 3-540-37988-6 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0060732/.

Zassenhaus:1973:EOM

[Zas73] Hans Zassenhaus. On the embedding of an order into a maximal order. Lecture Notes in Mathematics, 353:204–221, 1973. CODEN LNMAA2. ISBN 3-540-06518-0 (print), 3-540-37818-9 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0059271/.

Zeeman:1971:BW

[Zee71] E. C. Zeeman. Breaking of waves. Lecture Notes in Mathematics, 206:2–6, 1971. CODEN LNMAA2. ISBN 3-540-05495-2 (print), 3-540-36662-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070119/.

Zienkiewicz:1974:CVP

[Zie74] O. C. Zienkiewicz. Constrained variational principles and penalty function methods in finite element analysis. Lecture Notes in Mathematics, 363:207–214, 1974. CODEN LNMAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0069138/.

Zimmer:1972:CGC

[Zim72a] Horst G. Zimmer. Class groups and class fields of algebraic number fields. Lecture Notes in Mathematics, 262:51–56, 1972. CO-DEN LNMAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0058483/.

Zimmer:1972:CNA

[Zim72b] Horst G. Zimmer. Class numbers of algebraic number fields. Lecture Notes in Mathematics, 262:40–50, 1972. CODEN LN-MAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058482/.

Zimmer:1972:CPM

[Zim72c] Horst G. Zimmer. Computational Problems, Methods, and Results in Algebraic Number Theory, volume 262 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1972. CO-DEN LNMAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 103 pp. LCCN QA3 .L28 no. 262. URL http:/

/link.springer.com/book/10.1007/BFb0058472;   http://
www.springerlink.com/content/978-3-540-37466-4.

Zimmer:1972:CF

[Zim72d] Horst G. Zimmer. Continued fractions. Lecture Notes in Mathematics, 262:14–18, 1972. CODEN LNMAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058477/.

Zimmer:1972:DE

[Zim72e] Horst G. Zimmer. Diophantine equations. Lecture Notes in Mathematics, 262:56–70, 1972. CODEN LNMAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058484/.

Zimmer:1972:FP

[Zim72f] Horst G. Zimmer. Factorization of polynomials. Lecture Notes in Mathematics, 262:5–11, 1972. CODEN LNMAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058475/.

Zimmer:1972:FE

[Zim72g] Horst G. Zimmer. Field extensions. Lecture Notes in Mathematics, 262:18–24, 1972. CODEN LNMAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058478/.

Zimmer:1972:FF

[Zim72h] Horst G. Zimmer. Finite fields. Lecture Notes in Mathematics, 262:2–4, 1972. CODEN LNMAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058474/.

Zimmer:1972:GG

[Zim72i] Horst G. Zimmer. Galois groups. Lecture Notes in Mathematics, 262:11–14, 1972. CODEN LNMAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058476/.

Zimmer:1972:HPC

[Zim72j] Horst G. Zimmer. The Hasse principle for cubic surfaces. Lecture Notes in Mathematics, 262:70–71, 1972. CODEN LN-MAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0058485.

Zimmer:1972:I

[Zim72k] Horst G. Zimmer. Introduction. Lecture Notes in Mathematics, 262:1–2, 1972. CODEN LNMAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0058473.

Zimmer:1972:MO

[Zim72l] Horst G. Zimmer. Modules and orders. Lecture Notes in Mathematics, 262:24–29, 1972. CODEN LNMAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0058479/.

Zimmer:1972:PLF

[Zim72m] Horst G. Zimmer. Products of linear forms. Lecture Notes in Mathematics, 262:29–35, 1972. CODEN LNMAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058480/.

Zimmer:1972:UAN

[Zim72n] Horst G. Zimmer. Units in algebraic number fields. Lecture Notes in Mathematics, 262:35–40, 1972. CODEN LNMAA2. ISBN 3-540-05822-2 (print), 3-540-37466-3 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0058481/.

Zisman:1970:TEF

[Zis70] M. Zisman. K-théorie élémentaire. (French) [Elementary K theory]. Lecture Notes in Mathematics, 136:11–65, 1970. CODEN LNMAA2. ISBN 3-540-04925-8 (print), 3-540-36289-4 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059021/.

Zlamal:1974:FEM

[Zlá74] Miloš Zlámal. Finite element methods for parabolic equations. Lecture Notes in Mathematics, 363:215–221, 1974. CODEN LN-MAA2. ISBN 3-540-06617-9 (print), 3-540-37914-2 (e-book).

ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0069139/.

Zucker:1973:BM

[Zuc73a] J. I. Zucker. Back matter. Lecture Notes in Mathematics, 344: 454–485, 1973. CODEN LNMAA2. ISBN 3-540-06491-5 (print), 3-540-37806-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/content/pdf/bbm:
978-3-540-37806-8/1.pdf.

Zucker:1973:E

[Zuc73b] J. I. Zucker. Erratum. Lecture Notes in Mathematics, 344:487– 503, 1973. CODEN LNMAA2. ISBN 3-540-06491-5 (print), 3-540-37806-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
978-3-540-37806-8_7/.

Zucker:1973:IID

[Zuc73c] J. I. Zucker. Iterated inductive definitions, trees and ordinals. Lecture Notes in Mathematics, 344:392–453, 1973. CODEN LN-MAA2. ISBN 3-540-06491-5 (print), 3-540-37806-5 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0066745/.

Zieschang:1970:AEG

[ZVC70a] H. Zieschang, E. Vogt, and H.-D. Coldewey. Automorphismen ebener Gruppen. (German) [Automorphisms of plane groups]. Lecture Notes in Mathematics, 122:91–136, 1970. CODEN LN-MAA2. ISBN 3-540-04911-8 (print), 3-540-36256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059079/.

Zieschang:1970:EDG

[ZVC70b] H. Zieschang, E. Vogt, and H.-D. Coldewey. Ebene diskontinuierliche Gruppen. (German) [planar discontinuous groups]. Lecture Notes in Mathematics, 122:55–90, 1970. CODEN LN-MAA2. ISBN 3-540-04911-8 (print), 3-540-36256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059078/.

Zieschang:1970:FG

[ZVC70c] H. Zieschang, E. Vogt, and H.-D. Coldewey. Flächen. (German) []. Lecture Notes in Mathematics, 122:41–54, 1970. CODEN LNMAA2. ISBN 3-540-04911-8 (print), 3-540-36256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059077/.

Zieschang:1970:FED

[ZVC70d] H. Zieschang, E. Vogt, and H.-D. Coldewey. Flächen und ebene diskontinuierliche Gruppen, volume 122 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1970. CODEN LNMAA2. ISBN 3-540-04911-8 (print), 3-540-36256-8 (ebook). ISSN 0075-8434 (print), 1617-9692 (electronic). viii + 203 pp. LCCN QA3 .L28 no. 122. URL http:/

/link.springer.com/book/10.1007/BFb0059074;   http://
www.springerlink.com/content/978-3-540-36256-2.

Zieschang:1970:FGG

[ZVC70e] H. Zieschang, E. Vogt, and H.-D. Coldewey. Freie Gruppen und Graphen. (German) []. Lecture Notes in Mathematics, 122:1– 17, 1970. CODEN LNMAA2. ISBN 3-540-04911-8 (print), 3-540-36256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0059075/.

Zieschang:1970:KBG

[ZVC70f] H. Zieschang, E. Vogt, and H.-D. Coldewey. Kombinatorische Beschreibung von Gruppen und Flächenkomplexe. (German) []. Lecture Notes in Mathematics, 122:18–40, 1970. CODEN LN-MAA2. ISBN 3-540-04911-8 (print), 3-540-36256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0059076/.

Zieschang:1970:KAT

[ZVC70g] H. Zieschang, E. Vogt, and H.-D. Coldewey. Zur komplexanalytischen Theorie der Riemannschen Flächen und ebenen diskontinuierlichen Gruppen. (German) []. Lecture Notes in Mathematics, 122:137–166, 1970. CODEN LNMAA2. ISBN 3-540-04911-8 (print), 3-540-36256-8 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0059080/.

Zygmund:1971:ATF

[Zyg71a] Antoni Zygmund. Applications de la transformation de Fourier aux intégrales singulières. (French) []. Lecture Notes in Mathematics, 204:29–38, 1971. CODEN LNMAA2. ISBN 3-540-05491-X (print), 3-540-36660-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070449/.

Zygmund:1971:EDC

[Zyg71b] Antoni Zygmund. Étude du cas du noyau pair. (French) []. Lecture Notes in Mathematics, 204:39–46, 1971. CODEN LN-MAA2. ISBN 3-540-05491-X (print), 3-540-36660-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070450/.

Zygmund:1971:ISF

[Zyg71c] Antoni Zygmund. Intégrales Singulières. (French) [Singular integrals], volume 204 of Lecture Notes in Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1971. CODEN LNMAA2. ISBN 3-540-05491-X (print), 3-540-36660-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). iv + 53 pp. LCCN QA3 .L28 no. 204. URL http:/

/link.springer.com/book/10.1007/BFb0070445;   http://
www.springerlink.com/content/978-3-540-36660-7.

Zygmund:1971:IFI

[Zyg71d] Antoni Zygmund. Introduction. (French) [Introduction]. Lecture Notes in Mathematics, 204:1–6, 1971. CODEN LNMAA2. ISBN 3-540-05491-X (print), 3-540-36660-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://link.springer.

com/chapter/10.1007/BFb0070446/.

Zygmund:1971:OSE

[Zyg71e] Antoni Zygmund. Opérateurs singuliers et équations aux derivées partielles. (French) []. Lecture Notes in Mathematics, 204: 47–49, 1971. CODEN LNMAA2. ISBN 3-540-05491-X (print), 3-540-36660-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (elec-

tronic). URL http://link.springer.com/chapter/10.1007/
BFb0070451/.

Zygmund:1971:QPO

[Zyg71f] Antoni Zygmund. Quelques problémes ouverts. (French) []. Lecture Notes in Mathematics, 204:50–52, 1971. CODEN LN-MAA2. ISBN 3-540-05491-X (print), 3-540-36660-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL http://

link.springer.com/chapter/10.1007/BFb0070452/.

Zygmund:1971:THCb

[Zyg71g] Antoni Zygmund. Transformée de Hilbert dans le cas de plusieurs variables (noyaux impairs). (French) []. Lecture Notes in Mathematics, 204:26–28, 1971. CODEN LNMAA2. ISBN 3-540-05491-X (print), 3-540-36660-1 (e-book). ISSN 0075-8434 (print), 1617-

9692 (electronic). URL http://link.springer.com/chapter/
10.1007/BFb0070448/.

Zygmund:1971:THCa

[Zyg71h] Antoni Zygmund. Transformée de Hilbert dans le cas d’une variable. (French) []. Lecture Notes in Mathematics, 204:7–25, 1971. CODEN LNMAA2. ISBN 3-540-05491-X (print), 3-540-36660-1 (e-book). ISSN 0075-8434 (print), 1617-9692 (electronic). URL

http://link.springer.com/chapter/10.1007/BFb0070447/.